A Bibliography of Papers in Lecture Notes in Computer Science (2002) (Part 5 of 5)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

18 January 2025 Version 1.14

Title word cross-reference

(0, 1, ∞) [322]. (2^{m}) [834]. (λ) [721]. (t, n) [2171]. + [1488]. 1 [960, 3057, 2896, 19]. 1.5 [1327]. 2 [2439, 1861, 1894, 2665, 908, 2170, 959, 1911, 504, 1902, 1913, 1909, 1034, 1906, 971, 3422]. 2.5 [2404]. 27/26 [1332]. 3 [1799, 1325, 1032, 1022, 2274, 992, 2945, 1498, 1035, 205, 1897, 1894, 3003, 3310, 3059, 1868, 1798, 1893, 1911, 1881, 1847, 1849, 1816, 1054, 3445, 1055, 1751, 1823, 1902, 1882, 1828, 1913, 3042, 1909, 1867, 2254, 1907, 1928, 3041, 1917, 1020, 1820, 1906, 403, 1824, 1892]. 4 [1779]. 5n − o(n) [501]. 6 [1778]. ^{∗} [542]. ^{K} [617]. ^{n} [498]. ^{rd} [432]. _{F D} [2963]. _{CTL} [2551]. _{n} [506]. _{φ} [2640]. α [2988]. C [3066]. H [934]. PSPACE [478]. R [934]. χ [354]. Δ [2965]. F [2680]. GF(2^{m}) [114]. H [2491]. K [1101, 1685, 2692, 2717, 960, 131, 488, 1331, 495, 700, 724, 143, 2980]. K_{3,3} [2675]. K_{5} [2675]. l [1549]. λ [1617, 2292, 304, 946]. m [2180]. P [1517]. d/dt [343]. LDL [611]. M_{⟩\⌉} [611]. O(∞) [1337]. RT [1517]. SHOQ(D) [574]. X [1279]. μ [2122, 2563, 2125, 2121]. N [770, 19, 2363, 3305]. O(n log n) [3051]. O(pn + 1.151^{p}) [2674]. Ω [2568, 2112]. p [2995, 2674]. P − SHOQ(D) [574]. π [556, 531]. qMμ [2563]. QSI [2253]. \sqrt{2} [2568]. t [3407, 1651]. u [1549].

-Algorithm [2674]. -Approximation [1327, 1332]. -ary [1101]. -Automata [2112]. -Calculus [556, 2122, 2563, 2125, 2121]. -Click [2896]. -Colorable [488]. -Colorings [2491]. -Competitive [1337].

-Completeness [478]. -Conjunctions [504]. -Connected [960, 1617]. -Coteries [131]. -Counter [322]. -Cycle [1325]. -D [1909, 2439, 1022, 1847, 1816, 1055, 1751, 1909, 1907, 1034]. -D/ [1909]. -Depth [143]. -Dimensional [3305, 1779]. -DOF [1778]. -Edge-Connectivity [2665]. -Graph [2292]. -Median [2995]. -Minor-Free [2675]. -Nearest [2980]. -NN [700]. -OO [1517]. -Planarity [3066]. -Profit [2674]. -Reliable [2680]. -Resilient [2180]. -Round [959]. -Sat [495, 3445, 2692]. -Set [1685, 2363]. -Sided [3057]. -Spanners [1651]. -Term [724]. -Threshold [2171]. -to- [19]. -Trees [2965].

.NET [3201].

1-2-3 [440]. 1-out-of-n [2176]. 1.0 [1224]. 1.1 [44]. 1.1GHz [3112]. 12th [3346]. 1st [659].

2 [1535, 342, 2051, 1504]. 2.0 [339, 3334]. 2001 [3256]. 2002 [2316]. 21 [3413]. 2SAT [2668, 3380].

300 [1203]. 3G [2333, 2497].

4D [3415, 1826, 1831]. 4G [3332].

64b [3112].

802.11 [2466].

9000 [2484].

A-code [2154]. a-priori [1028]. A-Prolog [312]. AAA [87]. AAL2 [58]. abC [356]. Abdominal [1893, 1823]. Abduction [3196]. Abductive [271, 601]. ABEL [608]. Abelian [965, 969]. Abilities [665]. Ability [3426]. ABM [1964]. Abnormal [2268]. Abnormalities [1846]. Abox [2572]. Absence [3131]. Absorbing [2033, 2392].

Abstract [266, 2688, 2244, 532, 1551, 2401, 2680, 241, 3300, 554, 1628, 3465, 953, 972, 3297, 318, 2624, 2702, 966]. Abstracting [356, 2590]. Abstraction [557, 1371, 2630, 333, 316, 2629, 324, 2299, 2628, 1992, 322, 945, 1093, 2632]. Abstraction-Refinement [333]. Abstractions [3349]. abusiveness [1337]. Academic [805]. Academicians [2322]. Accelerated [448]. Accelerating [1087]. Accelerators [3095]. Acceptance [429, 465, 2707, 330]. Access [1962, 273, 1279, 2756, 413, 1273, 828, 412, 2579, 1408, 1409, 135, 1956, 1220, 174, 21, 2718, 2429, 2067, 2807, 2201, 84, 3262, 2577, 686, 3241, 1264, 815, 2467, 2419, 2853, 1366, 63, 1278]. Accessed [2356]. Accesses [2612]. Accessibility [484, 1281]. Accessible [23]. Accessing [3280, 3275, 2445, 2824]. ACCORD [653]. ACCORD/UML [653]. According [1938, 2076, 665]. Account [289, 2232, 1673]. Accounting [2046, 1174, 2533, 2534]. Accumulators [952]. Accuracy [228, 2590, 1281, 1926, 2956, 2096, 1590]. Accurate [3249, 1089, 108, 2094, 2624, 1892]. ACE [808]. Achieve [3146]. Achieving [2831]. Acid [1068]. ACK [53]. ACK-Based [53]. Acquiring [1443, 1472]. Acquisition [1585, 2274, 2933, 3417, 787, 1452, 1199, 527, 1446, 2486, 1054, 2786, 1933]. Acrobot^{T M} [1781]. Across [1838, 1658, 1045, 1161]. Act [2915]. Acting [1776]. Action [3035, 1114, 582, 308, 534, 3333, 785, 2632, 596]. Actionable [750]. Actions [598, 597, 274]. Activation [1714, 1931, 1722, 3425]. Activation-Related [1931]. Active [2058, 1028, 2542, 3337, 266, 2808, 3335, 958, 3324, 3327, 1057, 2061, 1222, 1772, 2807, 1054, 51, 1004, 1823, 3326, 818, 2278, 3340, 1523, 3332, 3329, 2806, 3341, 1370, 3328, 3333, 2604, 1166, 1822, 730, 1831, 3216]. Activities [2943, 297, 3037, 1274, 3254, 1699, 1860, 392, 1344]. Activity [156, 1794]. Actors [1372]. Actuated [1753]. Actuation [1761]. Actuators [1772]. Acyclic [3055, 1652, 1559]. Ad [105, 2354, 1389, 2077, 977, 192, 1238, 3265, 104, 2809, 189, 3258, 103, 3229, 1390, 188, 2480, 102, 2729, 101]. Ad-Hoc [105, 1389, 3265, 104, 1390, 977, 192, 189, 3229, 188, 2729]. AdaBoost.M1 [697]. Adaptable [2761, 1974, 422, 1765, 1975]. Adaptation [2955, 2064, 3416, 79, 2580, 1204, 2623, 2333, 3021, 1743, 2583, 3453, 397, 856]. Adapter [2747]. Adapting [2156, 1291, 3429, 374]. Adaptive [2955, 2989, 1731, 2262, 1715, 2913, 2338, 2767, 3412, 2083, 3268, 2738, 926, 3026, 3126, 357, 119, 2448, 1615, 137, 949, 2501, 3232, 2585, 3326, 2512, 996, 200, 879, 1618, 55, 1030, 383, 1019, 1499, 2456, 1999, 2480, 2447, 2515, 3426, 2063, 2374, 2960, 149, 1957, 3425, 2517, 2760, 848, 2984, 858]. Adaptivity [414]. Add [1486]. Add-on [1486]. Adding [2838, 2770]. Addition [1214, 2655]. Addition-Subtraction [1214]. Additivity [2387]. Address [29, 1381, 1383, 1384, 677]. Addressable [684]. Addressing [394, 439]. Adequacy [2248]. Adjoining [3019]. adjoint [301]. Adjusting [1507]. Adleman [2377]. ADMIN} [1156]. Administration [1154, 1156, 1152, 1140, 1116, 1128]. Administrations [1120, 1121, 1138, 1158, 1108]. Administrative [1124, 1180]. Admission [2446, 1338, 52, 2480, 75]. ADML [2504]. Advance [1338]. Advanced [870, 1989, 2056, 2210, 2209, 267]. Advances [3066, 213]. Advantage [2471]. Advantages [1042]. Adversaries [964, 2164]. Advisor [621]. AEQUITAS [1165]. AES [948]. Affability [1604]. Affairs [1145]. Affine [908, 1790]. Affordances [408]. After [2936, 1779]. Against [323, 2621, 958, 2181, 2165, 964, 2161, 1788].

Agenda [3009]. Agent [3206, 3200, 2024, 2473, 3007, 2719, 2783, 3209, 1211, 1212, 2507, 2844, 3004, 3010, 3207, 119, 3307, 2809, 2490, 2462, 3309, 140, 10, 2805, 2489, 370, 2508, 2817, 2825, 1206, 3024, 2806, 2814, 2823, 2800, 149, 2811, 2273, 2434, 1608, 2827, 2270, 3412, 568, 567, 577, 581, 1209, 2821, 2271, 1210, 1666, 307, 570, 2720, 3011]. Agent-Based [2507, 119, 2809, 2490, 10, 2805, 1206, 3024, 2811]. Agent-Mediated [2273]. Agent-Oriented [3309, 2827]. Agents [2810, 903, 3035, 2808, 2813, 2819, 3211, 1207, 2506, 2802, 2804, 2554, 2812, 3202, 2721, 2718, 3265, 2794, 169, 2815, 3006, 2716, 146, 2474, 3005, 3008, 3330, 394, 2272, 2803, 2820, 2816, 3203, 601, 2826, 921, 600, 2824, 3409, 2822, 1505]. Agglomerative [1090]. Agglutinative [1198]. Aggregate [2061, 760]. Aggregated [2454]. Aggregation [2951, 1184, 2595, 2855, 746]. Agile [2050, 426, 462, 452, 424, 425, 437, 467, 445, 419, 465, 436, 450, 455, 442, 471, 433]. Aglets [3204]. AGM [2170]. Agnostic [3126]. Agreement [1685, 2346, 2343, 3102, 2345, 1684]. Agreements [3125, 3251, 2325, 2062]. Ahead [16, 1143, 143]. AI [619]. Aid [411]. AIDA [1152]. Aided [1351, 2303]. Aimed [1264, 3426]. Air [1720, 3453]. Aircraft [802]. Airfoil [918]. Airfoils [849]. Airway [1850]. Alarm [2929, 2615]. Alert [1268, 2616]. Alerts [2614]. Alexandria [3287]. Algebra [257, 141, 222, 552]. Algebraic [480, 2174, 762, 933, 2308, 948, 259]. Algebras [1551, 170]. Algorithm [2523, 868, 1875, 2030, 2530, 904, 2871, 3094, 1434, 3077, 3027, 2970, 3209, 902, 3463, 871, 3427, 150, 3144, 3057, 235, 1213, 1215, 926, 78, 2665, 2986, 1332, 1709, 1209, 2425, 1956, 3161, 30, 861, 16, 2682, 865, 71, 2584, 2410, 1378, 51, 1337, 509, 53, 94, 132, 31, 876, 3174, 913, 914, 1801, 2995, 1075, 2971, 879, 3187, 1226, 2583, 717, 3115, 1600, 1819, 867, 880, 881, 2644, 769, 1080, 34, 2059, 2699, 179, 2437, 2957, 1003, 2956, 2403, 2674, 3426, 2722, 27, 2016, 1786, 3434, 2789]. Algorithm-Based [2584]. Algorithmic [1098, 335, 476]. Algorithms [2982, 1322, 895, 1539, 1324, 901, 2691, 1006, 740, 970, 480, 2668, 1325, 742, 2666, 1650, 551, 1326, 487, 1069, 850, 906, 1015, 2675, 1089, 129, 1660, 3233, 3189, 283, 2170, 2294, 495, 1074, 1647, 578, 1553, 2685, 1334, 789, 2988, 852, 499, 2348, 888, 841, 910, 999, 2698, 505, 29, 2118, 3424, 911, 897, 866, 842, 510, 632, 1339, 2417, 915, 2265, 916, 918, 1214, 1944, 2673, 662, 2983, 1065, 2991, 847, 133, 2405, 889, 1623, 2633, 1341, 920, 3444, 859, 2389, 882, 1577, 1924, 2678, 3180, 2980]. Algorithms-Based [3189]. ALiCE [3383, 1471]. Aligning [2433]. Alignment [2083, 2084, 1931, 1250, 1881, 2094, 1264, 1081, 1864, 1266, 1080, 2086]. Alignments [2085, 1077]. Alinari [1278]. All-Interval-Series [2258]. All-Optical [38]. All-Pairs [2655]. All-to-all [1534]. Alleviating [2062]. Allocation [3227, 2483, 2053, 3467, 3388, 1675, 2273, 50, 3252]. Allow [1111]. Almost [515, 2707, 778, 734]. ALOHA [63]. Alphabet [538]. Alphabet-Based [538]. Alphabets [1630]. Alterations [2960]. Altering [1746]. Alternating [704, 2120, 2720, 2121]. Alternating-Offers [2720]. Alternation [2115, 620]. Alternative [120, 1056, 1265, 1216, 1571]. Alternatives [1376]. Alzheimer [1844]. Ambient [2525]. Ambients [547, 546]. Ambiguity [3190, 1092]. AMBO [2353]. AMC [357, 137]. Amino [1068]. Amino-Acid [1068]. AMnet [3334]. Among [3276, 2163, 3092, 3386, 3005, 2062, 217]. amongst [1120, 1158]. Amperometric [1010]. Amplified [2165]. Amplitude [99]. Amygdala [3431]. Analog [346]. Analogies [3181]. Analogy [3188]. Analyses [2693, 710]. Analysing [630, 2196, 799, 271]. Analysis [1097, 2248, 559, 630, 340, 739, 1714, 2943, 325, 1610, 3372, 628, 1611, 1682, 800, 968, 801, 1261, 558, 1569, 1494, 3282, 2926, 957, 1550, 487, 2630, 1989, 177, 1707, 1064, 1203, 1000, 303, 3278, 2578, 2003, 1486, 1447, 1292, 1193, 1812, 1239, 1798, 1368, 410, 852, 153, 314, 1605, 1800, 12, 2939, 841, 41, 58, 1729, 2928, 3384, 2382, 1345, 46, 3450, 1503, 1881, 71, 1769, 2168, 2505, 762, 3258, 93, 790, 1902, 842, 116, 1742, 2151, 219, 510, 3382, 830, 3039, 3421, 515, 1417, 820, 1031, 2971, 2463, 2018, 993, 1909, 961, 732, 776, 809]. Analysis [1711, 115, 2565, 1638, 1794, 3172, 3197, 272, 47, 1827, 1547, 2633, 1810, 1235, 1469, 3389, 2004, 216, 1988, 1537, 3093, 1841, 1366, 1010, 1258, 2392, 1170, 1831, 328, 2873, 1602, 2200]. Analyst [2229]. Analytic [1058]. Analytical [1539, 3412, 1511]. Analyze [1277]. Analyzing [2336, 1604, 43, 1664, 2614, 2309, 1231]. Anaphoric [2441]. Anatomical [1922, 1887, 1883]. Ancestral [1066]. Aneurysms [1800, 1801]. Angiography [2274, 1882, 1899]. Angle [1038]. Animal [1784]. Anisotropic [1025]. ANN [884]. Annealing [2481, 3234]. Annotation [3296, 2782, 1280, 1275, 1052, 1456, 1238, 1250]. Annotations [2544, 2792]. Anomalous [2612]. Anomaly [2338, 2613]. Anomaly-Based [2613]. Anonymizing [670, 669]. Anonymous [952, 2155, 2815, 2776, 2482]. ANQL [3329]. Answer [603, 293, 582, 621, 587, 606, 602]. Answering [2735, 251, 2841, 611, 770, 2857]. Answers [2834]. Ant [2030, 2269, 922, 925, 606, 421, 1216]. Ant-Based [925]. ANTARCTICA [3266]. Antenna [109]. Anticipating [404]. Ants [2869]. Any [1521, 2016]. AOP [834]. Aortic [1908]. API [3238]. Apnea [2940]. Apparel [1205]. Appearance [1000, 1055, 2283]. Appearance-Based [1055]. Appliance [2413]. Appliances [61]. Applicable [888]. Application [227, 229, 1085, 974, 1299, 630, 1006, 1493, 299, 2756, 952, 3367, 1488, 2891, 1683, 2038, 1491, 1725, 1857, 261, 1540, 3003, 3126, 3456, 1927, 1238, 910, 863, 401, 1503, 2330, 2501, 26, 2065, 87, 3215, 2498, 3387, 2508, 1024, 1910, 1886, 391, 917, 1516, 2371, 3155, 2449, 1619, 1078, 945, 3332, 2002, 2339, 2456, 1837, 1787, 2213, 2395, 1295]. Application-Aware [3387]. Application-Centric [2501]. Application-Level [2449]. Application-Oriented [26]. Application-Specific [2756, 261]. Applications [1097, 3217, 3331, 2354, 2529, 2740, 1094, 3210, 482, 402, 977, 1971, 1494, 1990, 1377, 389, 1498, 831, 3076, 1873, 2770, 2758, 1015, 1513, 3221, 3268, 2818, 3079, 3325, 2337, 2582, 2759, 1356, 1074, 1271, 1072, 1496, 2580, 21, 3265, 1357, 475, 1939, 3428, 2747, 1532, 654, 1616, 510, 1504, 374, 2778, 2012, 2583, 973, 2825, 1866, 2765, 2696, 19, 2596, 367, 1305, 1421, 1794, 2566, 167, 881, 2069, 735, 779, 1537, 2217, 3122, 3154, 3263, 144, 2760]. Applicator [1865]. Applied [801, 1657, 3148, 919, 1165, 527, 495, 830, 3185, 3012, 2980]. Apply [426]. Applying [1309, 3071, 463, 2462, 332, 618, 2991, 954, 2537]. Approach [2058, 1304, 2261, 1578, 1450, 1714, 3081, 1314, 1610, 2446, 922, 816, 378, 2813, 3083, 3016, 2966, 3458, 3385, 3423, 2985, 2839, 1136, 747, 1029, 205, 3177, 1199, 1689, 1894, 3189, 3429, 3299, 1146, 1552, 3145, 3055, 317, 1613, 1604, 1179, 3432, 875, 888, 2787, 2013, 3365, 2271, 1505, 1357, 1847, 189, 3302, 1663, 2092, 2805, 3220, 1463, 1980, 2695, 1856, 1937, 2886, 1898, 301, 805, 3015, 716, 2778, 2445, 2519, 1618, 1037, 384, 2371, 2979, 312, 1580, 1620, 2615, 1355, 2961, 3466, 2743, 1672, 812, 271, 1674, 2930, 385, 2032, 734, 778, 1018, 1622, 1623, 1576, 442]. Approach [1810, 2020, 2814, 3453, 2604, 2499, 3036, 1148, 1639, 1170, 397, 1397, 1398, 3, 5, 1602, 1609, 3009]. Approaches [3007, 2976, 828, 2433, 3464, 2110, 1806, 3354, 649, 2037, 1686]. Appropriate [10, 821]. Approximability [2685, 476, 2687]. Approximate [1649, 2839, 2683, 2476, 1538, 1589, 772, 6]. Approximated [14]. Approximately [3129]. Approximating [2692, 2288, 1086, 1330, 1329, 1331, 3169, 2854]. Approximation [1322, 2691, 321, 2668, 1325, 2000, 1326, 1069, 1327, 1572, 2665, 259, 1332, 495, 1334, 1877, 2667, 2669, 1336, 1339, 1648, 886, 1201, 1809, 1591, 3312, 525]. Approximations [478, 1550, 1617, 309, 1614, 754, 1556, 1562]. Apriori [1600]. APSEE} [2304]. Arabic [1635]. Arbitrary [965, 3075, 259, 3366]. Arc [235, 236, 234]. Architectural [811, 1736, 826, 1689, 1146, 514, 230, 2477]. Architecture [2058, 1696, 154, 1389, 1935, 3337, 3237, 2808, 2046, 2763, 1490, 2270, 1717, 2714, 651, 2758, 2802, 388, 1528, 184, 3334, 79, 2752, 2326, 1531, 2772, 16, 46, 3323, 2410, 114, 26, 1268, 3258, 2732, 175, 1508, 146, 2602, 95, 1390, 656, 805, 1021, 2534, 3332, 92, 2460, 3228, 806, 1422, 601, 23, 1354, 1965, 127, 1988, 149, 1713, 1182, 2434, 1258, 3029]. Architectures [2289, 1997, 817, 1540, 2821, 834, 1693, 893, 1536, 1131, 2816]. Archive [1283, 1246, 2594, 1200, 1231]. Archiver [3236]. Archives [1231, 1240, 774]. Archiving [1230]. Arcs [561]. Area [3051, 1132, 3075, 1771, 2466, 2483, 3254, 97, 2862, 232]. Areas [439]. Argument [625, 3420]. Argumentation [569, 2958]. Argumentative [608]. Arguments [1179]. ARISTA [3184]. Arithmetic [269, 320, 2638, 834, 917, 2637]. Arity [2562]. Arizona [3256]. Arm [2274, 1856]. Arrangements [3128]. Array [1097, 1096, 109, 764]. Arrays [1629, 1038, 2370, 2016]. Arrhythmias [2941]. Art [3315, 205, 1149, 1923]. ArtemisExpress [393]. Artery [1853]. Arthroplasty [1791, 1790, 1779]. Arthroscopy [1870]. Articles [2097, 164]. Artificial [2418, 2945, 3458, 2509, 3428, 2882, 1606, 3187]. Artificially [3461, 2014]. ary [1101]. Ascent [691]. ASL [514]. Aspartame [3176]. Aspect [1735, 1739, 1740, 3075, 655, 176]. Aspect-Oriented [1740, 655]. Aspect-Specific [1735]. Aspects [581, 1536, 2331]. aspps} [615]. ASPs [2471]. Assembly [1347, 205, 2315, 2669, 2101, 3033]. Assertion [808]. Assess [2997]. Assessing [1121, 1790, 1929]. Assessment [2471, 1682, 796, 1119, 1168, 824, 795, 3001, 2101, 2994]. Assets [2076]. Assignment [1069, 78, 2355, 64, 912, 913]. Assignment-Based [2355]. Assist [3395]. Assistance [1774, 378, 406]. Assistant [1150, 2793]. Assistants [387]. Assisted [1888, 872, 2071, 1778, 2928, 1753]. Assisting [1471, 1908, 1847, 17]. Associated [1133]. Association [745, 2840, 753, 755, 3192, 111, 770]. Associations [772]. Associative [2992, 3002, 3419, 3000]. Associativity [507]. Assumptions [589, 3091]. Assured [2062, 2515]. Astrolabe [685]. ASUR [376]. Asymmetric [2951, 103]. Asymmetries [1811, 2899]. Asymptotic [2667, 1562, 3143]. Asynchronous [895, 3104, 3372, 924, 3367, 2748, 1998, 1996, 3365, 2385, 37, 2598, 876, 537, 3371, 2502, 326, 2727, 2164, 2014, 927, 3366, 3368, 2646]. Atlas [1930, 1796, 1815, 1828, 3169]. Atlas-Based [1796, 1828]. Atlases [1790]. ATM [106, 11, 2449, 1387, 1388, 2531, 34, 32]. Atom [2300]. Atom-Replacement [2300]. Atomic [2342, 686, 2353]. Atrial [2938, 2980]. Atrium [2938]. Atrophy [1844]. Atrophy-Behavior [1844].

Attached [2054, 1841]. Attack

[682, 2165, 1485, 3410]. Attackers [949]. Attacks [958, 2181, 2193, 2621, 2618, 201, 1387, 1388]. Attending [2902]. Attention [1057, 2903, 1021, 2900, 2904, 2901]. Attentional [2905]. Attract [851]. Attraction [875]. Attribute [1402, 1594, 702, 254, 1745, 1469, 1568, 1644, 1642]. Attribute-Efficient [702]. Attributed [2298]. Attributes [2756, 2603, 745, 1557, 1545, 2972, 731]. Attrition [1605]. Auction [2189, 2537]. Audience [1358]. Audiences [130]. Audio [1250, 3281, 2906]. Audio-Oculomotor [2906]. Audio-Visual [1250]. Audio/Video [3281]. Augmentation [2665]. Augmented [2952, 1854, 3433, 2847, 1867, 384, 1863]. Augmenting [1217]. Austrian [1231]. Authenticated [972, 2182]. Authentication [814, 2542, 2196, 2195, 2155, 110, 1411, 2034, 978, 1383, 1384, 98, 822, 815, 2192]. Authenticity [829]. Author [74, 151, 1229, 1284, 1319, 1342, 1380, 1415, 1441, 1477, 1541, 1625, 1654, 1679, 1703, 1728, 1749, 1848, 1934, 1978, 2009, 2044, 2072, 2080, 2111, 2187, 2204, 2246, 172, 2284, 2321, 2341, 2366, 2393, 2510, 2543, 2574, 2609, 2627, 2651, 2709, 2799, 2828, 2858, 2924, 2948, 3045, 3089, 3116, 3149, 3199, 3213, 3242, 3255, 3272, 204, 3290, 3313, 3342, 3363, 3373, 3394, 3469, 238, 248, 268, 313, 365, 418, 472, 526, 566, 623, 658, 690, 736, 780, 813, 837, 928, 947, 987, 1062, 1104, 1183]. Authorities [831]. Authority [1401]. Authorization [835, 2542, 3231, 3228, 822]. Autobahn [2039]. Autoconcurrency [553]. Automata [2028, 2011, 2446, 2468, 2130, 2041, 269, 2128, 924, 484, 2040, 2038, 1440, 536, 2112, 2015, 3109, 2380, 2013, 475, 2120, 2115, 761, 508, 2385, 2035, 2019, 2026, 2031, 3311, 916, 2018, 2034, 2012, 1430, 2119, 2707, 2042, 2124, 2114, 2023, 802, 2116, 2025, 2373, 2036, 524, 2014, 927, 1429, 2389, 2037, 2043, 2121, 1439]. Automated [832, 2630, 337, 2377, 303, 2553, 369, 700, 430, 3464, 1250, 113, 1849, 221, 1826, 1282, 2327, 1809, 1988, 1818, 1850, 2720, 2822, 1788]. Automatic [3103, 1280, 1275, 1032, 1052, 2081, 1198, 1813, 1707, 1363, 884, 2487, 1718, 2912, 1998, 324, 1846, 153, 1496, 2584, 707, 1751, 1359, 1856, 632, 1795, 2962, 3021, 145, 1834, 2504, 2947, 1841, 3182, 328, 1473, 2097, 1475]. Automatically [2236]. Automating [2571, 1303]. Automation [1205, 405]. Automaton [2033, 3113, 1546, 1428]. Autonomous [903, 3040, 2733, 3005, 3008, 3413]. Autoreducibility [490]. Availability [131, 2728, 2851, 2601]. Available [2530, 1146]. AVANTI [1160]. Average [2693, 2679]. Average-Case [2693, 2679]. Aviation [792]. Avionics [1989]. AVISS [340]. Avoiding [2581, 2697]. Aware [378, 402, 1990, 1211, 1964, 202, 2587, 2809, 31, 3387, 2053, 3268]. Awareness [379]. Awareness-Enhanced [379]. Away [909]. Axe [457]. Axes [1883]. Axiomatising [593, 599]. Axiomatizing [552]. aZIMAS [3200].

B2B [1371, 2433, 182]. Back [1128]. Backbone [3374]. Backbones [39]. Backfill [3244, 3246, 3248]. Backfilling [3247]. Background [2916]. Backoff [2466]. Backtracking [319]. Backward [351, 1025]. Bag [578]. Bagging [696]. Balance [1715, 3221]. Balanced [34]. Balancing [1496, 57, 3220, 1538, 2470, 1515]. Bandwidth [2052, 2530, 2518, 2483, 2438, 45, 913, 127, 2592]. Bar [415, 702]. Bargaining [2720]. Barter [170]. Bartering [2272]. Base [2955, 697, 1460, 498, 979, 157]. Based [2247, 1629, 2523, 1139, 2982, 2248, 2439, 2829, 559, 2734, 2530, 1950, 1352, 1935, 3035, 2984, 147, 1977, 2028, 2808, 970, 3303, 2521, 3335, 2468, 2175, 1691, 299, 2833, 1519, 2763, 968, 1255, 2813, 641, 339, 335, 2819, 2336, 2835, 1971, 2958, 413, 1784, 2783, 995, 1205, 2270, 1650, 207, 351, 1524, 251, 957, 2516, 828, 3261, 1550, 68, 2401, 1367, 277, 2630, 831, 3427, 2040, 22, 1211, 382, 72, 744, 2603, 311, 1456, 1839, 333, 1808, 1529, 177, 2987, 3423, 2770, 2758, 1683, 1796, 192, 3356, 2158, 2629, 609, 990, 1089, 2264, 129, 252, 180]. Based [1000, 205, 1897, 2507, 983, 1854, 1857, 1013, 2068, 791, 664, 1155, 2642, 2351, 3189, 3386, 991, 1862, 1825, 934, 2047, 2184, 1163, 2334, 1122, 1812, 3299, 372, 2621, 2088, 3465, 1222, 1552, 807, 1709, 2659, 3456, 1877, 925, 1919, 119, 1368, 3118, 3150, 410, 1702, 2721, 852, 2174, 228, 1911, 1614, 703, 66, 1765, 874, 2479, 2355, 83, 2908, 2409, 2911, 1351, 2580, 1588, 798, 2495, 2511, 3028, 1424, 2781, 2772, 138, 1253, 2162, 1847, 1939, 181, 2809, 1912, 2476, 2584, 69, 2412, 2090, 57, 2490, 3408, 113, 2462, 11, 1448, 1983, 707, 846, 1232, 169, 2201, 1816, 1054, 1982]. Based [10, 1462, 2849, 1055, 110, 2065, 710, 711, 2732, 2091, 1508, 53, 116, 94, 2805, 123, 146, 3271, 2539, 3215, 3226, 1666, 2585, 1844, 1558, 1856, 1317, 3044, 170, 1828, 3039, 712, 765, 223, 2417, 2305, 3326, 2251, 1281, 1886, 3193, 258, 663, 1913, 2512, 2053, 2317, 3467, 521, 2345, 2591, 2034, 2278, 2445, 1909, 1561, 2182, 3195, 2406, 1598, 384, 1030, 62, 99, 304, 2825, 2272, 1264, 1021, 2929, 1945, 803, 1206, 115, 1842, 54, 2238, 2070, 2640, 1952, 2615, 1636, 1355, 3228, 2764, 786, 815, 1672, 3024, 203, 271, 3197, 3114, 2514, 1745, 1201, 843, 657]. Based [613, 1969, 3338, 3393, 1423, 1622, 1582, 1623, 3244, 3328, 2403, 1917, 3353, 2784, 2920, 1635, 2414, 2901, 1234, 3230, 2311, 1645, 350, 2613, 1034, 2004, 858, 592, 2477, 3391, 75, 1591, 3392, 2811, 1577, 3033, 2255, 538, 3036, 3013, 2257, 1903, 1182, 2923, 2434, 194, 2191, 2202, 1924, 1899, 220, 210, 3260, 32, 2774, 2402, 144, 2919, 647, 1023, 2183, 1051, 182, 900, 2592, 2917, 2980, 1853, 3216, 187, 681, 2526]. Bases [569, 1472, 2774]. Basic [2379, 2300, 1581]. Basics [2131]. Basing [2157]. Basis [2934, 2993]. Bayes [2952, 2949, 2927, 3026]. Bayesian [1040, 907, 1725, 3418, 3419, 2879, 1378, 866, 1898, 2281, 1589, 3036, 1592, 2967]. BD [1360]. BDD [351]. BDD-Based [351]. BDDs [2556]. BDI [2551, 2554, 583]. Be [3411, 1981, 2286, 2862, 456, 1715]. Bead [2376]. Bead-Sort [2376]. Beam [3193]. Beam-Wise [3193]. BECN [11]. Behavior [2932, 2783, 1371, 577, 161, 839, 2013, 134, 43, 1844, 2513, 1270]. Behavior-Based [2783]. Behavioral [338, 1981, 2286, 2650]. Behaviors [2919]. Behaviour [903, 2288, 1212, 3421, 1607, 3223, 2306, 162, 185]. Behavioural [2647]. Beijing [1114]. Belief [794, 609, 3399, 570]. Beliefs [2488]. Benchmarking [186, 1520]. Bendixon [2565]. Benefit [796]. Beobots [2914]. Beowulf [1500]. Berica [1132]. Bernstein [2151]. Best [1136, 1132, 1612, 2988, 3159, 1164, 1153]. Better [493, 3189, 1628, 2669, 2093, 819]. Between [1393, 2518, 531, 2471, 2375, 540, 3414, 2711, 1740, 2769, 1846, 3181, 1169, 1912, 1663, 763, 1465, 3018, 2567, 2595, 3288, 65, 1012, 102, 1860, 2098, 3208]. Beyond [469, 452, 2007, 2153, 390, 171]. Bézier [1917]. Bias [923, 742, 1931, 862]. Biased [2652, 857]. Bibliographic [1247, 1235]. Bibliography [3320, 1626]. Bicriteria [1321]. Bidirectional [347, 3168]. Big [454]. Bigraphs [2287]. Bilayer [3058]. Bilingual [2083, 2094, 2097]. Bin [2667, 2405]. Binary [1732, 1657, 3075, 474, 506, 100, 3405, 1579, 847, 1622, 2260]. Binding [984, 2552, 2558, 1729]. Binding-Time [1729]. Binocular [2871]. Bioinformatics [1074, 1628]. Biological [275, 2835, 2874, 2878, 1544, 2882, 1706, 3155, 1727, 2983, 2897]. Biologically [2860, 2886, 2900, 1690]. Biology [1421]. Biomechanical [1775, 1905, 1837]. Biometric [814]. Biometrics [815]. Biometrics-Based [815]. Biomolecular [510, 1093]. Biopsy [1774, 1761, 1863]. Biosensor [1010]. Bipartite [1332, 2297, 3074]. Biplane [1799]. Bipolarity [2842, 3300]. Birthday [966]. Bisection [476]. Bisimilarity [563, 564]. Bisimulation [551, 550, 565]. Bit [3448, 834, 2189, 975, 917]. Bit-Serial [834, 917]. Bit-Slice [2189]. Bits [490]. Black [965]. Black-Box [968, 965]. Blended [1842]. Blind [2183]. Bloat [877]. Blob [2379]. Block [968, 2167, 2585, 950, 2467, 1066, 1510, 232]. Block-Cipher-Based [968]. Block-Matching [1510]. Block-Wise [1066]. Blocking [40, 2465, 2266]. Blocks [3365, 3264]. Blockwise [949]. Blockwise-Adaptive [949]. BlueGene [3245]. BlueGene/L [3245]. Bluetooth [1412, 2496, 2482]. BLX [2988]. BLX- [2988]. Body [1352, 3383, 994]. Bon [2450]. Bone [1352, 1778, 1773]. Bone-Based [1352]. Bonferroni [754]. Bonferroni-Type [754]. Book [3061]. Books [1282, 1266, 217, 1233]. Boolean [3123, 2639, 593, 982, 2180, 905, 501, 3129, 2695, 315]. Boomerang [2165]. Boosting [744, 184, 3126, 703, 756, 711, 717, 726]. Bootstrapping [2098]. Border [1096, 1159]. Both [631, 729]. Bottleneck [1007]. Bottom [3138, 2900]. Bottom-Up [3138, 2900]. Bound [482, 2180, 2839, 490, 501, 504, 2198]. Boundaries [2877, 2422]. Boundary [540, 2278, 2403]. Bounded [3105, 2688, 334, 2661, 1431, 2642, 497, 2159, 964, 1426, 1432, 2699, 2664]. Bounded-Degree [2688]. Bounded-Storage [2159]. Bounding [2361, 2947]. Bounds [2357, 3424, 2356, 2154, 2259]. Box [968, 3423, 965, 2947, 2998]. Box-Jenkins [2998]. Boxed [547]. Boxes [2179]. BPA [508]. BPM [1683]. BPP [563]. Brachytherapy [1782]. Brain [1811, 2943, 1817, 1789, 1813, 1785, 1857, 1807, 1797, 1815, 1805, 3384, 2885, 1816, 1795, 1810, 1860, 1788]. BRAINCHILD [1108]. Brains [1796, 2874]. Brainstem [2945]. Branch [3055]. Branch-and-Cut [3055]. Branching [482]. Branchpoint [1850]. Breaking [2193]. Breakpoint [1102]. Breakpoints [1086]. Brelstaff [2888]. Bridging [1740, 1160]. Brief [306]. Broad [2891]. Broadcast [2342, 2351, 951]. Broadcasting [2681, 3391, 85]. Brocade [661]. Bronchial [1849]. Brownian [1918]. Browsers [372, 1253]. Browsing [3278, 2334, 3279, 3192, 2283, 1951]. Büchi [3109, 2115, 2114]. Büchi-Automata [2114]. BuddyWeb [194]. Buffer [106, 2624]. Buffering [1660, 37]. Buffers [28]. Building [2207, 1735, 180, 2843, 1140, 1521, 3432, 1243, 1462, 367, 1895, 1018, 2214, 859, 3108, 2098, 3264, 1108, 1445]. Building/Using [1462]. Bulletin [3179]. Bully [132]. Burstiness [106]. Bus [1670, 47]. Business [3308, 2207, 2770, 2711, 1291, 1125, 452, 1375, 1159, 716, 2322, 1466, 2532, 2206, 1373, 2729, 1683, 1159]. Buy [819]. Buying [2474]. byte [1641]. Bytecode [1746]. Byzantine [3102, 2362, 2356].

C [2006, 2274, 356, 3106, 2487, 1702, 1856, 808]. C-Arm [2274, 1856]. C88 [514]. CA [2017, 2032]. CAA [2034]. Cable [1752]. Cable-Driven [1752]. Cache [2671, 1964, 3262, 122, 178]. Caches [2195, 1242]. Caching [2754, 2233, 89, 175, 3382, 176, 677, 3381, 646, 2055, 194]. CAD [206]. CafeOBJ} [929]. Calculating [1205]. Calculation [2375]. Calculator [1224].

Calculi [939, 532]. Calculus [2122, 548, 556, 3403, 244, 1431, 545, 938, 2563, 1433, 2125, 2567, 946, 592, 525, 2121]. Calculuses [2843]. Calibration [3463, 1856, 1753, 1049]. Call [2446, 2292, 25, 2465, 2480, 3238]. Call-by-Value [2292]. Calligraphic [211]. Callosum [1896]. Camera [990, 1750, 2279, 1863, 3084]. Cameras [1061]. Campaigning [1149]. Can [468, 2160, 3411, 1521, 1981, 2286, 667, 2862, 1715]. Canal [1895]. Cancer [1603]. Candidates [1068]. Canonical [362, 345]. Can’t [1405]. Canton [1174]. Cantor [2565]. Cantor-Bendixon [2565]. Capabilities [2619, 1596, 2996]. Capacity [2576, 14, 2622, 2060]. Capture [636]. Capturing [1447]. Car [3038, 864, 1026]. Car-Like [3038]. Card [59, 1155]. Cardiac [1830, 1825, 1882, 1828, 1826, 1834, 1838, 1837, 1827, 1780, 1831]. Cardinality [2668, 524, 2191]. Cardinality-Based [2191]. Cardiovascular [2936]. Cardiovascularity [2987]. Cardioversion [2938]. Cards [1176]. Care [2464, 3159, 3452]. Carefully [749]. Caribbean [1119]. Carlo [1792, 3215, 2459, 3041]. Carotid [1853]. carrier [2436]. Cartographic [219]. CAs [3422]. Cascade [3170]. Case [2247, 2955, 1362, 2877, 1736, 2169, 1873, 2705, 1452, 2693, 1132, 621, 1294, 3004, 1223, 3456, 2679, 3424, 398, 292, 393, 1449, 2251, 955, 2272, 718, 675, 1952, 529, 3395, 3024, 635, 431, 216, 253, 2255, 2257]. Case-Based [2247, 2251, 2272, 2255, 2257]. Cases [3097, 873, 1122]. Casting [2676]. Catalog [2336]. Catalogues [1247]. Categories [1662, 395]. Categorization [2909, 2885, 1378, 2859, 3180]. Categorizing [2076]. Category [1234]. Catheter [1772]. Causal [128, 695, 3166, 3184, 584]. Causality [2295, 561]. Causation [598]. Cause [433]. CB [1159]. CB-BUSINESS [1159]. CBC [949]. CBR [1443, 2959]. CCS [515, 531]. CDMA [100, 2463, 99, 91, 2436, 2467]. CDPD [3270]. Cell [2887, 1773, 2867]. Cells [2861]. Cellular [895, 2030, 2461, 2022, 2033, 2024, 2028, 2011, 2446, 924, 2040, 2038, 2015, 2380, 2495, 2013, 2385, 2035, 2019, 2026, 916, 2018, 2034, 2012, 2996, 2042, 2023, 2025, 2036, 2014, 927, 2389, 2037, 2043]. Censorship [669]. Centered [398, 2884, 419]. Centers [448, 2172]. Central [2157, 1008]. Centralised [1132]. Centralization [1106]. Centric [1347, 2501, 141, 1369, 152]. cerevisiae [1100]. Certificate [672]. Certificates [1401, 1402]. Certification [811, 812]. Certifying [2570]. Cervantes [1277]. CFS [678]. Chain [1204, 1113]. Chains [549, 3372, 706, 1214, 2610]. Challenge [476]. Challenges [3273, 270, 2285, 3301, 792, 3375, 1979, 1344]. Chance [298]. Change [1014, 3399, 451, 1461, 1621, 2977, 2789]. Change-Detection [2977]. Changes [2220, 1474, 1748]. Changing [3048, 697, 454]. Channel [2468, 1529, 78, 64, 123, 517, 913, 2436, 2467, 63, 2049]. Channel-Based [1529]. Channels [2182]. Chaos [867]. Chaotic [2017, 2421]. Chapter [3319, 3314, 3316, 3318, 3315, 3317]. Character [381, 1641]. Characterisation [3163]. Characteristic [2170, 381, 971]. Characteristics [174, 230]. Characterization [1896, 161, 309, 1843, 2182]. Characterizations [1681]. Characterizing [160, 1208, 2672, 2220, 723, 523]. Charge [845]. Charging [2542, 2533, 2534, 203]. Charts [415, 539]. Cheaper [538]. Checkable [794]. Checker [354, 327, 357, 2109, 353]. Checking [535, 360, 338, 3105, 321, 331, 325, 1682, 334, 816, 3094, 2196, 348, 277, 323, 2630, 342, 3403, 252, 748, 2642, 2638, 324, 807, 67, 346, 1985, 314, 336, 1438, 542, 3293, 344, 361, 543, 332, 2198, 1430, 276, 1432, 326, 802, 345, 808, 350, 2634, 1439, 2646]. Checkpointing [147, 3377, 138]. Checks [983]. Chek [354]. Children [2940]. Chinese [2368, 3177, 1199, 2090]. Choice [1549, 2935, 1142]. Choices [1301]. Cholangiography [1852]. Cholangiopancreatography [1933]. Chomsky [486]. Chomsky-Like [486]. Choose [749]. Chop [543]. Chord [678]. Chord/CFS [678]. Chordal [2666, 2358]. Chosen [976]. Chosen-Plaintext [976]. Chromatic [1332]. CHRs [274]. Church [1036]. Ciliates [2315]. CIM [2330]. Cipher [968]. Ciphers [980, 2167, 950, 981]. Circle [434]. Circuit [3367, 2189, 2151, 917, 3368, 2646]. Circuits [3372, 501, 3371, 2631]. Circulants [2494]. Circular [3047]. Citation [160, 1154]. CITI [1162, 1161]. Cities [1161]. Citizen [1154, 1172, 1145, 1150, 1176, 1142, 1134]. Citizens [1156]. Citizenship [1111]. City [403]. CJD [1814]. Claims [3452]. Clash [433]. Class [2833, 626, 2949, 982, 1358, 2485, 701, 854, 59, 1420, 537, 2019, 713, 3160]. Class-Based [2833]. Class-Conditional [2949]. Classes [2656, 250, 3128, 2297, 2935, 3135, 756, 3130, 31, 763, 3169, 3140]. classical [2863]. Classification [3443, 3183, 742, 207, 1813, 1785, 747, 749, 699, 2907, 1063, 1271, 2951, 3145, 3161, 1613, 1605, 3464, 2928, 2918, 1847, 3127, 3405, 1595, 2092, 764, 3147, 3421, 2953, 2417, 713, 3396, 716, 1598, 1835, 1900, 767, 1620, 1320, 2281, 2267, 2956, 1576, 1234, 1793, 2268, 728, 2255, 729, 1903, 3436, 1584]. Classifications [709]. Classificatory [1623]. Classifier [2992, 890, 891, 2927, 2015, 700, 894, 1618, 893, 2980]. Classifiers [2952, 3183, 697, 701, 733, 777, 1582, 1589]. Classifying [2424, 2902]. Clause [694, 3402, 2964, 3444]. Cleaning [352]. Clearing [2703]. Clearinghouses [1677].

Click [2896]. Client [177, 196, 1607, 423, 183, 3263]. Client-Based [177]. Client-Server [3263]. Client-Side [183]. Clients [822]. Climbing [3138]. Clinical [1452, 2934, 1781, 1866, 771]. Clip [1155]. Clipping [99, 397]. Clique [2694, 2656]. Clique-Width [2656]. CLM [1345]. Clock [1084]. Clocks [2295]. Closed [1761]. Closeness [1578, 1580]. Closing [2947]. Closure [259, 1514]. Cloth [1500]. Cloud [2953]. Clouds [1915]. CLP [277, 310, 288, 303]. Cluster [1977, 1512, 1022, 1500, 2733, 1531, 1534]. Clustered [3066]. Clustering [3407, 2829, 2030, 2261, 1101, 2970, 1958, 1013, 926, 751, 1007, 925, 3419, 3175, 1575, 712, 765, 915, 3195, 3168, 725, 727, 85, 1951, 1571]. Clusters [1539, 3377, 189, 1571]. CM [1526]. CMA [878]. CMA-Evolution [878]. CMAC [3423]. CMAC-Based [3423]. CMM [439]. CMMI [433]. CMPC [122]. Co [868, 1817, 3252, 3282, 807, 861, 865, 889, 3009]. Co-allocation [3252]. Co-evolutionary [861, 865]. Co-evolving [889]. Co-generative [807]. Co-occurrence [3282]. Co-operating [868]. Co-operative [861, 865]. Co-ordination [3009]. Co-registration [1817]. Coach [456]. Coaching [2407, 425]. CoAgnostic [3125]. Coalgebraic [556]. Coalition [2010, 599, 3025]. Coalitions [1372]. COBA [609]. Code [797, 3324, 389, 3221, 3201, 1225, 697, 449, 2334, 358, 430, 100, 2198, 2463, 3222, 3205, 1297, 1020, 2154, 99]. Coded [2988, 852, 914, 2467]. Coder [123]. Codes [2346]. Coding [2589, 2408, 1721, 2424, 1712, 689]. Coevolution [862]. Cognitive [2989, 3283]. Coherence [158]. Cohesion [2455, 2485]. Coinductive [939]. Collaboration [1345, 3229, 2803, 862]. Collaborations [1312, 680, 1438]. Collaborative [2740, 640, 1372, 2934, 1647, 707, 710, 2732, 384, 1565, 155, 2727, 2726, 194, 3267].

Collations [1277]. Collecting [2731]. Collection [282, 281, 1540, 1994]. Collections [1243, 1936, 1270]. Collective [2013, 2032, 1520]. Collision [1875]. Colon [1920]. Colonies [2269]. Colonoscopy [1927, 1920]. Colony [2030, 922, 606, 2869]. Color [219, 1030, 1025, 1020, 2947]. Color-Based [1030]. Colorable [488]. Colorimetric [2960]. Coloring [1323, 488, 1330, 1332]. Colorings [2491]. Colour [1839, 3417, 474, 2277]. Colouring [3379, 474]. Combat [3453]. Combination [1442, 1682, 2170, 2912, 375, 1933]. Combinations [2556, 1194]. Combinative [1833]. Combinatorial [2985, 1099, 900]. Combined [1208, 1756, 123, 28, 1905, 1025]. Combiner [981]. Combining [3183, 321, 692, 1042, 700, 2089, 336, 897, 1823, 2998, 1871, 2484, 3111, 1077, 1668]. Command [1564, 417]. Commerce [2958, 2211, 2826, 1377, 1376, 1374, 2322, 3259]. Commercial [2099, 2619]. Commit [139]. Commitment [1455, 2472, 984, 2158, 1353]. Commitments [2163, 983]. Committed [1918]. Committees [3164]. committing [955]. Common [1073, 3137, 812, 2786, 2613, 433]. Communicating [557, 3107, 370, 90]. Communication [2581, 1519, 3349, 2719, 3261, 2361, 1529, 1528, 3386, 2076, 2762, 119, 379, 2679, 83, 2383, 2809, 124, 3400, 3387, 2481, 2497, 3332, 1536, 2764, 1526, 179, 2388, 2164, 2771, 1534]. Communication-Optimal [2679]. Communications [2461, 2078, 1165, 2409, 3448, 15, 2815, 1138, 3098]. Communities [192, 153, 2522, 1129, 2723, 165]. Community [1243]. Community-Led [1243]. commutative [2559]. commutativity [579]. Comorphism [521]. Comorphism-Based [521]. Compact [1752, 1650, 3062, 2358, 1652, 1762]. Compactness [3140]. Company [2100, 2101]. Comparative [76, 177, 2337, 2229, 1900, 1079, 2957, 900]. Compare [2359]. Compare-and-Swap [2359]. Comparing [895, 3035, 3007, 1193, 756, 678]. Comparison [896, 2714, 2170, 2478, 3004, 1647, 3022, 789, 1911, 752, 1605, 3464, 1665, 2953, 1719, 1913, 2253, 2929, 2655, 1838, 188, 1196, 689]. Comparison-Addition [2655]. Comparisons [2997, 2638]. Compass [2844]. Compatibility [2649, 348]. Compatible [2582, 1331, 1762, 1764, 1760]. Compensation [2279, 1928, 1510]. Competition [1044]. Competitive [2471, 487, 3439, 2693, 1337, 1721]. Competitors [810]. Compilation [572, 3103, 1995, 1993, 2570, 2097]. Compiler [2006, 942, 2005]. Compiling [169, 2648]. Complement [1331]. Complementary [851, 1961]. Complementation [2115, 2116]. Complete [550, 2657, 507]. Completeness [478, 279]. Complex [426, 1682, 989, 2839, 2959, 681, 336, 2742, 708, 1257, 3042, 2467]. Complexity [2547, 1724, 2666, 2958, 2589, 2367, 2599, 577, 3082, 2555, 578, 2685, 3161, 502, 3131, 3130, 508, 1575, 516, 3311, 1340, 955, 2655, 2124, 2390, 2678]. Compliant [1212, 2194, 821]. Component [1714, 2943, 1519, 3348, 339, 347, 1211, 2770, 746, 1000, 1491, 2479, 1424, 2772, 3450, 222, 732, 776, 1711, 3222, 3228, 3292, 1745, 3328, 217, 2216, 1182, 144]. Component-Based [1519, 339, 1211, 1424, 1745, 3328, 144]. Component-Expert [1491]. Component-Oriented [3348]. Componentization [1272]. Components [2247, 1312, 1518, 1276, 2551, 2795, 2582, 161, 2773, 1521, 3223, 795, 2647, 639]. Composable [1735, 984, 1976]. Compose [1741]. Composed [2325]. Composing [3364, 2752]. Composite [2731, 1788]. Composition [2557, 853, 3351, 2765, 3341]. Compositional [2955, 2548, 1986, 3110, 2427, 2199, 3191]. Compositions [2347]. Compound [3427, 787, 3071, 2894]. Comprehensible [3165]. Comprehension [1271]. Comprehensive [1647, 1502]. Compressed [2442, 1195, 1929, 1641]. Compressibility [158, 3195]. Compressible [1560]. Compressing [361]. Compression [1977, 1955, 226, 1798, 1185, 3305, 1719]. Computable [536]. Computation [974, 479, 1633, 1512, 960, 2361, 619, 959, 2334, 2343, 2379, 587, 506, 2383, 2385, 3468, 2519, 2370, 1514, 1093, 2476]. Computational [2188, 692, 2547, 1478, 3189, 2315, 502, 2382, 3005, 3387, 1426, 2996, 2867, 2936, 1899, 3431]. Computations [2375]. Computed [2276]. Computer [1783, 903, 781, 3283, 1888, 2891, 1832, 991, 2621, 782, 1351, 1778, 2928, 1626, 1887, 1883, 2374, 1690, 1397, 1398, 2303]. Computer-Aided [1351]. Computer-Assisted [1778]. Computerised [2000]. Computers [84, 90, 2390]. Computing [3053, 2877, 1519, 3052, 1497, 1022, 1498, 3224, 1211, 2369, 3219, 1478, 243, 3250, 2076, 1481, 1486, 3266, 135, 2315, 498, 499, 3301, 137, 2381, 1483, 2382, 3262, 3220, 1579, 292, 2728, 2074, 3360, 1381, 2386, 400, 810, 3212, 3238, 2388, 1522, 3214, 971, 3235, 3264, 1584, 85, 1686, 2501]. Concast [3338]. Concast-Based [3338]. Concave [1977]. Concealment [2414]. Concentration [1720]. Concept [2248, 630, 3443, 1572, 3128, 2379, 1501, 1351, 3130, 3400, 3395]. Conception [410]. Concepts [1309, 1656, 2210, 3459, 707, 3345, 2456, 735, 779, 1166]. Conceptual [575, 2215, 2225, 2836, 2242, 2791, 2711, 2243, 2934, 2208, 2210, 2219, 2779, 1460, 2230, 1662, 633, 24, 1943, 627, 2778, 2961, 2239, 2311, 2213, 2205]. Concerning [2957]. ConCert [3224]. ConChord [672]. Concise [474]. Concluding [1413]. Conclusion [3319].

Conclusions [8]. Concrete [3349, 1747]. Concurrency [541, 3369]. Concurrent [958, 275, 263, 2713, 2699, 2641, 3093]. Condensed [755]. Condition [2355, 3437, 2345, 944]. Condition-Based [2345]. Conditional [580, 2949, 1586, 1587, 598, 701]. Conditionally [1190]. Conditioner [2516]. Conditions [1574, 2251]. Conductive [2186]. Conference [1318]. Conferences [2592]. Conferencing [1967, 127]. Confidence [1048, 852, 719, 722, 3143]. Confidentiality [2190, 1413, 2199]. Confidentiality-Preserving [2199]. Configurability [810]. Configurable [2811]. Configural [2922]. Configuration [1472, 1289, 864, 17, 86]. Configurations [1599, 1672]. Configuring [2771]. Conflict [2630, 2837, 1557]. Conflicting [569, 1219]. Conflicts [3307]. Confluence [363, 2298]. Conformance [650, 3092, 3291, 1435, 326]. Confusion [784]. Congestion [2052, 2529, 48, 11, 51, 2070, 2069, 125, 2536, 2060]. Congruent [3092]. Conjecture [2663]. Conjugate [2409]. Conjunction [2791]. Conjunctions [504]. Conjunctive [250, 2855]. Connected [960, 1617, 2743, 1341]. Connection [2610]. Connectionist [3016]. Connections [43, 2202]. Connectivity [2684, 39, 368, 2665, 1336, 1806]. Conquer [1513]. Conscious [1218, 179]. Consensus [2363, 2357, 1697, 1557, 3256, 3012]. Consequences [1082, 1301]. Conservative [1511]. Conserving [2380]. Consideration [890, 1462]. Considerations [811, 2505, 683, 3263]. Considering [56, 2597, 2593]. Consistency [2831, 323, 2645, 609, 1302, 1573, 3132, 3460, 933, 2851, 2493]. Consistency-Based [609]. Consistency-Preserving [1302]. Consistent [128, 2834, 599, 1810, 395, 1669]. Consolidating [1077]. Constant [984, 99, 769]. Constant-Memory [769].

Constellations [2500]. Constituency [1108]. Constrained [1323, 510, 3292, 35]. Constraint [479, 2545, 2830, 3303, 2668, 275, 2976, 277, 311, 1189, 610, 3433, 3145, 997, 3432, 910, 57, 223, 933, 2260, 2395, 840]. Constraint-Based [3303, 277, 57, 223]. Constraints [269, 992, 2599, 2770, 1725, 1300, 934, 2558, 2425, 2302, 279, 933, 3467, 2564, 855, 276, 2437, 274, 3447, 600, 328]. Constructing [1440, 1235]. Construction [1352, 3317, 473, 1090, 982, 2350, 2773, 1431, 2179, 500, 2846, 519, 1618, 973, 1349, 107, 1837, 2216, 3091]. Constructions [968]. Constructive [910]. Constructivism [891]. consulting [1161]. Contact [1777]. Containment [251]. Contemporary [1283, 823]. Content [1950, 1280, 1275, 3261, 1367, 687, 177, 192, 2578, 372, 2590, 2591, 1744, 2601, 1264, 1952, 2961, 1672, 684, 1645, 646, 178, 3298, 397, 2541]. Content-Based [1950, 3261, 1367, 2591, 1264, 1952, 1672, 1645]. Contention [2347]. Contention-Resolution [2347]. Contents [79, 1368, 3376]. Contest [3086, 236, 234]. Context [2829, 402, 995, 1586, 1211, 2234, 1930, 2488, 3268, 2724, 2809, 374, 2739, 399]. Context-Aware [402, 1211, 2809, 3268]. Context-Specific [1586]. Contexts [585]. Contextual [402, 1585, 2724, 2093, 217, 2977]. Contingency [725]. Continual [1187]. Continuity [2867]. Continuous [3407, 549, 2529, 2599, 3037, 1213, 122, 2954, 2930, 1852]. Continuous-Time [549]. Continuous-Valued [3407]. Contour [2881, 2863, 2864, 1004, 1898, 2866, 1023]. Contours [1823, 1822]. Contract [2218]. Contracting [2060]. Contrast [2861, 1834, 1787, 1899]. Contrast-Enhanced [1787]. Contribute [468]. Contribution [1497, 153]. Control [2052, 2058, 1299, 2529, 273, 3335, 2446, 1279, 2756, 2719, 801, 2396, 828, 2000, 3416, 2715, 3148, 241, 2001, 2328, 1854, 1408, 1409, 3038, 807, 2299, 2533, 861, 112, 17, 11, 1983, 26, 2201, 51, 53, 61, 1091, 1338, 52, 1426, 3404, 366, 2519, 809, 2070, 1546, 2069, 2480, 2447, 2419, 125, 405, 546, 2477, 1988, 75, 2191, 2536, 3431, 2625]. Control-Dominated [1299]. Control-Theoretical [2519]. Controllability [2043]. Controlled [2190, 2109, 1565, 3395]. Controller [3416, 2588, 894, 3028, 1423]. Controllers [537, 3034]. Controlling [2429, 1766, 867]. Controls [821]. Convergence [41, 3258, 3421]. Convergency [3442]. Convergent [691, 996]. Conversations [2800]. Conversion [226, 1232, 3395]. Convex [2705, 3310, 2990]. Convolution [1925]. Cooking [2750]. Cool [2323]. Cooperating [2869, 353]. Cooperation [193, 1456, 1209, 1116, 2475]. Cooperative [672, 2735, 2361, 2721, 134, 1210, 112, 175, 2716, 1959, 2737, 13, 2278, 2272, 675, 136, 674, 1963, 2055, 2725, 862, 165]. Coordinate [1829]. Coordinated [147]. Coordinating [3251, 2825, 2722]. Coordination [3007, 199, 2715, 2713, 186, 2919]. Coordinator [132]. Coplanarity [992]. Coprocessors [2192]. COPS [2064, 2335]. Copy [163, 2415, 2489]. Copying [282]. COR [2978]. CORAS [791]. CORBA [1309, 2751, 2813, 2767, 2758, 2748, 2759, 2056, 66, 2747, 1693, 3222, 2764, 2426, 2757]. CORBA-Based [2758, 66, 2764]. Core [2519, 3388, 1288, 1568]. Core-Stateless [2519]. Corners [3069, 1003]. Coronary [1882, 2282, 1824]. Corpora [153, 2094]. Corporate [1960]. Corpus [1203, 1896, 2096]. Corral [3324]. Correcting [2346]. Correction [1778, 214, 1899]. Correctness [2635, 2773, 935, 1375]. Correlated [2436, 770]. Correlates [1844]. Correlation [2616, 2614, 2615, 2202]. Correlations [2402]. Correspondence [2911, 1027]. Correspondences [2712, 2797].

Corresponding [1019]. Cortex

[2866, 3431]. Cortical

[1802, 2872, 1804, 2867, 1915, 1623]. Cosine [768]. Cost [2244, 1174, 756, 57, 2492, 713, 2100, 2710, 2437, 1049, 3160]. Cost-Sensitive [756, 713, 3160]. Costs [1205, 582, 2685, 179, 2063]. Coteries [131]. COTS [426, 800, 640, 1700, 2769]. COTS/SOUP [800]. Counter [320, 2661, 2629, 322, 3292]. Counter-Constrained [3292]. Counter-Example [2629]. Counters [2516]. Counters-Based [2516]. Counting [2691, 3058, 2170, 518]. Country [1137, 1132, 2660]. Counts [2707]. Couple [2988]. Coupling [3222]. Course [3023]. Covariant [625]. COVAX [1283]. Cover [1324, 1329, 1334, 2674]. Coverage [797, 3017, 1590]. Covering [1213, 2458, 2667, 3111]. Covers [1325]. Cow Suite [1314]. CPM [2714]. CPS [938]. Cramer [2152]. Crawling [1620]. Crawls [1236]. Creak [1298]. CREAM [1473]. Create [1401]. Created [1608]. Creating [1450, 1402, 3265, 1200]. CREAtion [1473]. Creativity [441]. Credentials [952, 818]. Creole [3187]. CRESS [3100]. Criteria [2336, 812, 918]. Criterion [712]. Critical [1688, 428, 788, 1689, 2309, 795, 2323]. Criticality [800]. Cross [3058, 2715, 2084, 411, 1159, 1038, 1194, 1810, 2097]. Cross-Border [1159]. Cross-echo [1038]. Cross-Language [2084, 1194]. Cross-Lingually [2097]. Cross-Modal [411]. Cross-Organizational [2715]. Crossconnects [38]. CrossGrid [1494, 1489, 1490, 1491]. Crossing [2700, 1808, 1327, 3057, 3072, 3071, 3063]. Crossover [851, 2988, 852, 853, 855]. Crowd [2031]. Crowds [677]. CRT [963]. CRT-Exponent [963]. Cryptanalysis [2166, 980, 2167, 963, 2161]. Crypto [515, 2185]. Crypto-CCS [515]. Crypto-integrity [2185]. Cryptographic [982, 2157, 1197, 2193, 533, 2409, 560]. Cryptography [2184, 2168, 2421]. Cryptology [969]. Cryptosystem [2410, 2152]. Cryptosystems [970, 831, 2162]. CSP [2966, 2259]. CSPs [309, 3433]. CT [1029, 1847, 1849, 1858, 1910, 1906, 1841]. CTA [1908]. CTL [542, 528, 352]. Cube [2658]. Cubes [2191]. Cubic [1872]. Cues [2889, 406]. Cull [1801]. Cultural [1276, 1278]. Culture [1283, 404, 433]. Current [2458, 1190, 2451, 213, 3254, 1495]. Curvature [3174]. Curve [2659, 106, 122]. Curved [3414]. Curves [1053, 2170, 2703, 971, 1010]. Customer [452, 1605]. Customizable [2765]. Customized [2243, 1974]. Customizing [1305]. Cut [943, 3055, 715]. Cuts [1871]. Cutting [1878, 2660]. CVC [353]. Cycle [1325, 2319, 1607, 1957]. Cycles [716]. Cylindrical [1932].

D [1911, 1913, 1906, 2439, 1799, 1861, 1032, 2404, 1022, 2274, 992, 2945, 1498, 1035, 205, 1897, 1894, 3003, 908, 3310, 3059, 1868, 1798, 1893, 1911, 1881, 1847, 1849, 1816, 1054, 1055, 1751, 1823, 1902, 1882, 1828, 1913, 3042, 1909, 1867, 2254, 1907, 1928, 3041, 1917, 1020, 1034, 1820, 1906, 403, 1824, 1892, 3422]. D- [1911, 1913]. D-CT [1849]. D-Image [908]. D-Reconstruction [1032]. D-Visualization [1798]. D/ [1909, 1906]. DAC [1513]. DACA [2058]. DAD [2581]. Daffodil [1274]. DAGs [2967]. DAMPVM [1513]. DAMPVM/DAC [1513]. Dark [2893, 432]. DASG [1631]. Data [737, 2344, 3407, 2925, 1962, 739, 154, 1935, 1732, 2735, 2943, 1443, 1830, 1101, 1817, 3374, 1228, 642, 1350, 2834, 742, 1971, 1657, 286, 995, 199, 1658, 255, 2237, 251, 3211, 1717, 2231, 2233, 2933, 1218, 744, 2838, 1035, 1813, 1707, 2963, 1225, 252, 787, 1491, 1379, 2236, 1705, 748, 1013, 1807, 1894, 527, 3107, 1843, 2478, 751, 611, 2378, 3266, 206, 1612, 1642, 2672, 1922, 2364, 3167, 2442, 680, 3301, 2718, 2847, 1940, 1246, 141, 2398, 2890, 3119, 3151, 560, 2577, 2850, 1595, 1204, 2400, 763, 1359, 764, 2223, 2753, 686, 2728, 1257, 1118, 1596, 1910, 1886, 3193]. Data [2737, 254, 2229, 1102, 2616, 2851, 2445, 1651, 91, 54, 1542, 1834, 1599, 2241, 2232, 1794, 1320, 1355, 822, 3295, 1949, 1173, 2653, 613, 1270, 2853, 2739, 1547, 1965, 249, 3381, 3133, 1970, 2477, 3101, 2213, 3240, 729, 1677, 2191, 3093, 1406, 1407, 3452, 1583, 1933, 1755, 1608, 1603, 182, 2224, 1668, 1644, 3173, 685, 1256]. Data-Driven [199]. Data-Handlers [1491]. Database [1950, 1352, 2780, 266, 2244, 1656, 2933, 1857, 261, 260, 1185, 3225, 2479, 759, 140, 24, 139, 2203, 1545, 1360, 3295, 2743, 2443, 450, 152, 2973, 2214, 2216, 1592, 3263, 2855]. Database-Centric [152]. Databases [1955, 273, 2835, 1570, 264, 270, 1068, 2733, 259, 3415, 2558, 752, 278, 1190, 2429, 2845, 1976, 2476, 169, 2848, 579, 1946, 1364, 1947, 305, 589, 766, 2793, 1944, 3155, 1948, 3297, 769, 1188, 2856, 771, 772, 1951, 2832, 1669]. Datalog [512]. Datascape [3170]. DataScript [1732]. Dataset [3163, 1571]. Datasets [1039, 3164, 762]. Datatype [2777, 2798]. Date [1036, 818]. DAWNING [1525]. DAWNING-3000 [1525]. Day [447]. DB [265]. DBLP [1626]. DBMS [1952, 1184]. DBMSs [1949]. DDH [985]. Dealing [1344]. Death [434]. Debris [2038, 3449]. Debugger [20]. Debugging [3404, 312, 3378]. Decentral [164]. Decentralised [3009]. Decentralized [2285, 2722, 3214]. Decidability [2201, 560, 563, 588, 2123, 1427]. Decidable [540, 485, 537, 544]. Deciding [2639, 2638, 329]. Decision [378, 551, 3412, 2643, 2974, 3189, 1447, 1708, 1573, 1553, 1574, 704, 2394, 705, 506, 2251, 1597, 879, 1564, 723, 3114, 1576, 2571, 3406, 3453, 771, 3398, 1568, 1668]. Decision-Making [3398].

Decision-Theoretic [378]. Decision-Tree [705]. Decisions [713]. Decks [2067]. Declarative [286, 2962]. Decomposing [2290, 1675]. Decomposition [2976, 3306, 1595, 1056, 1623, 3368, 2634]. Decreasing [1994]. Dedicated [2686, 1496]. Deduction [2964, 2567]. Deductions [3462]. Deductive [1208, 273, 278, 579, 305, 589, 528]. Deep [1857, 1771, 1770]. Default [597]. Defeasible [295, 3465, 2775]. Defect [1029, 3450]. Defects [1845]. Defense [2193, 1485]. Definability [2555, 1543]. Defined [2444, 97]. Defining [1304, 2795, 1975]. Definite [3014]. Definition [417]. Definitional [943]. Definitions [1379, 2441]. Deformable [1875, 1805, 1009, 1782, 1816, 1890, 3036]. Deformation [1896, 1500, 1825, 1905, 1876, 1810]. Deformities [1778]. Degradation [1697]. Degree [2688, 2491, 1604]. Delaunay [2659]. Delay [2611, 3386, 1220, 56, 45, 2496, 2596, 2437, 2202]. Delays [179]. Delegate [1403, 1404]. DELI [2001]. Deliberative [1144, 1143]. Delineation [1829]. Delivering [142, 2541, 10, 2333]. Delivery [2332, 1130, 2342, 1971, 687, 22, 177, 2986, 1109, 1106, 1127, 19, 18, 3376, 178]. Delta [702]. Delta-Bar-Delta [702]. Demand [2471, 2544, 2576, 1964, 2704, 2541, 2703, 33, 3381]. Demarcation [1801]. Dementia [1832, 3464, 1844]. Democracy [1178, 1144, 1151, 1143]. Democratic [1112]. Demonstration [1279, 3085]. Deniability [832]. Deniable [978]. Denial [1383, 1384, 815, 3410]. Denial-of-Service [1383]. Denoising [2440, 1012]. Dense [1033]. Densities [65]. Density [490, 1074, 1711, 726, 2014, 927]. Dependability [1696, 781, 1702, 782, 1693, 810, 783, 1991, 1699, 785]. Dependable [1424, 1686]. Depended [2872]. Dependencies [2218, 2551, 753, 1557, 1561, 3292]. Dependency [757]. Dependent [239, 995, 1212, 591, 2946, 1425, 1059, 85, 2942]. Deploying [2587, 184, 2752, 668]. Deployment [3327, 1225, 1169, 118, 2765, 3328]. Deployments [421]. Depth [3094, 997, 143]. derivable [743]. Derivation [1737, 511, 328]. Derivational [3396]. Derivatives [3176]. Derived [1545, 1810]. Deriving [1314, 2095, 774]. Descartes [3142]. Descendants [940]. Describe [3098]. Described [1965]. Describing [2302, 1287]. Description [575, 573, 2712, 636, 2557, 2772, 2441, 3014, 2884, 942, 3114, 2572, 1988]. Descriptions [2420, 1285, 221, 2648, 3395]. Descriptive [2095]. Descriptors [2283]. Design [1680, 932, 1687, 634, 3335, 355, 1094, 402, 1527, 1656, 1301, 1174, 892, 3367, 3457, 2933, 72, 1700, 2768, 1064, 2748, 376, 2582, 1740, 2487, 156, 449, 1376, 206, 1877, 1525, 70, 1296, 3365, 60, 1096, 864, 3459, 25, 112, 2747, 1982, 1336, 2065, 398, 87, 116, 2753, 836, 2481, 1360, 655, 384, 62, 2534, 115, 2449, 419, 2002, 2755, 678, 1863, 1526, 2647, 648, 2311, 1313, 3391, 1348, 3366, 3368, 178, 32, 3263, 399, 395, 2646]. Designated [3069]. Designing [2955, 2810, 1455, 3370, 3336, 1022, 2245, 2843, 1980, 130, 1100, 393, 419, 377, 2973, 1272, 1233]. Desires [2488]. Desk [1132]. Desktop [1274]. Detail [3317]. Detect [297, 67]. Detecting [2612, 1586, 2611, 1927, 2590, 2950, 3172, 3410, 1003, 1845, 2610, 2625]. Detection [1875, 738, 1811, 1714, 1014, 2338, 1716, 2927, 1035, 2619, 2357, 2611, 1739, 1029, 1897, 1807, 2071, 2360, 2975, 2294, 3002, 2863, 2622, 2898, 2860, 997, 2864, 2617, 88, 2918, 757, 3450, 112, 2490, 113, 1461, 1001, 2716, 2203, 2623, 2626, 1898, 3340, 408, 2895, 2862, 1078, 1794, 1929, 2403, 2352, 2613, 2624, 2977, 3204, 2947, 1841, 2789]. Detections [2678]. Detector [2502].

Detectors [2365]. Determinacy [2117]. Determination [2930]. Determining [1566, 3380]. Deterministic [848, 2575, 853, 2890]. Determinization [2114, 2116]. Detour [103]. Develop [1309, 1379]. Developing [3316, 2645, 3384, 3256, 2710, 1668]. Development [1734, 1752, 473, 428, 1784, 1488, 3457, 1683, 1513, 462, 2507, 826, 653, 2420, 457, 1770, 1285, 1756, 2617, 1316, 401, 1217, 108, 1757, 124, 2505, 3006, 61, 2805, 432, 1779, 419, 799, 1310, 657, 2568, 2785, 1760, 1780, 649, 2213, 3422]. Developments [2546, 1855, 3257]. Deviation [1084]. Deviations [1797]. Device [1774, 411, 369, 386, 391, 1254, 23, 395]. Device-Independent [391]. Devices [3210, 413, 79, 373, 396, 366, 383, 405, 210, 417]. DH [985]. DH-DDH [985]. DHT [681]. DHT-based [681]. DHTs [662]. DiaGen [2307]. Diagnose [1619]. Diagnosing [1832]. Diagnosis [2966, 2975, 1814, 1766, 3023, 1760, 1429, 3025, 3029, 2980]. Diagnostic [2619, 1616, 771]. Diagram [212, 1315, 156, 2696]. Diagrams [1655, 1802, 2643, 506, 2228, 2307, 3088, 3114]. Dialogs [79]. Dialogue [3016, 3020]. Diameter [1333, 1335]. Diaries [369]. Diary [368]. DIAS [1268]. Dictionary [2181, 1185]. Diff [2789, 2060]. Diff-Serv [2060]. Difference [2375, 2585, 993, 3114, 531, 458]. Difference-Histograms [993]. Differences [2769, 763]. Different [828, 281, 79, 1846, 2496, 1681, 1019, 1837, 2405, 2256]. Differential [2166, 2161, 1099, 3173]. Differential-Linear [2166]. Differentiated [25, 2483, 2514, 2063, 2534]. Differentiation [2526, 56, 2527]. Diffie [2181]. Diffracting [1840]. DiffServ [2058, 52, 2595, 3388]. Diffusion [1803, 1041, 1022, 1808, 1806, 1025]. Digging [1294]. Digit [498, 1618]. Digital [3287, 1252, 1236, 1279, 801, 1237, 1261, 1280, 1275, 3273, 2452, 1273, 3274, 3282, 3285, 1248, 1251, 3289, 2603, 387, 1276, 1152, 1274, 1643, 1239, 1271, 1800, 58, 1253, 1246, 2476, 2462, 1268, 2451, 1244, 2590, 393, 1150, 1257, 1281, 818, 1160, 2730, 3286, 1266, 1270, 3288, 1249, 1003, 1272, 2421, 2255, 2213, 1265, 1258, 3122, 3154]. Digitisation [1282]. Digraphs [3061]. Dihedral [2684]. Dimension [2599, 3128, 259, 3130, 3129, 3196, 1562]. Dimensional [738, 1977, 1955, 2338, 3064, 2017, 3148, 3050, 1063, 3065, 1956, 1506, 2442, 1757, 3305, 766, 1779, 1320, 1865, 1020, 1860, 2016, 2, 3]. Dimensioning [2579, 106]. Diophantine [2381]. Diploidy [1215]. Dipoles [1860]. Direct [845, 375, 1555, 2910, 3186, 1056]. Directed [1085, 1733, 2712, 3064, 2553, 3055, 1652, 1559]. Direction [3256, 2916, 3084]. Directional [2676, 1885, 1933]. Directions [1663]. Directory [1366, 3236]. Disagree [440]. Disambiguation [3015]. Disaster [1139]. Discard [1931]. Discernibility [1602]. Discourse [2261, 1144, 1170]. Discovering [2796, 695, 3159, 2397, 2400, 3186, 3155, 2726, 2824]. Discovery [737, 2925, 692, 741, 2933, 2629, 3156, 750, 3181, 2724, 3190, 3225, 189, 1378, 3184, 3188, 3179, 3226, 3196, 3326, 715, 3178, 2310, 3198, 2766, 2266, 3185, 3197, 769, 3339, 946, 2434, 3121, 3153]. Discrete [2262, 2169, 2320, 346, 762, 842, 768, 1673, 330]. Discrete-Valued [762]. Discretization [2984]. Discriminant [1193]. Discrimination [1718]. Discriminative [1812, 725]. Discussion [1386, 1413, 1394, 1412, 1402, 1404, 1409, 1392, 1396, 1405, 1390, 1411, 1384, 1400, 1388, 1382, 1410, 1407, 1398]. Discussions [237]. Disease [1844, 1794, 1760]. Diseases [2987]. Disembodied [390]. Disjoint [3056, 789]. Disjunction [603, 2555, 2260]. Disjunctive [291, 2558, 2560, 290, 296, 2634]. Disk [1218, 2640, 3240]. Disparity [1017]. Display [1859, 380, 1860]. Dissection [1887]. Dissipative [2023]. Distance [1649, 3146, 744, 2697, 1914, 2264, 496, 2683, 10, 1221, 1822, 5]. Distance-Based [2264]. Distinguishing [1656]. Distortion [2414]. Distributed [2163, 835, 1962, 901, 154, 2754, 438, 3349, 3007, 2763, 2719, 2712, 2396, 3209, 902, 150, 688, 1700, 3251, 2172, 1225, 787, 2337, 2731, 983, 156, 244, 664, 2346, 2195, 2294, 1709, 2379, 2724, 1375, 854, 137, 3377, 1505, 2772, 2845, 201, 2490, 2747, 2301, 169, 140, 1268, 1532, 132, 139, 2438, 513, 686, 537, 1464, 2736, 420, 2591, 2793, 1743, 2825, 2765, 3115, 145, 2502, 3096, 1492, 2443, 2535, 2470, 678, 1999, 1969, 3410, 2437, 683, 3239, 2340, 464, 127, 899, 1676, 3214, 2725, 144, 165, 2060, 2760, 3009]. Distributed-Verifier [2163]. Distributing [371]. Distribution [3123, 3146, 744, 2172, 2578, 3128, 988, 110, 1910, 2814, 3160]. Distribution-Sensitive [744]. Distributional [1442]. Distributions [3407, 850, 2873]. Divergences [2084, 2089]. Diversity [870, 1215, 804, 882]. Diversity-Guided [882]. Divide [1513, 1160]. Divide-and-Conquer [1513]. DKS [1565]. DLV [617, 616]. DM [2932]. DMPI [1525]. DNA [2315, 1096, 2382, 1091, 3174, 2370]. DNA-Tract [3174]. DNF [724]. DNS [673, 148]. Do [2680, 461, 1981, 2286, 2212, 3312, 417]. Do-All [2680]. Docking [1076]. Document [227, 229, 44, 1367, 1456, 224, 1961, 1646, 141, 2444, 1448, 1936, 1369, 222, 1959, 3396, 2961, 2743, 2419, 1235, 3277, 1957]. Document-Centric [141]. Documentation [1141]. Documents [1938, 1972, 1368, 643, 2230, 2430, 1232, 761, 1946, 2790, 209, 1200, 3276, 2591, 208, 265, 1966, 1258, 2789]. Does [1413, 2885, 2586, 1298]. DOF [1778]. DOGMA [2787]. Dokumentation [1141].

Dokuz [1221]. DOLCE [1459]. Dolev [559]. Domain [479, 1309, 2712, 3283, 986, 309, 2351, 934, 1162, 2505, 148, 3014, 3032, 274, 859, 774, 33, 2528]. Domain-Based [2351]. Domain-Independent [3032]. Domain-Specific [2505, 774]. Domains [306, 2959, 1448, 2266, 3032, 1287]. Dominance [1552]. Dominance-Based [1552]. Dominated [1299]. Dominating [505]. Don’t [3159]. DoS-Resistant [1385, 1386]. DPLL [592]. DPLL-Based [592]. DPS [826]. Drawing [3054, 3087, 3053, 3051, 3086, 3046, 3064, 3076, 3049, 3065, 3169, 3074, 230, 225, 3088]. Drawings [3078, 3052, 3062, 3050, 205, 235, 3075, 3047, 3069, 2708, 231, 3068]. Drilling [1773]. Drive [461, 1026, 3414]. Driven [1752, 482, 3046, 199, 1057, 651, 1692, 3411, 2752, 2208, 2210, 457, 1772, 1463, 1317, 1281, 656, 2512, 2778, 432, 1468, 1970, 2528, 649, 1475, 185]. DRR [1932]. DRT [107]. Drug [2036]. DSA [1911]. DSDM [448]. DSP [2006]. DSpace [1269]. DSRC [94]. DSRC-Based [94]. DTD [2430, 774]. DTDs [1938]. DTI [1925]. Dual [2701, 1834, 47, 1341]. Duality [529]. Duplication [1070]. Duration [1431]. During [2938, 1861, 1791, 1776, 1751]. DUSTer [2084]. Dutch [1126, 1109]. Dynamic [2064, 2342, 2468, 3218, 1938, 2763, 3210, 275, 989, 2472, 943, 952, 3027, 3040, 3211, 687, 3280, 2001, 129, 839, 1215, 2250, 1043, 2788, 2762, 3080, 2364, 1815, 2742, 2466, 841, 137, 3377, 134, 1187, 2476, 26, 660, 1241, 2353, 912, 2469, 1910, 715, 1031, 818, 1648, 2710, 3001, 1546, 1310, 3341, 570, 35, 2062, 3388, 2437, 1354, 3243, 3398, 1748, 844, 2060, 1480, 3078]. Dynamic-Cut [943]. Dynamic-Sized [2364]. Dynamical [161]. Dynamically [3323, 684, 2214]. Dynamics [838, 1637, 2028, 2011, 2207, 2926, 2027, 2909, 1027, 42, 2032, 2025, 1899]. DYNAMO [3236].

E-A [24]. E-Barter [170]. E-business [1683]. e-Citizenship [1111]. E-Commerce [2826, 1377, 1376, 1374, 2322, 3259]. E-content [2541]. e-Contract [2218]. e-Democracy [1151, 1143]. E-GOV [1114]. e-Governance [1175, 1147]. e-Government [1174, 1137, 1157, 1165, 1119, 1168, 1132, 1171, 1146, 1121, 1169, 1179, 1162, 1117, 1107, 1150, 1164, 1176, 1153, 1181, 1133, 1167, 1131, 1129, 1148, 1170]. e-Learning [1174, 1117, 1309, 2458, 2456]. E-MuniS [1136]. e-Negotiation [2218]. E-ONIGOCO [13]. e-Practices [1112]. E-R [2221, 1345, 24]. e-Society [1173]. e-Voting [1112]. E/S [1348]. EA [858]. Each [665]. Eager [1420]. eal [934]. Ear [1895, 506]. Early [2342, 556, 2071, 206, 1781, 2866, 1383, 1384, 3253]. Early-Delivery [2342]. Ears [1777, 1877]. Earth [3287, 1246]. Earth-Observation [1246]. Earthquakes [1718]. Easily [3135]. Easy [3275, 1431]. eBook [1232]. ebXML [2433]. ECG [2933, 2939]. echo [1038]. Echocardiographic [1833]. Echographic [1766]. echography [1766, 1767]. Eco [1661]. Eco-ontologies [1661]. Ecological [2881]. Economic [3250, 2540]. Economical [500]. Ecosystem [2509, 2806]. Ecosystem-Inspired [2806]. Edge [2401, 2643, 1330, 2665, 474, 645, 2861, 37, 2689, 2862, 2403, 859, 3455]. Edge-Colouring [474]. Edge-Preserving [3455]. Edge-Valued [2643]. Edges [481, 2877, 3]. Edit [3281]. Editors [155, 3117]. EDMIS [2428]. Education [1271, 1223, 1221, 442]. Educational [413]. EER [2242]. Effect [2831, 1768, 2626, 855, 3160]. Effective [460, 2951, 519, 3468, 230, 2310, 2803]. Effectiveness [1192, 1791, 3451]. Effects [2461, 2938, 240, 938, 1109, 1681, 862]. Efficacy [438]. Efficiency [1174, 156, 2382, 793, 840]. Efficient [535, 2058, 1139, 2344, 128, 1385, 974, 338, 1685, 1633, 147, 1977, 1955, 970, 3058, 1386, 641, 207, 952, 2681, 907, 2839, 3144, 1225, 1076, 3057, 2350, 664, 1540, 702, 2159, 2276, 2429, 2067, 1054, 3262, 146, 510, 2753, 2736, 2995, 1226, 1835, 276, 2972, 2653, 1929, 2059, 2699, 3393, 1008, 1003, 36, 2016, 2499, 2086, 2789, 63, 3, 5, 1070, 315]. Efficiently [770, 772]. EFSMs [3090]. Egocentric [2916]. eGOV [1118]. Eigenspectrum [3120, 3152]. Elastic [2526]. Elastography [1885]. Elbtunnel [809]. Electing [132]. Election [2502]. Electric [3001, 1502, 2257]. Electromagnetic [1774]. Electronic [1130, 2958, 1178, 1136, 411, 1163, 2211, 1127, 1159, 1105, 37, 1200, 1149, 13, 2273, 1233]. Electrophysiological [1857]. Elektronische [1141]. Element [1889, 1827]. Elementary [2017]. Elements [1393, 2788, 1557, 217, 232, 3435]. Eliminating [1437]. Elimination [2963, 585]. EliMT [2102]. Ellipses [2694]. Elliptic [2170, 1806]. Elusiveness [804]. EM-TCP [94]. Email [3180]. Embedded [2008, 3210, 1995, 1738, 1993, 1869, 1985, 1424, 2747, 1982, 1980, 1992, 1743, 1979, 2002, 657, 1999, 3124, 1991, 1730]. Embedder [3085]. Embedding [295, 1691, 2701, 3065, 2380, 2494, 2016]. Embeddings [3061]. Embolism [1791]. Emergence [2022, 3187, 2023, 2021]. Emergent [2726, 3025, 668]. Emerging [740, 2873]. Émigré [388]. Emissions [2942]. Empirical [2261, 695, 581, 749, 2769, 437, 467, 380, 2240, 2994, 3277, 168]. Empirically [996]. Employing [909, 1253, 89]. Empowered [3008]. Empowering [3203]. Emptiness [2661, 2634]. Enabled [3209, 412, 3385, 396]. Enabling [3217, 2756, 1116, 280, 821]. Enactment [2455]. Encapsulating [1395, 1396]. Encephalomyopathies [1619]. Enclaves [1410]. Encoding [695, 2642, 712, 3090]. Encodings [2258, 3062, 846]. Encounters [1293]. Encrypted [2202]. Encryption [951, 2155, 956, 949, 2182, 961, 955, 964]. End [166, 2759, 71, 2513, 91, 3240]. End-to-End [2759, 71, 2513, 91, 3240]. End-User [166]. Ended [560]. Endmember [3002]. Endoscope [1752, 1764, 1753]. Endoscopic [1759, 1757, 1750, 1755]. Endoscopy [1851, 1872]. Endovascular [1799, 1861]. Energy [1990, 2681, 2005]. Energy-Aware [1990]. Energy-Efficient [2681]. Enforcement [2190, 1189]. Enforcing [2324]. Engagement [1222]. Engine [864, 2732, 3230]. Engineering [2314, 692, 72, 3356, 652, 1291, 235, 1458, 2285, 3010, 2787, 1395, 1396, 181, 298, 2805, 830, 1504, 3354, 2522, 35, 3389, 2528, 1495]. Engines [1208, 1192, 644, 758, 136, 2973]. English [3023]. Enhance [2056]. Enhanced [1097, 1955, 3233, 379, 2807, 11, 15, 1882, 1858, 1509, 3098, 1787]. Enhancement [992, 2861, 91, 133]. Enhancing [2024, 2166, 55, 1636, 121, 2005]. Enough [2388]. Enriched [529]. Enriching [3004, 1671]. Enrichment [3014]. Ensemble [2969, 3165, 3439, 699, 773]. Ensembles [1704]. Ensuring [1470]. Entanglement [2662, 2387]. Enterprise [2050, 117, 654, 2508, 2778, 2538, 674, 1373]. Enterprises [2730]. Entertainment [3392]. Entire [1100]. Entities [370, 1019, 127]. Entity [3017, 3177, 1199, 1948, 265, 2217]. Entity-Relationship [1948, 265, 2217]. Entries [1036, 2097]. Entropy [1047, 2086, 2977]. Entry [381]. Enumerate [1077]. Envelope [2884]. Environment [1450, 2763, 339, 1174, 2082, 1551, 2099, 581, 2407, 1501, 2448, 66, 854, 1505, 3384, 3231, 1462, 3271, 3220, 130, 1138, 391, 1538, 1485, 1516, 1349, 400, 145, 2710, 2304, 2531, 1515, 808, 2021, 446, 1502, 1522, 3214, 168, 427, 2457, 3234, 2873, 2827]. Environmental [1498, 3428, 1668]. Environments [2461, 154, 2581, 3336, 1032, 76, 1215, 3386, 3022, 2762, 135, 1496, 137, 2501, 2476, 3262, 1244, 3229, 3345, 805, 55, 2766, 2764, 1499, 3212, 2771, 1495, 2055, 2434, 2824, 844, 3264, 3011]. Enzymatic [2037]. Enzymes [1092]. Epidemic [2513]. Epilepsy [1903]. Episodes [3452]. EPNI [2401]. EPSB [37]. Equal [2197]. Equations [1058, 2167, 489, 3173]. Equilibrium [3430]. Equivalence [556, 288, 607, 3293, 246]. Equivalence-Checking [3293]. Equivalences [508]. Erasing [3140]. Erasure [689]. erbrand [934]. Ergonomic [1348]. Ergonomics [1351]. Ergotracer} [162]. Erlang [942]. Erlass [1141]. EROOS [2215]. ERP [426, 2750]. Error [1053, 2346, 2436, 214, 2414]. Error-Correcting [2346]. Error-Correction [214]. Errors [1034]. ERX [1948, 265]. ERX-QL [265]. Essence [930]. Essential [1376, 948, 434]. EST [1077]. Establish [2729]. Establishing [424]. Establishment [1168]. Estimate [2478, 701]. Estimated [3419, 1860]. Estimates [292]. Estimating [797, 1084, 3033, 3452]. Estimation [2052, 2530, 869, 1041, 2875, 1043, 1836, 1615, 2910, 713, 996, 1905, 1056, 1806, 1711, 2596, 1019, 2402]. Estimator [227]. Estimators [726]. EU [2458]. euclidean [1902]. Euler [2297]. EURO [1162, 1161]. EURO-CITI [1162, 1161]. Europe [1489, 1149]. European [1165, 1120, 1140, 1121, 1153, 1161]. Evaluating [2265, 2240, 918, 383]. Evaluation [1696, 2576, 1362, 266, 1733, 2943, 2190, 1255, 378, 1261, 1048, 995, 2242, 382, 76, 2619, 1660, 242, 2579, 2812, 1777, 1927, 789, 1702, 1756, 1694, 96, 3448, 1448, 1532, 1244, 28, 467, 1038, 2229, 1750, 1349, 942, 211, 1923, 2467, 2514, 1864, 1515, 90, 821, 1850, 3277, 101, 1755, 2632, 2941]. Event [371, 1588, 3179, 3382, 529, 271, 133, 3114, 613, 1166, 185, 144, 330].

Event-Based [271, 3114, 613]. Event-driven [185]. Event-State [529]. Events [747, 1718, 2882, 2950]. Evidence [1656, 702, 350, 840]. Evidence-Based [350]. Evoked [1623]. Evolution [903, 838, 2984, 1736, 924, 3423, 1089, 872, 1302, 78, 2319, 875, 3309, 3175, 3350, 1626, 660, 885, 2475, 1277, 878, 2371, 3333, 1468, 2784, 927]. Evolutionary [868, 904, 901, 892, 902, 905, 850, 906, 907, 839, 908, 888, 841, 2013, 846, 842, 876, 917, 883, 918, 2710, 867, 1546, 1623, 920, 921, 882, 2980, 861, 865]. Evolutive [887, 496]. Evolvable [2740]. Evolve [3333]. Evolved [886, 2014, 3422]. Evolving [571, 1938, 2015, 2285, 422, 1190, 1973, 920, 772, 2917, 889]. eVoting [1142]. Exact [1629, 3302, 2398, 1562, 1591]. Examination [2722]. Examine [1844]. Example [2629, 2912, 2088, 1222, 498, 1847, 398, 2091, 1645]. Example-Based [2088, 1847, 2091]. Examples [2954]. Exception [1602]. Exchange [1385, 1386, 2181, 957, 1457, 972, 718, 182]. Exchanges [371]. Excluding [1327]. Exclusion [3209]. Executable [650, 1296]. Executed [1981, 2286]. Executing [2804]. Execution [3336, 568, 2001, 303, 1540, 308, 2304, 2624]. Executions [2731]. Exhaust [864]. Existing [2758, 449, 217]. Exit [2178]. Exit-Polls [2178]. Exophthalmia [1889]. Expand [1524, 1142]. Expanding [976]. Expansion [2829, 984, 707, 1194]. Expectation [1819, 1786]. Expectation-Maximization [1786]. Expected [3146, 488, 850, 1086, 2707]. Experience [1230, 1683, 1738, 3201, 1781, 3384, 2772, 2498, 1147, 2100, 1948, 2749, 1129, 458, 3011]. Experiences [1301, 2099, 1291, 2194, 2607, 3253, 1382, 471, 668, 1295, 1445]. Experiment [1064, 2377, 1931, 124, 2091, 1766, 435]. Experimental [2988, 1694, 211, 899, 1260].

Experimentation [1073, 1599].

Experiments [3162, 893, 1767, 3234, 3397]. Expert [3027, 1491, 750, 708, 2257, 1786, 1788]. Expert-Guided [750]. Explaining [909, 773]. Explanations [290, 1254]. Explicit [294, 2773, 2555, 501, 3323, 48, 1364]. Exploitation [2882]. Exploiting [338, 2375, 2611, 2640, 216]. Exploits [2613]. Exploration [1593, 355, 3110, 3042]. Explorations [3287]. Exploring [1085, 438, 1273, 678]. Exponent [963]. Exponential [2675, 711]. Exponentially [538]. Exponentiation [1226]. Expressing [753]. Expression [1101, 871, 884, 764, 915, 1099, 3173]. Expressions [320, 2713, 2910, 509, 511, 1561, 296]. Expressive [2237, 2557, 579, 530, 2125, 2572]. Extend [2087]. Extended [1549, 2688, 2680, 2027, 3300, 554, 1628, 3465, 953, 1663, 509, 972, 2508, 1710, 3069, 2500, 3141, 3297, 3380, 318, 2217, 2702, 966, 234, 2720, 731, 2855]. Extending [1028, 1442, 2548, 156, 1316, 1307, 2777, 2798, 1952, 2853, 1428]. Extensible [3323, 2825, 648, 1258]. Extension [1308, 2551, 574, 2559, 1436, 1596, 2567, 75, 1957]. Extensions [693, 2174, 148, 2222, 1566]. External [2381, 2282]. eXtinct [468]. Extra [1212, 3196, 2923]. Extra-Theoretical [3196]. Extract [1823]. Extracting [1073, 1227, 2891, 3106, 2236, 2399, 643, 2230, 1195, 2430, 1345, 3196, 1075, 2445]. Extraction [3017, 3280, 1456, 1363, 2840, 884, 3158, 206, 3132, 1368, 3167, 1554, 3437, 3450, 3438, 2444, 761, 2890, 1947, 1359, 3178, 208, 2993, 230, 1256, 217]. Extractors [964, 1644]. Extrapolating [359]. Extremal [1027]. Extreme [3279, 463, 470, 420, 443, 2484, 436, 442, 427]. Extremes [441]. Extremity [1881].

Exudate [1900]. Eye [2895, 2919]. Eye-Head-Hand [2919]. Eyes [2894, 1061]. Eylul [1221].

F [493]. Face [2893, 1045, 2275, 2908, 2909, 1055, 2922]. Faces [2907, 2923]. Facial [1716, 2895]. Facilitating [2218, 2823]. Facility [1322, 1324, 1309, 2001, 1333, 1339, 1341]. Fact [1412]. Factor [2455, 984, 2266, 3121, 3153]. Factoring [2175, 935, 2162]. Factorization [2275, 2151, 1034]. Factorization-Based [1034]. Factorizations [869]. Factors [1168, 3178, 732, 776, 1344]. Faculty [1223]. Fading [2436, 2467]. Failover [2339]. Failure [2365, 2697, 2357, 2360, 941, 2502, 328]. Failures [2360]. Fair [2061, 2518, 364, 410, 517, 2519]. Fairness [557, 49, 562]. False [1718]. Familiar [2922]. Family [2106, 3243]. Far [2467]. Fashion [2788]. FASME [1131]. Fast [2365, 738, 740, 480, 1094, 1535, 1089, 1697, 1528, 1074, 1893, 997, 1220, 999, 88, 16, 1009, 84, 1004, 132, 139, 2598, 2094, 2583, 98, 2436, 3180]. Fast/Slow [2436]. Faster [1630, 2665, 1217, 2673]. Fastest [3414]. Fault [1696, 2365, 1687, 1682, 38, 1692, 2738, 1689, 2975, 138, 16, 803, 2443, 684, 2340, 1429, 1690, 2646]. Fault-Tolerant [2365, 138, 2443, 684, 1690]. Faults [797]. Faulty [3172, 1406, 1407]. FAUST [1075]. Favorite [424]. FD [303]. FDMA [107]. FDMA-Based [2495]. Feasibility [1375]. Feasible [2958]. Featural [2922]. Feature [2263, 2550, 1057, 749, 1063, 3190, 3437, 3450, 3194, 2882, 3449, 1616, 766, 1913, 2265, 3340, 2990, 2972]. Feature-Based [1913]. Feature-Driven [1057]. Features [3443, 1716, 1989, 703, 1195, 2451, 2953, 2897, 3036, 2942]. Featuring [2955]. FEC [2577, 53]. Federated [2046, 2712, 1274]. Federation [2452]. Fedora [1259]. Feed [884]. Feed-Forward [884]. Feedback [2064, 1768, 2898, 1758, 2409, 1769, 1751, 2673, 1254]. Feedforward [3148]. Feeding [1460]. Feeds [3211]. Fest [466]. Fewer [3444]. Feyerabend [3360]. FH [107]. FH-FDMA [107]. Fiber [1808, 1925, 1845]. Fibrillation [2938, 2980]. Fibring [583]. Fiction [1412]. Field [2863, 2865, 1503, 1764, 2953, 1991, 1502]. Fields [1045, 1898, 1133, 1033, 971]. Fighting [877]. Figures [223]. File [1524, 163, 135, 2690, 59, 171, 1509, 2413]. File-I [1509]. File-I/O [1509]. Files [825, 2743, 1641, 1260]. Filter [990, 2746, 1030, 981, 1008]. Filter-Combiner [981]. Filtering [1647, 1615, 1885, 2954, 2267, 1025]. Filters [1022, 2862]. Filtrons [2373]. Find [1571]. Finding [1393, 745, 1074, 2847, 2220, 3424, 732, 776, 1078, 1067, 1083, 2729, 1639, 1852, 647]. Findings [437]. Fine [3231, 3205]. Fine-Grain [3231]. Fine-Grained [3205]. finger [1771]. Fingerprints [2203]. Fingers [2186]. Finite [306, 479, 269, 3146, 491, 309, 475, 28, 1889, 3440, 3311, 2707, 2124, 3292, 522, 1827, 524, 274, 971]. FIPA [2819, 1212]. FIPA-Based [2819]. FIPA-Compliant [1212]. Fire [2864]. Firewall [2339]. Firewalls [2411]. Firing [1635, 2016]. First [1230, 2545, 1175, 331, 3113, 3094, 556, 498, 30, 2669, 893, 435]. First-Order [2545, 331, 3113]. Fish [2029]. Fishbowl [470]. Fisher [1045, 1591, 1903]. Fit [2660]. Fitness [850, 839, 873, 874, 867, 843, 2374]. Fitting [1053]. Fixed [2130, 2675, 3057, 845, 2131, 543, 913, 2673, 1065]. Fixed-Parameter [2675, 1065]. Fixpoint [2642, 544]. fJDBC [2853]. FLAME [3331]. Flash [677]. Flat [2684]. Flat-State [2684]. Flaws [954]. FlexCC2 [2006]. Flexibility [2838, 1124, 1532]. Flexible [2244, 3211, 264, 2842, 2843, 1253, 2432, 2849, 1890, 3262, 2562, 2594, 1138, 1753, 2851, 2730, 1773, 1599, 2653, 3341, 3338, 2727, 1522, 1722, 2855]. Flight [527]. Flipping [3174]. Flocking [926]. Flood [1501]. Flora [3286]. Flow [2523, 2888, 1041, 1042, 1326, 3463, 2872, 2503, 1553, 997, 2781, 26, 1337, 52, 3404, 2480, 2447, 1860, 3389, 1899, 2625]. Flow/Depth [997]. Flows [2877, 1498, 2038, 2588, 2518, 513, 2454]. Fluctuations [856]. Fluent [2102]. Fluid [1916, 807, 1899]. Fluoroscopic [1908]. Fluoroscopy [1856, 1909]. Fluoroscopy-Based [1856]. Flux [3416]. Flux-Oriented [3416]. Fly [1927]. Fly-Through [1927]. Flying [2917]. fMRI [2943, 1931]. Focused [1236, 1865]. Focussed [1646]. Following [1245, 434]. Fontan [2936]. Force [1777, 1776, 1769, 1755]. Forceps [1763]. Forces [1769]. Forcing [1172]. Ford [2101]. Forecasting [2959, 1501, 2998, 1720, 886, 2257]. Foreword [823]. Form [207, 1500, 208, 1365]. Formal [2546, 1989, 3097, 3361, 929, 2621, 829, 1460, 2235, 336, 2787, 2850, 3099, 170, 830, 1449, 2631, 2616, 1421, 1422, 808, 431, 1469, 2312, 3012, 1366, 2048, 2200, 334]. Formalising [1375, 2306]. Formalism [3401]. Formalization [596]. Format [408]. Formation [2033, 2087, 1038, 2387, 3025]. Forming [2010]. Formula [2171, 1952]. Formulae [224, 3109, 352]. Formulas [331, 593, 329, 3380]. Formulation [3405, 2856, 584]. Forum [3254]. Forward [351, 884, 1941, 1025]. Forwarding [89]. Found [382]. Foundation [287, 215, 3455]. Foundations [3401, 1248, 652, 2219, 2779, 2308]. Founded [294, 512, 937, 2774]. Founder [1083]. Four [474, 424, 2821]. Four-Colour [474]. Foveated [2871, 2865]. Fox [1245]. FP6 [783, 1699]. FPGA [2421]. Fractal [1562]. Fractional [2930, 3063]. Fragment [3174, 1432, 2567]. Fragmentation [497]. Frame [208, 2774, 596]. Frame-Based [2774]. Frames [588]. Framework [569, 2455, 901, 2740, 558, 1736, 2645, 1808, 3082, 1739, 1212, 2554, 2507, 2974, 2843, 791, 2975, 3266, 2621, 829, 2762, 66, 3225, 21, 2617, 637, 2398, 1369, 1172, 2498, 1833, 170, 3387, 1360, 2039, 3137, 1559, 2075, 2240, 2730, 2746, 2825, 2532, 1492, 3228, 2066, 1827, 2853, 3304, 3230, 1479, 1676, 2811, 1287, 3214, 1567, 2664, 1202, 2060, 3191, 1644, 3216, 1464]. Frameworks [630, 638, 1961]. Free [2363, 1498, 2675, 1500, 2379, 2297, 2364, 506, 632, 913, 914, 3018, 522, 2349]. Free-Form [1500]. Freeing [193]. French [1230]. Frequencies [1197]. Frequency [3386, 912]. Frequent [743, 745, 3441, 3155, 421]. Fresh [136]. Friction [1776]. Friendly [163, 125]. Front [1128, 1153, 1240]. Frontal [1051]. Frontotemporal [1844]. FT [1693]. FT-CORBA [1693]. Fujitsu [1535]. Full [557, 2494]. Fully [3209, 2194, 2667, 1648, 1850]. Function [968, 1533, 57, 876, 973, 2993, 855, 886, 1081, 881, 1827, 102, 726, 1722, 3425]. Functional [932, 1789, 1082, 2945, 2551, 1807, 1922, 1756, 3384, 1912, 1243, 3369, 941, 1561, 245, 2930, 2647, 946, 126]. Functionality [2406]. Functionally [1075]. Functions [3123, 1512, 320, 2367, 982, 2180, 976, 839, 3134, 2276, 3294, 760, 3129, 985, 867, 3124, 1917, 971]. Fundamental [1462, 1173]. Fundus [2947]. Funnel [2671]. Fusion [1737, 554, 997, 853, 2850, 2256, 1583]. FUTURA [2257]. Future [2073, 2451, 1108]. Fuzziness [2853]. Fuzzy [2989, 2829, 2978, 2976, 2270, 1551, 2987, 2328, 3300, 1660, 2085, 887, 2560, 1209, 1588, 1191, 1210, 374, 1720, 893, 3034, 2973, 3453, 1548, 1577, 302, 2573, 2977, 1852, 2929].

G1 [2485]. G2 [2485]. GA [3454, 2953, 881]. Gabor [2870]. Gain [850]. GALA [3366].

Galician [1634]. Game [3275, 708, 13, 2538, 2904, 2021, 2814, 3398]. Game-of-Life [2021]. Game-Theoretic [2814]. Games [903, 691, 2128, 3131, 2383, 2120, 2118, 708, 2117, 543, 2113, 2563, 318]. Gaming [2041]. Gap [1740]. GARA [835]. Garbage [282, 281, 1994]. Gas [875, 2031, 3029, 848, 853, 1575]. Gastric [1603]. Gastro [1758]. Gastro-Intestinal [1758]. Gate [2981]. Gates [931]. Gateway [12, 43, 87, 62, 2329, 2755, 2339]. Gathering [1123]. Gaussian [3407, 2999, 1047, 2280]. Gaze [2872, 408, 2895]. Gaze-Depended [2872]. GBML [2262]. Geeks [451]. GEM [949]. Gender [2907]. Gene [1101, 1098, 2315, 764, 915, 1102, 1079, 1078, 3157, 1099, 3173]. Gene-Order [1102]. General [691, 131, 3107, 2219, 3432, 3130, 3129, 2171, 2386, 2786]. General-Purpose [2386]. General-Sum [691]. Generalisation [631, 1669]. Generalised [881]. Generalization [1723, 628, 624, 3426, 216]. Generalized [324, 1553, 1554, 3188, 1711, 2454, 2537, 2960, 2389, 966, 2632]. Generally [888]. Generate [1374, 1667]. Generated [2017, 646]. Generating [1068, 750, 2430, 654, 221, 1297, 232, 1644]. Generation [2439, 974, 559, 2207, 389, 2643, 1813, 2089, 2235, 2381, 1729, 3265, 2584, 396, 124, 222, 1616, 1463, 1241, 632, 3112, 2100, 917, 2504, 1874, 2527, 431, 1748, 1932, 3182]. Generation-Heavy [2089]. Generative [1734, 3344, 1738, 2879, 1730, 807]. Generator [2793]. Generators [3448, 1743]. Generic [2982, 262, 2156, 1000, 1376, 224, 46, 2805, 2026, 515]. Genes [871]. Genetic [2982, 895, 1006, 2984, 896, 870, 3463, 871, 905, 2987, 851, 884, 1213, 1215, 3189, 2986, 873, 2425, 2988, 852, 861, 910, 865, 2584, 1378, 3424, 897, 3421, 913, 914, 916, 879, 1214, 2983, 2991, 880, 847, 899, 3453, 859, 1577, 3434, 3435]. Genetics [916]. GenGED [2289]. Genome [1097, 1094, 2371, 1071, 1077]. Genomic [1064]. Genotype [3454]. Genotypes [3422]. Gentle [2905]. Geo [1661]. Geo- [1661]. GeoCosm [1355]. Geodata [2797]. Geodesic [1802]. Geodesics [1806]. Geographic [1659, 3285, 1662, 1676]. Geographical [1656, 2213]. Geographically [1657, 2470]. Geography [763]. Geometric [1053, 3060, 3056, 1333, 910, 2398, 2695, 3073, 2239, 730]. Geometrical [3036]. Geometries [1977]. Geometry [2274, 2880, 3284]. Geospatial [1656, 1658, 1355]. GeoVIBE [3285]. GeoVSM [1659]. German [1129]. Germany [1107]. Gesture [1051]. Get [436]. Getting [2100, 783]. GF [834]. Giganet [1534]. Giotto [1983]. Giotto-Based [1983]. GIS [1204, 2213]. GISs [1660]. Given [1086, 3128]. Glass [1808]. Global [2455, 3374, 473, 1042, 3252, 839, 2235, 1506, 2100, 2743, 3254]. Global-Local [839]. Globalization [412]. Globally [1535, 1998, 3366]. GlobData [2443]. Glucose [2932]. Glucose-Insulin [2932]. GMPLS [46]. Gnutella [202, 667, 666]. Goal [1237, 2226, 2903]. Goal-Oriented [1237]. Goals [2488, 3030, 715]. Gödel [2565]. Goes [1143]. Gold [1460, 1906]. GOLOG [267]. Good [679]. Google [1638]. GOV [1114]. Governance [1175, 1125, 1147]. Government [1174, 1137, 1157, 1165, 1119, 1168, 1132, 1171, 1122, 1146, 1121, 1169, 1179, 1109, 1162, 1128, 1117, 1107, 1123, 1105, 1150, 1164, 1176, 1153, 1181, 1133, 1167, 1131, 1129, 1148, 1170]. GPRS [1681, 50]. GPS [2575]. GQ [958]. GraCAD [2311]. Graded [1663]. Gradient [691, 3427, 1814]. Grading [1853]. Gradual [753]. Grafting [1908]. Grain [3231]. Grained [3205]. Gram [3120, 3152]. Grammar [2621, 214]. Grammar-Based [2621]. Grammars [3019, 311, 2308, 1745]. Granular [1604, 1579, 1583, 1584].

Granularity [588, 1510]. Granules [1578, 1558, 1580]. Graph [160, 3087, 2288, 2314, 2291, 482, 2318, 3086, 3046, 3076, 2485, 2320, 1327, 3079, 3082, 2701, 2292, 2294, 2295, 2296, 3065, 2315, 1002, 3055, 2298, 2174, 2299, 2301, 2302, 2201, 2492, 3043, 223, 2305, 3193, 2317, 1559, 2308, 917, 3074, 2309, 857, 2961, 2304, 3111, 214, 2311]. Graph-Based [2174, 2201, 2305, 3193, 2317, 2311]. Graph-Drawing [3086]. Graph-Driven [482]. Graphic [209, 210]. Graphical [1440, 909, 3119, 3151]. Graphics [229, 218, 215]. Graphs [3054, 2694, 478, 3078, 2290, 481, 2688, 3051, 2656, 2548, 2666, 3064, 486, 3402, 488, 2791, 3049, 1327, 2675, 2491, 3048, 2657, 2358, 3050, 3071, 3059, 1332, 1553, 3066, 2297, 497, 1652, 158, 2444, 2126, 2689, 3069, 3073, 2708, 3068, 3067, 3406, 159, 2664]. GRASP [2970]. grasping [1763]. Grayvalue [996]. GREC’2001 [237]. Greedy [698, 2347]. Greenstone [1273]. Grey [1024, 1892]. Grey-Level [1024]. Grid [3050, 3383, 2276, 3215, 3226, 3254, 3301, 3235, 3217, 3237, 1493, 3218, 1497, 1494, 3224, 3221, 3219, 1478, 3250, 2701, 1486, 2752, 1487, 1483, 3384, 3231, 3220, 3229, 3223, 1485, 3375, 3228, 3238, 3239, 3240, 3435, 3234, 1480, 3216]. Grid-Based [3215]. GridLab [3217]. GridRM [3237]. GridRPC [3238]. Grids [3211, 3374, 3387, 3230]. Grothendieck [521]. Ground [209, 592, 1788]. Ground-Truthing [209]. Group [3054, 142, 1434, 2181, 368, 1707, 2156, 2155, 83, 1694, 2474, 1410, 898, 2482, 155, 2757, 107]. Group-Based [83]. Group-Theoretic [3054]. Grouping [2881, 1594, 1187, 1024, 1942, 3338, 216]. Groups [977, 965, 2158, 1664, 2152]. Growing [430]. Growth [2040, 1796, 3441, 122]. Growth-Curve [122]. GSM [823, 50]. GSM/GPRS [50].

GSMP [82, 16]. GSPN [75]. Guarantee [2466]. Guaranteed [2520]. Guarantees [2526, 2575, 3091]. Guard [2468, 1570]. Guarded [2130, 1564, 2129]. Guessed [490]. Guessing [633]. GUHA [3176]. Guidance [1761, 1857, 2918, 2903, 2916, 1864]. Guide [1879, 1799, 1861, 410, 181, 1141, 377, 840]. Guided [351, 750, 2899, 1763, 1855, 2571, 882]. Guided-Prioritized [351]. Guidelines [1233]. Guiding [2785]. Gummy [2186]. Gyral [1802].

H.323 [62]. Habituation [3039]. Habitude [77]. HAL [931, 272]. Halting [2372]. Hamming [2193, 496]. Hand [2919]. Handheld [393]. handiMessenger [379]. Handlers [1491]. Handling [2830, 2735, 311, 1707, 2748, 2089, 3190, 594, 81, 881, 1622]. Handoff [80, 64, 2495, 2598, 98, 2459]. Handover [2580, 2480]. Hands [1051]. Handsign [991]. Handwriting [1048]. Handwritten [1618, 3281]. Handy [853]. Haploid [1066]. Haplotyping [1065]. Haptic [1768, 1758]. Hard [701, 1994, 3447]. Harder [565]. Hardness [2169, 1336]. Hardware [2193, 2379, 2648, 1917, 2021]. Hardware-Based [1917]. Hardware-Free [2379]. Hardware-Implemented [2021]. Harm [1413]. Harmonic [2276]. Harmonics [1932]. Harness [1479]. Harvesting [1773]. Hash [968, 986, 686, 683]. Hash-Function [968]. Having [485]. Hazard [1417]. HCI [394]. Head [1035, 1854, 940, 944, 2919]. Heading [1105]. Heal [1311]. Health [2464]. Healthcare [297, 1296]. Healthy [2937]. Heap [2671]. Heart [2926, 1760]. Heartbeat [2937]. Heaviest [510]. Heavy [2089]. Height [809]. Held [3256]. Helicopter [1983]. Hellman [2181]. Helps [1337]. hemispheric [1811]. Hensel [2169]. Hereditary [2656]. Heritage [1276, 1278].

HERMES [2413]. Heterogeneity [2473, 667, 1256]. Heterogeneous [1539, 3370, 642, 1225, 1496, 3220, 414, 1538, 883, 2445, 2793, 2238, 3253, 1499, 1522, 3214, 2055, 143]. Heuristic [896, 3031]. Heuristics [3074]. Hewlett [1269]. Hexagonal [2500]. Heyting [1551]. HG [62]. HG-iPhone [62]. HGV [3067]. Hidden [2158, 2440, 2489, 975, 1359, 3015, 732, 776, 65, 920, 1010, 2416]. Hidden-Markov-Modells [1010]. Hiding [984, 2158]. Hierarchical [1444, 2913, 2184, 3171, 2911, 2067, 2594, 632, 95, 115, 2595, 1582, 2901, 771, 2313, 1543, 1023, 3431]. Hierarchically [478, 2340]. Hierarchies [3040, 1358, 1290, 633, 627, 3067, 3070, 624]. Hierarchy [2122, 338, 486, 3412, 2677, 536, 372, 123, 1020, 576, 2395]. High [738, 1977, 1955, 1519, 1022, 1873, 3148, 1478, 1274, 1507, 1063, 2622, 2326, 3065, 1956, 1506, 2442, 1881, 108, 124, 1859, 539, 766, 1281, 2100, 1320, 1865, 2424, 34, 2059, 3101, 1933, 2453, 2, 3, 2517, 1892]. High-Accurate [108]. High-Dimensional [1977, 1955, 3148, 1063, 3065, 1956, 1506, 766, 2, 3]. High-Level [1274, 539]. High-Performance [1519]. High-Resolution [1859, 1933]. High-Speed [2326, 34, 2059]. Higher [940, 555, 287, 1745, 3425]. Higher-Order [940, 555, 287, 1745]. Highlighting [155]. Highlights [3281]. Highly [708, 3396, 2789]. Hill [3138]. Hill-Climbing [3138]. HillClimbers [897]. Hints [2065, 1600]. Hip [1791, 1779]. HiPE [942]. HiPE/x86 [942]. HiperLAN [2051]. HiperLAN/2 [2051]. Histogram [859]. Histograms [993]. Histological [1817]. Histology [1839]. Historical [1036]. History [2672, 1066, 1070]. HIV [2371, 2036]. HLR [2435, 2745]. HMAT [2460]. HNM [3302]. HNOWS [1540]. Hoc [105, 2354, 1389, 2077, 1238, 3265, 104, 3258, 103, 1390, 2480, 102, 101, 977, 192, 2809, 189, 3229, 188, 2729]. HOL [1301]. HOL-OCL} [1301]. Holds [1142]. Holistic [1179, 1131]. Holographic [3415]. Holter [2939]. Home [61, 391, 405, 3392, 2413, 62]. Homepage [1639]. Homing [2920]. Homologous [1079]. Homology [1094]. Hop [2680, 2698]. Hopfield [1709]. HOPLA [555]. Horn [300, 2964]. Horowitz [3121, 3153]. Hospital [2944, 2738]. Host [2597]. Hosting [2765]. Hosts [102]. Hot [2339]. Hot-Failover [2339]. Hozo [1462]. HTML [2236, 429, 1959]. HTTP [44]. HTTP/1.1 [44]. Hull [2990, 2633]. Human [1352, 1817, 1350, 781, 1280, 1275, 3283, 1347, 991, 2912, 2907, 2721, 782, 1351, 2880, 708, 3006, 1891, 3178, 916, 1750, 1779, 1895, 1067, 994, 1354, 1860, 2868, 1850, 2947, 1344, 2919, 2892]. Human-Body [1352]. Human-Centric [1347]. Human-Computer [781, 3283, 782]. Human-Computer-Interaction [991]. Human-Machine [1750]. Humane [1677]. Hungarian [3302]. HVS [2892]. Hybrid [341, 343, 540, 275, 2813, 2082, 1488, 1707, 3177, 2959, 317, 863, 138, 2465, 1867, 1834, 1536, 1081, 802, 1623, 2257, 3434, 1824, 1260, 2416]. Hybridization [888, 1091]. Hybrids [2982, 839]. Hydrating [3156]. hydraulic [1772]. HyGLEAM [888]. Hyper [1771, 2492, 964]. Hyper-encryption [964]. Hyper-finger [1771]. Hyper-Star [2492]. Hyperbolic [502]. Hyperclasses [639]. Hypercomputation [2368]. Hypercube [2492]. Hyperedge [2300]. Hyperelliptic [971]. Hypergraph [1323, 619]. Hyperheuristics [919]. Hypermedia [2955, 648]. Hyperonyms [633]. Hyperspectral [2999, 3002]. Hypertext [3274, 2233, 152]. Hypertextualization [1281]. Hypotheses [3139]. Hypothesis [1085, 512, 305, 3169]. Hypothesis-Founded [512]. Hysteresis [231].

I-PETER [3023]. I/O [3113, 1600, 1509, 1522]. IACBC [949]. i}Boost [711]. ICA [2263, 2949, 3384]. ICARO [807]. ICP [999]. ICSpace [2214]. ICT [1182, 417]. ID [2215, 2184, 110, 2183]. ID-Based [2184, 110, 2183]. ID-Logic [2215]. Ideal [2862]. Idempotency [507]. Identification [958, 1611, 694, 1068, 2487, 887, 1063, 1846, 42, 3034]. Identifying [2626]. Identity [2462, 818, 2776]. Identity-Based [2462]. IDS [2616]. IEEE [2466]. IF [339]. IF-2.0 [339]. IFO [493]. II [2136, 2144, 291, 541]. III [2137, 2145]. IKE [957]. IKRAFT [1446]. Ill [2932, 2266]. Ill-Structured [2266]. Illegal [2489]. Illumination [1045, 1016]. Illusion [2888]. ILOG [3079]. ILP [333, 2004]. ILP-Based [2004]. Image [1006, 1799, 1830, 3317, 1014, 1060, 2401, 1792, 1914, 2999, 1916, 1958, 1761, 908, 2440, 1007, 1812, 3415, 1765, 1505, 2879, 1005, 1047, 2412, 3175, 2890, 3305, 2732, 1508, 123, 1833, 2277, 1858, 3042, 1027, 1038, 2278, 2869, 1509, 1923, 994, 1907, 3172, 2424, 404, 1645, 2902, 1059, 3455, 1786, 2870, 2892]. Image-Based [1812]. Images [229, 1839, 1785, 1029, 1897, 1011, 2893, 3002, 1614, 1836, 3306, 1195, 2928, 1847, 1912, 1849, 1816, 1001, 1823, 219, 1891, 1667, 1828, 3043, 1885, 1024, 1826, 1866, 2279, 1819, 225, 231, 2983, 1907, 2282, 3172, 1929, 1837, 1622, 1025, 1003, 1809, 1020, 1906, 1845, 2947, 1892, 1908]. Imaging [1825, 1869, 1909, 230, 1865, 1788]. iMinMax [6]. Immanent [1133]. Immobile [1381]. Immune [1687]. Immunological [2985]. Impact [2365, 3249, 2540, 43, 2496, 2615, 1969]. Impaired [2458]. Impersonation [958]. Implementation [650, 3318, 3335, 2944, 1416, 1527, 1524, 2933, 72, 2768, 1740, 1521, 1121, 2194, 70, 60, 25, 2065, 1508, 1147, 3112, 1021, 2534, 2449, 2816, 2916, 2514, 1526, 2059, 3391, 2421, 253, 1577, 1713, 1202]. Implementations [868, 1482, 2389].

Implemented [1696, 2913, 2021].

Implementing

[448, 2376, 603, 68, 2099, 202, 2758, 1106]. Implications [2511, 2025, 2674, 1182]. Implicit [2547, 1004]. Importance [1574]. Importing [1258]. Imposing [1356]. Impossibilities [2157]. Improve [1919, 2728, 2586, 1078, 2096, 2499, 1677, 1960]. Improved [1322, 2668, 44, 1069, 2697, 2358, 1215, 1807, 3334, 2863, 2179, 1334, 925, 505, 3450, 2682, 706, 110, 509, 1339, 1885, 383, 3163, 2259, 1534]. Improvement [1136, 80, 3028, 94, 1343, 101]. Improvements [961]. Improving [3206, 2978, 2544, 2581, 2949, 2783, 3077, 449, 3109, 1851, 228, 1532, 2492, 49, 2408, 3241, 2996, 2447, 2956, 3426, 2255, 1143, 793, 1115]. IMPULSE [1181]. iMVC [1224]. In-Flight [527]. In-Situ [2221]. In-vivo [1836, 1840]. Inage [1852]. Including [2461]. Inclusion [1580, 1562]. Inclusionof [703]. Incomplete [1595]. Inconsistency [594, 2716]. Incorporate [1028, 1674]. Incorporating [419, 1819]. Incorporation [2826]. Increasing [801, 1706, 878, 1727]. Incremental [2952, 331, 1444, 2839, 702, 2310]. Indeed [2322, 2964]. Independence [1585, 1586, 1587, 804]. Independent [1714, 2943, 2686, 2688, 3219, 2578, 591, 1709, 2672, 2654, 391, 732, 776, 1711, 3032, 2426, 2722]. Index [2829, 74, 151, 1229, 1284, 1319, 1342, 1380, 1415, 1441, 1477, 1541, 1625, 1654, 1679, 1703, 1728, 1749, 1848, 1934, 1978, 2009, 2044, 2072, 2080, 2111, 2187, 2204, 2246, 172, 2284, 2321, 2341, 2366, 2393, 2510, 2543, 2574, 2609, 2627, 2651, 2709, 2799, 2828, 2858, 2924, 2948, 3045, 3089, 3116, 3149, 3199, 3213, 3242, 3255, 3272, 204, 3290, 3313, 3342, 3363, 3373, 3394, 3469, 238, 248, 268, 313, 365, 418, 472, 526, 566, 623, 658, 690, 736, 780, 813, 837, 928, 947, 987, 1062, 1104]. Index [1183, 2134, 3321, 2133, 2938, 2589, 263, 500, 3281, 775]. Indexing [3296, 1280, 1275, 1198, 1064, 3415, 2628, 2591, 1653, 1672, 1676, 3013, 3298, 2, 3, 5]. Indicators [1177]. Indirection [676]. Individual [1065]. Indoor [1032, 3043]. Induced [1694]. Induction [3416, 907, 1594, 700, 705, 756, 761, 764, 3193, 1576, 1590, 1584, 3422]. Inductive [1208, 3138, 3134, 3451, 719, 3141]. Inductively [3133]. Indulgent [1697]. Industrial [1497, 1738, 1356, 1290, 3028, 1616, 2281, 392]. Industry [1205]. Inequalities [483, 754]. Infection [2944, 2036]. Inferable [3133]. Inference [694, 1553, 3134, 2879, 515, 3141, 3406, 3453, 2191, 1852]. Inferencing [1241, 2785]. Inferring [906, 1066, 1070, 3173]. Infinite [291, 486, 485, 277, 3293, 2113, 1562, 604, 345, 318]. Infinite-State [3293, 345]. Infinite-Valued [604]. Inflectional [3396]. Inflectional/Derivational [3396]. Influence [1822]. Info [1351]. Info-Ergonomics [1351]. Informal [1460, 1449, 431]. Information [1578, 160, 1154, 2455, 1110, 2830, 634, 1350, 626, 2207, 1192, 2719, 1135, 2712, 641, 2221, 3283, 2404, 1659, 2401, 3280, 2234, 1456, 1363, 1916, 2841, 2795, 1136, 1156, 590, 1452, 1290, 2360, 1447, 3004, 1007, 3299, 3132, 2227, 643, 498, 3418, 371, 2533, 2908, 1263, 2430, 2781, 637, 1253, 69, 824, 386, 761, 1268, 3262, 61, 3271, 1666, 885, 1558, 2408, 2744, 374, 1200, 1118, 663, 2648, 2737, 818, 1560, 2371, 2825, 3286, 2238, 2241, 1580, 1562, 1563, 1636, 1355, 1467, 1566, 1194, 2470, 136, 152, 409, 2922, 783, 2739, 1699, 1915, 1601, 3284, 1249, 217]. Information [1860, 121, 1818, 1676, 403, 2213, 2977, 2702, 2413, 1678, 730, 3208, 1603, 1960, 3397]. Information-Flow-Based [2781]. INFOSEC [2615]. Infrared [2893]. Infrastructure [105, 814, 1493, 3374, 1350, 1519, 199, 3261, 2761, 3268, 827, 1140, 1974, 3265, 824, 3256, 676, 3267]. Infrastructure-Less [105]. Infrastructures [1688, 1689]. Inheritance [3352, 2935, 2742, 2774]. Inhibition [2863]. Inhibitor [561]. Initial [459, 892, 1119, 889, 2722]. Initialization [2262, 1420]. Initiative [1699, 1240]. Injection [1696]. Innovative [2041, 1140, 3112]. Input [2649, 31, 28, 380, 2707, 3091]. Input-Queued [31]. Input/Output [2649]. Inputs [2363]. Inputting [210]. Insect [2875, 2894]. Insect-Inspired [2875]. Insecurity [3103]. Inseparable [881]. Inserted [1890]. Insertion [3019, 1880, 1768, 1782]. Insights [2888]. Inspection [2281]. Inspiration [1706, 1727]. Inspired [2875, 2900, 2806, 1690]. Instability [41]. Instance [710, 711]. Instance-Based [710, 711]. Instances [1656, 495, 3424]. Instantiation [638]. Institution [2273]. Institutional [1269]. Institutions [580]. Instructions [1671]. Instrument [1765]. Instruments [1751, 1890]. Insulin [2932]. Integer [2158, 1675]. Integers [523, 847]. Integral [1859, 1925]. Integrals [1569]. Integrate [2864]. Integrate-and-Fire [2864]. Integrated [1783, 1830, 2834, 3327, 1659, 2082, 1457, 1894, 1274, 1351, 759, 1106, 3262, 2602, 1167, 1865, 2531, 1235, 1161, 1148, 2457]. Integrating [1299, 2270, 611, 308, 24, 836, 2252, 2254, 1948, 2273, 1678, 2757]. Integration [3358, 2830, 2888, 642, 299, 1371, 2237, 1044, 652, 2975, 2433, 1662, 3006, 2744, 2866, 2793, 1466, 1355, 1467, 2239, 2086, 1899]. Integrity [1470, 1189, 2056, 2185]. Intelligence [2104, 1606]. Intelligent [2830, 2808, 2396, 3457, 1207, 1958, 611, 2721, 2821, 2794, 222, 1241, 2744, 1453, 3461, 2497, 2054, 1343, 1608]. Intensity [1911, 1814, 1913, 1865, 1810].

Intensity-Based [1913]. Intensive [1451, 2614]. Intention [3465, 3409]. Inter [1811, 1165, 1922, 1373, 2923, 2202]. Inter-enterprise [1373]. Inter-extra-Ortho [2923]. Inter-hemispheric [1811]. Inter-Packet [2202]. Inter-registrar [1165]. Interacting [2200]. Interaction [2741, 3283, 3083, 1312, 390, 991, 1574, 2713, 375, 2287, 3340, 2329, 881, 182]. Interactions [402, 389, 2550, 1739, 67, 3080, 2399, 1031]. Interactive [2576, 3087, 1494, 892, 1880, 2611, 2974, 1446, 1447, 1868, 2486, 755, 2662, 2964, 3286, 1262, 91, 1492, 2566, 2021, 2856, 3277, 3298, 2610, 2163, 3276]. Interchange [1935, 1315, 2428, 1242]. Interconnecting [2445]. Interconnection [2492]. Interdomain [2522]. Interest [2891]. Interested [1372]. Interface [407, 1950, 3282, 3285, 348, 1471, 390, 410, 3006, 3021, 1750, 3281, 3101, 210]. Interfaces [2008, 212, 3273, 347, 3289, 637, 386, 393, 391, 211, 409, 403, 3269, 1677, 1987]. Interference [2436, 2467, 2188]. Interim [3121, 3153]. Interlaced [3205]. interleaving [30, 31]. Intermediaries [3208]. Intermediary [1136]. Intermediate [2555, 3030]. Intermediation [1159, 2826]. Internal [2938, 3440, 2282]. International [659]. Internationalized [148]. Internet [2955, 1385, 1386, 1658, 412, 76, 1119, 827, 3266, 174, 41, 644, 48, 108, 1973, 1177, 87, 61, 1149, 2538, 42, 2586, 2596, 2803, 162, 2605, 2447, 23, 2606, 676, 3391, 1676, 159, 1143, 178]. Interoperability [2711, 1120, 3207, 1509, 2749, 1240]. Interoperation [2797]. Interplane [993]. Interpolation [586, 1665, 596]. Interpretation [227, 1006, 266, 989, 1839, 3420, 206, 221, 220]. Interpreters [939]. Interpreting [2887, 223, 225]. Interprocedural [1227, 2004]. InterQoS [2538]. Interrupt [1433]. Intersecting [3147, 3073]. Intersection [2694, 3441]. Intersections [483]. Interstate [1131]. Interval [2258, 2262, 3403, 1708, 852, 2400, 588, 2960, 1224]. Intervals [2984, 2262, 2540]. Intervention [1882, 1824]. Interventional [1871]. Interventions [1879, 1799, 1861]. InterViewer [3080]. Interworkflow [1181]. Intestinal [1851, 1758]. Intra [1894, 1868, 1882, 1905, 1920, 2528]. Intra-domain [2528]. Intra-operative [1894, 1868, 1882, 1905]. Intra-patient [1920]. Intracranial [1800]. Intractable [498]. Intranet [1116]. Intraoperative [1791, 1869, 1893, 1908]. IntraText [1281]. Intrathoracic [1850]. Introducing [2221, 871, 670, 423, 2625]. Introduction [2135, 2136, 2137, 2138, 2139, 2140, 2141, 2142, 2314, 3314, 239, 532, 3117, 245, 2129, 1]. Intrusion [2619, 2622, 979, 2617, 112, 2490, 113, 2203, 2623, 2626, 2614, 1410, 2613, 3204, 2625]. Intrusion-Resilient [979]. Intrusion-Tolerant [1410]. Intrusions [2620]. Intuitive [991]. Invalid [1190]. Invalidation [335]. Invalidation-Based [335]. Invariant [2872, 2861, 2942]. Invariants [2549, 534, 2309]. Invasive [1879, 1771, 1762, 1882, 1783, 1775]. Invention [3198]. Inverse [2042]. Inversion [1885, 1102]. Inversions [1086]. Inverted [1260]. Investigating [1549, 1655, 2570]. Investigation [1263, 3445, 1883, 899]. Investigations [1840, 889]. Invited [1087, 1093]. Involvement [1172]. Involving [3450, 760]. IO [1524]. IP [2523, 72, 39, 166, 2047, 2326, 70, 88, 29, 71, 396, 51, 95, 2417, 2817, 2607, 111, 786, 1382, 195, 2822, 2453, 2517, 2525]. IP-over-WDM [39]. IP-Phone [396]. iPhone [62]. IPsec [118, 2335]. IPTPS’02 [659]. IPv6 [1383, 1384, 2586, 3257, 92, 2460]. IR [1368].

IR-Based [1368]. Irrational [2568]. Irreducible [2677]. Irregular [1513, 2685]. IRTL [2533]. IS-IS [32]. Isabelle [2570]. Island [1020]. ISO [2484]. Isomorphic [3053]. Isomorphism [497, 2567]. Isosurfaces [1926]. Issue [2374]. Issues [1146, 1626, 209, 3375, 2957, 1793, 1344, 1445]. IST-Project [1152]. IT-Architectures [1131]. IT-Support [1348]. Italian [1132]. Itemsets [743, 3441]. Iterated [869]. Iterative [744, 2861]. Iteratively [766]. Itineraries [2820]. ItLv [3276]. iTutor [392]. IV [2138, 2146].

J2EE [463]. J2ME [389]. Java [2763, 68, 1995, 2768, 1993, 3361, 1643, 460, 60, 59, 1420, 2765, 145, 1746]. Java-like [3361]. JCOD [1993]. Jenkins [2998]. Jini [60]. Jittered [2611]. JMX [146]. Job [2291, 3249, 3245, 914, 3246, 3243]. Job-Shop [914]. Jobs [1507, 3248]. Join [548, 244, 2845]. Joining [1597]. Joins [1364, 3406]. Joint [782, 1887, 1910, 1821]. Journals [2451, 1282]. Journey [1670]. JPEG [1195]. JPEG-2000 [1195]. jRate [2768]. Judicial [1165]. Junction [2860]. JUnit [431]. Just [386]. Just-in-Place [386]. JViews [3079].

K-Calculuses [2843]. K-Means [1575]. K-Nearest [3183]. Kademlia [663]. Kalman [990]. KANTOO [2109]. KDSM [2266]. Kernel [3210, 2927, 1040, 1700, 2415, 3408, 734, 778, 3322, 1903]. Kernel-Based [3408]. Kernels [3429]. Key [1385, 2207, 1386, 2181, 957, 2172, 2654, 2410, 110, 972, 2310, 786, 2482, 439, 1343]. Key-Exchange [957]. Keynote [1381]. Keys [2176, 869, 256, 1406, 1407]. Keystrokes [381]. Keyword [1234]. Keyword-Based [1234]. KF [2789]. KF-Diff [2789]. Kidney [2928, 1898, 1818]. KiMPA [1201]. Kind [1293]. Kinematic [3038]. Kinematics [1201, 1827].

Kinematics-Based [1201]. Kinetic [2020]. Kiosks [1130]. KIWI [1140]. KMsim [1450]. Knee [1781, 1790, 1887, 1883, 1870]. KNN [5, 6]. Knobs [425]. Knowledge [1028, 2925, 569, 2332, 1450, 1443, 692, 1451, 2835, 995, 207, 2933, 2791, 2963, 3356, 3411, 1452, 1472, 1289, 2250, 1140, 750, 1163, 1446, 3004, 206, 1460, 308, 3190, 3028, 3307, 3184, 3196, 1449, 1453, 3467, 3178, 2095, 1206, 2310, 225, 2266, 3185, 570, 2786, 1180, 2784, 157, 2774, 1960, 1108]. Knowledge-Based [2835, 995, 207, 3356, 1163, 3028, 3467]. Knowledge-Driven [3411]. Knowledge-Intensive [1451]. Known [976, 1910]. Known-Plaintext [976]. Komm [1129]. Korean [2090]. Korean-Chinese [2090]. Kruppa [1058]. KSPC [381]. Kullback [3146, 1914].

L [2663, 493, 3245]. L-Conjecture [2663]. Labeled [3052, 2954, 729]. Labeling [1805, 1849]. Labelling [1018]. Labels [3014]. Laboratories [1269]. Lambda [320, 936, 938]. Lambda-Lifting [936]. LAMGAC [1504]. LAN [2535]. Landmark [2270, 661]. Landmark-Based [2270]. Landmarks [1671]. Landscapes [843]. Language [2545, 904, 1732, 2105, 2833, 1280, 1275, 2452, 286, 1248, 3016, 1198, 264, 567, 631, 1197, 2711, 2084, 180, 2843, 244, 2775, 2792, 1317, 258, 3396, 2109, 1286, 3021, 555, 1945, 400, 2504, 1564, 247, 3395, 3329, 1194, 253, 2634]. Languages [638, 1735, 255, 3343, 3355, 2337, 3432, 2300, 1268, 3351, 552, 3141, 3133, 3140, 2098, 1303]. Laparoscopic [1752, 1754, 1758, 1756, 1751, 1867]. Laparoscopy [1893]. Large [741, 1688, 428, 742, 2630, 72, 2369, 2491, 462, 3164, 153, 3192, 3127, 762, 120, 45, 130, 2753, 2229, 878, 2513, 1698, 1095, 666, 769, 2956, 2096, 3070, 2389, 668, 1951]. Large-Scale [72, 2513, 1698, 1095, 666, 668].

Larger [982]. Largest [3128]. Laser [1864]. Lasers [2409]. Last [1414]. Latency [82, 135]. Latent [1788]. Lateral [1809]. Lattice [2031, 2020]. Lattices [3443, 2310, 3395]. Law [821]. Layer [76, 73, 1212, 670, 2533, 2065, 1845, 66, 189]. Layered [3071, 2071, 2419]. Layering [319, 3055]. Layout [3079, 2981]. Lazy [1594, 2493, 596]. Lazy-Formalization [596]. Lead [1460]. Leaf [1101]. Learn [3007]. Learnability [3136]. Learnable [3135]. Learned [2912]. Learner [702]. Learners [10]. Learning [2952, 3123, 2984, 3443, 2446, 2468, 3007, 1715, 742, 1174, 3125, 890, 2270, 2081, 694, 720, 871, 333, 905, 3148, 3166, 3156, 2997, 746, 1572, 2921, 1725, 3162, 2969, 614, 1705, 1199, 3138, 441, 3003, 3126, 3022, 1222, 2968, 3118, 3150, 705, 894, 3171, 854, 3464, 2618, 758, 1210, 1117, 707, 3130, 3129, 10, 708, 1026, 709, 2203, 885, 1046, 2896, 285, 3137, 3451, 3440, 715, 2093, 3345, 1559, 2971, 2998, 2915, 1037, 718, 400, 1726, 3001, 3024, 724, 734, 778, 3124, 725, 735, 779, 3426, 3160, 3140, 1704, 3430, 729, 2725, 3455, 3122, 3154, 1639, 3180, 730]. Learning [3442, 3431, 775, 2967, 1309, 2458, 3163, 2456]. Learnt [1715]. Least [1918, 1033]. Leave [1943]. Leaves [1331]. Lecture [124, 1087, 1093]. Led [1243]. Left [1836, 1826]. Legacy [430, 1190, 795, 446]. Legal [1126, 1174, 2617, 1115]. Legislation [2458]. Legislature [1178]. Leibler [3146, 1914]. Lempel [1653]. Length [3454, 3420, 3159, 1096, 3405, 37, 510, 712, 1088, 3073, 3380, 576, 63]. Length-Constrained [510]. Lengths [1507, 3063]. Lens [2503]. LEO [109]. Lesions [2945, 1839]. Less [105]. Level [1873, 2758, 3251, 2325, 2084, 1897, 1274, 206, 1833, 539, 1024, 99, 2449, 225, 2062, 3447, 3101, 1924, 1899, 2625, 1853, 2578, 278, 1118, 1954, 731]. Level-Set [1853]. Levelized [2444]. Levels [2541, 1510].

Levels-of-Service [2541]. Levenshtein [1073]. Leveraging [2758, 2742, 399]. Lexical [1442, 1281, 2109]. Lexicon [2097, 2086, 2098]. Lexicons [2087]. Libraries [1236, 1279, 1237, 3273, 3274, 3282, 3285, 1248, 3289, 3280, 3275, 1276, 1478, 1274, 1253, 1268, 1244, 2590, 2648, 3288, 2853, 1269, 1258]. Library [1252, 1261, 1280, 1275, 2452, 1273, 1251, 1529, 1528, 1643, 1239, 1271, 1521, 1246, 1816, 393, 1257, 1281, 3067, 1526, 1270, 1272, 1522, 1265, 3122, 3154, 164]. LICO [1529]. Life [2319, 1607, 434, 2021, 1166]. Life-Cycle [2319]. Life-Event [1166]. LifeCycle [897]. Lifting [2169, 936]. Light [1045, 432, 1115]. Light-Fields [1045]. Lightpath [36]. Lightweight [1993, 3325, 1411, 3330, 3288, 1366]. Like [486, 3038, 2161, 3284, 3380, 3409, 2127, 3361]. Limitations [469]. Limited [2081, 2662, 913]. Limiting [873]. Limits [300, 3269]. Line [903, 3414, 212, 1048, 413, 1347, 3050, 697, 3075, 3110, 964, 714, 1925, 1153, 879, 803, 1620, 225, 231, 2453, 487]. Line-Crawling [1620]. Lineage [3422]. Linear [2166, 3462, 3077, 484, 1640, 980, 983, 3128, 3399, 1043, 3075, 259, 2912, 1193, 702, 938, 3167, 344, 1719, 2308, 2161, 1864, 2569, 2637, 1924, 85, 2941, 2418]. Linearizability [2493]. Linearly [938]. Lines [226, 3049, 3144, 1141]. Lingo [610]. Lingually [2097]. Linguistic [2978, 832, 632]. Link [2697, 1754, 2293, 106, 3389]. Link-Failure [2697]. Linkage [871, 2264]. Linkages [2684]. Linked [3452]. Linking [3324, 1551]. Links [84, 123]. Linux [2415, 54, 2339, 3322]. Lipton [2377]. Liquid [3457]. List [3379]. Lists [2652, 1393]. Literary [1277]. Literature [2132, 1969]. Lithuania [1175]. Live [426, 1840]. Lived [2588]. Liveness [322, 3312]. Liver [1774, 1762, 1901, 1910, 1855, 1867]. ’Living [2739, 1346, 1966, 152]. Load [3221, 1496, 2821, 57, 3220, 1538, 2470, 1515, 1970, 2257]. Load-Balance [3221]. Loading [3205]. LOC [2887]. Local [896, 473, 1135, 1042, 1657, 3252, 22, 293, 839, 1132, 1594, 1109, 2466, 158, 2397, 543, 3188, 1147, 1671, 724, 2405, 1129, 2569, 1921, 3447, 2967, 1108]. Locality [174, 2640]. Localization [319, 2945, 1035, 3041, 2340, 1624]. Localizing [2902]. Locally [1535, 1998, 497, 3366]. Locating [680, 510]. Location [1322, 1324, 378, 382, 73, 1213, 1333, 410, 83, 89, 2435, 3271, 2745, 370, 1339, 1883, 97, 90, 1341, 178, 3260, 85]. Location-Aware [378]. Location-Based [382, 410]. Lock [2364]. Lock-Free [2364]. Locomotion [2916, 2873]. Log [2871, 825, 1239, 497, 618]. log-Bounded [497]. Log-Polar [2871]. Logarithm [2169]. Logarithmic [1002]. Logging [128, 1239]. Logic [595, 571, 612, 295, 932, 2545, 575, 2639, 273, 2215, 794, 299, 2833, 3113, 1735, 2958, 2472, 320, 541, 297, 2963, 567, 2320, 270, 2804, 2552, 3300, 619, 309, 610, 3139, 2975, 259, 598, 3465, 599, 308, 3202, 607, 1191, 3365, 579, 543, 221, 298, 2126, 1558, 941, 937, 300, 1449, 285, 3137, 2441, 301, 304, 296, 276, 287, 945, 3098, 2124, 1701, 2125, 604, 601, 2007, 544, 2569, 602, 1428, 1567, 3398, 596]. Logic-Based [299, 2958]. Logical [2215, 2225, 289, 311, 283, 2295, 3167, 2223, 312, 3212, 284]. Logically [2380]. Logics [580, 2130, 573, 2712, 2555, 583, 2557, 594, 588, 521, 2565, 530, 2129, 570, 2572, 2965]. Logistical [3374]. LogoMedia [2103]. Logs [3186]. LoL [377]. Long [2937, 1633, 2588, 758, 3378, 2610]. Long-Lived [2588]. Long-Running [3378]. Long-Term [2937, 758]. Longest [3191]. Longitudinal [1794]. Look [281, 16, 143]. Look-Ahead [16, 143]. Looking [2153, 1941]. Lookmark [2404]. Lookup [2589, 673, 664, 29, 2417, 2453]. Looping [1964]. Loops [284, 2021]. Loosely [1997]. Lord [2344]. Loss [2513, 3389]. Lossy [2448, 517, 1929]. Lossy-Compressed [1929]. Lost [382]. LOTOS [3108]. Low [1611, 742, 2589, 82, 39, 206, 2100, 477, 2069, 2437, 406, 1049]. Low-Complexity [2589]. Low-Cost [2100, 2437]. Low-Level [206]. Low-Paying [1611]. Low-Resolution [406]. Low-Weight [2069]. Lower [482, 2839, 2357, 501, 504, 1881, 3424]. LSD [951]. LSP [14]. LTL [3109]. LUC [975]. Luminous [1503]. Lung [1849, 1841]. LV [1842]. LV-RV [1842].

m [2206, 3265]. m-Business [2206]. M-services [3265]. M2D2 [2616]. Ma [1279]. MAC [47]. Machine [2102, 2104, 3416, 333, 3403, 3156, 2107, 3003, 554, 1481, 2088, 2968, 2089, 3464, 1503, 2090, 709, 2091, 2092, 2035, 714, 3451, 2093, 2998, 2100, 1750, 3001, 2339, 2101, 1864, 3143, 2923, 2098, 1639, 3180, 60]. Machine-Checking [3403]. Machines [241, 492, 2874, 1193, 2939, 2669, 1438, 3405, 1900, 2990, 2372, 719, 722, 3292, 728, 1704, 2200]. Macro [720, 1741, 81, 2023]. Macro-Operators [720]. Macroblock [2414]. Macros [1741]. Macroscopic [2038, 666, 1884]. Made [535, 310]. MAF [2801]. MAFRA [1464]. Magic [1191]. Magnetic [1825, 1912, 1809, 1933]. Magneto [1867]. Magneto-optic [1867]. Magnitude [2585]. Magnitudes [2870]. Main [139]. Mainstream [1483]. Maintaining [626, 3047, 1200]. Maintenance [2344, 2546, 2231, 1463, 2109, 392]. Maiorana [982]. Make [697, 667]. Makespan [1340]. Making [3412, 1447, 1708, 1483, 879, 3398]. Malformed [1777, 1877]. Malicious [2612]. Mamdani [3453]. Mammograms [2255]. Mammography [1787]. Man [1864, 2374, 2923]. Man-Machine [1864].

Manage [1135, 1937]. Managed [3411]. Management [262, 426, 2332, 1450, 2808, 3335, 1451, 2046, 2336, 68, 1964, 2933, 82, 2603, 73, 1137, 3343, 3251, 2061, 2045, 3082, 161, 1125, 1401, 1402, 1403, 1404, 791, 1140, 1974, 2047, 2211, 2056, 66, 83, 3301, 137, 2821, 1531, 3231, 1246, 2501, 2807, 69, 1448, 1369, 1946, 116, 146, 2435, 2745, 1484, 1596, 1453, 3330, 2817, 2851, 2406, 3254, 1744, 111, 2066, 2764, 2820, 2331, 2806, 1410, 90, 1354, 3328, 1468, 2784, 3230, 2822, 2049, 1960, 2592, 3009, 3011, 2048, 685, 1108, 2801, 1134]. Manager [2050, 1162]. Managerial [2730]. Managing [1470, 2737, 2323, 1748, 624]. Maneuverability [3028]. Manhattan [2682]. Manifold [1878, 864]. Manipulation [1769, 1746]. Manipulator [1752, 1756, 1762, 1764, 1763, 1773, 1760]. Manual [1768]. Manufacturing [1351, 1346, 2696, 1343, 1344]. Manuscript [233]. Many [2160, 1428]. Many-Valued [1428]. Map [1578, 413, 3047, 2476, 3175, 1666, 2900, 3284, 216, 1669, 1821]. Map-Based [413]. Map-Like [3284]. Mapped [1915]. Mapping [3083, 2872, 2788, 2431, 2781, 1232, 2223, 666, 2016, 1464]. Mappings [1548]. Maps [2989, 849, 218, 474, 221, 1925, 880, 3041, 1601, 1810, 220]. Margin [3128, 703, 3127]. Margin-Based [703]. Marginal [869]. MARIAN [1643]. MARISM [2816]. MARISM-A [2816]. Mark [282]. Markers [2261]. Market [2539, 3179, 2703]. Market-Based [2539]. Marking [2061, 48, 2062, 2454, 2515, 2063, 2517]. Markov [549, 3372, 2440, 706, 1204, 1898, 3015, 723, 920, 1010, 2416]. Markup [2711, 1475]. Marrow [1773]. MASA [3006]. Masking [980, 2193]. Mass [1068, 1876, 138]. Master [1756]. Match [1072]. Matches [2735]. Matching [1629, 1649, 207, 1640, 1393, 1894, 1630, 1628, 578, 1642, 1869, 2679, 1026, 3043, 223, 1538, 2403, 1921, 1641, 1510, 1850, 1023]. Material [3023, 1827]. Materialization [1228]. Materials [10]. Mathematical [1723, 1064, 2843, 224, 3459]. Mathematics [3118, 3150, 1287]. MATLAB [1512]. MATLAB-MPI [1512]. Matrices [757, 1056, 232]. Matrix [1041, 1224, 2275, 2695, 2981, 3120, 3152, 1602]. Matroid [483]. Matter [1803, 1806]. Matters [1127]. Maturity [798]. Maude [3108]. Max [2668, 2966, 1709, 1514, 2696, 3447]. Max-2SAT [2668]. Max-Closure [1514]. Max-CSP [2966]. MAX-SAT [3447]. Maximal [3181, 2408]. Maximization [1819, 1786]. Maximize [131]. Maximizing [3125, 2705, 494]. Maximum [1227, 2611, 1330, 1331, 1074, 1047, 1337, 2086]. Maximum-Density [1074]. MB [1203]. McCain [584]. McFarland [982]. MCM [1345]. MDA [1309, 2293, 1295]. Meaning [3438, 3223]. Meaningful [2846, 2702, 1571]. Meanings [633]. Means [2978, 2831, 1189, 3167, 1575, 2500, 2947]. Measure [1073, 1053, 158, 1910, 1907]. Measurement [2521, 1862, 1775, 70, 1776, 2535, 674, 2760, 2892]. Measurement-Based [2521]. Measurements [1784, 1804, 2511, 108, 2607, 127]. Measures [2547, 1048, 870, 1552, 1911, 1912, 1580, 2266, 844]. Measuring [2836, 2330, 1465, 1313, 840]. Mechanical [3113, 1784, 2644]. Mechanics [1843, 2020]. Mechanism [38, 2364, 1778, 2847, 2807, 113, 17, 2512, 81]. Mechanisms [2542, 2888, 2533, 1759, 3382, 2055]. MEDIA} [1129, 2052, 2529, 1964, 2599, 21, 12, 1263, 122, 2608, 2755, 18, 2605, 126]. Media-on-Demand [1964]. Medial [2880]. Median [2995, 1008]. Medians [1102]. Mediated [1466, 2273]. Mediating [1254]. Mediation [2227, 102]. Mediator [2734, 644, 1976, 2736, 1668]. Medical [2925, 1792, 1873, 752, 1614, 1449, 2929, 1923, 1820, 771, 1924, 3452, 2892]. Meet [422, 1220]. Meets [433]. MEGA [2568]. MEGICO [1453]. Meiosis [1215]. Melanoma [1612]. MELL [2559]. Mellon [1259]. Membership [2354, 626, 1434, 1694, 509]. Membrane [2388]. Memetic [915, 889]. Memoization [1733]. Memories [3002]. Memory [323, 2645, 1507, 3415, 2415, 139, 2353, 2356, 1994, 769, 2699, 2005]. Memory-Efficient [2699]. Memoryless [2117, 981]. Mendelian [2935]. Mental [3317, 3047]. Menu [416]. Mere [2007]. Mereological [503, 1555, 1624]. Merging [569, 487, 2488, 864, 2091, 2727]. Merkle [1393, 1394]. Mesh [2439, 1805]. Meshes [1878]. Message [128, 2548, 2350, 3420, 3405, 1530, 539, 2100, 2513, 326]. Message-Passing [2350]. Messaging [2724, 404, 2823]. Meta [1450, 2218, 1356, 3010, 2794, 2501, 3163, 136, 1256, 1469, 1747, 731]. Meta-computing [2501]. Meta-data [1256]. Meta-learning [3163]. Meta-level [731]. Meta-model [2218]. Meta-modeling [1356]. Meta-modelling [1450, 3010]. Meta-programming [1747]. Meta-properties [1469]. Meta-search [2794]. Metabolic [1082]. Metacomputer [1480]. Metacomputers [3253]. Metacomputing [1479]. Metadata [1935, 1473, 2428, 765, 1257, 254, 2793, 1676, 3277, 1258]. MetaDL [1257]. Metae [1282]. Metaheuristics [898]. Metalevel [388]. Metamodel [1304, 1308, 872, 1292, 1286]. Metamodel-Assisted [872]. Metamodeling [1287]. Metamodels [2209]. Metaphors [1471, 390]. Metaprogramming [1735]. Metasearch [1196]. Metering [2154]. Method [3054, 2439, 117, 1228, 1090, 1586, 207, 294, 2891, 1707, 919, 2748, 1291, 2084, 1125, 1897, 3386, 1862, 224, 1007, 700, 1612, 1223, 1927, 1702, 1185, 1191, 2429, 1435, 2476, 84, 3110, 2438, 3451, 1056, 1598, 98, 230, 768, 1078, 3111, 1566, 1201, 431, 33, 77, 1034, 232, 1818, 1077, 126, 85]. Methodological [1360]. Methodologies [422, 954]. Methodology [1696, 2978, 1362, 2336, 3438, 1453, 3034, 2259, 2266, 2785, 1373]. Methods [426, 801, 1042, 1585, 349, 3156, 3165, 929, 2638, 3003, 1116, 1927, 1605, 1345, 29, 3408, 1665, 1667, 437, 1719, 332, 2229, 1088, 2329, 1079, 1421, 436, 1066, 2957, 862, 1070, 2967]. Metric [1649, 2423, 3433, 2968, 1339, 663, 1651]. Metrics [1754, 749, 3433, 798, 2330, 1936, 425, 1902, 725, 1313]. Metrology [1015]. Metropolis [2008, 1986]. MIB [2057]. MICO [2746]. Micro [1770, 1772, 1763, 81, 1966]. Micro-grasping [1763]. Micro-hydraulic [1772]. Micro-robotic [1770]. Microaneurysms [2947]. Microarray [1094, 3172]. Microarrays [1095]. MicroBill [3393]. Microcalcification [2927]. Microcalcifications [2255]. Microcircuit [2886]. Microeconomic [170]. Microorganisms [1039]. Microprocessor [2635, 3112]. Microprocessors [2636]. Microscope [1854]. Microscopy [2928]. Microsoft [2107]. Microsurgery [1770]. Microwave [1855]. MIDAS [1360]. MIDAS/BD [1360]. Midbit [3124]. Middleware [2756, 2761, 2582, 2769, 2752, 2762, 2533, 1296, 1246, 3271, 2602, 2074, 2806, 1999, 2606, 2771, 3392, 3264, 144]. Middleware1 [640]. Migration [2812, 135, 1267, 2504, 2029, 1480]. Migratory [388]. Millennium [2209, 2608, 1142]. Min [1329, 2696, 2692]. Min-Max [2696]. Min-sum [1329]. Mincut [1103]. Mind [2786]. Mine [3441]. Miner [613]. Mini [1783]. Mini-invasive [1783]. Minimal [1085, 2258, 2177, 2657, 587, 2384, 2362, 1214, 806, 3435]. Minimally [1879, 1771, 1762, 1882]. Minimax [3146, 1210]. Minimised [3139]. Minimising [1914]. Minimization [2700, 3057, 364, 1096, 885, 1340, 1034]. Minimizing [1337, 2063]. Minimum [2658, 1089, 3420, 1333, 1335, 476, 2682, 712, 2689, 380, 2370, 3380]. Minimum-Diameter [1333]. Minimum-Evolution [1089]. Minimum-Model [2370]. Mining [2925, 739, 154, 740, 743, 255, 2840, 748, 611, 3441, 1612, 755, 759, 3119, 3151, 2400, 1545, 1959, 766, 1607, 1596, 3186, 3193, 254, 768, 1599, 613, 2739, 771, 1577, 772, 1608, 3191, 1603, 685]. Ministerial [1116]. Minor [1327, 2675]. MIPv6 [2581]. MIRA [2535]. MIRAGE [3316, 3318]. Missing [2293]. Mission [428, 2615, 3413]. Mission-Impact-Based [2615]. Mistakes [424]. Misuse [1680, 2490]. Mitochondrial [1619]. Mixed [3019, 376, 1043, 934, 1263, 2164]. Mixed-Media [1263]. Mixing [351, 2178]. Mixture [3144, 2280, 735, 779]. MLP [1713]. MLR [731]. MML [3407, 3406]. Mnemosyne [203, 671]. MnM [1475]. Mobile [407, 2810, 105, 3206, 1962, 2461, 3200, 2717, 2446, 2763, 3210, 2813, 1032, 2913, 413, 3359, 3209, 3211, 1211, 382, 76, 3201, 2802, 2804, 376, 2582, 161, 244, 369, 545, 2812, 3207, 135, 2399, 119, 3202, 410, 379, 2448, 64, 137, 2718, 3265, 88, 401, 2476, 2067, 2490, 2462, 386, 396, 140, 3256, 84, 2815, 3262, 2716, 103, 94, 123, 146, 2435, 3271, 2745, 2498, 95, 2728, 1411, 374, 2026, 2074, 2287, 2198, 2817, 366, 1381, 1037, 384, 55, 394, 2825, 383, 2534, 2597, 367, 2532, 3257, 893, 377, 3332, 92, 2460, 1670, 822, 2820, 2816, 1382, 2806, 3203, 409]. Mobile [188, 1975, 90, 2527, 404, 3205, 2814, 77, 102, 27, 405, 75, 2811, 3269, 2434, 2824, 3267, 2822, 397, 3263, 85, 3208, 399]. Mobile-Agent [3207]. Mobile-Agent-Based [2825]. Mobility [82, 3265, 186, 2807, 547, 81, 2597, 167, 3212, 3333]. MobiMine [757]. Modal [2122, 532, 411, 2125, 585, 584, 2121, 1039, 1914, 1982, 2882, 1866]. Modalities [1837]. Mode [784, 1533, 2384]. Model [535, 360, 262, 1539, 338, 559, 2461, 1352, 3105, 147, 2741, 173, 3316, 321, 794, 1444, 1549, 1611, 1682, 334, 2244, 923, 275, 3094, 2913, 2196, 1784, 556, 1659, 389, 1488, 277, 323, 2630, 354, 2231, 2233, 342, 651, 1796, 1916, 1437, 2156, 2377, 252, 1379, 1302, 2959, 2250, 1431, 791, 2642, 2638, 3107, 2788, 653, 324, 2752, 2620, 3299, 3415, 2173, 807, 357, 67, 346, 1702, 2721, 3161, 1904, 314, 2394, 2159, 2908, 2380, 336, 798, 2935, 278, 607, 137, 1756, 1557, 2330, 1912, 141, 69, 2490, 3309, 1438, 2899, 1816, 140, 542, 2849, 897, 1890]. Model [2910, 344, 361, 1221, 1244, 56, 2850, 1901, 543, 116, 24, 3226, 1606, 2223, 1317, 1889, 912, 2728, 2031, 1887, 3169, 332, 656, 2287, 1905, 3015, 2229, 2616, 2018, 2866, 2903, 2370, 3170, 1779, 1021, 1876, 803, 2900, 1842, 1181, 1895, 276, 1432, 3222, 2232, 1636, 2304, 3292, 345, 981, 657, 1929, 2029, 1837, 821, 2036, 1511, 3353, 648, 350, 2868, 771, 859, 649, 1577, 3033, 2213, 2312, 2994, 2416, 900, 2646, 2218]. Model-Based [791, 3299, 1702, 1816, 803, 657, 3353, 900]. Model-Building [859]. Model-Checker [354]. Model-Checking [323, 252, 807, 336]. Model-Driven [1317, 656, 649]. Modeling [3370, 2008, 2932, 2207, 2318, 320, 3463, 2242, 1873, 2933, 1989, 2715, 1930, 629, 2325, 2711, 1000, 625, 2243, 1740, 1452, 156, 1376, 1356, 1290, 3003, 1986, 2219, 346, 2911, 2412, 3175, 2505, 1662, 3369, 2636, 24, 1891, 1317, 1307, 2228, 1693, 1118, 627, 3404, 2778, 2252, 1286, 2894, 42, 1842, 2532, 1305, 2459, 1674, 3410, 2936, 1423, 2414, 1884, 1303, 3260, 1960, 2760, 2048]. Modelization [1785]. Modelling [1306, 868, 1723, 2473, 3303, 801, 2274, 2038, 1880, 1291, 3417, 1825, 2208, 2210, 1996, 2209, 1357, 1124, 1544, 109, 391, 3023, 2427, 307, 3446, 1373, 75, 3100, 2205, 2892, 2303, 1445, 1450, 3010]. Modells [1010]. Models [2931, 1028, 2978, 1680, 3358, 650, 692, 2215, 2225, 1350, 1135, 291, 1715, 992, 1888, 323, 2677, 2552, 3106, 3165, 1894, 3439, 909, 2440, 988, 2226, 2779, 2935, 1296, 949, 863, 1345, 2879, 1047, 267, 2201, 1268, 3119, 3151, 1004, 654, 790, 292, 2896, 1937, 2886, 414, 2240, 2012, 2254, 1581, 3096, 2206, 2280, 1948, 2710, 326, 3001, 3157, 2239, 2867, 1675, 735, 779, 920, 2477, 899, 3036, 2086, 1287, 772, 2043, 3431, 1892]. Modern [1643]. Modification [1557, 2432, 1008]. Modifications [3081, 892, 308]. Modified [1039, 2516, 73, 57, 1103, 2967]. Modifiers [2960]. Modular [3104, 811, 1737, 1518, 1993, 327, 1209, 2747, 1226, 1310, 3436]. Modular-Fuzzy [1209]. Modulation [99]. Module [1057, 54]. Modules [348]. Modulo [974, 507, 592]. Molecular [1084, 1840, 1515]. Molecule [1093]. Molecules [1076]. MON [2326]. Monadic [2127, 485, 938, 2126, 2125]. Monads [240]. Monge [480]. Monitor [20]. Monitoring [2064, 3331, 2612, 3237, 1493, 2813, 2057, 1718, 2326, 2355, 3437, 186, 3428, 2607, 2940, 803, 2327]. Monkey [2885]. Monocular [1032]. Monogenic [1017]. Monolingual [2096]. Monotone [3123, 255, 2708]. Monotonic [1573, 612]. Monte [1792, 3215, 2459, 3041]. Moorea [2804]. Morphable [2908, 2910]. Morphisms [522]. Morphological [3002, 2939, 2424]. Morphometry [1844]. Most [436, 770]. Motifs [1073]. Motion [1352, 3414, 1044, 2875, 997, 1765, 1351, 1615, 2878, 1664, 2585, 2252, 1056, 1061, 2279, 1838, 1928, 2414, 1510, 2902, 2402, 2873]. Motions [2684]. Motivated [2860]. Motivation [3273, 3332]. Motivational [3035]. Motor [3148, 2101]. Mounted [1854]. Mouth [371]. Move [88]. Movement [2911, 77]. Movements [2912, 2874]. MOver [3236]. Moving [3127, 1664, 3260]. MPEG [2583]. MPEG-4 [2583]. MPI [1512, 1535, 1527, 1533, 1524, 1488, 1518, 1521, 1487, 1482, 1504, 1538, 1509, 1523, 1526, 1515, 1534, 1537, 1520]. MPI-2 [1535, 1504]. MPI-IO [1524]. MPICH [1526]. MPICH-CM [1526]. MPICH2 [1482]. MPLS [2530, 72, 82, 14, 16, 57, 52, 2460, 2531, 3389, 32]. MPLS-ATM [2531]. MPLS-Based [72]. MR [1830, 1817, 1808, 1785, 1869, 1762, 1816, 1823, 1891, 1828, 1885, 1763, 1855, 1834, 1819, 1907, 1760, 1787, 1906, 1852, 1788]. MR-Compatible [1762]. MR-Guided [1763, 1855]. MR-Images [1816]. MR-Mammography [1787]. MRA [1911]. MRCP [1933]. MRD [20]. MRF [1821]. MRI [1803, 1813, 1761, 1807, 1843, 1869, 1836, 1764, 1835, 1806, 1794]. mRNA [1077]. MS [1794]. MSR [2107]. MSR-MT [2107]. MT [2081, 2099, 2107, 2087, 2109, 2095]. Mu [532, 525]. Mu-Calculi [532]. Mu-Calculus [525]. Much [1570, 679]. Multi [2635, 2523, 2984, 2024, 2473, 1039, 2270, 354, 3412, 1914, 568, 1529, 567, 1968, 577, 581, 2578, 3233, 2969, 614, 1011, 1209, 1642, 2359, 66, 853, 278, 2821, 863, 2271, 1505, 104, 865, 1210, 100, 189, 760, 2662, 542, 1982, 2882, 866, 842, 1666, 1859, 1833, 2801, 95, 2728, 1118, 301, 2962, 918, 99, 1866, 107, 3466, 867, 2436, 2467, 307, 570, 1954, 2069, 2524, 2600, 33, 648, 1860, 1641, 3406, 1373, 2771, 1548, 1748, 2720, 1583, 1397, 1398, 2303, 2980, 3011, 395, 1669]. Multi-adaptive [2984]. Multi-adjoint [301]. Multi-agent [2270, 3412, 568, 567, 577, 581, 1209, 2821, 2271, 1210, 1666, 307, 570, 2720, 3011]. Multi-agents [1505]. Multi-attribute [1642]. Multi-byte [1641]. Multi-carrier [2436]. Multi-code [100, 99]. Multi-criteria [918]. Multi-device [395]. Multi-domain [33]. Multi-group [107]. Multi-layer [66, 189]. Multi-level [2578, 278, 1118, 1954]. Multi-management [2801]. Multi-modal [1039, 1914, 1982, 2882, 1866]. Multi-objective [863, 865, 866, 842, 867, 2980]. Multi-OS [1397, 1398]. Multi-paradigm [2962, 2303]. Multi-path [2523, 95, 2467, 2524]. Multi-perspective [1373]. Multi-platform [1529, 648]. Multi-projector [1859]. Multi-prover [2662]. Multi-purpose [614]. Multi-relational [760]. Multi-resolution [2600]. Multi-robot [3466]. Multi-scale [1833, 1669]. Multi-sender [2069]. Multi-sensor [1011]. Multi-site [3233]. Multi-source [104, 1583]. Multi-sourced [1860]. Multi-stage [1748]. Multi-step [2635, 853]. Multi-trees [2969]. Multi-user [2771]. Multi-valued [354, 542, 1548]. Multi-vehicle [1982]. Multi-version [2728]. Multi-way [3406]. Multi-word [2359]. Multiagent [3266, 3307, 3006, 2475, 718, 149]. Multiagents [134, 3020]. Multicamera [1049]. Multicarrier [2500]. Multicast [105, 22, 2057, 2518, 2071, 104, 17, 15, 120, 20, 3099, 3387, 2075, 19, 2597, 2070, 18, 86, 35, 47, 2437, 33, 77, 27, 3391]. Multicasting [38, 2435, 2449]. Multicasts [2537]. Multichannel [47]. Multiclass [3145, 704]. Multicluster [3252]. Multicolorings [516]. Multiconstraint [3390]. Multicriteria [1708]. Multidatabase [2852, 2710]. Multidimensional [1248, 1307, 2228, 1937]. Multifractal [173]. Multigraphs [1330]. Multilayer [902]. Multilayered [2024, 2026]. Multilevel [1322]. Multilingual [2452, 3182]. Multimedia [2810, 1955, 2046, 1218, 2045, 2578, 2582, 2047, 369, 1967, 119, 2580, 2067, 2415, 2465, 2594, 1241, 2602, 2583, 2601, 2596, 54, 3297, 1954, 2605, 125, 127, 75, 2604, 1963, 1957, 2413, 3298, 2049]. Multimeme [911]. Multimodal [1916, 988, 1046, 392]. Multinational [1139]. Multinomial [693].

Multiobjective [905, 864].

Multiparadigm [3355]. Multiparticle [2031]. Multiparty [959, 2592]. Multiple [3183, 628, 413, 890, 831, 150, 1631, 1290, 2425, 2988, 1772, 2541, 2397, 84, 3247, 31, 2946, 1081, 2467, 1194, 2239, 1080, 1809, 1645, 195, 921, 2965]. Multiple-Access [2467]. Multiple-Queue [3247]. Multiple-Segments [1772]. Multiple-Valued [2965]. Multiplication [2695]. Multiplicity [511]. Multiprocessor [2686, 494, 2687]. MultiProt [1080]. Multireader [2349]. Multirelational [720, 2217]. Multiresolution [1805, 1876, 3295]. Multiscale [1060, 2611, 908, 752]. Multiset [1427]. Multisource [3295]. Multispectral [993]. Multistrategy [716]. Multithreaded [327]. Multivalued [3000]. Multivariable [2930]. Municipal [1136, 1128]. Municipalities [1127]. MuniS [1136]. Mur [2640]. Muscle [1714, 1780]. Music [233, 1264, 3185]. Musical [1613, 3188, 3121, 3153]. Musicology [3185]. Mutability [191]. Mutation [848, 850, 846, 857, 860]. Mutation-Based [846]. Mutative [856]. Mutual [3209, 1916, 885, 1915]. MVC [2222]. Myocardial [1829]. Myth [2324].

n [2176, 3383]. N-Body [3383]. NADIA [23]. Naive [2952, 2949]. Nakagami [2467]. Name [672, 559]. Named [3017, 3177, 1199]. Names [148]. Naming [2790, 191]. Narratives [2234]. Narrow [1770]. Narrowing [941]. nasal [1764]. Nash [599, 3430]. Nash-Consistent [599]. Natural [904, 2104, 2843, 2898, 2792, 3021, 2567, 2873]. Nature [2315]. Navigating [1247]. Navigation [1799, 1774, 2270, 382, 1854, 1862, 2844, 1856, 1790, 1865, 1863, 1624, 1963]. Navigational [402, 406]. Near [2467]. Near/Far [2467]. Nearest [1977, 3183, 2839, 2267, 1515, 2980]. Nearest-Neighbor [2839, 1515]. Nearly [2862]. Nearness [3401]. Necessary [2360]. Necessity [1486, 1552]. NED [2744]. Need [742, 387, 369, 1628, 2212, 3312]. Needed [2322, 1111]. Needle [1880, 1768, 1765, 1776, 1782, 1863]. Needs [1276, 1291, 422, 394, 665]. Neem [2740]. Negation [294, 512, 604]. Negative [3141, 2275]. Negotiating [3251, 601]. Negotiation [2218, 2802, 1408, 1409, 3005, 2538, 2739, 2720]. Negotiator [2819]. Neighbor [1977, 3183, 2839, 1204, 1515, 1822]. Neighborhood [2401, 875, 2967]. NeighborNet [1090]. Neighbour [3326, 2267]. Neighbourhood [700, 2951]. Neighbours [2980]. Nerve [1845]. NESPOLE [2108]. Nested [3094, 491, 2299, 296, 2224]. Net [1253, 362, 301, 3236]. Net-Based [1253]. Netherlands [1106]. Netlets [2818]. Nets [2418, 807, 561, 245, 3101, 2303]. Network [2064, 2332, 2808, 3335, 2813, 2945, 68, 3458, 3427, 72, 82, 3385, 2045, 2818, 3325, 166, 670, 2540, 2579, 2334, 2812, 2621, 2622, 2326, 66, 1109, 1220, 2821, 40, 16, 3450, 3448, 2682, 534, 71, 112, 69, 1336, 3232, 2492, 146, 20, 45, 2626, 2602, 2417, 2481, 1411, 3000, 2995, 805, 2075, 2406, 2607, 91, 2996, 1621, 2820, 666, 2806, 3388, 2059, 65, 3376, 23, 1589, 3333, 674, 3426, 127, 102, 2054, 2604, 1713, 773, 3264, 668, 2664, 143, 101, 2416, 2049, 3011]. Network-Design [1336]. Network-Integrated [2602]. Networked [160, 3386, 2415, 2851]. Networking [2077, 3261, 1128, 26, 675, 3333, 2525]. Networks [2523, 2058, 2050, 2439, 105, 2530, 117, 1389, 3337, 2521, 2446, 1724, 960, 2976, 3327, 1090, 2681, 38, 131, 2680, 39, 3423, 907, 3148, 2999, 695, 1725, 3334, 2047, 1709, 2056, 2326, 681, 3270, 2724, 3167, 83, 64, 2466, 2495, 190, 3265, 2698, 104, 3438, 679, 3428, 2067, 2807, 57, 11, 189, 89, 51, 3258, 2577, 1706, 103, 2483, 2465, 2435, 2745, 1338, 660, 513, 2353, 1390, 1710, 2817, 2519, 2993, 3340, 1900, 1727, 2449, 2597, 1721, 1698, 1492, 1681, 3332, 1387, 1388, 1067, 822, 886, 3329, 2991, 2079, 684, 3168, 2595, 3142, 188, 2480, 47, 2527, 3339, 3328, 2340, 3084, 33, 36]. Networks [3389, 2063, 216, 1712, 2811, 1722, 195, 2499, 1577, 3259, 3436, 2037, 178, 1592, 50, 3425, 2822, 661, 2967, 3173, 1445]. Neural [2439, 2992, 2418, 2888, 1724, 890, 891, 2945, 3458, 1717, 3427, 3423, 3148, 1725, 3167, 875, 3450, 3438, 3448, 3428, 2168, 1706, 2886, 2417, 1710, 301, 2995, 1900, 1727, 1721, 1492, 1621, 1067, 2991, 3168, 3142, 2868, 1712, 1722, 1713, 773, 3436, 3425, 2416]. Neuro [2929, 3453, 1577]. Neuro-Fuzzy [1577, 2929]. Neuro-Genetic [3453]. Neuroanatomical [1846]. Neurocomputation [1581]. Neurocomputing [1620, 1581, 1582]. NeuroGrid [190]. Neuroimaging [1930]. Neurointerventions [1913]. Neuromorphic [2914]. Neurons [2864]. Neurosurgery [1789, 1857, 1764, 1763]. Neurovascular [1798]. Neutral [1667]. News [2576, 3296, 3280, 2097]. News-On-Demand [2576]. Next [3349, 2207, 3265, 124]. NF [1365]. NF-SS [1365]. NFA [2661]. NFAs [499]. NFS [1524]. NG [2326]. NG-MON [2326]. NLP [3013]. NLP-Based [3013]. NMR [1069]. NN [3454, 1715, 700]. No [3073]. NodeOS [3322]. Nodule [1847]. Nodules [1841]. Noise [1904, 1091]. Noisy [2954]. Nomadic [2819, 2824]. Nomenclature [1849]. NoMoRe [612, 605]. Non [2163, 229, 612, 2418, 2188, 1399, 743, 1916, 492, 2863, 2559, 2659, 2275, 1775, 1919, 30, 2276, 2013, 579, 1337, 1902, 31, 1828, 305, 589, 1918, 955, 1400, 1819, 1838, 2593, 2260, 2482, 2426, 1917, 1787, 2390, 1852, 3422]. Non-abusiveness [1337]. Non-anonymous [2482]. Non-binary [2260]. Non-classical [2863]. Non-committing [955].

Non-commutative [2559]. Non-commutativity [579]. Non-continuous [1852]. Non-CORBA [2426]. Non-Delaunay-Based [2659]. Non-derivable [743]. Non-euclidean [1902]. Non-interactive [2163]. Non-interference [2188]. Non-interleaving [30, 31]. Non-invasive [1775]. Non-linear [2418]. Non-monotonic [612]. Non-negative [2275]. Non-planar [229]. Non-preemptive [492]. Non-real-time [2593]. Non-regular [2276]. Non-repudiation [1399, 1400]. Non-rigid [1916, 1919, 1828, 1918, 1819, 1838, 1917, 1787]. Non-trivial [2013]. Non-uniform [305, 589, 2390, 3422]. Nondeterminism [1530]. Nondeterministic [482, 2119]. Nonlinear [2931, 1041, 1022, 2179, 346, 3028, 3408, 2996, 2994, 3455, 2941]. Nonmonotonic [586]. Nonparametric [877]. Nonrigid [1910, 1921]. Normal [603, 937, 1883, 1365]. Normalcy [1797]. Normalization [242, 940, 1196]. Normative [2506]. North [2039]. Notation [376]. Notations [2209, 1281]. Note [2547]. Notes [2143, 2144, 2145, 2146, 2147, 2148, 2149, 2150, 1530, 3281, 1240]. Notices [1165]. Notification [2332, 48]. Notion [2773]. Notions [933]. Novel [1774, 816, 1418, 1471, 3450, 56, 768, 1907, 1864, 1686]. NP [1431, 507]. NP-complete [507]. NP-Easy [1431]. NSF [3256]. NTIF [3107]. NTRU [961]. Nuanced [590]. Number [2180, 1086, 873, 2557, 2425, 2380, 975, 3451, 1597, 380]. Number-Conserving [2380]. Numbering [1940]. Numbers [2105, 243, 934, 65, 3063]. Numeric [2972, 2260]. Numerical [1443, 1478, 3222, 1499, 1495]. Nurse [919]. NuSMV [342].

O [3113, 1600, 1509, 1522]. OAI [1242]. OBGP [41]. Object [630, 2754, 2780, 634, 1228, 628, 2215, 2225, 3349, 2756, 3210, 2244, 638, 3359, 651, 629, 3356, 3355, 990, 1000, 3357, 1974, 3003, 2788, 2898, 2422, 2429, 2918, 644, 1217, 2879, 1047, 2747, 759, 2890, 1946, 790, 2946, 1046, 2896, 627, 3345, 1031, 2746, 1037, 1030, 3346, 3354, 1184, 2443, 1949, 2897, 2973, 2901, 2771, 1747, 2905, 2457, 2760, 2757]. Object-Based [2901]. Object-Orientation [3357]. Object-Oriented [630, 634, 628, 2225, 2244, 638, 651, 629, 3356, 3355, 2788, 2429, 644, 790, 3345, 3354, 2973, 2771, 2457]. Object-Relational [1228, 759, 1946, 1184, 1949]. Object-Specific [2897]. objective [863, 865, 866, 842, 867, 2980]. Objectives [890, 1219]. Objects [1875, 2763, 989, 2891, 1967, 1009, 1663, 1664, 2915, 2743, 2426, 210, 1265, 3260, 1583]. Oblivious [2671, 2159]. Observation [3171, 1246]. Observational [473, 556]. Observations [660, 1514]. Observatory [1157]. Observer [1929]. Obstacle [997, 1620]. Obtain [265]. Obtained [226, 1887]. OCaml [247]. Occluded [3306]. Occlusion [2889]. Occupancy [1759]. occurrence [3282]. OCL} [1301, 1289, 1300]. October [3256]. Octree [1924]. Oculogyration [1832]. Oculomotor [2906]. Off [903, 2590, 795, 46]. Off-Line [903]. Off-the-Shelf [795]. Offender [2364]. Offering [1146]. Offers [2720]. Offices [1128]. Official [1141]. Officials [1121]. Offline [1185]. Offspring [2988]. OIL [1372]. OKE [3324]. OLAP [2428, 254, 253]. Oldest [30]. Oligonucleotide [1100, 1095]. OLS [1043]. OLS-TLS [1043]. Olympic [2514]. Omni [3414, 1933]. Omni-directional [1933]. Omni-drive [3414]. On-Demand [2544, 2541, 33]. On-Line [212, 1048, 413, 964, 714, 879, 803, 487]. ON-OFF [46]. Once [482]. One [3064, 2017, 2157, 489, 2661, 1327, 697, 1122, 1146, 2174, 498, 1128, 70, 1123, 2016, 773].

One-Dimensional [3064, 2016]. One-Stop [1122, 1146, 1128, 1123]. One-Time [2174]. One-Way [2157, 70]. ONIGOCO [13]. Online [1130, 3416, 3427, 1471, 698, 1109, 1243, 3424, 1337, 1172, 393, 1281, 1876, 1231, 3023, 3001, 3032, 3029, 1278, 492]. Only [697, 2191]. onto [2016]. OntoEdit [2785]. OntoLog [1238]. Ontological [1353, 2219, 2779, 1469]. Ontologically [2236]. Ontologies [1470, 1455, 2833, 2791, 1471, 1372, 2843, 1459, 574, 2235, 1238, 2220, 1973, 1462, 1464, 1465, 2212, 1467, 1661]. Ontology [1454, 1442, 2734, 1457, 1212, 1458, 3004, 2208, 2210, 2775, 2433, 2742, 2787, 2781, 1461, 1463, 765, 2212, 1466, 2777, 2798, 2206, 1468, 2784, 2785, 1475, 2205]. Ontology-Based [2734, 765, 2784]. Ontology-Driven [2208, 2210, 1463]. Ontology-Mediated [1466]. OO [1517, 1444, 2209, 1313]. OPAC [1643]. Open [3331, 3337, 2666, 2974, 1179, 16, 1764, 560, 2602, 662, 2786, 1240]. Open-Ended [560]. OpenDLib [1251]. OpenGL [1875]. Opening [1156]. OpenMP [1488]. OpenSource [342]. Operating [3210, 1418, 1854, 3357, 868]. Operation [2359, 3044]. Operational [246, 1676]. Operations [120, 1520]. operative [1789, 1894, 1868, 861, 865, 1882, 1905]. Operator [2988, 854, 855, 3120, 3152, 858]. Operators [720, 845, 3192, 1944, 860]. Opinion [2033]. Opinions [3012]. Opportunities [270]. Opposites [851]. Optic [1041, 1042, 1867]. Optical [1817, 38, 39, 2872, 41, 1757, 2876, 37, 2079, 36]. Optimal [1629, 1101, 1094, 131, 2680, 965, 990, 492, 700, 2575, 2679, 504, 506, 2408, 2481, 1909, 725, 27, 3430]. Optimality [2654, 3143]. Optimally [3126]. Optimally-Smooth [3126]. Optimisation [2523, 902, 919, 897, 918, 165, 182]. Optimised [127]. Optimising [266]. Optimism [2774]. Optimistic [2178, 2757]. Optimization [3454, 642, 922, 869, 849, 3064, 2985, 1968, 250, 3072, 2558, 2425, 645, 1506, 863, 864, 59, 866, 842, 876, 1027, 2962, 883, 606, 855, 1600, 1536, 2710, 1059, 860, 1216, 2391, 900, 2980]. Optimizations [2005]. Optimized [737, 1533, 696, 1764]. Optimizing [2006, 2328, 3408, 1744]. Optimum [838]. Options [2800]. OptorSim [3218]. Or-Sets [1561]. Oracle [2156, 2173, 955]. Oracles [2697, 2683, 1684]. Orange [3423]. Orbit [1891]. Orbital [1889]. Orbitofrontal [3431]. Orchestrating [548]. Order [2545, 331, 3113, 3092, 556, 2158, 336, 940, 706, 2882, 2636, 2152, 520, 1102, 555, 3115, 287, 2125, 1745, 585, 3425, 1439]. Ordered [603, 2308, 1547, 2956, 3133, 602, 1603, 3173]. Ordering [1101, 1684, 2313]. Orderings [484]. Orderly [3076]. Ordinary [785]. ordination [3009]. Organ [1862, 2273]. Organisation [3324]. Organization [2938, 2396, 1109, 454, 215]. Organizational [2715]. Organizations [1451, 2506, 1346]. Organizing [160, 1109, 3466, 1546, 665, 849, 200, 880]. Organs [1893, 1823]. Orientation [3357, 1909, 2254, 176]. Oriented [1450, 630, 2780, 634, 628, 2215, 2225, 2244, 1237, 3348, 638, 3416, 651, 629, 3356, 3355, 1740, 748, 2788, 3010, 2486, 1531, 2429, 637, 644, 3309, 26, 790, 1172, 1937, 627, 3223, 655, 3345, 2903, 2520, 111, 3347, 3346, 3354, 1305, 3020, 812, 1354, 2973, 2784, 2771, 2457, 2827]. Origami [217]. Ortho [2923]. Orthognatic [1888]. Orthogonal [3078, 3052, 3046, 3062, 3310, 3059, 99, 734, 778, 2973]. Orthoses [1777]. Orthosis [1877]. Oscillation [2931]. OSGi [2582]. Osteotomy [1881]. Other [610, 2638]. Otoacoustic [2942]. OTS [2426]. Our [424]. Out-of-Order [336, 2636]. Outages [1681]. Outbound [2192]. Outdoor [2918]. Outer [3051]. Outer-Planar [3051]. Outlier [738]. Outlines [404]. Outlook [3319]. Outperform [1102]. Output [2649, 2381, 28, 3091]. Outs [544]. Overdefined [2167]. Overflow [2624]. Overhead [2745, 1994, 806]. Overlap [522]. Overlap-Free [522]. Overlapping [3166]. Overlay [2521, 679, 1858, 1753, 3253, 661]. Overload [185]. Overloading [310]. Overpassing [14]. Overriding [1112]. Overview [306, 269, 1422, 3238]. Ownership [1383, 1384].

P [2376, 2381, 2384, 197]. P2P [192, 2533, 2577, 1526, 194]. P2P-Based [194]. P3P [2194]. PAC [3144]. Pack [1968]. Package [1308, 2228]. Packages [432]. Packard [1269]. Packet [3331, 30, 48, 56, 31, 37, 2417, 2436, 2593, 2059, 65, 2202, 2517]. Packet-Aware [31]. Packets [63]. Packing [481, 3423, 2405]. Padding [962]. Paddings [961]. Page [1637]. PagePrompter [1608]. PageRank [1638]. Pages [387, 1959, 121]. Paging [2497, 97]. Pair [438, 2997, 435, 439, 464, 458]. Pair-Wise [2997]. Pairing [970, 831]. Pairing-Based [970]. Pairings [2183]. Pairs [1930, 2611, 474, 440, 2655]. Pairwise [983, 699]. Palpation [1780]. PAMR [3090]. Pan [1121, 2279]. Pan-European [1121]. Pan-Tilt [2279]. Panel [237]. Panorama [1872]. Panoramas [2279]. Panoramic [2920]. Paper [650, 3268, 205, 1395, 408, 3264]. PAPR [2500]. Paraconsistent [595, 593, 2857]. Paradigm [1299, 1687, 629, 653, 1217, 1423, 2457, 2962, 2303]. Parallel [848, 2030, 904, 1512, 2976, 1524, 1498, 1488, 850, 3227, 2733, 3061, 1506, 2679, 1778, 863, 2845, 1503, 2410, 3247, 1508, 1859, 876, 1484, 1517, 916, 878, 1226, 1516, 2386, 197, 1514, 1599, 1600, 3222, 1509, 2467, 1515, 1511, 3378, 1510, 1495, 1522, 1577, 3248, 1439]. Parallel-I [1522]. Parallel-I/O [1522]. Parallelism [1227, 2564]. Parameter [3416, 2675, 990, 3057, 3227, 861, 913, 914, 2673, 1065, 1033]. Parameter-Free [913, 914]. Parameterisation [2280]. Parameterized [2691, 505, 518, 1842]. Parameters [227, 2910, 3444]. Parametric [1053, 1434, 310, 589, 2631]. Parametrised [2648]. Paraphrasing [3397]. Parcellation [1802]. Parcelling [1807]. Parenthesis [2679]. Pareto [849, 3430]. Parity [2120, 2118, 2117]. Paroxysmal [2980]. PARS [2106]. Parsimonious [173]. Parsimony [877]. Parsing [3019, 904]. Part [2135, 2136, 2137, 2138, 2139, 2140, 2141, 2142, 2143, 2144, 2145, 2146, 2147, 2148, 2149, 2150, 541]. Partial [2735, 266, 1733, 3092, 207, 986, 242, 3433, 1334, 1368, 498, 2908, 1532, 589, 1835, 3115, 2239, 1439]. Partial-Domain [986]. Partially [3306, 37]. Participation [1145, 2426, 1343]. Participative [1353]. Particle [897, 1030]. Partition [998, 1260]. Partitioned [3083]. Partitioning [628, 2355, 2704, 1005]. Partitions [1650]. Partners [810, 2729]. Partnerships [1135]. Party [432]. Passengers [1670]. Passing [2350, 1530, 326]. Passive [2326]. Password [816, 972]. Password-Authenticated [972]. Pasta [191]. Patch [1001]. Patching [2626]. Path [2530, 3414, 250, 3050, 1013, 2276, 2476, 2026, 3073, 1945, 27, 2004, 3390, 1439, 2523, 95, 2467, 2524]. Path-Planner [2026]. Path-Vector [3390]. Path-Width [3050]. PathFinder [355]. Pathological [2937, 1796]. Pathologies [1874]. Pathology [3138]. Paths [3414, 2643, 2655, 843]. Pathway [2913]. Pathways [1082]. Patient [1779, 1920]. Patients [2932, 2929, 1809]. Pattern [2992, 638, 1640, 1199, 2015, 3441, 1642, 2479, 3428, 1054, 425, 1026, 3188, 3179, 1667, 2753, 2035, 722, 3141, 769, 1641, 232, 3140, 1016]. Pattern-Specific [1054]. Patterns [932, 740, 1261, 1197, 3278, 2487, 2076, 2912, 752, 2479, 3159, 174, 1374, 2397, 2847, 2090, 3193, 3440, 3155, 1542, 734, 778, 635, 1601, 2477, 2902]. Paying [1611]. Payload [2624]. PBL [1223]. PC [1854, 1924]. PC-Based [1854, 1924]. PCA [693, 2953]. PCS [89]. PCTL* [550]. PDA [372, 398]. PDAs [2823]. PDEs [1806]. Peak [1069]. Pebbles [491]. Pedestrian [2032]. Peer [199, 659, 688, 673, 2073, 198, 664, 670, 163, 671, 681, 3225, 680, 854, 190, 196, 1976, 201, 171, 660, 663, 200, 197, 666, 203, 678, 188, 683, 3339, 677, 2723, 3214]. Peer-to-Peer [199, 659, 688, 673, 198, 664, 670, 163, 671, 681, 3225, 680, 854, 190, 196, 1976, 201, 171, 660, 663, 200, 666, 203, 678, 188, 683, 3339, 677, 2723, 3214]. Peering [2073]. Peers [2722, 195]. Pen [380]. Penetration [1776]. People [2893, 439, 2283]. People-CMM [439]. Peptide [1068]. Peptides [1067]. Perception [2721, 2879]. Perceptions [445]. Perceptrons [902]. Perceptual [2740, 1024, 1721, 215, 216]. Percutaneous [1880, 1858]. Perfect [984]. Perform [3035]. Performance [1578, 3331, 2808, 2521, 3372, 1519, 2813, 1022, 1494, 3249, 76, 2768, 2748, 1478, 1372, 1215, 184, 2579, 1486, 2051, 66, 12, 70, 58, 186, 96, 46, 29, 2747, 1532, 3369, 93, 94, 2623, 28, 45, 292, 1538, 2100, 878, 55, 1078, 942, 1305, 2755, 806, 2467, 847, 2424, 1864, 1929, 2447, 47, 1479, 3389, 1793, 1537, 2499, 2994, 1059, 844, 101, 7, 4, 2335, 2760]. Performance-Oriented [1305]. Performances [1264]. Performed [1956, 1858]. Performing [2453]. Perfusion [1811, 1830, 1825]. Period [14]. Periodic [53]. Periodicities [741]. Periods [2926]. Permutation [869, 859]. Permutations [2157]. Persian [2450]. Persistence [2973]. Persistency [1217]. Persistent [1974, 3111, 2426]. Personal [1570, 387, 83, 2483, 370, 2481, 2497, 2374]. Personalisation [2532]. Personalised [3023]. Personalizable [2794]. Personalization [2955, 1377, 2066]. Personalized [10, 18, 3376, 1234]. Perspective [1130, 2314, 2041, 444, 186, 2229, 225, 1592, 730, 1373]. Perspectives [1626, 213, 1265]. Perturbed [2658]. Perturbing [2025]. Pervasive [297, 2580, 1296, 2766, 810, 3264]. Petascale [1481]. PETER [3023]. Petri [807, 1253, 362, 561, 3101, 2303]. Petri-Nets [2303]. Petroleum [1340]. PGI [1064]. PGV [968]. Phantom [1774, 1863]. Phase [2033, 1691, 459, 1017, 2409, 756, 2505, 3302, 724, 1899]. Phase-Conjugate [2409]. Phased [109]. Phases [716]. PhD [3346]. Phenomena [2038, 1544, 2032]. Phenomenon [1170]. Phenotype [851]. Pheromone [2971]. Philological [1281]. Philological/Textual [1281]. Phone [396]. Phonemes [767]. Phones [415, 367]. Photo [3279]. Photographic [1278]. Photometric [2277]. Phylogenetic [1082, 1090, 906, 1088]. Phylogeny [1089, 1102, 1071]. Physical [828, 2223, 3212, 2007]. Physically [1784]. Physics [1839]. Piconet [2496, 2482]. Piecewise [3167]. Piecewise-Linear [3167]. Pigmented [1839]. Piled [1009]. PIM [1391, 1392]. Pinna [2888]. Pinna-Brelstaff [2888]. Pipeline [1539, 1340]. Pipelines [2003, 336, 1744]. Pipelining [3241]. Pitfalls [2205]. Pixel [2401, 2910]. PKI [833, 786]. PLACE [370, 386, 432]. Placeholder [3253]. Placement [687, 2478, 1765, 2648, 179]. Plaintext [976]. Plan [1114, 308, 1147]. Planar [478, 3078, 2688, 3051, 1090, 3062, 2701, 474, 505, 1056, 2708, 3068, 229]. Planarity [3066]. Plane [502, 3069]. Planner [581, 2026]. Planning [3103, 1789, 1314, 2550, 1888, 582, 617, 1857, 308, 453, 2276, 2508, 3468, 1670, 2566, 2101, 3032, 3030, 3413, 576, 3033, 1824]. Plans [2504, 2566]. Plant [807, 3044, 2268]. Plaster [3156]. Platform [2740, 3335, 1512, 2914, 1152, 1163, 60, 59, 1982, 146, 2535, 2826, 1161, 1529, 648]. Platform-Based [1982]. Platforms [2008, 1376, 2801, 414]. Plausible [832]. Play [2650, 708]. Play-out [2650]. Playback [2599]. Playing [2041, 13]. PLSR [1825]. Plus [2436, 1116, 572]. PN [2259]. Pneumocephalus [1789]. Point [2130, 1655, 2131, 3061, 988, 543, 1664, 1027, 1915]. Point-Set [3061]. Pointer [3093]. Pointers [89]. Points [3414, 2423, 2170, 1919]. Polar [2871, 1829]. Polarimetric [1845]. Polder [1480]. Policies [2754, 2190, 177, 1408, 1409, 2411, 2302, 2065, 2496, 1635, 1143]. Policy [3035, 2337, 2047, 3465, 69, 113, 118, 116, 122, 3243, 3230, 2528, 3235, 2822]. Policy-Based [2047, 3465, 69, 113, 116]. Policy-Driven [2528]. Political [1113]. Polling [1523]. Polls [2178]. Pollutants [1720]. Polycategorical [728]. Polydioptric [1061]. Polygon [998, 1201]. Polygonal [341]. Polygons [2696]. Polyhedron [2676]. Polymatroid [483]. Polymorphic [944]. Polymorphism [310]. Polynomial [488, 536, 2667, 3130, 1742, 296, 2930, 3133, 576]. Polynomial-Length [576]. Polynomial-Time [536]. Polyominoes [3310]. Pomset [3033]. Pomset-Based [3033]. Pomsets [557, 553]. Pooled [1064]. Popularity [2512, 2053]. Popularity-Based [2053]. Populated [2442]. Population [2248, 1950, 870, 1857, 2981]. Populations [868, 2028, 2027, 1846, 878]. Portable [2747]. Portal [1166]. Portals [2955, 1527]. Porting [3201, 1528]. Pose [1045, 1019]. Posed [421]. Position [650, 3268, 1395, 3084, 3264]. Positioning [1751, 1750]. Positive [3133, 3011]. Possibilistic [3300, 705, 2855]. Possibilities [1492, 600, 2205]. Possibility [1552]. Possible [2891, 3128]. Post [3192]. Post-processing [3192]. Poster [3362]. Posteriori [1789]. Posture [1715]. Potential [1969, 1113, 1933]. Potentially [1562]. Potentials [1623]. Power [1324, 801, 2237, 2599, 1218, 114, 2384, 2483, 3044, 530, 3001, 2125, 1115]. POWER-Light [1115]. Power-Sum [114]. Powerful [2843]. Practical [835, 349, 310, 1683, 2359, 181, 833, 1856, 1267, 3253, 786, 1065, 2674, 3013, 1295]. Practice [1175, 1136, 441, 1132, 3384, 1164, 1153, 247]. Practices [1112, 452, 454, 445, 434]. Practitioner [2749]. Pre [1789, 1953]. Pre-operative [1789]. Pre-viewing [1953]. Precise [1287]. Precision [1549, 1611]. Preclusivity [1550]. Predicate [2629, 2360]. Predict [1612]. Predictable [3330]. Predicting [756]. Prediction [2418, 3458, 1995, 2987, 747, 2085, 887, 3419, 29, 911, 2277, 1079, 1621]. Predictions [3240, 773]. Predictive [3131, 727]. Preemption [3244]. preemptive [492]. Preference [2841, 1705, 3419, 1378, 1547]. Preference-Ordered [1547]. Preferences [572, 2256]. Preferred [3396, 602]. Prefetching [1204]. Prefix [1632, 3441, 344, 2126]. Prefix-Recognizable [344, 2126]. Prefixes [362]. Preliminary [739, 2550, 1781, 1448, 3405, 296, 2983, 930]. Premium [2514, 2525]. Preoperative [1908, 1909]. Preprocessor [3085]. Presburger [2638, 2637]. Presence [124]. Presentation [1253, 416, 383, 408]. Presentations [1250, 1241, 3297]. Presenting [2266]. Preservation [2831, 1118, 1267]. Preserving [1302, 1215, 2306, 1184, 2199, 1787, 3455]. Pressure [923, 368, 201, 877, 1115]. Preventing [1791]. Prevention [827]. Price [2538]. Prices [3219]. Pricing [2540, 2541, 2474, 2539, 2537, 2060]. Primal [1341]. Primal-Dual [1341]. Prime [974]. Primitives [2157, 2911, 1023]. Principal [746, 1000, 3450]. Principle [1089, 2977, 840]. Principles [1984, 1864]. Prior [2662, 1918]. priori [1028]. Priorities [3247]. Prioritized [351, 597]. Priority [1324, 2671, 2423, 2495]. Privacy [1570, 827, 748, 2194, 1177, 821]. Privacy-Oriented [748]. Private [960, 1135, 1112]. Proactive [816, 2524]. Probabilistic [2375, 306, 2188, 2033, 2011, 922, 1650, 551, 1930, 2264, 1000, 2975, 574, 106, 1436, 2698, 706, 2019, 3311, 1698, 1018, 859, 774, 330, 1788, 1445]. Probabilities [701]. Probability [2375, 1586, 2657, 1588, 40, 713, 2563, 2436, 3389]. Probe [430, 2370]. Probes [1100, 1099]. Problem [2258, 2694, 1685, 2830, 2269, 922, 1326, 2680, 3379, 3423, 2839, 2661, 845, 2986, 1222, 1709, 1334, 3055, 2364, 1501, 3383, 854, 863, 2698, 476, 2669, 2682, 1337, 509, 842, 975, 513, 913, 2981, 2995, 916, 2042, 1065, 812, 2405, 921, 3012, 966, 596]. Problem-Based [1222]. Problem-Oriented [812]. Problems [1322, 3103, 2691, 2291, 483, 2666, 2452, 1418, 3289, 1137, 2985, 1328, 619, 1213, 610, 3439, 2693, 78, 1628, 1506, 2690, 2935, 1336, 1106, 3445, 1339, 914, 518, 2687, 3467, 2345, 883, 1514, 2502, 1684, 857, 1321, 2260, 843, 421, 3446, 1341, 27, 3447, 157, 2664, 2395, 1216, 2391]. Procedural [2560]. Procedure [1770, 3238, 2936, 793]. Procedures [1880, 2571, 1669]. Process [1454, 1110, 1450, 2741, 939, 3412, 2040, 788, 1125, 791, 2486, 1351, 1374, 3028, 2505, 132, 170, 1484, 3186, 2240, 443, 555, 419, 799, 2304, 2101, 3197, 613, 2739, 439, 2477, 2098, 3442, 182, 2303, 613]. Process-Based [182]. Process-Oriented [1450, 2486]. Processes [426, 2008, 2207, 2336, 557, 3092, 3411, 462, 2250, 1458, 3107, 1043, 1121, 1375, 1124, 467, 2356, 723, 2427, 2726, 564, 1180, 1373, 2273, 2729]. Processing [1014, 1060, 76, 2843, 2733, 2845, 1976, 25, 140, 208, 1509, 1565, 1231, 3020, 1969, 3239, 1641, 2868, 3435, 3192]. Processor [3252, 2003, 806]. Processors [2006, 1990, 1481, 2002, 2059, 143]. Product [869, 789, 2832]. Production [1205, 3249, 2986, 1998, 2508, 1354].

Products [974, 1376, 1555, 1340, 2281]. Profile [1306, 1971, 1300, 1931, 1869, 1557, 1606, 3174, 183]. Profile-Based [1971]. Profiles [1082, 1238, 915, 1969, 1234]. Profiling [1976, 1099]. Profit [163, 2674]. Program [1737, 349, 2001, 303, 309, 610, 1729, 1742, 2306, 246, 1748]. Programmable [3334, 2056, 2059, 2192, 3339]. Programming [2545, 1731, 438, 1734, 273, 275, 3348, 870, 1488, 297, 488, 905, 567, 631, 3344, 1738, 3355, 2804, 851, 884, 625, 198, 244, 621, 873, 308, 3202, 463, 3365, 1005, 1984, 1741, 298, 941, 3421, 420, 301, 443, 2484, 1516, 606, 276, 287, 945, 436, 435, 601, 1745, 2007, 1675, 2853, 439, 464, 1730, 899, 1567, 3398, 458, 427, 2457, 1747]. Programming/Agile [436]. Programs [595, 571, 612, 295, 482, 291, 603, 1995, 288, 3361, 3106, 2487, 327, 2560, 1985, 607, 1530, 512, 937, 1484, 300, 3404, 933, 2962, 304, 145, 296, 604, 808, 3378, 602]. Progress [617, 1240]. Project [426, 459, 1136, 2047, 2685, 422, 1245, 1123, 1277, 1282, 1259, 1161, 2725, 3121, 3153, 1165, 1152, 1129, 1480]. Projection [3162, 1869, 1907, 1841]. Projections [3310]. Projective [1034]. projector [1859]. Projects [426, 1314, 462, 2486]. Prolog [304, 302, 2573, 2744, 312]. PROLOG/RDBMS [2744]. PromethOS [3323]. Promise [1142]. Promoting [636]. Prone [1920]. Proof [1399, 289, 2550, 568, 986, 748, 474, 2662, 520, 1400, 287, 528, 2566, 2568, 954]. Proof-System [568]. Proof-Theoretic [287]. Proofs [2163, 958, 2552, 358, 955, 2571]. Propagation [1879, 2461, 2558, 2490, 2880, 2716, 279, 803, 721, 1928, 1820]. Properties [2937, 480, 2958, 1784, 1587, 250, 2003, 2555, 1775, 344, 2946, 2884, 1562, 666, 1020, 3312, 1427, 596, 1469]. Property [325, 2238, 1827]. Property-Based [2238]. Proportional [56]. Proposal [1455, 1126, 78, 1588, 18]. Proposed [1546]. Propositional [2639, 591, 525]. Propositionalization [3194]. Prospects [3270]. Prostate [1897, 1761, 1782, 1905]. Protection [117, 827, 163, 2194, 71, 824, 2489, 821, 185]. Protein [3083, 1068, 3080, 2399, 911, 1075, 1067, 1081, 1080, 3436]. Protein-Protein [3080, 2399]. Protocol [105, 142, 3103, 340, 1734, 794, 1997, 2196, 957, 323, 3251, 1178, 2448, 1694, 104, 15, 110, 2065, 87, 103, 139, 2539, 830, 2053, 3288, 47, 3390, 3095, 1366, 2720, 32]. Protocols [1385, 1386, 558, 335, 2645, 337, 76, 2680, 2773, 533, 2713, 2347, 3258, 560, 1449, 2345, 3292, 86, 1272, 1406, 1407, 126]. Prototype [1279, 1122, 3408, 2406, 3287]. Prototypes [2874, 2267]. Prototyping [1252]. Provability [829]. Provable [2174]. Provably [953, 949]. Proven [166]. prover [2662]. Provider [397]. Provider-Specified [397]. Providing [2542, 2531, 2739]. Proving [2163, 363, 349, 288, 2553, 336, 2562]. Provision [818, 92, 2460, 1158]. Provisioning [2541, 2511, 2528, 1182]. Provisions [2458]. Proximity [1650, 1636, 2470]. Proxy [2763, 122, 2055]. Proxy-Caching [2055]. PRS [3409]. PRS-Like [3409]. Prudent [1395, 1396]. Prüfer [2425]. Pruning [3462]. Pseudonymizing [825]. Pseudorandom [976, 3448, 973]. Pseudospectra [1512]. Psychology [3283]. Public [1110, 1130, 1135, 1120, 1140, 2410, 1138, 1145, 822, 2786, 1158, 1166, 1170]. Public-Key [2410]. Public-Private [1135]. Publish [3281]. Publishing [1972]. Pulmonary [1843, 1791, 1847]. Pulse [99, 2940]. Purely [289]. Purpose [2386, 614]. Purposes [3097]. Pursuit [1832, 1542]. Pushdown [2128, 508, 2561, 564]. Pushing [387]. Putamen [1814]. Puzzles [1393, 1394, 610].

PVM

[1518, 1528, 1507, 1505, 1532, 1508, 1523]. PVM-Based [1508]. PWB [1347]. PWD [147]. Pyramid [1002]. Python [3335]. Python-Based [3335].

Q [1210, 715, 2971, 721]. Q&A [3456]. Q-Cut [715]. Q-Learning [1210, 2971]. Q-values [721]. QL [265]. QoS [1306, 2064, 2819, 2045, 2818, 3219, 2587, 2056, 2466, 2511, 2538, 1743, 2583, 2601, 81, 2534, 1681, 92, 2460, 2324, 2531, 1970, 3390, 3391, 75]. QoS-differentiated [2534]. QoS-Driven [1970]. Quad [1712]. Quad-Tree [1712]. Quadratic [936]. Qualitative [580, 1255, 590, 2249, 2253, 1673, 3157, 2883, 2256]. Qualitatively [2252]. Quality [2576, 1637, 992, 3463, 2242, 2328, 2047, 2737, 2240, 2817, 2241, 1923, 2331, 121, 2528, 1786, 2517, 1892, 1278, 1115]. Quantification [1789]. Quantified [2639, 593, 934]. Quantifier [585]. Quantitative [580, 1550, 3399, 1927, 2563, 1750, 3354, 1850, 689]. Quantization [2424]. Quantum [2375, 479, 1379, 2383, 2662, 3311, 2386, 2372, 2020, 2391, 2392]. Quasi [78, 2968, 1523]. Quasi- [1523]. Quasi-Metric [2968]. Quasi-Solution [78]. QUBOS [2639]. Queaps [2670]. Queries [575, 2423, 1367, 2838, 905, 2839, 250, 2974, 260, 3132, 681, 1956, 190, 644, 1187, 3440, 1698, 2854, 1645, 7, 4, 647, 2832]. Query [2829, 1950, 2735, 642, 2190, 2835, 286, 251, 1248, 264, 1968, 2841, 611, 2558, 2846, 1976, 2847, 707, 2445, 1944, 1945, 2710, 1565, 3329, 1194, 1969, 3239, 1256, 2856, 253, 2857, 1059, 85, 647, 3397]. Query-Dependent [1059]. Querying [2734, 2717, 2836, 2837, 2842, 2731, 2442, 2848, 2849, 2398, 1364, 1943, 2736, 258, 2852, 265, 2961, 1949, 1627, 2224, 2855]. Quest [315]. Questions [1149, 662]. Queue [2671, 2061, 3247]. Queued [31, 28]. Queues [1437]. Queuing [2465].

QuickTime [1872]. QuickTime-VR [1872]. Quo [1169].

R [2221, 1345, 24, 2353]. Racing [3164]. RADAR [3326]. Radial [2993]. Radical [1288]. Radio [2461, 2579, 2698, 65, 3101, 50]. Radiocoloring [478]. Radiotherapy [3095]. Random [2989, 838, 481, 869, 2156, 2706, 490, 2071, 495, 2173, 757, 985, 1898, 1710, 519, 967, 955, 856]. Randomized [3102]. Range [896, 1717, 3042, 3, 6]. Ranges [3191]. Ranking [2689, 2957, 727]. RANVEC [236]. Rapid [383, 408, 1095, 1310, 2859]. Rare [756]. Raster [226, 219]. Raster-Scanned [219]. Raster-to-Vector [226]. Rate [2926, 2588, 11, 2515]. Rate-Based [11]. Rates [850]. Ratio [3075]. Rational [511]. Rationale [3268]. Rationality [3409]. Ray [1907, 1820, 1906, 1029]. RBAC [2419]. RBF [1886, 886, 3142, 3426, 3436]. RBF-Based [1886]. RC4 [967]. RCPS [2269]. RDBMS [2744]. RDF [1474, 2961, 1965]. RDF-Described [1965]. Re [2476, 2923]. Re-computation [2476]. Re-visiting [2923]. Reach [3430]. Reaching [1535]. Reaction [2037]. Reactions [552]. Reactive [2649, 3367, 2804, 279, 307]. Read [482]. Read-Once-Branching [482]. Readability [383]. Reader [3274]. Reading [2865]. Ready [1533, 783]. Ready-Mode [1533]. Real [2365, 448, 903, 2581, 2717, 339, 989, 1418, 2367, 3027, 2767, 1050, 3458, 1347, 1700, 2768, 1440, 1419, 625, 243, 1761, 1300, 653, 3270, 997, 2988, 852, 2479, 1758, 894, 2271, 1435, 1984, 2747, 124, 1001, 1890, 2623, 2594, 914, 1992, 1743, 1021, 1866, 1876, 54, 1721, 276, 3034, 1994, 1701, 2535, 167, 1999, 2897, 155, 3304, 921, 1428, 1932, 2593]. Real-Coded [2988, 852, 914]. Real-Time [2365, 903, 2581, 2717, 339, 989, 1418, 3027, 2767, 1050, 3458, 1347, 1700, 2768, 1440, 1419, 1761, 1300, 653, 3270, 2479, 1758, 2271, 1984, 2747, 124, 1890, 2623, 2594, 1992, 1743, 1866, 54, 1721, 276, 1994, 1701, 2535, 167, 2593, 1999, 155, 3304, 1428, 1932]. Real-World [1001]. Realistic [3337, 2886, 1449, 3446]. Reality [1854, 1858, 1867, 384, 2324, 1863]. Realizability [539]. Rearrangement [1071]. Reasoning [2247, 2955, 586, 3104, 612, 575, 593, 291, 256, 2791, 1457, 2377, 590, 2226, 608, 3456, 1558, 3400, 2272, 2777, 2798, 1542, 733, 777, 597, 2572, 274, 2255, 1428, 2257, 584]. Reasons [2251]. Reassignment [102]. Rebuilding [1110]. Receive [1533]. Receiver [2512]. Receiver-Driven [2512]. Receivers [2177]. Receptive [2863]. Recognition [229, 212, 1048, 3037, 1000, 3177, 1199, 2893, 991, 224, 1045, 2275, 2479, 1047, 3428, 1055, 3449, 219, 1667, 213, 2896, 2035, 1618, 1037, 722, 215, 2922, 2897, 2883, 232, 2923, 2905, 1051]. Recognizable [344, 2126, 523]. Recognize [214]. Recognizer [211]. Recognizing [1797, 2910, 3043]. Recombinants [1083]. Recommendation [3419, 1378, 710, 2957]. Recommendations [2194]. Reconciliation [2238]. Reconciling [2759]. Reconfigurable [1211, 2353]. Reconfiguration [3324]. Reconfiguring [1113]. Reconstructed [992]. Reconstructing [3144]. Reconstruction [1799, 319, 1032, 2274, 1089, 205, 3310, 1868, 2659, 2908, 2889, 3174, 1102, 994, 1018, 3084, 1034, 1824, 2870]. Record [2264]. Recording [1718]. Recordings [2939]. Recovering [2422]. Recovery [128, 1717, 1893, 16, 1009, 201, 1516, 2513, 786, 2414]. Rectangular [3069, 3068, 232]. Rectification [229, 223]. Recurrent [3454, 884, 3000, 1067, 1712, 1722]. Recursion [944]. Recursive [1655, 2367, 3134, 2494, 2832]. RED [43, 51]. Redefining [3360]. Redistribution [2522].

Reduce [2465]. Reduced [2165, 2572]. Reduced-Round [2165]. Reducing [135, 542, 2065, 2745, 1597, 856]. Reduction [3103, 321, 319, 363, 3094, 1058, 720, 3072, 309, 3071, 2415, 3408, 2438, 1889, 3115, 2500, 1288, 3031, 2965]. Reductions [2571]. Reducts [3183, 1565, 1571]. Redundancy [2963, 53]. Redundant [2258, 796, 787, 846]. Reengineering [2305]. Refactoring [1736, 449]. Reference [1470, 1964, 2485, 2625]. Reference-Aware [1964]. References [9, 2950]. Referential [1470]. Refinement [323, 2630, 333, 527, 1072, 2199, 562]. Refinement-Robust [562]. Reflection [1731]. Reflections [1123]. Reflective [3134, 2762, 1999, 144]. Reflexive [640]. Refutation [3380]. Regard [2576]. Region [2705, 123, 280, 2278]. Region-Boundary [2278]. Regional [1843, 2043]. Regions [1927, 2278, 998]. Registering [2227]. Registers [2357, 1036, 2349]. registrar [1165]. Registration [1830, 1914, 1916, 1011, 1862, 908, 1919, 1798, 1869, 1922, 1911, 1908, 999, 1828, 1910, 1913, 1027, 1920, 1909, 1918, 1829, 1866, 97, 1819, 1838, 1907, 2282, 1915, 1917, 1921, 1787, 1906, 1821, 1817]. Registry [2612]. Regression [698, 3439, 714, 719, 881]. Regular [535, 360, 481, 250, 553, 509, 2564, 2025, 3140, 2276]. Regularities [2890]. Regularization [1044]. Regularized [1803, 1921, 3455]. Regulating [2730]. Regulation [2338, 3157, 1171]. Regulatory [2455, 3173]. Rehearsal [1782]. Reification [2776]. Reinforcement [3007, 3148, 1026, 715, 3431]. Relabelling [2294, 2296]. Relate [441]. Related [483, 619, 492, 1931]. Relating [2635]. Relation [2346, 2400, 600]. Relational [1304, 1228, 264, 2431, 503, 2432, 759, 1946, 1364, 1947, 1596, 254, 1561, 1948, 1184, 1949, 2853, 1592, 760]. Relations [2163, 1705, 703, 1555, 134, 2432, 1232, 1579, 1563, 2973, 1012, 1860]. Relationship [2471, 1462, 2154, 1948, 265, 3120, 3152, 65, 2217]. Relationships [1294, 3176, 2230, 3445, 2930]. Relative [1664, 2648, 1265, 5]. Relatives [1489]. Relativistic [2390]. Relaxed [1956, 3248]. Relays [2604]. Release [453]. Releasing [1570]. Relevance [1254]. Relevant [3139, 1063, 1075, 2097]. Reliability [2759, 789, 795, 2255, 2664]. Reliable [22, 2680, 688, 2045, 2351, 17, 709, 3099, 86, 2537, 3101, 195]. Religion [858]. Religion-Based [858]. Relocation [179]. Remaining [1149]. Remarks [2077, 2663]. Remeshing [1517]. Remote [995, 2057, 1761, 2334, 1967, 1771, 1770, 1531, 124, 2753, 3241, 3238, 405]. Remotely [61]. Removal [208]. Renaming [1308]. Renardus [1245]. Rendering [1508, 2438, 2946, 1924]. Rendezvous [2657]. Rendezvousing [368]. Rental [2693]. Repeat [2364]. Repeated [2266]. Repeats [496]. Replacement [2754, 1781, 2300, 122]. Replacing [2743]. Replanning [1202]. Replay [3378]. Replica [687, 2600]. Replicas [142]. Replicated [3206, 177, 2601, 2443, 167]. Replicating [2021]. Replication [1085, 2332, 3218, 2759, 689]. Replication-Directed [1085]. Replications [2851]. Replicators [2022]. Report [3086, 2550, 2099, 659, 1683, 1738, 617, 3256, 1242, 296, 2983, 1240, 3121, 3153]. Reports [1153]. Repositories [1474]. Repository [1269, 149]. Representation [2262, 2401, 1050, 2837, 1186, 1457, 2963, 845, 1289, 2250, 3399, 1446, 2442, 2742, 1886, 3451, 2229, 2631, 1206, 2608, 1467, 859]. Representational [862]. Representations [1437, 755, 2890, 3400, 847, 345, 2883]. Representatives [3056]. Representing [269, 3128, 2961, 600, 2224]. Reproducing [2380]. Reprojection [1034]. repudiation [1399, 1400]. Reputation [202].

Reputation-Aware [202]. Request [1267, 1134]. Requested [3249]. Requests [142]. Required [380]. Requirement [410, 1220, 2049]. Requirements [1237, 1440, 3219, 1376, 1507, 1981, 2286, 2650, 1243, 1123, 1088, 419, 271, 1158, 1258]. Resampling [1013]. Rescue [176]. Research [739, 2107, 2194, 3256, 1626, 1257, 2322, 3375, 427, 1108]. Reservation [2539, 3246]. Reservations [3247, 1338]. Reservoir [879]. Reset [1692]. Reset-Driven [1692]. Residual [797]. Residue [3139]. Residues [2424]. Resilient [2180, 2179, 979, 200]. Resistant [1385, 1386, 669]. Resolution [672, 2549, 3402, 1739, 504, 2347, 1859, 3014, 406, 3380, 1933, 2600]. Resonance [1825, 1912, 1809, 1933]. Resource [3237, 3303, 2081, 688, 3343, 3251, 2738, 3219, 3227, 2724, 874, 3225, 2533, 2541, 137, 3231, 3232, 2539, 3226, 2053, 3467, 1992, 2198, 3254, 2766, 3292, 2595, 3388, 3339, 50]. Resource-Based [874]. Resource-Dependencies [3292]. Resource-Limited [2081]. Resources [2796, 413, 545, 1426, 2736, 1194, 546, 3182]. Respiratory [1928]. Respiring [1774]. Response [1139, 44, 2478, 2330, 113, 49]. Responses [2887]. Restart [841]. Restoration [71, 2983, 3455]. Restricted [2363, 2377, 1335, 2965]. Restricting [1092]. Restriction [1092]. Restrictions [2557, 1566]. Result [3141]. Results [2681, 1136, 495, 1781, 3293, 1262, 893, 1129]. Retaining [3457]. Retargetable [2006]. Retina [2947]. Retinal [1845]. Retino [2872]. Retino-Cortical [2872]. Retransmission [22, 86]. Retrieval [1977, 1443, 1192, 1255, 1261, 1659, 1367, 1218, 192, 252, 3299, 1263, 1646, 1666, 2753, 2408, 2825, 1264, 3018, 1952, 3185, 2653, 1194, 136, 152, 1954, 409, 1249, 1965, 1645, 1953, 1059, 1957, 3298, 3397]. Retrieving [2591, 1676]. Reusable [1212]. Reuse [634, 640, 636, 635, 3353]. Reusing [2800]. Reversal [1085, 2661]. Reversal-Bounded [2661]. Reversals [1085]. Reverse [2430]. Reversible [2385]. Review [1726]. Revising [692, 3157]. Revision [609]. Revisited [2714, 1393, 1394, 2683, 1106]. Revisiting [949]. Revocation [2177, 952]. Revolutionize [3185]. Reward [3418]. Rewrite [940]. Rewriting [294, 2292, 1427]. REYES [1508]. Rhine [2039]. RichProlog [285]. Ridgeline [1862]. Right [2938, 2865, 1826, 1173]. Rights [2603]. Rigid [1076, 1916, 1919, 1828, 1918, 1819, 1838, 1917, 1787]. Rigor [1112]. Rigorous [784]. Rijndael [2160, 2161]. Rijndael-Like [2161]. Ring [977, 3326, 978, 2183]. Ring-Based [3326]. Rings [2344, 3070]. RIONA [700]. Risk [1168, 791, 820]. River [216]. RLC [96]. Road [2476, 216, 2525]. Roadmapping [1699]. ROADRUNNER [1363]. Roaming [2068, 833, 87]. Robin [150]. RoboCup [2407]. ROBODOC [1791]. Robot [3414, 2913, 2270, 3040, 894, 1766, 1753, 1620, 2916, 1767, 3466]. Robotic [1783, 1759, 1882, 3021, 1750, 1202, 3431, 1770]. Robotics [2914, 893]. Robotized [1751]. Robots [2026, 1037, 2917]. Robust [2022, 2361, 919, 1683, 2172, 2045, 908, 2860, 1904, 1615, 3438, 1856, 1027, 208, 717, 2154, 1866, 2596, 1820, 562]. Robustness [710, 1479]. Role [2262, 638, 2017, 629, 2880, 1462, 116, 1176, 2922, 2897, 1113]. Role-Based [116]. Roles [2557]. Roof [1018]. Room [2368]. Root [2644]. ROSA [3337]. Rotation [227, 2872, 1883]. Rotation-Invariant [2872]. Rough [1578, 1593, 1610, 1549, 1611, 1569, 1551, 1550, 1617, 1572, 1552, 1553, 1613, 1614, 1588, 1554, 1556, 1544, 1616, 1575, 1558, 1559, 1598, 1542, 1580, 1620, 1581, 1562, 1621, 1564, 1622, 1582, 1547, 1623, 1624, 1590, 771, 1591, 1548, 1567, 1571]. Rough-Set-Theoretic [1562]. Round [150, 959, 2165, 42]. Round-Robin [150].

Round-Trip [42]. Rounding [1328]. Routed [2340, 36]. Router [166, 3323, 17]. Routers [29, 3326, 3341, 3388, 3391]. Routing [2523, 2291, 38, 2358, 2812, 41, 190, 863, 104, 3323, 57, 103, 20, 1338, 3226, 37, 2070, 662, 133, 188, 35, 2524, 2437, 36, 27, 3390, 921, 32, 661]. Routing-Based [2070]. Routing-Transferring [3226]. RSA [2175, 2169, 962, 956, 963]. RSVP [92]. RT [1437]. RTIPA [2047]. RTOS [1989]. RTP [2057]. RU [493]. Rubber [2186]. Rule [2984, 2262, 2081, 2987, 2840, 2711, 884, 3138, 1594, 700, 1554, 755, 756, 2490, 169, 1616, 764, 258, 2993, 2309, 3142, 1635, 1590]. Rule-Based [2984, 2987, 169, 258, 1635]. Rules [266, 311, 745, 2770, 3300, 1356, 2087, 1553, 1574, 3167, 753, 3192, 1395, 1396, 2302, 2400, 3147, 1597, 1564, 770, 2214, 1576, 771, 1584]. Ruminations [2351]. Run [3077]. Running [842, 3378]. Runs [1298]. Runtime [3324, 1496, 2760]. Runtimes [3249]. RV [1842].

S [1100, 1474, 2961, 2179, 1473, 1348]. S-Boxes [2179]. S-CREAM [1473]. S1S [2123]. S2S [2123]. Saarland [1116]. Safe [974, 1570, 568, 3343, 1980, 539, 3044, 1523, 802]. Safe-Prime [974]. Safety [1680, 3331, 1682, 801, 789, 1985, 358, 1772, 2201, 790, 1417, 809, 2309, 799, 795, 3312, 3093, 1427]. Safety-Active [1772]. Safety-Analysis [3093]. Safety-Critical [2309, 795]. Salesman [922, 1337]. Saliency [1021, 2900]. Saliency-Based [1021]. Sampling [2424, 769, 330]. Sanctity [1112]. Sat [495, 2258, 3103, 2692, 331, 2630, 333, 3445, 332, 3468, 3446, 329, 2637, 3444, 3447]. SAT-Compilation [3103]. Satellite [896, 106, 109]. Satisfaction [479, 3433, 3467, 2260, 2374, 840]. Satisfiability [2639, 331, 591, 1430, 2572, 592, 315]. Satisfy [647]. Satisfying [2773]. Saving [2483]. Say [417]. Says [417]. SBH [1092]. SBQ [2518]. Scalability [3206, 2586, 1479]. Scalable [2354, 3374, 1990, 2589, 687, 3385, 2578, 3225, 186, 1187, 175, 1484, 667, 2512, 1598, 2535, 769, 3341, 2054, 403, 685]. Scale [227, 1688, 72, 2369, 2513, 1698, 1095, 666, 2960, 2389, 668, 1892, 1833, 1669]. SCALEA [1537]. Scaled [746]. Scaling [462, 703]. Scan [222]. Scan-to-XML [222]. Scanned [219]. Scanner [1761, 1858]. Scans [1835]. Scattered [484]. SCC [2633]. Scenario [384]. Scenarios [2068, 1122, 3332]. Scene [2884, 2859]. Scene-Centered [2884]. Scenes [2898, 3042, 1031]. SCG [115]. Schedulability [1988]. Scheduler [2518, 56, 1423, 3243]. Schedulers [3252]. Scheduling [1323, 2686, 896, 3303, 2291, 3027, 2767, 44, 3249, 1347, 1218, 150, 1989, 919, 2738, 3219, 492, 3233, 3250, 2986, 2769, 1507, 2685, 2575, 30, 2669, 3245, 3247, 3232, 31, 1425, 513, 914, 2687, 2496, 3254, 3253, 2593, 898, 3413, 34, 2059, 1511, 3246, 3304, 858, 1219, 3031, 157, 3248, 143, 3435, 3234]. Schema [262, 3040, 2235, 2431, 1939, 2432, 2444, 304, 1365]. Schema-Based [304]. Schemas [575, 2244, 2242, 2243, 1937, 1370]. Schemata [2217]. Schematics [2221]. Scheme [2177, 147, 2529, 2468, 3416, 1964, 73, 2158, 2358, 951, 135, 2860, 83, 1220, 2409, 64, 2466, 2495, 2667, 138, 1940, 2415, 2462, 11, 2171, 89, 3405, 711, 120, 37, 2408, 2463, 19, 2447, 2062, 2454, 2016, 63]. Schemes [958, 986, 962, 695, 984, 2174, 3408, 2801, 2182, 2154, 613, 677, 954, 2216]. Schizophrenia [739, 1896]. Schnorr [958]. Scholarly [1281]. School [3434]. SCI [1502]. Science [1271, 1626]. Scientific [1519, 2227, 680, 3295, 3197, 1965, 946, 164]. Sclerosis [1809]. Scores [1264, 2929]. Scoring [1081]. Scottsdale [3256]. Screen [1273, 373, 210]. Scribble [211]. Scripting [1732, 180]. Scrum [455]. SCSPs [3433]. SDSI [672]. Seamless [1377, 1159, 98, 81, 2826]. Search [1955, 896, 1094, 1192, 3094, 289, 720, 192, 695, 293, 1274, 1612, 2479, 644, 758, 2794, 760, 846, 2899, 2732, 2585, 1262, 1320, 3466, 724, 136, 2405, 2868, 3447, 3031, 3269, 3121, 3153, 3, 5, 900, 2967]. Search-Space [720]. Searched [840]. Searches [3227, 1087, 6, 647]. Searching [2549, 1650, 373, 1651, 3018, 1672, 1627, 1676, 1609]. Second [2882, 298, 2125, 585]. Second-Order [2125, 585]. Secret [974, 965, 983, 2171]. Secrets [1416]. Sectional [1810]. Sector [1110]. Secure [1385, 3308, 2780, 3210, 1386, 2181, 1972, 2172, 1178, 2337, 826, 959, 2343, 2155, 3270, 1179, 953, 278, 949, 1316, 3448, 2462, 833, 87, 3400, 1138, 3229, 2182, 98, 2816, 2482, 3393, 2192, 2164, 2335]. SecureUML [1317]. Security [340, 814, 1389, 3337, 958, 2756, 1688, 558, 957, 337, 2603, 1413, 986, 1700, 2758, 788, 976, 1412, 787, 2047, 1391, 1392, 2174, 1162, 2411, 956, 949, 1395, 1396, 112, 69, 118, 790, 116, 975, 1317, 2408, 830, 1390, 1485, 805, 2198, 2075, 1381, 823, 2161, 111, 786, 3001, 1382, 1173, 821, 683, 2200]. Seed [1801]. Seeing [2893]. Seen [706]. Segment [1072, 1819]. Segmentation [1717, 1044, 2999, 1796, 235, 1036, 1011, 1797, 236, 1614, 1765, 1804, 1505, 2584, 2412, 1816, 1823, 1833, 1828, 3043, 1886, 3042, 1826, 1795, 2278, 2869, 1834, 1020, 1820, 2868, 1818, 1850, 1786, 1899, 234, 1821, 1788, 1853]. Segmenting [1039, 1852]. Segments [1074, 1772, 510]. Segregation [2888]. Seismic [1006]. Seismological [1718]. Selected [2953]. Selecting [3462, 766]. Selection [2530, 3035, 923, 2263, 3402, 990, 851, 129, 1491, 749, 3386, 700, 2951, 2988, 3161, 3437, 100, 3194, 1204, 2092, 1616, 2265, 1095, 3023, 2972, 821, 27, 2994, 2905]. Selections [2221]. Selective [1995, 3246]. Selectivity [2901]. Selectorecombinative [847]. Self [848, 160, 1731, 849, 746, 1372, 2875, 1171, 2724, 2380, 2348, 1976, 200, 3073, 3466, 880, 856, 1546, 2021, 3243, 1049, 858, 665]. Self-adaptation [856]. Self-adaptive [848, 858]. Self-aggregation [746]. Self-calibration [1049]. Self-Interested [1372]. Self-intersecting [3073]. Self-Motion [2875]. Self-Organizing [160, 3466, 1546, 665, 849, 200, 880]. Self-Profiling [1976]. Self-regulation [1171]. Self-Replicating [2021]. Self-Reproducing [2380]. Self-Stabilizing [2348]. Self-Structuring [2724]. Self-Tuning [3243]. Semantic [3296, 1455, 2782, 642, 3462, 1371, 2233, 3402, 3299, 3022, 2227, 508, 219, 1241, 2095, 2238, 1018, 2214, 2797, 1256, 1475, 1470, 2796, 270, 1472, 574, 1476, 2792, 765, 2456]. Semantic-Based [3299]. Semantical [2304, 860]. Semantically [190]. Semantics [1442, 2782, 3462, 294, 583, 2560, 278, 607, 561, 1662, 512, 937, 515, 2778, 246, 1355, 2566, 604, 1965, 1288, 1746, 2312, 1287, 3108, 2217, 525, 2776]. Semantics-Based [1355]. Semantics-Driven [2778]. Semi [737, 3078, 334, 492, 259, 1473, 643, 80, 1641, 2097, 1475]. Semi-algebraic [259]. Semi-automatic [1473, 2097, 1475]. Semi-dynamic [3078]. Semi-formal [334]. Semi-online [492]. Semi-Soft [80]. Semi-structured [737, 643, 1641]. Semiconductor [2409]. Semidefinite [488, 1005]. Semiqualitative [2979]. Semiquantitative [2979]. Semirandom [488]. Semiring [3433]. Semistructured [2735, 286, 251, 249, 2856, 1365, 2224]. Sender [2520, 2069]. Sense [3015, 2786]. Senses [3178]. Sensing [995]. Sensitive [2829, 744, 756, 1257, 713, 803, 3160]. Sensitivity [1807]. Sensor [3211, 1038, 2920, 2811, 1011]. Sentence [2083, 2094]. Separability [775]. Separating [3060, 955]. Separation [2931, 218, 416, 985, 2991, 329]. Septic [2929]. Sequence [2548, 1014, 1068, 1074, 1869, 3448, 2562, 510, 539, 1088, 2370, 3033].

Sequence-Length [1088]. Sequencers [2360]. Sequences [1085, 1799, 1052, 747, 2912, 2911, 1800, 2397, 3175, 2890, 2371, 767, 3155, 107, 1079, 1067, 994, 1066, 2902, 1083, 3436, 3122, 3154]. Sequencing [1091]. Sequencing-by-Hybridization [1091]. Sequent [2567]. Sequential [3107, 2493, 2646]. Sequentiality [289]. Sequents [241]. Serial [1843, 834, 917, 878, 383, 408, 2266, 2467]. Serial/Parallel [2467]. Serializable [263]. Series [2937, 2258, 741, 2976, 2338, 1931, 3061, 887, 752, 3190, 2847, 1912, 2998, 1720, 2253, 886, 1577]. Series-Parallel [2976, 3061]. Serv [2060]. Servent [202]. Server [2576, 1964, 72, 129, 3386, 1217, 108, 61, 2470, 179, 2419, 2054, 3263]. Server-Based [129]. Servers [193, 1524, 3385, 493, 184, 1966, 2600, 3381, 185]. Service [2810, 1208, 2576, 1130, 2754, 2542, 2741, 2046, 573, 3336, 3327, 2336, 1251, 2767, 382, 673, 3251, 2325, 2328, 2804, 2731, 2047, 1961, 67, 2526, 1109, 3225, 1245, 2541, 12, 2415, 189, 1106, 87, 1127, 2498, 2333, 2353, 2817, 1383, 1384, 62, 2601, 2520, 91, 2066, 2331, 2514, 3341, 3338, 3410, 2062, 2515, 1975, 2527, 2528, 3235, 1182, 2434, 144, 399, 2757, 2048, 3216]. Service-Based [3216]. Service-Level [2325]. Service-Specific [3336]. Services [3206, 2751, 2717, 177, 2325, 2045, 1136, 166, 2587, 2047, 1163, 2752, 3266, 1146, 2056, 67, 2533, 1162, 1531, 2067, 3232, 3271, 1159, 836, 2481, 391, 2534, 2449, 2532, 1167, 3098, 2339, 2327, 2514, 1969, 176, 3393, 2749, 3328, 2323, 1158, 2063, 2477, 1676, 3100, 1148, 1113, 2824, 3260, 417, 3265]. Serving [673]. Session [2053, 3362, 2335, 2048]. Session-Aware [2053]. Sessions [2512, 2610]. Set [1085, 1578, 559, 1685, 1324, 2363, 1593, 1610, 1691, 1938, 603, 582, 1897, 1329, 621, 3061, 578, 1552, 1709, 1613, 1614, 505, 100, 1616, 1833, 1559, 2345, 606, 1542, 1580, 1562, 1636, 2673, 2214, 1622, 1547, 1590, 1591, 1083, 1899, 731, 1853]. Set-Based [559, 1636]. Sets [1549, 1611, 2688, 742, 1551, 1617, 3166, 293, 490, 553, 1554, 1191, 2381, 3192, 1544, 2753, 519, 1561, 1598, 1562, 1564, 3111, 522, 523, 1623, 2956, 771, 1548, 602, 1567]. Sets-Based [1623]. Setups [3033]. SEuS [3158]. Several [2213]. SHACAL [2165]. Shading [2877]. Shape [1028, 1044, 1896, 1894, 988, 1893, 1195, 999, 2412, 1901, 1902, 2026, 1842, 1895, 1810, 1809, 1793, 1903, 1016]. Shaped [1845]. Shapely [2299]. Shapes [3059]. Shaping [2575]. Share [2519]. Shareable [1379]. Shared [974, 323, 2645, 2969, 84, 37, 2356, 2743]. Sharing [965, 983, 2518, 163, 874, 2171, 171, 1959, 3208]. Sharp [490, 2422, 2634]. Sharpening [457]. Shear [1924]. Shear-Warp [1924]. Shedding [1970]. Shelf [795]. Shifters [1691]. Ship [454]. Shock [2929]. Shop [2291, 914, 898]. Shoppers [1471]. Shopping [378, 827]. Short [2937, 2943, 2173, 1726, 2266, 1099]. Short- [2937]. Shortcut [3378]. Shortest [2643, 30, 2476, 2655, 843]. Shot [1951]. Shots [2590]. ShotWeave [1951]. Should [417]. Shoup [2152]. Shrinkage [1012]. Shuffles [967]. SiBIR [979]. Side [183]. Sided [3057, 3074]. SIENA [133]. Sign [817]. Sign-On [817]. Signal [2418, 1569, 1719, 1067, 1929, 2873]. Signaling [16, 2053, 2520, 2595]. Signature [957, 986, 2174, 2462, 954, 2183]. Signature-Based [957]. Signatures [2176, 1442, 2175, 977, 2173, 2174, 979, 985]. Signed [593]. Signer [979]. Signer-Base [979]. Signs [3142]. Silicone [2186]. Silk [3322]. SIMBA [2271]. SIMD [1021, 1713]. Similar [1633, 647]. Similarity [1955, 2836, 1550, 2423, 2838, 1956, 1911, 1912, 2849, 2398, 1465, 1910, 1563, 1907, 1577, 7, 6, 2892]. Similarity-Based [2849, 1577]. Simple [535, 2978, 3058, 3016, 2518, 11, 2861, 842, 712, 2182, 2699, 2349, 3]. Simpler [523]. Simplified [2520]. Simplifying [2631]. Simulate [408]. Simulated [3309, 2481, 3234]. Simulating [2011, 2234, 1500, 371]. Simulation [1879, 1352, 549, 2028, 3218, 640, 2041, 286, 1498, 1888, 1792, 1873, 1880, 1878, 2579, 1353, 2621, 364, 2486, 2407, 1351, 824, 1782, 267, 565, 109, 1882, 1889, 2469, 2039, 520, 1871, 3410, 1511, 1502, 1495, 1924]. Simulations [1450, 536, 1499, 1515]. Simulator [1750, 2386]. Simulators [640, 1874]. Simultaneous [2701, 1063]. Single [2635, 2680, 817, 1015, 2698, 3042, 1065, 1863, 1016]. Single-Camera [1863]. Single-Hop [2680, 2698]. Single-Step [2635]. Single-View [1015]. Sinus [1755]. SIP [2068, 3100]. SIQL [2852]. Site [1362, 1604, 1463, 1078, 1278, 3233]. Sites [1358, 1177, 3382, 423, 1310, 2097]. Situ [2221]. Situated [2024]. Situational [1291]. Situations [3008, 2268]. Size [368, 1631, 1127, 380]. Sized [2364, 60]. Skeleton [228]. Skeleton-Based [228]. Skeletonization [1850]. Sketch [3046, 210]. Sketch-Based [210]. Sketch-Driven [3046]. Ski [2693]. Ski-Rental [2693]. Skill [1755]. Skills [1451, 1754, 3171]. Skin [1839, 1001]. Skip [2652]. SLA [2327]. Slave [1756]. Sleep [2940]. Slice [2189]. Slices [1507]. Sliding [1652]. Sloppy [223]. Slotted [63]. Slow [45, 2436]. Small [1273, 680, 499, 373, 60, 1759, 963, 380, 210]. Small-Sized [60]. Small-World [680]. Smallest [3128, 2660]. Smart [407, 1154, 2650, 1253, 3264, 1155]. Smartcard [828]. Smartcard-Based [828]. Smartcards [1131]. SmartGov [1163]. SMEs [2471]. SMILES [614]. Smith [1087]. Smith-Waterman [1087]. Smooth [1832, 3126]. Smoothing [226, 706, 168]. SMP [184, 1534]. SMV [2646]. Snake [1054]. SNAP [3251]. Snippets [3364].

SNMP [2329]. SNMPv3 [116]. Snow [2876, 3322]. SNP [1065]. SOAP [902, 2972]. Soccer [2407]. Social [2455]. Society [2455, 1173, 783, 1699]. Socio [3273]. Socio-technical [3273]. Sociotechnical [1348]. SOFA [2772]. Soft [1784, 1775, 80, 1889, 1871, 1876, 2459, 1012, 3447]. Soft-Tissue [1784]. Software [1696, 2612, 2546, 1935, 2008, 2314, 800, 1135, 428, 1205, 2966, 348, 316, 1738, 3356, 652, 462, 2507, 826, 1311, 2285, 79, 3010, 2486, 2319, 314, 1285, 503, 1424, 3350, 792, 1982, 2505, 833, 222, 298, 1980, 2626, 836, 1417, 3347, 799, 3354, 3096, 326, 1621, 2304, 3203, 1298, 1991, 635, 3353, 1988, 649, 3093, 1303]. SOL [587]. Solar [3044]. Solo [1750]. Solution [448, 1685, 78, 836, 1719, 3036]. Solutions [1418, 412, 387, 1628, 786, 345]. Solve [2360, 421]. Solvers [2639, 603, 315]. Solving [2452, 2128, 2839, 1213, 610, 3439, 934, 2986, 1501, 3383, 910, 3307, 1504, 916, 2502, 1684, 2405, 2637, 921]. Some [2691, 3081, 3076, 2663, 745, 1137, 2555, 1458, 2131, 2010, 3310, 495, 949, 3131, 662, 2957, 2389]. Sometimes [2199]. Sort [2376]. Sorting [2030, 1632, 925, 373, 843]. SOS [2520]. Sound [550, 416, 1722]. Soundness [2571]. SOUP [800]. Source [3221, 1043, 123, 33, 104, 1583]. Source-Channel [123]. sourced [1860]. Sources [2734, 1962, 642, 2834, 2712, 1225, 2227, 46, 2753, 2445, 1840, 2238, 2991, 646, 1644]. Space [1053, 160, 363, 2263, 720, 746, 1725, 1770, 500, 2412, 3110, 964, 2252, 1651, 2894, 1806, 2990, 2279, 1673, 2653, 678, 1235, 3031, 2967, 840]. Space-Bounded [964]. Space-Economical [500]. Space-Efficient [2653]. Space-Variant [2894, 2279]. Spaces [738, 2024, 2423, 2630, 1572, 2378, 1063, 3136, 2724, 2743]. Spam [117]. SPAMM [1826]. Spanish [3017]. Spanners [2666, 1651]. Spanning [3076, 1333, 1335, 2689, 477, 1321]. Spans [576]. Sparse [1035, 698, 1719]. Sparsely [2442]. Spatial [2011, 3274, 541, 2369, 926, 259, 2911, 174, 1778, 2845, 1912, 759, 1667, 416, 1648, 2884, 2862, 1672, 2023, 1675, 1622, 1669]. Spatial-Temporal [2911]. Spatially [3220]. Spatio [260, 3175, 2848, 1665, 1885, 2213, 1678]. Spatio-Temporal [260, 3175, 2848, 2213, 1665, 1885]. Spatio-Thematic [1678]. Spatiotemporal [2943, 1188]. Speaker [2394]. SpecC [3114]. Specialisation [631]. Specialised [1448]. SPEcialization [624, 628, 625, 145]. SPEcialization/Generalization [624, 628]. Specialized [2230]. Specific [2756, 3336, 1735, 1586, 1137, 261, 1765, 1054, 2505, 1100, 2002, 2897, 774]. Specification [3308, 1732, 2289, 1237, 1312, 2645, 187, 2377, 2337, 1862, 2411, 637, 1425, 1979, 3304, 3012]. Specification-based [187]. Specification-Oriented [637]. Specifications [3104, 3113, 503, 2561, 514, 1701, 271, 3091]. Specified [478, 2236, 397]. Specifying [2488]. SPECT [1811]. Spectral [1069]. Spectrogram [3039]. Spectrometry [1068]. Spectrum [159]. Speculative [3468]. Speech [2418, 2878, 3302, 2108, 1721]. Speech-to-Speech [2108]. Speed [2622, 2326, 2865, 2577, 2590, 345, 34, 2059, 2453]. SPeeDI [341]. Speedup [1995, 2675]. Sphere [1060, 3423]. Sphere-Packing [3423]. Spherical [2894, 2279]. Spike [2931]. Spikes [2859]. Spin [1808]. Spiral [434]. Splice [1077]. Spline [1877, 1712]. Split [588]. Split-Frames [588]. Splitting [2554]. SPM [1922]. Spoken [2110, 417]. Spontaneous [2022, 2010]. Sporting [3382]. Sports [3296, 3461]. Spots [3172]. Spreading [2467]. Spring [3085, 1876]. SQL [1945]. SQL4X [264]. Squad [2016]. Square [2644, 2660]. Squares [1033].

Squashing [744]. SRTP [125]. SS [1365]. SSL [794]. SSL/TLS [794]. SST [2170]. Stability [2347]. Stabilizing [2348]. Stable [291, 2010, 607, 2347, 292, 998]. Stack [1734]. Stacking [731]. Stage [3458, 1717, 1931, 2669, 1748, 2405]. Stages [206]. Staggering [3377]. Stampi [1522]. Stampi-I [1522]. Stampi-I/O [1522]. Standard [1239, 2194, 1906, 417, 1126]. Standardized [1800]. Standards [2608]. Star [2492]. Start [459, 871, 1482, 2618, 45, 1078]. State [2684, 3315, 363, 2630, 205, 491, 78, 3175, 1438, 3110, 3293, 1149, 803, 1923, 529, 3292, 2339, 345, 3031, 2200]. Statecharts [1436, 2312, 2303]. Stateless [2061, 2519, 2352]. Statements [2635]. States [2033, 3109, 3307]. Static [44, 1640, 2003, 27]. Stationary [2926]. Stations [65]. Statistical [2937, 1014, 1930, 2084, 1894, 1846, 2089, 789, 1904, 2890, 1901, 2091, 1902, 1905, 1895, 1726, 1794, 1235, 1818, 1591, 2086, 1822, 1788]. Statistically [2158]. Statistically-Hiding [2158]. Statistics [1657, 2881, 873, 2882]. Status [20]. Steerable [3227]. Steering [454]. Steganographic [671, 203]. Steganography [953]. Stegonagraphy [832]. Steiner [2425]. STeLLa} [3030]. Stem [1773]. Stemmer [1635, 2450]. Stemming [1634, 1198]. Stenosis [1853]. Stent [1908]. Step [1171, 552, 1953, 2635, 853]. Stepping [2611, 2202]. Stepping-Stone [2611]. Steps [1175, 473, 3111]. Stereo [2871, 997, 2422, 2890, 2277, 3041, 2920]. Stereoscopic [1859, 1753, 1933]. Stereotactic [1857]. Stereotypes [1292]. Stereotypical [1293]. Steuer [1141]. Stick [223]. Still [1628]. Stochastic [904, 1803, 2620, 2899, 2018, 724, 1827]. Stock [3179]. Stone [2611]. Stones [2202]. Stop [1122, 1146, 1128, 1123]. Storage [672, 2177, 671, 2159, 2442, 2362, 191, 3241, 183, 203, 2419, 3381, 2054]. Stories [3181, 453]. Storing [2432, 1949]. Story [1951]. Straight [3414, 3050, 3075]. Straight-Line [3414, 3050, 3075]. Strain [1836]. Strategie [2471]. Strategies [838, 3218, 695, 3221, 872, 3071, 908, 2769, 2812, 460, 1169, 841, 1538, 1153, 718, 423, 3246, 165]. Strategy [2955, 2544, 788, 875, 430, 48, 2465, 2497, 878, 2538, 2515, 1515, 3248, 793]. Stratified [2546, 277]. Stream [2196, 487, 980, 981, 18, 3381]. Streaming [2052, 2529, 2598, 2605, 125, 195, 2055, 2413, 2517]. Streams [2611, 1642, 1627, 1970]. Streett [2116]. Strengthening [628, 1933]. Stress [1836]. Stretch [477]. Strictness [2122]. String [1629, 1649, 1630, 1628, 500, 3305, 1538, 3155]. Strings [1235]. Stripe [1016]. Striping [3377, 2600]. STRIPS [3032]. Strong [3448, 563, 964]. Stronger [923]. Structural [229, 1028, 1006, 325, 1724, 1940, 945, 2647, 1080, 2224]. Structure [1637, 2221, 1367, 2838, 3166, 2921, 2236, 2503, 2378, 3176, 1368, 1556, 911, 520, 996, 948, 1651, 1061, 1079, 3034, 1020, 1921, 159, 220, 860, 3122, 3154, 3158]. Structure- [1367]. Structure-Sweetness [3176]. Structured [3443, 1646, 761, 416, 3193, 3440, 820, 3155, 2266, 737, 643, 1641]. Structures [2127, 2423, 3457, 263, 2320, 1401, 1402, 495, 2672, 1956, 2364, 500, 1890, 560, 1075, 2161, 2023, 2653, 2419, 1820, 3093]. Structuring [1358, 2724, 153, 3168, 774]. Stuck [2646]. Stuck-at [2646]. Student [444, 1222, 1271, 445]. Students [1223, 3346]. Studies [1812, 2981, 3024]. Study [1362, 580, 1736, 982, 1347, 368, 695, 2337, 1452, 749, 3189, 621, 3004, 1223, 2988, 2013, 40, 59, 1244, 398, 393, 1449, 1538, 380, 2371, 2979, 111, 1952, 2454, 2957, 3160, 253, 168, 7, 4, 85, 2335, 900]. Studying [10]. Style [826]. Sub [715]. Sub-goals [715].

Subclasses [3135]. Subgraph [497]. Subgraphs [3053]. Subgroup [750, 759]. Subject [1922, 2942]. Subject-Dependent [2942]. Subjective [134, 2889]. Subjects [2937, 1838]. Submodule [3091]. Subscription [3393]. Subsequences [1633]. Subset [2263, 1331]. Subsets [766, 665]. Subshifts [2017]. Subspace [3408, 3195]. Substitution [2300]. Substitutivity [1208]. Substructure [737]. Subtraction [1214]. Subtyping [310, 547]. Succeeding [454]. Success [1168]. Successors [3186]. Sufficient [2360]. Suffix [1097, 1629, 706]. Suffixes [500]. Suggested [2886]. Suggestions [2846]. Sugiyama [3081]. Suites [1314]. Sulcal [1804]. Sulci [1930]. Sum [691, 1332, 789, 114, 993, 2191, 1329]. Sum- [993]. Sum-of-Disjoint-Product [789]. Sum-Only [2191]. Summaries [3158]. Summarization [3280, 369]. Summary [237, 234]. Super [1630]. Super-Alphabets [1630]. Superimposed [2221]. Superiority [2865]. Superquadrics [1717, 1009]. Supersingular [969]. Supertrees [1103]. Supervised [1199, 3438, 775]. Supine [1920]. Supplementary [406]. Supply [2471, 2704, 2703]. Support [1443, 2944, 2763, 3210, 1784, 3268, 2818, 698, 2974, 1193, 1144, 2951, 2319, 2939, 754, 1253, 1243, 3405, 52, 95, 280, 1959, 714, 655, 13, 2993, 1900, 2990, 2777, 2798, 1206, 1151, 91, 1619, 1952, 3295, 167, 2480, 3333, 1272, 1760, 2771, 3453, 728, 392, 2604, 1348, 1475, 3277, 2413, 85, 1668]. Supported [1059, 1077]. Supporting [1139, 3211, 1367, 1458, 1274, 2844, 422, 2364, 1162, 3323, 25, 1973, 148, 3229, 414, 420, 2597, 3024, 1180, 2723, 3259, 2725, 3208]. Supports [2711]. Surface [1802, 1498, 2677, 2038, 1897, 1862, 1805, 3449, 1915]. Surfaces [1053, 1919, 2889, 1831]. Surfing [1480]. Surgery [1783, 1888, 1758, 1771, 1781, 1756, 1762, 1759, 1751, 1753, 1867, 1871, 1750, 1755]. Surgical [1854, 1862, 1770, 1762, 1751, 1859, 1870, 1864, 1874, 1760, 1755]. Surrogates [1265]. Surveillance [2944, 3466]. Survey [1121, 445, 3170, 1581, 2316]. Survivability [1688, 788, 1991]. Survivable [1389, 1390]. Survival [1610]. Suture [1769]. SVM [3429, 767]. SW [419]. Swap [2359]. Swarm [897]. Sweep [3110]. Sweep-Line [3110]. Sweetening [1459]. Sweetness [3176]. SWIFI [1695]. Switch [31, 28, 37, 34]. Switching [2297, 30, 58, 46, 3243]. Switzerland [1174, 1170]. Sybil [682]. Syllables [767]. Symbol [2133, 213]. Symbolic [650, 321, 319, 2926, 2643, 342, 3107, 2089, 67, 317, 515, 332, 2628, 2633, 3108, 3093, 2632]. Symbols [2562, 633, 214]. Symmetric [3093]. Symmetrically [3054]. Symmetries [2290, 2700]. Symmetry [321, 3094]. Symptoms [2625]. Synapse [2999]. Synchronisation [538]. Synchronization [2713, 2864, 3114, 2016, 2389, 126]. Synchronized [1242, 1920, 1049]. Synchronous [895, 1299, 347, 1998, 1996, 552, 981]. Synchrony [2757]. Syndrome [1798]. Synergy [3018]. Syntactic [1943]. Syntactical [860]. Syntax [1747]. Synthesis [3367, 1792, 2912, 1424, 3302, 917, 1743, 3096, 1423, 1722]. Synthesizing [909]. System [2102, 835, 1783, 1139, 3200, 1687, 2984, 2262, 1593, 2944, 2104, 299, 1279, 2046, 1682, 3210, 2719, 801, 781, 428, 2404, 2835, 995, 1524, 2396, 2274, 1251, 3016, 68, 1367, 3457, 1057, 1570, 1211, 2933, 568, 1456, 2099, 688, 2987, 1363, 2045, 609, 2377, 2795, 615, 617, 161, 636, 614, 2934, 1832, 2733, 1643, 3266, 3038, 807, 163, 1791, 2051, 3456, 2326, 2399, 1877, 2479, 2355, 894, 1770, 1781, 503, 1555, 3437, 1296, 3419, 1505, 3384, 2772, 2918, 637, 196, 112, 2490, 100, 2462, 1782, 1983, 1757, 124, 10, 3245, 1369, 2108, 120, 2428, 2805, 2203, 20, 2474, 616, 1046, 2408, 3044]. System [2469, 2744, 1147, 1959, 1200, 2251, 1858, 1766, 663, 1538, 618, 2109, 2198, 2463, 917, 13, 1867, 3021, 2406, 1829, 1744, 2095, 42, 809, 1779, 2929, 1206, 1151, 107, 942, 3096, 893, 528, 2764, 2820, 2436, 2467, 3024, 203, 1865, 1864, 1863, 2936, 90, 1354, 3393, 2569, 1049, 1760, 2613, 1780, 3391, 3453, 1988, 646, 149, 1348, 157, 3204, 2257, 194, 2413, 1755, 32, 3431, 1960, 2892, 2107, 3383, 62]. System-Z [618]. Systematic [1741, 431]. Systems [2247, 2050, 2365, 3206, 3308, 586, 3104, 1110, 1299, 2576, 2376, 341, 343, 540, 3105, 1935, 1687, 2780, 3370, 2288, 2024, 2473, 634, 811, 2871, 593, 3349, 2207, 2106, 275, 641, 339, 320, 3359, 3252, 890, 891, 2338, 199, 1418, 2237, 3027, 2649, 3367, 796, 568, 2234, 1707, 2987, 567, 2167, 3251, 2619, 577, 2711, 652, 180, 1156, 2027, 376, 2350, 1452, 156, 2974, 3056, 2843, 3291, 1290, 3357, 653, 1998, 2294, 887, 1144, 3207, 1209, 1996, 2622, 346, 2298, 317, 358, 1702, 782, 2580, 1374, 58, 2718, 2381, 2271, 1316, 940, 3307, 201, 1435, 1210, 792]. Systems [2662, 140, 3256, 1982, 2384, 2882, 3293, 344, 1346, 3006, 3247, 132, 139, 2438, 2623, 1666, 3400, 1425, 830, 2736, 2305, 515, 1596, 517, 1719, 2475, 2737, 2356, 520, 2497, 1992, 200, 2100, 1560, 1743, 1750, 99, 2979, 2309, 2852, 2502, 2241, 1979, 1562, 1563, 795, 815, 666, 3001, 1701, 1546, 1566, 2653, 307, 2470, 678, 657, 1999, 1954, 669, 1093, 2727, 3288, 2726, 2641, 2784, 3205, 1884, 1730, 405, 2477, 75, 1690, 403, 2312, 2857, 1963, 1748, 3248, 3267, 2720, 1216, 1397, 1398, 330, 1668, 3011, 659, 3346]. Systolic [114].

T1 [1785]. T1-Weighted [1785]. Table [2589, 1568, 3216]. Tabled [276, 287]. Tables [1586, 2236, 686, 1945, 725, 683, 2224]. Tabular [610, 721]. Tabulation [277]. Tactical [3453]. Tactile [406]. Tag [13].

Tailoring [428]. Take [1431, 1405]. Takes [3035]. Taking [390]. Talking [1266]. Tandem [496, 1070]. Tapering [1002]. Target [747, 3220, 1910]. Targets [3127, 380]. Tarzan [670]. Task [1372, 2197, 3220, 2687, 3020, 2014, 927, 1219, 143, 1639, 1480]. Tasks [2686, 568, 3219, 494, 2899]. Tax [1141, 257]. Taxes [1135]. Taxonomies [632]. Taxonomy [117]. Taylor [2644]. TCLP [310]. TCP [2588, 2526, 2448, 80, 96, 43, 93, 94, 49, 45, 836, 55, 91, 2447, 2454, 125, 2499]. TCP-based [2526]. TCP-Friendly [125]. TCP-Vegas [49, 45]. TCS [894]. TCTL [1432]. TD [721]. TDMA [2495]. TDMA/FDMA [2495]. TDMA/FDMA-Based [2495]. Teaching [1223, 443, 442, 2457]. Team [2407, 424, 3430]. Teams [468, 1098, 461, 434, 2725]. Teamwork [2827]. Teaq [198]. Technical [227, 1707, 1242, 1855, 3273]. Technically [1172]. Technically-Oriented [1172]. Technique [482, 720, 1807, 699, 372, 2481, 1024, 1267, 2426, 3070, 3378, 1439, 1951]. Techniques [3087, 2735, 1977, 1955, 2318, 2270, 333, 1785, 2061, 3361, 1725, 3291, 2478, 2089, 3450, 3448, 762, 3293, 3400, 381, 3112, 3467, 2607, 152, 1196, 2641, 3095, 2994, 1608, 2005]. Technologies [2464, 1252, 2078, 1458, 1146, 1169, 2508, 1948, 1668]. Technology [2058, 1280, 1275, 1993, 1491, 1109, 2333, 2817, 1160, 3347]. TechSat [3413]. Tele [1766, 1161, 1767]. Tele-consulting [1161]. Tele-echography [1766, 1767]. Tele-voting [1161]. Telecommunication [2804, 67, 2305]. Teleoperated [1768]. Teleoperation [1769]. Telephone [58]. Tell [426]. Temperature [807]. Template [2218, 1896, 1928]. Templates [1189, 1075, 1297]. Temporal [575, 532, 1014, 2549, 2040, 2837, 1186, 747, 260, 3279, 1238, 358, 752, 2911, 174, 2397, 1912, 3175, 2848, 2400, 588, 1944, 2253, 2950, 530, 1701, 253, 2213, 2402, 1665, 1885]. Temporal-Safety [358]. Temporally [3220]. Ten [3289]. Tendencies [1270]. Tennis [1052]. Tension [1133]. Tensions [2248]. Tensor [1803, 1024, 1806, 1810]. Tensor-Warped [1806]. Tensors [996]. TER [1767]. Term [2937, 2943, 3282, 1043, 758, 707, 724, 3133]. Term-Based [707]. TERMGRAPH [2316]. Terminal [2610]. Terminals [374]. Termination [349, 2294, 1729, 1742, 2352]. Terminological [2854, 2572]. Terminology [3018]. Terms [2829, 485, 1640, 934, 578, 2846]. Ternary [2494]. Terrain [1674]. Terrestrial [2579]. Territorial [3466]. Territory [3099]. Test [1085, 1314, 1691, 461, 3097, 1440, 457, 432, 3096, 522, 435, 2956, 431, 1591]. Test-Driven [432]. Test-Sets [522]. Testbed [1499]. Testers [468, 461]. Testing [187, 3291, 460, 3066, 607, 424, 1435, 429, 514, 2240, 1895, 465, 819, 3095, 126]. Tests [650, 430]. Tetrahedralization [2678]. Texas [3286]. Text [3183, 218, 192, 2840, 749, 2085, 1446, 153, 632, 381, 1653, 768, 3018, 2653, 1627, 1641, 3013, 774]. Text-Based [192]. Text/Graphics [218]. Texts [1634, 1631, 3022, 3184, 2095, 1641]. Textual [1281, 1277, 2283]. Texture [1904, 2277, 993, 214, 2867, 1915, 2868]. THD [1543]. Their [2190, 2003, 1458, 3448, 2882, 2308, 2154, 1860, 2389, 1265]. Them [2322]. Thematic [1678]. Themes [2213]. Theorem [2553, 474, 336, 2562, 2644, 524]. Theorem-Proving [336]. Theorems [586, 2965]. Theoretic [3054, 378, 278, 955, 2371, 287, 1562, 2814]. Theoretical [581, 852, 3220, 3196, 2519, 1638, 3455]. Theories [2215, 2225, 651, 2556, 3460, 2964, 592]. Theory [1073, 262, 1175, 239, 2127, 2000, 485, 2843, 3399, 260, 1614, 170, 1726, 2904, 247, 3020, 2916, 3197, 597, 3398, 584, 2390, 1543]. Therapy [2396, 1858, 1855, 2036]. Thermochemical [3216]. Thermocoagulation [1855]. Thesaurus [1255, 3018]. Thesaurus-Based [1255]. Theta [3059]. Thickness [3060, 1804]. Thing [679]. Third [1293, 3112, 2527]. Third-Generation [3112]. Thoracic [1847]. Thread [3325, 1523]. Thread-Safe [1523]. Three [3049, 3050, 184, 1757, 594, 3073, 1865, 1020, 1860, 899]. Three-Dimensional [3050, 1757, 1865, 1020, 1860]. Three-Tier [184]. Three-Valued [594]. Threshold [977, 2394, 2162, 2171, 972, 2689, 973]. Thresholding [231]. throughout [2319]. Throughput [494, 80, 3241, 3240]. Thwarting [1387, 1388]. Thyself [1311]. Tibial [1881]. Tickets [1155]. Tid [3441]. Tier [184]. TIF2VEC [235]. Tight [2356, 3084]. Tiled [1204]. Tiling [1094]. Tilt [390, 2279]. Time [1696, 2365, 2937, 903, 2581, 549, 2717, 794, 1997, 741, 339, 989, 3077, 2338, 1418, 3027, 2767, 44, 1050, 3458, 1347, 82, 488, 368, 1700, 2768, 936, 1440, 3219, 1419, 536, 1761, 1300, 1931, 2986, 653, 2478, 1507, 887, 2295, 3270, 997, 1985, 2174, 752, 3190, 2479, 1758, 2667, 1729, 2271, 2847, 1435, 1984, 108, 2747, 124, 1337, 1890, 56, 842, 49, 2623, 1742, 2594, 1575, 1425, 1943, 1992, 2998, 2463, 1720, 2252, 1743, 42, 1021, 1866, 2940, 1876, 54, 1721, 276, 3034, 1994, 886, 1701, 2535, 167, 1422, 1999, 155, 544, 3304, 3133, 1577, 1428, 773, 1932, 3173]. Time [1115, 2593]. Time-Dependent [1425]. Time-Ordered [3173]. Time-Outs [544]. Time-Safety [1985]. Time-Series [2937, 2338, 887, 752, 886]. Time-Triggered [1696, 1997, 1422]. Time-Varying [3219]. Timed [3105, 3403, 3291, 1438, 563, 3371, 1430, 3090, 1429, 1428, 1427, 1439, 1987]. Timeline [3276]. Timely [1992]. Times [2330, 2496, 2392]. Timestamping [2699]. Timetabling [3434, 2395]. Timing [1387, 1388, 328]. TINMAN [2198]. Tip [1776]. Tipper [3461]. Tissue [1784, 1839, 1813, 1785, 2384, 1910, 1871, 1835, 1876, 2273]. Tissues [1775, 1889, 1840]. TLS [794, 1043, 956, 2335]. Tokens [833]. Tolerable [2611]. Tolerance [38, 1692, 1563, 1576]. Tolerant [2365, 1687, 2738, 138, 2443, 684, 1410, 1690, 2646]. Tolerating [2164]. Tomographic [1926]. Too [679]. Tool [2471, 1154, 340, 341, 343, 1734, 355, 1494, 3274, 342, 180, 2579, 1447, 1239, 2621, 1770, 70, 3437, 1790, 1870, 655, 2109, 3276, 613, 1695, 2311, 1348, 1143]. Toolkit [1517]. Tools [1135, 387, 1476, 463, 1758, 451, 1243, 2317, 3345, 2002, 1864, 3108]. Toolset [3106]. Top [2717, 3289, 1502]. Topics [2950]. Topographic [221]. Topological [2378, 2379]. Topologically [1669]. Topologically-Consistent [1669]. Topology [3137, 2025, 920, 159]. TOPSIS [1708]. Total [1791, 1781, 1779, 1033, 1012]. Tourist [377]. Tower [3044]. Trace [975]. Traceable [2536]. Tracer [162]. Traceroute [2511]. Traces [557, 303, 2731, 647]. Tracing [233, 2202]. Track [1249]. Trackable [1664]. Tracker [1867, 1863]. Tracking [838, 1861, 989, 1050, 1808, 1040, 990, 827, 988, 1757, 109, 2876, 1828, 1030, 1474, 435, 1051]. Tract [1851, 3174, 1925]. Tractability [1065]. Tractable [3057, 2673]. Tractography [1803]. Trade [2590]. Trade-Off [2590]. Tradeoffs [2595]. Trader [2755]. Trading [688, 1707, 2729]. Traditional [152]. Traffic [173, 2516, 72, 2061, 1155, 2526, 2575, 46, 2065, 2745, 912, 2039, 2522, 2593, 35, 3389, 2528, 168]. Trailing [272]. Trainable [211]. Trainers [1754]. Training [3427, 1813, 1782, 1870, 1492, 2267, 1874, 1780, 2086, 1755]. Trajectories [802]. Trajectory [2632]. Trans [1764]. Trans-nasal [1764]. Transaction [3308, 76, 2533, 1162, 267, 140, 2728, 203, 1975]. Transaction-Based [203]. Transactional [2231, 1163, 751, 1109, 1148]. Transactions [548, 2759, 2426]. Transcoding [372]. Transcript [1386, 1394, 1412, 1402, 1404, 1409, 1392, 1396, 1405, 1390, 1411, 1384, 1400, 1388, 1382, 1410, 1407, 1398]. Transcriptome [1100]. Transducers [491]. Transductive [722, 3143]. Transfer [1512, 1136, 2159, 2690]. Transferring [3226]. Transfers [3240]. Transform [1714, 1039, 3428, 768]. Transformation [3358, 2288, 2314, 2215, 2225, 2318, 286, 2293, 2295, 2296, 2315, 2298, 2299, 935, 2301, 3305, 1056, 2309, 2310, 2304, 3284, 2906]. Transformations [1304, 2291, 359, 1227, 2320, 3221, 1302, 938, 2319, 2348, 1790, 2628, 2306, 304, 1303]. Transforming [2289, 1137]. Transient [43]. Transit [2940]. Transition [3175, 2640, 2312]. Transitions [2033, 3303, 3109, 361, 724]. Transitive [2175, 259, 1514]. Transitivity [2017]. TRANSLATE^{T M} [2103]. Translating [1361, 2124]. Translation [331, 2104, 2105, 2106, 3113, 2082, 2107, 3109, 2088, 2089, 2090, 2091, 2108, 2092, 3451, 2093, 2110, 2100, 2329, 296, 1078, 2101, 2096, 1256, 2098]. Transmission [3270, 56, 120, 123, 2463, 99, 54, 2063, 3101]. Transmitting [1658]. Transparency [2455]. Transparent [2818, 1158]. Transplantation [1773]. Transport [76, 2587, 15]. Transportation [845, 1340, 2012]. Transpositions [1632]. Transrectal [1761]. Transversal [619]. Trap [1169]. Travel [411, 1206]. Traveling [922, 1337]. Traversal [2641]. Traversals [351]. Treatment [1689, 1777, 1256, 1760, 2268]. TREC2001 [1249]. Tree [360, 2952, 3019, 2461, 2127, 1682, 2976, 359, 989, 2666, 2128, 1640, 3441, 1333, 2425, 705, 2690, 257, 17, 2120, 1849, 761, 516, 3440, 2564, 2119, 1321, 86, 3380, 3133, 1712, 2121]. Tree-Decomposition [2976]. Tree-Like [3380, 2127]. TreeITL [3441]. TreeITL-Mine [3441]. Trees [1649, 943, 3076, 906, 1968, 3189, 1331, 3075, 474, 1573, 1335, 704, 2704, 2928, 706, 2494, 1943, 2689, 1648, 477, 727, 1850, 2965, 2969]. Treewidth [1327, 497, 2664]. TRELLIS [1447]. Trends [1495]. Triad [2222, 678]. Trial [124, 1863]. Trials [2934]. Triangle [2297, 1892]. Triangle-Free [2297]. Triangulation [2677, 1592]. Trie [664, 1653]. Tries [2706]. Trigger [1189]. Triggered [1696, 1997, 1422]. Triggers [441]. Trip [42]. trivial [2013]. Trunking [58]. Trunking/Switching [58]. Trust [2472, 831, 1403, 1404, 3460, 3400]. Trusted [1397, 1398]. Trustful [1138]. Trustless [3224]. Truth [1788, 2855]. Truthing [209]. TTP [1702]. TTP/C [1702]. Tube [1852]. Tubular [1890]. Tumor [1774, 1796, 1797, 1765, 1795]. Tumors [1855, 1788]. Tuning [1513, 2861, 3243, 3036]. Tunneling [1236]. Tunnelling [3325]. Tuples [1190]. Tupling [1737]. Turbine [801]. Turbines [3029]. Turbo [2463]. Turing [2372]. Turkish [1192, 1197, 1203]. Turner [584]. Turning [1460, 425]. Tutorial [2314, 931, 2315, 3294, 2607, 2608, 2079, 2605, 2606, 318]. Tutorials [622, 1318]. Tutoring [3457]. Tweakable [950]. Two [3414, 2033, 1325, 2338, 2128, 1717, 3049, 2661, 491, 492, 1931, 1961, 789, 2988, 756, 2669, 763, 3074, 2419, 3447, 2016]. Two-Dimensional [2016]. Two-Layered [2419]. Two-Phase [756]. Two-Sided [3074]. Two-Stage [1931, 2669]. Two-Way [2128, 2661, 491]. Type [1733, 239, 754, 2570, 2571]. Type-Certifying [2570]. Type-Directed [1733]. Type2 [2932]. Typechecking [249]. Typed [2298, 1433]. Types [533, 3294, 531]. Typhoon [3175]. Typical [1328]. Typing [547]. Tyranny [193].

Ubiquitous [1350, 2285, 1357, 400, 3212]. UCLID [2636]. UDP [2065, 836, 2454]. UK [1168, 1172]. Ultra [2859]. Ultra-Rapid [2859]. Ultrasonic [1765, 1038]. Ultrasound [1873, 1897, 1868, 1757, 1898, 1867, 1865, 2892]. UML [1306, 1299, 3358, 1314, 1315, 1312, 2242, 1437, 1291, 1419, 1472, 1289, 156, 1300, 1290, 653, 1292, 2420, 2779, 1436, 1316, 1438, 1317, 2222, 1307, 2228, 391, 3096, 1305, 1310, 1288, 1297, 1313, 2312, 1287, 1303, 1318]. UML-Based [1317]. UML/OCL [1289]. UMLinux [1695]. UMLsec [1316]. UMTS [2579, 377, 3101]. Unanticipated [3350]. Unavailability [1694]. Unawareness [1186]. Unbalanced [963]. Unbound [1076]. Unbounded [337, 332]. Uncertainty [2676, 305, 589, 937, 2251, 1674, 1622, 1188]. Unconditionally [2172, 2155]. Undecidability [564]. Undecidable [540]. Under-Approximation [321]. Undermining [2613]. Understanding [402, 3016, 2040, 2234, 1124, 247, 216, 2919]. Undirected [2673]. Undisplayed [1927]. Unfairness [2062, 2515]. Unfamiliar [2922]. Unfoldings [362]. Unicast [2518, 195]. Unification [286, 507, 2849]. Unified [3225, 1286, 1760]. Uniform [3123, 3128, 2319, 2707, 2424, 305, 589, 2390, 3422]. Unifying [1011]. Unigesture [385]. Uninterpreted [320]. Union [1165]. Unions [3440, 3140]. Unique [985]. Unit [2658, 1854, 527, 514, 1773]. UNITE [2827]. Units [2301]. Universal [2547, 68, 962, 2350, 379, 2380, 1432]. Universality [2019]. Universally [984, 2347]. Universe [463]. University [1244, 1221]. Unix [825, 3011]. Unknown [3146, 2236, 2618, 2152]. Unlabeled [729]. Unparsing [2307]. Unraveling [2084, 247]. Unrestricted [2646]. Unsatisfiable [495]. Unsplittable [1326]. Unstructured [3404, 220]. Unsuitable [1169]. Unsupervised [3443, 746, 2921, 1007, 1005, 3043, 2869, 2280]. Untangle [1808]. Untrained [3415]. Up-to-Date [818]. Update [2831]. Updates [572, 263, 308]. Updating [2394, 258]. Upgraded [795]. Upgrading [1991]. Upload [3210]. upon [2982]. UPPAAL [1416]. Upper [2180]. Upright [1715]. Urban [2461, 2040, 2042]. USA [621]. USA-Advisor [621]. Usability [382, 1677, 1265]. Usable [367]. Usage [2046, 2783, 419, 1600, 404, 2784, 3230, 127]. Usage-Centered [419]. Usage-Oriented [2784]. Use [1680, 2464, 3462, 892, 2099, 1294, 1122, 1521, 1223, 1758, 2928, 3098, 3395, 867, 1254, 635, 1917, 1099, 3277, 399, 2420]. Used [938, 3112, 2862, 1969]. User [1130, 212, 1261, 2783, 1456, 2841, 166, 827, 2844, 453, 410, 375, 3419, 637, 2444, 1378, 2065, 398, 1606, 3229, 414, 97, 1681, 162, 1969, 1270, 1468, 403, 2723, 210, 1265, 2771]. User-Defined [2444]. User-Driven [1468]. User-System [1456]. Users [2810, 3386, 369, 379, 2534, 3208]. Using [2052, 1696, 3407, 1875, 2064, 1053, 2925, 2734, 1110, 479, 630, 1352, 2418, 2639, 1734, 1649, 3372, 580, 1512, 794, 1039, 1611, 3210, 1803, 2719, 275, 1275, 1715, 1569, 3324, 989, 1032, 2949, 603, 2196, 2549, 320, 891, 1802, 992, 2927, 3064, 1358, 902, 832, 1050, 3463, 2630, 3412, 2057, 2485, 82, 297, 2643, 333, 2791, 2987, 906, 673, 631, 1796, 1916, 2061, 1968, 3156, 2818, 2328, 2377, 1000, 1896, 1379, 1528, 2969, 1472, 1289, 1401, 1402, 1807, 2556, 3439, 1931, 1290, 3189, 1832, 224, 2087, 259, 3420, 653, 2440, 3158, 873, 324, 2752, 1007, 3415]. Using [3441, 1552, 496, 2056, 2425, 2235, 701, 67, 1919, 1368, 1238, 2713, 997, 1922, 2575, 1185, 2479, 875, 2533, 2276, 1800, 1778, 106, 754, 1805, 755, 3419, 3384, 16, 3450, 3438, 3448, 1009, 1939, 2067, 57, 2444, 3408, 11, 267, 2115, 2505, 761, 3424, 3302, 1462, 1751, 711, 222, 1902, 103, 2489, 2435, 1859, 1844, 2277, 632, 1828, 416, 2222, 1889, 2481, 374, 1898, 1910, 3451, 1801, 2093, 3042, 1027, 1038, 2631, 2463, 2817, 993, 1867, 1653, 1214, 1750, 1840, 2993, 2095, 42, 732, 776, 3170, 1900, 2869, 767, 2940, 803, 2852, 230, 768, 276, 1599, 2950, 1636, 1387]. Using [1388, 92, 2460, 2522, 2904, 1838, 1067, 2304, 247, 3020, 802, 880, 345, 1194, 3041, 3114, 2470, 2424, 136, 2930, 2644, 1310, 1018, 2482, 2454, 1266, 2647, 421, 2419, 2214, 1915, 3339, 2957, 2956, 3333, 2096, 1921, 77, 2054, 2613, 1820, 546, 920, 858, 1872, 3390, 3160, 232, 3453, 1818, 75, 771, 859, 646, 3100, 921, 2255, 3240, 2977, 2257, 1071, 3436, 1932, 1924, 3248, 1010, 1406, 1407, 3452, 1824, 1583, 3180, 1608, 3234, 330, 85, 165, 2224, 2967, 3173, 1571, 685, 395, 2646]. Utilities [1221]. Utility [2334, 196, 846, 2626]. Utilization [3273, 2595, 3389]. Utilizing [3416, 2890, 408].

V [2139, 2147]. V1 [2862]. v2.0 [3030]. Vacuum [352]. Vadis [1169]. Vague [2839]. Valid [2245]. Validate [127]. Validating [2264]. Validation [1789, 339, 1785, 3306, 798, 1787, 3036, 1786, 1260, 1788]. Validity [3097, 353]. Value [2292, 1594, 3419, 1113]. Valued [3407, 2643, 594, 762, 604, 1576, 1428, 2965, 354, 542, 1548]. Values [1549, 2855, 721]. Vantage [2423]. Variability [1028]. Variable [3454, 1549, 1611, 2218, 489, 3166, 1573, 3159, 706, 3445, 37, 232, 63]. Variable-Length-Don’t-Care [3159]. Variables [591, 2562, 3440, 2930]. Variance [696]. Variant [2152, 2894, 2279]. Variants [3276, 1277, 1077, 2224]. Variation [845, 43, 1012]. Variational [693, 1044, 3429]. Variations [2888]. Varieties [969]. Variety [2176]. Varying [848, 3219]. Vascular [1879, 1884]. VBR [168]. VCC [3369]. VCR [120]. VCRM [1134]. Vector [226, 1224, 1658, 3427, 698, 1193, 3136, 2939, 3405, 714, 1900, 2990, 230, 2424, 1008, 2414, 3390, 728, 220].

Vectorization [228]. Vectors [1615, 633, 2993]. Vegas [49, 45]. Vegetable [2028]. Vehicle [2291, 863, 2101, 1988, 1982]. Vehicles [1032, 3038]. Velocity [2249, 1297]. Venn [3088]. Ventral [2913]. Ventricle [1836]. Ventricles [1826]. Ventricular [1809, 2941]. Venture [3099]. Verb [2090]. Verifiable [1178, 985]. Verification [3308, 341, 343, 2546, 794, 1434, 335, 351, 323, 2645, 337, 316, 2377, 527, 2622, 2394, 336, 798, 2411, 2636, 3112, 3371, 2631, 1430, 1979, 1422, 3114, 2644, 808, 33, 318, 330, 793]. Verifier [2163, 2640]. Verifying [320, 3102, 517, 3101]. Versa [247, 166]. Versatile [1783, 1376, 1695]. Version [2103, 1593, 1572, 930, 1115, 2728]. Versioning [1461]. versus [960, 3252, 1575, 3311, 1381]. Vertex [1334, 1336, 2673, 2674, 2313]. Vertex-Connectivity [1336]. Vertical [2580, 1764]. Very [745, 2369, 2266, 2453]. Vessel [2282]. Vessels [1841]. Vga [1213]. VI [2140, 2148]. via [3103, 325, 593, 473, 2926, 982, 323, 1792, 1737, 488, 1068, 3433, 2088, 308, 2794, 2408, 3169, 588, 1430, 2614, 1829, 1092, 2797, 2571, 1746, 2624, 3025, 1528]. Vibrational [1213]. Vice [247, 166]. Vickrey [2537]. Video [1950, 2576, 2589, 1052, 3280, 991, 3270, 1800, 124, 2848, 1055, 2590, 2591, 2070, 1952, 3376, 1249, 2600, 3281, 3381, 2054, 1712, 1953, 2055, 1265, 2517, 1951, 2592]. Video-Based [991]. Videography [1859]. Videos [3296]. Vienna [1134]. View [784, 251, 263, 1015, 2716, 1606, 2923, 1051]. View-Based [251, 2923]. Viewer [3276]. Viewing [1933, 1953]. Viewpoints [3188]. Views [2344, 2245, 2223, 3067, 1872]. VII [2141, 2149]. VIII [2142, 2150]. Violation [1177]. ViRoom [1049]. Virtual [1283, 1784, 14, 73, 1207, 2078, 3386, 1481, 1851, 1927, 2721, 60, 1503, 833, 130, 3042, 1992, 1920, 1349, 2214, 2726, 1872, 2757]. Virtualized [679]. Viscous [1916]. Vision [1057, 2891, 2914, 2865, 2880, 3044, 2886, 2894, 3041, 2897, 2920, 1051, 2917]. Vision-Based [3044, 2920, 1051, 2917]. visiting [2923]. Visual [3104, 3287, 3317, 2888, 2877, 3273, 2318, 2913, 3285, 3289, 2921, 3278, 1768, 3299, 2865, 1250, 2847, 2885, 2171, 2899, 2890, 2882, 1751, 2866, 383, 408, 1021, 2900, 2904, 2916, 3281, 2901, 2868, 2283, 1287, 2892, 1156]. Visualisations [1174]. Visualization [3283, 2404, 3280, 3082, 1854, 1868, 3080, 1798, 1886, 1925, 1277, 1867, 3286, 1926, 1923, 880, 3288, 3284, 1182, 3277]. Visualizations [3283]. Visualize [3276]. Visualizing [415, 2399, 2946, 3070]. Visually [2458]. vivo [1836, 1898, 1840, 1775]. VLSI [2696]. VM [946]. Vocabularies [2256]. Vocabulary [1059, 417]. Vocabulary-Supported [1059]. VOD [120]. Voice [412, 12]. Voice-Enabled [412]. VoIPv6 [25]. Volcanic [1718]. Volume [2658, 1910, 1835, 1926, 1923, 1787, 1924, 1841]. Volume-Preserving [1787]. Volumetric [1039, 1878, 1886, 1871, 1876]. VOP [2584]. Voronoï [1802, 1655, 2705, 2696]. Vote [1149]. Voter [2018]. Voting [3035, 1112, 1024, 1142, 2592, 1161]. Voting-Based [3035]. Vowels [3009]. Voxel [1919, 1844, 1909]. Voxel-Based [1919, 1844, 1909]. VPN [836, 2534, 115]. VPRSM [1609]. VQ [2589]. VR [1872]. Vs [689, 1550, 2590, 3198, 2868]. VSS [983]. Vulnerabilities [2626]. Vulnerability [824].

Wait [2363, 2349]. Wait-Free [2363, 2349]. Wajsberg [1551]. Wakeup [2698]. Walker [3077]. Walking [2918]. Walks [2392]. WAM [282, 281, 280]. WAN [2535]. WAN/LAN [2535]. Want [1981, 2286]. WAP [415, 2463]. Wapplet [21]. Warehouse [1935, 1228, 2231, 1668]. Warehouses [2344]. Warp [645, 1924].

Warp-Edge [645]. Warped [1806]. Warps [1918]. Waste [2268]. Watchdog [806]. Water [3463, 3458, 2268]. Waterman [1087]. Watershed [1039, 1823]. Watersheds [1823]. Wave [2859]. Wavelength [38, 36]. Wavelength-Routed [36]. Wavelet [1714, 1050, 3450, 3428, 1012]. Way [2128, 2157, 2661, 491, 70, 833, 1542, 2905, 3406]. Wayfinding [1671]. Ways [2160]. WCDMA [96]. WCET [2003]. WDM [39, 40, 71]. Weak [2547, 741, 3092, 1585, 1587, 323, 2645, 550, 697, 1684, 564]. Weakest [1754, 2502]. Weaknesses [2156]. Wear [3449]. Wearable [21]. Weaver [2105]. Web [1362, 1637, 2751, 1205, 1531, 2222, 3347, 2419, 2749, 153, 1609, 1230, 1139, 1470, 2796, 1208, 3200, 1279, 641, 2404, 1971, 2783, 1358, 187, 177, 387, 1363, 270, 180, 1472, 184, 2844, 1961, 2195, 2750, 574, 2752, 1476, 2088, 643, 1604, 2479, 158, 1357, 2792, 758, 181, 2794, 1448, 1461, 2732, 175, 1204, 1463, 1666, 1359, 3382, 1937, 765, 2469, 1607, 1360, 2777, 2798, 768, 2241, 1231, 3185, 2327, 2470, 2456, 678, 1310, 152, 176, 2323, 121, 2477, 3391, 646, 3392, 1113, 185, 194, 3182, 397, 1608, 2224, 1644]. Web-Based [1139, 641, 180, 2479, 1448, 2732, 1666, 2477, 3391, 3392]. Web-ERP [2750]. Web-Oriented [3347, 1937]. WebODE [1457]. Website [3278]. Websites [1107, 2223]. Wedge [1845]. Weighbor [1071]. Weight [1725, 390, 2193, 857, 2069]. Weight-Biased [857]. Weighted [2978, 1657, 1785, 493, 561, 1008, 1921]. Weighting [711, 1954, 3444]. Weights [1557, 3168]. Well [294, 937, 2774]. Well-Founded [294, 937, 2774]. Were [2322]. Westphalia [2039]. WFMS [1369]. WFSTs [1266]. Wheelchair [407]. Where [1105, 2322, 531]. Whether [1403, 1404]. Which [1628, 3133]. while [1570]. White [1803, 1806]. Who [417]. Wiccap [2223]. Wide [3017, 1038, 1231]. Wide-Coverage [3017]. Wideband [1721]. Widening [540].

Width [2656, 3050, 933, 99]. wif [1174]. Will [3185]. Window [1652, 2476, 4]. Window-Based [2476]. Windows [150, 3072, 2986, 2612]. Winner [3035]. Winner-Takes-All [3035]. WIP [2498]. Wire [1879, 1799, 1861]. Wireless [2050, 413, 2681, 76, 2680, 3266, 1220, 64, 2466, 2067, 3256, 84, 3258, 93, 87, 2483, 2465, 1766, 91, 367, 2764, 2480, 2447, 90, 392, 2499, 3269, 3259, 3267, 168, 101, 2049]. Wise [2997, 3193, 1066]. within [1140, 1292, 1116, 861, 1664, 948, 1161, 2086]. Without [2292, 2343, 163, 1364, 1523]. Wizardry [451]. WML [2067]. Wolf [3106]. WOnDA [648]. Word [489, 1203, 2084, 2087, 371, 1652, 633, 2092, 3015, 3178, 1266, 1414, 2359]. Word-Level [2084]. Word-of-Mouth [371]. WordNet [3014]. Words [3417, 749]. Work [2361, 388, 697, 2730, 2726]. Workbench [1457]. Workcell [3033]. Workers [1351, 1349]. Workflow [2719, 3079, 2211, 1369, 2764, 2427]. Workflows [2715, 2713]. Working [2384, 1349]. Workplaces [1611]. Works [1613, 1277]. Workshop [3352, 2318, 659, 2320, 2319, 3256, 2317, 3346, 2316]. Workshops [1318]. World [448, 625, 680, 1001, 1383, 1384, 2897, 921, 1231]. Worlds [1207]. Worst [3424, 292]. Worst-Case [3424, 292]. Would [1981, 2286]. Wrapper [3182]. Wrapping [1701]. Write [2160]. Writer [2934]. Writing [441]. WUML [1357]. WWW [173, 84, 3197, 2097].

X [1029, 2844, 1907, 1906]. X-Compass [2844]. X-Ray [1907, 1906, 1029]. X2RMap [2432]. x86 [942]. XCS [892]. XDoC [1369]. XDoC-WFMS [1369]. XGuide [181]. XHTML [401]. XML [2734, 1935, 1938, 640, 1126, 1283, 286, 256, 1972, 68, 1367, 2245, 2838, 264, 2243, 2788, 2844, 1239, 1368, 2431, 257, 2430, 1939, 181, 1940, 141, 2432, 2444, 1936, 2398, 1946, 222, 2790, 1947, 1943, 1937, 1281, 258, 2445, 1944, 2329, 1945, 1942, 1948, 265, 1949, 1370, 930, 1627, 3230, 646, 157, 774, 2789]. XML-Based [1935, 68, 1939, 181, 3230]. XML-Driven [1281]. XML-Schema [2431]. XML-SQL [1945]. XML/SNMP [2329]. XOR [663]. XP [468, 428, 459, 461, 469, 447, 453, 463, 452, 425, 434, 423, 466, 446]. XPath [645, 1941]. XQuery [1361]. XSLT [1361, 1937, 646]. XTR [2153, 975].

Yao [559]. Yoix [180].

Z [618]. Z-log [618]. Zamin [2509]. Zero [1697]. Zeta [971]. zetoc [1252]. Ziv [1653]. Zurich [1174].

References

Yu:2002:I

[1] Cui Yu. 1. introduction. Lecture Notes in Computer Science, 2341: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                              springer.de/link/service/series/
                                              0558/bibs/2341/23410001.htm;   http:
                                              //link.springer.de/link/service/
                                              series/0558/papers/2341/23410001.
                                              pdf.

Yu:2002:HDI

[2] Cui Yu. 2. high-dimensional indexing. Lecture Notes in Computer Science, 2341:9–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                              //link.springer.de/link/service/
                                              series/0558/bibs/2341/23410009.
                                              htm;   http://link.springer.de/
                                              link/service/series/0558/papers/
                                              2341/23410009.pdf.

Yu:2002:IES

[3] Cui Yu. 3. indexing the edges — A simple and yet efficient approach to high-dimensional range search. Lecture Notes in Computer Science, 2341:37–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2341/23410037.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2341/23410037.pdf.

Yu:2002:PSW

[4] Cui Yu. 4. performance study of window queries. Lecture Notes in Computer Science, 2341:65–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2341/23410065.htm;   http://link.
springer.de/link/service/series/
0558/papers/2341/23410065.pdf.

Yu:2002:IRD

[5] Cui Yu. 5. indexing the relative distance — an efficient approach to KNN search. Lecture Notes in Computer Science, 2341:85–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2341/23410085.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2341/23410085.pdf.

Yu:2002:SRA

[6] Cui Yu. 6. similarity range and approximate KNN searches with iMin-Max. Lecture Notes in Computer Science, 2341:109–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2341/23410109.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2341/23410109.pdf.

Yu:2002:PSS

[7] Cui Yu. 7. performance study of similarity queries. Lecture Notes in Computer Science, 2341:123– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2341/23410123.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2341/23410123.
pdf.

Yu:2002:C

[8] Cui Yu. 8. conclusions. Lecture Notes in Computer Science, 2341: 141–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2341/23410141.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2341/23410141.
pdf.

Anonymous:2002:R

[9] Anonymous. References. Lecture Notes in Computer Science, 2341: 145–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2341/2341refe.pdf.

Koyama:2002:ABP

[10] Akio Koyama, Leonard Barolli, Zixue Cheng, and Norio Shiratori. An agentbased personalized distance learning system for delivering appropriate studying materials to learners. Lecture Notes in Computer Science, 2343:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430003.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430003.pdf.

Kim:2002:SRB

[11] Ik-Kyun Kim, Koohong Kang, Zyn-Oh Kim, Jong-Soo Jang, and Youngsik Baek. A simple rate-based congestion control scheme using an enhanced BECN in ATM networks. Lecture Notes in Computer Science, 2343:17–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430017.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430017.pdf.

Jang:2002:PAM

[12] Suk ki Jang, Hoon Kim, Jong-Dae Park, Sang-Sig Nam, and Kwang-Che Park. Performance analysis of media gateway for voice service. Lecture Notes in Computer Science, 2343:28–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430028.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2343/23430028.pdf.

Munemori:2002:OEP

[13] Jun Munemori, Tomohiro Muta, and Takashi Yoshino. E-ONIGOCO: Electronic playing tag support system for cooperative game. Lecture Notes in Computer Science, 2343: 40–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2343/23430040.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2343/23430040.
                                        pdf.

Chong:2002:AOP

[14] Ilyoung Chong and Sungchang Lee. Approximated overpassing period to virtual capacity of LSP in MPLS. Lecture Notes in Computer Science, 2343:50–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2343/23430050.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2343/23430050.pdf.

Koh:2002:ECT

[15] Seok Joo Koh, Juyoung Park, Eunsook Kim, and Shin-Gak Kang. Enhanced communications transport protocol for multicast transport. Lecture Notes in Computer Science, 2343:64–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2343/23430064.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430064.pdf.

Kang:2002:LAA

[16] Hyun Joo Kang, Seong Gon Choi, Jun Kyun Choi, and Kyou Ho Lee. A look-ahead algorithm for fast fault recovery in MPLS network using GSMP open signaling architecture. Lecture Notes in Computer Science, 2343:75–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430075.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430075.pdf.

Kim:2002:RAC

[17] Eunsook Kim, Seok Joo Koh, Juyoung Park, Shin-Gak Kang, and Jongwon Choe. A router assisting control tree configuration mechanism for reliable multicast. Lecture Notes in Computer Science, 2343:84–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430084.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430084.pdf.

Sato:2002:PMP

[18] Katsuhiko Sato and Michiaki Katsumoto. A proposal of multicast for personalized media stream delivery. Lecture Notes in Computer Science, 2343:94–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2343/23430094.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2343/23430094.pdf.

Park:2002:NDS

[19] Juyoung Park, Seok Joo Koh, Eunsook Kim, Shin-Gak Kang, and Dae Young Kim. A new delivery scheme for 1-to-N multicast applications. Lecture Notes in Computer Science, 2343:109–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2343/23430109.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2343/23430109.pdf.

Lee:2002:MRD

[20] Jaeyoung Lee, Heonkyu Park, and Woohyung Choi. Multicast routing debugger (MRD) — A system to monitor the status of multicast network. Lecture Notes in Computer Science, 2343:119–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2343/23430119.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2343/23430119.pdf.

Iwamoto:2002:WMA

[21] Takeshi Iwamoto, Nobuhiko Nishio, and Hideyuki Tokuda. Wapplet: a media access framework for wearable applications. Lecture Notes in Computer Science, 2343:127–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430127.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430127.pdf.

Cheon:2002:RMD

[22] Jung-Hoon Cheon, Jin-Woo Jung, and Hyun-Kook Kahng. Reliable multicast delivery based on local retransmission. Lecture Notes in Computer Science, 2343:138–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430138.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430138.pdf.

Shinomiya:2002:NNA

[23] Shunsuke Shinomiya, Yutaka Kidawara, Takeshi Sakurada, Seiji Tsuchiike, and Shin ichi Nakagawa. NADIA: Network accessible device on the Internet architecture. Lecture Notes in Computer Science, 2343:146–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430146.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430146.pdf.

Lee:2002:DCM

[24] Jae-Woo Lee, Jong-Ha Hong, and Doo-Kwon Baik. Database conceptual modeling for integrating E-A model and E-R model. Lecture Notes in Computer Science, 2343:157–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2343/23430157.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2343/23430157.pdf.

Kim:2002:DIV

[25] Chinchol Kim, Byounguk Choi, Yichul Kang, Keecheon Kim, and Sunyoung Han. Design and implementation of the VoIPv6 supporting the differentiated call processing. Lecture Notes in Computer Science, 2343:169–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2343/23430169.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2343/23430169.pdf.

Kitagata:2002:AOF

[26] Gen Kitagata, Takuo Suganuma, and Norio Shiratori. Application-oriented flow control in dynamic networking architecture. Lecture Notes in Computer Science, 2343:180–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2343/23430180.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2343/23430180.pdf.

Tajima:2002:OPS

[27] Shigeto Tajima, Masakazu Fujii, Nobuo Funabiki, Tokumi Yokohira, Kazufumi Tsunemura, and Teruo Higashino. An optimal path selection algorithm for static and mobile multicast routing problems. Lecture Notes in Computer Science, 2343:193–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430193.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430193.pdf.

Lee:2002:PEC

[28] Tsern-Huei Lee and Ying-Che Kuo. Performance evaluation of combined input output queued switch with finite input and output buffers. Lecture Notes in Computer Science, 2343:203–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430203.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430203.pdf.

Kawabe:2002:PPM

[29] Ryo Kawabe, Shingo Ata, and Masayuki Murata. Performance prediction methods for address lookup algorithms of IP routers. Lecture Notes in Computer Science, 2343: 215–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2343/23430215.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2343/23430215.
pdf.

Hong:2002:SOF

[30] Young-Ki Hong, Sungchang Lee, and Hyeong-Ho Lee. Shortest and oldest first non-interleaving packet switching scheduling algorithm. Lecture Notes in Computer Science, 2343:226–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2343/23430226.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2343/23430226.pdf.

Lee:2002:PAN

[31] Sungchang Lee and Ilyoung Chong. A packet-aware non-interleaving scheduling algorithm with multiple classes for input-queued switch. Lecture Notes in Computer Science, 2343: 237–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2343/23430237.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2343/23430237.
                                        pdf.

Yang:2002:DRP

[32] Mijeong Yang, Byoungchun Jeon, and Youkyoung Lee. A design of the IS-IS routing protocol for the ATM based MPLS system. Lecture Notes in Computer Science, 2343:248–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2343/23430248.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2343/23430248.pdf.

Suh:2002:NDS

[33] Chang jin Suh and Doo hyun Han. A new on-demand source verification method in multi-domain multicast networks. Lecture Notes in Computer Science, 2343:257–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430257.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430257.pdf.

Sheu:2002:BSA

[34] Shiann-Tsong Sheu, Bih-Hwang Lee, Hsuen-Wen Tseng, Chiu-Yun Ko, and Wei-Zhen Liang. A balanced scheduling algorithm for high-speed ATM switch. Lecture Notes in Computer Science, 2343:268–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430268.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430268.pdf.

Seok:2002:DCT

[35] Yongho Seok, Youngseok Lee, Yanghee Choi, and Changhoon Kim. Dynamic constrained traffic engineering for multicast routing. Lecture Notes in Computer Science, 2343: 278–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2343/23430278.htm;   http:

                                        //link.springer.de/link/service/
                                        series/0558/papers/2343/23430278.
                                        pdf.

Taira:2002:ELR

[36] Koji Taira, Yongbing Zhang, Hideaki Takagi, and Sajal K. Das. Efficient lightpath routing in wavelengthrouted optical networks. Lecture Notes in Computer Science, 2343: 291–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2343/23430291.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2343/23430291.
                                        pdf.

Lim:2002:EEP

[37] Huhnkuk Lim and Chang-Soo Park. EPSB (electronic partially shared buffering): a buffering scheme for asynchronous and variable length optical routing for the edge optical packet switch. Lecture Notes in Computer Science, 2343:305–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2343/23430305.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2343/23430305.pdf.

Chang:2002:MOC

[38] Chi-Yuan Chang and Sy-Yen Kuo. Multicasting optical crossconnects with fault tolerance mechanism and wavelength routing in all-optical networks. Lecture Notes in Computer Science, 2343:313–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430313.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430313.pdf.

Coelho:2002:OBL

[39] Rui M. F. Coelho and Mário M. Freire. Optical backbones with low connectivity for IP-over-WDM networks. Lecture Notes in Computer Science, 2343:327–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430327.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430327.pdf.

Jung:2002:SBP

[40] Kiyoung Jung and JunKyun Choi. A study of blocking probability in WDM network. Lecture Notes in Computer Science, 2343:337–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430337.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430337.pdf.

Jeong:2002:IAO

[41] Sangjin Jeong and Chan-Hyun Youn. Instability analysis for OBGP routing convergence in optical Internet. Lecture Notes in Computer Science, 2343:346–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2343/23430346.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2343/23430346.pdf.

Ohsaki:2002:MRT

[42] Hiroyuki Ohsaki, Mitsushige Morita, and Masayuki Murata. On modeling round-trip time dynamics of the Internet using system identification. Lecture Notes in Computer Science, 2343:359–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2343/23430359.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2343/23430359.pdf.

Kisimoto:2002:AIT

[43] Motohisa Kisimoto, Hiroyuki Ohsaki, and Masayuki Murata. Analyzing the impact of TCP connections variation on transient behavior of RED gateway. Lecture Notes in Computer Science, 2343:372–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2343/23430372.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2343/23430372.pdf.

Cha:2002:SDS

[44] Hojung Cha, Jiho Bang, and Rhan Ha. Static document scheduling with improved response time in HTTP/ 1.1. Lecture Notes in Computer Science, 2343:384–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430384.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430384.pdf.

Lee:2002:TVS

[45] Soo hyeong Lee, Byung G. Kim, and Yanghee Choi. TCP-Vegas slow start performance in large bandwidth delay network. Lecture Notes in Computer Science, 2343:394–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430394.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430394.pdf.

Kao:2002:PAG

[46] Ling-Chih Kao and Zsehong Tsai. Performance analysis of a generic GMPLS switching architecture under ON-OFF traffic sources. Lecture Notes in Computer Science, 2343: 407–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2343/23430407.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2343/23430407.
pdf.

Sheu:2002:PAM

[47] Shiann-Tsong Sheu and Jenhui Chen. Performance analysis of multicast MAC protocol for multichannel dual bus networks. Lecture Notes in Computer Science, 2343:419–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.de/link/service/
                                        series/0558/bibs/2343/23430419.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2343/23430419.pdf.

Kim:2002:NPM

[48] Byung-Chul Kim, Young-Soo Choi, and You-Ze Cho. A new packet marking strategy for explicit congestion notification in the Internet. Lecture Notes in Computer Science, 2343:431–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2343/23430431.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2343/23430431.pdf.

Lee:2002:IFR

[49] Soo hyeong Lee, Byung G. Kim, and Yanghee Choi. Improving the fairness and the response time of TCP-Vegas. Lecture Notes in Computer Science, 2343:443–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2343/23430443.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2343/23430443.pdf.

Wu:2002:RRA

[50] Jean-Lien C. Wu, Wei-Yeh Chen, and Hung-Huan Liu. Radio resource allocation in GSM/GPRS networks. Lecture Notes in Computer Science, 2343:457–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430457.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430457.pdf.

Koo:2002:NAR

[51] Jahon Koo, Kwangsue Chung, Hwasung Kim, and Hyukjoon Lee. A new active RED algorithm for congestion control in IP networks. Lecture Notes in Computer Science, 2343:469–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430469.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430469.pdf.

Lim:2002:FAC

[52] Ji-Young Lim and Ki-Joon Chae. Flow admission control for MPLS support of DiffServ. Lecture Notes in Computer Science, 2343:480–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430480.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430480.pdf.

Lee:2002:ABR

[53] So-Hyun Lee, Myoung-Kyoung Ji, Tae-Uk Choi, and Ki-Dong Chung. An ACK-based redundancy control algorithm in periodic FEC. Lecture Notes in Computer Science, 2343:494–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2343/23430494.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2343/23430494.pdf.

Park:2002:RTM

[54] Nam-Sup Park, Sang-Jun Nam, and Tai-Yun Kim. Real-time multimedia data transmission module based on Linux. Lecture Notes in Computer Science, 2343:504–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2343/23430504.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2343/23430504.pdf.

Niu:2002:ATE

[55] Ying Xia Niu, Choong Seon Hong, and Byung-Deok Chung. An adaptive TCP for enhancing the performance of TCP in mobile environments. Lecture Notes in Computer Science, 2343:516–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2343/23430516.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2343/23430516.pdf.

Lai:2002:NSP

[56] Yuan-Cheng Lai, Wei-Hsi Li, and Arthur Chang. A novel scheduler for the proportional delay differentiation model by considering packet transmission time. Lecture Notes in Computer Science, 2343:527–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430527.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430527.pdf.

Kim:2002:LBU

[57] Sung-Dae Kim, Hahnearl Jeon, Jai-Yong Lee, Yong-Joon Kim, and Hyung-Taek Kim. Load balancing using modified cost function for constraint-based routing over MPLS networks. Lecture Notes in Computer Science, 2343:539–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430539.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430539.pdf.

Jeong:2002:PAA

[58] Seong-Ho Jeong and John Copeland. Performance analysis of AAL2 trunking/switching for digital telephone switching systems. Lecture Notes in Computer Science, 2343:549– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2343/23430549.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2343/23430549.
pdf.

Kim:2002:SOC

[59] Do-Woo Kim and Min-Soo Jung. A study on the optimization of class file for Java card platform. Lecture Notes in Computer Science, 2343:563–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2343/23430563.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2343/23430563.pdf.

Jung:2002:DIS

[60] Jun-Young Jung and Min-Soo Jung. Design and implementation of smallsized Java Virtual Machine on Java platform Jini. Lecture Notes in Computer Science, 2343:571– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2343/23430571.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2343/23430571.
                                        pdf.

Lee:2002:DIH

[61] Jonghwan Lee, Yongho Kim, and Kyongsok Kim. Development of Internet home server to control information appliances remotely. Lecture Notes in Computer Science, 2343: 581–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2343/23430581.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2343/23430581.
                                        pdf.

Oh:2002:DHI

[62] YeunJoo Oh, KwangRoh Park, Kyung-Shik Lim, and KyoungRok Cho. Design of HG-iPhone service based on H.323 in the Home Gateway System. Lecture Notes in Computer Science, 2343:589–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430589.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430589.pdf.

Yaakob:2002:EAS

[63] Nor Diana Yaakob, Tanabe Hidehiko, and Umeda Hiroyuki. An efficient access scheme for variable length packets in slotted ALOHA channel. Lecture Notes in Computer Science, 2343:598–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430598.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430598.pdf.

Jang:2002:CAS

[64] Kil-Woong Jang and Ki-Jun Han. A channel assignment scheme for handoff in wireless mobile networks. Lecture Notes in Computer Science, 2343:609–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430609.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430609.pdf.

Shigeyasu:2002:RBD

[65] Tetsuya Shigeyasu, Hideyuki Ishinaka, and Hiroshi Matsuno. A relationship between densities and numbers of hidden stations of packet radio network. Lecture Notes in Computer Science, 2343:618–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430618.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430618.pdf.

Hong:2002:CBP

[66] Daniel Won-Kyu Hong, Choong Seon Hong, Yoo Jae Hyoung, Dong-Sik Yun, and Woo-Sung Kim. A CORBAbased performance management framework for multi-layer network environment. Lecture Notes in Computer Science, 2343:631–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430631.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430631.pdf.

Hamada:2002:USM

[67] Takayuki Hamada, Tatsuhiro Tsuchiya, Masahide Nakamura, and Tohru Kikuno. Using symbolic model checking to detect service interactions in telecommunication services. Lecture Notes in Computer Science, 2343:641–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430641.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430641.pdf.

Cha:2002:IXB

[68] Si-Ho Cha, Jae-Oh Lee, Young-Keun Choi, and Kook-Hyun Cho. Implementing an XML-based universal network management system in Java. Lecture Notes in Computer Science, 2343:652–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430652.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430652.pdf.

Kim:2002:IMP

[69] Sook-Yeon Kim, Myung-Eun Kim, Kiyoung Kim, and Jongsoo Jang. Information model for policy-based network security management. Lecture Notes in Computer Science, 2343: 662–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2343/23430662.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2343/23430662.
pdf.

Jeong:2002:DIO

[70] Jaehoon Jeong, Seungyun Lee, Yongjin Kim, and Yanghee Choi. Design and implementation of oneway IP performance measurement tool. Lecture Notes in Computer Science, 2343:673–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430673.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2343/23430673.pdf.

Kim:2002:AEE

[71] Young Ae Kim and Jun Kyun Choi. Analysis of the end-to-end protection and restoration algorithm in the IP over WDM network. Lecture Notes in Computer Science, 2343:687–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2343/23430687.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2343/23430687.pdf.

Choi:2002:DIT

[72] Taesang Choi, Seunghyun Yoon, Hyungseok Chung, Changhoon Kim, Jungsook Park, Bungjoon Lee, and Taesoo Jeong. Design and implementation of traffic engineering server for a large-scale MPLS-based IP network. Lecture Notes in Computer Science, 2343:699–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2343/23430699.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2343/23430699.pdf.

Choo:2002:MVL

[73] Hyunseung Choo, Hee Yong Youn, and Daewoo Chung. Modified virtual layer scheme for location management. Lecture Notes in Computer Science, 2343:712–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2343/23430712.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2343/23430712.pdf.

Anonymous:2002:AIa

[74] Anonymous. Author index. Lecture Notes in Computer Science, 2343: 725–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2343/2343auth.pdf.

Tsang:2002:MQB

[75] Tony Tsang. Modelling QoS based admission control in mobile multimedia systems using an extension of GSPN. Lecture Notes in Computer Science, 2344:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440003.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440003.pdf.

Choi:2002:CPE

[76] Yoonsuk Choi, Howon Jung, and Kyungshik Lim. A comparative performance evaluation of the transport layer protocols for wireless transaction processing in mobile Internet environments. Lecture Notes in Computer Science, 2344:11–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440011.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2344/23440011.pdf.

Sung:2002:MMM

[77] Su lyun Sung, Yongtae Shin, Jongwon Choe, and Kwangsue Chung. Mobile multicast method using the habitude of movement. Lecture Notes in Computer Science, 2344:22–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440022.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440022.pdf.

Funabiki:2002:PQS

[78] Nobuo Funabiki, Toru Nakanishi, Tokumi Yokohira, Shigeto Tajima, and Teruo Higashino. A proposal of a quasi-solution state evolution algorithm for channel assignment problems. Lecture Notes in Computer Science, 2344:32–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440032.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440032.pdf.

Goebel:2002:SAA

[79] Steffen Goebel, Sven Buchholz, Thomas Ziegert, and Alexander Schill. Software architecture for the adaptation of dialogs and contents to different devices. Lecture Notes in Computer Science, 2344:42–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

series/0558/bibs/2344/23440042.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440042.pdf.

Hong:2002:TTI

[80] Tae-Chul Hong, In-Ho Roh, and Young-Yong Kim. TCP throughput improvement for semi-soft & soft handoff. Lecture Notes in Computer Science, 2344:52–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440052.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440052.pdf.

Pack:2002:SQH

[81] Sangheon Pack and Yanghee Choi. Seamless QoS handling mechanism for macro and micro mobility. Lecture Notes in Computer Science, 2344: 62–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2344/23440062.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2344/23440062.
pdf.

Choi:2002:MML

[82] Seong Gon Choi, Hyun Joo Kang, and Jun Kyun Choi. Mobility management with low latency time by using the GSMP on the MPLS network. Lecture Notes in Computer Science, 2344:74–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2344/23440074.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440074.pdf.

Huh:2002:GBL

[83] Yungoo Huh and Cheeha Kim. Groupbased location management scheme in personal communication networks. Lecture Notes in Computer Science, 2344:81–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440081.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440081.pdf.

Konishi:2002:FWA

[84] Yousuke Konishi, Susumu Ishihara, and Tadanori Mizuno. A fast WWW access method for multiple wireless links shared by mobile computers. Lecture Notes in Computer Science, 2344:91–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440091.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440091.pdf.

Yu:2002:SBM

[85] Young ho Yu, Il dong Jung, and Kyongsok Kim. A study of broadcasting method using linear clustering to support location dependent query in mobile computing. Lecture Notes in Computer Science, 2344:104–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440104.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440104.pdf.

Sato:2002:RTC

[86] Fumiaki Sato and Tadanori Mizuno. Retransmission tree configuration for reliable multicast protocols. Lecture Notes in Computer Science, 2344: 113–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2344/23440113.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2344/23440113.
pdf.

Lee:2002:AAP

[87] ByungGil Lee, HyunGon Kim, and KilHoum Park. An AAA application protocol design and service for secure wireless Internet gateway roaming. Lecture Notes in Computer Science, 2344:123–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440123.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440123.pdf.

Jung:2002:FMD

[88] Jin-Woo Jung, Jung-Hoon Cheon, Sangwoo Nam, Youn-Joo Lee, and Hyun-Kook Kahng. Fast move detection over mobile IP. Lecture Notes in Computer Science, 2344: 136–??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2344/23440136.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2344/23440136.
                                        pdf.

Kong:2002:NLC

[89] Ki-Sik Kong, Joon-Min Gil, Youn-Hee Han, Ui-Sung Song, and Chong-Sun Hwang. A new location caching scheme employing forwarding pointers in PCS networks. Lecture Notes in Computer Science, 2344:143–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440143.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440143.pdf.

Shimada:2002:ELM

[90] Hideki Shimada, Shigeaki Tagashira, Tsuneo Nakanishi, and Akira Fukuda. Evaluation of a location management system for wireless communicating mobile computers. Lecture Notes in Computer Science, 2344: 154–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2344/23440154.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2344/23440154.
                                        pdf.

Park:2002:EEE

[91] Insoo Park and Yongjin Park. The enhancement of end-to-end TCP on wireless CDMA network to support interactive data service. Lecture Notes in Computer Science, 2344:165–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440165.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440165.pdf.

Qu:2002:QPAa

[92] Zhaowei Qu and Choong Seon Hong. A QoS provision architecture for mobile IPv6 using RSVP. Lecture Notes in Computer Science, 2344:176–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440176.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440176.pdf.

Lai:2002:PAW

[93] Yuan-Cheng Lai, Shuen-Chich Tsai, and Arthur Chang. Performance analysis of wireless TCP. Lecture Notes in Computer Science, 2344: 188–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2344/23440188.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2344/23440188.
pdf.

Lee:2002:ETA

[94] KyeongJin Lee, SeungYun Lee, YongJin Kim, ByongChul Kim, and JaeYong Lee. EM-TCP: An algorithm for mobile TCP performance improvement in a DSRC-based ITS. Lecture Notes in Computer Science, 2344:200–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440200.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440200.pdf.

Ma:2002:MPS

[95] Youngsik Ma, Donghyun Chae, Wonjong Noh, Jiyoung Lee, Yeonjoong Kim, and Sunshin An. A multi-path support for mobile IP with the hierarchical architecture. Lecture Notes in Computer Science, 2344:212–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440212.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440212.pdf.

Kang:2002:PET

[96] Ji-Eun Kang, Dong-Min Kim, Yong-Chul Shin, Hee-Yoon Park, and Jai-Yong Lee. Performance evaluation of TCP over WCDMA RLC. Lecture Notes in Computer Science, 2344:221–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440221.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440221.pdf.

Park:2002:UDL

[97] Kwang-Youn Park, Sang-Jun Park, In-Sook Cho, and Byung-Gi Kim. User defined location area for registration and paging. Lecture Notes in Computer Science, 2344:229–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440229.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440229.pdf.

Ohzahata:2002:FAM

[98] Satoshi Ohzahata, Shigetomo Kimura, and Yoshihiko Ebihara. A fast authentication method for secure and seamless handoff. Lecture Notes in Computer Science, 2344:243–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440243.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440243.pdf.

Oh:2002:OMC

[99] Jong Kap Oh, Sung Pil Kim, Myoung Jin Kim, Ho Seong Ahn, and Seung Moon Ryu. Orthogonal multicode CDMA systems with constant amplitude transmission based on level clipping and pulse width modulation. Lecture Notes in Computer Science, 2344:253–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440253.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2344/23440253.pdf.

Kim:2002:SCS

[100] Sung Pil Kim, Jong Kap Oh, Myoung Jin Kim, Ho Seong Ahn, and Seung Moon Ryu. On the selection of code set for binary multi-code CDMA system. Lecture Notes in Computer Science, 2344:264–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440264.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440264.pdf.

Yamamoto:2002:PEI

[101] Takayuki Yamamoto, Masashi Sugano, Masayuki Murata, Takaaki Hatauchi, and Yohei Hosooka. Performance evaluation and improvement of an ad hoc wireless network. Lecture Notes in Computer Science, 2344:276–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440276.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440276.pdf.

Suzuki:2002:RMF

[102] Yoshinari Suzuki, Susumu Ishihara, and Tadanori Mizuno. Reassignment of mediation function between hosts on mobile ad hoc network. Lecture Notes in Computer Science, 2344:288–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440288.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440288.pdf.

Lee:2002:ARP

[103] Jong-Hoon Lee, Joo-Hwan Seo, Jung-Seok Lee, Won Ryu, and Ki-Jun Han. An asymmetric routing protocol using detour in mobile ad hoc networks. Lecture Notes in Computer Science, 2344:300–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440300.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440300.pdf.

Kang:2002:MSM

[104] Hyun jeong Kang and Mee jeong Lee. A multi-source multicast routing protocol for ad-hoc networks. Lecture Notes in Computer Science, 2344:309–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440309.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440309.pdf.

Ahn:2002:ILM

[105] Sanghyun Ahn, Yujin Lim, Wonjun Lee, and EunyongHa. Infrastructureless multicast protocol in mobile ad-hoc networks. Lecture Notes in Computer Science, 2344:321–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440321.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440321.pdf.

Jang:2002:BDA

[106] Yeong M. Jang. Buffer dimensioning of ATM satellite link using probabilistic burstiness curve. Lecture Notes in Computer Science, 2344:333–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440333.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440333.pdf.

Park:2002:CSM

[107] Seong-Bok Park, Kwang-Eog Lee, Young-Kyun Choi, and Soon-Ja Kim. Construction of sequences for multigroup FH-FDMA in DRT system. Lecture Notes in Computer Science, 2344:342–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440342.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440342.pdf.

Kitaguchi:2002:DHA

[108] Yoshiaki Kitaguchi, Haruo Okazawa, Syunsuke Shinomiya, Yutaka Kidawara, Kaisuya Hakozaki, and Shin ichi Nakagawa. Development of a high-accurate time server for measurements of the Internet. Lecture Notes in Computer Science, 2344:351–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440351.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440351.pdf.

Kyun:2002:MSP

[109] Ng Chee Kyun, Ashraf Gasim Elsid Abdalla, Nor Kamariah Noordin, Sabira Khatun, Borhanuddin Mohd Ali, and Ratna Kalos Zakiah Sahbuddin. Modelling and simulation of phased array antenna for LEO satellite tracking. Lecture Notes in Computer Science, 2344:359–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440359.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440359.pdf.

Ku:2002:IIB

[110] Wei-Chi Ku. An improved ID-based authentication and key distribution protocol. Lecture Notes in Computer Science, 2344:375–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440375.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440375.pdf.

Park:2002:SSA

[111] Wonjoo Park, Jaehoon Nah, and Sungwon Sohn. A study of security association management oriented to IP security. Lecture Notes in Computer Science, 2344:381–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440381.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440381.pdf.

Kim:2002:DNS

[112] Byoung-Koo Kim, Jong-Su Jang, and Tai M. Chung. Design of network security control system for cooperative intrusion detection. Lecture Notes in Computer Science, 2344:389–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440389.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440389.pdf.

Kim:2002:PBI

[113] Jinoh Kim, Kiyoung Kim, and Jongsoo Jang. Policy-based intrusion detection and automated response mechanism. Lecture Notes in Computer Science, 2344:399–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440399.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440399.pdf.

Kim:2002:PSS

[114] Nam-Yeun Kim, Hyun-Sung Kim, Won-Ho Lee, and Kee-Young Yoo. A power-sum systolic architecture in GF(2^{m}). Lecture Notes in Computer Science, 2344:409–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440409.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440409.pdf.

Park:2002:DAH

[115] Chan-Woo Park, Jaehoon Nah, JongHyun Jang, and Chimoon Han. Design and analysis of hierarchical VPN based on SCG. Lecture Notes in Computer Science, 2344: 418–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2344/23440418.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2344/23440418.
pdf.

Lee:2002:DAR

[116] HyungHyo Lee and BongNam Noh. Design and analysis of role-based security model in SNMPv3 for policy-based security management. Lecture Notes in Computer Science, 2344:430–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440430.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440430.pdf.

Asami:2002:TSP

[117] Tohru Asami, Takahiro Kikuchi, Kenji Rikitake, Hiroshi Nagata, Tatsuaki Hamai, and Yoshinori Hatori. A taxonomy of spam and a protection method for enterprise networks. Lecture Notes in Computer Science, 2344:442–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440442.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440442.pdf.

Kim:2002:SPD

[118] Geon Woo Kim, So Hee Park, Jae Hoon Nah, and Sung Won Sohn. Security policy deployment in IPsec. Lecture Notes in Computer Science, 2344:453–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440453.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440453.pdf.

Hashimoto:2002:MAB

[119] Koji Hashimoto, Yoshitaka Shibata, and Norio Shiratori. Mobile agentbased adaptive multimedia communication. Lecture Notes in Computer Science, 2344:467–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440467.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440467.pdf.

Lee:2002:AMT

[120] Seung-Won Lee, Ki-Wan Kim, Jung-Min Choi, and Ki-Dong Chung. An alternative multicast transmission scheme for VCR operations in a large VOD system. Lecture Notes in Computer Science, 2344: 479–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2344/23440479.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2344/23440479.
pdf.

Tavanapong:2002:EIQ

[121] Wallapak Tavanapong, Wenlu Li, and Ki-Hwan Kim. Enhancing information quality for Web pages. Lecture Notes in Computer Science, 2344:491–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440491.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440491.pdf.

Park:2002:RPG

[122] Seongho Park, Yongwoon Park, Hyeonok Hong, and Kidong Chung. A replacement policy with growth-curve on the continuous media proxy cache (CMPC). Lecture Notes in Computer Science, 2344:501–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440501.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440501.pdf.

Lee:2002:HRB

[123] ByungGil Lee and KilHoum Park. Hierarchy and region based sourcechannel combined coder for image transmission over mobile links. Lecture Notes in Computer Science, 2344:512–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440512.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440512.pdf.

Komine:2002:DHP

[124] Takahiro Komine, Akihiko Machizawa, Shin ichi Nakagawa, Fumito Kubota, and Yasuo Tan. Development of “High Presence” video communication system — trial experiment of the next generation real-time remote lecture — . Lecture Notes in Computer Science, 2344:521–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440521.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440521.pdf.

Song:2002:STF

[125] Byunghun Song, Kwangsue Chung, and Yongtae Shin. SRTP: TCPfriendly congestion control for multimedia streaming. Lecture Notes in Computer Science, 2344:529– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2344/23440529.htm;   http:

//link.springer.de/link/service/
series/0558/papers/2344/23440529.
pdf.

Yamada:2002:MFT

[126] Makoto Yamada, Takanori Mori, Atsushi Fukada, Akio Nakata, and Teruo Higashino. A method for functional testing of media synchronization protocols. Lecture Notes in Computer Science, 2344:539–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440539.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440539.pdf.

Sureswaran:2002:MVO

[127] R. Sureswaran and O. Abouabadalla. Measurements to validate optimised bandwidth usage by the distributed network entities architecture for multimedia conferencing. Lecture Notes in Computer Science, 2344: 551–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2344/23440551.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2344/23440551.
pdf.

Ahn:2002:CER

[128] JinHo Ahn, Sung-Gi Min, and Chong-Sun Hwang. Consistent and efficient recovery for causal message logging. Lecture Notes in Computer Science, 2344:565–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440565.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440565.pdf.

Dong:2002:SBD

[129] Yingfei Dong, Zhi-Li Zhang, and Y. Thomas Hou. Server-based dynamic server selection algorithms. Lecture Notes in Computer Science, 2344:575–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440575.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440575.pdf.

Liesenborgs:2002:DVE

[130] Jori Liesenborgs, Peter Quax, Wim Lamotte, and Frank Van Reeth. Designing a virtual environment for large audiences. Lecture Notes in Computer Science, 2344:585–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440585.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440585.pdf.

Choi:2002:OCM

[131] Eun Hye Choi, Tatsuhiro Tsuchiya, and Tohru Kikuno. Optimal k-coteries that maximize availability in general networks. Lecture Notes in Computer Science, 2344:596–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2344/23440596.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440596.pdf.

Lee:2002:FBA

[132] Seok-Hyoung Lee and Hoon Choi. The fast bully algorithm: For electing a coordinator process in distributed systems. Lecture Notes in Computer Science, 2344:609–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440609.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440609.pdf.

Ryu:2002:ESE

[133] Ki-Yeol Ryu and Jung-Tae Lee. An enhancement of SIENA event routing algorithms. Lecture Notes in Computer Science, 2344:623–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440623.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440623.pdf.

Katoh:2002:DBM

[134] Takashi Katoh, Tetsuo Kinoshita, and Norio Shiratori. Dynamic behavior of multiagents with subjective cooperative relations. Lecture Notes in Computer Science, 2344: 634–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2344/23440634.htm;   http:

                                        //link.springer.de/link/service/
                                        series/0558/papers/2344/23440634.
                                        pdf.

Han:2002:FMS

[135] Moon Seog Han, Sang Yun Park, and Young Ik Eom. A file migration scheme for reducing file access latency in mobile computing environments. Lecture Notes in Computer Science, 2344:644–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440644.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440644.pdf.

Sato:2002:FIR

[136] Nobuyoshi Sato, Minoru Uehara, Yoshifumi Sakai, and Hideki Mori. Fresh information retrieval using cooperative meta search engines. Lecture Notes in Computer Science, 2344:656–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440656.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440656.pdf.

Jeong:2002:AAM

[137] Dong Won Jeong and Doo-Kwon Baik. AMC: An adaptive mobile computing model for dynamic resource management in distributed computing environments. Lecture Notes in Computer Science, 2344:671–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440671.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440671.pdf.

Kang:2002:HFT

[138] YunHee Kang, HoSang Ham, and ChongSun Hwang. A hybrid faulttolerant scheme based on checkpointing in MASs. Lecture Notes in Computer Science, 2344:679–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440679.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440679.pdf.

Lee:2002:FCP

[139] Inseon Lee and Heon Y. Yeom. A fast commit protocol for distributed main memory database systems. Lecture Notes in Computer Science, 2344:691–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440691.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440691.pdf.

Komiya:2002:MAM

[140] Takao Komiya, Hiroyuki Ohshida, and Makoto Takizawa. Mobile agent model for transaction processing in distributed database systems. Lecture Notes in Computer Science, 2344:703–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440703.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440703.pdf.

Kim:2002:DMA

[141] Sung Wan Kim, Pan Seop Shin, Youn Hee Kim, Jaeho Lee, and Hae Chull Lim. A data model and algebra for document-centric XML document. Lecture Notes in Computer Science, 2344:714–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440714.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440714.pdf.

Arai:2002:GPD

[142] Keijirou Arai, Katsuya Tanaka, and Makoto Takizawa. Group protocol for delivering requests to replicas. Lecture Notes in Computer Science, 2344:724–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440724.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440724.pdf.

Woo:2002:DLA

[143] Namyoon Woo and Heon Y. Yeom. k-depth look-ahead task scheduling in network of heterogeneous processors. Lecture Notes in Computer Science, 2344:736–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440736.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440736.pdf.

Yoon:2002:RES

[144] Eun young Yoon and Yong-Ik Yoon. Reflective event service middleware for distributed component-based applications. Lecture Notes in Computer Science, 2344:746–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440746.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440746.pdf.

Park:2002:ASJ

[145] Jung Gyu Park and Myong-Soon Park. Automatic specialization of Java programs in the distributed environment. Lecture Notes in Computer Science, 2344:757–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440757.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440757.pdf.

Lee:2002:JBE

[146] Jung Hwan Lee and Choong Seon Hong. A JMX based efficient mobile agents platform architecture for network management. Lecture Notes in Computer Science, 2344: 768–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2344/23440768.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2344/23440768.
                                        pdf.

Baik:2002:ECC

[147] MaengSoon Baik, JinGon Shon, Kibom Kim, JinHo Ahn, and Chong-Sun Hwang. An efficient coordinated checkpointing scheme based on PWD model. Lecture Notes in Computer Science, 2344:780–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2344/23440780.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2344/23440780.pdf.

Lee:2002:EDS

[148] Dongman Lee, Hyewon Shin, Soon J. Hyun, Younghee Lee, Myoungchurl Kim, and Hee Yong Youn. Extensions to DNS for supporting internationalized domain names. Lecture Notes in Computer Science, 2344: 791–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2344/23440791.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2344/23440791.
                                        pdf.

Uchiya:2002:AAR

[149] Takahiro Uchiya, Takashi Katoh, Takuo Suganuma, Tetsuo Kinoshita, and Norio Shiratori. An architecture of agent repository for adaptive multiagent system. Lecture Notes in Computer Science, 2344: 802–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2344/23440802.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2344/23440802.
pdf.

Choi:2002:RRS

[150] Jin Seek Choi, Bum Sik Bae, Hyeong Ho Lee, and Hyeung Sub Lee. Round-robin scheduling algorithm with multiple distributed windows. Lecture Notes in Computer Science, 2344:814–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2344/23440814.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2344/23440814.pdf.

Anonymous:2002:AIb

[151] Anonymous. Author index. Lecture Notes in Computer Science, 2344: 821–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2344/2344auth.pdf.

Schimkat:2002:LHW

[152] Ralf-Dieter Schimkat, Wolfgang Küchlin, and Frank Nestel. Living hypertext — Web retrieval techniques for traditional database-centric information. Lecture Notes in Computer Science, 2346:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2346/23460001.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2346/23460001.pdf.

Heyer:2002:AAL

[153] Gerhard Heyer, Uwe Quasthoff, and Christian Wolff. Automatic analysis of large text corpora — A contribution to structuring WEB communities. Lecture Notes in Computer Science, 2346:15–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2346/23460015.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2346/23460015.pdf.

Ashrafi:2002:DMA

[154] Mafruz Zaman Ashrafi, David Taniar, and Kate Smith. A data mining architecture for distributed environments. Lecture Notes in Computer Science, 2346:27–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2346/23460027.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2346/23460027.pdf.

Shen:2002:CHR

[155] Haifeng Shen and Chengzheng Sun. Collaborative highlighting for realtime group editors. Lecture Notes in Computer Science, 2346:39–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.

de/link/service/series/0558/bibs/
2346/23460039.htm;   http://link.
springer.de/link/service/series/
0558/papers/2346/23460039.pdf.

Fengler:2002:EME

[156] Olga Fengler, Wolfgang Fengler, and Veselka Duridanova. Extending the modeling efficiency of the UML activity diagram for the design of distributed systems. Lecture Notes in Computer Science, 2346:51–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2346/23460051.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2346/23460051.pdf.

Varela:2002:XKB

[157] Leonilde Rocha Varela, Joaquim Nunes Aparício, and Sílvio Carmo Silva. An XML knowledge base system for scheduling problems. Lecture Notes in Computer Science, 2346: 63–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2346/23460063.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2346/23460063.
pdf.

Janssen:2002:CML

[158] Jeannette C. M. Janssen and Nauzer Kalyaniwalla. Compressibility as a measure of local coherence in Web graphs. Lecture Notes in Computer Science, 2346:75–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2346/23460075.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2346/23460075.pdf.

Vukadinovic:2002:SSI

[159] Danica Vukadinović, Polly Huang, and Thomas Erlebach. On the spectrum and structure of Internet topology graphs. Lecture Notes in Computer Science, 2346:83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2346/23460083.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2346/23460083.pdf.

An:2002:CCG

[160] Yuan An, Jeannette C. M. Janssen, and Evangelos E. Milios. Characterizing the citation graph as a self-organizing networked information space. Lecture Notes in Computer Science, 2346:97–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2346/23460097.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2346/23460097.pdf.

Erfurth:2002:CMD

[161] Christian Erfurth and Wilhelm Rossak. Characterization and management of dynamical behavior in a system with mobile components. Lecture Notes in Computer Science, 2346: 109–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2346/23460109.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2346/23460109.
pdf.

Rodriguez:2002:EIU

[162] Carlos Gregorio Rodríguez and Pedro Palao Gostanza. Ergotracer: An Internet user behaviour tracer. Lecture Notes in Computer Science, 2346:121–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2346/23460121.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2346/23460121.pdf.

Grimm:2002:FPP

[163] Rüdiger Grimm and Jürgen Nützel. A friendly peer-to-peer file sharing system with profit but without copy protection. Lecture Notes in Computer Science, 2346:133–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2346/23460133.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2346/23460133.pdf.

Wulff:2002:DLS

[164] Markus Wulff. A decentral library for scientific articles. Lecture Notes in Computer Science, 2346: 143–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/

                                        0558/bibs/2346/23460143.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2346/23460143.
                                        pdf.

Yuen:2002:ODC

[165] Saiho Yuen, Peter Kropf, and Gilbert Babin. Optimisation of distributed communities using cooperative strategies. Lecture Notes in Computer Science, 2346:153–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2346/23460153.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2346/23460153.pdf.

Eichler:2002:PIN

[166] Gerald Eichler, Anne Thomas, and Ralf Widera. Proven IP network services: From end-user to router and vice versa. Lecture Notes in Computer Science, 2346:169–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2346/23460169.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2346/23460169.pdf.

Roth:2002:MSR

[167] Jörg Roth. Mobility support for replicated real-time applications. Lecture Notes in Computer Science, 2346:181–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2346/23460181.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2346/23460181.pdf.

Won:2002:ESV

[168] Youjip Won and Bowie Shim. Empirical study of VBR traffic smoothing in wireless environment. Lecture Notes in Computer Science, 2346: 193–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2346/23460193.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2346/23460193.
pdf.

Ko:2002:CRB

[169] K. W. Ko, I. T. Kim, S. B. Yoo, K. C. Kim, Y. S. Lee, S. D. Kim, and J. S. Kim. Compiling rule-based agents for distributed databases. Lecture Notes in Computer Science, 2346:205–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2346/23460205.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2346/23460205.pdf.

Lopez:2002:FFB

[170] Natalia López, Manuel Núñez, Ismael Rodríguez, and Fernando Rubio. A formal framework for E-barter based on microeconomic theory and process algebras. Lecture Notes in Computer Science, 2346:217–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2346/23460217.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2346/23460217.pdf.

Lechner:2002:PPB

[171] Ulrike Lechner. Peer-to-peer beyond file sharing. Lecture Notes in Computer Science, 2346:229–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2346/23460229.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2346/23460229.pdf.

Anonymous:2002:AIc

[172] Anonymous. Author index. Lecture Notes in Computer Science, 2346: 251–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2346/2346pref.pdf.

Balamash:2002:PMM

[173] Abdullah Balamash and Marwan Krunz. A parsimonious multifractal model for WWW traffic. Lecture Notes in Computer Science, 2376:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2376/23760001.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2376/23760001.pdf.

Ishibashi:2002:CTS

[174] Keisuke Ishibashi, Masaki Aida, and Makoto Imase. Characteristics of temporal and spatial locality of Internet access patterns. Lecture Notes in Computer Science, 2376:15–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2376/23760015.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2376/23760015.pdf.

Lancellotti:2002:SAC

[175] Riccardo Lancellotti, Bruno Ciciani, and Michele Colajanni. A scalable architecture for cooperative Web caching. Lecture Notes in Computer Science, 2376:29–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2376/23760029.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2376/23760029.pdf.

Segura-Devillechaise:2002:CWS

[176] Marc Ségura-Devillechaise and Jean-Marc Menaud. Caching Web services: Aspect orientation to the rescue. Lecture Notes in Computer Science, 2376:42–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2376/23760042.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2376/23760042.pdf.

Conti:2002:RWS

[177] Marco Conti, Enrico Gregori, and Willy Lapenna. Replicated Web services: a comparative analysis of client-based content delivery policies. Lecture Notes in Computer Science, 2376:53–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2376/23760053.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2376/23760053.pdf.

Wierzbicki:2002:ICL

[178] Adam Wierzbicki. Internet cache location and design of content delivery networks. Lecture Notes in Computer Science, 2376:69–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2376/23760069.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2376/23760069.pdf.

Shim:2002:SPA

[179] Junho Shim, Taehee Lee, and Sang goo Lee. A server placement algorithm conscious of communication delays and relocation costs. Lecture Notes in Computer Science, 2376:83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2376/23760083.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2376/23760083.pdf.

Drechsler:2002:YSL

[180] Richard L. Drechsler and John M. Mocenigo. The Yoix scripting language as a tool for building Web-based systems. Lecture Notes in Computer Science, 2376:90–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2376/23760090.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2376/23760090.pdf.

Kerer:2002:XPG

[181] Clemens Kerer, Engin Kirda, and Christopher Krügel. XGuide — A practical guide to XML-based Web engineering. Lecture Notes in Computer Science, 2376:104–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2376/23760104.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2376/23760104.pdf.

Zirpins:2002:PBO

[182] Christian Zirpins, Kevin Schütt, and Giacomo Piccinelli. Process-based optimisation of data exchange for B2B interaction. Lecture Notes in Computer Science, 2376:118–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2376/23760118.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2376/23760118.pdf.

Riche:2002:CSP

[183] Stéphanie Riché, Gavin Brebner, and Mickey Gittler. Client-side profile storage. Lecture Notes in Computer Science, 2376:127–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2376/23760127.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2376/23760127.pdf.

Foglia:2002:BPT

[184] Pierfrancesco Foglia, Roberto Giorgi, and Cosimo Antonio Prete. Boosting the performance of three-tier Web servers deploying SMP architecture. Lecture Notes in Computer Science, 2376:134–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2376/23760134.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2376/23760134.pdf.

Voigt:2002:OBP

[185] Thiemo Voigt. Overload behaviour and protection of event-driven Web servers. Lecture Notes in Computer Science, 2376:147–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2376/23760147.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2376/23760147.pdf.

Jun:2002:SBP

[186] Kyungkoo Jun, Dan C. Marinescu, Yongchang Ji, and Gabriela M. Marinescu. Scalable benchmarking and performance monitoring: a coordination and mobility perspective. Lecture Notes in Computer Science, 2376:158–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2376/23760158.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2376/23760158.pdf.

Chen:2002:TSB

[187] J. Chen and S. Chovanec. Towards specification-based Web testing. Lecture Notes in Computer Science, 2376:165–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2376/23760165.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2376/23760165.pdf.

Schollmeier:2002:RMA

[188] Rüdiger Schollmeier, Ingo Gruber, and Michael Finkenzeller. Routing in mobile ad-hoc and peerto-peer networks A comparison. Lecture Notes in Computer Science, 2376:172–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2376/23760172.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2376/23760172.pdf.

Klein:2002:MLC

[189] Michael Klein and Birgitta König-Ries. Multi-layer clusters in ad-hoc networks — an approach to service discovery. Lecture Notes in Computer Science, 2376:187–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2376/23760187.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2376/23760187.pdf.

Joseph:2002:NSR

[190] Sam Joseph. NeuroGrid: Semantically routing queries in peer-to-peer networks. Lecture Notes in Computer Science, 2376:202–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2376/23760202.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2376/23760202.pdf.

Moreton:2002:SMN

[191] Tim D. Moreton, Ian A. Pratt, and Timothy L. Harris. Storage, mutability and naming in pasta. Lecture Notes in Computer Science, 2376:215–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2376/23760215.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2376/23760215.pdf.

Cuenca-Acuna:2002:TBC

[192] Francisco Matias Cuenca-Acuna and Thu D. Nguyen. Text-based content search and retrieval in adhoc P2P communities. Lecture Notes in Computer Science, 2376: 220–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2376/23760220.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2376/23760220.
pdf.

Balzarotti:2002:FCS

[193] Davide Balzarotti, Carlo Ghezzi, and Mattia Monga. Freeing cooperation from servers tyranny. Lecture Notes in Computer Science, 2376: 235–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2376/23760235.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2376/23760235.
pdf.

Wang:2002:BPB

[194] XiaoYu Wang, WeeSiong Ng, and BengChin Ooi. BuddyWeb: a P2Pbased collaborative Web caching system. Lecture Notes in Computer Science, 2376:247–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2376/23760247.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2376/23760247.pdf.

Unterkircher:2002:RUS

[195] Florian Unterkircher and Michael Welzl. Reliable unicast streaming with multiple peers in IP networks. Lecture Notes in Computer Science, 2376:252–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2376/23760252.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2376/23760252.pdf.

Karp:2002:CUP

[196] Alan Karp and Vana Kalogeraki. The client utility as a peer-to-peer system. Lecture Notes in Computer Science, 2376:260–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2376/23760260.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2376/23760260.pdf.

Oliveira:2002:PPP

[197] Licínio Oliveira, Luís Lopes, and Fernando Silva. P: Parallel peer to peer. Lecture Notes in Computer Science, 2376:274–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2376/23760274.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2376/23760274.pdf.

Evans:2002:PPP

[198] Huw Evans and Peter Dickman. Peer-to-peer programming with teaq. Lecture Notes in Computer Science, 2376:289–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2376/23760289.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2376/23760289.pdf.

Busi:2002:TDD

[199] Nadia Busi, Cristian Manfredini, Alberto Montresor, and Gianluigi Zavattaro. Towards a data-driven coordination infrastructure for peer-to-peer systems. Lecture Notes in Computer Science, 2376:295–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2376/23760295.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2376/23760295.pdf.

Montresor:2002:TAR

[200] Alberto Montresor, Hein Meling, and Özalp Babaoğlu. Towards adaptive, resilient and self-organizing peer-to-peer systems. Lecture Notes in Computer Science, 2376:300–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2376/23760300.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2376/23760300.pdf.

Keyani:2002:PPD

[201] Pedram Keyani, Brian Larson, and Muthukumar Senthil. Peer pressure: Distributed recovery from attacks in peer-to-peer systems. Lecture Notes in Computer Science, 2376:306–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2376/23760306.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2376/23760306.pdf.

Cornelli:2002:IRA

[202] Fabrizio Cornelli, Ernesto Damiani, Sabrina De Capitani di Vimercati, Stefano Paraboschi, and Pierangela Samarati. Implementing a reputationaware gnutella servent. Lecture Notes in Computer Science, 2376: 321–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2376/23760321.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2376/23760321.
                                        pdf.

Roscoe:2002:TBC

[203] Timothy Roscoe and Steven Hand. Transaction-based charging in mnemosyne: a peer-to-peer steganographic storage system. Lecture Notes in Computer Science, 2376:335–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2376/23760335.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2376/23760335.pdf.

Anonymous:2002:AId

[204] Anonymous. Author index. Lecture Notes in Computer Science, 2376: 351–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2376/2376auth.pdf.

El-Mejbri:2002:RPB

[205] El-Fathi El-Mejbri, Hans Grabowski, Harald Kunze, Ralf-Stefan Lossack, and Arno Michelis. 3D reconstruction of paper based assembly drawings: State of the art and approach. Lecture Notes in Computer Science, 2390:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2390/23900001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2390/23900001.pdf.

Grabowski:2002:ILL

[206] Hans Grabowski, Harald Kunze, Ralf Lossack, and Arno Michelis. Interpretation of low-level CAD data for knowledge extraction in early design stages. Lecture Notes in Computer Science, 2390:13–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2390/23900013.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2390/23900013.pdf.

Byun:2002:KBP

[207] Yungcheol Byun, Joongbae Kim, Yeongwoo Choi, Gyeonghwan Kim, and Yillbyung Lee. Knowledgebased partial matching: An efficient form classification method. Lecture Notes in Computer Science, 2390:25–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2390/23900025.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2390/23900025.pdf.

Nishiwaki:2002:RFE

[208] Daisuke Nishiwaki, Masato Hayashi, and Atsushi Sato. Robust frame extraction and removal for processing form documents. Lecture Notes in Computer Science, 2390:36–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2390/23900036.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2390/23900036.pdf.

Lopresti:2002:IGT

[209] Daniel Lopresti and George Nagy. Issues in ground-truthing graphic documents. Lecture Notes in Computer Science, 2390:46–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2390/23900046.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2390/23900046.pdf.

Wenyin:2002:SBU

[210] Liu Wenyin, Xiangyu Jin, and Zhengxing Sun. Sketch-based user interface for inputting graphic objects on small screen devices. Lecture Notes in Computer Science, 2390: 67–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2390/23900067.htm;   http:
                                        //link.springer.de/link/service/

series/0558/papers/2390/23900067.
pdf.

Pimentel:2002:EET

[211] César F. Pimentel, Manuel J. da Fonseca, and Joaquim A. Jorge. Experimental evaluation of a trainable scribble recognizer for calligraphic interfaces. Lecture Notes in Computer Science, 2390:81–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2390/23900081.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2390/23900081.pdf.

Blostein:2002:UIL

[212] Dorothea Blostein, Ed Lank, Arlis Rose, and Richard Zanibbi. User interfaces for on-line diagram recognition. Lecture Notes in Computer Science, 2390:92–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2390/23900092.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2390/23900092.pdf.

Llados:2002:SRC

[213] Josep Lladós, Ernest Valveny, Gemma Sánchez, and Enric Martí. Symbol recognition: Current advances and perspectives. Lecture Notes in Computer Science, 2390:104–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2390/23900104.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2390/23900104.pdf.

Sanchez:2002:ECG

[214] Gemma Sánchez, Josep Lladós, and Karl Tombre. An error-correction graph grammar to recognize texture symbols. Lecture Notes in Computer Science, 2390:128–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2390/23900128.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2390/23900128.pdf.

Saund:2002:POF

[215] Eric Saund, James Mahoney, David Fleet, and Daniel Larner. Perceptual organization as a foundation for graphics recognition. Lecture Notes in Computer Science, 2390: 139–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2390/23900139.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2390/23900139.
                                        pdf.

Thomson:2002:EPG

[216] Robert C. Thomson and Rupert Brooks. Exploiting perceptual grouping for map analysis, understanding and generalization: The case of road and river networks. Lecture Notes in Computer Science, 2390:148–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2390/23900148.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2390/23900148.pdf.

Suzuki:2002:ECI

[217] Takeyuki Suzuki, Jien Kato, and Toyohide Watanabe. Extraction of contextual information existing among component elements of origami books. Lecture Notes in Computer Science, 2390:158–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2390/23900158.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2390/23900158.pdf.

Cao:2002:TGS

[218] Ruini Cao and Chew Lim Tan. Text/graphics separation in maps. Lecture Notes in Computer Science, 2390:167–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2390/23900167.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2390/23900167.pdf.

Levachkine:2002:SAR

[219] Serguei Levachkine, Aurelio Velázquez, Victor Alexandrov, and Mikhail Kharinov. Semantic analysis and recognition of raster-scanned color cartographic images. Lecture Notes in Computer Science, 2390:178–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2390/23900178.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2390/23900178.pdf.

Weindorf:2002:SBI

[220] Manuel Weindorf. Structure based interpretation of unstructured vector maps. Lecture Notes in Computer Science, 2390:190–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2390/23900190.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2390/23900190.pdf.

Lanza:2002:GLD

[221] Antonietta Lanza, Donato Malerba, Francesca A. Lisi, Annalisa Appice, and Michelangelo Ceci. Generating logic descriptions for the automated interpretation of topographic maps. Lecture Notes in Computer Science, 2390:200–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2390/23900200.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2390/23900200.pdf.

Lamiroy:2002:SXU

[222] Bart Lamiroy and Laurent Najman. Scan-to-XML: Using software component algebra for intelligent document generation. Lecture Notes in Computer Science, 2390:211–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2390/23900211.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2390/23900211.pdf.

Mahoney:2002:ISS

[223] James V. Mahoney and Markus P. J. Fromherz. Interpreting sloppy stick figures by graph rectification and constraint-based matching. Lecture Notes in Computer Science, 2390:222–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2390/23900222.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2390/23900222.pdf.

Garcia:2002:UGD

[224] Pascal Garcia and Bertrand Coüasnon. Using a generic document recognition method for mathematical formulae recognition. Lecture Notes in Computer Science, 2390:236–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2390/23900236.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2390/23900236.pdf.

Pridmore:2002:ILD

[225] Tony P. Pridmore, Ahmed Darwish, and Dave Elliman. Interpreting line drawing images: a knowledge level perspective. Lecture Notes in Computer Science, 2390:245–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2390/23900245.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2390/23900245.pdf.

Bodansky:2002:SCL

[226] Eugene Bodansky, Alexander Gribov, and Morakot Pilouk. Smoothing and compression of lines obtained by raster-to-vector conversion. Lecture Notes in Computer Science, 2390:256–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2390/23900256.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2390/23900256.pdf.

Adam:2002:SRP

[227] Sebastien Adam, Jean-Marc Ogier, Claude Cariou, and Joel Gardes. A scale and rotation parameters estimator application to technical document interpretation. Lecture Notes in Computer Science, 2390: 266–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2390/23900266.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2390/23900266.
                                        pdf.

Hilaire:2002:IAS

[228] Xavier Hilaire and Karl Tombre. Improving the accuracy of skeletonbased vectorization. Lecture Notes in Computer Science, 2390:273– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2390/23900273.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2390/23900273.
pdf.

Agam:2002:SRN

[229] Gady Agam and Changhua Wu. Structural rectification of non-planar document images: Application to graphics recognition. Lecture Notes in Computer Science, 2390:289–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2390/23900289.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2390/23900289.pdf.

Park:2002:EVE

[230] Young-Jun Park and Young-Bin Kwon. An effective vector extraction method on architectural imaging using drawing characteristics. Lecture Notes in Computer Science, 2390:299–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2390/23900299.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2390/23900299.pdf.

Pridmore:2002:TIL

[231] Tony P. Pridmore. Thresholding images of line drawings with hysteresis. Lecture Notes in Computer Science, 2390:310–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2390/23900310.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2390/23900310.pdf.

Toshihiro:2002:RMM

[232] Kanahori Toshihiro and Suzuki Masakazu. A recognition method of matrices by using variable block pattern elements generating rectangular area. Lecture Notes in Computer Science, 2390:320–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2390/23900320.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2390/23900320.pdf.

Ng:2002:MMT

[233] Kia Ng. Music manuscript tracing. Lecture Notes in Computer Science, 2390:330–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2390/23900330.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2390/23900330.pdf.

Wenyin:2002:ESA

[234] Liu Wenyin, Jian Zhai, and Dov Dori. Extended summary of the arc segmentation contest. Lecture Notes in Computer Science, 2390: 343–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2390/23900343.htm;   http:

//link.springer.de/link/service/
series/0558/papers/2390/23900343.
pdf.

Elliman:2002:TAA

[235] Dave Elliman. TIF2VEC, an algorithm for arc segmentation in engineering drawings. Lecture Notes in Computer Science, 2390:350–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2390/23900350.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2390/23900350.pdf.

Hilaire:2002:RAS

[236] Xavier Hilaire. RANVEC and the arc segmentation contest. Lecture Notes in Computer Science, 2390: 359–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2390/23900359.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2390/23900359.
pdf.

Chhabra:2002:SPD

[237] Atul Chhabra and Karl Tombre. Summary of panel discussions at GREC’2001. Lecture Notes in Computer Science, 2390:365–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/link/service/
series/0558/bibs/2390/23900365.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2390/23900365.pdf.

Anonymous:2002:AIe

[238] Anonymous. Author index. Lecture Notes in Computer Science, 2390: 367–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2390/2390ind.pdf.

Barthe:2002:IDT

[239] Gilles Barthe and Thierry Coquand. An introduction to dependent type theory. Lecture Notes in Computer Science, 2395:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2395/23950001.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2395/23950001.pdf.

Benton:2002:ME

[240] Nick Benton, John Hughes, and Eugenio Moggi. Monads and effects. Lecture Notes in Computer Science, 2395:42–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2395/23950042.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2395/23950042.pdf.

Curien:2002:AMC

[241] Pierre-Louis Curien. Abstract machines, control, and sequents. Lecture Notes in Computer Science, 2395: 123–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2395/23950123.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2395/23950123.
pdf.

Dybjer:2002:NPE

[242] Peter Dybjer and Andrzej Filinski. Normalization and partial evaluation. Lecture Notes in Computer Science, 2395:137–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2395/23950137.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2395/23950137.pdf.

Edalat:2002:CRN

[243] Abbas Edalat and Reinhold Heckmann. Computing with real numbers. Lecture Notes in Computer Science, 2395:193–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2395/23950193.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2395/23950193.pdf.

Fournet:2002:JCL

[244] Cédric Fournet and Georges Gonthier. The join calculus: a language for distributed mobile programming. Lecture Notes in Computer Science, 2395:268–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2395/23950268.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2395/23950268.pdf.

Odersky:2002:IFN

[245] Martin Odersky. An introduction to functional nets. Lecture Notes in Computer Science, 2395: 333–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2395/23950333.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2395/23950333.
                                        pdf.

Pitts:2002:OSP

[246] Andrew M. Pitts. Operational semantics and program equivalence. Lecture Notes in Computer Science, 2395:378–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2395/23950378.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2395/23950378.pdf.

Remy:2002:UUU

[247] Didier Rémy. Using, understanding, and unraveling the OCaml language. from practice to theory and vice versa. Lecture Notes in Computer Science, 2395:413–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2395/23950413.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2395/23950413.pdf;   http://link.
                                        springer.de/link/service/series/

0558/papers/2395/2395a413.pdf;
http://link.springer.de/link/service/
series/0558/papers/2395/2395b413.
pdf;   http://link.springer.de/
link/service/series/0558/papers/
2395/2395c413.pdf;   http://link.
springer.de/link/service/series/
0558/papers/2395/2395d413.pdf;
http://link.springer.de/link/service/
series/0558/papers/2395/2395e413.
pdf;   http://link.springer.de/
link/service/series/0558/papers/
2395/2395f413.pdf;   http://link.
springer.de/link/service/series/
0558/papers/2395/2395g413.pdf;
http://link.springer.de/link/service/
series/0558/papers/2395/2395h413.
pdf;   http://link.springer.de/
link/service/series/0558/papers/
2395/2395i413.pdf;   http://link.
springer.de/link/service/series/
0558/papers/2395/2395j413.pdf;
http://link.springer.de/link/service/
series/0558/papers/2395/2395k413.
pdf;   http://link.springer.de/
link/service/series/0558/papers/
2395/2395l413.pdf;   http://link.
springer.de/link/service/series/
0558/papers/2395/2395m413.pdf;
http://link.springer.de/link/service/
series/0558/papers/2395/2395n413.
pdf.

Anonymous:2002:AIf

[248] Anonymous. Author index. Lecture Notes in Computer Science, 2395: 537–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2395/2395auth.pdf.

Suciu:2002:TSD

[249] Dan Suciu. Typechecking for semistructured data. Lecture Notes in Computer Science, 2397:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2397/23970001.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2397/23970001.pdf.

Deutsch:2002:OPC

[250] Alin Deutsch and Val Tannen. Optimization properties for classes of conjunctive regular path queries. Lecture Notes in Computer Science, 2397:21–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2397/23970021.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2397/23970021.pdf.

Calvanese:2002:VBQ

[251] Diego Calvanese, Giuseppe De Giacomo, Maurizio Lenzerini, and Moshe Y. Vardi. View-based query answering and query containment over semistructured data. Lecture Notes in Computer Science, 2397:40–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2397/23970040.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2397/23970040.pdf.

Dovier:2002:MCB

[252] Agostino Dovier and Elisa Quintarelli. Model-checking based data retrieval. Lecture Notes in Computer Science, 2397:62–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2397/23970062.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2397/23970062.pdf.

Vaisman:2002:TQL

[253] Alejandro A. Vaisman and Alberto O. Mendelzon. A temporal query language for OLAP: Implementation and a case study. Lecture Notes in Computer Science, 2397:78–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2397/23970078.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2397/23970078.pdf.

Merrett:2002:AMR

[254] T. H. Merrett. Attribute metadata for relational OLAP and data mining. Lecture Notes in Computer Science, 2397:97–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2397/23970097.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2397/23970097.pdf.

Calders:2002:MDM

[255] Toon Calders and Jef Wijsen. On monotone data mining languages.

Lecture Notes in Computer Science, 2397:119–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2397/23970119.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2397/23970119.pdf.

Buneman:2002:RAK

[256] Peter Buneman, Susan Davidson, Wenfei Fan, Carmem Hara, and Wang-Chiew Tan. Reasoning about keys for XML. Lecture Notes in Computer Science, 2397:133–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2397/23970133.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2397/23970133.pdf.

Jagadish:2002:TTA

[257] H. V. Jagadish, Laks V. S. Lakshmanan, Divesh Srivastava, and Keith Thompson. TAX: a tree algebra for XML. Lecture Notes in Computer Science, 2397:149–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2397/23970149.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2397/23970149.pdf.

May:2002:RBQ

[258] Wolfgang May. A rule-based querying and updating language for XML. Lecture Notes in Computer Science, 2397:165–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2397/23970165.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2397/23970165.pdf.

Geerts:2002:LAS

[259] Floris Geerts. Linear approximation of semi-algebraic spatial databases using transitive closure logic, in arbitrary dimension. Lecture Notes in Computer Science, 2397:182–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2397/23970182.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2397/23970182.pdf.

Geerts:2002:TST

[260] Floris Geerts, Sofie Haesevoets, and Bart Kuijpers. A theory of spatio-temporal database queries. Lecture Notes in Computer Science, 2397:198–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2397/23970198.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2397/23970198.pdf.

Fisher:2002:ASD

[261] Kathleen Fisher, Colin Goodall, Karin Högstedt, and Anne Rogers. An application-specific database. Lecture Notes in Computer Science, 2397: 213–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2397/23970213.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2397/23970213.
                                        pdf.

Alagic:2002:MTG

[262] Suad Alagić and Philip A. Bernstein. A model theory for generic schema management. Lecture Notes in Computer Science, 2397:228–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2397/23970228.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2397/23970228.pdf.

Cosmadakis:2002:VSU

[263] Stavros Cosmadakis, Kleoni Ioannidou, and Stergios Stergio. View serializable updates of concurrent index structures. Lecture Notes in Computer Science, 2397:247–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2397/23970247.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2397/23970247.pdf.

Cohen:2002:SFQ

[264] Sara Cohen, Yaron Kanza, and Yehoshua Sagiv. SQL4X: a flexible query language for XML and relational databases. Lecture Notes in Computer Science, 2397:263–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2397/23970263.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2397/23970263.pdf.

Psaila:2002:EQQ

[265] Giuseppe Psaila. ERX-QL: Querying an entity-relationship DB to obtain XML documents. Lecture Notes in Computer Science, 2397: 281–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2397/23970281.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2397/23970281.
pdf.

Bailey:2002:OAD

[266] James Bailey, Alexandra Poulovassilis, and Simon Courtenage. Optimising active database rules by partial evaluation and abstract interpretation. Lecture Notes in Computer Science, 2397:300–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2397/23970300.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2397/23970300.pdf.

Kiringa:2002:SAT

[267] Iluju Kiringa. Simulation of advanced transaction models using GOLOG. Lecture Notes in Computer Science, 2397:318–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2397/23970318.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2397/23970318.pdf.

Anonymous:2002:AIg

[268] Anonymous. Author index. Lecture Notes in Computer Science, 2397: 343–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2397/2397auth.pdf.

Boigelot:2002:RAC

[269] Bernard Boigelot and Pierre Wolper. Representing arithmetic constraints with finite automata: An overview. Lecture Notes in Computer Science, 2401:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2401/24010001.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2401/24010001.pdf.

Decker:2002:LDS

[270] Stefan Decker. Logic databases on the Semantic Web: Challenges and opportunities. Lecture Notes in Computer Science, 2401:20–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2401/24010020.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2401/24010020.pdf.

Russo:2002:AAA

[271] Alessandra Russo, Rob Miller, Bashar Nuseibeh, and Jeff Kramer. An abductive approach for analysing event-based requirements specifications. Lecture Notes in Computer Science, 2401:22–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2401/24010022.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2401/24010022.pdf.

Schrijvers:2002:TAH

[272] Tom Schrijvers, Maria García de la Banda, and Bart Demoen. Trailing analysis for HAL. Lecture Notes in Computer Science, 2401: 38–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2401/24010038.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2401/24010038.
pdf.

Barker:2002:ACD

[273] Steve Barker. Access control for deductive databases by logic programming. Lecture Notes in Computer Science, 2401:54–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2401/24010054.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2401/24010054.pdf.

Thielscher:2002:RAA

[274] Michael Thielscher. Reasoning about actions with CHRs and finite domain constraints. Lecture Notes in Computer Science, 2401:70–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2401/24010070.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2401/24010070.pdf.

Bockmayr:2002:UHC

[275] Alexander Bockmayr and Arnaud Courtois. Using hybrid concurrent constraint programming to model dynamic biological systems. Lecture Notes in Computer Science, 2401:85–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2401/24010085.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2401/24010085.pdf.

Pemmasani:2002:ERT

[276] Giridhar Pemmasani, C. R. Ramakrishnan, and I. V. Ramakrishnan. Efficient real-time model checking using tabled logic programming and constraints. Lecture Notes in Computer Science, 2401:100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2401/24010100.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2401/24010100.pdf.

Charatonik:2002:CBI

[277] Witold Charatonik, Supratik Mukhopadhyay, and Andreas Podelski. Constraintbased infinite model checking and tabulation for stratified CLP. Lecture Notes in Computer Science, 2401: 115–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2401/24010115.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2401/24010115.
pdf.

Jamil:2002:MTS

[278] Hasan M. Jamil and Gillian Dobbie. A model theoretic semantics for multi-level secure deductive databases. Lecture Notes in Computer Science, 2401:130–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2401/24010130.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2401/24010130.pdf.

Maher:2002:PCR

[279] Michael J. Maher. Propagation completeness of reactive constraints. Lecture Notes in Computer Science, 2401:148–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2401/24010148.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2401/24010148.pdf.

Makholm:2002:EWR

[280] Henning Makholm and Konstantinos Sagonas. On enabling the WAM with region support. Lecture Notes in Computer Science, 2401:

163–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2401/24010163.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2401/24010163.
                                        pdf.

Demoen:2002:DLG

[281] Bart Demoen. A different look at garbage collection for the WAM. Lecture Notes in Computer Science, 2401:179–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2401/24010179.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2401/24010179.pdf.

Demoen:2002:CGC

[282] Bart Demoen, Phuong-Lan Nguyen, and Ruben Vandeginste. Copying garbage collection for the WAM: To mark or not to mark? Lecture Notes in Computer Science, 2401: 194–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2401/24010194.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2401/24010194.
                                        pdf.

Ganzinger:2002:LA

[283] Harald Ganzinger and David McAllester. Logical algorithms. Lecture Notes in Computer Science, 2401:209– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2401/24010209.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2401/24010209.
pdf.

Schimpf:2002:LL

[284] Joachim Schimpf. Logical loops. Lecture Notes in Computer Science, 2401:224–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2401/24010224.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2401/24010224.pdf.

Martin:2002:LLR

[285] Eric Martin, Phuong Nguyen, Arun Sharma, and Frank Stephan. Learning in logic with RichProlog. Lecture Notes in Computer Science, 2401: 239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2401/24010239.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2401/24010239.
pdf.

Bry:2002:TDQ

[286] François Bry and Sebastian Schaffert. Towards a declarative query and transformation language for XML and semistructured data: Simulation unification. Lecture Notes in Computer Science, 2401:255–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2401/24010255.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2401/24010255.pdf.

Pientka:2002:PTF

[287] Brigitte Pientka. A proof-theoretic foundation for tabled higher-order logic programming. Lecture Notes in Computer Science, 2401:271– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2401/24010271.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2401/24010271.
                                        pdf.

Craciunescu:2002:PEC

[288] Sorin Craciunescu. Proving the equivalence of CLP programs. Lecture Notes in Computer Science, 2401: 287–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2401/24010287.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2401/24010287.
                                        pdf.

Bruscoli:2002:PLA

[289] Paola Bruscoli. A purely logical account of sequentiality in proof search. Lecture Notes in Computer Science, 2401:302–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2401/24010302.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2401/24010302.pdf.

Inoue:2002:DE

[290] Katsumi Inoue and Chiaki Sakama. Disjunctive explanations. Lecture Notes in Computer Science, 2401: 317–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2401/24010317.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2401/24010317.
pdf.

Bonatti:2002:RIS

[291] Piero A. Bonatti. Reasoning with infinite stable models II: Disjunctive programs. Lecture Notes in Computer Science, 2401:333–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2401/24010333.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2401/24010333.pdf.

Lonc:2002:CSM

[292] Zbigniew Lonc and Mirosław Truszczynski. Computing stable models: Worst-case performance estimates. Lecture Notes in Computer Science, 2401:347–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2401/24010347.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2401/24010347.pdf.

Dimopoulos:2002:TLS

[293] Yannis Dimopoulos and Andreas Sideris. Towards local search for answer sets. Lecture Notes in Computer Science, 2401:363–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2401/24010363.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2401/24010363.pdf.

Cabalar:2002:RMW

[294] Pedro Cabalar. A rewriting method for well-founded semantics with explicit negation. Lecture Notes in Computer Science, 2401:378–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2401/24010378.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2401/24010378.pdf.

Antoniou:2002:EDL

[295] Grigoris Antoniou and Michael J. Maher. Embedding defeasible logic into logic programs. Lecture Notes in Computer Science, 2401:393– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2401/24010393.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2401/24010393.
                                        pdf.

Pearce:2002:PTL

[296] David Pearce, Vladimir Sarsakov, Torsten Schaub, Hans Tompits, and Stefan Woltran. A polynomial translation of logic programs with nested expressions into disjunctive logic programs: Preliminary report. Lecture Notes in Computer Science, 2401:405–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2401/24010405.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2401/24010405.pdf.

Christensen:2002:ULP

[297] Henrik Bærbak Christensen. Using logic programming to detect activities in pervasive healthcare. Lecture Notes in Computer Science, 2401:421–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2401/24010421.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2401/24010421.pdf.

Lau:2002:LPS

[298] Kung-Kiu Lau and Michel Vanden Bossche. Logic programming for software engineering: a second chance. Lecture Notes in Computer Science, 2401:437–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2401/24010437.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2401/24010437.pdf.

Benko:2002:LBS

[299] Tamás Benko, Péter Krauth, and Péter Szeredi. A logic-based system for application integration. Lecture Notes in Computer Science, 2401: 452–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2401/24010452.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2401/24010452.
                                        pdf.

Ma:2002:LHL

[300] Shilong Ma, Yuefei Sui, and Ke Xu. The limits of Horn logic programs. Lecture Notes in Computer Science, 2401:467–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2401/24010467.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2401/24010467.pdf.

Medina:2002:MAL

[301] Jesús Medina, Enrique Mérida-Casermeiro, and Manuel Ojeda-Aciego. Multi-adjoint logic programming: a neural net approach. Lecture Notes in Computer Science, 2401:468–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2401/24010468.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2401/24010468.pdf.

Vaucheret:2002:FPa

[302] Claudio Vaucheret, Sergio Guadarrama, and Susana Muñoz. Fuzzy prolog. Lecture Notes in Computer Science, 2401:469–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2401/24010469.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2401/24010469.pdf.

Ducasse:2002:AAC

[303] Mireille Ducassé and Ludovic Langevine. Automated analysis of CLP(FD) program execution traces. Lecture Notes in Computer Science, 2401: 470–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2401/24010470.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2401/24010470.
pdf.

Olmer:2002:SBT

[304] Petr Olmer and Petr Štepánek. Schema-based transformations of logic programs in λprolog. Lecture Notes in Computer Science, 2401: 472–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2401/24010472.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2401/24010472.
pdf.

Loyer:2002:NUH

[305] Yann Loyer and Umberto Straccia. Non-uniform hypothesis in deductive databases with uncertainty. Lecture Notes in Computer Science, 2401:473–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2401/24010473.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2401/24010473.pdf.

Angelopoulos:2002:PFD

[306] Nicos Angelopoulos. Probabilistic finite domains: a brief overview. Lecture Notes in Computer Science, 2401:475–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2401/24010475.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2401/24010475.pdf.

Sampath:2002:MMA

[307] Prahladavaradan Sampath. Modelling multi-agent reactive systems. Lecture Notes in Computer Science, 2401:476–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2401/24010476.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2401/24010476.pdf.

Hayashi:2002:IPA

[308] Hisashi Hayashi, Kenta Cho, and Akihiko Ohsuga. Integrating planning, action execution, knowledge updates and plan modifications via logic programming. Lecture Notes in Computer Science, 2401:477–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2401/24010477.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2401/24010477.pdf.

Ferrand:2002:LPC

[309] Gérard Ferrand and Arnaud Lallouet. A logic program characterization of domain reduction approximations in finite domain CSPs. Lecture Notes in Computer Science, 2401:478–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2401/24010478.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2401/24010478.pdf.

Coquery:2002:TOS

[310] Emmanuel Coquery and François Fages. TCLP: Overloading, subtyping and parametric polymorphism made practical for CLP. Lecture Notes in Computer Science, 2401:480–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2401/24010480.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2401/24010480.pdf.

Christiansen:2002:LGB

[311] Henning Christiansen. Logical grammars based on constraint handling rules. Lecture Notes in Computer Science, 2401:481–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2401/24010481.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2401/24010481.pdf.

Osorio:2002:DPL

[312] Mauricio Osorio, Juan Antonio Navarro, and José Arrazola. Debugging in A-Prolog: a logical approach. Lecture Notes in Computer Science, 2401:482–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2401/24010482.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2401/24010482.pdf.

Anonymous:2002:AIh

[313] Anonymous. Author index. Lecture Notes in Computer Science, 2401: 485–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2401/2401ind.pdf.

Holzmann:2002:SAM

[314] Gerard J. Holzmann. Software analysis and model checking. Lecture Notes in Computer Science, 2404: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2404/24040001.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2404/24040001.
pdf.

Zhang:2002:QEB

[315] Lintao Zhang and Sharad Malik. The quest for efficient Boolean satisfiability solvers. Lecture Notes in Computer Science, 2404:17–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2404/24040017.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2404/24040017.pdf.

Cousot:2002:ASV

[316] Patrick Cousot and Radhia Cousot. On abstraction in software verification. Lecture Notes in Computer Science, 2404:37–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2404/24040037.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2404/24040037.pdf.

Henzinger:2002:SAH

[317] Thomas A. Henzinger. The symbolic approach to hybrid systems. Lecture Notes in Computer Science, 2404:57–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2404/24040057.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2404/24040057.pdf.

Thomas:2002:IGV

[318] Wolfgang Thomas. Infinite games and verification (extended abstract of a tutorial). Lecture Notes in Computer Science, 2404:58–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2404/24040058.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2404/24040058.pdf.

Barner:2002:SLR

[319] Sharon Barner, Daniel Geist, and Anna Gringauze. Symbolic localization reduction with reconstruction layering and backtracking. Lecture Notes in Computer Science, 2404:65–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2404/24040065.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2404/24040065.pdf.

Bryant:2002:MVS

[320] Randal E. Bryant, Shuvendu K. Lahiri, and Sanjit A. Seshia. Modeling and verifying systems using a logic of counter arithmetic with lambda expressions and uninterpreted functions. Lecture Notes in Computer Science, 2404:78–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2404/24040078.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2404/24040078.pdf.

Barner:2002:CSR

[321] Sharon Barner and Orna Grumberg. Combining symmetry reduction and under-approximation for symbolic model checking. Lecture Notes in Computer Science, 2404: 93–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2404/24040093.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2404/24040093.
pdf.

Pnueli:2002:LCA

[322] Amir Pnueli, Jessie Xu, and Lenore Zuck. Liveness with (0, 1, ∞)- counter abstraction. Lecture Notes in Computer Science, 2404:107– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2404/24040107.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2404/24040107.
pdf.

Chatterjee:2002:SMC

[323] Prosenjit Chatterjee, Hemanthkumar Sivaraj, and Ganesh Gopalakrishnan. Shared memory consistency protocol verification against weak memory models: Refinement via modelchecking. Lecture Notes in Computer Science, 2404:123–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2404/24040123.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2404/24040123.pdf.

Godefroid:2002:AAU

[324] Patrice Godefroid and Radha Jagadeesan. Automatic abstraction using generalized model checking. Lecture Notes in Computer Science, 2404:137–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2404/24040137.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2404/24040137.pdf.

Baumgartner:2002:PCS

[325] Jason Baumgartner, Andreas Kuehlmann, and Jacob Abraham. Property checking via structural analysis. Lecture Notes in Computer Science, 2404:151–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2404/24040151.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2404/24040151.pdf.

Rajamani:2002:CCM

[326] Sriram K. Rajamani and Jakob Rehof. Conformance checking for models of asynchronous message passing software. Lecture Notes in Computer Science, 2404:166–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2404/24040166.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2404/24040166.pdf.

Flanagan:2002:MCM

[327] Cormac Flanagan, Shaz Qadeer, and Sanjit A. Seshia. A modular checker for multithreaded programs. Lecture Notes in Computer Science, 2404:180–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2404/24040180.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2404/24040180.pdf.

Yoneda:2002:ADT

[328] Tomohiro Yoneda, Tomoya Kitai, and Chris Myers. Automatic derivation of timing constraints by failure analysis. Lecture Notes in Computer Science, 2404:195–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2404/24040195.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2404/24040195.pdf.

Strichman:2002:DSF

[329] Ofer Strichman, Sanjit A. Seshia, and Randal E. Bryant. Deciding separation formulas with SAT. Lecture Notes in Computer Science, 2404:209–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2404/24040209.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2404/24040209.pdf.

Younes:2002:PVD

[330] Håkan L. S. Younes and Reid G. Simmons. Probabilistic verification of discrete event systems using acceptance sampling. Lecture Notes in Computer Science, 2404:223–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2404/24040223.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2404/24040223.pdf.

Barrett:2002:CSF

[331] Clark W. Barrett, David L. Dill, and Aaron Stump. Checking satisfiability of first-order formulas by incremental translation to SAT. Lecture Notes in Computer Science, 2404:236–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2404/24040236.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2404/24040236.pdf.

McMillan:2002:ASM

[332] Ken L. McMillan. Applying SAT methods in unbounded symbolic model checking. Lecture Notes in Computer Science, 2404:250–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2404/24040250.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2404/24040250.pdf.

Clarke:2002:SBA

[333] Edmund Clarke, Anubhav Gupta, James Kukula, and Ofer Strichman. SAT based abstraction-refinement using ILP and machine learning techniques. Lecture Notes in Computer Science, 2404:265–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2404/24040265.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2404/24040265.pdf.

Bingham:2002:SFB

[334] Jesse D. Bingham and Alan J. Hu. Semi-formal bounded model checking. Lecture Notes in Computer Science, 2404:280–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2404/24040280.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2404/24040280.pdf.

Bozzano:2002:AVI

[335] Marco Bozzano and Giorgio Delzanno. Algorithmic verification of invalidationbased protocols. Lecture Notes in Computer Science, 2404:295– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2404/24040295.htm;   http:
//link.springer.de/link/service/

                                        series/0558/papers/2404/24040295.
                                        pdf.

Jacobi:2002:FVC

[336] Christian Jacobi. Formal verification of complex out-of-order pipelines by combining model-checking and theoremproving. Lecture Notes in Computer Science, 2404:309–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2404/24040309.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2404/24040309.pdf.

Chevalier:2002:AUV

[337] Yannick Chevalier and Laurent Vigneron. Automated unbounded verification of security protocols. Lecture Notes in Computer Science, 2404:324–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2404/24040324.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2404/24040324.pdf.

Alur:2002:EBH

[338] Rajeev Alur, Michael McDougall, and Zijiang Yang. Exploiting behavioral hierarchy for efficient model checking. Lecture Notes in Computer Science, 2404:338–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2404/24040338.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2404/24040338.pdf.

Bozga:2002:IVE

[339] Marius Bozga, Susanne Graf, and Laurent Mounier. IF-2.0: a validation environment for componentbased real-time systems. Lecture Notes in Computer Science, 2404: 343–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2404/24040343.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2404/24040343.
pdf.

Armando:2002:ASP

[340] Alessandro Armando, David Basin, Mehdi Bouallagui, Yannick Chevalier, Luca Compagna, Sebastian Mödersheim, Michael Rusinowitch, Mathieu Turuani, Luca Viganò, and Laurent Vigneron. The AVISS security protocol analysis tool. Lecture Notes in Computer Science, 2404: 349–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2404/24040349.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2404/24040349.
pdf.

Asarin:2002:SVT

[341] Eugene Asarin, Gordon Pace, Gerardo Schneider, and Sergio Yovine. SPeeDI — A verification tool for polygonal hybrid systems. Lecture Notes in Computer Science, 2404: 354–??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2404/24040354.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2404/24040354.
                                        pdf.

Cimatti:2002:NOT

[342] Alessandro Cimatti, Edmund Clarke, Enrico Giunchiglia, Fausto Giunchiglia, Marco Pistore, Marco Roveri, Roberto Sebastiani, and Armando Tacchella. NuSMV 2: An OpenSource tool for symbolic model checking. Lecture Notes in Computer Science, 2404: 359–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2404/24040359.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2404/24040359.
                                        pdf.

Asarin:2002:TVH

[343] Eugene Asarin, Thao Dang, and Oded Maler. The d/dt tool for verification of hybrid systems. Lecture Notes in Computer Science, 2404:365–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2404/24040365.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2404/24040365.pdf.

Kupferman:2002:MCL

[344] Orna Kupferman, Nir Piterman, and Moshe Y. Vardi. Model checking linear properties of prefix-recognizable systems. Lecture Notes in Computer Science, 2404:371–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2404/24040371.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2404/24040371.pdf.

Rybina:2002:UCR

[345] Tatiana Rybina and Andrei Voronkov. Using canonical representations of solutions to speed up infinite-state model checking. Lecture Notes in Computer Science, 2404:386–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2404/24040386.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2404/24040386.pdf.

Hartong:2002:DMM

[346] Walter Hartong, Lars Hedrich, and Erich Barke. On discrete modeling and model checking for nonlinear analog systems. Lecture Notes in Computer Science, 2404:401–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2404/24040401.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2404/24040401.pdf.

Chakrabarti:2002:SBC

[347] Arindam Chakrabarti, Luca de Alfaro, Thomas A. Henzinger, and Freddy Y. C. Mang. Synchronous and bidirectional component interfaces. Lecture Notes in Computer Science, 2404:414–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2404/24040414.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2404/24040414.pdf.

Chakrabarti:2002:ICC

[348] Arindam Chakrabarti, Luca de Alfaro, Thomas A. Henzinger, Marcin Jurdzinski, and Freddy Y. C. Mang. Interface compatibility checking for software modules. Lecture Notes in Computer Science, 2404:428–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2404/24040428.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2404/24040428.pdf.

Colon:2002:PMP

[349] Michael A. Colón and Henny B. Sipma. Practical methods for proving program termination. Lecture Notes in Computer Science, 2404: 442–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2404/24040442.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2404/24040442.
                                        pdf.

Tan:2002:EBM

[350] Li Tan and Rance Cleaveland. Evidence-based model checking. Lecture Notes in Computer Science, 2404:

455–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2404/24040455.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2404/24040455.
pdf.

Cabodi:2002:MFB

[351] Gianpiero Cabodi, Sergio Nocco, and Stefano Quer. Mixing forward and backward traversals in guided-prioritized BDD-based verification. Lecture Notes in Computer Science, 2404:471–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2404/24040471.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2404/24040471.pdf.

Purandare:2002:VCC

[352] Mitra Purandare and Fabio Somenzi. Vacuum cleaning CTL formulae. Lecture Notes in Computer Science, 2404:485–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2404/24040485.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2404/24040485.pdf.

Stump:2002:CCV

[353] Aaron Stump, Clark W. Barrett, and David L. Dill. CVC: a cooperating validity checker. Lecture Notes in Computer Science, 2404: 500–??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2404/24040500.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2404/24040500.
                                        pdf.

Chechik:2002:MVM

[354] Marsha Chechik, Arie Gurfinkel, and Benet Devereux. χChek: a multi-valued model-checker. Lecture Notes in Computer Science, 2404: 505–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2404/24040505.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2404/24040505.
                                        pdf.

Ben-David:2002:PTD

[355] Shoham Ben-David, Anna Gringauze, Baruch Sterin, and Yaron Wolfsthal. PathFinder: a tool for design exploration. Lecture Notes in Computer Science, 2404:510–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2404/24040510.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2404/24040510.pdf.

Dams:2002:ACA

[356] Dennis Dams, William Hesse, and Gerard Holzmann. Abstracting C with abC. Lecture Notes in Computer Science, 2404:515–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2404/24040515.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2404/24040515.pdf.

Groce:2002:AAM

[357] Alex Groce, Doron Peled, and Mihalis Yannakakis. AMC: An adaptive model checker. Lecture Notes in Computer Science, 2404:521–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2404/24040521.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2404/24040521.pdf.

Henzinger:2002:TSP

[358] Thomas A. Henzinger, Ranjit Jhala, Rupak Majumdar, George C. Necula, Grégoire Sutre, and Westley Weimer. Temporal-safety proofs for systems code. Lecture Notes in Computer Science, 2404:526–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2404/24040526.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2404/24040526.pdf.

Bouajjani:2002:ETT

[359] Ahmed Bouajjani and Tayssir Touili. Extrapolating tree transformations. Lecture Notes in Computer Science, 2404:539–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2404/24040539.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2404/24040539.pdf.

Abdulla:2002:RTM

[360] Parosh Aziz Abdulla, Bengt Jonsson, Pritha Mahata, and Julien d’Orso. Regular tree model checking. Lecture Notes in Computer Science, 2404:555–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2404/24040555.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2404/24040555.pdf.

Kurshan:2002:CTM

[361] Robert Kurshan, Vladimir Levin, and Hüsnü Yenigün. Compressing transitions for model checking. Lecture Notes in Computer Science, 2404:569–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2404/24040569.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2404/24040569.pdf.

Khomenko:2002:CPP

[362] Victor Khomenko, Maciej Koutny, and Walter Vogler. Canonical prefixes of petri net unfoldings. Lecture Notes in Computer Science, 2404:582–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2404/24040582.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2404/24040582.pdf.

Blom:2002:SSR

[363] Stefan Blom and Jaco van de Pol. State space reduction by proving confluence. Lecture Notes in Computer Science, 2404:596–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2404/24040596.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2404/24040596.pdf.

Gurumurthy:2002:FSM

[364] Sankar Gurumurthy, Roderick Bloem, and Fabio Somenzi. Fair simulation minimization. Lecture Notes in Computer Science, 2404:610– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2404/24040610.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2404/24040610.
pdf.

Anonymous:2002:AIi

[365] Anonymous. Author index. Lecture Notes in Computer Science, 2404: 625–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2404/2404ind.pdf.

Myers:2002:MDC

[366] B. A. Myers. Mobile devices for control. Lecture Notes in Computer Science, 2411:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2411/24110001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2411/24110001.pdf.

Passani:2002:BUW

[367] L. Passani. Building usable wireless applications for mobile phones. Lecture Notes in Computer Science, 2411:9–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2411/24110009.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2411/24110009.pdf.

Colbert:2002:DSR

[368] M. Colbert. A diary study of rendezvousing: Group size, time pressure and connectivity. Lecture Notes in Computer Science, 2411: 21–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2411/24110021.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2411/24110021.
pdf.

Gelgon:2002:AMD

[369] M. Gelgon and K. Tilhou. Automated multimedia diaries of mobile device users need summarization. Lecture Notes in Computer Science, 2411:36–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2411/24110036.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2411/24110036.pdf.

Lin:2002:PLA

[370] J. Lin, R. Laddaga, and H. Naito. Personal location agent for communicating entities (PLACE). Lecture Notes in Computer Science, 2411: 45–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2411/24110045.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2411/24110045.
pdf.

Huang:2002:DEI

[371] E. M. Huang, M. Terry, E. Mynatt, K. Lyons, and A. Chen. Distributing event information by simulating word-of-mouth exchanges. Lecture Notes in Computer Science, 2411:60–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2411/24110060.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2411/24110060.pdf.

Gonzalez-Castano:2002:NTT

[372] F. J. González-Castaño, L. Anido-Rifón, and E. Costa-Montenegro. A new transcoding technique for PDA browsers, based on content hierarchy. Lecture Notes in Computer Science, 2411:69–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2411/24110069.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2411/24110069.pdf.

Jones:2002:SSS

[373] M. Jones, G. Buchanan, and H. Thimbleby. Sorting out searching on small screen devices. Lecture Notes in Computer Science, 2411: 81–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2411/24110081.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2411/24110081.
                                        pdf.

Mantyjarvi:2002:AAM

[374] J. Mäntyjärvi and T. Seppänen. Adapting applications in mobile terminals using fuzzy context information. Lecture Notes in Computer Science, 2411:95–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2411/24110095.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2411/24110095.pdf.

Holland:2002:DCN

[375] S. Holland, D. R. Morse, and H. Gedenryd. Direct combination: a new user interaction. Lecture Notes in Computer Science, 2411:108–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2411/24110108.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2411/24110108.pdf.

Dubois:2002:ADN

[376] E. Dubois, P. Gray, and L. Nigay. ASUR++: a design notation for mobile mixed systems. Lecture Notes in Computer Science, 2411:123–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2411/24110123.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2411/24110123.pdf.

Pospischil:2002:DLM

[377] G. Pospischil, M. Umlauft, and E. Michlmayr. Designing LoL@, a mobile tourist guide for UMTS. Lecture Notes in Computer Science, 2411:140–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2411/24110140.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2411/24110140.pdf.

Bohnenberger:2002:LAS

[378] T. Bohnenberger, A. Jameson, A. Krüger, and A. Butz. Location-aware shopping assistance: Evaluation of a decision-theoretic approach. Lecture Notes in Computer Science, 2411:155–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2411/24110155.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2411/24110155.pdf.

Hibino:2002:HAE

[379] S. Hibino and A. Mockus. handiMessenger: Awareness-enhanced universal communication for mobile users. Lecture Notes in Computer Science, 2411:170–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2411/24110170.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2411/24110170.pdf.

Mizobuchi:2002:ESM

[380] S. Mizobuchi, K. Mori, X. Ren, and Y. Michiaki. An empirical study of the minimum required size and the minimum number of targets for pen input on the small display. Lecture Notes in Computer Science, 2411:184–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2411/24110184.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2411/24110184.pdf.

MacKenzie:2002:KKP

[381] I. S. MacKenzie. KSPC (keystrokes per character) as a characteristic of text entry techniques. Lecture Notes in Computer Science, 2411: 195–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2411/24110195.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2411/24110195.
pdf.

Chincholle:2002:LFU

[382] D. Chincholle, M. Goldstein, M. Nyberg, and M. Eriksson. Lost or found? A usability evaluation of a mobile navigation and location-based service. Lecture Notes in Computer Science, 2411:211–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2411/24110211.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2411/24110211.pdf.

Oquist:2002:TIR

[383] G. Öquist and M. Goldstein. Towards an improved readability on mobile devices: Evaluating adaptive rapid serial visual presentation. Lecture Notes in Computer Science, 2411:225–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2411/24110225.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2411/24110225.pdf.

Nigay:2002:MCA

[384] L. Nigay, P. Salembier, T. Marchand, P. Renevier, and L. Pasqualetti. Mobile and collaborative augmented reality: a scenario based design approach. Lecture Notes in Computer Science, 2411:241–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2411/24110241.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2411/24110241.pdf.

Sazawal:2002:UA

[385] V. Sazawal, R. Want, and G. Borriello. The unigesture approach. Lecture Notes in Computer Science, 2411:256–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2411/24110256.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2411/24110256.pdf.

Kjeldskov:2002:JPI

[386] J. Kjeldskov. “Just-in-Place” information for mobile device interfaces. Lecture Notes in Computer Science, 2411:271–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2411/24110271.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2411/24110271.pdf.

Costa-Montenegro:2002:PWP

[387] E. Costa-Montenegro, F. J. González-Castaño, J. García-Reinoso, J. Vales-Alonso, and L. Anido-Rifón. Pushing Web pages into personal digital assistants: Need, tools and solutions. Lecture Notes in Computer Science, 2411:276–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2411/24110276.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2411/24110276.pdf.

Dourish:2002:EMA

[388] P. Dourish and A. van der Hoek. Émigré: Metalevel architecture and migratory work. Lecture Notes in Computer Science, 2411:281– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2411/24110281.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2411/24110281.
pdf.

Carboni:2002:IMC

[389] D. Carboni, S. Sanna, S. Giroux, and G. Paddeu. Interactions model and code generation for J2ME applications. Lecture Notes in Computer Science, 2411:286–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2411/24110286.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2411/24110286.pdf.

Fallman:2002:IWT

[390] D. Fallman. An interface with weight: Taking interaction by tilt beyond disembodied metaphors. Lecture Notes in Computer Science, 2411: 291–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2411/24110291.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2411/24110291.
                                        pdf.

Mayora-Ibarra:2002:UMD

[391] O. Mayora-Ibarra, E. Cambranes-Martínez, C. Miranda-Palma, A. Fuentes-Penna, and O. De la Paz-Arroyo. UML modelling of device-independent interfaces and services for a home environment application. Lecture Notes in Computer Science, 2411: 296–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2411/24110296.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2411/24110296.
                                        pdf.

Tummolini:2002:IWM

[392] L. Tummolini, A. Lorenzon, G. Bo, and R. Vaccaro. iTutor: a wireless and multimodal support to industrial maintenance activities. Lecture Notes in Computer Science, 2411:302–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2411/24110302.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2411/24110302.pdf.

Luchini:2002:ACS

[393] K. Luchini, C. Quintana, and E. Soloway. ArtemisExpress: a case study in designing handheld interfaces for an online digital library. Lecture Notes in Computer Science, 2411:306–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2411/24110306.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2411/24110306.pdf.

OHare:2002:AMH

[394] G. O’Hare and M. O’Grady. Addressing mobile HCI needs through agents. Lecture Notes in Computer Science, 2411:311–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2411/24110311.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2411/24110311.pdf.

vanWelie:2002:CMD

[395] M. van Welie and B. de Groot. Consistent multi-device design using device categories. Lecture Notes in Computer Science, 2411:315–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2411/24110315.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2411/24110315.pdf.

Kohno:2002:NGI

[396] M. Kohno and J. Rekimoto. New generation of IP-phone enabled mobile devices. Lecture Notes in Computer Science, 2411:319–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2411/24110319.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2411/24110319.pdf.

Yang:2002:CPS

[397] S. Yang, H. Lee, K. Chung, and H. Kim. A content providerspecified Web clipping approach for mobile content adaptation. Lecture Notes in Computer Science, 2411: 324–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2411/24110324.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2411/24110324.
                                        pdf.

Lamberts:2002:CSP

[398] H. Lamberts. Case study: a PDA example of user centered design. Lecture Notes in Computer Science, 2411:329–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2411/24110329.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2411/24110329.pdf.

deGroot:2002:LCU

[399] B. de Groot and M. van Welie. Leveraging the context of use in mobile service design. Lecture Notes in Computer Science, 2411:334–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2411/24110334.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2411/24110334.pdf.

Paredes:2002:UCE

[400] M. Paredes, M. Ortega, P. P. Sánchez-Villalón, and J.Á. Velázquez-Iturbide. A ubiquitous computing environment for language learning. Lecture Notes in Computer Science, 2411:339–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2411/24110339.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2411/24110339.pdf.

Kaikkonen:2002:XMA

[401] A. Kaikkonen and V. Roto. XHTML in mobile application development. Lecture Notes in Computer Science, 2411:344–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2411/24110344.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2411/24110344.pdf.

Bradley:2002:UCI

[402] N. A. Bradley and M. D. Dunlop. Understanding contextual interactions to design navigational context-aware applications. Lecture Notes in Computer Science, 2411:349–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2411/24110349.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2411/24110349.pdf.

Vainio:2002:TSU

[403] T. Vainio, O. Kotala, I. Rakkolainen, and H. Kupila. Towards scalable user interfaces in 3D city information systems. Lecture Notes in Computer Science, 2411:354–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2411/24110354.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2411/24110354.pdf.

Soronen:2002:MIM

[404] A. Soronen and V. Tuomisto. Mobile image messaging — anticipating the outlines of the usage culture. Lecture Notes in Computer Science, 2411:359–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2411/24110359.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2411/24110359.pdf.

Tarrini:2002:RCH

[405] L. Tarrini, R. Bianchi Bandinelli, V. Miori, and G. Bertini. Remote control of home automation systems with mobile devices. Lecture Notes in Computer Science, 2411:364–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2411/24110364.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2411/24110364.pdf.

Sokoler:2002:LRS

[406] T. Sokoler, L. Nelson, and E. R. Pedersen. Low-resolution supplementary tactile cues for navigational assistance. Lecture Notes in Computer Science, 2411:369–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2411/24110369.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2411/24110369.pdf.

Abascal:2002:MIS

[407] J. Abascal, D. Cagigas, N. Garay, and L. Gardeazabal. Mobile interface for a smart wheelchair. Lecture Notes in Computer Science, 2411:373–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2411/24110373.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2411/24110373.pdf.

Oquist:2002:UGD

[408] G. Öquist, S. Björk, and M. Goldstein. Utilizing gaze detection to simulate the affordances of paper in the rapid serial visual presentation format. Lecture Notes in Computer Science, 2411:378–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2411/24110378.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2411/24110378.pdf.

Schofield:2002:IMI

[409] E. Schofield and G. Kubin. On interfaces for mobile information retrieval. Lecture Notes in Computer Science, 2411:383–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2411/24110383.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2411/24110383.pdf.

Hermann:2002:URA

[410] F. Hermann and F. Heidmann. User requirement analysis and interface conception for a mobile, location-based fair guide. Lecture Notes in Computer Science, 2411:388–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2411/24110388.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2411/24110388.pdf.

Fontana:2002:CME

[411] F. Fontana, A. Fusiello, M. Gobbi, V. Murino, D. Rocchesso, L. Sartor, and A. Panuccio. A crossmodal electronic travel aid device. Lecture Notes in Computer Science, 2411:393–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2411/24110393.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2411/24110393.pdf.

Chesta:2002:GVE

[412] C. Chesta. Globalization of voiceenabled Internet access solutions. Lecture Notes in Computer Science, 2411:398–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2411/24110398.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2411/24110398.pdf.

Brusilovsky:2002:MBA

[413] P. Brusilovsky and R. Rizzo. Mapbased access to multiple educational on-line resources from mobile wireless devices. Lecture Notes in Computer Science, 2411:404–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2411/24110404.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2411/24110404.pdf.

Marucci:2002:SAH

[414] L. Marucci and F. Paternò. Supporting adaptivity to heterogeneous platforms through user models. Lecture Notes in Computer Science, 2411: 409–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2411/24110409.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2411/24110409.
pdf.

Chittaro:2002:VBC

[415] L. Chittaro and A. Camaggio. Visualizing bar charts on WAP phones. Lecture Notes in Computer Science, 2411:414–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2411/24110414.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2411/24110414.pdf.

Lorho:2002:SMP

[416] G. Lorho, J. Hiipakka, and J. Marila. Structured menu presentation using spatial sound separation. Lecture Notes in Computer Science, 2411:419–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2411/24110419.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2411/24110419.pdf.

vonNiman:2002:DSW

[417] B. von Niman, C. Chaplin, J.-A. Collado-Vega, L. Groh, S. McGlashan, W. Mellors, and D. van Leeuwen. “Do as I Say! But Who Says What I Should Say — or Do?” on the definition of a standard spoken command vocabulary for ICT devices and services. Lecture Notes in Computer Science, 2411:425–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2411/24110425.
htm;   http://link.springer.de/
link/service/series/0558/papers/

2411/24110425.pdf. ETSI Special

Task Force STF182.

Anonymous:2002:AIj

[418] Anonymous. Author index. Lecture Notes in Computer Science, 2411: 431–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2411/2411auth.pdf.

Patton:2002:DRI

[419] Jeff Patton. Designing requirements: Incorporating usage-centered design into an agile SW development process. Lecture Notes in Computer Science, 2418:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180001.pdf.

Maurer:2002:SDE

[420] Frank Maurer. Supporting distributed extreme programming. Lecture Notes in Computer Science, 2418: 13–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2418/24180013.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2418/24180013.
pdf.

Shaw:2002:UAS

[421] Steve Shaw. Using ant to solve problems posed by frequent deployments. Lecture Notes in Computer Science, 2418:23–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180023.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180023.pdf.

Henninger:2002:SAM

[422] Scott Henninger, Aditya Ivaturi, Krishna Nuli, and Ashok Thirunavukkaras. Supporting adaptable methodologies to meet evolving project needs. Lecture Notes in Computer Science, 2418:33–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180033.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180033.pdf.

Rasmusson:2002:SIX

[423] Jonathan Rasmusson. Strategies for introducing XP to new client sites. Lecture Notes in Computer Science, 2418:45–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180045.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180045.pdf.

Johansen:2002:EAT

[424] Kay Johansen and Anthony Perkins. Establishing an agile testing team: Our four favorite “Mistakes”. Lecture Notes in Computer Science, 2418: 52–??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2418/24180052.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2418/24180052.
pdf.

Krebs:2002:TKC

[425] William Krebs. Turning the knobs: a coaching pattern for XP through agile metrics. Lecture Notes in Computer Science, 2418:60–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180060.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180060.pdf.

Alleman:2002:APM

[426] Glen B. Alleman. Agile project management methods for ERP: How to apply agile processes to complex COTS projects and live to tell about it. Lecture Notes in Computer Science, 2418:70–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180070.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180070.pdf.

Wood:2002:EPR

[427] William A. Wood and William L. Kleb. Extreme programming in a research environment. Lecture Notes in Computer Science, 2418:89–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2418/24180089.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2418/24180089.pdf.

Bowers:2002:TXL

[428] Jason Bowers, John May, Erik Melander, Matthew Baarman, and Azeem Ayoob. Tailoring XP for large system mission critical software development. Lecture Notes in Computer Science, 2418:100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2418/24180100.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2418/24180100.pdf.

Kitiyakara:2002:ATH

[429] Narti Kitiyakara. Acceptance testing HTML. Lecture Notes in Computer Science, 2418:112–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2418/24180112.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2418/24180112.pdf.

Jalis:2002:PTS

[430] Asim Jalis. Probe tests: a strategy for growing automated tests around legacy code. Lecture Notes in Computer Science, 2418:122–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2418/24180122.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2418/24180122.pdf.

Stotts:2002:IFM

[431] David Stotts, Mark Lindsey, and Angus Antley. An informal formal method for systematic JUnit test case generation. Lecture Notes in Computer Science, 2418:131–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180131.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180131.pdf.

Newkirk:2002:LDP

[432] James Newkirk. A light in a dark place: Test-driven development with 3^{rd} party packages. Lecture Notes in Computer Science, 2418:144–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180144.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180144.pdf.

Turner:2002:AMC

[433] Richard Turner and Apurva Jain. Agile meets CMMI: Culture clash or common cause? Lecture Notes in Computer Science, 2418:153–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180153.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180153.pdf.

Ramachandran:2002:CLS

[434] Vinay Ramachandran and Anuja Shukla. Circle of life, spiral of death: Are XP teams following the essential practices? Lecture Notes in Computer Science, 2418:166–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2418/24180166.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2418/24180166.pdf.

Rostaher:2002:TTF

[435] Matevz Rostaher and Marjan Hericko. Tracking test first pair programming — an experiment. Lecture Notes in Computer Science, 2418:174–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2418/24180174.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2418/24180174.pdf.

Reifer:2002:HGM

[436] Donald J. Reifer and President. How to get the most out of extreme programming/agile methods. Lecture Notes in Computer Science, 2418:185–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2418/24180185.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2418/24180185.pdf.

Lindvall:2002:EFA

[437] Mikael Lindvall, Vic Basili, Barry Boehm, Patricia Costa, Kathleen Dangle, Forrest Shull, Roseanne Tesoriero, Laurie Williams, and Marvin Zelkowitz. Empirical findings in agile methods. Lecture Notes in Computer Science, 2418:197–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180197.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180197.pdf.

Baheti:2002:EED

[438] Prashant Baheti, Edward Gehringer, and David Stotts. Exploring the efficacy of distributed pair programming. Lecture Notes in Computer Science, 2418:208–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180208.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180208.pdf.

Srinivasa:2002:PPA

[439] Gopal Srinivasa and Prasanth Ganesan. Pair programming: Addressing key process areas of the people-CMM. Lecture Notes in Computer Science, 2418:221–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180221.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180221.pdf.

Miller:2002:WPD

[440] Roy W. Miller. When pairs disagree, 1-2-3. Lecture Notes in Computer Science, 2418:231–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2418/24180231.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2418/24180231.pdf.

Gabriel:2002:TPH

[441] Richard P. Gabriel. Triggers and practice: How extremes in writing relate to creativity and learning. Lecture Notes in Computer Science, 2418:237–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2418/24180237.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2418/24180237.pdf.

Steinberg:2002:ETA

[442] Daniel Steinberg. Extreme teaching — an agile approach to education. Lecture Notes in Computer Science, 2418:238–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2418/24180238.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2418/24180238.pdf.

Naveda:2002:EPT

[443] J. Fernando Naveda, Kent Beck, Richard P. Gabriel, Jorge Diaz Herrera, Watts Humphrey, Michael Mc-Cracken, and Dave West. Extreme programming as a teaching process. Lecture Notes in Computer Science, 2418:239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180239.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180239.pdf.

Caristi:2002:SP

[444] James Caristi, Frank Maurer, and Michael Rettig. From the student’s perspective. Lecture Notes in Computer Science, 2418:240–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180240.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180240.pdf.

Melnik:2002:PAP

[445] Grigori Melnik and Frank Maurer. Perceptions of agile practices: a student survey. Lecture Notes in Computer Science, 2418:241–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180241.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180241.pdf.

Thomas:2002:XLE

[446] Kuryan Thomas and Arlen Bankston. XP in a legacy environment. Lecture Notes in Computer Science, 2418:

251–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2418/24180251.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2418/24180251.
                                        pdf.

Grenning:2002:XD

[447] James Grenning. XP for a day. Lecture Notes in Computer Science, 2418:253–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2418/24180253.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2418/24180253.pdf.

Airth:2002:ASC

[448] Alan Airth. Accelerated solution centers — implementing DSDM in the real world. Lecture Notes in Computer Science, 2418:254–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2418/24180254.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2418/24180254.pdf.

Fowler:2002:RID

[449] Martin Fowler. Refactoring: Improving the design of existing code. Lecture Notes in Computer Science, 2418:256–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2418/24180256.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180256.pdf.

Sadalage:2002:AD

[450] Pramod Sadalage and Peter Schuh. The agile database. Lecture Notes in Computer Science, 2418:257– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2418/24180257.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2418/24180257.
pdf.

Kerievsky:2002:CWT

[451] Joshua Kerievsky and Diana Larsen. Change wizardry — tools for geeks. Lecture Notes in Computer Science, 2418:259–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180259.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180259.pdf.

Hodgetts:2002:BCA

[452] Paul Hodgetts. Beyond the customer: Agile business practices for XP. Lecture Notes in Computer Science, 2418:261–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180261.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180261.pdf.

Hendrickson:2002:XRP

[453] Chet Hendrickson, Ann Anderson, and Ron Jeffries. XP release planning and user stories. Lecture Notes in Computer Science, 2418:263–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2418/24180263.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2418/24180263.pdf.

Lindstrom:2002:SBS

[454] Lowell Lindstrom. Steering the big ship: Succeeding in changing an organization’s practices. Lecture Notes in Computer Science, 2418: 264–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2418/24180264.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2418/24180264.
                                        pdf.

Schwaber:2002:SA

[455] Ken Schwaber. Scrum and agile 101. Lecture Notes in Computer Science, 2418:266–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2418/24180266.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2418/24180266.pdf.

Wake:2002:HC

[456] William Wake and Ron Jeffries. How to be a coach. Lecture Notes in Computer Science, 2418:268–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180268.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180268.pdf.

Hill:2002:SAT

[457] Michael Hill. Sharpening the axe for test driven development. Lecture Notes in Computer Science, 2418:269–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180269.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180269.pdf.

Williams:2002:PPE

[458] Laurie Williams and Robert Kessler. Pair programming: Experience the difference. Lecture Notes in Computer Science, 2418:271–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180271.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180271.pdf.

Breitling:2002:HSX

[459] Holger Breitling and Martin Lippert. How to start an XP project: The initial phase. Lecture Notes in Computer Science, 2418:273–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2418/24180273.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2418/24180273.pdf.

Goodsen:2002:EJT

[460] John Goodsen. Effective Java testing strategies. Lecture Notes in Computer Science, 2418:275–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2418/24180275.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2418/24180275.pdf.

Crispin:2002:TDT

[461] Lisa Crispin. Test drive for testers: What, when, and how testers do for XP teams. Lecture Notes in Computer Science, 2418:277–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2418/24180277.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2418/24180277.pdf.

Eckstein:2002:SAP

[462] Jutta Eckstein. Scaling agile processes: Agile software development in large projects. Lecture Notes in Computer Science, 2418:279–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2418/24180279.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2418/24180279.pdf.

Hightower:2002:AXT

[463] Richard Hightower. Applying XP tools to J2EE for the extreme programming universe. Lecture Notes in Computer Science, 2418:281–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180281.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180281.pdf.

Stotts:2002:DPP

[464] David Stotts and Laurie Williams. Distributed pair programming. Lecture Notes in Computer Science, 2418: 283–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2418/24180283.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2418/24180283.
pdf.

Pettichord:2002:AAT

[465] Bret Pettichord and Brian Marick. Agile acceptance testing. Lecture Notes in Computer Science, 2418:284–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180284.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180284.pdf.

Talbott:2002:XF

[466] Nathaniel Talbott and Duff O’Melia. XP fest. Lecture Notes in Computer Science, 2418:285–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180285.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180285.pdf.

Melnik:2002:EEA

[467] Grigori Melnik, Laurie Williams, and Adam Geras. Empirical evaluation of agile processes. Lecture Notes in Computer Science, 2418: 286–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2418/24180286.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2418/24180286.
pdf.

Auer:2002:TEH

[468] Ken Auer, Ron Jeffries, Jeff Canna, Glen B. Alleman, Lisa Crispin, and Janet Gregory. Are testers eXtinct? how can testers contribute to XP teams? Lecture Notes in Computer Science, 2418:287–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180287.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180287.pdf.

Fraser:2002:XBL

[469] Steven Fraser, Rachel Reinitz, Ken Auer, Barry Boehm, Ward Cunningham, and Rob Mee. XP — beyond limitations? Lecture Notes in Computer Science, 2418:288–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2418/24180288.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2418/24180288.pdf.

Johansen:2002:EF

[470] Zhon Johansen, Ken Auer, Brian Button, Alistair Cockburn, James Grenning, Kay Johansen, and Duff O’Melia. Extreme fishbowl. Lecture Notes in Computer Science, 2418: 289–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2418/24180289.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2418/24180289.
pdf.

Thomas:2002:AE

[471] Dave Thomas, John Manzo, Narti Kitiyakara, Russell Stay, and Aldo Dagnino. Agile experiences. Lecture Notes in Computer Science, 2418: 290–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2418/24180290.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2418/24180290.
pdf.

Anonymous:2002:AIk

[472] Anonymous. Author index. Lecture Notes in Computer Science, 2418: 291–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2418/2418auth.pdf.

Bidoit:2002:GDL

[473] Michel Bidoit, Donald Sannella, and Andrzej Tarlecki. Global development via local observational construction steps. Lecture Notes in Computer Science, 2420:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2420/24200001.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2420/24200001.pdf.

Gibbons:2002:ECP

[474] Alan Gibbons and Paul Sant. Edgecolouring pairs of binary trees: Towards a concise proof of the fourcolour theorem of planar maps. Lecture Notes in Computer Science, 2420:25–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2420/24200025.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2420/24200025.pdf.

Karhumaki:2002:AFA

[475] Juhani Karhumäki. Applications of finite automata. Lecture Notes in Computer Science, 2420:40–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2420/24200040.htm;   http://link.

springer.de/link/service/series/
0558/papers/2420/24200040.pdf.

Karpinski:2002:AMB

[476] Marek Karpinski. Approximability of the minimum bisection problem: An algorithmic challenge. Lecture Notes in Computer Science, 2420:59–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2420/24200059.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2420/24200059.pdf.

Peleg:2002:LSS

[477] David Peleg. Low stretch spanning trees. Lecture Notes in Computer Science, 2420:68–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2420/24200068.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2420/24200068.pdf.

Andreou:2002:RHS

[478] Maria I. Andreou, Dimitris A. Fotakis, Sotiris E. Nikoletseas, Vicky G. Papadopoulou, and Paul G. Spirakis. On radiocoloring hierarchically specified planar graphs: PSPACEcompleteness and approximations. Lecture Notes in Computer Science, 2420:81–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2420/24200081.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2420/24200081.pdf.

Angelsmark:2002:FDC

[479] Ola Angelsmark, Vilhelm Dahllöf, and Peter Jonsson. Finite domain constraint satisfaction using quantum computation. Lecture Notes in Computer Science, 2420: 93–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2420/24200093.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2420/24200093.
                                        pdf.

Bein:2002:FAA

[480] Wolfgang W. Bein, Peter Brucker, Lawrence L. Larmore, and James K. Park. Fast algorithms with algebraic Monge properties. Lecture Notes in Computer Science, 2420: 104–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2420/24200104.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2420/24200104.
                                        pdf.

Beis:2002:PER

[481] Mihalis Beis, William Duckworth, and Michele Zito. Packing edges in random regular graphs. Lecture Notes in Computer Science, 2420: 118–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2420/24200118.htm;   http:

//link.springer.de/link/service/
series/0558/papers/2420/24200118.
pdf.

Bollig:2002:LBT

[482] Beate Bollig and Philipp Woelfel. A lower bound technique for nondeterministic graph-driven read-once-branching programs and its applications. Lecture Notes in Computer Science, 2420:131–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2420/24200131.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2420/24200131.pdf.

Boros:2002:MIP

[483] Endre Boros, Khaled Elbassioni, Vladimir Gurvich, and Leonid Khachiyan. Matroid intersections, polymatroid inequalities, and related problems. Lecture Notes in Computer Science, 2420:143–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2420/24200143.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2420/24200143.pdf.

Carton:2002:AAS

[484] Olivier Carton. Accessibility in automata on scattered linear orderings. Lecture Notes in Computer Science, 2420:155–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2420/24200155.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2420/24200155.pdf.

Caucal:2002:ITH

[485] Didier Caucal. On infinite terms having a decidable monadic theory. Lecture Notes in Computer Science, 2420:165–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2420/24200165.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2420/24200165.pdf.

Caucal:2002:CLH

[486] Didier Caucal and Teodor Knapik. A Chomsky-like hierarchy of infinite graphs. Lecture Notes in Computer Science, 2420:177–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2420/24200177.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2420/24200177.pdf.

Chan:2002:CAL

[487] Wun-Tat Chan, Tak-Wah Lam, Hing-Fung Ting, and Prudence W. H. Wong. Competitive analysis of on-line stream merging algorithms. Lecture Notes in Computer Science, 2420:188–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2420/24200188.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2420/24200188.pdf.

Coja-Oghlan:2002:CCS

[488] Amin Coja-Oghlan. Coloring kcolorable semirandom graphs in polynomial expected time via semidefinite programming. Lecture Notes in Computer Science, 2420:201– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2420/24200201.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2420/24200201.
pdf.

Dabrowski:2002:WEO

[489] Robert Dabrowski and Wojtek Plandowski. On word equations in one variable. Lecture Notes in Computer Science, 2420:212–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2420/24200212.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2420/24200212.pdf.

Ebert:2002:ARS

[490] Todd Ebert and Wolfgang Merkle. Autoreducibility of random sets: a sharp bound on the density of guessed bits. Lecture Notes in Computer Science, 2420:221–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2420/24200221.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2420/24200221.pdf.

Engelfriet:2002:TWF

[491] Joost Engelfriet and Sebastian Maneth. Two-way finite state transducers with nested pebbles. Lecture Notes in Computer Science, 2420:234– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2420/24200234.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2420/24200234.
                                        pdf.

Epstein:2002:ONP

[492] Leah Epstein and Lene M. Favrholdt. Optimal non-preemptive semi-online scheduling on two related machines. Lecture Notes in Computer Science, 2420:245–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2420/24200245.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2420/24200245.pdf.

Epstein:2002:MWS

[493] Leah Epstein, Csanád Imreh, and Rob van Stee. More on weighted servers or F IFO is better than L RU. Lecture Notes in Computer Science, 2420:257–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2420/24200257.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2420/24200257.pdf.

Fishkin:2002:MTM

[494] Aleksei V. Fishkin and Guochuan Zhang. On maximizing the throughput of multiprocessor tasks. Lecture Notes in Computer Science, 2420: 269–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2420/24200269.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2420/24200269.
pdf.

Goerdt:2002:SRR

[495] Andreas Goerdt and Tomasz Jurdziński. Some results on random unsatisfiable k-sat instances and approximation algorithms applied to random structures. Lecture Notes in Computer Science, 2420:280–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2420/24200280.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2420/24200280.pdf.

Groult:2002:ETR

[496] Richard Groult, Martine Léonard, and Laurent Mouchard. Evolutive tandem repeats using Hamming distance. Lecture Notes in Computer Science, 2420:292–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2420/24200292.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2420/24200292.pdf.

Hajiaghayi:2002:SIL

[497] MohammadTaghi Hajiaghayi and Naomi Nishimura. Subgraph isomorphism, log-bounded fragmentation and graphs of (locally) bounded treewidth. Lecture Notes in Computer Science, 2420:305–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2420/24200305.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2420/24200305.pdf.

Hirvensalo:2002:CPI

[498] Mika Hirvensalo and Juhani Karhumäki. Computing partial information out of intractable one — the first digit of 2^{n} at base 3 as an example. Lecture Notes in Computer Science, 2420:319–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2420/24200319.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2420/24200319.pdf.

Ilie:2002:ACS

[499] Lucian Ilie and Sheng Yu. Algorithms for computing small NFAs. Lecture Notes in Computer Science, 2420:328–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2420/24200328.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2420/24200328.pdf.

Inenaga:2002:SEC

[500] Shunsuke Inenaga, Ayumi Shinohara, Masayuki Takeda, Hideo Bannai, and Setsuo Arikawa. Spaceeconomical construction of index structures for all suffixes of a string. Lecture Notes in Computer Science, 2420:341–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2420/24200341.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2420/24200341.pdf.

Iwama:2002:ELB

[501] Kazuo Iwama and Hiroki Morizumi. An explicit lower bound of 5n − o(n) for Boolean circuits. Lecture Notes in Computer Science, 2420: 353–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2420/24200353.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2420/24200353.
pdf.

Iwamoto:2002:CCH

[502] Chuzo Iwamoto, Takeshi Andou, Kenichi Morita, and Katsunobu Imai. Computational complexity in the hyperbolic plane. Lecture Notes in Computer Science, 2420: 365–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/

                                        0558/bibs/2420/24200365.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2420/24200365.
                                        pdf.

Janicki:2002:MSR

[503] Ryszard Janicki. On a mereological system for relational software specifications. Lecture Notes in Computer Science, 2420:375–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2420/24200375.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2420/24200375.pdf.

Johannsen:2002:OLB

[504] Jan Johannsen and N. S. Narayanaswamy. An optimal lower bound for resolution with 2-conjunctions. Lecture Notes in Computer Science, 2420: 387–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2420/24200387.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2420/24200387.
                                        pdf.

Kanj:2002:IPA

[505] Iyad A. Kanj and Ljubomir Perković. Improved parameterized algorithms for planar dominating set. Lecture Notes in Computer Science, 2420: 399–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2420/24200399.htm;   http:
                                        //link.springer.de/link/service/

series/0558/papers/2420/24200399.
pdf.

Kara:2002:OFB

[506] Jan Kára and Daniel Král’. Optimal free binary decision diagrams for computation of EAR_{n}. Lecture Notes in Computer Science, 2420: 411–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2420/24200411.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2420/24200411.
pdf.

Klima:2002:UMA

[507] Ondřej Klíma. Unification modulo associativity and idempotency is NPcomplete. Lecture Notes in Computer Science, 2420:423–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2420/24200423.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2420/24200423.pdf.

Kucera:2002:CSE

[508] Antonín Kučera and Richard Mayr. On the complexity of semantic equivalences for pushdown automata and BPA. Lecture Notes in Computer Science, 2420:433–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2420/24200433.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2420/24200433.pdf.

Kupferman:2002:IAM

[509] Orna Kupferman and Sharon Zuhovitzky. An improved algorithm for the membership problem for extended regular expressions. Lecture Notes in Computer Science, 2420:446– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2420/24200446.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2420/24200446.
                                        pdf.

Lin:2002:EAL

[510] Yaw-Ling Lin, Tao Jiang, and Kun-Mao Chao. Efficient algorithms for locating the length-constrained heaviest segments, with applications to biomolecular sequence analysis. Lecture Notes in Computer Science, 2420:459–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2420/24200459.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2420/24200459.pdf.

Lombardy:2002:DRE

[511] Sylvain Lombardy and Jacques Sakarovitch. Derivation of rational expressions with multiplicity. Lecture Notes in Computer Science, 2420:471– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2420/24200471.htm;   http:
                                        //link.springer.de/link/service/

series/0558/papers/2420/24200471.
pdf.

Loyer:2002:HFS

[512] Yann Loyer and Nicolas Spyratos. Hypothesis-founded semantics for datalog programs with negation. Lecture Notes in Computer Science, 2420:483–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2420/24200483.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2420/24200483.pdf.

Lucking:2002:PSF

[513] Thomas Lücking, Burkhard Monien, and Manuel Rode. On the problem of scheduling flows on distributed networks. Lecture Notes in Computer Science, 2420:495–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2420/24200495.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2420/24200495.pdf.

Machado:2002:UTC

[514] Patricia D. L. Machado and Donald Sannella. Unit testing for C88 ASL architectural specifications. Lecture Notes in Computer Science, 2420:506–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2420/24200506.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2420/24200506.pdf.

Martinelli:2002:SSA

[515] Fabio Martinelli. Symbolic semantics and analysis for crypto-CCS with (almost) generic inference systems. Lecture Notes in Computer Science, 2420:519–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2420/24200519.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2420/24200519.pdf.

Marx:2002:CTM

[516] Dániel Marx. The complexity of tree multicolorings. Lecture Notes in Computer Science, 2420: 532–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2420/24200532.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2420/24200532.
                                        pdf.

Masson:2002:VFL

[517] Benoît Masson and Ph. Schnoebelen. On verifying fair lossy channel systems. Lecture Notes in Computer Science, 2420:543–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2420/24200543.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2420/24200543.pdf.

McCartin:2002:PCP

[518] Catherine McCartin. Parameterized counting problems. Lecture Notes in Computer Science, 2420: 556–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2420/24200556.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2420/24200556.
pdf.

Merkle:2002:CER

[519] Wolfgang Merkle and Nenad Mihailovic. On the construction of effective random sets. Lecture Notes in Computer Science, 2420: 568–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2420/24200568.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2420/24200568.
pdf.

Messner:2002:SSO

[520] Jochen Messner. On the structure of the simulation order of proof systems. Lecture Notes in Computer Science, 2420:581–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2420/24200581.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2420/24200581.pdf.

Mossakowski:2002:CBG

[521] Till Mossakowski. Comorphismbased grothendieck logics. Lecture Notes in Computer Science, 2420: 593–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2420/24200593.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2420/24200593.
pdf.

Richomme:2002:FTS

[522] Gwenael Richomme and Francis Wlazinski. Finite test-sets for overlap-free morphisms. Lecture Notes in Computer Science, 2420: 605–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2420/24200605.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2420/24200605.
pdf.

Rigo:2002:CSR

[523] Michel Rigo. Characterizing simpler recognizable sets of integers. Lecture Notes in Computer Science, 2420:615–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2420/24200615.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2420/24200615.pdf.

Tantau:2002:TCT

[524] Till Tantau. Towards a cardinality theorem for finite automata. Lecture Notes in Computer Science, 2420:625–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2420/24200625.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2420/24200625.pdf.

Villemaire:2002:ASP

[525] Roger Villemaire. An approximation semantics for the propositional mucalculus. Lecture Notes in Computer Science, 2420:637–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2420/24200637.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2420/24200637.pdf.

Anonymous:2002:AIl

[526] Anonymous. Author index. Lecture Notes in Computer Science, 2420: 651–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2420/2420auth.pdf.

Fokkink:2002:RVA

[527] Wan Fokkink, Natalia Ioustinova, Ernst Kesseler, Jaco van de Pol, Yaroslav S. Usenko, and Yuri A. Yushtein. Refinement and verification applied to an in-flight data acquisition unit. Lecture Notes in Computer Science, 2421:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2421/24210001.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2421/24210001.pdf.

Pnueli:2002:DPS

[528] Amir Pnueli and Yonit Kesten. A deductive proof system for CTL. Lecture Notes in Computer Science, 2421:24–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2421/24210024.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2421/24210024.pdf.

Pratt:2002:ESD

[529] Vaughan R. Pratt. Event-state duality: The enriched case. Lecture Notes in Computer Science, 2421: 41–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2421/24210041.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2421/24210041.
                                        pdf.

Rabinovich:2002:EPT

[530] Alexander Rabinovich. Expressive power of temporal logics. Lecture Notes in Computer Science, 2421: 57–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2421/24210057.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2421/24210057.
                                        pdf.

Sangiorgi:2002:TWD

[531] Davide Sangiorgi. Types, or: Where’s the difference between CCS and π ? Lecture Notes in Computer Science, 2421:76–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2421/24210076.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2421/24210076.pdf.

Bradfield:2002:IMT

[532] Julian C. Bradfield. Introduction to modal and temporal mu-calculi (abstract). Lecture Notes in Computer Science, 2421:98–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2421/24210098.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2421/24210098.pdf.

Gordon:2002:TCP

[533] Andrew D. Gordon. Types for cryptographic protocols. Lecture Notes in Computer Science, 2421: 99–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2421/24210099.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2421/24210099.
pdf.

Kesten:2002:NIA

[534] Yonit Kesten, Amir Pnueli, Elad Shahar, and Lenore Zuck. Network invariants in action. Lecture Notes in Computer Science, 2421: 101–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2421/24210101.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2421/24210101.
pdf.

Abdulla:2002:RMC

[535] Parosh Aziz Abdulla, Bengt Jonsson, Marcus Nilsson, and Julien d’Orso. Regular model checking made simple and efficient. Lecture Notes in Computer Science, 2421: 116–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2421/24210116.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2421/24210116.
pdf.

Etessami:2002:HPT

[536] Kousha Etessami. A hierarchy of polynomial-time computable simulations for automata. Lecture Notes in Computer Science, 2421: 131–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2421/24210131.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2421/24210131.
pdf.

Madhusudan:2002:DCA

[537] P. Madhusudan and P. S. Thiagarajan. A decidable class of asynchronous distributed controllers.

Lecture Notes in Computer Science, 2421:145–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2421/24210145.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2421/24210145.pdf.

Valmari:2002:ABS

[538] Antti Valmari and Antti Kervinen. Alphabet-based synchronisation is exponentially cheaper. Lecture Notes in Computer Science, 2421: 161–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2421/24210161.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2421/24210161.
pdf.

Lohrey:2002:SRH

[539] Markus Lohrey. Safe realizability of high-level message sequence charts. Lecture Notes in Computer Science, 2421:177–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2421/24210177.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2421/24210177.pdf.

Asarin:2002:WBB

[540] Eugene Asarin and Gerardo Schneider. Widening the boundary between decidable and undecidable hybrid systems. Lecture Notes in Computer Science, 2421:193–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2421/24210193.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2421/24210193.pdf.

Caires:2002:SLC

[541] Luís Caires and Luca Cardelli. A spatial logic for concurrency (Part II). Lecture Notes in Computer Science, 2421:209–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2421/24210209.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2421/24210209.pdf.

Konikowska:2002:RMC

[542] Beata Konikowska and Wojciech Penczek. Reducing model checking from multi-valued CTL^{∗} to CTL^{∗}. Lecture Notes in Computer Science, 2421:226–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2421/24210226.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2421/24210226.pdf.

Lange:2002:LMC

[543] Martin Lange. Local model checking games for fixed point logic with chop. Lecture Notes in Computer Science, 2421:240–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2421/24210240.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2421/24210240.pdf.

Sorea:2002:DFL

[544] Maria Sorea. A decidable fixpoint logic for time-outs. Lecture Notes in Computer Science, 2421: 255–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2421/24210255.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2421/24210255.
pdf.

Godskesen:2002:CMR

[545] Jens Chr. Godskesen, Thomas Hildebrandt, and Vladimiro Sassone. A calculus of mobile resources. Lecture Notes in Computer Science, 2421:272–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2421/24210272.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2421/24210272.pdf.

Teller:2002:UAC

[546] David Teller, Pascal Zimmer, and Daniel Hirschkoff. Using ambients to control resources. Lecture Notes in Computer Science, 2421: 288–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2421/24210288.htm;   http:
//link.springer.de/link/service/

                                        series/0558/papers/2421/24210288.
                                        pdf.

Merro:2002:TSM

[547] Massimo Merro and Vladimiro Sassone. Typing and subtyping mobility in boxed ambients. Lecture Notes in Computer Science, 2421: 304–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2421/24210304.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2421/24210304.
                                        pdf.

Bruni:2002:OTJ

[548] Roberto Bruni, Cosimo Laneve, and Ugo Montanari. Orchestrating transactions in join calculus. Lecture Notes in Computer Science, 2421:321–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2421/24210321.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2421/24210321.pdf.

Baier:2002:SCT

[549] Christel Baier, Joost-Pieter Katoen, Holger Hermanns, and Boudewijn Haverkort. Simulation for continuoustime Markov chains. Lecture Notes in Computer Science, 2421:338– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2421/24210338.htm;   http:
                                        //link.springer.de/link/service/

series/0558/papers/2421/24210338.
pdf.

Desharnais:2002:WBS

[550] Josée Desharnais, Vineet Gupta, Radha Jagadeesan, and Prakash Panangaden. Weak bisimulation is sound and complete for PCTL*. Lecture Notes in Computer Science, 2421:355–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2421/24210355.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2421/24210355.pdf.

Cattani:2002:DAP

[551] Stefano Cattani and Roberto Segala. Decision algorithms for probabilistic bisimulation. Lecture Notes in Computer Science, 2421:371– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2421/24210371.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2421/24210371.
pdf.

Luttgen:2002:AAS

[552] Gerald Lüttgen and Michael Mendler. Axiomatizing an algebra of step reactions for synchronous languages. Lecture Notes in Computer Science, 2421:386–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2421/24210386.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2421/24210386.pdf.

Fanchon:2002:RSP

[553] Jean Fanchon and Rémi Morin. Regular sets of pomsets with autoconcurrency. Lecture Notes in Computer Science, 2421:402–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2421/24210402.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2421/24210402.pdf.

Gardner:2002:FME

[554] Philippa Gardner, Cosimo Laneve, and Lucian Wischik. The fusion machine (extended abstract). Lecture Notes in Computer Science, 2421:418–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2421/24210418.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2421/24210418.pdf.

Nygaard:2002:HHO

[555] Mikkel Nygaard and Glynn Winskel. HOPLA—A higher-order process language. Lecture Notes in Computer Science, 2421:434–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2421/24210434.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2421/24210434.pdf.

Buscemi:2002:FOC

[556] Maria Grazia Buscemi and Ugo Montanari. A first order coalgebraic model of π-calculus early observational equivalence. Lecture Notes in Computer Science, 2421:449–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2421/24210449.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2421/24210449.pdf.

Brookes:2002:TPF

[557] Stephen Brookes. Traces, pomsets, fairness and full abstraction for communicating processes. Lecture Notes in Computer Science, 2421:466–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2421/24210466.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2421/24210466.pdf.

Boreale:2002:FAS

[558] Michele Boreale and Maria Grazia Buscemi. A framework for the analysis of security protocols. Lecture Notes in Computer Science, 2421: 483–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2421/24210483.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2421/24210483.
pdf.

Amadio:2002:NGS

[559] Roberto M. Amadio and Witold Charatonik. On name generation and set-based analysis in the Dolev–Yao model. Lecture Notes in Computer Science, 2421:499–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2421/24210499.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2421/24210499.pdf.

Kusters:2002:DCP

[560] Ralf Küsters. On the decidability of cryptographic protocols with open-ended data structures. Lecture Notes in Computer Science, 2421:515–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2421/24210515.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2421/24210515.pdf.

Kleijn:2002:CSP

[561] H. C. M. Kleijn and M. Koutny. Causality semantics of petri nets with weighted inhibitor arcs. Lecture Notes in Computer Science, 2421: 531–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2421/24210531.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2421/24210531.
                                        pdf.

Volzer:2002:RRF

[562] Hagen Völzer. Refinement-robust fairness. Lecture Notes in Computer Science, 2421:547–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2421/24210547.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2421/24210547.pdf.

Lasota:2002:DSB

[563] Sławomir Lasota. Decidability of strong bisimilarity for timed BPP. Lecture Notes in Computer Science, 2421:562–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2421/24210562.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2421/24210562.pdf.

Srba:2002:UWB

[564] Jirí Srba. Undecidability of weak bisimilarity for pushdown processes. Lecture Notes in Computer Science, 2421:579–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2421/24210579.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2421/24210579.pdf.

Kucera:2002:WSH

[565] Antonín Kucera and Richard Mayr. Why is simulation harder than bisimulation? Lecture Notes in Computer Science, 2421:594–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2421/24210594.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2421/24210594.pdf.

Anonymous:2002:AIm

[566] Anonymous. Author index. Lecture Notes in Computer Science, 2421: 611–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2421/2421ind.pdf.

Costantini:2002:LPL

[567] Stefania Costantini and Arianna Tocchio. A logic programming language for multi-agent systems. Lecture Notes in Computer Science, 2424:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2424/24240001.pdf.

Ciampolini:2002:PSS

[568] A. Ciampolini, E. Lamma, P. Mello, and P. Torroni. A proof-system for the safe execution of tasks in multi-agent systems. Lecture Notes in Computer Science, 2424:14–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240014.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2424/24240014.pdf.

Amgoud:2002:AFM

[569] Leila Amgoud and Simon Parsons. An argumentation framework for merging conflicting knowledge bases. Lecture Notes in Computer Science, 2424:27–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240027.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2424/24240027.pdf.

Schmidt:2002:MAL

[570] Renate A. Schmidt and Dmitry Tishkovsky. Multi-agent logics of dynamic belief and knowledge. Lecture Notes in Computer Science, 2424:38–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240038.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2424/24240038.pdf.

Alferes:2002:ELP

[571] José Júlio Alferes, Antonio Brogi, João Alexandre Leite, and Luís Moniz Pereira. Evolving logic programs. Lecture Notes in Computer Science, 2424:50–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240050.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2424/24240050.pdf.

Alferes:2002:CUP

[572] José Júlio Alferes, Pierangelo Dell’Acqua, and Luís Moniz Pereira. A compilation of updates plus preferences. Lecture Notes in Computer Science, 2424:62–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2424/24240062.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2424/24240062.pdf.

Bonatti:2002:TSD

[573] Piero A. Bonatti. Towards service description logics. Lecture Notes in Computer Science, 2424: 74–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2424/24240074.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2424/24240074.
                                        pdf.

Giugno:2002:PEP

[574] Rosalba Giugno and Thomas Lukasiewicz. P − SHOQ(D): a probabilistic extension of SHOQ(D) for probabilistic ontologies in the Semantic Web. Lecture Notes in Computer Science, 2424:86–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2424/24240086.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2424/24240086.pdf.

Artale:2002:TDL

[575] Alessandro Artale, Enrico Franconi, Frank Wolter, and Michael Zakharyaschev. A temporal description logic for reasoning over conceptual schemas and queries. Lecture Notes in Computer Science, 2424:98–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240098.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2424/24240098.pdf.

Turner:2002:PLP

[576] Hudson Turner. Polynomial-length planning spans the polynomial hierarchy. Lecture Notes in Computer Science, 2424:111–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240111.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2424/24240111.pdf.

Dekhtyar:2002:CMA

[577] Michael Dekhtyar, Alexander Dikovsky, and Mars Valiev. Complexity of multi-agent systems behavior. Lecture Notes in Computer Science, 2424:125–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240125.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2424/24240125.pdf.

Greco:2002:CAM

[578] Gianluigi Greco and Ester Zumpano. Complexity and algorithms for the matching of bag and set terms. Lecture Notes in Computer Science, 2424:137–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2424/24240137.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2424/24240137.pdf.

Krajci:2002:NCE

[579] S. Krajči, R. Lencses, J. Medina, M. Ojeda-Aciego, A. Valverde, and P. Vojtáš. Non-commutativity and expressive deductive logic databases. Lecture Notes in Computer Science, 2424:149–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2424/24240149.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2424/24240149.pdf.

Beierle:2002:UIS

[580] Christoph Beierle and Gabriele Kern-Isberner. Using institutions for the study of qualitative and quantitative conditional logics. Lecture Notes in Computer Science, 2424: 161–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2424/24240161.htm;   http:

//link.springer.de/link/service/
series/0558/papers/2424/24240161.
pdf.

Dix:2002:TEA

[581] Jürgen Dix, Hector Munoz-Avila, Dana Nau, and Lingling Zhang. Theoretical and empirical aspects of a planner in a multi-agent environment. Lecture Notes in Computer Science, 2424:173–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240173.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2424/24240173.pdf.

Eiter:2002:ASP

[582] Thomas Eiter, Wolfgang Faber, Nicola Leone, Gerald Pfeifer, and Axel Polleres. Answer set planning under action costs. Lecture Notes in Computer Science, 2424:186–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240186.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2424/24240186.pdf.

Governatori:2002:FSB

[583] Guido Governatori, Vineet Padmanabhan, and Abdul Sattar. On fibring semantics for BDI logics. Lecture Notes in Computer Science, 2424:198–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240198.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2424/24240198.pdf.

White:2002:MFM

[584] Graham White. A modal formulation of McCain and Turner’s theory of causal reasoning. Lecture Notes in Computer Science, 2424: 211–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2424/24240211.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2424/24240211.
                                        pdf.

Szalas:2002:SOQ

[585] Andrzej Szalas. Second-order quantifier elimination in modal contexts. Lecture Notes in Computer Science, 2424:223–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2424/24240223.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2424/24240223.pdf.

Amir:2002:ITN

[586] Eyal Amir. Interpolation theorems for nonmonotonic reasoning systems. Lecture Notes in Computer Science, 2424:233–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2424/24240233.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2424/24240233.pdf.

Iwanuma:2002:MAC

[587] Koji Iwanuma and Katsumi Inoue. Minimal answer computation and SOL. Lecture Notes in Computer Science, 2424:245–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240245.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2424/24240245.pdf.

Montanari:2002:DIT

[588] Angelo Montanari, Guido Sciavicco, and Nicola Vitacolonna. Decidability of interval temporal logics over split-frames via granularity. Lecture Notes in Computer Science, 2424:259–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240259.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2424/24240259.pdf.

Loyer:2002:UPN

[589] Yann Loyer and Umberto Straccia. Uncertainty and partial non-uniform assumptions in parametric deductive databases. Lecture Notes in Computer Science, 2424:271–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240271.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2424/24240271.pdf.

El-Sayed:2002:QRN

[590] Mazen El-Sayed and Daniel Pacholczyk. A qualitative reasoning with nuanced information. Lecture Notes in Computer Science, 2424: 283–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2424/24240283.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2424/24240283.
                                        pdf.

Giunchiglia:2002:DIV

[591] Enrico Giunchiglia, Marco Maratea, and Armando Tacchella. Dependent and independent variables in propositional satisfiability. Lecture Notes in Computer Science, 2424: 296–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2424/24240296.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2424/24240296.
                                        pdf.

Tinelli:2002:DBC

[592] Cesare Tinelli. A DPLL-based calculus for ground satisfiability modulo theories. Lecture Notes in Computer Science, 2424:308–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2424/24240308.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2424/24240308.pdf.

Besnard:2002:PRQ

[593] Philippe Besnard, Torsten Schaub, Hans Tompits, and Stefan Woltran. Paraconsistent reasoning via quantified Boolean formulas, I: Axiomatising signed systems. Lecture Notes in Computer Science, 2424: 320–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2424/24240320.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2424/24240320.
pdf.

Konieczny:2002:TVL

[594] Sébastien Konieczny and Pierre Marquis. Three-valued logics for inconsistency handling. Lecture Notes in Computer Science, 2424: 332–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2424/24240332.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2424/24240332.
pdf.

Alcantara:2002:PLP

[595] João Alcântara, Carlos Viegas Damásio, and Luís Moniz Pereira. Paraconsistent logic programs. Lecture Notes in Computer Science, 2424: 345–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2424/24240345.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2424/24240345.
pdf.

Zhang:2002:IPA

[596] Dongmo Zhang and Norman Foo. Interpolation properties of action logic: Lazy-formalization to the frame problem. Lecture Notes in Computer Science, 2424:357–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2424/24240357.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2424/24240357.pdf.

Son:2002:RAA

[597] Tran Cao Son and Enrico Pontelli. Reasoning about actions in prioritized default theory. Lecture Notes in Computer Science, 2424: 369–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2424/24240369.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2424/24240369.
                                        pdf.

Giordano:2002:TCL

[598] Laura Giordano and Camilla Schwind. Towards a conditional logic of actions and causation. Lecture Notes in Computer Science, 2424:382– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2424/24240382.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2424/24240382.
                                        pdf.

Hansen:2002:ANC

[599] Helle Hvid Hansen and Marc Pauly. Axiomatising Nash-consistent coalition logic. Lecture Notes in Computer Science, 2424:394–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240394.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2424/24240394.pdf.

Wallace:2002:RPR

[600] Richard J. Wallace. Representing possibilities in relation to constraints and agents. Lecture Notes in Computer Science, 2424:407–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240407.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2424/24240407.pdf.

Sadri:2002:ALP

[601] Fariba Sadri, Francesca Toni, and Paolo Torroni. An abductive logic programming architecture for negotiating agents. Lecture Notes in Computer Science, 2424:419–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240419.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2424/24240419.pdf.

VanNieuwenborgh:2002:PAS

[602] Davy Van Nieuwenborgh and Dirk Vermeir. Preferred answer sets for ordered logic programs. Lecture Notes in Computer Science, 2424: 432–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2424/24240432.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2424/24240432.
pdf.

Brewka:2002:IOD

[603] Gerhard Brewka, Ilkka Niemelä, and Tommi Syrjänen. Implementing ordered disjunction using answer set solvers for normal programs. Lecture Notes in Computer Science, 2424:444–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240444.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2424/24240444.pdf.

Rondogiannis:2002:IVS

[604] Panos Rondogiannis and William W. Wadge. An infinite-valued semantics for logic programs with negation. Lecture Notes in Computer Science, 2424:456–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240456.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2424/24240456.pdf.

Linke:2002:MN

[605] Thomas Linke, Christian Anger, and Kathrin Konczak. More on noMoRe.

Lecture Notes in Computer Science, 2424:468–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240468.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2424/24240468.pdf.

Nicolas:2002:ASP

[606] Pascal Nicolas, Frédéric Saubion, and Igor Stéphan. Answer set programming by ant colony optimization. Lecture Notes in Computer Science, 2424:481–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240481.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2424/24240481.pdf.

Janhunen:2002:TEL

[607] Tomi Janhunen and Emilia Oikarinen. Testing the equivalence of logic programs under stable model semantics. Lecture Notes in Computer Science, 2424:493–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240493.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2424/24240493.pdf.

Haenni:2002:ARA

[608] Rolf Haenni. Argumentative reasoning with ABEL. Lecture Notes in Computer Science, 2424:505– ??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2424/24240505.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2424/24240505.
                                        pdf.

Delgrande:2002:CCB

[609] James P. Delgrande, Aaron Hunter, and Torsten Schaub. COBA: a consistency-based belief revision system. Lecture Notes in Computer Science, 2424:509–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2424/24240509.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2424/24240509.pdf.

Finkel:2002:CLP

[610] Raphael Finkel, Victor W. Marek, and Mirosław Truszczyński. Constraint lingo: a program for solving logic puzzles and other tabular constraint problems. Lecture Notes in Computer Science, 2424:513–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2424/24240513.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2424/24240513.pdf.

Giannotti:2002:IDM

[611] Fosca Giannotti and Giuseppe Manco. LDL-M_{⟩\⌉}: Integrating data mining with intelligent query answering. Lecture Notes in Computer Science, 2424:517–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240517.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2424/24240517.pdf.

Anger:2002:NNM

[612] Christian Anger, Kathrin Konczak, and Thomas Linke. NoMoRe: Nonmonotonic reasoning with logic programs. Lecture Notes in Computer Science, 2424:521–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240521.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2424/24240521.pdf.

Schimm:2002:PMT

[613] Guido Schimm. Process Miner — A tool for mining process schemes from event-based data. Lecture Notes in Computer Science, 2424: 525–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2424/24240525.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2424/24240525.
pdf.

Estruch:2002:SMP

[614] V. Estruch, C. Ferri, J. Hernández, and M. J. Ramírez. SMILES: a multi-purpose learning system. Lecture Notes in Computer Science, 2424:529–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2424/24240529.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2424/24240529.pdf.

East:2002:AS

[615] Deborah East and Mirosław Truszczyński. The aspps system. Lecture Notes in Computer Science, 2424:533– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2424/24240533.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2424/24240533.
                                        pdf.

Leone:2002:DS

[616] Nicola Leone, Gerald Pfeifer, Wolfgang Faber, Francesco Calimeri, Tina Dell’Armi, Thomas Eiter, Georg Gottlob, Giovambattista Ianni, Giuseppe Ielpa, Christoph Koch, Simona Perri, and Axel Polleres. The DLV system. Lecture Notes in Computer Science, 2424:537–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2424/24240537.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2424/24240537.pdf.

Eiter:2002:DPS

[617] Thomas Eiter, Wolfgang Faber, Nicola Leone, Gerald Pfeifer, and Axel Polleres. The DLV^{K} planning system: Progress report. Lecture Notes in Computer Science, 2424:

541–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2424/24240541.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2424/24240541.
pdf.

Minock:2002:ZLA

[618] Michael Minock and Hansi Kraus. Z-log: Applying system-Z. Lecture Notes in Computer Science, 2424: 545–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2424/24240545.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2424/24240545.
pdf.

Eiter:2002:HTC

[619] Thomas Eiter and Georg Gottlob. Hypergraph transversal computation and related problems in logic and AI. Lecture Notes in Computer Science, 2424:549–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2424/24240549.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2424/24240549.pdf.

Vardi:2002:A

[620] Moshe Y. Vardi. Alternation. Lecture Notes in Computer Science, 2424:565–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2424/24240565.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2424/24240565.pdf.

Gelfond:2002:UAC

[621] Michael Gelfond. The USA-Advisor: a case study in answer set programming. Lecture Notes in Computer Science, 2424:566–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2424/24240566.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2424/24240566.pdf.

Anonymous:2002:T

[622] Anonymous.  Tutorials.  Lecture

Notes in Computer Science, 2424: 569–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2424/24240569.pdf.

Anonymous:2002:AIn

[623] Anonymous. Author index. Lecture Notes in Computer Science, 2424: 571–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2424/2424auth.pdf.

Huchard:2002:MSG

[624] Marianne Huchard, Hernan Astudillo, and Petko Valtchev. MAnaging SPEcialization/Generalization HIerarchies. Lecture Notes in Computer Science, 2426:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2426/24260001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2426/24260001.pdf.

Ducournau:2002:RWA

[625] Roland Ducournau. “real world” as an argument for covariant specialization in programming and modeling. Lecture Notes in Computer Science, 2426:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2426/24260003.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2426/24260003.pdf.

Berry:2002:MCM

[626] Anne Berry and Alain Sigayret. Maintaining class membership information. Lecture Notes in Computer Science, 2426:13–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2426/24260013.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2426/24260013.pdf.

Marcos:2002:HOO

[627] Esperanza Marcos and Jose María Cavero. Hierarchies in object oriented conceptual modeling. Lecture Notes in Computer Science, 2426: 24–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/

                                        0558/bibs/2426/24260024.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2426/24260024.
                                        pdf.

Bekaert:2002:SGO

[628] Pieter Bekaert, Geert Delanote, Frank Devos, and Eric Steegmans. Specialization/generalization in objectoriented analysis: Strengthening and multiple partitioning. Lecture Notes in Computer Science, 2426: 34–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2426/24260034.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2426/24260034.
                                        pdf.

Coulondre:2002:TNR

[629] Stéphane Coulondre and Thérèse Libourel. Towards a new role paradigm for object-oriented modeling. Lecture Notes in Computer Science, 2426:44–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2426/24260044.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2426/24260044.pdf.

Arevalo:2002:AOO

[630] Gabriela Arévalo and Tom Mens. Analysing object-oriented application frameworks using concept analysis. Lecture Notes in Computer Science, 2426:53–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2426/24260053.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2426/24260053.pdf.

Crescenzo:2002:UBS

[631] Pierre Crescenzo and Philippe Lahire. Using both specialisation and generalisation in a programming language: Why and how? Lecture Notes in Computer Science, 2426: 64–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2426/24260064.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2426/24260064.
pdf.

Llorens:2002:AGH

[632] Juan Lloréns and Hernán Astudillo. Automatic generation of hierarchical taxonomies from free text using linguistic algorithms. Lecture Notes in Computer Science, 2426: 74–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2426/24260074.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2426/24260074.
pdf.

Lafourcade:2002:GHS

[633] Mathieu Lafourcade. Guessing hierarchies and symbols for word meanings through hyperonyms and conceptual vectors. Lecture Notes in Computer Science, 2426:84–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2426/24260084.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2426/24260084.pdf.

Bardou:2002:ROO

[634] Daniel Bardou, Agnès Conte, and Liz Kendall. Reuse in objectoriented information systems design. Lecture Notes in Computer Science, 2426:94–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2426/24260094.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2426/24260094.pdf.

Silveira:2002:SRU

[635] Maria Clara Silveira and Raul Moreira Vidal. Software reuse with use case patterns. Lecture Notes in Computer Science, 2426:96–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2426/24260096.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2426/24260096.pdf.

Estrella:2002:PRT

[636] Florida Estrella, Sebastien Gaspard, Zsolt Kovacs, Jean-Marie Le Goff, and Richard McClatchey. Promoting reuse through the capture of system description. Lecture Notes in Computer Science, 2426:101–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2426/24260101.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2426/24260101.pdf.

Kantorowitz:2002:SOF

[637] Eliezer Kantorowitz and Sally Tadmor. A specification-oriented framework for information system user interfaces. Lecture Notes in Computer Science, 2426:112–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2426/24260112.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2426/24260112.pdf.

Braga:2002:RPL

[638] Rosana T. V. Braga and Paulo Cesar Masiero. The role of pattern languages in the instantiation of object-oriented frameworks. Lecture Notes in Computer Science, 2426:122–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2426/24260122.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2426/24260122.pdf.

Turki:2002:CH

[639] Slim Turki and Michel Léonard. IS components with hyperclasses. Lecture Notes in Computer Science, 2426:132–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2426/24260132.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2426/24260132.pdf.

Blanc:2002:CSR

[640] Mathieu Blanc, Fabien Costantini, Sébastien Dubois, Manuel Forget, Olivier Francillon, and Christian Toinard. Collaborative simulation by reuse of COTS simulators with a reflexive XML middleware1. Lecture Notes in Computer Science, 2426:142–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2426/24260142.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2426/24260142.pdf.

Boucelma:2002:EWB

[641] Omar Boucelma and Zoé Lacroix. Efficient Web-based information systems. Lecture Notes in Computer Science, 2426:152–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2426/24260152.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2426/24260152.pdf.

Beneventano:2002:SIQ

[642] Domenico Beneventano, Sonia Bergamaschi, Silvana Castano, Valeria De Antonellis, Alfio Ferrara, Francesco Guerra, Federica Mandreoli, Giorgio Carlo Ornetti, and Maurizio Vincini. Semantic integration and query optimization of heterogeneous data sources. Lecture Notes in Computer Science, 2426:154–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/link/service/
series/0558/bibs/2426/24260154.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2426/24260154.pdf.

Hemnani:2002:EIS

[643] Ajay Hemnani and Stephane Bressan. Extracting information from semi-structured Web documents. Lecture Notes in Computer Science, 2426: 166–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2426/24260166.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2426/24260166.
pdf.

Katchaounov:2002:OOM

[644] Timour Katchaounov, Tore Risch, and Simon Zürcher. Object-oriented mediator queries to Internet search engines. Lecture Notes in Computer Science, 2426:176–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2426/24260176.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2426/24260176.pdf.

He:2002:WEO

[645] Haiyun He and Curtis Dyreson. Warp-edge optimization in XPath. Lecture Notes in Computer Science, 2426:187–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2426/24260187.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2426/24260187.pdf.

Turau:2002:CSW

[646] Volker Turau. A caching system for Web content generated from XML sources using XSLT. Lecture Notes in Computer Science, 2426:197–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2426/24260197.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2426/24260197.pdf.

Zaiane:2002:FSQ

[647] Osmar R. Zaïane and Alexander Strilets. Finding similar queries to satisfy searches based on query traces. Lecture Notes in Computer Science, 2426:207–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2426/24260207.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2426/24260207.pdf.

Synodinos:2002:WEM

[648] Dionysios G. Synodinos and Paris Avgeriou. WOnDA: An extensible multi-platform hypermedia design model. Lecture Notes in Computer Science, 2426:217–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2426/24260217.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2426/24260217.pdf.

Turk:2002:MDA

[649] Dan Turk, Robert France, Bernhard Rumpe, and Geri Georg. Modeldriven approaches to software development. Lecture Notes in Computer Science, 2426:229–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2426/24260229.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2426/24260229.pdf.

Baar:2002:ESC

[650] Thomas Baar. Executable and symbolic conformance tests for implementation models (position paper). Lecture Notes in Computer Science, 2426:231–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2426/24260231.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2426/24260231.pdf.

Clark:2002:OOT

[651] Tony Clark, Andy Evans, and Robert France. Object-oriented theories for model driven architecture. Lecture Notes in Computer Science, 2426:235–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2426/24260235.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2426/24260235.pdf.

Denno:2002:SEF

[652] Peter Denno and Allison Barnard Feeney. Systems engineering foundations of software systems integration. Lecture Notes in Computer Science, 2426:245–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2426/24260245.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2426/24260245.pdf.

Gerard:2002:UMP

[653] Sébastien Gérard, François Terrier, and Yann Tanguy. Using the model paradigm for real-time systems development: ACCORD/UML. Lecture Notes in Computer Science, 2426:260–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2426/24260260.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2426/24260260.pdf.

Kulkarni:2002:GEA

[654] Vinay Kulkarni, R. Venkatesh, and Sreedhar Reddy. Generating enterprise applications from models. Lecture Notes in Computer Science, 2426:270–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2426/24260270.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2426/24260270.pdf.

Mekerke:2002:TSA

[655] François Mekerke, Geri Georg, and Robert France. Tool support for aspect-oriented design. Lecture Notes in Computer Science, 2426: 280–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2426/24260280.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2426/24260280.
pdf.

Mellor:2002:MDA

[656] Stephen J. Mellor, Kendall Scott, Axel Uhl, and Dirk Weise. Model-driven architecture. Lecture Notes in Computer Science, 2426:290–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2426/24260290.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2426/24260290.pdf.

Schatz:2002:MBD

[657] B. Schätz, A. Pretschner, F. Huber, and J. Philipps. Modelbased development of embedded systems. Lecture Notes in Computer Science, 2426:298–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2426/24260298.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2426/24260298.pdf.

Anonymous:2002:AIo

[658] Anonymous. Author index. Lecture Notes in Computer Science, 2426: 313–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2426/2426ind.pdf.

Clayton:2002:WRI

[659] Richard Clayton. Workshop report for IPTPS’02: 1st International Workshop on Peer-to-Peer Systems. Lecture Notes in Computer Science, 2429:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2429/24290001.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2429/24290001.pdf.

Liben-Nowell:2002:ODE

[660] David Liben-Nowell, Hari Balakrishnan, and David Karger. Observations on the dynamic evolution of peer-to-peer networks. Lecture Notes in Computer Science, 2429: 22–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2429/24290022.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2429/24290022.
                                        pdf.

Zhao:2002:BLR

[661] Ben Y. Zhao, Yitao Duan, Ling Huang, Anthony D. Joseph, and John D. Kubiatowicz. Brocade: Landmark routing on overlay networks. Lecture Notes in Computer Science, 2429:34–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2429/24290034.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2429/24290034.pdf.

Ratnasamy:2002:RAD

[662] Sylvia Ratnasamy, Ion Stoica, and Scott Shenker. Routing algorithms for DHTs: Some open questions. Lecture Notes in Computer Science, 2429:45–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2429/24290045.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2429/24290045.pdf.

Maymounkov:2002:KPP

[663] Petar Maymounkov and David Mazières. Kademlia: a peer-to-peer information system based on the XOR metric. Lecture Notes in Computer Science, 2429:53–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2429/24290053.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2429/24290053.pdf.

Freedman:2002:EPP

[664] Michael J. Freedman and Radek Vingralek. Efficient peer-to-peer lookup based on a distributed trie. Lecture Notes in Computer Science, 2429:66–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2429/24290066.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2429/24290066.pdf.

Vahdat:2002:SOS

[665] Amin Vahdat, Jeff Chase, Rebecca Braynard, Dejan Kostic, Patrick Reynolds, and Adolfo Rodriguez. Selforganizing subsets: From each according to his abilities, to each according to his needs. Lecture Notes in Computer Science, 2429:76–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2429/24290076.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2429/24290076.pdf.

Ripeanu:2002:MGN

[666] Matei Ripeanu and Ian Foster. Mapping the gnutella network: Macroscopic properties of large-scale peer-to-peer systems. Lecture Notes in Computer Science, 2429:85–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2429/24290085.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2429/24290085.pdf.

Lv:2002:CHM

[667] Qin Lv, Sylvia Ratnasamy, and Scott Shenker. Can heterogeneity make gnutella scalable? Lecture Notes in Computer Science, 2429: 94–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2429/24290094.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2429/24290094.
pdf.

Wilcox-OHearn:2002:EDL

[668] Bryce Wilcox-O’Hearn. Experiences deploying a large-scale emergent network. Lecture Notes in Computer Science, 2429:104–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2429/24290104.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2429/24290104.pdf.

Serjantov:2002:ACR

[669] Andrei Serjantov. Anonymizing censorship resistant systems. Lecture Notes in Computer Science, 2429: 111–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2429/24290111.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2429/24290111.
pdf.

Freedman:2002:ITP

[670] Michael J. Freedman, Emil Sit, Josh Cates, and Robert Morris. Introducing tarzan, a peer-to-peer anonymizing network layer. Lecture Notes in Computer Science, 2429: 121–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2429/24290121.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2429/24290121.
pdf.

Hand:2002:MPP

[671] Steven Hand and Timothy Roscoe. Mnemosyne: Peer-to-peer steganographic storage. Lecture Notes in Computer Science, 2429:130– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2429/24290130.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2429/24290130.
pdf.

Ajmani:2002:CCS

[672] Sameer Ajmani, Dwaine E. Clarke, Chuang-Hue Moh, and Steven Richman. ConChord: Cooperative SDSI certificate storage and name resolution. Lecture Notes in Computer Science, 2429:141–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2429/24290141.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2429/24290141.pdf.

Cox:2002:SDU

[673] Russ Cox, Athicha Muthitacharoen, and Robert T. Morris. Serving DNS using a peer-to-peer lookup service. Lecture Notes in Computer Science, 2429:155–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2429/24290155.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2429/24290155.pdf.

Srinivasan:2002:NMC

[674] Sridhar Srinivasan and Ellen Zegura. Network measurement as a cooperative enterprise. Lecture Notes in Computer Science, 2429:166–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2429/24290166.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2429/24290166.pdf.

Padmanabhan:2002:CCN

[675] Venkata N. Padmanabhan and Kunwadee Sripanidkulchai. The case for cooperative networking. Lecture Notes in Computer Science, 2429: 178–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2429/24290178.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2429/24290178.
pdf.

Stoica:2002:III

[676] Ion Stoica, Daniel Adkins, Sylvia Ratnasamy, Scott Shenker, Sonesh Surana, and Shelley Zhuang. Internet indirection infrastructure. Lecture Notes in Computer Science, 2429: 191–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2429/24290191.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2429/24290191.
                                        pdf.

Stading:2002:PPC

[677] Tyron Stading, Petros Maniatis, and Mary Baker. Peer-to-peer caching schemes to address flash crowds. Lecture Notes in Computer Science, 2429:203–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2429/24290203.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2429/24290203.pdf.

Saroiu:2002:EDS

[678] Stefan Saroiu, P. Krishna Gummadi, and Steven D. Gribble. Exploring the design space of distributed and peer-to-peer systems: Comparing the Web, TRIAD, and Chord/ CFS. Lecture Notes in Computer Science, 2429:214–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2429/24290214.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2429/24290214.pdf.

Keleher:2002:VON

[679] Pete Keleher, Bobby Bhattacharjee, and Bujor Silaghi. Are virtualized overlay networks too much of a good thing? Lecture Notes in Computer Science, 2429:225–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2429/24290225.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2429/24290225.pdf.

Iamnitchi:2002:LDS

[680] Adriana Iamnitchi, Matei Ripeanu, and Ian Foster. Locating data in (small-world?) peer-to-peer scientific collaborations. Lecture Notes in Computer Science, 2429: 232–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2429/24290232.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2429/24290232.
pdf.

Harren:2002:CQD

[681] Matthew Harren, Joseph M. Hellerstein, Ryan Huebsch, Boon Thau Loo, Scott Shenker, and Ion Stoica. Complex queries in DHT-based peer-to-peer networks. Lecture Notes in Computer Science, 2429:242–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2429/24290242.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2429/24290242.pdf.

Douceur:2002:SA

[682] John R. Douceur. The sybil attack. Lecture Notes in Computer Science, 2429:251–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2429/24290251.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2429/24290251.pdf.

Sit:2002:SCP

[683] Emil Sit and Robert Morris. Security considerations for peer-to-peer distributed hash tables. Lecture Notes in Computer Science, 2429: 261–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2429/24290261.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2429/24290261.
                                        pdf.

Saia:2002:DFT

[684] Jared Saia, Amos Fiat, Steve Gribble, Anna R. Karlin, and Stefan Saroiu. Dynamically faulttolerant content addressable networks. Lecture Notes in Computer Science, 2429:270–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2429/24290270.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2429/24290270.pdf.

vanRenesse:2002:SMD

[685] Robbert van Renesse, Kenneth Birman, Dan Dumitriu, and Werner Vogels. Scalable management and data mining using astrolabe. Lecture Notes in Computer Science, 2429: 280–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2429/24290280.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2429/24290280.
pdf.

Lynch:2002:ADA

[686] Nancy Lynch, Dahlia Malkhi, and David Ratajczak. Atomic data access in distributed hash tables. Lecture Notes in Computer Science, 2429:295–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2429/24290295.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2429/24290295.pdf.

Chen:2002:DRP

[687] Yan Chen, Randy H. Katz, and John D. Kubiatowicz. Dynamic replica placement for scalable content delivery. Lecture Notes in Computer Science, 2429:306–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2429/24290306.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2429/24290306.pdf.

Cooper:2002:PPR

[688] Brian F. Cooper and Hector Garcia-Molina. Peer-to-peer resource trading in a reliable distributed system. Lecture Notes in Computer Science, 2429:319–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2429/24290319.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2429/24290319.pdf.

Weatherspoon:2002:ECV

[689] Hakim Weatherspoon and John D. Kubiatowicz. Erasure coding vs. replication: a quantitative comparison. Lecture Notes in Computer Science, 2429:328–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2429/24290328.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2429/24290328.pdf.

Anonymous:2002:AIp

[690] Anonymous. Author index. Lecture Notes in Computer Science, 2429: 339–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2429/2429indx.pdf.

Banerjee:2002:CGA

[691] Bikramjit Banerjee and Jing Peng. Convergent gradient ascent in generalsum games. Lecture Notes in Computer Science, 2430:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2430/24300001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2430/24300001.pdf.

Bay:2002:REM

[692] Stephen D. Bay, Daniel G. Shapiro, and Pat Langley. Revising engineering models: Combining computational discovery with knowledge. Lecture Notes in Computer Science, 2430:10–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2430/24300010.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2430/24300010.pdf.

Buntine:2002:VEE

[693] Wray Buntine. Variational extensions to EM and multinomial PCA. Lecture Notes in Computer Science, 2430:23–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2430/24300023.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2430/24300023.pdf.

Carreras:2002:LIC

[694] Xavier Carreras, Lluís Màrquez, Vasin Punyakanok, and Dan Roth. Learning and inference for clause identification. Lecture Notes in Computer Science, 2430:35–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2430/24300035.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2430/24300035.pdf.

Dai:2002:ESE

[695] Honghua Dai, Gang Li, and Yiqing Tu. An empirical study of encoding schemes and search strategies in discovering causal networks. Lecture Notes in Computer Science, 2430:48–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2430/24300048.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2430/24300048.pdf.

Derbeko:2002:VOB

[696] Philip Derbeko, Ran El-Yaniv, and Ron Meir. Variance optimized bagging. Lecture Notes in Computer Science, 2430:60–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2430/24300060.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2430/24300060.pdf.

Eibl:2002:HMA

[697] Günther Eibl and Karl Peter Pfeiffer. How to make AdaBoost.M1 work for weak base classifiers by changing only one line of the code. Lecture Notes in Computer Science, 2430:72–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2430/24300072.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2430/24300072.pdf.

Engel:2002:SOG

[698] Yaakov Engel, Shie Mannor, and Ron Meir. Sparse online greedy support vector regression. Lecture Notes in Computer Science, 2430: 84–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2430/24300084.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2430/24300084.
pdf.

Furnkranz:2002:PCE

[699] Johannes Fürnkranz. Pairwise classification as an ensemble technique. Lecture Notes in Computer Science, 2430:97–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2430/24300097.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2430/24300097.pdf.

Gora:2002:RCC

[700] Grzegorz Góra and Arkadiusz Wojna. RIONA: a classifier combining rule induction and k-NN method with automated selection of optimal neighbourhood. Lecture Notes in Computer Science, 2430:111–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2430/24300111.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2430/24300111.pdf.

Halck:2002:UHC

[701] Ole Martin Halck. Using hard classifiers to estimate conditional class probabilities. Lecture Notes in Computer Science, 2430:124–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2430/24300124.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2430/24300124.pdf.

Harris:2002:EID

[702] Harlan D. Harris. Evidence that incremental delta-bar-delta is an attributeefficient linear learner. Lecture Notes in Computer Science, 2430: 135–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2430/24300135.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2430/24300135.
                                        pdf.

Hoche:2002:SBM

[703] Susanne Hoche and Stefan Wrobel. Scaling boosting by marginbased inclusionof features and relations. Lecture Notes in Computer Science, 2430:148–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2430/24300148.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2430/24300148.pdf.

Holmes:2002:MAD

[704] Geoffrey Holmes, Bernhard Pfahringer, Richard Kirkby, Eibe Frank, and Mark Hall. Multiclass alternating decision trees. Lecture Notes in Computer Science, 2430:161–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2430/24300161.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2430/24300161.pdf.

Hullermeier:2002:PID

[705] Eyke Hüllermeier. Possibilistic induction in decision-tree learning. Lecture Notes in Computer Science, 2430:173–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2430/24300173.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2430/24300173.pdf.

Kermorvant:2002:ISP

[706] Christopher Kermorvant and Pierre Dupont. Improved smoothing for probabilistic suffix trees seen as variable order Markov chains. Lecture Notes in Computer Science, 2430: 185–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2430/24300185.htm;   http:
//link.springer.de/link/service/

                                        series/0558/papers/2430/24300185.
                                        pdf.

Klink:2002:CLT

[707] Stefan Klink, Armin Hust, Markus Junker, and Andreas Dengel. Collaborative learning of term-based concepts for automatic query expansion. Lecture Notes in Computer Science, 2430:195–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2430/24300195.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2430/24300195.pdf.

Kraakenes:2002:LPH

[708] Tony Kråkenes and Ole Martin Halck. Learning to play a highly complex game from human expert games. Lecture Notes in Computer Science, 2430:207–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2430/24300207.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2430/24300207.pdf.

Kukar:2002:RCM

[709] Matjaz Kukar and Igor Kononenko. Reliable classifications with machine learning. Lecture Notes in Computer Science, 2430:219–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2430/24300219.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2430/24300219.pdf.

Kushmerick:2002:RAI

[710] Nicholas Kushmerick. Robustness analyses of instance-based collaborative recommendation. Lecture Notes in Computer Science, 2430: 232–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2430/24300232.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2430/24300232.
pdf.

Kwek:2002:BBU

[711] Stephen Kwek and Chau Nguyen. iboost: Boosting using an instancebased exponential weighting scheme. Lecture Notes in Computer Science, 2430:245–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2430/24300245.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2430/24300245.pdf.

Ludl:2002:TSC

[712] Marcus-Christopher Ludl and Gerhard Widmer. Towards a simple clustering criterion based on minimum length encoding. Lecture Notes in Computer Science, 2430:258–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2430/24300258.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2430/24300258.pdf.

Margineantu:2002:CPE

[713] Dragos D. Margineantu. Class probability estimation and cost-sensitive classification decisions. Lecture Notes in Computer Science, 2430: 270–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2430/24300270.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2430/24300270.
pdf.

Martin:2002:LSV

[714] Mario Martin. On-line support vector machine regression. Lecture Notes in Computer Science, 2430: 282–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2430/24300282.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2430/24300282.
pdf.

Menache:2002:QCD

[715] Ishai Menache, Shie Mannor, and Nahum Shimkin. Q-cut — dynamic discovery of sub-goals in reinforcement learning. Lecture Notes in Computer Science, 2430:295–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2430/24300295.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2430/24300295.pdf.

Morik:2002:MAC

[716] Katharina Morik and Stefan Rüping. A multistrategy approach to the classification of phases in business cycles. Lecture Notes in Computer Science, 2430:307–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2430/24300307.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2430/24300307.pdf.

Nock:2002:RBA

[717] Richard Nock and Patrice Lefaucheur. A robust boosting algorithm. Lecture Notes in Computer Science, 2430: 319–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2430/24300319.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2430/24300319.
pdf.

Ontanon:2002:CES

[718] Santiago Ontañón and Enric Plaza. Case exchange strategies in multiagent learning. Lecture Notes in Computer Science, 2430:331–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2430/24300331.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2430/24300331.pdf.

Papadopoulos:2002:ICM

[719] Harris Papadopoulos, Kostas Proedrou, Volodya Vovk, and Alex Gammerman. Inductive confidence machines for regression. Lecture Notes in Computer Science, 2430: 345–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2430/24300345.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2430/24300345.
pdf.

Castillo:2002:MOM

[720] Lourdes Peña Castillo and Stefan Wrobel. Macro-operators in multirelational learning: a search-space reduction technique. Lecture Notes in Computer Science, 2430:357–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2430/24300357.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2430/24300357.pdf.

Preux:2002:PQV

[721] Philippe Preux. Propagation of Qvalues in tabular TD(λ). Lecture Notes in Computer Science, 2430: 369–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2430/24300369.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2430/24300369.
pdf.

Proedrou:2002:TCM

[722] Kostas Proedrou, Ilia Nouretdinov, Volodya Vovk, and Alex Gammerman. Transductive confidence machines for pattern recognition. Lecture Notes in Computer Science, 2430:381–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2430/24300381.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2430/24300381.pdf.

Ratitch:2002:CMD

[723] Bohdana Ratitch and Doina Precup. Characterizing Markov decision processes. Lecture Notes in Computer Science, 2430:391–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2430/24300391.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2430/24300391.pdf.

Ruckert:2002:PTS

[724] Ulrich Rückert, Stefan Kramer, and Luc De Raedt. Phase transitions and stochastic local search in k-term DNF learning. Lecture Notes in Computer Science, 2430:405–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2430/24300405.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2430/24300405.pdf.

Sinkkonen:2002:DCO

[725] Janne Sinkkonen, Samuel Kaski, and Janne Nikkilä. Discriminative clustering: Optimal contingency tables by learning metrics.

Lecture Notes in Computer Science, 2430:418–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2430/24300418.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2430/24300418.pdf.

Thollard:2002:BDF

[726] Franck Thollard, Marc Sebban, and Philippe Ezequel. Boosting density function estimators. Lecture Notes in Computer Science, 2430:431–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2430/24300431.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2430/24300431.pdf.

Todorovski:2002:RPC

[727] Ljupco Todorovski, Hendrik Blockeel, and Sašo Dzeroski. Ranking with predictive clustering trees. Lecture Notes in Computer Science, 2430:444–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2430/24300444.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2430/24300444.pdf.

Tsochantaridis:2002:SVM

[728] Ioannis Tsochantaridis and Thomas Hofmann. Support vector machines for polycategorical classification. Lecture Notes in Computer Science, 2430:456–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2430/24300456.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2430/24300456.pdf.

Vittaut:2002:LCB

[729] Jean-Noël Vittaut, Massih-Reza Amini, and Patrick Gallinari. Learning classification with both labeled and unlabeled data. Lecture Notes in Computer Science, 2430:468–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2430/24300468.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2430/24300468.pdf.

Yeang:2002:IGP

[730] Chen-Hsiang Yeang. An information geometric perspective on active learning. Lecture Notes in Computer Science, 2430:480–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2430/24300480.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2430/24300480.pdf.

Zenko:2002:SES

[731] Bernard Zenko and Sašo Dzeroski. Stacking with an extended set of meta-level attributes and MLR. Lecture Notes in Computer Science, 2430:493–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2430/24300493.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2430/24300493.pdf.

Oja:2002:FHFa

[732] Erkki Oja. Finding hidden factors using independent component analysis. Lecture Notes in Computer Science, 2430:505–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2430/24300505.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2430/24300505.pdf.

Roth:2002:RCa

[733] Dan Roth. Reasoning with classifiers. Lecture Notes in Computer Science, 2430:506–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2430/24300506.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2430/24300506.pdf.

Scholkopf:2002:KALa

[734] Bernhard Schölkopf, Jason Weston, Eleazar Eskin, Christina Leslie, and William Stafford Noble. A kernel approach for learning from almost orthogonal patterns. Lecture Notes in Computer Science, 2430: 511–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/

0558/bibs/2430/24300511.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2430/24300511.
pdf.

Smyth:2002:LMMa

[735] Padhraic Smyth. Learning with mixture models: Concepts and applications. Lecture Notes in Computer Science, 2430:529–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2430/24300529.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2430/24300529.pdf.

Anonymous:2002:AIq

[736] Anonymous. Author index. Lecture Notes in Computer Science, 2430: 531–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2430/2430ind.pdf.

Abe:2002:OSD

[737] Kenji Abe, Shinji Kawasoe, Tatsuya Asai, Hiroki Arimura, and Setsuo Arikawa. Optimized substructure discovery for semi-structured data. Lecture Notes in Computer Science, 2431:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2431/24310001.pdf.

Angiulli:2002:FOD

[738] Fabrizio Angiulli and Clara Pizzuti. Fast outlier detection in high dimensional spaces. Lecture Notes in Computer Science, 2431: 15–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2431/24310015.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2431/24310015.
pdf.

Arnborg:2002:DMS

[739] Stefan Arnborg, Ingrid Agartz, Håkan Hall, Erik Jönsson, Anna Sillén, and Göran Sedvall. Data mining in schizophrenia research — preliminary analysis. Lecture Notes in Computer Science, 2431:27–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310027.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2431/24310027.pdf.

Bailey:2002:FAM

[740] James Bailey, Thomas Manoukian, and Kotagiri Ramamohanarao. Fast algorithms for mining emerging patterns. Lecture Notes in Computer Science, 2431:39–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310039.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2431/24310039.pdf.

Berberidis:2002:DWP

[741] Christos Berberidis, Ioannis Vlahavas, Walid G. Aref, Mikhail Atallah, and Ahmed K. Elmagarmid. On the discovery of weak periodicities in large time series. Lecture Notes in Computer Science, 2431:51–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310051.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2431/24310051.pdf.

Brain:2002:NLB

[742] Damien Brain and Geoffrey I. Webb. The need for low bias algorithms in classification learning from large data sets. Lecture Notes in Computer Science, 2431:62–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310062.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2431/24310062.pdf.

Calders:2002:MAN

[743] Toon Calders and Bart Goethals. Mining all non-derivable frequent itemsets. Lecture Notes in Computer Science, 2431:74–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310074.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2431/24310074.pdf.

Choki:2002:IDS

[744] Yuta Choki and Einoshin Suzuki. Iterative data squashing for boosting based on a distribution-sensitive distance. Lecture Notes in Computer Science, 2431:86–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310086.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2431/24310086.pdf.

Coenen:2002:FAR

[745] Frans Coenen and Paul Leng. Finding association rules with some very frequent attributes. Lecture Notes in Computer Science, 2431: 99–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2431/24310099.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2431/24310099.
pdf.

Ding:2002:ULS

[746] Chris Ding, Xiaofeng He, Hongyuan Zha, and Horst Simon. Unsupervised learning: Self-aggregation in scaled principal component space. Lecture Notes in Computer Science, 2431:112–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310112.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2431/24310112.pdf.

Domeniconi:2002:CAP

[747] Carlotta Domeniconi, Chang shing Perng, Ricardo Vilalta, and Sheng Ma. A classification approach for prediction of target events in temporal sequences. Lecture Notes in Computer Science, 2431:125–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310125.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2431/24310125.pdf.

Felty:2002:POD

[748] Amy Felty and Stan Matwin. Privacyoriented data mining by proof checking. Lecture Notes in Computer Science, 2431:138–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310138.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2431/24310138.pdf.

Forman:2002:CYW

[749] George Forman. Choose your words carefully: An empirical study of feature selection metrics for text classification. Lecture Notes in Computer Science, 2431:150–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310150.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2431/24310150.pdf.

Gamberger:2002:GAK

[750] Dragan Gamberger and Nada Lavrac. Generating actionable knowledge by expert-guided subgroup discovery. Lecture Notes in Computer Science, 2431:163–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2431/24310163.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2431/24310163.pdf.

Giannotti:2002:CTD

[751] Fosca Giannotti, Cristian Gozzi, and Giuseppe Manco. Clustering transactional data. Lecture Notes in Computer Science, 2431:175–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2431/24310175.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2431/24310175.pdf.

Hirano:2002:MCT

[752] Shoji Hirano and Shusaku Tsumoto. Multiscale comparison of temporal patterns in time-series medical databases. Lecture Notes in Computer Science, 2431:188–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2431/24310188.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2431/24310188.pdf.

Hullermeier:2002:ARE

[753] Eyke Hüllermeier. Association rules for expressing gradual dependencies. Lecture Notes in Computer Science, 2431:200–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310200.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2431/24310200.pdf.

Jaroszewicz:2002:SAU

[754] Szymon Jaroszewicz and Dan A. Simovici. Support approximations using Bonferroni-type inequalities. Lecture Notes in Computer Science, 2431:212–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310212.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2431/24310212.pdf.

Jeudy:2002:UCR

[755] Baptiste Jeudy and Jean-François Boulicaut. Using condensed representations for interactive association rule mining. Lecture Notes in Computer Science, 2431:225–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310225.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2431/24310225.pdf.

Joshi:2002:PRC

[756] Mahesh V. Joshi, Ramesh C. Agarwal, and Vipin Kumar. Predicting rare classes: Comparing two-phase rule induction to cost-sensitive boosting. Lecture Notes in Computer Science, 2431:237–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310237.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2431/24310237.pdf.

Kargupta:2002:DDM

[757] Hillol Kargupta, Krishnamoorthy Sivakumar, and Samiran Ghosh. Dependency detection in MobiMine and random matrices. Lecture Notes in Computer Science, 2431: 250–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2431/24310250.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2431/24310250.
pdf.

Kemp:2002:LTL

[758] Charles Kemp and Kotagiri Ramamohanarao. Long-term learning for Web search engines. Lecture Notes in Computer Science, 2431: 263–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2431/24310263.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2431/24310263.
pdf.

Klosgen:2002:SSM

[759] Willi Klösgen and Michael May. Spatial subgroup mining integrated in an object-relational spatial database. Lecture Notes in Computer Science, 2431:275–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310275.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2431/24310275.pdf.

Knobbe:2002:IAF

[760] Arno J. Knobbe, Arno Siebes, and Bart Marseille. Involving aggregate functions in multi-relational search. Lecture Notes in Computer Science, 2431:287–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310287.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2431/24310287.pdf.

Kosala:2002:IES

[761] Raymond Kosala, Jan Van den Bussche, Maurice Bruynooghe, and Hendrik Blockeel. Information extraction in structured documents using tree automata induction. Lecture Notes in Computer Science, 2431: 299–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2431/24310299.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2431/24310299.
pdf.

Koyuturk:2002:ATA

[762] Mehmet Koyutürk, Ananth Grama, and Naren Ramakrishnan. Algebraic techniques for analysis of large discrete-valued datasets. Lecture Notes in Computer Science, 2431: 311–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2431/24310311.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2431/24310311.
                                        pdf.

Li:2002:GDB

[763] Jinyan Li and Limsoon Wong. Geography of differences between two classes of data. Lecture Notes in Computer Science, 2431:325–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2431/24310325.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2431/24310325.pdf.

Liden:2002:RIC

[764] Per Lidén, Lars Asker, and Henrik Boström. Rule induction for classification of gene expression array data. Lecture Notes in Computer Science, 2431:338–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2431/24310338.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2431/24310338.pdf.

Maedche:2002:COB

[765] Alexander Maedche and Valentin Zacharias. Clustering ontologybased metadata in the Semantic Web. Lecture Notes in Computer Science, 2431:348–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310348.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2431/24310348.pdf.

Mamitsuka:2002:ISF

[766] Hiroshi Mamitsuka. Iteratively selecting feature subsets for mining from high-dimensional databases. Lecture Notes in Computer Science, 2431:361–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310361.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2431/24310361.pdf.

Paass:2002:SCU

[767] Gerhard Paaß, Edda Leopold, Martha Larson, Jörg Kindermann, and Stefan Eickeler. SVM classification using sequences of phonemes and syllables. Lecture Notes in Computer Science, 2431:373–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310373.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2431/24310373.pdf.

Park:2002:NWT

[768] Laurence A. F. Park, Marimuthu Palaniswami, and Kotagiri Ramamohanarao. A novel Web text mining method using the discrete cosine transform. Lecture Notes in Computer Science, 2431:385–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2431/24310385.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2431/24310385.pdf.

Scheffer:2002:SCM

[769] Tobias Scheffer and Stefan Wrobel. A scalable constant-memory sampling algorithm for pattern discovery in large databases. Lecture Notes in Computer Science, 2431:397–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2431/24310397.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2431/24310397.pdf.

Sese:2002:AMC

[770] Jun Sese and Shinichi Morishita. Answering the most correlated N association rules efficiently. Lecture Notes in Computer Science, 2431: 410–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2431/24310410.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2431/24310410.
                                        pdf.

Tsumoto:2002:MHD

[771] Shusaku Tsumoto. Mining hierarchical decision rules from clinical databases using rough sets and medical diagnostic model. Lecture Notes in Computer Science, 2431:423–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310423.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2431/24310423.pdf.

Veloso:2002:EMA

[772] Adriano Veloso, Bruno Gusmão, Wagner Meira Jr., Marcio Carvalho, Srini Parthasarathy, and Mohammed Zaki. Efficiently mining approximate models of associations in evolving databases. Lecture Notes in Computer Science, 2431:435–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310435.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2431/24310435.pdf.

Wall:2002:EPN

[773] Robert Wall, Pádraig Cunningham, and Paul Walsh. Explaining predictions from a neural network ensemble one at a time. Lecture Notes in Computer Science, 2431:449–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310449.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2431/24310449.pdf.

Winkler:2002:SDS

[774] Karsten Winkler and Myra Spiliopoulou. Structuring domain-specific text archives by deriving a probabilistic XML DTD. Lecture Notes in Computer Science, 2431:461–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2431/24310461.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2431/24310461.pdf.

Zighed:2002:SIS

[775] Djamel A. Zighed, Stéphane Lallich, and Fabrice Muhlenbach. Separability index in supervised learning. Lecture Notes in Computer Science, 2431:475–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2431/24310475.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2431/24310475.pdf.

Oja:2002:FHFb

[776] Erkki Oja. Finding hidden factors using independent component analysis. Lecture Notes in Computer Science, 2431:488–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2431/24310488.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2431/24310488.pdf.

Roth:2002:RCb

[777] Dan Roth. Reasoning with classifiers. Lecture Notes in Computer Science, 2431:489–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310489.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2431/24310489.pdf.

Scholkopf:2002:KALb

[778] Bernhard Schölkopf, Jason Weston, Eleazar Eskin, Christina Leslie, and William Stafford Noble. A kernel approach for learning from almost orthogonal patterns. Lecture Notes in Computer Science, 2431: 494–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2431/24310494.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2431/24310494.
pdf.

Smyth:2002:LMMb

[779] Padhraic Smyth. Learning with mixture models: Concepts and applications. Lecture Notes in Computer Science, 2431:512–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2431/24310512.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2431/24310512.pdf.

Anonymous:2002:AIr

[780] Anonymous. Author index. Lecture Notes in Computer Science, 2431: 513–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2431/2431ind.pdf.

Bologna:2002:HCS

[781] Sandro Bologna and Erik Hollnagel. Human-computer system dependability. Lecture Notes in Computer Science, 2434:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2434/24340001.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2434/24340001.pdf.

Hollnagel:2002:DJH

[782] Erik Hollnagel. Dependability of joint human-computer systems. Lecture Notes in Computer Science, 2434: 4–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2434/24340004.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2434/24340004.
                                        pdf.

Servida:2002:DIS

[783] Andrea Servida. Dependability in the information society: Getting ready for the FP6. Lecture Notes in Computer Science, 2434:10–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2434/24340010.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2434/24340010.pdf.

Bredereke:2002:RVM

[784] Jan Bredereke and Axel Lankenau. A rigorous view of mode confusion. Lecture Notes in Computer Science, 2434:19–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2434/24340019.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2434/24340019.pdf.

Voss:2002:DOA

[785] Alexander Voß, Roger Slack, Rob Procter, Robin Williams, Mark Hartswood, and Mark Rouncefield. Dependability as ordinary action. Lecture Notes in Computer Science, 2434:32–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2434/24340032.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2434/24340032.pdf.

Rhee:2002:PSK

[786] Yoon-Jung Rhee and Tai-Yun Kim. Practical solutions to key recovery based on PKI in IP security. Lecture Notes in Computer Science, 2434:44–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2434/24340044.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2434/24340044.pdf.

Droste:2002:RDA

[787] Thomas Droste. Redundant data acquisition in a distributed security compound. Lecture Notes in Computer Science, 2434:53–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2434/24340053.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2434/24340053.pdf.

Dafelmair:2002:SSS

[788] Ferdinand J. Dafelmair. Survivability strategy for a security critical process. Lecture Notes in Computer Science, 2434:61–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2434/24340061.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2434/24340061.pdf.

Heidtmann:2002:SCT

[789] Klaus Heidtmann. Statistical comparison of two sum-of-disjoint-product algorithms for reliability and safety evaluation. Lecture Notes in Computer Science, 2434:70–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2434/24340070.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2434/24340070.pdf.

Lano:2002:SSA

[790] Kevin Lano, David Clark, and Kelly Androutsopoulos. Safety and security analysis of object-oriented models. Lecture Notes in Computer Science, 2434:82–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2434/24340082.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2434/24340082.pdf.

Fredriksen:2002:CFM

[791] Rune Fredriksen, Monica Kristiansen, Bjørn Axel Gran, Ketil Stølen, Tom Arthur Opperud, and Theo Dimitrakos. The CORAS framework for a model-based risk management process. Lecture Notes in Computer Science, 2434:94–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2434/24340094.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2434/24340094.pdf.

Knight:2002:SCA

[792] John C. Knight. Software challenges in aviation systems. Lecture Notes in Computer Science, 2434: 106–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2434/24340106.htm;   http:
//link.springer.de/link/service/

                                        series/0558/papers/2434/24340106.
                                        pdf.

Zhang:2002:SIE

[793] Wenhui Zhang. A strategy for improving the efficiency of procedure verification. Lecture Notes in Computer Science, 2434:113–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2434/24340113.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2434/24340113.pdf.

Benerecetti:2002:VST

[794] Massimo Benerecetti, Maurizio Panti, Luca Spalazzi, and Simone Tacconi. Verification of the SSL/TLS protocol using a model checkable logic of belief and time. Lecture Notes in Computer Science, 2434:126–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2434/24340126.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2434/24340126.pdf.

Popov:2002:RAL

[795] Peter Popov. Reliability assessment of legacy safety-critical systems upgraded with off-the-shelf components. Lecture Notes in Computer Science, 2434:139–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2434/24340139.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2434/24340139.pdf.

Chen:2002:ABR

[796] Luping Chen, John May, and Gordon Hughes. Assessment of the benefit of redundant systems. Lecture Notes in Computer Science, 2434: 151–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2434/24340151.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2434/24340151.
pdf.

Bishop:2002:ERF

[797] Peter G. Bishop. Estimating residual faults from code coverage. Lecture Notes in Computer Science, 2434:163–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2434/24340163.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2434/24340163.pdf.

Jacobs:2002:TMB

[798] Jef Jacobs and Jos Trienekens. Towards a metrics based verification and validation maturity model. Lecture Notes in Computer Science, 2434:175–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2434/24340175.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2434/24340175.pdf.

Paynter:2002:ASS

[799] Stephen E. Paynter and Bob W. Born. Analysing the safety of a software development process. Lecture Notes in Computer Science, 2434:186–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2434/24340186.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2434/24340186.pdf.

Bishop:2002:SCA

[800] Peter Bishop, Robin Bloomfield, Tim Clement, and Sofia Guerra. Software criticality analysis of COTS/ SOUP. Lecture Notes in Computer Science, 2434:198–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2434/24340198.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2434/24340198.pdf.

Bobbio:2002:MIM

[801] Andrea Bobbio, Ester Ciancamerla, Giuliana Franceschinis, Rossano Gaeta, Michele Minichino, and Luigi Portinale. Methods of increasing modelling power for safety analysis, applied to a turbine digital control system. Lecture Notes in Computer Science, 2434:212–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2434/24340212.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2434/24340212.pdf.

RomanideOliveira:2002:CST

[802] Ítalo Romani de Oliveira and Paulo Sérgio Cugnasca. Checking safe trajectories of aircraft using hybrid automata. Lecture Notes in Computer Science, 2434:224–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2434/24340224.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2434/24340224.pdf.

Papadopoulos:2002:MBL

[803] Yiannis Papadopoulos. Model-based on-line monitoring using a state sensitive fault propagation model. Lecture Notes in Computer Science, 2434:236–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2434/24340236.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2434/24340236.pdf.

Littlewood:2002:DEI

[804] Bev Littlewood. On diversity, and the elusiveness of independence. Lecture Notes in Computer Science, 2434:249–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2434/24340249.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2434/24340249.pdf.

Mohajerani:2002:ANN

[805] MahdiReza Mohajerani and Ali Moeini. An approach to a new network security architecture for academic environments. Lecture Notes in Computer Science, 2434:252–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2434/24340252.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2434/24340252.pdf.

Rodriguez:2002:WPA

[806] Francisco Rodríguez, José Carlos Campelo, and Juan José Serrano. A watchdog processor architecture with minimal performance overhead. Lecture Notes in Computer Science, 2434:261–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2434/24340261.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2434/24340261.pdf.

Gribaudo:2002:MCB

[807] Marco Gribaudo, A. Horváth, A. Bobbio, Enrico Tronci, Ester Ciancamerla, and Michele Minichino. Modelchecking based on fluid petri nets for the temperature control system of the ICARO co-generative plant. Lecture Notes in Computer Science, 2434:273–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2434/24340273.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2434/24340273.pdf.

Sharma:2002:ACE

[808] Babita Sharma, S. D. Dhodapkar, and S. Ramesh. Assertion checking environment (ACE) for formal verification of C programs. Lecture Notes in Computer Science, 2434:284–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2434/24340284.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2434/24340284.pdf.

Ortmeier:2002:SAH

[809] Frank Ortmeier, Gerhard Schellhorn, Andreas Thums, Wolfgang Reif, Bernhard Hering, and Helmut Trappschuh. Safety analysis of the height control system for the elbtunnel. Lecture Notes in Computer Science, 2434:296–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2434/24340296.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2434/24340296.pdf.

Saridakis:2002:DCP

[810] Titos Saridakis. Dependability and configurability: Partners or competitors in pervasive computing? Lecture Notes in Computer Science, 2434:309–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2434/24340309.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2434/24340309.pdf.

Bate:2002:ACC

[811] Iain Bate and Tim Kelly. Architectural considerations in the certification of modular systems. Lecture Notes in Computer Science, 2434:321–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2434/24340321.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2434/24340321.pdf.

Rottke:2002:POA

[812] Thomas Rottke, Denis Hatebur, Maritta Heisel, and Monika Heiner. A problem-oriented approach to common criteria certification. Lecture Notes in Computer Science, 2434: 334–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2434/24340334.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2434/24340334.
                                        pdf.

Anonymous:2002:AIs

[813] Anonymous. Author index. Lecture Notes in Computer Science, 2434: 347–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2434/2434ind.pdf.

Armington:2002:BAI

[814] John Armington, Purdy Ho, Paul Koznek, and Richard Martinez. Biometric authentication in infrastructure security. Lecture Notes in Computer Science, 2437:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2437/24370001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2437/24370001.pdf.

Rila:2002:DAB

[815] Luciano Rila. Denial of access in biometrics-based authentication systems. Lecture Notes in Computer Science, 2437:19–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2437/24370019.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2437/24370019.pdf.

Blundo:2002:NAP

[816] Carlo Blundo, Paolo D’Arco, Alfredo De Santis, and Clemente Galdi. A novel approach to proactive password checking. Lecture Notes in Computer Science, 2437:30–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2437/24370030.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2437/24370030.pdf.

Clercq:2002:SSA

[817] Jan De Clercq. Single sign-on architectures. Lecture Notes in Computer Science, 2437:40–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2437/24370040.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2437/24370040.pdf.

Mont:2002:ADC

[818] Marco Casassa Mont and Richard Brown. Active digital credentials: Dynamic provision of up-to-date identity information. Lecture Notes in Computer Science, 2437: 59–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2437/24370059.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2437/24370059.
                                        pdf.

Schechter:2002:HBB

[819] Stuart Schechter. How to buy better testing. Lecture Notes in Computer Science, 2437:73–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2437/24370073.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2437/24370073.pdf.

McEvoy:2002:SRA

[820] Neil McEvoy and Andrew Whitcombe. Structured risk analysis. Lecture Notes in Computer Science, 2437:88–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2437/24370088.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2437/24370088.pdf.

Siougle:2002:MEL

[821] E. S. Siougle and V. C. Zorkadis. A model enabling law compliant privacy protection through the selection and evaluation of appropriate security controls. Lecture Notes in Computer Science, 2437:104–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2437/24370104.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2437/24370104.pdf.

Reddy:2002:AAM

[822] Prakash Reddy, Venky Krishnan, Kan Zhang, and Devaraj Das. Authentication and authorization of mobile clients in public data networks. Lecture Notes in Computer Science, 2437:115–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2437/24370115.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2437/24370115.pdf.

Pagliusi:2002:CFG

[823] Paulo S. Pagliusi. A contemporary foreword on GSM security.

Lecture Notes in Computer Science, 2437:129–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2437/24370129.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2437/24370129.pdf.

Kim:2002:VAS

[824] HyungJong Kim, KyungHee Koh, DongHoon Shin, and HongGeun Kim. Vulnerability assessment simulation for information infrastructure protection. Lecture Notes in Computer Science, 2437:145–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2437/24370145.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2437/24370145.pdf.

Flegel:2002:PUL

[825] Ulrich Flegel. Pseudonymizing Unix log files. Lecture Notes in Computer Science, 2437:162–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2437/24370162.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2437/24370162.pdf.

Fenkam:2002:DAS

[826] Pascal Fenkam, Harald Gall, Mehdi Jazayeri, and Christopher Kruegel. DPS: An architectural style for development of secure software. Lecture Notes in Computer Science, 2437:180–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2437/24370180.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2437/24370180.pdf.

Enzmann:2002:NIU

[827] Matthias Enzmann, Thomas Kunz, and Markus Schneider. A new infrastructure for user tracking prevention and privacy protection in Internet shopping. Lecture Notes in Computer Science, 2437:199–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2437/24370199.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2437/24370199.pdf.

Canovas:2002:DSB

[828] Óscar Cánovas, Antonio F. Gómez, Humberto Martínez, and Gregorio Martínez. Different Smartcard-based approaches to physical access control. Lecture Notes in Computer Science, 2437:214–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2437/24370214.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2437/24370214.pdf.

Gurgens:2002:APF

[829] Sigrid Gürgens, Peter Ochsenschläger, and Carsten Rudolph. Authenticity and provability — A formal framework. Lecture Notes in Computer Science, 2437:227–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2437/24370227.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2437/24370227.pdf.

Lopez:2002:PEA

[830] Javier Lopez, Juan J. Ortega, and Jose M. Troya. Protocol engineering applied to formal analysis of security systems. Lecture Notes in Computer Science, 2437:246–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2437/24370246.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2437/24370246.pdf.

Chen:2002:AMT

[831] L. Chen, K. Harrison, D. Soldera, and N. P. Smart. Applications of multiple trust authorities in pairing based cryptosystems. Lecture Notes in Computer Science, 2437:260–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2437/24370260.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2437/24370260.pdf.

Chapman:2002:PDU

[832] Mark Chapman and George Davida. Plausible deniability using automated linguistic stegonagraphy. Lecture Notes in Computer Science, 2437:

276–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2437/24370276.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2437/24370276.
pdf.

Kwon:2002:VST

[833] Taekyoung Kwon. Virtual software tokens — A practical way to secure PKI roaming. Lecture Notes in Computer Science, 2437:288–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2437/24370288.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2437/24370288.pdf.

Kim:2002:BSA

[834] Hyun-Sung Kim and Kee-Young Yoo. Bit-serial AOP arithmetic architectures over GF(2^{m}). Lecture Notes in Computer Science, 2437: 303–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2437/24370303.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2437/24370303.
pdf.

Adamson:2002:PDA

[835] William A. Adamson and Olga Kornievskaia. A practical distributed authorization system for GARA. Lecture Notes in Computer Science, 2437:314–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2437/24370314.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2437/24370314.pdf.

Lopez:2002:DVS

[836] Javier Lopez, Jose A. Montenegro, Rodrigo Roman, and Jorge Davila. Design of a VPN software solution integrating TCP and UDP services. Lecture Notes in Computer Science, 2437:325–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2437/24370325.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2437/24370325.pdf.

Anonymous:2002:AIt

[837] Anonymous. Author index. Lecture Notes in Computer Science, 2437: 339–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2437/2437auth.pdf.

Arnold:2002:RDO

[838] Dirk V. Arnold and Hans-Georg Beyer. Random dynamics optimum tracking with evolution strategies. Lecture Notes in Computer Science, 2439:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390003.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2439/24390003.pdf.

Eriksson:2002:BEG

[839] Roger Eriksson and Björn Olsson. On the behavior of evolutionary globallocal hybrids with dynamic fitness functions. Lecture Notes in Computer Science, 2439:13–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390013.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390013.pdf.

vanHemert:2002:MSS

[840] Jano I. van Hemert and Thomas Bäck. Measuring the searched space to guide efficiency: The principle and evidence on constraint satisfaction. Lecture Notes in Computer Science, 2439:23–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390023.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390023.pdf.

Jansen:2002:ADR

[841] Thomas Jansen. On the analysis of dynamic restart strategies for evolutionary algorithms. Lecture Notes in Computer Science, 2439:33–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390033.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2439/24390033.pdf.

Laumanns:2002:RTA

[842] Marco Laumanns, Lothar Thiele, Eckart Zitzler, Emo Welzl, and Kalyanmoy Deb. Running time analysis of multi-objective evolutionary algorithms on a simple discrete optimization problem. Lecture Notes in Computer Science, 2439: 44–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2439/24390044.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2439/24390044.
                                        pdf.

Scharnow:2002:FLB

[843] Jens Scharnow, Karsten Tinnefeld, and Ingo Wegener. Fitness landscapes based on sorting and shortest paths problems. Lecture Notes in Computer Science, 2439:54–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390054.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390054.pdf.

Weicker:2002:PMD

[844] Karsten Weicker. Performance measures for dynamic environments. Lecture Notes in Computer Science, 2439:64–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390064.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390064.pdf.

Eckert:2002:DRV

[845] Christoph Eckert and Jens Gottlieb. Direct representation and variation operators for the fixed charge transportation problem. Lecture Notes in Computer Science, 2439: 77–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2439/24390077.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2439/24390077.
pdf.

Knowles:2002:URE

[846] Joshua D. Knowles and Richard A. Watson. On the utility of redundant encodings in mutation-based evolutionary search. Lecture Notes in Computer Science, 2439:88–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390088.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390088.pdf.

Rothlauf:2002:BRI

[847] Franz Rothlauf. Binary representations of integers and the performance of selectorecombinative genetic algorithms. Lecture Notes in Computer Science, 2439:99–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390099.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390099.pdf.

Aguirre:2002:PVM

[848] Hernán E. Aguirre and Kiyoshi Tanaka. Parallel varying mutation in deterministic and self-adaptive GAs. Lecture Notes in Computer Science, 2439:111–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390111.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390111.pdf.

Buche:2002:SOM

[849] Dirk Büche, Gianfranco Guidati, Peter Stoll, and Petros Koumoutsakos. Self-organizing maps for Pareto optimization of airfoils. Lecture Notes in Computer Science, 2439: 122–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2439/24390122.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2439/24390122.
                                        pdf.

Corne:2002:FDE

[850] David W. Corne, Martin J. Oates, and Douglas B. Kell. On fitness distributions and expected fitness gain of mutation rates in parallel evolutionary algorithms. Lecture Notes in Computer Science, 2439:132–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2439/24390132.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390132.pdf.

Dolin:2002:OAC

[851] Brad Dolin, M. G. Arenas, and J. J. Merelo. Opposites attract: Complementary phenotype selection for crossover in genetic programming. Lecture Notes in Computer Science, 2439:142–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390142.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390142.pdf.

Hervas-Martinez:2002:TAC

[852] C. Hervás-Martínez, D. Ortiz-Boyer, and N. García-Pedrajas. Theoretical analysis of the confidence interval based crossover for real-coded genetic algorithms. Lecture Notes in Computer Science, 2439:153–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390153.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390153.pdf.

Ikeda:2002:DMS

[853] Kokolo Ikeda and Shigenobu Kobayashi. Deterministic multi-step crossover fusion: a handy crossover composition for GAs. Lecture Notes in Computer Science, 2439:162–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390162.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390162.pdf.

Jelasity:2002:OLP

[854] Márk Jelasity, Mike Preuß, and A. E. Eiben. Operator learning for a problem class in a distributed peer-to-peer environment. Lecture Notes in Computer Science, 2439:172–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390172.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390172.pdf.

Ortiz-Boyer:2002:COE

[855] D. Ortiz-Boyer, C. Hervás-Martínez, and N. García-Pedrajas. Crossover operator effect in function optimization with constraints. Lecture Notes in Computer Science, 2439: 184–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2439/24390184.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2439/24390184.
                                        pdf.

Runarsson:2002:RRF

[856] Thomas Philip Runarsson. Reducing random fluctuations in mutative selfadaptation. Lecture Notes in Computer Science, 2439:194–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2439/24390194.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390194.pdf.

Raidl:2002:WBM

[857] Günther R. Raidl, Gabriele Kodydek, and Bryant A. Julstrom. On weight-biased mutation for graph problems. Lecture Notes in Computer Science, 2439:204–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390204.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390204.pdf.

Thomsen:2002:SAO

[858] René Thomsen and Thiemo Krink. Self-adaptive operator scheduling using the religion-based EA. Lecture Notes in Computer Science, 2439: 214–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2439/24390214.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2439/24390214.
pdf.

Tsutsui:2002:PMB

[859] Shigeyoshi Tsutsui. Probabilistic model-building genetic algorithms in permutation representation domain using edge histogram. Lecture Notes in Computer Science, 2439: 224–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/

                                        0558/bibs/2439/24390224.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2439/24390224.
                                        pdf.

Wiesmann:2002:SSM

[860] Dirk Wiesmann. From syntactical to semantical mutation operators for structure optimization. Lecture Notes in Computer Science, 2439:234–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390234.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390234.pdf.

Iorio:2002:PCW

[861] Antony Iorio and Xiaodong Li. Parameter control within a co-operative co-evolutionary genetic algorithm. Lecture Notes in Computer Science, 2439:247–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390247.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390247.pdf.

Wiegand:2002:ERB

[862] R. Paul Wiegand, William C. Liles, and Kenneth A. De Jong. The effects of representational bias on collaboration methods in cooperative coevolution. Lecture Notes in Computer Science, 2439:257–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2439/24390257.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390257.pdf.

Jozefowiez:2002:PHM

[863] Nicolas Jozefowiez, Frédéric Semet, and El-Ghazali Talbi. Parallel and hybrid models for multiobjective optimization: Application to the vehicle routing problem. Lecture Notes in Computer Science, 2439:271–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390271.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390271.pdf.

Kanazaki:2002:MDO

[864] Masahiro Kanazaki, Masashi Morikaw, Shigeru Obayashi, and Kazuhiro Nakahashi. Multiobjective design optimization of merging configuration for an exhaust manifold of a car engine. Lecture Notes in Computer Science, 2439:281–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390281.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390281.pdf.

Keerativuttitumrong:2002:MOC

[865] Nattavut Keerativuttitumrong, Nachol Chaiyaratana, and Vara Varavithya. Multi-objective co-operative co-evolutionary genetic algorithm. Lecture Notes in Computer Science, 2439:288–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390288.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390288.pdf.

Laumanns:2002:BOA

[866] Marco Laumanns and Jiri Ocenasek. Bayesian optimization algorithms for multi-objective optimization. Lecture Notes in Computer Science, 2439:298–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390298.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390298.pdf.

Richter:2002:EAC

[867] Hendrik Richter. An evolutionary algorithm for controlling chaos: The use of multi-objective fitness functions. Lecture Notes in Computer Science, 2439:308–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390308.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390308.pdf.

Adamidis:2002:MEA

[868] Panagiotis Adamidis and Vasilios Petridis. On modelling evolutionary algorithm implementations through co-operating populations. Lecture Notes in Computer Science, 2439:321–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390321.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390321.pdf.

Bosman:2002:POI

[869] Peter A. N. Bosman and Dirk Thierens. Permutation optimization by iterated estimation of random keys marginal product factorizations. Lecture Notes in Computer Science, 2439:331–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390331.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390331.pdf.

Burke:2002:APD

[870] Edmund Burke, Steven Gustafson, Graham Kendall, and Natalio Krasnogor. Advanced population diversity measures in genetic programming. Lecture Notes in Computer Science, 2439:341–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390341.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390341.pdf.

Chen:2002:ISE

[871] Ying ping Chen and David E. Goldberg. Introducing start expression genes to the linkage learning genetic algorithm. Lecture Notes in Computer Science, 2439:351–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390351.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390351.pdf.

Emmerich:2002:MAE

[872] Michael Emmerich, Alexios Giotis, Mutlu Özdemir, Thomas Bäck, and Kyriakos Giannakoglou. Metamodelassisted evolution strategies. Lecture Notes in Computer Science, 2439: 361–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2439/24390361.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2439/24390361.
                                        pdf.

Giacobini:2002:LNF

[873] Mario Giacobini, Marco Tomassini, and Leonardo Vanneschi. Limiting the number of fitness cases in genetic programming using statistics. Lecture Notes in Computer Science, 2439:371–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390371.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390371.pdf.

Horn:2002:RBF

[874] Jeffrey Horn. Resource-based fitness sharing. Lecture Notes in Computer Science, 2439:381–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390381.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390381.pdf.

Huhse:2002:ESN

[875] Jutta Huhse, Thomas Villmann, Peter Merz, and Andreas Zell. Evolution strategy with neighborhood attraction using a neural gas approach. Lecture Notes in Computer Science, 2439:391–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390391.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390391.pdf.

Liu:2002:NAP

[876] Pu Liu, Francis Lau, Michael J. Lewis, and Cho li Wang. A new asynchronous parallel evolutionary algorithm for function optimization. Lecture Notes in Computer Science, 2439:401–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390401.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390401.pdf.

Luke:2002:FBN

[877] Sean Luke and Liviu Panait. Fighting bloat with nonparametric parsimony pressure. Lecture Notes in Computer Science, 2439:411–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390411.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390411.pdf.

Muller:2002:ISP

[878] Sibylle D. Müller, Nikolaus Hansen, and Petros Koumoutsakos. Increasing the serial and the parallel performance of the CMAevolution strategy with large populations. Lecture Notes in Computer Science, 2439:422–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390422.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390422.pdf.

Munteanu:2002:ARG

[879] Cristian Munteanu and Agostinho Rosa. Adaptive reservoir genetic algorithm with on-line decision making. Lecture Notes in Computer Science, 2439:432–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390432.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390432.pdf.

Romero:2002:GAV

[880] G. Romero, J. J. Merelo, P. A. Castillo, J. G. Castellano, and M. G. Arenas. Genetic algorithm visualization using self-organizing maps. Lecture Notes in Computer Science, 2439:442–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390442.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390442.pdf.

Roy:2002:GRG

[881] Rajkumar Roy and Ashutosh Tiwari. Generalised regression GA for handling inseparable function interaction: Algorithm and applications. Lecture Notes in Computer Science, 2439:452–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390452.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390452.pdf.

Ursem:2002:DGE

[882] Rasmus K. Ursem. Diversity-guided evolutionary algorithms. Lecture Notes in Computer Science, 2439: 462–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2439/24390462.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2439/24390462.
pdf.

Munoz:2002:EOH

[883] Lluís A. Belanche Muñoz. Evolutionary optimization of heterogeneous problems. Lecture Notes in Computer Science, 2439:475–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390475.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390475.pdf.

Dorado:2002:ARF

[884] Julian Dorado, Juan R. Rabuñal, Antonino Santos, Alejandro Pazos, and Daniel Rivero. Automatic recurrent and feed-forward ANN rule and expression extraction with genetic programming. Lecture Notes in Computer Science, 2439:485–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390485.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390485.pdf.

Liu:2002:LEM

[885] Yong Liu and Xin Yao. Learning and evolution by minimization of mutual information. Lecture Notes in Computer Science, 2439: 495–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2439/24390495.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2439/24390495.
                                        pdf.

Rivas:2002:ERN

[886] V. M. Rivas, P. A. Castillo, and J. J. Merelo. Evolved RBF networks for time-series forecasting and function approximation. Lecture Notes in Computer Science, 2439: 505–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2439/24390505.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2439/24390505.
pdf.

Gonzalez:2002:EIF

[887] Jesús González, Ignacio Rojas, and Héctor Pomares. Evolutive identification of fuzzy systems for time-series prediction. Lecture Notes in Computer Science, 2439:517–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390517.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390517.pdf.

Jakob:2002:HAG

[888] Wilfried Jakob. HyGLEAM — an approach to generally applicable hybridization of evolutionary algorithms. Lecture Notes in Computer Science, 2439:527–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390527.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390527.pdf.

Smith:2002:CEM

[889] Jim Smith. Co-evolving memetic algorithms: Initial investigations. Lecture Notes in Computer Science, 2439:537–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390537.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390537.pdf.

Bull:2002:CMO

[890] Larry Bull and Matt Studley. Consideration of multiple objectives in neural learning classifier systems. Lecture Notes in Computer Science, 2439:549–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390549.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390549.pdf.

Bull:2002:UCN

[891] Larry Bull. On using constructivism in neural classifier systems. Lecture Notes in Computer Science, 2439:558–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390558.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390558.pdf.

Bull:2002:IMX

[892] Larry Bull, David Wyatt, and Ian Parmee. Initial modifications to XCS for use in interactive evolutionary design. Lecture Notes in Computer Science, 2439:568–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390568.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390568.pdf.

Pipe:2002:FRE

[893] A. G. Pipe and B. Carse. First results from experiments in fuzzy classifier system architectures for mobile robotics. Lecture Notes in Computer Science, 2439:578–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390578.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390578.pdf.

Hurst:2002:TLC

[894] Jacob Hurst, Larry Bull, and Chris Melhuish. TCS learning classifier system controller on a real robot. Lecture Notes in Computer Science, 2439:588–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390588.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390588.pdf.

Alba:2002:CSA

[895] Enrique Alba, Mario Giacobini, Marco Tomassini, and Sergio Romero. Comparing synchronous and asynchronous cellular genetic algorithms. Lecture Notes in Computer Science, 2439:601–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390601.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390601.pdf.

Barbulescu:2002:SRS

[896] L. Barbulescu, A. E. Howe, J. P. Watson, and L. D. Whitley. Satellite range scheduling: a comparison of genetic, heuristic and local search. Lecture Notes in Computer Science, 2439:611–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390611.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390611.pdf.

Krink:2002:LMC

[897] Thiemo Krink and Morten Løvbjerg. The LifeCycle model: Combining particle swarm optimisation, genetic algorithms and HillClimbers. Lecture Notes in Computer Science, 2439:621–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390621.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390621.pdf.

Sampels:2002:MGS

[898] Michael Sampels, Christian Blum, Monaldo Mastrolilli, and Olivia Rossi-Doria. Metaheuristics for group shop scheduling. Lecture Notes in Computer Science, 2439:631–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390631.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390631.pdf.

Tomassini:2002:EIT

[899] Marco Tomassini, Leonardo Vanneschi, Francisco Fernández, and Germán Galeano. Experimental investigation of three distributed genetic programming models. Lecture Notes in Computer Science, 2439:641–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390641.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390641.pdf.

Zlochin:2002:MBS

[900] Mark Zlochin and Marco Dorigo. Model-based search for combinatorial optimization: a comparative study. Lecture Notes in Computer Science, 2439:651–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390651.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390651.pdf.

Arenas:2002:FDE

[901] M. G. Arenas, Pierre Collet, A. E. Eiben, Márk Jelasity, J. J. Merelo, Ben Paechter, Mike Preuß, and Marc Schoenauer. A framework for distributed evolutionary algorithms. Lecture Notes in Computer Science, 2439:665–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390665.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390665.pdf.

Castillo:2002:OMP

[902] P. A. Castillo, M. G. Arenas, J. G. Castellano, J. J. Merelo, V. M. Rivas, and G. Romero. Optimisation of multilayer perceptrons using a distributed evolutionary algorithm with SOAP. Lecture Notes in Computer Science, 2439:676–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390676.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390676.pdf.

Anderson:2002:LEB

[903] Eike Falk Anderson. Off-line evolution of behaviour for autonomous agents in real-time computer games. Lecture Notes in Computer Science, 2439:689–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390689.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390689.pdf.

Araujo:2002:PEA

[904] Lourdes Araujo. A parallel evolutionary algorithm for stochastic natural language parsing. Lecture Notes in Computer Science, 2439: 700–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2439/24390700.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2439/24390700.
pdf.

Cordon:2002:ELB

[905] Oscar Cordón, Enrique Herrera-Viedma, and María Luque. Evolutionary learning of Boolean queries by multiobjective genetic programming. Lecture Notes in Computer Science, 2439:710–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390710.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390710.pdf.

Cotta:2002:IPT

[906] Carlos Cotta and Pablo Moscato. Inferring phylogenetic trees using evolutionary algorithms. Lecture Notes in Computer Science, 2439: 720–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2439/24390720.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2439/24390720.
pdf.

Cotta:2002:TME

[907] Carlos Cotta and Jorge Muruzábal. Towards a more efficient evolutionary induction of Bayesian networks. Lecture Notes in Computer Science, 2439:730–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390730.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390730.pdf.

Garcia:2002:RMA

[908] Héctor Fernando Gómez García, Arturo González Vega, Arturo Hernández Aguirre, José Luis Marroquín Zaleta, and Carlos Coello Coello. Robust multiscale affine 2D-image registration through evolutionary strategies. Lecture Notes in Computer Science, 2439:740–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390740.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390740.pdf.

Garionis:2002:SGM

[909] Ralf Garionis. Synthesizing graphical models employing explaining away. Lecture Notes in Computer Science, 2439:749–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390749.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390749.pdf.

Joan-Arinyo:2002:CGC

[910] R. Joan-Arinyo, M. V. Luzón, and A. Soto. Constructive geometric constraint solving: a new application of genetic algorithms. Lecture Notes in Computer Science, 2439:759–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390759.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390759.pdf.

Krasnogor:2002:MAP

[911] N. Krasnogor, B. P. Blackburne, E. K. Burke, and J. D. Hirst. Multimeme algorithms for protein structure prediction. Lecture Notes in Computer Science, 2439:769–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390769.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390769.pdf.

Mabed:2002:DTM

[912] Hakim Mabed, Alexandre Caminada, Jin-Kao Hao, and Denis Renaud. A dynamic traffic model for frequency assignment. Lecture Notes in Computer Science, 2439:779–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390779.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390779.pdf.

Matsui:2002:PFG

[913] Shouichi Matsui, Isamu Watanabe, and Ken ichi Tokoro. A parameter-free genetic algorithm for a fixed channel assignment problem with limited bandwidth. Lecture Notes in Computer Science, 2439:789–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390789.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390789.pdf.

Matsui:2002:RCP

[914] Shouichi Matsui, Isamu Watanabe, and Ken ichi Tokoro. Realcoded parameter-free genetic algorithm for job-shop scheduling problems. Lecture Notes in Computer Science, 2439:800–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390800.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390800.pdf.

Merz:2002:CGE

[915] Peter Merz and Andreas Zell. Clustering gene expression profiles with memetic algorithms. Lecture Notes in Computer Science, 2439:811– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2439/24390811.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2439/24390811.
pdf.

Moore:2002:CAG

[916] Jason H. Moore and Lance W. Hahn. Cellular automata and genetic algorithms for parallel problem solving in human genetics. Lecture Notes in Computer Science, 2439:821–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390821.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390821.pdf.

Motegi:2002:EGG

[917] Makoto Motegi, Naofumi Homma, Takafumi Aoki, and Tatsuo Higuchi. Evolutionary graph generation system and its application to bitserial arithmetic circuit synthesis. Lecture Notes in Computer Science, 2439:831–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390831.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390831.pdf.

Naujoks:2002:EMC

[918] Boris Naujoks, Lars Willmes, Thomas Bäck, and Werner Haase. Evaluating multi-criteria evolutionary algorithms for airfoil optimisation. Lecture Notes in Computer Science, 2439:841–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390841.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2439/24390841.pdf.

Cowling:2002:HRO

[919] Peter Cowling, Graham Kendall, and Eric Soubeiga. Hyperheuristics: a robust optimisation method applied to nurse scheduling. Lecture Notes in Computer Science, 2439: 851–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2439/24390851.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2439/24390851.
                                        pdf.

Thomsen:2002:ETH

[920] René Thomsen. Evolving the topology of hidden Markov models using evolutionary algorithms. Lecture Notes in Computer Science, 2439:861–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390861.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390861.pdf.

Urquhart:2002:SRW

[921] Neil Urquhart, Peter Ross, Ben Paechter, and Ken Chisholm. Solving a real world routing problem using multiple evolutionary agents. Lecture Notes in Computer Science, 2439:871–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390871.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390871.pdf.

Bianchi:2002:ACO

[922] Leonora Bianchi, Luca Maria Gambardella, and Marco Dorigo. An ant colony optimization approach to the probabilistic traveling salesman problem. Lecture Notes in Computer Science, 2439:883–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390883.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2439/24390883.pdf.

Blum:2002:WMB

[923] Christian Blum and Michael Sampels. When model bias is stronger than selection pressure. Lecture Notes in Computer Science, 2439: 893–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2439/24390893.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2439/24390893.
pdf.

Capcarrere:2002:EAC

[924] Mathieu S. Capcarrere. Evolution of asynchronous cellular automata. Lecture Notes in Computer Science, 2439:903–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2439/24390903.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2439/24390903.pdf.

Handl:2002:IAB

[925] Julia Handl and Bernd Meyer. Improved ant-based clustering and sorting. Lecture Notes in Computer Science, 2439:913–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390913.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390913.pdf.

Folino:2002:AFA

[926] Gianluigi Folino and Giandomenico Spezzano. An adaptive flocking algorithm for spatial clustering. Lecture Notes in Computer Science, 2439:924–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390924.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390924.pdf.

Tomassini:2002:EAC

[927] Marco Tomassini and Mattias Venzi. Evolution of asynchronous cellular automata for the density task. Lecture Notes in Computer Science, 2439:934–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2439/24390934.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2439/24390934.pdf.

Anonymous:2002:AIu

[928] Anonymous. Author index. Lecture Notes in Computer Science, 2439: 945–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2439/2439auth.pdf.

Futatsugi:2002:FMC

[929] Kokichi Futatsugi. Formal methods in cafeobj . Lecture Notes in Computer Science, 2441:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2441/24410001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2441/24410001.pdf.

Simeon:2002:EXP

[930] Jérôme Siméon and Philip Wadler. The essence of XML (preliminary version). Lecture Notes in Computer Science, 2441:21–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2441/24410021.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2441/24410021.pdf.

GarciadelaBanda:2002:GHH

[931] María García de la Banda, Bart Demoen, Kim Marriott, and Peter J. Stuckey. To the Gates of HAL: a HAL tutorial. Lecture Notes in Computer Science, 2441:47–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2441/24410047.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2441/24410047.pdf.

Antoy:2002:FLD

[932] Sergio Antoy and Michael Hanus. Functional logic design patterns. Lecture Notes in Computer Science, 2441:67–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2441/24410067.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2441/24410067.pdf.

Meer:2002:CWN

[933] Klaus Meer. On consistency and width notions for constraint programs with algebraic constraints. Lecture Notes in Computer Science, 2441:88–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2441/24410088.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2441/24410088.pdf.

Garcia-Diaz:2002:SMQ

[934] Miguel García-Díaz and Susana Nieva. Solving mixed quantified constraints over a domain based on R eal numbers and H erbrand terms. Lecture Notes in Computer Science, 2441:103–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

series/0558/bibs/2441/24410103.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2441/24410103.pdf.

Iranzo:2002:CFT

[935] Pascual Julián Iranzo. On the correctness of the factoring transformation. Lecture Notes in Computer Science, 2441:119–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2441/24410119.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2441/24410119.pdf.

Danvy:2002:LLQ

[936] Olivier Danvy and Ulrik P. Schultz. Lambda-lifting in quadratic time. Lecture Notes in Computer Science, 2441:134–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2441/24410134.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2441/24410134.pdf.

Loyer:2002:WFS

[937] Yann Loyer and Umberto Straccia. The well-founded semantics in normal logic programs with uncertainty. Lecture Notes in Computer Science, 2441:152–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2441/24410152.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2441/24410152.pdf.

Hasegawa:2002:LUE

[938] Masahito Hasegawa. Linearly used effects: Monadic and CPS transformations into the linear lambda calculus. Lecture Notes in Computer Science, 2441:167–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2441/24410167.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2441/24410167.pdf.

Barbosa:2002:CIP

[939] Luís Soares Barbosa and José Nuno Oliveira. Coinductive interpreters for process calculi. Lecture Notes in Computer Science, 2441:183– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2441/24410183.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2441/24410183.
                                        pdf.

Kasuya:2002:DHN

[940] Hideto Kasuya, Masahiko Sakai, and Kiyoshi Agusa. Descendants and head normalization of higher-order rewrite systems. Lecture Notes in Computer Science, 2441:198–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2441/24410198.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2441/24410198.pdf.

Lopez-Fraguas:2002:NFF

[941] Francisco Javier López-Fraguas and Jaime Sánchez-Hernández. Narrowing failure in functional logic programming. Lecture Notes in Computer Science, 2441:212–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2441/24410212.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2441/24410212.pdf.

Pettersson:2002:HXE

[942] Mikael Pettersson, Konstantinos Sagonas, and Erik Johansson. The HiPE/x86 Erlang compiler: System description and performance evaluation. Lecture Notes in Computer Science, 2441:228–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2441/24410228.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2441/24410228.pdf.

Caballero:2002:DCD

[943] Rafael Caballero and Francisco Javier López-Fraguas. Dynamic-cut with definitional trees. Lecture Notes in Computer Science, 2441:245– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2441/24410245.htm;   http:
//link.springer.de/link/service/

                                        series/0558/papers/2441/24410245.
                                        pdf.

Smaus:2002:HCP

[944] Jan-Georg Smaus. The head condition and polymorphic recursion. Lecture Notes in Computer Science, 2441:259–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2441/24410259.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2441/24410259.pdf.

Porto:2002:SAA

[945] António Porto. Structural abstraction and application in logic programming. Lecture Notes in Computer Science, 2441:275–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2441/24410275.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2441/24410275.pdf.

Sumii:2002:VFC

[946] Eijiro Sumii and Hideo Bannai. VMλ: a functional calculus for scientific discovery. Lecture Notes in Computer Science, 2441:290–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2441/24410290.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2441/24410290.pdf.

Anonymous:2002:AIv

[947] Anonymous. Author index. Lecture Notes in Computer Science, 2441: 305–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2441/2441ind.pdf.

Murphy:2002:EAS

[948] Sean Murphy and Matthew J. B. Robshaw. Essential algebraic structure within the AES. Lecture Notes in Computer Science, 2442: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2442/24420001.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2442/24420001.
pdf.

Joux:2002:BAA

[949] Antoine Joux, Gwenaëlle Martinet, and Frédéric Valette. Blockwiseadaptive attackers: Revisiting the (in)security of some provably secure encryption models: CBC, GEM, IACBC. Lecture Notes in Computer Science, 2442:17–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420017.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420017.pdf.

Liskov:2002:TBC

[950] Moses Liskov, Ronald L. Rivest, and David Wagner. Tweakable block ciphers. Lecture Notes in Computer Science, 2442:31–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420031.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420031.pdf.

Halevy:2002:LBE

[951] Dani Halevy and Adi Shamir. The LSD broadcast encryption scheme. Lecture Notes in Computer Science, 2442:47–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420047.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420047.pdf.

Camenisch:2002:DAA

[952] Jan Camenisch and Anna Lysyanskaya. Dynamic accumulators and application to efficient revocation of anonymous credentials. Lecture Notes in Computer Science, 2442:61–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420061.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420061.pdf.

Hopper:2002:PSS

[953] Nicholas J. Hopper, John Langford, and Luis von Ahn. Provably secure steganography (extended abstract). Lecture Notes in Computer Science, 2442:77–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420077.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420077.pdf.

Stern:2002:FAP

[954] Jacques Stern, David Pointcheval, John Malone-Lee, and Nigel P. Smart. Flaws in applying proof methodologies to signature schemes. Lecture Notes in Computer Science, 2442:93–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420093.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420093.pdf.

Nielsen:2002:SRO

[955] Jesper Buus Nielsen. Separating random oracle proofs from complexity theoretic proofs: The noncommitting encryption case. Lecture Notes in Computer Science, 2442: 111–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2442/24420111.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2442/24420111.
pdf.

Jonsson:2002:SRE

[956] Jakob Jonsson and Burton S. Kaliski Jr. On the security of RSA encryption in TLS. Lecture Notes in Computer Science, 2442:127–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2442/24420127.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2442/24420127.pdf.

Canetti:2002:SAI

[957] Ran Canetti and Hugo Krawczyk. Security analysis of IKE’s signaturebased key-exchange protocol. Lecture Notes in Computer Science, 2442: 143–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2442/24420143.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2442/24420143.
                                        pdf.

Bellare:2002:GSI

[958] Mihir Bellare and Adriana Palacio. GQ and Schnorr identification schemes: Proofs of security against impersonation under active and concurrent attacks. Lecture Notes in Computer Science, 2442:162–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2442/24420162.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2442/24420162.pdf.

Gennaro:2002:RSM

[959] Rosario Gennaro, Yuval Ishai, Eyal Kushilevitz, and Tal Rabin. On 2-round secure multiparty computation. Lecture Notes in Computer Science, 2442:178–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420178.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420178.pdf.

Blaser:2002:PCC

[960] Markus Bläser, Andreas Jakoby, Maciej Liśkiewicz, and Bodo Siebert. Private computation — k-connected versus 1-connected networks. Lecture Notes in Computer Science, 2442: 194–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2442/24420194.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2442/24420194.
pdf.

Nguyen:2002:AIN

[961] Phong Q. Nguyen and David Pointcheval. Analysis and improvements of NTRU encryption paddings. Lecture Notes in Computer Science, 2442:210– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2442/24420210.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2442/24420210.
pdf.

Coron:2002:UPS

[962] Jean-Sébastien Coron, Marc Joye, David Naccache, and Pascal Paillier. Universal padding schemes for RSA. Lecture Notes in Computer Science, 2442:226–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2442/24420226.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2442/24420226.pdf.

May:2002:CUR

[963] Alexander May. Cryptanalysis of unbalanced RSA with small CRTexponent. Lecture Notes in Computer Science, 2442:242–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2442/24420242.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2442/24420242.pdf.

Lu:2002:HEA

[964] Chi-Jen Lu. Hyper-encryption against space-bounded adversaries from online strong extractors. Lecture Notes in Computer Science, 2442: 257–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2442/24420257.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2442/24420257.
                                        pdf.

Cramer:2002:OBB

[965] Ronald Cramer and Serge Fehr. Optimal black-box secret sharing over arbitrary Abelian groups. Lecture Notes in Computer Science, 2442: 272–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/

0558/bibs/2442/24420272.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2442/24420272.
pdf.

Wagner:2002:GBP

[966] David Wagner. A generalized birthday problem (extended abstract). Lecture Notes in Computer Science, 2442:288–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420288.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420288.pdf.

Mironov:2002:RSR

[967] Ilya Mironov. (not so) random shuffles of RC4. Lecture Notes in Computer Science, 2442:304–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420304.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420304.pdf.

Black:2002:BBA

[968] John Black, Phillip Rogaway, and Thomas Shrimpton. Black-box analysis of the block-cipher-based hashfunction constructions from PGV. Lecture Notes in Computer Science, 2442:320–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420320.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2442/24420320.pdf.

Rubin:2002:SAV

[969] Karl Rubin and Alice Silverberg. Supersingular Abelian varieties in cryptology. Lecture Notes in Computer Science, 2442:336–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420336.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420336.pdf.

Barreto:2002:EAP

[970] Paulo S. L. M. Barreto, Hae Y. Kim, Ben Lynn, and Michael Scott. Efficient algorithms for pairing-based cryptosystems. Lecture Notes in Computer Science, 2442:354–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420354.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420354.pdf.

Vercauteren:2002:CZF

[971] Frederik Vercauteren. Computing zeta functions of hyperelliptic curves over finite fields of characteristic 2. Lecture Notes in Computer Science, 2442:369–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420369.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420369.pdf.

MacKenzie:2002:TPA

[972] Philip MacKenzie, Thomas Shrimpton, and Markus Jakobsson. Threshold password-authenticated key exchange (extended abstract). Lecture Notes in Computer Science, 2442: 385–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2442/24420385.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2442/24420385.
pdf.

Nielsen:2002:TPF

[973] Jesper Buus Nielsen. A threshold pseudorandom function construction and its applications. Lecture Notes in Computer Science, 2442: 401–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2442/24420401.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2442/24420401.
pdf.

Algesheimer:2002:ECM

[974] Joy Algesheimer, Jan Camenisch, and Victor Shoup. Efficient computation modulo a shared secret with application to the generation of shared safe-prime products. Lecture Notes in Computer Science, 2442:417–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420417.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2442/24420417.pdf.

Li:2002:HNP

[975] Wen-Ching W. Li, Mats Näslund, and Igor E. Shparlinski. Hidden number problem with the trace and bit security of XTR and LUC. Lecture Notes in Computer Science, 2442:433–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2442/24420433.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2442/24420433.pdf.

Damgaard:2002:EPF

[976] Ivan Damgård and Jesper Buus Nielsen. Expanding pseudorandom functions; or: From known-plaintext security to chosen-plaintext security. Lecture Notes in Computer Science, 2442:449–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2442/24420449.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2442/24420449.pdf.

Bresson:2002:TRS

[977] Emmanuel Bresson, Jacques Stern, and Michael Szydlo. Threshold ring signatures and applications to ad-hoc groups. Lecture Notes in Computer Science, 2442:465–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2442/24420465.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420465.pdf.

Naor:2002:DRA

[978] Moni Naor. Deniable ring authentication. Lecture Notes in Computer Science, 2442:481–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420481.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420481.pdf.

Itkis:2002:SSB

[979] Gene Itkis and Leonid Reyzin. SiBIR: Signer-base intrusion-resilient signatures. Lecture Notes in Computer Science, 2442:499–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420499.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420499.pdf.

Coppersmith:2002:CSC

[980] Don Coppersmith, Shai Halevi, and Charanjit Jutla. Cryptanalysis of stream ciphers with linear masking. Lecture Notes in Computer Science, 2442:515–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420515.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420515.pdf.

Sarkar:2002:FCM

[981] Palash Sarkar. The filter-combiner model for memoryless synchronous stream ciphers. Lecture Notes in Computer Science, 2442:533– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2442/24420533.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2442/24420533.
                                        pdf.

Carlet:2002:LCC

[982] Claude Carlet. A larger class of cryptographic Boolean functions via a study of the Maiorana–McFarland construction. Lecture Notes in Computer Science, 2442:549–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2442/24420549.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2442/24420549.pdf.

Fehr:2002:LVD

[983] Serge Fehr and Ueli Maurer. Linear VSS and distributed commitments based on secret sharing and pairwise checks. Lecture Notes in Computer Science, 2442:565–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2442/24420565.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2442/24420565.pdf.

Damgaard:2002:PHP

[984] Ivan Damgård and Jesper Buus Nielsen. Perfect hiding and perfect binding universally composable commitment schemes with constant expansion factor. Lecture Notes in Computer Science, 2442:581–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420581.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420581.pdf.

Lysyanskaya:2002:USV

[985] Anna Lysyanskaya. Unique signatures and verifiable random functions from the DH-DDH separation. Lecture Notes in Computer Science, 2442:597–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420597.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420597.pdf.

Coron:2002:SPP

[986] Jean-Sébastien Coron. Security proof for partial-domain hash signature schemes. Lecture Notes in Computer Science, 2442:613–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420613.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420613.pdf.

Anonymous:2002:AIw

[987] Anonymous. Author index. Lecture Notes in Computer Science, 2442: 627–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2442/2442auth.pdf.

Giebel:2002:MST

[988] J. Giebel and D. M. Gavrila. Multimodal shape tracking with point distribution models. Lecture Notes in Computer Science, 2449: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2449/24490001.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2449/24490001.
                                        pdf.

Brandner:2002:RTT

[989] M. Brandner and A. Pinz. Realtime tracking of complex objects using dynamic interpretation tree. Lecture Notes in Computer Science, 2449:9–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490009.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2449/24490009.pdf.

Denzler:2002:OCP

[990] J. Denzler, M. Zobel, and H. Niemann. On optimal camera parameter selection in Kalman filter based object tracking. Lecture Notes in Computer Science, 2449:17–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490017.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490017.pdf.

Funck:2002:VBH

[991] S. Funck. Video-based handsign recognition for intuitive human-computer-interaction. Lecture Notes in Computer Science, 2449:26–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490026.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490026.pdf.

Cantzler:2002:QER

[992] H. Cantzler, R. B. Fisher, and M. Devy. Quality enhancement of reconstructed 3D models using coplanarity and constraints. Lecture Notes in Computer Science, 2449:34–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490034.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490034.pdf.

Munzenmayer:2002:MTA

[993] C. Münzenmayer, H. Volk, C. Küblbeck, K. Spinnler, and T. Wittenberg. Multispectral texture analysis using interplane sum- and differencehistograms. Lecture Notes in Computer Science, 2449:42–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490042.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2449/24490042.pdf.

Remondino:2002:HBR

[994] F. Remondino. Human body reconstruction from image sequences. Lecture Notes in Computer Science, 2449:50–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490050.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2449/24490050.pdf.

Buckner:2002:KBS

[995] J. Bückner, M. Pahl, O. Stahlhut, and C.-E. Liedtke. A knowledgebased system for context dependent evaluation of remote sensing data. Lecture Notes in Computer Science, 2449:58–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490058.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2449/24490058.pdf.

Middendorf:2002:ECA

[996] M. Middendorf and H.-H. Nagel. Empirically convergent adaptive estimation of grayvalue structure tensors. Lecture Notes in Computer Science, 2449:66–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490066.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490066.pdf.

Heinrich:2002:RTF

[997] S. Heinrich. Real time fusion of motion and stereo using flow/depth constraint for fast obstacle detection. Lecture Notes in Computer Science, 2449:75–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490075.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490075.pdf.

Rocha:2002:PPS

[998] J. Rocha. Polygon partition into stable regions. Lecture Notes in Computer Science, 2449:83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490083.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490083.pdf.

Jost:2002:FIA

[999] T. Jost and H. Hügli. Fast ICP algorithms for shape registration. Lecture Notes in Computer Science, 2449:91–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490091.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2449/24490091.pdf.

Drexler:2002:ABG

[1000] C. Drexler, F. Mattern, and J. Denzler. Appearance based generic object modeling and recognition using probabilistic principal component analysis. Lecture Notes in Computer Science, 2449:100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490100.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2449/24490100.pdf.

Kruppa:2002:SPD

[1001] H. Kruppa, M. A. Bauer, and B. Schiele. Skin patch detection in real-world images. Lecture Notes in Computer Science, 2449: 109–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2449/24490109.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2449/24490109.
                                        pdf.

Haxhimusa:2002:LTG

[1002] Y. Haxhimusa, R. Glantz, M. Saib, G. Langs, and W. G. Kropatsch. Logarithmic tapering graph pyramid. Lecture Notes in Computer Science, 2449:117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490117.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2449/24490117.pdf.

Sojka:2002:NEA

[1003] E. Sojka. A new and efficient algorithm for detecting the corners in digital images. Lecture Notes in Computer Science, 2449:125–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490125.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490125.pdf.

Kuhne:2002:FIA

[1004] G. Kühne, J. Weickert, M. Beier, and W. Effelsberg. Fast implicit active contour models. Lecture Notes in Computer Science, 2449: 133–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2449/24490133.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2449/24490133.
pdf.

Keuchel:2002:UIP

[1005] J. Keuchel, C. Schnörr, C. Schellewald, and D. Cremers. Unsupervised image partitioning with semidefinite programming. Lecture Notes in Computer Science, 2449:141–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490141.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490141.pdf.

Aurnhammer:2002:AGA

[1006] M. Aurnhammer and K. Tönnies. The application of genetic algorithms in structural seismic image interpretation. Lecture Notes in Computer Science, 2449:150–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490150.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2449/24490150.pdf.

Goldberger:2002:UIC

[1007] J. Goldberger, H. Greenspan, and S. Gordon. Unsupervised image clustering using the information bottleneck method. Lecture Notes in Computer Science, 2449:158–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490158.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2449/24490158.pdf.

Smolka:2002:EMC

[1008] B. Smolka. Efficient modification of the central weighted vector median filter. Lecture Notes in Computer Science, 2449:166–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490166.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2449/24490166.pdf.

Katsoulas:2002:FRP

[1009] D. Katsoulas and A. Jaklič. Fast recovery of piled deformable objects using superquadrics. Lecture Notes in Computer Science, 2449: 174–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2449/24490174.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2449/24490174.
pdf.

Weitzenberg:2002:AAB

[1010] J. Weitzenberg, S. Posch, and M. Rost.

Analysis of amperometric biosensor curves using hidden-Markov-modells. Lecture Notes in Computer Science, 2449:182–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490182.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490182.pdf.

Flach:2002:URS

[1011] B. Flach, E. Kask, D. Schlesinger, and A. Skulish. Unifying registration and segmentation for multi-sensor images. Lecture Notes in Computer Science, 2449:190–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490190.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490190.pdf.

Steidl:2002:RBS

[1012] G. Steidl and J. Weickert. Relations between soft wavelet shrinkage and total variation denoising. Lecture Notes in Computer Science, 2449:198–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490198.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2449/24490198.pdf.

Fischer:2002:DRP

[1013] B. Fischer and J. M. Buhmann. Data resampling for path based clustering. Lecture Notes in Computer Science, 2449:206–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490206.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2449/24490206.pdf.

Brocke:2002:SIS

[1014] M. Brocke. Statistical image sequence processing for temporal change detection. Lecture Notes in Computer Science, 2449:215–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490215.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2449/24490215.pdf.

Criminisi:2002:SVM

[1015] A. Criminisi. Single-view metrology: Algorithms and applications.

Lecture Notes in Computer Science, 2449:224–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490224.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490224.pdf.

Winkelbach:2002:SSS

[1016] S. Winkelbach and F. M. Wahl. Shape from single stripe pattern illumination. Lecture Notes in Computer Science, 2449:240–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490240.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490240.pdf.

Felsberg:2002:DMP

[1017] M. Felsberg. Disparity from monogenic phase. Lecture Notes in Computer Science, 2449:248–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490248.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490248.pdf.

Scholze:2002:PAB

[1018] S. Scholze, T. Moons, and L. Van Gool.

A probabilistic approach to building roof reconstruction using semantic labelling. Lecture Notes in Computer Science, 2449:257–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490257.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2449/24490257.pdf.

Rosenhahn:2002:APE

[1019] B. Rosenhahn and G. Sommer. Adaptive pose estimation for different corresponding entities. Lecture Notes in Computer Science, 2449: 265–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2449/24490265.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2449/24490265.
                                        pdf.

Sturm:2002:PTD

[1020] P. Sturm and L. Priese. Properties of a three-dimensional island hierarchy for segmentation of 3D images with the color structure code. Lecture Notes in Computer Science, 2449:274–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490274.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2449/24490274.pdf.

Ouerhani:2002:RTI

[1021] N. Ouerhani, H. Hügli, P.-Y. Burgi, and P.-F. Ruedi. A real time implementation of the saliency-based model of visual attention on a SIMD architecture. Lecture Notes in Computer Science, 2449:282–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490282.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490282.pdf.

Bruhn:2002:DDN

[1022] A. Bruhn, T. Jakob, M. Fischer, T. Kohlberger, J. Weickert, U. Brüning, and C. Schnörr. Designing 3-D nonlinear diffusion filters for high performance cluster computing. Lecture Notes in Computer Science, 2449:290–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490290.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490290.pdf.

Zhang:2002:HPB

[1023] X. Zhang and H. Burkhardt. Hierarchical primitives based contour matching. Lecture Notes in Computer Science, 2449:298–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490298.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490298.pdf.

Massad:2002:ATV

[1024] A. Massad, M. Babós, and B. Mertsching.

Application of the tensor voting technique for perceptual grouping to grey-level images. Lecture Notes in Computer Science, 2449:306– ??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2449/24490306.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2449/24490306.
                                        pdf.

Smolka:2002:CFB

[1025] B. Smolka. Combined forward and backward anisotropic diffusion filtering of color images. Lecture Notes in Computer Science, 2449: 314–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2449/24490314.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2449/24490314.
                                        pdf.

Krodel:2002:RLD

[1026] M. Krödel and K.-D. Kuhnert. Reinforcement learning to drive a car by pattern matching. Lecture Notes in Computer Science, 2449: 322–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2449/24490322.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2449/24490322.
                                        pdf.

Meshoul:2002:RPC

[1027] S. Meshoul and M. Batouche. Robust point correspondence for image registration using optimization with extremal dynamics. Lecture Notes in Computer Science, 2449: 330–??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2449/24490330.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2449/24490330.
pdf.

Al-Zubi:2002:EAS

[1028] S. Al-Zubi and K. Tönnies. Extending active shape models to incorporate apriori knowledge about structural variability. Lecture Notes in Computer Science, 2449:338–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490338.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490338.pdf.

Eisele:2002:NAD

[1029] H. Eisele and F. A. Hamprecht. A new approach for defect detection in X-ray CT images. Lecture Notes in Computer Science, 2449:345–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490345.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490345.pdf.

Nummiaro:2002:OTA

[1030] K. Nummiaro, E. Koller-Meier, and L. Van Gool. Object tracking with an adaptive color-based particle filter. Lecture Notes in Computer Science, 2449:353–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490353.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2449/24490353.pdf.

Moller:2002:AOI

[1031] B. Möller and S. Posch. Analysis of object interactions in dynamic scenes. Lecture Notes in Computer Science, 2449:361–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490361.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2449/24490361.pdf.

Brauer-Burchardt:2002:AMR

[1032] C. Bräuer-Burchardt. Automatic monocular 3D-reconstruction of indoor environments using mobile vehicles. Lecture Notes in Computer Science, 2449:370–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490370.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2449/24490370.pdf.

Spies:2002:DPF

[1033] H. Spies and C. S. Garbe. Dense parameter fields from total least squares. Lecture Notes in Computer Science, 2449:379–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490379.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2449/24490379.pdf.

Tang:2002:FBM

[1034] W. K. Tang and Y. S. Hung. A factorization-based method for projective reconstruction with minimization of 2-D reprojection errors. Lecture Notes in Computer Science, 2449:387–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490387.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490387.pdf.

Clabian:2002:HDL

[1035] M. Clabian, H. Rötzer, H. Bischof, and W. Kropatsch. Head detection and localization from sparse 3D data. Lecture Notes in Computer Science, 2449:395–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490395.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490395.pdf.

Feldbach:2002:SDE

[1036] M. Feldbach and K. D. Tönnies.

Segmentation of the date in entries of historical church registers. Lecture Notes in Computer Science, 2449:403–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490403.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2449/24490403.pdf.

Nguyen:2002:AOR

[1037] M.-C. Nguyen and B. Radig. Approach to object recognition by learning mobile robots. Lecture Notes in Computer Science, 2449:411–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490411.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2449/24490411.pdf.

Miro:2002:UIF

[1038] A. Di Miro, B. Wirnitzer, C. Zott, and R. Klinnert. Ultrasonic image formation using wide angle sensor arrays and cross-echo evaluation. Lecture Notes in Computer Science, 2449:421–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490421.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2449/24490421.pdf.

Bergner:2002:SMM

[1039] S. Bergner, R. Pohle, S. Al-Zubi, K. Tönnies, A. Eitner, and T. R. Neu. Segmenting microorganisms in multi-modal volumetric datasets using a modified watershed transform. Lecture Notes in Computer Science, 2449:429–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490429.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490429.pdf.

Comaniciu:2002:BKT

[1040] D. Comaniciu. Bayesian kernel tracking. Lecture Notes in Computer Science, 2449:438–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490438.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490438.pdf.

Brox:2002:NMD

[1041] T. Brox and J. Weickert. Nonlinear matrix diffusion for optic flow estimation. Lecture Notes in Computer Science, 2449:446–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490446.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490446.pdf.

Bruhn:2002:CAL

[1042] A. Bruhn, J. Weickert, and C. Schnörr.

Combining the advantages of local and global optic flow methods. Lecture Notes in Computer Science, 2449:454–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490454.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490454.pdf.

Garbe:2002:MOT

[1043] C. S. Garbe, H. Spies, and B. Jähne. Mixed OLS-TLS for the estimation of dynamic processes with a linear source term. Lecture Notes in Computer Science, 2449:463–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490463.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2449/24490463.pdf.

Cremers:2002:MCV

[1044] D. Cremers and C. Schnörr. Motion competition: Variational integration of motion segmentation and shape regularization. Lecture Notes in Computer Science, 2449:472–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490472.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2449/24490472.pdf.

Gross:2002:FLF

[1045] R. Gross, I. Matthews, and S. Baker. Fisher light-fields for face recognition across pose and illumination. Lecture Notes in Computer Science, 2449:481–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490481.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2449/24490481.pdf.

Lomker:2002:MSO

[1046] F. Lömker and G. Sagerer. A multimodal system for object learning. Lecture Notes in Computer Science, 2449:490–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490490.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490490.pdf.

Keysers:2002:MEG

[1047] D. Keysers, F. J. Och, and H. Ney.

Maximum entropy and Gaussian models for image object recognition. Lecture Notes in Computer Science, 2449:498–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490498.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490498.pdf.

Brakensiek:2002:ECM

[1048] A. Brakensiek, A. Kosmala, and G. Rigoll. Evaluation of confidence measures for on-line handwriting recognition. Lecture Notes in Computer Science, 2449:507–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490507.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490507.pdf.

Svoboda:2002:VLC

[1049] T. Svoboda, H. Hug, and L. Van Gool. ViRoom — low cost synchronized multicamera system and its selfcalibration. Lecture Notes in Computer Science, 2449:515–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490515.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490515.pdf.

Chateau:2002:RTT

[1050] T. Chateau, F. Jurie, M. Dhome, and X. Clady. Real-time tracking using wavelet representation. Lecture Notes in Computer Science, 2449:523–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490523.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490523.pdf.

Zieren:2002:HTF

[1051] J. Zieren, N. Unger, and S. Akyol. Hands tracking from frontal view for vision-based gesture recognition. Lecture Notes in Computer Science, 2449:531–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490531.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490531.pdf.

Calvo:2002:AAT

[1052] C. Calvo, A. Micarelli, and E. Sangineto.

Automatic annotation of tennis video sequences. Lecture Notes in Computer Science, 2449:540–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490540.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490540.pdf.

Ahn:2002:FPS

[1053] S. J. Ahn, W. Rauh, and E. Westkämper.

Fitting of parametric space curves and surfaces by using the geometric error measure. Lecture Notes in Computer Science, 2449:548–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490548.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490548.pdf.

Koninckx:2002:EAA

[1054] T. P. Koninckx and L. Van Gool.

Efficient, active 3D acquisition, based on a pattern-specific snake. Lecture Notes in Computer Science, 2449:557–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490557.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490557.pdf.

Kruger:2002:ABD

[1055] V. Krüger, R. Gross, and S. Baker. Appearance-based 3-D face recognition from video. Lecture Notes in Computer Science, 2449:566–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490566.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2449/24490566.pdf.

Navab:2002:DMM

[1056] N. Navab, Y. Genc, A. Khamene, and M. Mitschke. Direct method for motion estimation: An alternative to decomposition of planar transformation matrices. Lecture Notes in Computer Science, 2449:575–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490575.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2449/24490575.pdf.

Cheoi:2002:FDA

[1057] K. Cheoi and Y. Lee. A feature-driven attention module for an active vision system. Lecture Notes in Computer Science, 2449:583–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2449/24490583.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2449/24490583.pdf.

Canterakis:2002:ARK

[1058] N. Canterakis. Analytic reduction of the kruppa equations. Lecture Notes in Computer Science, 2449: 591–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2449/24490591.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2449/24490591.
pdf.

Vogel:2002:QDP

[1059] J. Vogel and B. Schiele. Querydependent performance optimization for vocabulary-supported image retrieval. Lecture Notes in Computer Science, 2449:600–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490600.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490600.pdf.

Bulow:2002:MIP

[1060] T. Bülow. Multiscale image processing on the sphere. Lecture Notes in Computer Science, 2449: 609–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2449/24490609.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2449/24490609.
pdf.

Neumann:2002:PCN

[1061] J. Neumann, C. Fermüller, and Y. Aloimonos. Polydioptric cameras: New eyes for structure from motion. Lecture Notes in Computer Science, 2449:618–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2449/24490618.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2449/24490618.pdf.

Anonymous:2002:AIx

[1062] Anonymous. Author index. Lecture Notes in Computer Science, 2449: 627–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2449/2449auth.pdf.

Grate:2002:SRF

[1063] L. R. Grate, C. Bhattacharyya, M. I. Jordan, and I. S. Mian. Simultaneous relevant feature identification and classification in high-dimensional spaces. Lecture Notes in Computer Science, 2452:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2452/24520001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2452/24520001.pdf.

Csuros:2002:PGI

[1064] Miklós Csűrös and Aleksandar Milosavljevic. Pooled genomic indexing (PGI): Mathematical analysis and experiment design. Lecture Notes in Computer Science, 2452:10–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2452/24520010.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2452/24520010.pdf.

Rizzi:2002:PAF

[1065] Romeo Rizzi, Vineet Bafna, Sorin Istrail, and Giuseppe Lancia. Practical algorithms and fixed-parameter tractability for the single individual SNP haplotyping problem. Lecture Notes in Computer Science, 2452:29–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2452/24520029.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2452/24520029.pdf.

Schwartz:2002:MIB

[1066] Russell Schwartz, Andrew G. Clark, and Sorin Istrail. Methods for inferring block-wise ancestral history from haploid sequences. Lecture Notes in Computer Science, 2452: 44–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2452/24520044.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2452/24520044.
pdf.

Reczko:2002:FSP

[1067] Martin Reczko, Petko Fiziev, Eike Staub, and Artemis Hatzigeorgiou. Finding signal peptides in human protein sequences using recurrent neural networks. Lecture Notes in Computer Science, 2452:60–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2452/24520060.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2452/24520060.pdf.

Edwards:2002:GPC

[1068] Nathan Edwards and Ross Lippert. Generating peptide candidates from amino-acid sequence databases for protein identification via mass spectrometry. Lecture Notes in Computer Science, 2452:68–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2452/24520068.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2452/24520068.pdf.

Chen:2002:IAA

[1069] Zhi-Zhong Chen, Tao Jiang, Guohui Lin, Jianjun Wen, Dong Xu, and Ying Xu. Improved approximation algorithms for NMR spectral peak assignment. Lecture Notes in Computer Science, 2452:82–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2452/24520082.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2452/24520082.pdf.

Zhang:2002:EMI

[1070] Louxin Zhang, Bin Ma, and Lusheng Wang. Efficient methods for inferring tandem duplication history.

Lecture Notes in Computer Science, 2452:97–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2452/24520097.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2452/24520097.pdf.

Wang:2002:GRP

[1071] Li-San Wang. Genome rearrangement phylogeny using weighbor. Lecture Notes in Computer Science, 2452:112–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2452/24520112.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2452/24520112.pdf.

Halpern:2002:SMR

[1072] Aaron L. Halpern, Daniel H. Huson, and Knut Reinert. Segment match refinement and applications. Lecture Notes in Computer Science, 2452:126–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2452/24520126.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2452/24520126.pdf.

Adebiyi:2002:ECM

[1073] Ezekiel F. Adebiyi and Michael Kaufmann. Extracting common motifs under the Levenshtein measure: Theory and experimentation. Lecture Notes in Computer Science, 2452:140–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2452/24520140.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2452/24520140.pdf.

Goldwasser:2002:FAF

[1074] Michael H. Goldwasser, Ming-Yang Kao, and Hsueh-I Lu. Fast algorithms for finding maximumdensity segments of a sequence with applications to bioinformatics. Lecture Notes in Computer Science, 2452:157–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2452/24520157.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2452/24520157.pdf.

Milik:2002:FAE

[1075] Mariusz Milik, Sandor Szalma, and Krzysztof A. Olszewski. FAUST: An algorithm for extracting functionally relevant templates from protein structures. Lecture Notes in Computer Science, 2452:172–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2452/24520172.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2452/24520172.pdf.

Duhovny:2002:EUD

[1076] Dina Duhovny, Ruth Nussinov, and Haim J. Wolfson. Efficient unbound docking of rigid molecules. Lecture Notes in Computer Science, 2452:185–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2452/24520185.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2452/24520185.pdf.

Wheeler:2002:MCC

[1077] Raymond Wheeler. A method of consolidating and combining EST and mRNA alignments to a genome to enumerate supported splice variants. Lecture Notes in Computer Science, 2452:201–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2452/24520201.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2452/24520201.pdf.

Pertea:2002:MIP

[1078] Mihaela Pertea and Steven L. Salzberg.

A method to improve the performance of translation start site detection and its application for gene finding. Lecture Notes in Computer Science, 2452:210–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2452/24520210.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2452/24520210.pdf.

Pedersen:2002:CMG

[1079] Christian N. S. Pedersen and Tejs Scharling. Comparative methods for gene structure prediction in homologous sequences. Lecture Notes in Computer Science, 2452: 220–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2452/24520220.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2452/24520220.
                                        pdf.

Shatsky:2002:MMP

[1080] Maxim Shatsky, Ruth Nussinov, and Haim J. Wolfson. MultiProt — A multiple protein structural alignment algorithm. Lecture Notes in Computer Science, 2452:235–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2452/24520235.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2452/24520235.pdf.

Rocke:2002:HSF

[1081] Emily Rocke. A hybrid scoring function for protein multiple alignment. Lecture Notes in Computer Science, 2452:251–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2452/24520251.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2452/24520251.pdf.

Bilu:2002:FCM

[1082] Yonatan Bilu and Michal Linial. Functional consequences in metabolic pathways from phylogenetic profiles. Lecture Notes in Computer Science, 2452:263–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2452/24520263.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2452/24520263.pdf.

Ukkonen:2002:FFS

[1083] Esko Ukkonen. Finding founder sequences from a set of recombinants. Lecture Notes in Computer Science, 2452:277–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2452/24520277.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2452/24520277.pdf.

Nakhleh:2002:EDM

[1084] Luay Nakhleh, Usman Roshan, Lisa Vawter, and Tandy Warnow. Estimating the deviation from a molecular clock. Lecture Notes in Computer Science, 2452:287–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2452/24520287.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2452/24520287.pdf.

Ajana:2002:ESA

[1085] Yasmine Ajana, Jean-François Lefebvre, Elisabeth R. M. Tillier, and Nadia El-Mabrouk. Exploring the set of all minimal sequences of reversals — an application to test the replication-directed reversal hypothesis. Lecture Notes in Computer Science, 2452:300–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2452/24520300.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2452/24520300.pdf.

Eriksen:2002:AEN

[1086] Niklas Eriksen. Approximating the expected number of inversions given the number of breakpoints. Lecture Notes in Computer Science, 2452:316–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2452/24520316.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2452/24520316.pdf.

Myers:2002:ILA

[1087] Gene Myers and Richard Durbin. Invited lecture — accelerating Smith-Waterman searches. Lecture Notes in Computer Science, 2452:331– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2452/24520331.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2452/24520331.
                                        pdf.

Moret:2002:SLR

[1088] Bernard M. E. Moret, Usman Roshan, and Tandy Warnow. Sequence-length requirements for phylogenetic methods. Lecture Notes in Computer Science, 2452:343–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2452/24520343.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2452/24520343.pdf.

Desper:2002:FAP

[1089] Richard Desper and Olivier Gascuel. Fast and accurate phylogeny reconstruction algorithms based on the minimum-evolution principle. Lecture Notes in Computer Science, 2452:357–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2452/24520357.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2452/24520357.pdf.

Bryant:2002:NAM

[1090] David Bryant and Vincent Moulton. NeighborNet: An agglomerative method for the construction of planar phylogenetic networks. Lecture Notes in Computer Science, 2452:375–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2452/24520375.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2452/24520375.pdf.

Leong:2002:CHN

[1091] H.-W. Leong, F. P. Preparata, W.- K. Sung, and H. Willy. On the control of hybridization noise in DNA sequencing-by-hybridization. Lecture Notes in Computer Science, 2452:392–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2452/24520392.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2452/24520392.pdf.

Skiena:2002:RSA

[1092] Steven Skiena and Sagi Snir. Restricting SBH ambiguity via restriction enzymes. Lecture Notes in Computer Science, 2452:404–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2452/24520404.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2452/24520404.pdf.

Shapiro:2002:ILM

[1093] Ehud Shapiro. Invited lecture — molecule as computation: Towards an abstraction of biomolecular systems. Lecture Notes in Computer Science, 2452:418–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2452/24520418.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2452/24520418.pdf.

Berman:2002:FOG

[1094] Piotr Berman, Paul Bertone, Bhaskar DasGupta, Mark Gerstein, Ming-Yang Kao, and Michael Snyder. Fast optimal genome tiling with applications to microarray design and homology search. Lecture Notes in Computer Science, 2452:419–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2452/24520419.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2452/24520419.pdf.

Rahmann:2002:RLS

[1095] Sven Rahmann. Rapid large-scale oligonucleotide selection for microarrays. Lecture Notes in Computer Science, 2452:434–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2452/24520434.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2452/24520434.pdf.

Kahng:2002:BLM

[1096] A. B. Kahng, I. I. Măndoiu, P. A.

Pevzner, S. Reda, and A. Z. Zelikovsky. Border length minimization in DNA array design. Lecture Notes in Computer Science, 2452:435–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2452/24520435.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2452/24520435.pdf.

Abouelhoda:2002:ESA

[1097] Mohamed Ibrahim Abouelhoda, Stefan Kurtz, and Enno Ohlebusch.

The enhanced suffix array and its applications to genome analysis. Lecture Notes in Computer Science, 2452:449–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2452/24520449.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2452/24520449.pdf.

Bergeron:2002:AGT

[1098] Anne Bergeron, Sylvie Corteel, and Mathieu Raffinot. The algorithmic of gene teams. Lecture Notes in Computer Science, 2452:464–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2452/24520464.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2452/24520464.pdf.

Warren:2002:CUS

[1099] Liling L. Warren and Ben Hui Liu. Combinatorial use of short probes for differential gene expression profiling. Lecture Notes in Computer Science, 2452:477–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2452/24520477.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2452/24520477.pdf.

Lipson:2002:DSO

[1100] Doron Lipson, Peter Webb, and Zohar Yakhini. Designing specific oligonucleotide probes for the entire S. cerevisiae transcriptome. Lecture Notes in Computer Science, 2452:491–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2452/24520491.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2452/24520491.pdf.

Bar-Joseph:2002:ACO

[1101] Ziv Bar-Joseph, Erik D. Demaine, David K. Gifford, Angèle M. Hamel, Tommi S. Jaakkola, and Nathan Srebro. K-ary clustering with optimal leaf ordering for gene expression data. Lecture Notes in Computer Science, 2452:506–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2452/24520506.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2452/24520506.pdf.

Moret:2002:IMO

[1102] Bernard M. E. Moret, Adam C. Siepel, Jijun Tang, and Tao Liu. Inversion medians outperform breakpoint medians in phylogeny reconstruction from geneorder data. Lecture Notes in Computer Science, 2452:521–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2452/24520521.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2452/24520521.pdf.

Page:2002:MMS

[1103] Roderic D. M. Page. Modified mincut supertrees. Lecture Notes in Computer Science, 2452:537–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2452/24520537.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2452/24520537.pdf.

Anonymous:2002:AIy

[1104] Anonymous. Author index. Lecture Notes in Computer Science, 2452: 553–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2452/2452auth.pdf.

Lenk:2002:EGW

[1105] Klaus Lenk and Roland Traunmüller. Electronic government: Where are we heading? Lecture Notes in Computer Science, 2456:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2456/24560001.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560001.pdf.

Kraaijenbrink:2002:CRP

[1106] Jeroen Kraaijenbrink. Centralization revisited? problems on implementing integrated service delivery in the netherlands. Lecture Notes in Computer Science, 2456: 10–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2456/24560010.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2456/24560010.
pdf.

Klumpp:2002:WGG

[1107] Dieter Klumpp. From Websites to e-government in Germany. Lecture Notes in Computer Science, 2456: 18–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2456/24560018.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2456/24560018.
pdf.

vanRossum:2002:BBC

[1108] Martin van Rossum, Daniele Chauvel, and Alasdair Mangham. BRAIN-CHILD, building a constituency for future research in knowledge management for local administrations. Lecture Notes in Computer Science, 2456:26–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560026.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560026.pdf.

Hoogwout:2002:OOS

[1109] Marcel Hoogwout. Organizing for online service delivery: The effects of network technology on the organization of transactional service delivery in Dutch local government. Lecture Notes in Computer Science, 2456:33–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2456/24560033.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560033.pdf.

Andersen:2002:PSP

[1110] Kim Viborg Andersen. Public sector process rebuilding using information systems. Lecture Notes in Computer Science, 2456:37–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2456/24560037.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560037.pdf.

Posch:2002:WNA

[1111] Reinhard Posch. What is needed to allow e-citizenship? Lecture Notes in Computer Science, 2456: 45–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2456/24560045.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2456/24560045.
                                        pdf.

Gronlund:2002:PSP

[1112] Åke Grönlund. Private sanctity — epractices overriding democratic rigor in e-voting. Lecture Notes in Computer Science, 2456:52–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2456/24560052.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560052.pdf.

Virili:2002:RPV

[1113] Francesco Virili and Maddalena Sorrentino. Reconfiguring the political value chain: The potential role of Web services. Lecture Notes in Computer Science, 2456:61–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560061.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560061.pdf.

Chen:2002:GAP

[1114] Xinxiang Chen. The E-GOV action plan in Beijing. Lecture Notes in Computer Science, 2456: 69–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2456/24560069.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2456/24560069.
pdf.

vanEngers:2002:PLV

[1115] Tom M. van Engers and Radboud A. W. Vanlerberghe. The POWER-light version: Improving legal quality under time pressure. Lecture Notes in Computer Science, 2456:75–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560075.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560075.pdf.

Gursch:2002:ISP

[1116] Benedikt Gursch, Christian Seel, and Öner Güngöz. Intranet “Saarland Plus“ — enabling new methods of cooperation within the ministerial administration. Lecture Notes in Computer Science, 2456:84–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2456/24560084.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560084.pdf.

Klein:2002:LG

[1117] Michel R. Klein and Jacques Dang. e-learning for e-government. Lecture Notes in Computer Science, 2456: 88–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2456/24560088.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2456/24560088.
                                        pdf.

Marciano:2002:MLI

[1118] Richard Marciano, Bertram Ludäscher, Ilya Zaslavsky, Reagan Moore, and Keith Pezzoli. Multi-level information modeling and preservation of eGOV data. Lecture Notes in Computer Science, 2456:93–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2456/24560093.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560093.pdf.

Durrant:2002:GIC

[1119] Fay Durrant. e-government and the Internet in the Caribbean: An initial assessment. Lecture Notes in Computer Science, 2456:101–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560101.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560101.pdf.

Fernandez:2002:TIA

[1120] Alejandro Fernández. Towards interoperability amongst European public administrations. Lecture Notes in Computer Science, 2456: 105–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2456/24560105.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2456/24560105.
pdf.

Heinderyckx:2002:AGI

[1121] François Heinderyckx. Assessing e-government implementation processes: a pan-European survey of administrations officials. Lecture Notes in Computer Science, 2456: 111–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2456/24560111.htm;   http:
//link.springer.de/link/service/

                                        series/0558/papers/2456/24560111.
                                        pdf.

Glassey:2002:OSG

[1122] Olivier Glassey. A one-stop government prototype based on use cases and scenarios. Lecture Notes in Computer Science, 2456:116–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2456/24560116.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560116.pdf.

Krenner:2002:RRG

[1123] Johanna Krenner. Reflections on the requirements gathering in an onestop government project. Lecture Notes in Computer Science, 2456: 124–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2456/24560124.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2456/24560124.
                                        pdf.

Klischewski:2002:UMF

[1124] Ralf Klischewski and Klaus Lenk. Understanding and modelling flexibility in administrative processes. Lecture Notes in Computer Science, 2456:129–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2456/24560129.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560129.pdf.

Falck:2002:BPM

[1125] Margrit Falck. Business process management — as a method of governance. Lecture Notes in Computer Science, 2456:137–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560137.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560137.pdf.

Boer:2002:PDL

[1126] Alexander Boer, Rinke Hoekstra, Radboud Winkels, Tom van Engers, and Frederik Willaert. Proposal for a Dutch legal XML Standard. Lecture Notes in Computer Science, 2456:142–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560142.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560142.pdf.

Leenes:2002:SME

[1127] Ronald Leenes and Jörgen Svensson. Size matters — electronic service delivery by municipalities? Lecture Notes in Computer Science, 2456:150–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560150.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560150.pdf.

Jacumeit:2002:ANM

[1128] Volker Jacumeit. Administration 2000 — networking municipal front and back offices for one-stop government. Lecture Notes in Computer Science, 2456:157–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560157.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560157.pdf.

Siegfried:2002:EGL

[1129] Tina Siegfried. The experience of German local communities with e-government — results of the MEDIA@Komm Project. Lecture Notes in Computer Science, 2456: 163–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2456/24560163.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2456/24560163.
pdf.

Ashford:2002:EPS

[1130] Ruth Ashford, Jennifer Rowley, and Frances Slack. Electronic public service delivery through online kiosks: The user’s perspective. Lecture Notes in Computer Science, 2456:169–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560169.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560169.pdf.

Riedl:2002:FSH

[1131] Reinhard Riedl and Nico Maibaum.

FASME — from Smartcards to holistic IT-architectures for interstate e-government. Lecture Notes in Computer Science, 2456:173– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2456/24560173.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2456/24560173.
pdf.

Gadda:2002:LGB

[1132] Lara Gadda and Alberto Savoldelli.

The local e-government best practice in Italian country: The case of the centralised desk of “Area Berica”. Lecture Notes in Computer Science, 2456:179–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560179.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560179.pdf.

Petrovic:2002:IFT

[1133] Otto Petrovic. The immanent fields of tension associated with e-government. Lecture Notes in Computer Science, 2456:187–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560187.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560187.pdf.

Wustinger:2002:VVC

[1134] Josef Wustinger, Gerhard Jakisch, Rolf Wohlmannstetter, and Rainer Riedel. VCRM — Vienna Citizen Request Management. Lecture Notes in Computer Science, 2456: 191–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2456/24560191.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2456/24560191.
pdf.

Bochicchio:2002:PPP

[1135] Mario A. Bochicchio and Antonella Longo. Public-private partnerships to manage local taxes: Information models and software tools. Lecture Notes in Computer Science, 2456:195–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560195.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560195.pdf.

Dobrev:2002:MEM

[1136] Bojil Dobrev, Mechthild Stoewer, Lambros Makris, and Eleonora Getsova. E-MuniS — electronic municipal information services — best practice transfer and improvement project: Project approach and intermediary results. Lecture Notes in Computer Science, 2456:199–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560199.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560199.pdf.

Costake:2002:SSG

[1137] Nicolae Costake. Some specific egovernment management problems in a transforming country. Lecture Notes in Computer Science, 2456: 207–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2456/24560207.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2456/24560207.
pdf.

Lopez:2002:TTF

[1138] J. Lopez, A. Maña, J. Montenegro, J. Ortega, and J. Troya. Towards a trustful and flexible environment for secure communications with public administrations. Lecture Notes in Computer Science, 2456:211–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560211.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560211.pdf.

Aedo:2002:SEM

[1139] Ignacio Aedo, Paloma Díaz, Camino Fernández, and Jorge de Castro. Supporting efficient multinational disaster response through a Web-based system. Lecture Notes in Computer Science, 2456:215–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2456/24560215.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560215.pdf.

Gadda:2002:KBI

[1140] Lara Gadda, Emilio Bugli Innocenti, and Alberto Savoldelli. KIWI: Building innovative knowledge management infrastructure within European public administration. Lecture Notes in Computer Science, 2456: 223–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2456/24560223.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2456/24560223.
                                        pdf.

Kocman:2002:ESE

[1141] Viktorija Kocman, Angela Stöger-Frank, and Simone Ulreich. Elektronische Steuer Erlass Dokumentation a documentation on official tax guide lines. Lecture Notes in Computer Science, 2456:230–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2456/24560230.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560230.pdf.

Watson:2002:VNM

[1142] Anthony Watson and Vincent Cordonnier. Voting in the new millennium: eVoting holds the promise to expand citizen choice. Lecture Notes in Computer Science, 2456:

234–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2456/24560234.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2456/24560234.
pdf.

Westholm:2002:DGA

[1143] Hilmar Westholm. e-democracy goes ahead. the Internet as a tool for improving deliberative policies? Lecture Notes in Computer Science, 2456:240–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560240.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560240.pdf.

Gordon:2002:DSS

[1144] Thomas F. Gordon and Gernot Richter. Discourse support systems for deliberative democracy. Lecture Notes in Computer Science, 2456:248–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560248.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560248.pdf.

Macintosh:2002:CPP

[1145] Ann Macintosh and Ella Smith.

Citizen participation in public affairs. Lecture Notes in Computer Science, 2456:256–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2456/24560256.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560256.pdf.

Gouscos:2002:AOO

[1146] Dimitris Gouscos, Giorgos Laskaridis, Dimitris Lioulias, Gregoris Mentzas, and Panagiotis Georgiadis. An approach to offering one-stop egovernment services — available technologies and architectural issues. Lecture Notes in Computer Science, 2456:264–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2456/24560264.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560264.pdf.

Maioli:2002:GLS

[1147] Cesare Maioli. e-governance for local system: a plan and implementation experience. Lecture Notes in Computer Science, 2456:272–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2456/24560272.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560272.pdf.

Vassilakis:2002:TGS

[1148] C. Vassilakis, G. Laskaridis, G. Lepouras, S. Rouvas, and P. Georgiadis. Transactional e-government services: An integrated approach. Lecture Notes in Computer Science, 2456:276–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560276.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560276.pdf.

Monnoyer-Smith:2002:EVI

[1149] Laurence Monnoyer-Smith and Eric Maigret. Electronic vote and Internet campaigning: State of the art in Europe and remaining questions. Lecture Notes in Computer Science, 2456:280–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560280.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560280.pdf.

Maibaum:2002:CDA

[1150] Nico Maibaum, Igor Sedov, and Clemens H. Cap. A citizen digital assistant for e-government. Lecture Notes in Computer Science, 2456: 284–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2456/24560284.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2456/24560284.
pdf.

Paralic:2002:SSD

[1151] Jan Paralic, Tomas Sabol, and Marian Mach. A system to support edemocracy. Lecture Notes in Computer Science, 2456:288–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2456/24560288.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560288.pdf.

Edl:2002:PAP

[1152] Anton Edl. IST-Project: AIDA — A platform for digital administration. Lecture Notes in Computer Science, 2456:292–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2456/24560292.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560292.pdf.

Millard:2002:GSB

[1153] Jeremy Millard. e-government strategies: Best practice reports from the European front line. Lecture Notes in Computer Science, 2456: 298–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2456/24560298.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2456/24560298.
                                        pdf.

Anagnostakis:2002:CCI

[1154] A. Anagnostakis, G. C. Sakellaris, M. Tzima, D. I. Fotiadis, and A. Likas. CITATION — citizen information tool in smart administration. Lecture Notes in Computer Science, 2456:307–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560307.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560307.pdf.

Frenkiel:2002:CCS

[1155] Michel Frenkiel, Paul Grison, and Philippe Laluyaux. Clip Card: Smart Card based traffic tickets. Lecture Notes in Computer Science, 2456:313–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560313.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560313.pdf.

Drion:2002:VAO

[1156] Benoît Drion and Norbert Benamou.

VISUAL ADMIN — opening administration information systems to citizens. Lecture Notes in Computer Science, 2456:319–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560319.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560319.pdf.

Dawkins:2002:GO

[1157] Freddie Dawkins. e-government observatory. Lecture Notes in Computer Science, 2456:326–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560326.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560326.pdf.

Tarabanis:2002:RTP

[1158] Konstantinos Tarabanis and Vassilios Peristeras. Requirements for transparent public services provision amongst public administrations. Lecture Notes in Computer Science, 2456:330–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2456/24560330.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560330.pdf.

Legal:2002:CBC

[1159] Maria Legal, Gregoris Mentzas, Dimitris Gouscos, and Panagiotis Georgiadis. CB-BUSINESS: Cross-border business intermediation through electronic seamless services. Lecture Notes in Computer Science, 2456:338–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2456/24560338.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560338.pdf.

Moussalli:2002:BDD

[1160] Antoinette Moussalli and Christopher Stokes. Bridging the digital divide with AVANTI technology. Lecture Notes in Computer Science, 2456:344–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560344.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560344.pdf.

Tambouris:2002:IPT

[1161] Efhimios Tambouris. An integrated platform for tele-voting and teleconsulting within and across European cities: The EURO-CITI project. Lecture Notes in Computer Science, 2456:350–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560350.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560350.pdf.

Ioannidis:2002:ECS

[1162] A. Ioannidis, M. Spanoudakis, G. Priggouris, C. Eliopoulou, S. Hadjiefthymiades, and L. Merakos. EURO-CITI security manager: Supporting transaction services in the e-government domain. Lecture Notes in Computer Science, 2456:358–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560358.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560358.pdf.

Georgiadis:2002:SKB

[1163] P. Georgiadis, G. Lepouras, C. Vassilakis, G. Boukis, E. Tambouris, S. Gorilas, E. Davenport, A. Macintosh, J. Fraser, and D. Lochhead. SmartGov: a knowledge-based platform for transactional electronic services. Lecture Notes in Computer Science, 2456:362–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2456/24560362.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560362.pdf.

Makolm:2002:BPG

[1164] Josef Makolm. Best practice in egovernment. Lecture Notes in Computer Science, 2456:370–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.de/link/service/
                                        series/0558/bibs/2456/24560370.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560370.pdf.

Diez:2002:GAJ

[1165] Carmen Diez and Javier Prenafeta. egovernment applied to judicial notices and inter-registrar communications in the European Union: The AEQUITAS Project. Lecture Notes in Computer Science, 2456:375–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2456/24560375.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560375.pdf.

Vintar:2002:CAL

[1166] Mirko Vintar and Anamarija Leben. The concepts of an active lifeevent public portal. Lecture Notes in Computer Science, 2456:383– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2456/24560383.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2456/24560383.
pdf.

Pfaff:2002:NST

[1167] Donovan Pfaff and Bernd Simon.

New services through integrated egovernment. Lecture Notes in Computer Science, 2456:391–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/link/service/
series/0558/bibs/2456/24560391.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560391.pdf.

Evangelidis:2002:RAS

[1168] A. Evangelidis, J. Akomode, A. Taleb-Bendiab, and M. Taylor. Risk assessment & success factors for egovernment in a UK establishment. Lecture Notes in Computer Science, 2456:395–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560395.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560395.pdf.

Hoegler:2002:QVG

[1169] Tamara Hoegler and Thilo Schuster. Quo vadis e-government? — A trap between unsuitable technologies and deployment strategies. Lecture Notes in Computer Science, 2456:403–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2456/24560403.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560403.pdf.

Yammine:2002:NAP

[1170] Anne Yammine. A new approach to the phenomenon of egovernment: Analysis of the public discourse on e-government in Switzerland. Lecture Notes in Computer Science, 2456:407–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2456/24560407.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560407.pdf.

Galindo:2002:SRG

[1171] Fernando Galindo. Self-regulation in e-government: a step more. Lecture Notes in Computer Science, 2456:411–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2456/24560411.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560411.pdf.

Leith:2002:UOF

[1172] Philip Leith and John Morison. UK online: Forcing citizen involvement into a technically-oriented framework? Lecture Notes in Computer Science, 2456:419–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560419.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560419.pdf.

Saarenpaa:2002:DSF

[1173] Ahti Saarenpää. Data security: a fundamental right in the e-society? Lecture Notes in Computer Science, 2456:424–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560424.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560424.pdf.

Brunschwig:2002:LDG

[1174] Colette Brunschwig. Legal design and e-government: Visualisations of cost & efficiency accounting in the wif! e-learning environment of the Canton of Zurich (Switzerland). Lecture Notes in Computer Science, 2456:430–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560430.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560430.pdf.

Augustinaitis:2002:FSG

[1175] Arūnas Augustinaitis and Rimantas Petrauskas. The first steps of egovernance in Lithuania: From theory to practice. Lecture Notes in Computer Science, 2456:438–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2456/24560438.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560438.pdf.

Menzel:2002:RCC

[1176] Thomas Menzel and Peter Reichstädter. The role of citizen cards in e-government. Lecture Notes in Computer Science, 2456: 446–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2456/24560446.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2456/24560446.
                                        pdf.

Klewitz-Hommelsen:2002:IPV

[1177] Sayeed Klewitz-Hommelsen. Indicators for privacy violation of Internet sites. Lecture Notes in Computer Science, 2456:456–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2456/24560456.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560456.pdf.

Desmedt:2002:VDP

[1178] Yvo Desmedt and Brian King. Verifiable democracy a protocol to secure an electronic legislature. Lecture Notes in Computer Science, 2456:460–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560460.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560460.pdf.

Hof:2002:AHO

[1179] Sonja Hof. Arguments for a holistic and open approach to secure e-government. Lecture Notes in Computer Science, 2456:464–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560464.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560464.pdf.

Staniszkis:2002:SAK

[1180] Witold Staniszkis. Supporting administrative knowledge processes. Lecture Notes in Computer Science, 2456:468–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560468.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2456/24560468.pdf.

Pasic:2002:IIM

[1181] Aljosa Pasic, Sara Diez, and Jose Antonio Espinosa. IMPULSE: Interworkflow model for e-government. Lecture Notes in Computer Science, 2456:472–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2456/24560472.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2456/24560472.pdf.

Wagenaar:2002:VIC

[1182] René Wagenaar and Marijn Janssen. Visualization of the implications of a component based ICT architecture for service provisioning. Lecture Notes in Computer Science, 2456:480–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2456/24560480.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2456/24560480.pdf.

Anonymous:2002:AIz

[1183] Anonymous. Author index. Lecture Notes in Computer Science, 2456: 485–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2456/2456auth.pdf.

Rahayu:2002:PAO

[1184] Johanna Wenny Rahayu and David Taniar. Preserving aggregation in an object-relational DBMS. Lecture Notes in Computer Science, 2457: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2457/24570001.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2457/24570001.
                                        pdf.

Hoque:2002:DCU

[1185] Abu Sayed Md. Latiful Hoque, Douglas McGregor, and John Wilson. Database compression using an offline dictionary method. Lecture Notes in Computer Science, 2457:11–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2457/24570011.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2457/24570011.pdf.

Chountas:2002:RTU

[1186] Panagiotis Chountas, Ilias Petrounias, Krassimir Atanassov, Vassilis Kodogiannis, and Elia El-Darzi. Representation of temporal unawareness. Lecture Notes in Computer Science, 2457:21–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2457/24570021.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2457/24570021.pdf.

Khan:2002:SDG

[1187] Sharifullah Khan and Peter L.

Mott. Scalable and dynamic grouping of continual queries. Lecture Notes in Computer Science, 2457: 31–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2457/24570031.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2457/24570031.
pdf.

Tossebro:2002:USD

[1188] Erlend Tøssebro and Mads Nygård. Uncertainty in spatiotemporal databases. Lecture Notes in Computer Science, 2457:43–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2457/24570043.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2457/24570043.pdf.

Dominguez:2002:ICE

[1189] Eladio Domínguez, Jorge Lloret, and María Antonia Zapata. Integrity constraint enforcement by means of trigger templates. Lecture Notes in Computer Science, 2457:54–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2457/24570054.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2457/24570054.pdf.

Jensen:2002:CLI

[1190] Ole Guttorm Jensen and Michael Böhlen. Current, legacy, and invalid tuples in conditionally evolving databases. Lecture Notes in Computer Science, 2457:65–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2457/24570065.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2457/24570065.pdf.

Jezek:2002:MSM

[1191] Karel Ježek and Martin Zíma.

Magic sets method with fuzzy logic. Lecture Notes in Computer Science, 2457:83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2457/24570083.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2457/24570083.pdf.

Bitirim:2002:IRE

[1192] Yiltan Bitirim, Yasar Tonta, and Hayri Sever. Information retrieval effectiveness of Turkish search engines. Lecture Notes in Computer Science, 2457:93–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2457/24570093.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2457/24570093.pdf.

Gokcen:2002:CLD

[1193] Ibrahim Gokcen and Jing Peng.

Comparing linear discriminant analysis and support vector machines. Lecture Notes in Computer Science, 2457:104–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2457/24570104.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2457/24570104.pdf.

Sadat:2002:CLI

[1194] Fatiha Sadat, Masatoshi Yoshikawa, and Shunsuke Uemura. Cross-language information retrieval using multiple resources and combinations for query expansion. Lecture Notes in Computer Science, 2457:114–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2457/24570114.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2457/24570114.pdf.

Jiang:2002:ESF

[1195] Jianmin Jiang, Baofeng Guo, and Pengjie Li. Extracting shape features in JPEG-2000 compressed images. Lecture Notes in Computer Science, 2457:123–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2457/24570123.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2457/24570123.pdf.

Sever:2002:CNT

[1196] Hayri Sever and Mehmet R. Tolun. Comparison of normalization techniques for metasearch. Lecture Notes in Computer Science, 2457: 133–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2457/24570133.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2457/24570133.
pdf.

Dalkilic:2002:CPF

[1197] Mehmet Emin Dalkiliç and Gökhan Dalkiliç. On the cryptographic patterns and frequencies in Turkish language. Lecture Notes in Computer Science, 2457:144–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2457/24570144.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2457/24570144.pdf.

Cho:2002:ASI

[1198] Sehyeong Cho and Seung-Soo Han.

Automatic stemming for indexing of an agglutinative language. Lecture Notes in Computer Science, 2457: 154–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2457/24570154.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2457/24570154.
pdf.

Fang:2002:PAC

[1199] Xiaoshan Fang and Huanye Sheng.

Pattern acquisition for Chinese named entity recognition: a supervised learning approach. Lecture Notes in Computer Science, 2457:166– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2457/24570166.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2457/24570166.
pdf.

Marchuk:2002:ISC

[1200] Alexander Marchuk, Andrey Nemov, Konstantin Fedorov, and Sergey Antyoufeev. The information system for creating and maintaining the electronic archive of documents. Lecture Notes in Computer Science, 2457:176–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2457/24570176.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2457/24570176.pdf.

Saykol:2002:KKB

[1201] Ediz Saykol, Gürcan Gülesir, Uğur Güdükbay, and Özgür Ulusoy. KiMPA: a kinematics-based method for polygon approximation. Lecture Notes in Computer Science, 2457:186– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2457/24570186.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2457/24570186.
pdf.

Yildirim:2002:IRR

[1202] Sule Yildirim and Turhan Tunali. The implementation of a robotic replanning framework. Lecture Notes in Computer Science, 2457:195–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2457/24570195.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2457/24570195.pdf.

Dalkilic:2002:MTC

[1203] Gökhan Dalkiliç and Yalçin Çebi. A 300 MB Turkish corpus and word analysis. Lecture Notes in Computer Science, 2457:205–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2457/24570205.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2457/24570205.pdf.

Lee:2002:ANS

[1204] Dong Ho Lee, Jung Sup Kim, Soo Duk Kim, Ki Chang Kim, Yoo-Sung Kim, and Jaehyun Park. Adaptation of a neighbor selection Markov chain for prefetching tiled Web GIS data. Lecture Notes in Computer Science, 2457:213–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2457/24570213.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2457/24570213.pdf.

Bulgun:2002:WBA

[1205] Ender Yazgan Bulgun, Alp Kut, Güngör Basser, and Mustafa Kasap. Web based automation software for calculating production costs in apparel industry. Lecture Notes in Computer Science, 2457:223–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2457/24570223.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2457/24570223.pdf.

Paprzycki:2002:KRA

[1206] Marcin Paprzycki, Austin Gilbert, and Minor Gordon. Knowledge representation in the agent-based travel support system. Lecture Notes in Computer Science, 2457:232–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2457/24570232.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2457/24570232.pdf.

Codognet:2002:IAV

[1207] Philippe Codognet. Intelligent agents in virtual worlds. Lecture Notes in Computer Science, 2457: 242–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2457/24570242.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2457/24570242.
                                        pdf.

Aragao:2002:CWS

[1208] Marcelo A. T. Aragão and Alvaro A. A. Fernandes. Characterizing Web service substitutivity with combined deductive and inductive engines. Lecture Notes in Computer Science, 2457:244–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2457/24570244.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2457/24570244.pdf.

Gultekin:2002:MFC

[1209] Irfan Gültekin and Ahmet Arslan. Modular-fuzzy cooperation algorithm for multi-agent systems. Lecture Notes in Computer Science, 2457:255–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2457/24570255.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2457/24570255.pdf.

Kilic:2002:MFQ

[1210] Alper Kilic and Ahmet Arslan. Minimax fuzzy Q-learning in cooperative multi-agent systems. Lecture Notes in Computer Science, 2457: 264–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2457/24570264.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2457/24570264.
pdf.

Chiao:2002:CBR

[1211] Hsin-Ta Chiao, Ming-Chun Cheng, Yue-Shan Chang, and Shyan-MingYuan. A component-based, reconfigurable mobile agent system for context-aware computing. Lecture Notes in Computer Science, 2457:273–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2457/24570273.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2457/24570273.pdf.

Erdur:2002:FCA

[1212] Riza Cenk Erdur and Oğuz Dikenelli. A FIPA-compliant agent framework with an extra layer for ontology dependent reusable behaviour. Lecture Notes in Computer Science, 2457:283–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2457/24570283.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2457/24570283.pdf.

Ermis:2002:VGA

[1213] Murat Ermis, Füsun Ülengin, and Abdurrahman Hacioğlu. Vibrational genetic algorithm (vga) for solving continuous covering location problems. Lecture Notes in Computer Science, 2457:293–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2457/24570293.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2457/24570293.pdf.

Nedjah:2002:MAS

[1214] Nadia Nedjah and Luiza de Macedo Mourelle. Minimal addition-subtraction chains using genetic algorithms. Lecture Notes in Computer Science, 2457: 303–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2457/24570303.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2457/24570303.
pdf.

Etaner-Uyar:2002:PDT

[1215] A. Sima Etaner-Uyar and A. Emre Harmanci. Preserving diversity through diploidy and meiosis for improved genetic algorithm performance in dynamic environments. Lecture Notes in Computer Science, 2457:314–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2457/24570314.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2457/24570314.pdf.

Yakhno:2002:ASA

[1216] Tatyana Yakhno and Emine Ekin.

Ant systems: Another alternative for optimization problems? Lecture Notes in Computer Science, 2457: 324–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2457/24570324.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2457/24570324.
pdf.

Kemikli:2002:AOP

[1217] Erdal Kemikli and Nadia Erdogan. Augmenting object persistency paradigm for faster server development. Lecture Notes in Computer Science, 2457:327–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2457/24570327.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2457/24570327.pdf.

Choi:2002:PCD

[1218] Jungwan Choi, Youjip Won, and S. W. Nam. Power conscious disk scheduling for multimedia data retrieval. Lecture Notes in Computer Science, 2457:336–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2457/24570336.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2457/24570336.pdf.

Topcuoglu:2002:TSC

[1219] Haluk Topcuoglu and Can Sevilmis. Task scheduling with conflicting objectives. Lecture Notes in Computer Science, 2457:346–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2457/24570346.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2457/24570346.pdf.

Hwang:2002:FAS

[1220] Sung-Ho Hwang and Ki-Jun Han. A fast access scheme to meet delay requirement for wireless access network. Lecture Notes in Computer Science, 2457:356–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2457/24570356.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2457/24570356.pdf.

Kut:2002:DEU

[1221] Alp Kut, Sen Çakir, and Ender Yazgan Bulgun. Dokuz Eylul University — distance education utilities model. Lecture Notes in Computer Science, 2457:366–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2457/24570366.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2457/24570366.pdf.

Grabinger:2002:PBL

[1222] Scott Grabinger and Joanna C.

Dunlap. Problem-based learning as an example of active learning and student engagement. Lecture Notes in Computer Science, 2457: 375–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2457/24570375.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2457/24570375.
pdf.

Gulsecen:2002:UPM

[1223] Sevinç Gülseçen. Use of PBL method in teaching IT to students from a faculty of education: a case study. Lecture Notes in Computer Science, 2457:385–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2457/24570385.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2457/24570385.pdf.

Bulgak:2002:IMV

[1224] Ayse Bulgak and Diliaver Eminov. Interval Matrix Vector Calculator — the iMVC 1.0. Lecture Notes in Computer Science, 2457:386–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2457/24570386.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2457/24570386.pdf.

Demir:2002:ECD

[1225] Deniz Demir, Haluk Topcuoglu, and Ergun Kavukcu. Efficient code deployment for heterogeneous distributed data sources. Lecture Notes in Computer Science, 2457: 395–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2457/24570395.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2457/24570395.
                                        pdf.

Nedjah:2002:EPM

[1226] Nadia Nedjah and Luiza de Macedo Mourelle. Efficient parallel modular exponentiation algorithm. Lecture Notes in Computer Science, 2457:405– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2457/24570405.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2457/24570405.
                                        pdf.

Chang:2002:ITE

[1227] Yu-Sug Chang, Hye-Jung Lee, Doo-Soon Park, and Im-Yeong Lee. Interprocedural transformations for extracting maximum parallelism. Lecture Notes in Computer Science, 2457: 415–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2457/24570415.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2457/24570415.
pdf.

Bebel:2002:MMO

[1228] Bartosz Bebel, Zbyszko Królikowski, and Robert Wrembel. On methods’ materialization in object-relational data warehouse. Lecture Notes in Computer Science, 2457:425– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2457/24570425.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2457/24570425.
pdf.

Anonymous:2002:AIba

[1229] Anonymous. Author index. Lecture Notes in Computer Science, 2457: 435–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2457/2457auth.pdf.

Abiteboul:2002:FEA

[1230] S. Abiteboul, G. Cobéna, J. Masanes, and G. Sedrati. A first experience in archiving the French Web.

Lecture Notes in Computer Science, 2458:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580001.pdf.

Rauber:2002:AOA

[1231] Andreas Rauber, Andreas Aschenbrenner, and Oliver Witvoet. Austrian online archive processing: Analyzing archives of the World Wide Web. Lecture Notes in Computer Science, 2458:16–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580016.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580016.pdf.

Ko:2002:CED

[1232] Seung-Kyu Ko, Myoung-Soo Kang, Won-Sung Sohn, Soon-Bum Lim, and Yoon-Chul Choy. Conversion of eBook documents based on mapping relations. Lecture Notes in Computer Science, 2458:32–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580032.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580032.pdf.

Wilson:2002:GDE

[1233] Ruth Wilson, Monica Landoni, and Forbes Gibb. Guidelines for designing electronic books. Lecture Notes in Computer Science, 2458: 47–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2458/24580047.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2458/24580047.
pdf.

Sun:2002:PCK

[1234] Aixin Sun, Ee-Peng Lim, and Wee-Keong Ng. Personalized classification for keyword-based category profiles. Lecture Notes in Computer Science, 2458:61–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580061.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580061.pdf.

Takasu:2002:SAB

[1235] Atsuhiro Takasu. Statistical analysis of bibliographic strings for constructing an integrated document space. Lecture Notes in Computer Science, 2458:75–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580075.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580075.pdf.

Bergmark:2002:FCT

[1236] Donna Bergmark, Carl Lagoze, and Alex Sbityakov. Focused crawls, tunneling, and digital libraries. Lecture Notes in Computer Science, 2458: 91–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2458/24580091.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2458/24580091.
                                        pdf.

Bolchini:2002:GOR

[1237] Davide Bolchini and Paolo Paolini. Goal-oriented requirements specification for digital libraries. Lecture Notes in Computer Science, 2458: 107–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2458/24580107.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2458/24580107.
                                        pdf.

Heggland:2002:OTA

[1238] Jon Heggland. OntoLog: Temporal annotation using ad hoc ontologies and application profiles. Lecture Notes in Computer Science, 2458:118–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2458/24580118.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2458/24580118.pdf.

Goncalves:2002:XLS

[1239] Marcos André Gonçalves, Ming Luo, Rao Shen, Mir Farooq Ali, and Edward A. Fox. An XML log standard and tool for digital library logging analysis. Lecture Notes in Computer Science, 2458:129–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580129.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580129.pdf.

VandeSompel:2002:NIF

[1240] Herbert Van de Sompel and Carl Lagoze. Notes from the interoperability front: a progress report on the open archives initiative. Lecture Notes in Computer Science, 2458:144–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580144.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580144.pdf.

Little:2002:DGI

[1241] Suzanne Little, Joost Geurts, and Jane Hunter. Dynamic generation of intelligent multimedia presentations through semantic inferencing. Lecture Notes in Computer Science, 2458:158–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580158.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580158.pdf.

Liu:2002:TRI

[1242] Xiaoming Liu, Kurt Maly, Mohammad Zubair, Rong Tang, Mohammed Imran Padshah, George Roncaglia, JoAnne Rocker, Michael Nelson, William von Ofenheim, Richard Luce, Jacqueline Stack, Frances Knudson, Beth Goldsmith, Irma Holtkamp, Miriam Blake, Jack Carter, Mariella Di Giacomo, Major Jerome Nutter, Susan Brown, Ron Montbrand, Sally Landenberger, Kathy Pierson, Vince Duran, and Beth Moser. Technical report interchange through synchronized OAI caches. Lecture Notes in Computer Science, 2458:176–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580176.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580176.pdf.

Khoo:2002:FRO

[1243] Michael Khoo, Holly Devaul, and Tamara Sumner. Functional requirements for online tools to support community-led collections building. Lecture Notes in Computer Science, 2458:190–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580190.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580190.pdf.

Kwak:2002:SEM

[1244] Byeong Heui Kwak, Woochun Jun, Le Gruenwald, and Suk-Ki Hong. A study on the evaluation model for university libraries in digital environments. Lecture Notes in Computer Science, 2458:204–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580204.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580204.pdf.

Huxley:2002:RFF

[1245] Lesly Huxley. Renardus: Following the fox from project to service. Lecture Notes in Computer Science, 2458:218–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580218.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580218.pdf.

Kiemle:2002:DAD

[1246] Stephan Kiemle. From digital archive to digital library — A middleware for earth-observation data management. Lecture Notes in Computer Science, 2458:230–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580230.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580230.pdf.

Aalberg:2002:NBC

[1247] Trond Aalberg. Navigating in bibliographic catalogues. Lecture Notes in Computer Science, 2458: 238–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/

                                        0558/bibs/2458/24580238.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2458/24580238.
                                        pdf.

Castelli:2002:FMQ

[1248] Donatella Castelli, Carlo Meghini, and Pasquale Pagano. Foundations of a multidimensional query language for digital libraries. Lecture Notes in Computer Science, 2458: 251–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2458/24580251.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2458/24580251.
                                        pdf.

Smeaton:2002:TVT

[1249] Alan F. Smeaton, Paul Over, Cash J. Costello, Arjen P. de Vries, David Doermann, Alexander Hauptmann, Mark E. Rorvig, John R. Smith, and Lide Wu. The TREC2001 video track: Information retrieval on digital video information. Lecture Notes in Computer Science, 2458: 266–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2458/24580266.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2458/24580266.
                                        pdf.

Jones:2002:AAA

[1250] Gareth J. F. Jones and Richard J. Edens. Automated alignment and annotation of audio-visual presentations. Lecture Notes in Computer Science, 2458:276–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580276.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580276.pdf.

Castelli:2002:ODL

[1251] Donatella Castelli and Pasquale Pagano. OpenDLib: a digital library service system. Lecture Notes in Computer Science, 2458: 292–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2458/24580292.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2458/24580292.
pdf.

Apps:2002:PDL

[1252] Ann Apps and Ross MacIntyre. Prototyping digital library technologies in zetoc. Lecture Notes in Computer Science, 2458:309–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580309.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580309.pdf.

Karadkar:2002:ESB

[1253] Unmil P. Karadkar, Jin-Cheon Na, and Richard Furuta. Employing smart browsers to support flexible information presentation in petri net-based digital libraries. Lecture Notes in Computer Science, 2458:324–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2458/24580324.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2458/24580324.pdf.

Ruthven:2002:UEM

[1254] Ian Ruthven. On the use of explanations as mediating device for relevance feedback. Lecture Notes in Computer Science, 2458:338–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2458/24580338.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2458/24580338.pdf.

Blocks:2002:QET

[1255] D. Blocks, C. Binding, D. Cunliffe, and D. Tudhope. Qualitative evaluation of thesaurus-based retrieval. Lecture Notes in Computer Science, 2458:346–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2458/24580346.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2458/24580346.pdf.

Strotgen:2002:MDE

[1256] Robert Strötgen. Meta-data extraction and query translation. treatment of semantic heterogeneity. Lecture Notes in Computer Science, 2458:362–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580362.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580362.pdf.

Makedon:2002:MDL

[1257] Fillia Makedon, James Ford, Li Shen, Tilmann Steinberg, Andrew Saykin, Heather Wishart, and Sarantos Kapidakis. MetaDL: a digital library of metadata for sensitive or complex research data. Lecture Notes in Computer Science, 2458: 374–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2458/24580374.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2458/24580374.
pdf.

Witten:2002:IDM

[1258] Ian H. Witten, David Bainbridge, Gordon Paynter, and Stefan Boddie. Importing documents and metadata into digital libraries: Requirements analysis and an extensible architecture. Lecture Notes in Computer Science, 2458:390–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580390.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580390.pdf.

Payette:2002:MFP

[1259] Sandra Payette and Thornton Staples. The mellon fedora project.

Lecture Notes in Computer Science, 2458:406–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2458/24580406.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2458/24580406.pdf.

Xi:2002:HPI

[1260] Wensi Xi, Ohm Sornil, Ming Luo, and Edward A. Fox. Hybrid partition inverted files: Experimental validation. Lecture Notes in Computer Science, 2458:422–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2458/24580422.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2458/24580422.pdf.

Bollen:2002:DLE

[1261] Johan Bollen, Somasekhar Vemulapalli, and Weining Xu. Digital library evaluation by analysis of user retrieval patterns. Lecture Notes in Computer Science, 2458:432–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2458/24580432.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2458/24580432.pdf.

Papadakis:2002:ISR

[1262] Ioannis Papadakis, Ioannis Andreou, and Vassileios Chrissikopoulos. Interactive search results. Lecture Notes in Computer Science, 2458:448–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580448.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580448.pdf.

Jones:2002:IMM

[1263] Gareth J. F. Jones and Adenike M.

Lam-Adesina. An investigation of mixed-media information retrieval. Lecture Notes in Computer Science, 2458:463–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580463.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580463.pdf.

Orio:2002:APS

[1264] Nicola Orio. Alignment of performances with scores aimed at content-based music access and retrieval. Lecture Notes in Computer Science, 2458:479–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580479.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580479.pdf.

Wildemuth:2002:ASV

[1265] Barbara M. Wildemuth, Gary Marchionini, Todd Wilkens, Meng Yang, Gary Geisler, Beth Fowler, Anthony Hughes, and Xiangming Mu. Alternative surrogates for video objects in a digital library: Users’ perspectives on their relative usability. Lecture Notes in Computer Science, 2458:493–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2458/24580493.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2458/24580493.pdf.

Serralheiro:2002:WAD

[1266] António Serralheiro, Diamantino Caseiro, Hugo Meinedo, and Isabel Trancoso. Word alignment in digital talking books using WFSTs. Lecture Notes in Computer Science, 2458:508–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2458/24580508.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2458/24580508.pdf.

Mellor:2002:MRP

[1267] Phil Mellor, Paul Wheatley, and Derek Sergeant. Migration on request, a practical technique for preservation. Lecture Notes in Computer Science, 2458:516–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2458/24580516.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2458/24580516.pdf.

Koubarakis:2002:IAD

[1268] M. Koubarakis, T. Koutris, C. Tryfonopoulos, and P. Raftopoulou. Information alert in distributed digital libraries: The models, languages, and architecture of DIAS. Lecture Notes in Computer Science, 2458:527–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580527.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580527.pdf.

Smith:2002:DIR

[1269] MacKenzie Smith. DSpace: An institutional repository from the MIT libraries and hewlett packard laboratories. Lecture Notes in Computer Science, 2458:543–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580543.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580543.pdf.

Sfakakis:2002:UBT

[1270] Michalis Sfakakis and Sarantos Kapidakis. User behavior tendencies on data collections in a digital library. Lecture Notes in Computer Science, 2458:550–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580550.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580550.pdf.

Greenberg:2002:SCC

[1271] Jane Greenberg, Kristen A. Bullard, M. L. James, Evelyn Daniel, and Peter White. Student comprehension of classification applications in a science education digital library. Lecture Notes in Computer Science, 2458:560–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2458/24580560.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2458/24580560.pdf.

Suleman:2002:DPS

[1272] Hussein Suleman and Edward A. Fox. Designing protocols in support of digital library componentization. Lecture Notes in Computer Science, 2458:568–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2458/24580568.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2458/24580568.pdf.

Buchanan:2002:ESS

[1273] George Buchanan, Matt Jones, and Gary Marsden. Exploring small screen digital library access with the greenstone digital library. Lecture Notes in Computer Science, 2458: 583–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2458/24580583.htm;   http:
                                        //link.springer.de/link/service/

series/0558/papers/2458/24580583.
pdf.

Fuhr:2002:DID

[1274] Norbert Fuhr, Claus-Peter Klas, André Schaefer, and Peter Mutschke. Daffodil: An integrated desktop for supporting high-level search activities in federated digital libraries. Lecture Notes in Computer Science, 2458:597–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580597.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580597.pdf.

Bontcheva:2002:UHL

[1275] Kalina Bontcheva, Diana Maynard, Hamish Cunningham, and Horacio Saggion. Using human language technology for automatic annotation and indexing of digital library content. Lecture Notes in Computer Science, 2458:613–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580613.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580613.pdf.

Crane:2002:CHD

[1276] Gregory Crane. Cultural heritage digital libraries: Needs and components. Lecture Notes in Computer Science, 2458:626–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2458/24580626.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2458/24580626.pdf.

Monroy:2002:VVT

[1277] Carlos Monroy, Rajiv Kochumman, Richard Furuta, Eduardo Urbina, Eréndira Melgoza, and Arpita Goenka. Visualization of variants in textual collations to analyze the evolution of literary works in the cervantes project. Lecture Notes in Computer Science, 2458:638–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2458/24580638.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2458/24580638.pdf.

dePolo:2002:AOA

[1278] Andrea de Polo and Sam Minelli. Alinari online: Access quality through a cultural photographic heritage site. Lecture Notes in Computer Science, 2458:654–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2458/24580654.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2458/24580654.pdf.

Bertino:2002:ACS

[1279] Elisa Bertino, Elena Ferrari, and Andrea Perego. An access control system for digital libraries and the Web: The maX prototype demonstration. Lecture Notes in Computer Science, 2458:656–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580656.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580656.pdf.

Bontcheva:2002:HLT

[1280] Kalina Bontcheva and Hamish Cunningham. Human language technology for automatic annotation and indexing of digital library content. Lecture Notes in Computer Science, 2458:658–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580658.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580658.pdf.

Mastidoro:2002:IDL

[1281] Nicola Mastidoro. The IntraText digital library: XML-driven online library based on high accessibility, lexical hypertextualization and scholarly accuracy in philological/textual notations. Lecture Notes in Computer Science, 2458:659–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2458/24580659.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2458/24580659.pdf.

Muhlberger:2002:MPA

[1282] Günter Mühlberger and Birgit Stehno.

The metae project — automated digitisation of books and journals.

Lecture Notes in Computer Science, 2458:660–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2458/24580660.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2458/24580660.pdf.

Bordoni:2002:CCC

[1283] Luciana Bordoni. COVAX: a contemporary culture virtual archive in XML. Lecture Notes in Computer Science, 2458:661–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2458/24580661.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2458/24580661.pdf.

Anonymous:2002:AIbb

[1284] Anonymous. Author index. Lecture Notes in Computer Science, 2458: 663–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2458/2458auth.pdf.

Jackson:2002:DSD

[1285] Michael Jackson. Descriptions in software development. Lecture Notes in Computer Science, 2460: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2460/24600001.htm;   http:
                                        //link.springer.de/link/service/

series/0558/papers/2460/24600001.
pdf.

Naumenko:2002:MUM

[1286] Andrey Naumenko and Alain Wegmann. A metamodel for the unified modeling language. Lecture Notes in Computer Science, 2460: 2–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2460/24600002.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2460/24600002.
pdf.

Varro:2002:MMP

[1287] Dániel Varró and András Pataricza.

Metamodeling mathematics: a precise and visual framework for describing semantics domains of UML models. Lecture Notes in Computer Science, 2460:18–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2460/24600018.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2460/24600018.pdf.

Steimann:2002:RRU

[1288] Friedrich Steimann and Thomas Kühne. A radical reduction of UML’s core semantics. Lecture Notes in Computer Science, 2460: 34–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2460/24600034.htm;   http:
//link.springer.de/link/service/

                                        series/0558/papers/2460/24600034.
                                        pdf.

Felfernig:2002:CKR

[1289] Alexander Felfernig, Gerhard Friedrich, Dietmar Jannach, and Markus Zanker. Configuration knowledge representation using UML/OCL. Lecture Notes in Computer Science, 2460: 49–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2460/24600049.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2460/24600049.
                                        pdf.

Frohlich:2002:UUI

[1290] Peter Fröhlich, Zaijun Hu, and Manfred Schoelzke. Using UML for information modeling in industrial systems with multiple hierarchies. Lecture Notes in Computer Science, 2460:63–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2460/24600063.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2460/24600063.pdf.

Dietzsch:2002:AUB

[1291] Andreas Dietzsch. Adapting the UML to business modelling’s needs — experiences in situational method engineering. Lecture Notes in Computer Science, 2460:73–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2460/24600073.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2460/24600073.pdf.

Gogolla:2002:AUS

[1292] Martin Gogolla and Brian Henderson-Sellers. Analysis of UML stereotypes within the UML metamodel. Lecture Notes in Computer Science, 2460:84–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2460/24600084.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2460/24600084.pdf.

Atkinson:2002:SET

[1293] Colin Atkinson, Thomas Kühne, and Brian Henderson-Sellers. Stereotypical encounters of the third kind. Lecture Notes in Computer Science, 2460:100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2460/24600100.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2460/24600100.pdf.

Genova:2002:DUC

[1294] Gonzalo Génova, Juan Llorens, and Víctor Quintana. Digging into use case relationships. Lecture Notes in Computer Science, 2460: 115–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2460/24600115.htm;   http:
//link.springer.de/link/service/

                                        series/0558/papers/2460/24600115.
                                        pdf.

deMiguel:2002:PEA

[1295] Miguel de Miguel, Jean Jourdan, and Serge Salicki. Practical experiences in the application of MDA. Lecture Notes in Computer Science, 2460:128–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2460/24600128.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2460/24600128.pdf.

Jorgensen:2002:EDM

[1296] Jens Bæk Jørgensen and Søren Christensen. Executable design models for a pervasive healthcare middleware system. Lecture Notes in Computer Science, 2460:140–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2460/24600140.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2460/24600140.pdf.

Sturm:2002:GCU

[1297] Thorsten Sturm, Jesco von Voss, and Marko Boger. Generating code from UML with velocity templates. Lecture Notes in Computer Science, 2460:150–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2460/24600150.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2460/24600150.pdf.

Selic:2002:DYS

[1298] Bran Selic. Does your software creak as it runs? Lecture Notes in Computer Science, 2460:162–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2460/24600162.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2460/24600162.pdf.

Andre:2002:ISP

[1299] Charles André, Marie-Agnès Peraldi-Frati, and Jean-Paul Rigault. Integrating the synchronous paradigm into UML: Application to controldominated systems. Lecture Notes in Computer Science, 2460:163– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2460/24600163.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2460/24600163.
pdf.

Flake:2002:UPR

[1300] Stephan Flake and Wolfgang Mueller.

A UML profile for real-time constraints with the OCL. Lecture Notes in Computer Science, 2460: 179–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2460/24600179.htm;   http:
//link.springer.de/link/service/

series/0558/papers/2460/24600179.
pdf.

Brucker:2002:HOE

[1301] Achim D. Brucker and Burkhart Wolff. HOL-OCL: Experiences, consequences and design choices. Lecture Notes in Computer Science, 2460:196–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2460/24600196.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2460/24600196.pdf.

Engels:2002:CPM

[1302] Gregor Engels, Reiko Heckel, Jochen M. Küster, and Luuk Groenewegen. Consistency-preserving model evolution through transformations. Lecture Notes in Computer Science, 2460: 212–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2460/24600212.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2460/24600212.
pdf.

Whittle:2002:TSM

[1303] Jon Whittle. Transformations and software modeling languages: Automating transformations in UML. Lecture Notes in Computer Science, 2460:227–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2460/24600227.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2460/24600227.pdf.

Akehurst:2002:RAD

[1304] David Akehurst and Stuart Kent.

A relational approach to defining transformations in a metamodel. Lecture Notes in Computer Science, 2460:243–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2460/24600243.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2460/24600243.pdf.

Pllana:2002:CUM

[1305] Sabri Pllana and Thomas Fahringer.

On customizing the UML for modeling performance-oriented applications. Lecture Notes in Computer Science, 2460:259–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2460/24600259.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2460/24600259.pdf.

Aagedal:2002:MQT

[1306] Jan Øyvind Aagedal and Earl F.

Ecklund, Jr. Modelling QoS: Towards a UML profile. Lecture Notes in Computer Science, 2460: 275–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2460/24600275.htm;   http:
//link.springer.de/link/service/

series/0558/papers/2460/24600275.
pdf.

Lujan-Mora:2002:EUM

[1307] Sergio Luján-Mora, Juan Trujillo, and Il-Yeol Song. Extending the UML for multidimensional modeling. Lecture Notes in Computer Science, 2460:290–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2460/24600290.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2460/24600290.pdf.

Clark:2002:MPE

[1308] Tony Clark, Andy Evans, and Stuart Kent. A metamodel for package extension with renaming. Lecture Notes in Computer Science, 2460:305–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2460/24600305.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2460/24600305.pdf.

Anido:2002:AMC

[1309] Luis Anido, Manuel Caeiro, Judith Rodríguez, and Juan Santos. Applying MDA concepts to develop a domain CORBA facility for Elearning. Lecture Notes in Computer Science, 2460:321–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2460/24600321.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2460/24600321.pdf.

Schattkowsky:2002:RDM

[1310] Tim Schattkowsky and Marc Lohmann.

Rapid development of modular dynamic Web sites using UML. Lecture Notes in Computer Science, 2460: 336–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2460/24600336.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2460/24600336.
pdf.

Garlan:2002:SHT

[1311] David Garlan. Software, heal thyself! Lecture Notes in Computer Science, 2460:351–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2460/24600351.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2460/24600351.pdf.

Cariou:2002:SUC

[1312] Eric Cariou and Antoine Beugnard.

The specification of UML collaborations as interaction components. Lecture Notes in Computer Science, 2460:352–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2460/24600352.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2460/24600352.pdf.

Tang:2002:MOD

[1313] Mei-Huei Tang and Mei-Hwa Chen. Measuring OO design metrics from UML. Lecture Notes in Computer Science, 2460:368–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2460/24600368.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2460/24600368.pdf.

Basanieri:2002:CAP

[1314] Francesca Basanieri, Antonia Bertolino, and Eda Marchetti. The Cow Suite approach to planning and deriving test suites in UML projects. Lecture Notes in Computer Science, 2460:383–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2460/24600383.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2460/24600383.pdf.

Boger:2002:DIU

[1315] Marko Boger, Mario Jeckle, Stefan Mueller, and Jens Fransson. Diagram interchange for UML. Lecture Notes in Computer Science, 2460: 398–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2460/24600398.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2460/24600398.
                                        pdf.

Jurjens:2002:UEU

[1316] Jan Jürjens. UMLsec: Extending UML for secure systems development. Lecture Notes in Computer Science, 2460:412–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2460/24600412.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2460/24600412.pdf.

Lodderstedt:2002:SUB

[1317] Torsten Lodderstedt, David Basin, and Jürgen Doser. SecureUML: a UML-based modeling language for model-driven security. Lecture Notes in Computer Science, 2460: 426–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2460/24600426.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2460/24600426.
pdf.

Whittle:2002:WTU

[1318] Jon Whittle. Workshops and tutorials at the UML 2002 Conference. Lecture Notes in Computer Science, 2460:442–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2460/24600442.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2460/24600442.pdf.

Anonymous:2002:AIbc

[1319] Anonymous. Author index. Lecture Notes in Computer Science, 2460: 449–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2460/2460ind.pdf.

Rabani:2002:SCH

[1320] Yuval Rabani. Search and classification of high dimensional data. Lecture Notes in Computer Science, 2462:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2462/24620001.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2462/24620001.pdf.

Ravi:2002:BST

[1321] R. Ravi. Bicriteria spanning tree problems. Lecture Notes in Computer Science, 2462:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2462/24620003.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2462/24620003.pdf.

Ageev:2002:IAA

[1322] Alexander Ageev. Improved approximation algorithms for multilevel facility location problems. Lecture Notes in Computer Science, 2462:5–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2462/24620005.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2462/24620005.pdf.

Ahuja:2002:CHC

[1323] Nitin Ahuja and Anand Srivastav. On constrained hypergraph coloring and scheduling. Lecture Notes in Computer Science, 2462:14–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2462/24620014.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2462/24620014.pdf.

Angelopoulos:2002:PPA

[1324] Spyros Angelopoulos and Allan Borodin.

On the power of priority algorithms for facility location and set cover. Lecture Notes in Computer Science, 2462:26–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2462/24620026.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2462/24620026.pdf.

Blaser:2002:TAA

[1325] Markus Bläser and Bodo Manthey. Two approximation algorithms for 3-cycle covers. Lecture Notes in Computer Science, 2462: 40–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2462/24620040.htm;   http:

                                        //link.springer.de/link/service/
                                        series/0558/papers/2462/24620040.
                                        pdf.

Chakrabarti:2002:AAU

[1326] Amit Chakrabarti, Chandra Chekuri, Anupam Gupta, and Amit Kumar. Approximation algorithms for the unsplittable flow problem. Lecture Notes in Computer Science, 2462:51–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2462/24620051.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2462/24620051.pdf.

Demaine:2002:ATG

[1327] Erik D. Demaine, MohammadTaghi Hajiaghayi, and Dimitrios M. Thilikos. 1.5-approximation for treewidth of graphs excluding a graph with one crossing as a minor. Lecture Notes in Computer Science, 2462: 67–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2462/24620067.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2462/24620067.
                                        pdf.

Doerr:2002:TRP

[1328] Benjamin Doerr. Typical rounding problems. Lecture Notes in Computer Science, 2462:81–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2462/24620081.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2462/24620081.pdf.

Feige:2002:AMS

[1329] Uriel Feige, László Lovász, and Prasad Tetali. Approximating min-sum set cover. Lecture Notes in Computer Science, 2462:94–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2462/24620094.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2462/24620094.pdf.

Feige:2002:AME

[1330] Uriel Feige, Eran Ofek, and Udi Wieder. Approximating maximum edge coloring in multigraphs. Lecture Notes in Computer Science, 2462:108–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2462/24620108.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2462/24620108.pdf.

Ganapathy:2002:ACM

[1331] Ganeshkumar Ganapathy and Tandy Warnow. Approximating the complement of the maximum compatible subset of leaves of k trees. Lecture Notes in Computer Science, 2462:122–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2462/24620122.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2462/24620122.pdf.

Giaro:2002:AAC

[1332] Krzysztof Giaro, Robert Janczewski, Marek Kubale, and Michał Małafiejski. A 27/26-approximation algorithm for the chromatic sum coloring of bipartite graphs. Lecture Notes in Computer Science, 2462:135–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2462/24620135.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2462/24620135.pdf.

Gudmundsson:2002:FLG

[1333] Joachim Gudmundsson, Herman Haverkort, Sang-Min Park, Chan-Su Shin, and Alexander Wolff. Facility location and the geometric minimum-diameter spanning tree. Lecture Notes in Computer Science, 2462:146–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2462/24620146.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2462/24620146.pdf.

Halperin:2002:IAA

[1334] Eran Halperin and Aravind Srinivasan. Improved approximation algorithms for the partial vertex cover problem. Lecture Notes in Computer Science, 2462:161–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2462/24620161.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2462/24620161.pdf.

Hassin:2002:MRD

[1335] Refael Hassin and Asaf Levin. Minimum restricted diameter spanning trees. Lecture Notes in Computer Science, 2462:175–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2462/24620175.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2462/24620175.pdf.

Kortsarz:2002:HAV

[1336] Guy Kortsarz, Robert Krauthgamer, and James R. Lee. Hardness of approximation for vertex-connectivity network-design problems. Lecture Notes in Computer Science, 2462: 185–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2462/24620185.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2462/24620185.
pdf.

Krumke:2002:NAH

[1337] Sven O. Krumke, Luigi Laura, Maarten Lipmann, Alberto Marchetti-Spaccamela, and Willem E. de Paepe. Non-abusiveness helps: An O(∞)- competitive algorithm for minimizing the maximum flow time in the online traveling salesman problem. Lecture Notes in Computer Science, 2462:200–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2462/24620200.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2462/24620200.pdf.

Lewin-Eytan:2002:RAC

[1338] Liane Lewin-Eytan, Joseph (Seffi) Naor, and Ariel Orda. Routing and admission control in networks with advance reservations. Lecture Notes in Computer Science, 2462:215–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2462/24620215.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2462/24620215.pdf.

Mahdian:2002:IAA

[1339] Mohammad Mahdian, Yinyu Ye, and Jiawei Zhang. Improved approximation algorithms for metric facility location problems. Lecture Notes in Computer Science, 2462:229–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2462/24620229.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2462/24620229.pdf.

Milidiu:2002:CMM

[1340] Ruy Luiz Milidiú, Artur Alves Pessoa, and Eduardo Sany Laber. Complexity of makespan minimization for pipeline transportation of petroleum products. Lecture Notes in Computer Science, 2462:243–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2462/24620243.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2462/24620243.pdf.

Swamy:2002:PDA

[1341] Chaitanya Swamy and Amit Kumar. Primal-dual algorithms for connected facility location problems. Lecture Notes in Computer Science, 2462:256–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2462/24620256.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2462/24620256.pdf.

Anonymous:2002:AIbd

[1342] Anonymous. Author index. Lecture Notes in Computer Science, 2462: 271–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2462/2462auth.pdf.

Vink:2002:PKI

[1343] Peter Vink and Frans M. van Eijnatten. Participation: The key to intelligent manufacturing improvement. Lecture Notes in Computer Science, 2465:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2465/24650001.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2465/24650001.pdf.

Yamada:2002:CDH

[1344] Shunji Yamada. Challenges in dealing with human factors issues in manufacturing activities. Lecture Notes in Computer Science, 2465:10–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2465/24650010.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2465/24650010.pdf.

Kamimura:2002:ERM

[1345] Manabu Kamimura, Naoyuki Kato, Ryo Kawabata, Satoshi Kumagai, and Kiyoshi Itoh. Extracting E-R models from collaboration analysis methods, MCM, and CLM. Lecture Notes in Computer Science, 2465:30–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2465/24650030.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2465/24650030.pdf.

Kurosu:2002:LMS

[1346] Noriaki Kurosu and Shunji Yamada. Living manufacturing systems with living organizations. Lecture Notes in Computer Science, 2465: 44–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2465/24650044.htm;   http:

//link.springer.de/link/service/
series/0558/papers/2465/24650044.
pdf.

Chikano:2002:SHC

[1347] Machiko Chikano, Yoshitaka Tomita, Yasuhiko Hiraide, and Eiji Arai. A study on human-centric realtime scheduling for PWB assembly line. Lecture Notes in Computer Science, 2465:57–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2465/24650057.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2465/24650057.pdf.

VanDeBovenkamp:2002:TIS

[1348] Martin Van De Bovenkamp, Ruben Jongkind, Gu Van Rhijn, Frans M. Van Eijnatten, Gudela Grote, Jouni Lehtelä, Timo Leskinen, Peter Vink, Scott Little, and Toni Wäfler. The E/S tool IT-support for ergonomic and sociotechnical system design. Lecture Notes in Computer Science, 2465:67–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2465/24650067.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2465/24650067.pdf.

Noro:2002:CVW

[1349] Kageyu Noro and Ryohei Tanaka. Construction of virtual working environment and evaluation of the workers. Lecture Notes in Computer Science, 2465:81–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2465/24650081.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2465/24650081.pdf.

Berkers:2002:HMD

[1350] Frank Berkers, Jan Goossenaerts, Dieter Hammer, and Hans Wortmann. Human models and data in the ubiquitous information infrastructure. Lecture Notes in Computer Science, 2465:91–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2465/24650091.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2465/24650091.pdf.

Imai:2002:MSH

[1351] Sayaka Imai, Takashi Tomii, and Hiroshi Arisawa. Motion simulation of the human workers for the integrated computer-aided manufacturing process simulation based on info-ergonomics concept. Lecture Notes in Computer Science, 2465:105–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2465/24650105.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2465/24650105.pdf.

Arisawa:2002:HBM

[1352] Hiroshi Arisawa, Takako Sato, and Takashi Tomii. Human-body motion simulation using bone-based human model and construction of motion database. Lecture Notes in Computer Science, 2465:115–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2465/24650115.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2465/24650115.pdf.

Goossenaerts:2002:OCP

[1353] Jan Goossenaerts and Christine Pelletier. Ontological commitment for participative simulation. Lecture Notes in Computer Science, 2465:127–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2465/24650127.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2465/24650127.pdf.

Shirase:2002:DMA

[1354] Keiichi Shirase, Hidefumi Wakamatsu, Akira Tsumaya, and Eiji Arai. Dynamic management architecture for human oriented production system. Lecture Notes in Computer Science, 2465:141–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2465/24650141.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2465/24650141.pdf.

Ram:2002:GSB

[1355] Sudha Ram, Vijay Khatri, Limin Zhang, and Daniel Dajun Zeng. Geo-Cosm: a semantics-based approach for information integration of geospatial data. Lecture Notes in Computer Science, 2465:152–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2465/24650152.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2465/24650152.pdf.

Frohlich:2002:IMR

[1356] Peter Fröhlich, Zaijun Hu, and Manfred Schoelzke. Imposing modeling rules on industrial applications through meta-modeling. Lecture Notes in Computer Science, 2465: 166–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2465/24650166.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2465/24650166.
pdf.

Kappel:2002:MUW

[1357] Gerti Kappel, B. Pröll, Werner Retschitzegger, and Wieland Schwinger. Modelling ubiquitous Web applications — the WUML approach. Lecture Notes in Computer Science, 2465:183–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2465/24650183.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2465/24650183.pdf.

Casteleyn:2002:SWS

[1358] Sven Casteleyn and Olga De Troyer.

Structuring Web sites using audience class hierarchies. Lecture Notes in Computer Science, 2465: 198–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2465/24650198.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2465/24650198.
pdf.

Liddle:2002:AED

[1359] Stephen W. Liddle, Sai Ho Yau, and David W. Embley. On the automatic extraction of data from the hidden Web. Lecture Notes in Computer Science, 2465:212–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2465/24650212.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2465/24650212.pdf.

Marcos:2002:MBM

[1360] E. Marcos, P. Cáceres, B. Vela, and J. M. Cavero. MIDAS/BD: a methodological framework for Web database design. Lecture Notes in Computer Science, 2465:227–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2465/24650227.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2465/24650227.pdf.

Lechner:2002:TXX

[1361] Stephan Lechner, Günter Preuner, and Michael Schrefl. Translating XQuery into XSLT. Lecture Notes in Computer Science, 2465:239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2465/24650239.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2465/24650239.pdf.

Atzeni:2002:WSE

[1362] Paolo Atzeni, Paolo Merialdo, and Giuseppe Sindoni. Web site evaluation: Methodology and case study. Lecture Notes in Computer Science, 2465:253–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2465/24650253.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2465/24650253.pdf.

Crescenzi:2002:AWI

[1363] Valter Crescenzi, Giansalvatore Mecca, and Paolo Merialdo. Automatic Web information extraction in the ROADRUNNER system. Lecture Notes in Computer Science, 2465: 264–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2465/24650264.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2465/24650264.
                                        pdf.

Lawrence:2002:QRD

[1364] Ramon Lawrence and Ken Barker.

Querying relational databases without explicit joins. Lecture Notes in Computer Science, 2465:278– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2465/24650278.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2465/24650278.
pdf.

Wu:2002:NSN

[1365] Xiaoying Wu, Tok Wang Ling, Sin Yeung Lee, Mong Li Lee, and Gillian Dobbie. NF-SS: a normal form for semistructured schema. Lecture Notes in Computer Science, 2465:292–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2465/24650292.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2465/24650292.pdf.

Wei:2002:FAL

[1366] Fang Wei and Georg Lausen. A formal analysis of the lightweight directory access protocol. Lecture Notes in Computer Science, 2465: 306–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2465/24650306.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2465/24650306.
pdf.

Chang:2002:XDR

[1367] Jae-Woo Chang. An XML document retrieval system supporting structure- and content-based queries. Lecture Notes in Computer Science, 2465:320–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2465/24650320.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2465/24650320.pdf.

Hatano:2002:EPX

[1368] Kenji Hatano, Hiroko Kinutani, Masatoshi Yoshikawa, and Shunsuke Uemura. Extraction of partial XML documents using IRbased structure and contents analysis. Lecture Notes in Computer Science, 2465:334–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2465/24650334.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2465/24650334.pdf.

Krishnan:2002:XWF

[1369] Rupa Krishnan, Lalitha Munaga, and Kamalakar Karlapalem. XDoC-WFMS: a framework for document centric workflow management system. Lecture Notes in Computer Science, 2465:348–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2465/24650348.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2465/24650348.pdf.

Schrefl:2002:AXS

[1370] Michael Schrefl and Martin Bernauer.

Active XML schemas. Lecture Notes in Computer Science, 2465: 363–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2465/24650363.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2465/24650363.
pdf.

Bussler:2002:BAS

[1371] Christoph Bussler. Behavior abstraction in semantic B2B integration. Lecture Notes in Computer Science, 2465:377–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2465/24650377.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2465/24650377.pdf.

Ermolayev:2002:OOC

[1372] Vadim Ermolayev, Natalya Keberle, and Vyachyslav Tolok. OIL ontologies for collaborative task performance in coalitions of self-interested actors. Lecture Notes in Computer Science, 2465:390–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2465/24650390.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2465/24650390.pdf.

Taveter:2002:MPM

[1373] Kuldar Taveter and Gerd Wagner. A multi-perspective methodology for modelling inter-enterprise business processes. Lecture Notes in Computer Science, 2465:403–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2465/24650403.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2465/24650403.pdf.

Jayaweera:2002:PPG

[1374] Prasad Jayaweera, Paul Johannesson, and Petia Wohed. Process patterns to generate E-commerce systems. Lecture Notes in Computer Science, 2465:417–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2465/24650417.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2465/24650417.pdf.

Janowski:2002:FFC

[1375] Tomasz Janowski and Adegboyega Ojo. Formalising feasibility and correctness of distributed business processes. Lecture Notes in Computer Science, 2465:432–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2465/24650432.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2465/24650432.pdf.

Frank:2002:MPV

[1376] Ulrich Frank. Modeling products for versatile E-commerce platforms — essential requirements and generic design alternatives. Lecture Notes in Computer Science, 2465: 444–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2465/24650444.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2465/24650444.
pdf.

Cappi:2002:SPC

[1377] Juan Cappi, Gustavo Rossi, Andrés Fortier, and Daniel Schwabe. Seamless personalization of E-commerce applications. Lecture Notes in Computer Science, 2465:457–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2465/24650457.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2465/24650457.pdf.

Ko:2002:DUP

[1378] SuJeong Ko and JungHyun Lee. Discovery of user preference through genetic algorithm and Bayesian categorization for recommendation. Lecture Notes in Computer Science, 2465: 471–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2465/24650471.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2465/24650471.
pdf.

Ellis:2002:UQD

[1379] Harry Ellis. Using the quantum data model to develop shareable definitions. Lecture Notes in Computer Science, 2465:485–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2465/24650485.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2465/24650485.pdf.

Anonymous:2002:AIbe

[1380] Anonymous. Author index. Lecture Notes in Computer Science, 2465: 499–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2465/2465auth.pdf.

Needham:2002:KAM

[1381] Roger Needham. Keynote address: Mobile computing versus immobile security. Lecture Notes in Computer Science, 2467:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2467/24670001.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2467/24670001.pdf.

Roe:2002:EMI

[1382] Michael Roe. Experiences of mobile IP security (transcript of discussion). Lecture Notes in Computer Science, 2467:4–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2467/24670004.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2467/24670004.pdf.

Nikander:2002:DSAa

[1383] Pekka Nikander. Denial-of-service, address ownership, and early authentication in the IPv6 world. Lecture Notes in Computer Science, 2467:12–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2467/24670012.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2467/24670012.pdf.

Nikander:2002:DSAb

[1384] Pekka Nikander. Denial of service, address ownership, and early authentication in the IPv6 world (transcript of discussion). Lecture Notes in Computer Science, 2467:22–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2467/24670022.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2467/24670022.pdf.

Aiello:2002:ERS

[1385] William Aiello, Steven M. Bellovin, Matt Blaze, Ran Canetti, John Ioannidis, Angelos D. Keromytis, and Omer Reingold. Efficient, DoS-resistant, secure key exchange for Internet protocols. Lecture Notes in Computer Science, 2467:27–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2467/24670027.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2467/24670027.pdf.

Blaze:2002:ERS

[1386] Matt Blaze. Efficient, DoS-resistant, secure key exchange for Internet protocols (transcript of discussion). Lecture Notes in Computer Science, 2467:40–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2467/24670040.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2467/24670040.pdf.

Price:2002:TTAa

[1387] Geraint Price. Thwarting timing attacks using ATM networks. Lecture Notes in Computer Science, 2467:49–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2467/24670049.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2467/24670049.pdf.

Price:2002:TTAb

[1388] Geraint Price. Thwarting timing attacks using ATM networks (transcript of discussion). Lecture Notes in Computer Science, 2467: 59–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/

0558/bibs/2467/24670059.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2467/24670059.
pdf.

Aura:2002:TSS

[1389] Tuomas Aura and Silja Mäki. Towards a survivable security architecture for ad-hoc networks. Lecture Notes in Computer Science, 2467: 63–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2467/24670063.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2467/24670063.
pdf.

Maki:2002:TSS

[1390] Silja Mäki. Towards a survivable security architecture for adhoc networks (transcript of discussion). Lecture Notes in Computer Science, 2467:74–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2467/24670074.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2467/24670074.pdf.

Gollmann:2002:PS

[1391] Dieter Gollmann. PIM security.

Lecture Notes in Computer Science, 2467:80–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2467/24670080.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2467/24670080.pdf.

Gollmann:2002:PST

[1392] Dieter Gollmann. PIM security (transcript of discussion). Lecture Notes in Computer Science, 2467: 82–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2467/24670082.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2467/24670082.
                                        pdf.

Christianson:2002:MPRa

[1393] Bruce Christianson and David Wheeler. Merkle puzzles revisited — finding matching elements between lists. Lecture Notes in Computer Science, 2467:87–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2467/24670087.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2467/24670087.pdf.

Christianson:2002:MPRb

[1394] Bruce Christianson. Merkle puzzles revisited (transcript of discussion). Lecture Notes in Computer Science, 2467:91–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2467/24670091.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2467/24670091.pdf.

Jurjens:2002:ERPa

[1395] Jan Jürjens. Encapsulating rules of prudent security engineering (position paper). Lecture Notes in Computer Science, 2467:95–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2467/24670095.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2467/24670095.pdf.

Jurjens:2002:ERPb

[1396] Jan Jürjens. Encapsulating rules of prudent security engineering (transcript of discussion). Lecture Notes in Computer Science, 2467: 102–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2467/24670102.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2467/24670102.
pdf.

Yoshiura:2002:MATa

[1397] Hiroshi Yoshiura, Kunihiko Miyazaki, Shinji Itoh, Kazuo Takaragi, and Ryoichi Sasaki. A multi-OS approach to trusted computer systems. Lecture Notes in Computer Science, 2467:107–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2467/24670107.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2467/24670107.pdf.

Yoshiura:2002:MATb

[1398] Hiroshi Yoshiura. A multi-OS approach to trusted computer systems (transcript of discussion). Lecture Notes in Computer Science, 2467: 115–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2467/24670115.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2467/24670115.
                                        pdf.

Bella:2002:PNR

[1399] Giampaolo Bella and Lawrence C. Paulson. A proof of non-repudiation. Lecture Notes in Computer Science, 2467:119–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2467/24670119.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2467/24670119.pdf.

Paulson:2002:PNR

[1400] Larry Paulson. A proof of nonrepudiation (transcript of discussion). Lecture Notes in Computer Science, 2467:126–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2467/24670126.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2467/24670126.pdf.

Firozabadi:2002:UACa

[1401] Babak Sadighi Firozabadi, Marek Sergot, and Olav Bandmann. Using authority certificates to create management structures. Lecture Notes in Computer Science, 2467: 134–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2467/24670134.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2467/24670134.
pdf.

Firozabadi:2002:UACb

[1402] Babak Sadighi Firozabadi. Using attribute certificates for creating management structures (transcript of discussion). Lecture Notes in Computer Science, 2467:146–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2467/24670146.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2467/24670146.pdf.

Foley:2002:TMWa

[1403] Simon N. Foley. Trust management and whether to delegate. Lecture Notes in Computer Science, 2467:151–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2467/24670151.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2467/24670151.pdf.

Foley:2002:TMWb

[1404] Simon N. Foley. Trust management and whether to delegate (transcript of discussion). Lecture Notes in Computer Science, 2467:158–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2467/24670158.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2467/24670158.pdf.

Lomas:2002:YCT

[1405] Mark Lomas. You can’t take it with you (transcript of discussion). Lecture Notes in Computer Science, 2467:166–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2467/24670166.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2467/24670166.pdf.

Wheeler:2002:PUKa

[1406] David Wheeler. Protocols using keys from faulty data. Lecture Notes in Computer Science, 2467: 170–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2467/24670170.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2467/24670170.
                                        pdf.

Wheeler:2002:PUKb

[1407] David Wheeler. Protocols using keys from faulty data (transcript of discussion). Lecture Notes in Computer Science, 2467:180–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2467/24670180.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2467/24670180.pdf.

Gligor:2002:NACa

[1408] Virgil D. Gligor, Himanshu Khurana, Radostina K. Koleva, Vijay G. Bharadwaj, and John S. Baras. On the negotiation of access control policies. Lecture Notes in Computer Science, 2467:188–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2467/24670188.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2467/24670188.pdf.

Gligor:2002:NACb

[1409] Virgil D. Gligor. Negotiation of access control policies (transcript of discussion). Lecture Notes in Computer Science, 2467:202–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2467/24670202.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2467/24670202.pdf.

Saidi:2002:ITG

[1410] Hassen Saïdi. Intrusion-tolerant group management in enclaves (transcript of discussion). Lecture Notes in Computer Science, 2467:213–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2467/24670213.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2467/24670213.pdf.

Malcolm:2002:LAM

[1411] James Malcolm. Lightweight authentication in a mobile network (transcript of discussion). Lecture Notes in Computer Science, 2467: 217–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2467/24670217.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2467/24670217.
                                        pdf.

Drabwell:2002:BSF

[1412] Peter Drabwell. Bluetooth security — fact or fiction? (transcript of discussion). Lecture Notes in Computer Science, 2467:221–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2467/24670221.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2467/24670221.pdf.

Christianson:2002:CDW

[1413] Bruce Christianson. Concluding discussion when does confidentiality harm security? Lecture Notes in Computer Science, 2467:229–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2467/24670229.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2467/24670229.pdf.

Thucydides:2002:LW

[1414] Thucydides. The last word. Lecture Notes in Computer Science, 2467: 239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2467/24670239.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2467/24670239.
pdf.

Anonymous:2002:AIbf

[1415] Anonymous. Author index. Lecture Notes in Computer Science, 2467: 241–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2467/2467auth.pdf.

Behrmann:2002:UIS

[1416] Gerd Behrmann, Johan Bengtsson, Alexandre David, Kim G. Larsen, Paul Pettersson, and Wang Yi. UPPAAL implementation secrets. Lecture Notes in Computer Science, 2469:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2469/24690003.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2469/24690003.pdf.

McDermid:2002:SHS

[1417] John McDermid. Software hazard and safety analysis. Lecture Notes in Computer Science, 2469: 23–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2469/24690023.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2469/24690023.
pdf.

Buttazzo:2002:RTO

[1418] Giorgio Buttazzo. Real-time operating systems: Problems and novel solutions. Lecture Notes in Computer Science, 2469:37–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2469/24690037.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2469/24690037.pdf.

Douglass:2002:RTU

[1419] Bruce Powel Douglass. Real-time UML. Lecture Notes in Computer Science, 2469:53–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2469/24690053.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2469/24690053.pdf;   http://link.
springer.de/link/service/series/
0558/papers/2469/2469053a.pdf;
http://link.springer.de/link/service/
series/0558/papers/2469/2469053b.
pdf.

Kozen:2002:ECI

[1420] Dexter Kozen and Matt Stillerman. Eager class initialization for Java. Lecture Notes in Computer Science, 2469:71–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2469/24690071.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2469/24690071.pdf.

Pnueli:2002:AFM

[1421] Amir Pnueli. Applications of formal methods in biology. Lecture Notes in Computer Science, 2469: 81–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2469/24690081.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2469/24690081.
pdf.

Rushby:2002:OFV

[1422] John Rushby. An overview of formal verification for the time-triggered architecture. Lecture Notes in Computer Science, 2469:83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2469/24690083.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2469/24690083.pdf.

Sifakis:2002:SMB

[1423] Joseph Sifakis. Scheduler modeling based on the controller synthesis paradigm. Lecture Notes in Computer Science, 2469:107–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2469/24690107.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2469/24690107.pdf.

Jhumka:2002:CBS

[1424] Arshad Jhumka, Martin Hiller, and Neeraj Suri. Component-based synthesis of dependable embedded software. Lecture Notes in Computer Science, 2469:111–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2469/24690111.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2469/24690111.pdf.

Lohr:2002:SST

[1425] Christophe Lohr and Jean-Pierre Courtiat. From the specification to the scheduling of time-dependent systems. Lecture Notes in Computer Science, 2469:129–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2469/24690129.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2469/24690129.pdf.

Maler:2002:CBC

[1426] Oded Maler, Bruce H. Krogh, and Moez Mahfoudh. On control with bounded computational resources. Lecture Notes in Computer Science, 2469:147–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2469/24690147.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2469/24690147.pdf.

Yamamoto:2002:DSP

[1427] Mitsuharu Yamamoto, Jean-Marie Cottin, and Masami Hagiya. Decidability of safety properties of timed multiset rewriting. Lecture Notes in Computer Science, 2469: 165–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2469/24690165.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2469/24690165.
pdf.

Vilas:2002:ETA

[1428] Ana Fernández Vilas, José J. Pazos Arias, and Rebeca P. Díaz Redondo. Extending timed automaton and real-time logic to many-valued reasoning. Lecture Notes in Computer Science, 2469:185–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2469/24690185.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2469/24690185.pdf.

Tripakis:2002:FDT

[1429] Stavros Tripakis. Fault diagnosis for timed automata. Lecture Notes in Computer Science, 2469: 205–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2469/24690205.htm;   http:
//link.springer.de/link/service/

                                        series/0558/papers/2469/24690205.
                                        pdf.

Niebert:2002:VTA

[1430] Peter Niebert, Moez Mahfoudh, Eugene Asarin, Marius Bozga, Oded Maler, and Navendu Jain. Verification of timed automata via satisfiability checking. Lecture Notes in Computer Science, 2469:225–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2469/24690225.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2469/24690225.pdf.

Franzle:2002:TIN

[1431] Martin Fränzle. Take it NP-Easy: Bounded model construction for duration calculus. Lecture Notes in Computer Science, 2469: 245–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2469/24690245.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2469/24690245.
                                        pdf.

Penczek:2002:TBM

[1432] Wojciech Penczek, Bożena Wożna, and Andrzej Zbrzezny. Towards bounded model checking for the universal fragment of TCTL. Lecture Notes in Computer Science, 2469: 265–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2469/24690265.htm;   http:

//link.springer.de/link/service/
series/0558/papers/2469/24690265.
pdf.

Palsberg:2002:TIC

[1433] Jens Palsberg and Di Ma. A typed interrupt calculus. Lecture Notes in Computer Science, 2469: 291–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2469/24690291.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2469/24690291.
pdf.

Bouajjani:2002:PVG

[1434] Ahmed Bouajjani and Agathe Merceron. Parametric verification of a group membership algorithm. Lecture Notes in Computer Science, 2469: 311–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2469/24690311.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2469/24690311.
pdf.

Khoumsi:2002:MTC

[1435] Ahmed Khoumsi. A method for testing the conformance of real time systems. Lecture Notes in Computer Science, 2469:331–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2469/24690331.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2469/24690331.pdf.

Jansen:2002:PEU

[1436] David N. Jansen, Holger Hermanns, and Joost-Pieter Katoen. A probabilistic extension of UML statecharts. Lecture Notes in Computer Science, 2469:355–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2469/24690355.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2469/24690355.pdf.

Damm:2002:EQR

[1437] Werner Damm and Bengt Jonsson. Eliminating queues from RT UML model representations. Lecture Notes in Computer Science, 2469:375–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2469/24690375.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2469/24690375.pdf.

Knapp:2002:MCT

[1438] Alexander Knapp, Stephan Merz, and Christopher Rauh. Model checking — timed UML state machines and collaborations. Lecture Notes in Computer Science, 2469:395–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2469/24690395.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2469/24690395.pdf.

Zhao:2002:POP

[1439] Jianhua Zhao, He Xu, Xuandong Li, Tao Zheng, and Guoliang Zheng. Partial order path technique for checking parallel timed automata. Lecture Notes in Computer Science, 2469:417–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2469/24690417.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2469/24690417.pdf.

Dierks:2002:CTA

[1440] Henning Dierks and Marc Lettrari. Constructing test automata from graphical real-time requirements. Lecture Notes in Computer Science, 2469:433–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2469/24690433.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2469/24690433.pdf.

Anonymous:2002:AIbg

[1441] Anonymous. Author index. Lecture Notes in Computer Science, 2469: 455–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2469/2469auth.pdf.

Alfonseca:2002:ELO

[1442] Enrique Alfonseca and Suresh Manandhar. Extending a lexical ontology by a combination of distributional semantics signatures. Lecture Notes in Computer Science, 2473:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2473/24730001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2473/24730001.pdf.

Bandini:2002:AKN

[1443] Stefania Bandini, Sara Manzoni, and Fabio Sartori. Acquiring knowledge and numerical data to support CBR retrieval. Lecture Notes in Computer Science, 2473:8–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2473/24730008.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2473/24730008.pdf.

Beydoun:2002:OMI

[1444] Ghassan Beydoun. An OO model for incremental hierarchical KA. Lecture Notes in Computer Science, 2473:14–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2473/24730014.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2473/24730014.pdf.

vanderGaag:2002:EMI

[1445] Linda C. van der Gaag and Eveline M. Helsper. Experiences with modelling issues in building probabilistic networks. Lecture Notes in Computer Science, 2473:21–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2473/24730021.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2473/24730021.pdf.

Gil:2002:IIK

[1446] Yolanda Gil and Varun Ratnakar.

IKRAFT: Interactive knowledge representation and acquisition from text. Lecture Notes in Computer Science, 2473:27–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2473/24730027.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2473/24730027.pdf.

Gil:2002:TIT

[1447] Yolanda Gil and Varun Ratnakar.

TRELLIS: An interactive tool for capturing information analysis and decision making. Lecture Notes in Computer Science, 2473:37–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2473/24730037.htm;   http://link.
springer.de/link/service/series/
0558/papers/2473/24730037.pdf.

Kim:2002:WBD

[1448] Mihye Kim and Paul Compton. Webbased document management for specialised domains: a preliminary evaluation. Lecture Notes in Computer Science, 2473:43–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2473/24730043.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2473/24730043.pdf.

Marcos:2002:IKF

[1449] Mar Marcos, Michael Balser, Annette ten Teije, and Frank van Harmelen. From informal knowledge to formal logic: a realistic case study in medical protocols. Lecture Notes in Computer Science, 2473:49–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2473/24730049.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2473/24730049.pdf.

Anjewierden:2002:KMM

[1450] Anjo Anjewierden, Irina Shostak, and Robert de Hoog. KMsim: a meta-modelling approach and environment for creating process-oriented knowledge management simulations. Lecture Notes in Computer Science, 2473:65–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2473/24730065.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2473/24730065.pdf.

Benjamins:2002:SMK

[1451] V. Richard Benjamins, José Manuel López Cobo, Jesús Contreras, Joaquín Casillas, Juan Blasco, Blanca de Otto, Juli García, Mercedes Blázquez, and Juan Manuel Dodero. Skills management in knowledge-intensive organizations. Lecture Notes in Computer Science, 2473:80–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2473/24730080.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2473/24730080.pdf.

Falkman:2002:KAM

[1452] Göran Falkman and Olof Torgersson.

Knowledge acquisition and modeling in clinical information systems: a case study. Lecture Notes in Computer Science, 2473:96–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2473/24730096.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2473/24730096.pdf.

Mate:2002:MIK

[1453] José Luis Maté, Luis Felipe Paradela, Juan Pazos, Alfonso Rodríguez-Patón, and Andrés Silva. MEGICO: An intelligent knowledge management methodology. Lecture Notes in Computer Science, 2473:102–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2473/24730102.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2473/24730102.pdf.

Aitken:2002:PO

[1454] Stuart Aitken and Jon Curtis. A process ontology. Lecture Notes in Computer Science, 2473:108– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2473/24730108.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2473/24730108.
                                        pdf.

Bachimont:2002:SCD

[1455] Bruno Bachimont, Antoine Isaac, and Raphaël Troncy. Semantic commitment for designing ontologies: a proposal. Lecture Notes in Computer Science, 2473:114–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2473/24730114.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2473/24730114.pdf.

Ciravegna:2002:USC

[1456] Fabio Ciravegna, Alexiei Dingli, Daniela Petrelli, and Yorick Wilks. User-system cooperation in document annotation based on information extraction. Lecture Notes in Computer Science, 2473:122–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2473/24730122.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2473/24730122.pdf.

Corcho:2002:WIW

[1457] Óscar Corcho, Mariano Fernández-López, Asunción Gómez-Pérez, and Óscar Vicente. WebODE: An integrated workbench for ontology representation, reasoning, and exchange. Lecture Notes in Computer Science, 2473:138–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2473/24730138.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2473/24730138.pdf.

Flett:2002:SOE

[1458] Alan Flett, Mariana Casella dos Santos, and Werner Ceusters. Some ontology engineering processes and their supporting technologies. Lecture Notes in Computer Science, 2473: 154–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2473/24730154.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2473/24730154.
pdf.

Gangemi:2002:SOD

[1459] Aldo Gangemi, Nicola Guarino, Claudio Masolo, Alessandro Oltramari, and Luc Schneider. Sweetening ontologies with DOLCE. Lecture Notes in Computer Science, 2473: 166–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2473/24730166.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2473/24730166.
pdf.

Hahn:2002:TLG

[1460] Udo Hahn and Stefan Schulz. Turning lead into gold? feeding a formal knowledge base with informal conceptual knowledge. Lecture Notes in Computer Science, 2473:182–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2473/24730182.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2473/24730182.pdf.

Klein:2002:OVC

[1461] Michel Klein, Dieter Fensel, Atanas Kiryakov, and Damyan Ognyanov. Ontology versioning and change detection on the Web. Lecture Notes in Computer Science, 2473:197–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2473/24730197.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2473/24730197.pdf.

Kozaki:2002:HEB

[1462] Kouji Kozaki, Yoshinobu Kitamura, Mitsuru Ikeda, and Riichiro Mizoguchi. Hozo: An environment for building/ using ontologies based on a fundamental consideration of “Role” and “Relationship”. Lecture Notes in Computer Science, 2473:213–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2473/24730213.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2473/24730213.pdf.

Lei:2002:ODA

[1463] Yuangui Lei, Enrico Motta, and John Domingue. An ontologydriven approach to Web site generation and maintenance. Lecture Notes in Computer Science, 2473: 219–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2473/24730219.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2473/24730219.
pdf.

Maedche:2002:MMF

[1464] Alexander Maedche, Boris Motik, Nuno Silva, and Raphael Volz. MAFRA – A MApping FRAmework for distributed ontologies. Lecture Notes in Computer Science, 2473: 235–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2473/24730235.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2473/24730235.
pdf.

Maedche:2002:MSB

[1465] Alexander Maedche and Steffen Staab.

Measuring similarity between ontologies. Lecture Notes in Computer Science, 2473:251–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2473/24730251.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2473/24730251.pdf.

Omelayenko:2002:OMB

[1466] Borys Omelayenko. Ontologymediated business integration. Lecture Notes in Computer Science, 2473: 264–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2473/24730264.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2473/24730264.
                                        pdf.

Reynaud:2002:ROI

[1467] Chantal Reynaud and Brigitte Safar. Representation of ontologies for information integration. Lecture Notes in Computer Science, 2473:270–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2473/24730270.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2473/24730270.pdf.

Stojanovic:2002:UDO

[1468] Ljiljana Stojanovic, Alexander Maedche, Boris Motik, and Nenad Stojanovic. User-driven ontology evolution management. Lecture Notes in Computer Science, 2473:285– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2473/24730285.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2473/24730285.
                                        pdf.

Tamma:2002:AMP

[1469] Valentina Tamma and Trevor J.

M. Bench Capon. Attribute metaproperties for formal ontological analysis. Lecture Notes in Computer Science, 2473:301–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2473/24730301.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2473/24730301.pdf.

Alani:2002:MRE

[1470] Harith Alani, Srinandan Dasmahapatra, Nicholas Gibbins, Hugh Glaser, Steve Harris, Yannis Kalfoglou, Kieron O’Hara, and Nigel Shadbolt. Managing reference: Ensuring referential integrity of ontologies for the Semantic Web. Lecture Notes in Computer Science, 2473:317–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2473/24730317.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2473/24730317.pdf.

Domingue:2002:AAO

[1471] John Domingue, Maria Martins, Jaicheng Tan, Arthur Stutt, and Helgi Pertusson. Alice: Assisting online shoppers through ontologies and novel interface metaphors. Lecture Notes in Computer Science, 2473:335–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2473/24730335.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2473/24730335.pdf.

Felfernig:2002:ACK

[1472] Alexander Felfernig, Gerhard Friedrich, Dietmar Jannach, Markus Stumptner, and Markus Zanker. Acquiring configuration knowledge bases in the Semantic Web using UML. Lecture Notes in Computer Science, 2473:352–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2473/24730352.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2473/24730352.pdf.

Handschuh:2002:CSA

[1473] Siegfried Handschuh, Steffen Staab, and Fabio Ciravegna. S-CREAM – semi-automatic CREAtion of metadata. Lecture Notes in Computer Science, 2473:358–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2473/24730358.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2473/24730358.pdf.

Ognyanov:2002:TCR

[1474] Damyan Ognyanov and Atanas Kiryakov. Tracking changes in RDF(S) repositories. Lecture Notes in Computer Science, 2473:373–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2473/24730373.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2473/24730373.pdf.

Vargas-Vera:2002:MOD

[1475] Maria Vargas-Vera, Enrico Motta, John Domingue, Mattia Lanzoni, Arthur Stutt, and Fabio Ciravegna. MnM: Ontology driven semi-automatic and automatic support for semantic markup. Lecture Notes in Computer Science, 2473:379–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2473/24730379.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2473/24730379.pdf.

Golbeck:2002:NTS

[1476] Jennifer Golbeck, Michael Grove, Bijan Parsia, Adtiya Kalyanpur, and James Hendler. New tools for the Semantic Web. Lecture Notes in Computer Science, 2473:392–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2473/24730392.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2473/24730392.pdf.

Anonymous:2002:AIbh

[1477] Anonymous. Author index. Lecture Notes in Computer Science, 2473: 401–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2473/2473indx.pdf.

Dongarra:2002:HPC

[1478] Jack Dongarra. High performance computing, computational Grid, and numerical libraries. Lecture Notes in Computer Science, 2474:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2474/24740001.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2474/24740001.pdf.

Sunderam:2002:PSR

[1479] Vaidy Sunderam. Performance, scalability, and robustness in the harness metacomputing framework. Lecture Notes in Computer Science, 2474:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2474/24740003.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2474/24740003.pdf.

vanAlbada:2002:SGD

[1480] Dick van Albada and Peter Sloot. Surfing the Grid — dynamic task migration in the Polder Metacomputer Project. Lecture Notes in Computer Science, 2474:4–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2474/24740004.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2474/24740004.pdf.

Geist:2002:PVM

[1481] Al Geist. Petascale virtual machine: Computing on 100,000 processors. Lecture Notes in Computer Science, 2474:6–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740006.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740006.pdf.

Gropp:2002:MNS

[1482] William Gropp. MPICH2: a new start for MPI implementations. Lecture Notes in Computer Science, 2474:7–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740007.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740007.pdf.

Juhasz:2002:MGC

[1483] Zoltan Juhasz. Making Grid computing mainstream. Lecture Notes in Computer Science, 2474:8–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2474/24740008.htm;   http://link.
springer.de/link/service/series/
0558/papers/2474/24740008.pdf.

Lusk:2002:PMS

[1484] Ewing Lusk. Process management for scalable parallel programs. Lecture Notes in Computer Science, 2474:9–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740009.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740009.pdf.

Miller:2002:SAD

[1485] Barton P. Miller. A security attack and defense in the Grid environment. Lecture Notes in Computer Science, 2474:10–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740010.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740010.pdf.

Gerndt:2002:PAN

[1486] Michael Gerndt. Performance analysis: Necessity or add-on in Grid computing. Lecture Notes in Computer Science, 2474:11–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740011.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740011.pdf.

Gropp:2002:MG

[1487] William Gropp and Ewing Lusk. MPI on the Grid. Lecture Notes in Computer Science, 2474:12–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2474/24740012.htm;   http://link.

springer.de/link/service/series/
0558/papers/2474/24740012.pdf.

Chapman:2002:PAD

[1488] Barbara Chapman. Parallel application development with the hybrid MPI + OpenMP programming model. Lecture Notes in Computer Science, 2474:13–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740013.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740013.pdf.

Bubak:2002:CRE

[1489] Marian Bubak and Michał Turała.

CrossGrid and its relatives in Europe. Lecture Notes in Computer Science, 2474:14–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740014.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740014.pdf.

Bubak:2002:TCA

[1490] Marian Bubak, Maciej Malawski, and Katarzyna Zajac. Towards the CrossGrid architecture. Lecture Notes in Computer Science, 2474: 16–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2474/24740016.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2474/24740016.
pdf.

Dutka:2002:ACE

[1491] Łukasz Dutka and Jacek Kitowski. Application of component-expert technology for selection of data-handlers in CrossGrid. Lecture Notes in Computer Science, 2474:25–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2474/24740025.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2474/24740025.pdf.

Ponce:2002:TNN

[1492] O. Ponce, J. Cuevas, A. Fuentes, J. Marco, R. Marco, C. Martínez-Rivero, R. Menéndez, and D. Rodríguez. Training of neural networks: Interactive possibilities in a distributed framework. Lecture Notes in Computer Science, 2474:33–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2474/24740033.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2474/24740033.pdf.

Balis:2002:IGA

[1493] Bartosz Balis, Marian Bubak, Włodzimierz Funika, Tomasz Szepieniec, and Roland Wismüller. An infrastructure for Grid application monitoring. Lecture Notes in Computer Science, 2474:41–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2474/24740041.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2474/24740041.pdf.

Bubak:2002:CPA

[1494] Marian Bubak, Włodzimierz Funika, and Roland Wismüller. The CrossGrid performance analysis tool for interactive Grid applications. Lecture Notes in Computer Science, 2474:50–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740050.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740050.pdf.

Trinitis:2002:CTN

[1495] Carsten Trinitis and Martin Schulz.

Current trends in numerical simulation for parallel engineering environments. Lecture Notes in Computer Science, 2474:61–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740061.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740061.pdf.

Hofinger:2002:ARL

[1496] Siegfried Höfinger. Automatic runtime load balancing of dedicated applications in heterogeneous environments. Lecture Notes in Computer Science, 2474:62–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740062.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2474/24740062.pdf.

Blaszczyk:2002:CIG

[1497] Andreas Blaszczyk and Axel Uhl. A contribution to industrial Grid computing. Lecture Notes in Computer Science, 2474:70–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2474/24740070.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2474/24740070.pdf.

Causin:2002:PCS

[1498] Paola Causin and Edie Miglio. Parallel computing for the simulation of 3D free surface flows in environmental applications. Lecture Notes in Computer Science, 2474:78–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2474/24740078.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2474/24740078.pdf.

Rotaru:2002:TAN

[1499] Tiberiu Rotaru and Hans-Heinrich Nägeli. Testbed for adaptive numerical simulations in heterogeneous environments. Lecture Notes in Computer Science, 2474:88–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2474/24740088.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2474/24740088.pdf.

Freitas:2002:SCF

[1500] Conceição Freitas, Luís Dias, and Miguel Dias. Simulating cloth freeform deformation with a Beowulf cluster. Lecture Notes in Computer Science, 2474:96–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740096.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740096.pdf.

Hluchy:2002:CPS

[1501] Ladislav Hluchy, Viet Dinh Tran, Ondrej Habala, Jan Astalos, Branislav Simo, and David Froehlich. Concept of a problem solving environment for flood forecasting. Lecture Notes in Computer Science, 2474: 105–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2474/24740105.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2474/24740105.
pdf.

Trinitis:2002:CEF

[1502] Carsten Trinitis, Martin Schulz, and Wolfgang Karl. A comprehensive electric field simulation environment on top of SCI. Lecture Notes in Computer Science, 2474:114–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740114.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740114.pdf.

Kasprzyk:2002:APV

[1503] Leszek Kasprzyk, Ryszard Nawrowski, and Andrzej Tomczewski. Application of a parallel virtual machine for the analysis of a luminous field. Lecture Notes in Computer Science, 2474:122–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2474/24740122.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2474/24740122.pdf.

Macias:2002:SEA

[1504] Elsa M. Macías and Alvaro Suárez. Solving engineering applications with LAMGAC over MPI-2. Lecture Notes in Computer Science, 2474: 130–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2474/24740130.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2474/24740130.
                                        pdf.

Kabir:2002:DIS

[1505] Yacine Kabir and A. Belhadj-Aissa. Distributed image segmentation system by a multi-agents approach (under PVM environment). Lecture Notes in Computer Science, 2474: 138–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2474/24740138.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2474/24740138.
                                        pdf.

Hofinger:2002:PGO

[1506] Siegfried Höfinger, Torsten Schindler, and András Aszódi. Parallel global optimization of high-dimensional problems. Lecture Notes in Computer Science, 2474:148–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740148.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740148.pdf.

Gine:2002:ALT

[1507] Francesc Giné, Francesc Solsona, Porfidio Hernández, and Emilio Luque. Adjusting the lengths of time slices when scheduling PVM jobs with high memory requirements. Lecture Notes in Computer Science, 2474:156–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740156.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740156.pdf.

Lazzarino:2002:PBP

[1508] Oscar Lazzarino, Andrea Sanna, Claudio Zunino, and Fabrizio Lamberti. A PVM-based parallel implementation of the REYES image rendering architecture. Lecture Notes in Computer Science, 2474:165–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740165.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2474/24740165.pdf.

Piriyakumar:2002:EFI

[1509] Douglas Antony Louis Piriyakumar, Paul Levi, and Rolf Rabenseifner. Enhanced file interoperability with parallel MPI file-I/O in image processing. Lecture Notes in Computer Science, 2474:174–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2474/24740174.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2474/24740174.pdf.

Tischler:2002:GLP

[1510] Florian Tischler and Andreas Uhl. Granularity levels in parallel blockmatching motion compensation. Lecture Notes in Computer Science, 2474: 183–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2474/24740183.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2474/24740183.
                                        pdf.

Song:2002:AMS

[1511] Ha Yoon Song, Junghwan Kim, and Kyun Rak Chong. An analytical model of scheduling for conservative parallel simulation. Lecture Notes in Computer Science, 2474:191–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2474/24740191.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2474/24740191.pdf.

Bekas:2002:PCP

[1512] Constantine Bekas, Efrosini Kokiopoulou, Efstratios Gallopoulos, and Valeria Simoncini. Parallel computation of pseudospectra using transfer functions on a MATLAB-MPI cluster platform. Lecture Notes in Computer Science, 2474:199–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740199.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740199.pdf.

Czarnul:2002:DTI

[1513] Pawel Czarnul. Development and tuning of irregular divide-and-conquer applications in DAMPVM/DAC. Lecture Notes in Computer Science, 2474: 208–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2474/24740208.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2474/24740208.
pdf.

Pagourtzis:2002:OPC

[1514] Aris Pagourtzis, Igor Potapov, and Wojciech Rytter. Observations on parallel computation of transitive and max-closure problems. Lecture Notes in Computer Science, 2474:217–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2474/24740217.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2474/24740217.pdf.

Serio:2002:ENN

[1515] Angela Di Serio and María B. Ibáñez. Evaluation of a nearest-neighbor load balancing strategy for parallel molecular simulations in MPI environment. Lecture Notes in Computer Science, 2474:226–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2474/24740226.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2474/24740226.pdf.

Nguyen:2002:ARP

[1516] Giang Thu Nguyen, Viet Dinh Tran, and Margareta Kotocova. Application recovery in parallel programming environment. Lecture Notes in Computer Science, 2474:234–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2474/24740234.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2474/24740234.pdf.

Malouin:2002:OPR

[1517] Éric Malouin, Julien Dompierre, François Guibault, and Robert Roy. P-OORT : a parallel remeshing toolkit. Lecture Notes in Computer Science, 2474:243–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2474/24740243.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740243.pdf.

Cotronis:2002:MMP

[1518] Yiannis Cotronis and Zacharias Tsiatsoulis. Modular MPI and PVM components. Lecture Notes in Computer Science, 2474:252–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740252.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740252.pdf.

Bernholdt:2002:CIH

[1519] David E. Bernholdt, Wael R. Elwasif, and James A. Kohl. Communication infrastructure in high-performance component-based scientific computing. Lecture Notes in Computer Science, 2474:260–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740260.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740260.pdf.

Worsch:2002:BCM

[1520] Thomas Worsch, Ralf Reussner, and Werner Augustin. On benchmarking collective MPI operations. Lecture Notes in Computer Science, 2474:271–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740271.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2474/24740271.pdf.

Gropp:2002:BLC

[1521] William Gropp. Building library components that can use any MPI implementation. Lecture Notes in Computer Science, 2474:280–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2474/24740280.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2474/24740280.pdf.

Tsujita:2002:SFP

[1522] Yuichi Tsujita, Toshiyuki Imamura, Hiroshi Takemiya, and Nobuhiro Yamagishi. Stampi-I/O: a flexible parallel-I/O library for heterogeneous computing environment. Lecture Notes in Computer Science, 2474:288–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2474/24740288.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2474/24740288.pdf.

Plachetka:2002:QTS

[1523] Tomas Plachetka. (quasi-) threadsafe PVM and (quasi-) threadsafe MPI without active polling. Lecture Notes in Computer Science, 2474:296–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2474/24740296.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740296.pdf.

Calderon:2002:IMI

[1524] Alejandro Calderón, Félix García, Jesús Carretero, Jose M. Pérez, and Javier Fernández. An implementation of MPI-IO on expand: a parallel file system based on NFS servers. Lecture Notes in Computer Science, 2474:306–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740306.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740306.pdf.

Huang:2002:DDD

[1525] Wei Huang, Zhe Wang, and Jie Ma. Design of DMPI on DAWNING-3000. Lecture Notes in Computer Science, 2474:314–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740314.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740314.pdf.

Selikhov:2002:MCC

[1526] Anton Selikhov, George Bosilca, Cecile Germain, Gilles Fedak, and Franck Cappello. MPICH-CM: a communication library design for a P2P MPI implementation. Lecture Notes in Computer Science, 2474: 323–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2474/24740323.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2474/24740323.
                                        pdf.

Brightwell:2002:DIM

[1527] Ron Brightwell, Arthur B. Maccabe, and Rolf Riesen. Design and implementation of MPI on portals 3.0. Lecture Notes in Computer Science, 2474:331–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2474/24740331.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2474/24740331.pdf.

Espenica:2002:PPA

[1528] Roberto Espenica and Pedro Medeiros. Porting PVM to the VIA architecture using a fast communication library. Lecture Notes in Computer Science, 2474:341–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2474/24740341.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2474/24740341.pdf.

Coli:2002:LMP

[1529] Moreno Coli, Paolo Palazzari, and Rodolfo Rughi. LICO: a multiplatform channel-based communication library. Lecture Notes in Computer Science, 2474:349–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2474/24740349.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740349.pdf.

Kranzlmuller:2002:NNM

[1530] Dieter Kranzlmüller and Martin Schulz. Notes on nondeterminism in message passing programs. Lecture Notes in Computer Science, 2474:357–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740357.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740357.pdf.

Jorba:2002:WRS

[1531] Josep Jorba, Rafael Bustos, Ángel Casquero, Tomàs Margalef, and Emilio Luque. Web remote services oriented architecture for cluster management. Lecture Notes in Computer Science, 2474:368–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740368.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740368.pdf.

Krysztop:2002:IFP

[1532] Bartosz Krysztop and Henryk Krawczyk.

Improving flexibility and performance of PVM applications by distributed partial evaluation. Lecture Notes in Computer Science, 2474: 376–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2474/24740376.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2474/24740376.
                                        pdf.

Brightwell:2002:RMR

[1533] Ron Brightwell. Ready-mode receive: An optimized receive function for MPI. Lecture Notes in Computer Science, 2474:385–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2474/24740385.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2474/24740385.pdf.

Traff:2002:IMA

[1534] Jesper Larsson Träff. Improved MPI all-to-all communication on a Giganet SMP cluster. Lecture Notes in Computer Science, 2474: 392–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2474/24740392.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2474/24740392.
                                        pdf.

Bisseling:2002:FMF

[1535] Georg Bißeling, Hans-Christian Hoppe, Alexander Supalov, Pierre Lagier, and Jean Latour. Fujitsu MPI-2: Fast locally, reaching globally. Lecture Notes in Computer Science, 2474:401–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2474/24740401.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740401.pdf.

Rabenseifner:2002:COA

[1536] Rolf Rabenseifner. Communication and optimization aspects on hybrid architectures. Lecture Notes in Computer Science, 2474:410–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740410.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740410.pdf.

Truong:2002:PAM

[1537] Hong-Linh Truong, Thomas Fahringer, Michael Geissler, and Georg Madsen. Performance analysis for MPI applications with SCALEA. Lecture Notes in Computer Science, 2474:421–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2474/24740421.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2474/24740421.pdf.

Michailidis:2002:PSL

[1538] Panagiotis D. Michailidis and Konstantinos G. Margaritis. A performance study of load balancing strategies for approximate string matching on an MPI heterogeneous system environment. Lecture Notes in Computer Science, 2474:432–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2474/24740432.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2474/24740432.pdf.

Almeida:2002:AMP

[1539] F. Almeida, D. González, L. M. Moreno, and C. Rodríguez. An analytical model for pipeline algorithms on heterogeneous clusters. Lecture Notes in Computer Science, 2474:441–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2474/24740441.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2474/24740441.pdf.

Furtado:2002:AEA

[1540] A. Furtado, A. Rebouças, J. R. de Souza, D. Rexachs, and E. Luque. Architectures for an efficient application execution in a collection of HNOWS. Lecture Notes in Computer Science, 2474:450–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2474/24740450.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2474/24740450.pdf.

Anonymous:2002:AIbi

[1541] Anonymous. Author index. Lecture Notes in Computer Science, 2474: 461–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2474/2474ind.pdf.

Pawlak:2002:PPD

[1542] Zdzisław Pawlak. In pursuit of patterns in data reasoning from data — the rough set way. Lecture Notes in Computer Science, 2475: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2475/24750001.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2475/24750001.
pdf.

Zadeh:2002:TTH

[1543] Lotfi A. Zadeh. Toward a theory of hierarchical definability (THD). Lecture Notes in Computer Science, 2475:10–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750010.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750010.pdf.

Komorowski:2002:MBP

[1544] Jan Komorowski. Modelling biological phenomena with rough sets. Lecture Notes in Computer Science, 2475:13–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750013.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750013.pdf.

Lin:2002:DMD

[1545] Tsau Young (”T. Y.”) Lin. Database mining on derived attributes. Lecture Notes in Computer Science, 2475: 14–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2475/24750014.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2475/24750014.
                                        pdf.

Russell:2002:PES

[1546] David W. Russell. A proposed evolutionary, self-organizing automaton for the control of dynamic systems. Lecture Notes in Computer Science, 2475:33–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750033.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750033.pdf.

Slowinski:2002:RSA

[1547] Roman Słowiński, Salvatore Greco, and Benedetto Matarazzo. Rough set analysis of preference-ordered data. Lecture Notes in Computer Science, 2475:44–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750044.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750044.pdf.

Turksen:2002:FSM

[1548] I. B. Türksen. Fuzzy sets, multivalued mappings, and rough sets. Lecture Notes in Computer Science, 2475:60–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750060.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750060.pdf.

Beynon:2002:ICV

[1549] Malcolm J. Beynon. Investigating the choice of l and u values in the extended variable precision rough sets model. Lecture Notes in Computer Science, 2475:61–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750061.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750061.pdf.

Cattaneo:2002:QAP

[1550] Gianpiero Cattaneo and Davide Ciucci.

A quantitative analysis of preclusivity vs. similarity based rough approximations. Lecture Notes in Computer Science, 2475:69–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750069.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750069.pdf.

Cattaneo:2002:HWA

[1551] Gianpiero Cattaneo and Davide Ciucci. Heyting wajsberg algebras as an abstract environment linking fuzzy and rough sets. Lecture Notes in Computer Science, 2475:77–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750077.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750077.pdf.

Greco:2002:DBR

[1552] Salvatore Greco, Masahiro Inuiguchi, and Roman Słowiński. Dominancebased rough set approach using possibility and necessity measures. Lecture Notes in Computer Science, 2475:85–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750085.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750085.pdf.

Greco:2002:GDA

[1553] Salvatore Greco, Zdzisław Pawlak, and Roman Słowiński. Generalized decision algorithms, rough inference rules, and flow graphs. Lecture Notes in Computer Science, 2475:93–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750093.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750093.pdf.

Inuiguchi:2002:GRS

[1554] Masahiro Inuiguchi and Tetsuzo Tanino. Generalized rough sets and rule extraction. Lecture Notes in Computer Science, 2475:105– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2475/24750105.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2475/24750105.
pdf.

Janicki:2002:TMS

[1555] Ryszard Janicki. Towards a mereological system for direct products and relations. Lecture Notes in Computer Science, 2475:113–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750113.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750113.pdf.

Jarvinen:2002:SRA

[1556] Jouni Järvinen. On the structure of rough approximations. Lecture Notes in Computer Science, 2475: 123–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2475/24750123.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2475/24750123.
pdf.

Katarzyniak:2002:MWC

[1557] Radoslaw Katarzyniak and Ngoc Thanh Nguyen. Modification of weights of conflict profile’s elements and dependencies of attributes in consensus model. Lecture Notes in Computer Science, 2475:131–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750131.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750131.pdf.

Liu:2002:RAI

[1558] Qing Liu and S. L. Jiang. Reasoning about information granules based on rough logic. Lecture Notes in Computer Science, 2475:139–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750139.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750139.pdf.

Midelfart:2002:RSF

[1559] Herman Midelfart and Jan Komorowski. A rough set framework for learning in a directed acyclic graph. Lecture Notes in Computer Science, 2475:144–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750144.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750144.pdf.

Moshkov:2002:CIS

[1560] Mikhail Moshkov. On compressible information systems. Lecture Notes in Computer Science, 2475:

156–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2475/24750156.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2475/24750156.
pdf.

Nakata:2002:FDR

[1561] Michinori Nakata and Tetsuya Murai. Functional dependencies in relational expressions based on orsets. Lecture Notes in Computer Science, 2475:161–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750161.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750161.pdf.

Polkowski:2002:APR

[1562] Lech Polkowski. On asymptotic properties of rough-set-theoretic approximations. fractal dimension, exact sets, and rough inclusion in potentially infinite information systems. Lecture Notes in Computer Science, 2475:167–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750167.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750167.pdf.

Pomykala:2002:ATS

[1563] J. A. Pomykała. About tolerance and similarity relations in information systems. Lecture Notes in Computer Science, 2475:175–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750175.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750175.pdf.

Ramsey:2002:RSG

[1564] Frederick V. Ramsey and James J. Alpigini. Rough sets, guarded command language, and decision rules. Lecture Notes in Computer Science, 2475:183–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750183.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750183.pdf.

Ras:2002:CQP

[1565] Zbigniew W. Ras and Agnieszka Dardzińska. Collaborative query processing in DKS controlled by reducts. Lecture Notes in Computer Science, 2475:189–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750189.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750189.pdf.

Rzasa:2002:NMD

[1566] Wojciech Rzasa and Zbigniew Suraj. A new method for determining of extensions and restrictions of information systems. Lecture Notes in Computer Science, 2475:197–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750197.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750197.pdf.

Vitoria:2002:LPF

[1567] Aida Vitória and Jan Małuszyński.

A logic programming framework for rough sets. Lecture Notes in Computer Science, 2475:205–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750205.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750205.pdf.

Wang:2002:ACD

[1568] G. Y. Wang. Attribute core of decision table. Lecture Notes in Computer Science, 2475:213–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750213.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750213.pdf.

Borkowski:2002:SAU

[1569] Maciej Borkowski. Signal analysis using rough integrals. Lecture Notes in Computer Science, 2475: 218–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2475/24750218.htm;   http:
//link.springer.de/link/service/

                                        series/0558/papers/2475/24750218.
                                        pdf.

Chiang:2002:HMP

[1570] Yi-Ting Chiang, Yu-Cheng Chiang, Tsan sheng Hsu, Churn-Jung Liau, and Da-Wei Wang. How much privacy? — A system to safe guard personal privacy while releasing databases. Lecture Notes in Computer Science, 2475:226–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750226.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750226.pdf.

doPrado:2002:RCA

[1571] Hércules Antonio do Prado, Paulo Martins Engel, and Homero Chaib Filho. Rough clustering: An alternative to find meaningful clusters by using the reducts from a dataset. Lecture Notes in Computer Science, 2475:234–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750234.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750234.pdf.

Dubois:2002:CLA

[1572] Vincent Dubois and Mohamed Quafafou. Concept learning with approximation: Rough version spaces. Lecture Notes in Computer Science, 2475: 239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2475/24750239.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2475/24750239.
pdf.

Giove:2002:VCM

[1573] Silvio Giove, Salvatore Greco, Benedetto Matarazzo, and Roman Słowiński. Variable consistency monotonic decision trees. Lecture Notes in Computer Science, 2475:247–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750247.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750247.pdf.

Greco:2002:IIC

[1574] Salvatore Greco, Benedetto Matarazzo, Roman Slowinski, and Jerzy Stefanowski. Importance and interaction of conditions in decision rules. Lecture Notes in Computer Science, 2475:255–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750255.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750255.pdf.

Lingras:2002:TCR

[1575] Pawan Lingras and Y. Y. Yao.

Time complexity of rough clustering: GAs versus K-means. Lecture Notes in Computer Science, 2475: 263–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2475/24750263.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2475/24750263.
                                        pdf.

Stefanowski:2002:IDR

[1576] Jerzy Stefanowski and Alexis Tsoukiàs. Induction of decision rules and classification in the valued tolerance approach. Lecture Notes in Computer Science, 2475:271–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750271.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750271.pdf.

Valdes:2002:TSM

[1577] Julio J. Valdés and Gabriel Mateescu. Time series model mining with similarity-based neurofuzzy networks and genetic algorithms: a parallel implementation. Lecture Notes in Computer Science, 2475:279–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750279.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750279.pdf.

Alpigini:2002:CPM

[1578] James J. Alpigini. Closeness of performance map information granules: a rough set approach. Lecture Notes in Computer Science, 2475: 289–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2475/24750289.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2475/24750289.
pdf.

Lin:2002:GCB

[1579] Tsau Young (”T. Y.”) Lin. Granular computing on binary relations. Lecture Notes in Computer Science, 2475:296–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750296.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750296.pdf.

Peters:2002:MIC

[1580] James F. Peters, Andrzej Skowron, Zbigniew Suraj, Maciej Borkowski, and Wojciech Rzasa. Measures of inclusion and closeness of information granules: a rough set approach. Lecture Notes in Computer Science, 2475:300–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750300.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750300.pdf.

Peters:2002:RNS

[1581] James F. Peters and Marcin S. Szczuka.

Rough neurocomputing: a survey of basic models of neurocomputation. Lecture Notes in Computer Science, 2475:308–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750308.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750308.pdf.

Skowron:2002:RNB

[1582] Andrzej Skowron, Jarosław Stepaniuk, and James F. Peters. Rough neurocomputing based on hierarchical classifiers. Lecture Notes in Computer Science, 2475:316–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750316.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750316.pdf.

Yager:2002:UGO

[1583] Ronald R. Yager. Using granular objects in multi-source data fusion. Lecture Notes in Computer Science, 2475:324–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750324.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750324.pdf.

Yao:2002:ICR

[1584] J. T. Yao and Y. Y. Yao. Induction of classification rules by granular computing. Lecture Notes in Computer Science, 2475:331–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750331.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750331.pdf.

Butz:2002:AMC

[1585] C. J. Butz and M. J. Sanscartier.

Acquisition methods for contextual weak independence. Lecture Notes in Computer Science, 2475: 339–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2475/24750339.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2475/24750339.
pdf.

Butz:2002:MDC

[1586] C. J. Butz and M. J. Sanscartier. A method for detecting context-specific independence in conditional probability tables. Lecture Notes in Computer Science, 2475:344–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750344.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750344.pdf.

Butz:2002:PWC

[1587] C. J. Butz and M. J. Sanscartier.

Properties of weak conditional independence. Lecture Notes in Computer Science, 2475:349–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750349.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2475/24750349.pdf.

Intan:2002:PPR

[1588] Rolly Intan and Masao Mukaidono. A proposal of probability of rough event based on probability of fuzzy event. Lecture Notes in Computer Science, 2475:357–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750357.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750357.pdf.

Slezak:2002:ABN

[1589] Dominik Slezak and Jakub Wróblewski. Approximate Bayesian network classifiers. Lecture Notes in Computer Science, 2475:365–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750365.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750365.pdf.

Tsumoto:2002:ACR

[1590] Shusaku Tsumoto. Accuracy and coverage in rough set rule induction. Lecture Notes in Computer Science, 2475:373–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750373.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750373.pdf.

Tsumoto:2002:STR

[1591] Shusaku Tsumoto. Statistical test for rough set approximation based on Fisher’s exact test. Lecture Notes in Computer Science, 2475: 381–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2475/24750381.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2475/24750381.
pdf.

Wong:2002:TBN

[1592] S. K. M. Wong, D. Wu, and C. J.

Butz. Triangulation of Bayesian networks: a relational database perspective. Lecture Notes in Computer Science, 2475:389–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750389.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750389.pdf.

Bazan:2002:NVR

[1593] Jan G. Bazan, Marcin S. Szczuka, and Jakub Wróblewski. A new version of rough set exploration system. Lecture Notes in Computer Science, 2475:397–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750397.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750397.pdf.

Gora:2002:LAV

[1594] Grzegorz Góra and Arkadiusz Wojna. Local attribute value grouping for lazy rule induction. Lecture Notes in Computer Science, 2475: 405–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2475/24750405.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2475/24750405.
pdf.

Latkowski:2002:IDD

[1595] Rafał Latkowski. Incomplete data decomposition for classification. Lecture Notes in Computer Science, 2475: 413–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2475/24750413.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2475/24750413.
pdf.

Martinez:2002:ERM

[1596] Juan F. Martinez, Anita Wasilewska, Michael Hadjimichael, Covadonga Fernandez, and Ernestina Menasalvas. Extension of relational management systems with data mining capabilities. Lecture Notes in Computer Science, 2475:421–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750421.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750421.pdf.

Mikolajczyk:2002:RND

[1597] Michał Mikołajczyk. Reducing number of decision rules by joining. Lecture Notes in Computer Science, 2475:425–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750425.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750425.pdf.

Nguyen:2002:SCM

[1598] Hung Son Nguyen. Scalable classification method based on rough sets. Lecture Notes in Computer Science, 2475:433–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750433.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750433.pdf.

Pena:2002:PDM

[1599] José M. Peña, F. Javier Crespo, Ernestina Menasalvas, and Victor Robles. Parallel data mining experimentation using flexible configurations. Lecture Notes in Computer Science, 2475:441–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750441.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750441.pdf.

Perez:2002:OAA

[1600] María S. Pérez, Ramón A. Pons, Félix García, Jesús Carretero, and María L. Córdoba. An optimization of apriori algorithm through the usage of parallel I/O and hints. Lecture Notes in Computer Science, 2475:449–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750449.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750449.pdf.

Skowron:2002:PIM

[1601] Andrzej Skowron and Piotr Synak. Patterns in information maps. Lecture Notes in Computer Science, 2475: 453–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2475/24750453.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2475/24750453.
pdf.

Zhao:2002:DMA

[1602] Min Zhao and Jue Wang. Discernibility matrix approach to exception analysis. Lecture Notes in Computer Science, 2475:461–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750461.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750461.pdf.

Zhong:2002:GCD

[1603] Ning Zhong, Ju-Zhen Dong, Y. Y.

Yao, and Setsuo Ohsuga. Gastric cancer data mining with ordered information. Lecture Notes in Computer Science, 2475:467–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750467.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750467.pdf.

Hochsztain:2002:GAA

[1604] Esther Hochsztain, Socorro Millán, and Ernestina Menasalvas. A granular approach for analyzing the degree of affability of a Web site. Lecture Notes in Computer Science, 2475:479–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750479.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750479.pdf.

Hu:2002:CCM

[1605] Xiaohua Hu. Comparison of classification methods for customer attrition analysis. Lecture Notes in Computer Science, 2475:487–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750487.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750487.pdf.

Li:2002:UPM

[1606] Yuefeng Li and Y. Y. Yao. User profile model: a view from artificial intelligence. Lecture Notes in Computer Science, 2475:493–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750493.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750493.pdf.

Marban:2002:MCL

[1607] Oscar Marban, Javier Segovia, Juan J. Cuadrado, and Cesar Montes. Mining the client’s life cycle behaviour in the Web. Lecture Notes in Computer Science, 2475:497–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750497.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750497.pdf.

Yao:2002:PIW

[1608] Y. Y. Yao, H. J. Hamilton, and Xuewei Wang. PagePrompter: An intelligent Web agent created using data mining techniques. Lecture Notes in Computer Science, 2475:506–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750506.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750506.pdf.

Ziarko:2002:VAW

[1609] Wojciech Ziarko and Xue Fei. VPRSM approach to WEB searching. Lecture Notes in Computer Science, 2475: 514–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2475/24750514.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2475/24750514.
pdf.

Bazan:2002:RSA

[1610] Jan Bazan, Antoni Osmólski, Andrzej Skowron, Dominik Slezak, Marcin Szczuka, and Jakub Wróblewski. Rough set approach to the survival analysis. Lecture Notes in Computer Science, 2475:522–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750522.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750522.pdf.

Beynon:2002:ILP

[1611] Malcolm J. Beynon. The identification of low-paying workplaces: An analysis using the variable precision rough sets model. Lecture Notes in Computer Science, 2475:530–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750530.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750530.pdf.

Grzymala-Busse:2002:SBD

[1612] Jerzy W. Grzymała-Busse and Zdzisław S. Hippe. A search for the best data mining method to predict melanoma. Lecture Notes in Computer Science, 2475:538–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750538.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750538.pdf.

Hippe:2002:TCM

[1613] Monika P. Hippe. Towards the classification of musical works: a rough set approach. Lecture Notes in Computer Science, 2475:546–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750546.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750546.pdf.

Hirano:2002:SMI

[1614] Shoji Hirano and Shusaku Tsumoto. Segmentation of medical images based on approximations in rough set theory. Lecture Notes in Computer Science, 2475:554–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750554.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750554.pdf.

Jang:2002:ARE

[1615] Seok-Woo Jang, Essam A. El-Kwae, and Hyung-Il Choi. Adaptive robust estimation for filtering motion vectors. Lecture Notes in Computer Science, 2475:564–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750564.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750564.pdf.

Lee:2002:RSF

[1616] Seungkoo Lee, Nicholas Propes, Guangfan Zhang, Yongshen Zhao, and George Vachtsevanos. Rough set feature selection and diagnostic rule generation for industrial applications. Lecture Notes in Computer Science, 2475:568–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750568.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750568.pdf.

Chen:2002:CAR

[1617] Li Chen. λ-connected approximations for rough sets. Lecture Notes in Computer Science, 2475: 572–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2475/24750572.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2475/24750572.
pdf.

Nguyen:2002:ACC

[1618] Tuan Trung Nguyen. Adaptive classifier construction: An approach to handwritten digit recognition. Lecture Notes in Computer Science, 2475:578–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750578.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750578.pdf.

Paszek:2002:ASD

[1619] Piotr Paszek and Alicja Wakulicz-Deja. The application of support diagnose in mitochondrial encephalomyopathies. Lecture Notes in Computer Science, 2475:586–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750586.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750586.pdf.

Peters:2002:OCL

[1620] James F. Peters, T. C. Ahn, and Maciej Borkowski. Obstacle classification by a line-crawling robot: a rough neurocomputing approach. Lecture Notes in Computer Science, 2475:594–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750594.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750594.pdf.

Ramanna:2002:RNN

[1621] Sheela Ramanna. Rough neural network for software change prediction. Lecture Notes in Computer Science, 2475:602–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750602.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750602.pdf.

Sinha:2002:HSU

[1622] D. Sinha and P. Laplante. Handling spatial uncertainty in binary images: a rough set based approach. Lecture Notes in Computer Science, 2475:610–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750610.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750610.pdf.

Smolinski:2002:EAR

[1623] Tomasz G. Smolinski, Grzegorz M. Boratyn, Mariofanna Milanova, Jacek M. Zurada, and Andrzej Wrobel. Evolutionary algorithms and rough setsbased hybrid approach to classificatory decomposition of cortical evoked potentials. Lecture Notes in Computer Science, 2475:621–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2475/24750621.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2475/24750621.pdf.

Szmigielski:2002:RML

[1624] Adam Szmigielski. Rough mereological localization and navigation. Lecture Notes in Computer Science, 2475:629–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2475/24750629.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2475/24750629.pdf.

Anonymous:2002:AIbj

[1625] Anonymous. Author index. Lecture Notes in Computer Science, 2475: 639–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2475/2475auth.pdf.

Ley:2002:DCS

[1626] Michael Ley. The DBLP computer science bibliography: Evolution, research issues, perspectives. Lecture Notes in Computer Science, 2476:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2476/24760001.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2476/24760001.pdf.

Suciu:2002:STQ

[1627] Dan Suciu. From searching text to querying XML streams. Lecture Notes in Computer Science, 2476: 11–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2476/24760011.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2476/24760011.
pdf.

Gonnet:2002:SMP

[1628] Gaston H. Gonnet. String matching problems from bioinformatics which still need better solutions extended abstract. Lecture Notes in Computer Science, 2476:27–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2476/24760027.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2476/24760027.pdf.

Abouelhoda:2002:OES

[1629] Mohamed Ibrahim Abouelhoda, Enno Ohlebusch, and Stefan Kurtz. Optimal exact string matching based on suffix arrays. Lecture Notes in Computer Science, 2476:31–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2476/24760031.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2476/24760031.pdf.

Fredriksson:2002:FSM

[1630] Kimmo Fredriksson. Faster string matching with super-alphabets. Lecture Notes in Computer Science, 2476: 44–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/

                                        0558/bibs/2476/24760044.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2476/24760044.
                                        pdf.

Crochemore:2002:SDM

[1631] Maxime Crochemore and Zdenek Tronícek. On the size of DASG for multiple texts. Lecture Notes in Computer Science, 2476:58–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2476/24760058.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2476/24760058.pdf.

Dias:2002:SPT

[1632] Zanoni Dias and João Meidanis. Sorting by prefix transpositions. Lecture Notes in Computer Science, 2476:65–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2476/24760065.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2476/24760065.pdf.

Arslan:2002:ECL

[1633] Abdullah N. Arslan and Ömer Egecioglu. Efficient computation of long similar subsequences. Lecture Notes in Computer Science, 2476: 77–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2476/24760077.htm;   http:
                                        //link.springer.de/link/service/

series/0558/papers/2476/24760077.
pdf.

Brisaboa:2002:SGT

[1634] Nieves R. Brisaboa, Carlos Callón, Juan-Ramón López, Ángeles S. Places, and Goretti Sanmartín. Stemming galician texts. Lecture Notes in Computer Science, 2476:91–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2476/24760091.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2476/24760091.pdf.

Sughaiyer:2002:FPA

[1635] Imad A. Al Sughaiyer and Ibrahim A. Al Kharashi. Firing policies for an Arabic rule-based stemmer. Lecture Notes in Computer Science, 2476:98–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2476/24760098.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2476/24760098.pdf.

Possas:2002:ESB

[1636] Bruno Pôssas, Nivio Ziviani, and Wagner Meira, Jr. Enhancing the setbased model using proximity information. Lecture Notes in Computer Science, 2476:104–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2476/24760104.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2476/24760104.pdf.

Baeza-Yates:2002:WSD

[1637] Ricardo Baeza-Yates, Felipe Saint-Jean, and Carlos Castillo. Web structure, dynamics and page quality. Lecture Notes in Computer Science, 2476:117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2476/24760117.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2476/24760117.pdf.

Pretto:2002:TAG

[1638] Luca Pretto. A theoretical analysis of Google’s PageRank. Lecture Notes in Computer Science, 2476: 131–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2476/24760131.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2476/24760131.
pdf.

Xi:2002:MLA

[1639] Wensi Xi, Edward A. Fox, Roy P. Tan, and Jiang Shu. Machine learning approach for homepage finding task. Lecture Notes in Computer Science, 2476:145–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2476/24760145.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2476/24760145.pdf.

Chauve:2002:TPM

[1640] Cedric Chauve. Tree pattern matching for linear static terms. Lecture Notes in Computer Science, 2476:160–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2476/24760160.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2476/24760160.pdf.

Takeda:2002:PTF

[1641] Masayuki Takeda, Satoru Miyamoto, Takuya Kida, Ayumi Shinohara, Shuichi Fukamachi, Takeshi Shinohara, and Setsuo Arikawa. Processing text files as is: Pattern matching over compressed texts, multi-byte character texts, and semi-structured texts. Lecture Notes in Computer Science, 2476:170–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2476/24760170.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2476/24760170.pdf.

Harada:2002:PMM

[1642] Lilian Harada. Pattern matching over multi-attribute data streams. Lecture Notes in Computer Science, 2476:187–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2476/24760187.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2476/24760187.pdf.

Goncalves:2002:JMO

[1643] Marcos André Gonçalves, Paul Mather, Jun Wang, Ye Zhou, Ming Luo, Ryan Richardson, Rao Shen, Liang Xu, and Edward A. Fox. Java MARIAN: From an OPAC to a modern digital library system. Lecture Notes in Computer Science, 2476:194–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2476/24760194.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2476/24760194.pdf.

deCastroReis:2002:FGA

[1644] Davi de Castro Reis, Robson Braga Araújo, Altigran S. da Silva, and Berthier A. Ribeiro-Neto. A framework for generating attribute extractors for Web data sources. Lecture Notes in Computer Science, 2476:210–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2476/24760210.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2476/24760210.pdf.

Tahaghoghi:2002:MEQ

[1645] Seyed M. M. Tahaghoghi, James A. Thom, and Hugh E. Williams. Multiple example queries in content-based image retrieval. Lecture Notes in Computer Science, 2476:227–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2476/24760227.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2476/24760227.pdf.

Kazai:2002:FSD

[1646] Gabrialla Kazai, Mounia Lalmas, and Thomas Roelleke. Focussed structured document retrieval. Lecture Notes in Computer Science, 2476: 241–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2476/24760241.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2476/24760241.
pdf.

Gonzalez-Caro:2002:TMC

[1647] Cristina N. González-Caro, Maritza L. Calderón-Benavides, José de J. Pérez-Alcázar, Juan C. García-Díaz, and Joaquin Delgado. Towards a more comprehensive comparison of collaborative filtering algorithms. Lecture Notes in Computer Science, 2476:248–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2476/24760248.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2476/24760248.pdf.

Navarro:2002:FDS

[1648] Gonzalo Navarro and Nora Reyes.

Fully dynamic spatial approximation trees. Lecture Notes in Computer Science, 2476:254–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2476/24760254.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2476/24760254.pdf.

Bartolini:2002:SMM

[1649] Ilaria Bartolini, Paolo Ciaccia, and Marco Patella. String matching with metric trees using an approximate distance. Lecture Notes in Computer Science, 2476:271–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2476/24760271.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2476/24760271.pdf.

Bustos:2002:PPS

[1650] Benjamin Bustos and Gonzalo Navarro. Probabilistic proximity searching algorithms based on compact partitions. Lecture Notes in Computer Science, 2476:284–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2476/24760284.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2476/24760284.pdf.

Navarro:2002:SDS

[1651] Gonzalo Navarro, Rodrigo Paredes, and Edgar Chávez. t-spanners as a data structure for metric space searching. Lecture Notes in Computer Science, 2476:298–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2476/24760298.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2476/24760298.pdf.

Inenaga:2002:CDA

[1652] Shunsuke Inenaga, Ayumi Shinohara, Masayuki Takeda, and Setsuo Arikawa. Compact directed acyclic word graphs for a sliding window. Lecture Notes in Computer Science, 2476:310–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2476/24760310.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2476/24760310.pdf.

Navarro:2002:ITU

[1653] Gonzalo Navarro. Indexing text using the Ziv–Lempel trie. Lecture Notes in Computer Science, 2476: 325–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2476/24760325.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2476/24760325.
pdf.

Anonymous:2002:AIbk

[1654] Anonymous. Author index. Lecture Notes in Computer Science, 2476: 337–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2476/2476ind.pdf.

Boots:2002:IRP

[1655] Barry Boots, Robert Feick, Narushige Shiode, and Steven Roberts. Investigating recursive point Voronoi diagrams. Lecture Notes in Computer Science, 2478:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2478/24780001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2478/24780001.pdf.

Brodaric:2002:DIE

[1656] Boyan Brodaric and Mark Gahegan. Distinguishing instances and evidence of geographical concepts for geospatial database design. Lecture Notes in Computer Science, 2478:22–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2478/24780022.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2478/24780022.pdf.

Brunsdon:2002:GWL

[1657] Chris Brunsdon, Stewart Fotheringham, and Martin Charlton. Geographically weighted local statistics applied to binary data. Lecture Notes in Computer Science, 2478: 38–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2478/24780038.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2478/24780038.
pdf.

Buttenfield:2002:TVG

[1658] Barbara P. Buttenfield. Transmitting vector geospatial data across the Internet. Lecture Notes in Computer Science, 2478:51–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2478/24780051.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2478/24780051.pdf.

Cai:2002:GIR

[1659] Guoray Cai. GeoVSM: An integrated retrieval model for geographic information. Lecture Notes in Computer Science, 2478:65–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2478/24780065.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2478/24780065.pdf.

Duff:2002:EBA

[1660] Damien Duff and Hans W. Guesgen. An evaluation of buffering algorithms in fuzzy GISs. Lecture Notes in Computer Science, 2478: 80–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2478/24780080.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2478/24780080.
pdf.

Fonseca:2002:GEO

[1661] Frederico Fonseca, James Martin, and M. Andrea Rodríguez. From geo- to eco-ontologies. Lecture Notes in Computer Science, 2478:

93–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2478/24780093.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2478/24780093.
                                        pdf.

Kuhn:2002:MSG

[1662] Werner Kuhn. Modeling the semantics of geographic categories through conceptual integration. Lecture Notes in Computer Science, 2478: 108–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2478/24780108.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2478/24780108.
                                        pdf.

Kulik:2002:GAD

[1663] Lars Kulik, Carola Eschenbach, Christopher Habel, and Hedda Rahel Schmidtke. A graded approach to directions between extended objects. Lecture Notes in Computer Science, 2478:119–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2478/24780119.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2478/24780119.pdf.

Laube:2002:ARM

[1664] Patrick Laube and Stephan Imfeld. Analyzing relative motion within groups of trackable moving point objects. Lecture Notes in Computer Science, 2478:132–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2478/24780132.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2478/24780132.pdf.

Li:2002:CST

[1665] Lixin Li and Peter Revesz. A comparison of spatio-temporal interpolation methods. Lecture Notes in Computer Science, 2478:145–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2478/24780145.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2478/24780145.pdf.

Li:2002:MAS

[1666] Maozhen Li, Sheng Zhou, and Christopher B. Jones. Multi-agent systems for Web-based map information retrieval. Lecture Notes in Computer Science, 2478:161–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2478/24780161.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2478/24780161.pdf.

Liebisch:2002:NMG

[1667] Niels Liebisch, Geoffrey Jacquez, Pierre Goovaerts, and Andreas Kaufmann. New methods to generate neutral images for spatial pattern recognition. Lecture Notes in Computer Science, 2478:181–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2478/24780181.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2478/24780181.pdf.

deCMoura:2002:CMD

[1668] Ana Maria de C. Moura, Marcio Victorino, and Asterio Tanaka. Combining mediator and data warehouse technologies for developing environmental decision support systems. Lecture Notes in Computer Science, 2478:196–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2478/24780196.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2478/24780196.pdf.

vanderPoorten:2002:TCM

[1669] P. M. van der Poorten, Sheng Zhou, and Christopher B. Jones. Topologically-consistent map generalisation procedures and multiscale spatial databases. Lecture Notes in Computer Science, 2478: 209–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2478/24780209.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2478/24780209.
                                        pdf.

Rainsford:2002:MJP

[1670] Desmond Rainsford and William A. Mackaness. Mobile journey planning for bus passengers. Lecture Notes in Computer Science, 2478: 228–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2478/24780228.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2478/24780228.
pdf.

Raubal:2002:EWI

[1671] Martin Raubal and Stephan Winter. Enriching wayfinding instructions with local landmarks. Lecture Notes in Computer Science, 2478:243–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2478/24780243.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2478/24780243.pdf.

Rodriguez:2002:CBA

[1672] M. Andrea Rodríguez and Francisco A.

Godoy. A content-based approach to searching and indexing spatial configurations. Lecture Notes in Computer Science, 2478:260–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2478/24780260.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2478/24780260.pdf.

Roy:2002:QAD

[1673] Anthony J. Roy and John G. Stell.

A qualitative account of discrete space. Lecture Notes in Computer Science, 2478:276–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2478/24780276.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2478/24780276.pdf.

Santos:2002:NAI

[1674] Jorge Santos, Weldon A. Lodwick, and Arnold Neumaier. A new approach to incorporate uncertainty in terrain modeling. Lecture Notes in Computer Science, 2478:291–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2478/24780291.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2478/24780291.pdf.

Shirabe:2002:DIP

[1675] Takeshi Shirabe and C. Dana Tomlin. Decomposing integer programming models for spatial allocation. Lecture Notes in Computer Science, 2478:300–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2478/24780300.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2478/24780300.pdf.

Tsou:2002:OMF

[1676] Ming-Hsiang Tsou. An operational metadata framework for searching, indexing, and retrieving distributed geographic information services on the Internet. Lecture Notes in Computer Science, 2478:313–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2478/24780313.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2478/24780313.pdf.

Walsh:2002:HII

[1677] Kuuipo A. Walsh, Cherri M. Pancake, Dawn J. Wright, Sally Haerer, and F. J. Hanus. “Humane” interfaces to improve the usability of data clearinghouses. Lecture Notes in Computer Science, 2478:333–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2478/24780333.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2478/24780333.pdf.

Worboys:2002:IST

[1678] Michael Worboys and Matt Duckham.

Integrating spatio-thematic information. Lecture Notes in Computer Science, 2478:346–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2478/24780346.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2478/24780346.pdf.

Anonymous:2002:AIbl

[1679] Anonymous. Author index. Lecture Notes in Computer Science, 2478: 363–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2478/2478auth.pdf.

Amalberti:2002:UMS

[1680] Rene Amalberti. Use and misuse of safety models in design. Lecture Notes in Computer Science, 2485:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2485/24850001.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2485/24850001.pdf.

Porcarelli:2002:EOQ

[1681] Stefano Porcarelli and Felicita Di Giandomenico. On the effects of outages on the QoS of GPRS networks under different user characterizations. Lecture Notes in Computer Science, 2485:2–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2485/24850002.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2485/24850002.pdf.

Bieber:2002:CFT

[1682] Pierre Bieber, Charles Castel, and Christel Seguin. Combination of fault tree analysis and model checking for safety assessment of complex system. Lecture Notes in Computer Science, 2485:19–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2485/24850019.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2485/24850019.pdf.

Csertan:2002:BBR

[1683] György Csertán, András Pataricza, P. Harang, Orsolya Dobán, Gabor Biros, András Dancsecz, and Ferenc Friedler. BPM based robust E-business application development (practical experience report). Lecture Notes in Computer Science, 2485:32–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2485/24850032.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2485/24850032.pdf.

Pedone:2002:SAP

[1684] Fernando Pedone, André Schiper, Péter Urbán, and David Cavin. Solving agreement problems with weak ordering oracles. Lecture Notes in Computer Science, 2485:44–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2485/24850044.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2485/24850044.pdf.

Anceaume:2002:ESS

[1685] Emmanuelle Anceaume, Michel Hurfin, and Philippe Raipin Parvedy. An efficient solution to the k-set agreement problem. Lecture Notes in Computer Science, 2485:62–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2485/24850062.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2485/24850062.pdf.

deLemos:2002:NAD

[1686] Rogério de Lemos. Novel approaches in dependable computing. Lecture Notes in Computer Science, 2485:79–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2485/24850079.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2485/24850079.pdf.

Avizienis:2002:ISP

[1687] Algirdas Avizienis. An immune system paradigm for the design of fault tolerant systems. Lecture Notes in Computer Science, 2485: 81–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2485/24850081.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2485/24850081.
                                        pdf.

Bigham:2002:SSL

[1688] John Bigham. Security and survivability of large scale critical infrastructures. Lecture Notes in Computer Science, 2485:84–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2485/24850084.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2485/24850084.pdf.

Fiadeiro:2002:AAF

[1689] José Luiz Fiadeiro. An architectural approach to fault treatment in critical infrastructures. Lecture Notes in Computer Science, 2485: 86–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2485/24850086.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2485/24850086.
pdf.

Tyrrell:2002:BIF

[1690] Andy Tyrrell. Biologically inspired fault-tolerant computer systems. Lecture Notes in Computer Science, 2485:88–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2485/24850088.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2485/24850088.pdf.

Bellos:2002:TSE

[1691] Maciej Bellos, Dimitri Kagaris, and Dimitris Nikolos. Test set embedding based on phase shifters. Lecture Notes in Computer Science, 2485:90–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2485/24850090.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2485/24850090.pdf.

Cunha:2002:RDF

[1692] João Carlos Cunha, António Correia, Jorge Henriques, Mário Zenha Rela, and João Gabriel Silva. Resetdriven fault tolerance. Lecture Notes in Computer Science, 2485: 102–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2485/24850102.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2485/24850102.
pdf.

Majzik:2002:TDM

[1693] István Majzik and Gábor Huszerl. Towards dependability modeling of FT-CORBA architectures. Lecture Notes in Computer Science, 2485:121–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2485/24850121.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2485/24850121.pdf.

Joshi:2002:EEU

[1694] Kaustubh R. Joshi, Michel Cukier, and William H. Sanders. Experimental evaluation of the unavailability induced by a group membership protocol. Lecture Notes in Computer Science, 2485:140–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2485/24850140.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2485/24850140.pdf.

Sieh:2002:UVS

[1695] Volkmar Sieh and Kerstin Buchacker.

UMLinux — A versatile SWIFI tool. Lecture Notes in Computer Science, 2485:159–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2485/24850159.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2485/24850159.pdf.

Ademaj:2002:MDE

[1696] Astrit Ademaj. A methodology for dependability evaluation of the time-triggered architecture using software implemented fault injection. Lecture Notes in Computer Science, 2485:172–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2485/24850172.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2485/24850172.pdf.

Dutta:2002:FIC

[1697] Partha Dutta and Rachid Guerraoui. Fast indulgent consensus with zero degradation. Lecture Notes in Computer Science, 2485: 191–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2485/24850191.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2485/24850191.
pdf.

Pedone:2002:PQL

[1698] Fernando Pedone, Nelson L. Duarte, and Mario Goulart. Probabilistic queries in large-scale networks. Lecture Notes in Computer Science, 2485:209–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2485/24850209.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2485/24850209.pdf.

Simoncini:2002:TIS

[1699] Luca Simoncini. Towards information society dependability initiative in FP6: Roadmapping activities in dependability. Lecture Notes in Computer Science, 2485:227–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2485/24850227.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2485/24850227.pdf.

Correia:2002:DCR

[1700] Miguel Correia, Paulo Veríssimo, and Nuno Ferreira Neves. The design of a COTS real-time distributed security kernel. Lecture Notes in Computer Science, 2485:234–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2485/24850234.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2485/24850234.pdf.

Rodriguez:2002:WRT

[1701] Manuel Rodríguez, Jean-Charles Fabre, and Jean Arlat. Wrapping real-time systems from temporal logic specifications. Lecture Notes in Computer Science, 2485:253–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2485/24850253.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2485/24850253.pdf.

Herout:2002:MBD

[1702] Pavel Herout, Stanislav Racek, and Jan Hlavicka. Model-based dependability evaluation method for TTP/C based systems. Lecture Notes in Computer Science, 2485:271–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2485/24850271.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2485/24850271.pdf.

Anonymous:2002:AIbm

[1703] Anonymous. Author index. Lecture Notes in Computer Science, 2485: 283–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2485/2485ind.pdf.

Valentini:2002:ELM

[1704] Giorgio Valentini and Francesco Masulli. Ensembles of learning machines. Lecture Notes in Computer Science, 2486:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2486/24860003.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2486/24860003.pdf.

Evgniou:2002:LPR

[1705] Theodoros Evgniou and Massimiliano Pontil. Learning preference relations from data. Lecture Notes in Computer Science, 2486:23–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2486/24860023.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2486/24860023.pdf.

Lauria:2002:IBI

[1706] Francesco E. Lauria. Increasing the biological inspiration of neural networks. Lecture Notes in Computer Science, 2486:35–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2486/24860035.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2486/24860035.pdf.

Corazza:2002:HAT

[1707] Marco Corazza, Paolo Vanni, and Umberto Loschi. Hybrid automatic trading systems: Technical analysis & group method of data handling. Lecture Notes in Computer Science, 2486:47–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2486/24860047.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2486/24860047.pdf.

Giove:2002:ITM

[1708] Silvio Giove. Interval TOPSIS for multicriteria decision making. Lecture Notes in Computer Science, 2486:56–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2486/24860056.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2486/24860056.pdf.

Grossi:2002:DAM

[1709] Giuliano Grossi and Roberto Posenato.

A distributed algorithm for max independent set problem based on Hopfield networks. Lecture Notes in Computer Science, 2486:64–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2486/24860064.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2486/24860064.pdf.

Martinelli:2002:ERN

[1710] G. Martinelli, F. M. Frattale Mascioli, M. Panella, and A. Rizzi. Extended random neural networks. Lecture Notes in Computer Science, 2486:75–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2486/24860075.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2486/24860075.pdf.

Palmieri:2002:GIC

[1711] Francesco Palmieri and Alessandra Budillon. Generalized independent component analysis as density estimation. Lecture Notes in Computer Science, 2486:83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2486/24860083.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2486/24860083.pdf.

Topi:2002:SRN

[1712] Lorenzo Topi, Raffaele Parisi, and Aurelio Uncini. Spline recurrent neural networks for quad-tree video coding. Lecture Notes in Computer Science, 2486:90–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2486/24860090.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2486/24860090.pdf.

Vitabile:2002:MNN

[1713] Salvatore Vitabile, Antonio Gentile, G. B. Dammone, and Filippo Sorbello. MLP neural network implementation on a SIMD architecture. Lecture Notes in Computer Science, 2486:99–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2486/24860099.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2486/24860099.pdf.

Azzerboni:2002:NAD

[1714] Bruno Azzerboni, Giovanni Finocchio, Maurizio Ipsale, Fabio La Foresta, and Francesco Carlo Morabito. A new approach to detection of muscle activation by independent component analysis and wavelet transform. Lecture Notes in Computer Science, 2486:109–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2486/24860109.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2486/24860109.pdf.

Borghese:2002:LBU

[1715] N. Alberto Borghese. Learning to balance upright posture: What can be learnt using adaptive NN models? Lecture Notes in Computer Science, 2486:117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2486/24860117.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2486/24860117.pdf.

Campadelli:2002:DFF

[1716] Paola Campadelli, Elena Casiraghi, and Raffaella Lanzarotti. Detection of facial features. Lecture Notes in Computer Science, 2486: 124–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2486/24860124.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2486/24860124.
                                        pdf.

Chella:2002:TSN

[1717] Antonio Chella and Roberto Pirrone. A two stage neural architecture for segmentation and superquadrics recovery from range data. Lecture Notes in Computer Science, 2486:132–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2486/24860132.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2486/24860132.pdf.

Ezin:2002:ADE

[1718] E. C. Ezin, Flora Giudicepietro, Simona Petrosino, S. Scarpetta, and A. Vanacore. Automatic discrimination of earthquakes and false events in seismological recording for volcanic monitoring. Lecture Notes in Computer Science, 2486:140–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2486/24860140.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2486/24860140.pdf.

Mattera:2002:CSC

[1719] Davide Mattera, Francesco Palmieri, and Michele Di Monte. A comparison of signal compression methods by sparse solution of linear systems. Lecture Notes in Computer Science, 2486:146–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2486/24860146.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2486/24860146.pdf.

Morabito:2002:FTS

[1720] Francesco Carlo Morabito and Mario Versaci. Fuzzy time series for forecasting pollutants concentration in the air. Lecture Notes in Computer Science, 2486:152–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2486/24860152.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2486/24860152.pdf.

Pasero:2002:RTP

[1721] Eros Pasero and Alfonso Montuori.

Real-time perceptual coding of wideband speech by competitive neural networks. Lecture Notes in Computer Science, 2486:160–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2486/24860160.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2486/24860160.pdf.

Uncini:2002:SSF

[1722] Aurelio Uncini. Sound synthesis by flexible activation function recurrent neural networks. Lecture Notes in Computer Science, 2486: 168–??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2486/24860168.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2486/24860168.
                                        pdf.

Anthony:2002:MMG

[1723] Martin Anthony. Mathematical modelling of generalization. Lecture Notes in Computer Science, 2486: 181–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2486/24860181.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2486/24860181.
                                        pdf.

Bertoni:2002:SCN

[1724] Alberto Bertoni and Beatrice Palano. Structural complexity and neural networks. Lecture Notes in Computer Science, 2486:190–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2486/24860190.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2486/24860190.pdf.

Eleuteri:2002:BLT

[1725] A. Eleuteri, R. Tagliaferri, L. Milano, F. Acernese, and M. De Laurentiis. Bayesian learning techniques: Application to neural networks with constraints on weight space. Lecture Notes in Computer Science, 2486:216–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2486/24860216.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2486/24860216.pdf.

Pontil:2002:SRS

[1726] Massimiliano Pontil. A short review of statistical learning theory. Lecture Notes in Computer Science, 2486:233–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2486/24860233.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2486/24860233.pdf.

Parisi:2002:IBI

[1727] Domenico Parisi. Increasing the biological inspiration of neural networks. Lecture Notes in Computer Science, 2486:243–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2486/24860243.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2486/24860243.pdf.

Anonymous:2002:AIbn

[1728] Anonymous. Author index. Lecture Notes in Computer Science, 2486: 253–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2486/2486ind.pdf.

Jones:2002:PGT

[1729] Neil D. Jones and Arne J. Glenstrup. Program generation, termination, and binding-time analysis. Lecture Notes in Computer Science, 2487:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2487/24870001.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2487/24870001.pdf.

Sztipanovits:2002:GPE

[1730] Janos Sztipanovits and Gabor Karsai. Generative programming for embedded systems. Lecture Notes in Computer Science, 2487:32–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2487/24870032.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2487/24870032.pdf.

Attardi:2002:SRA

[1731] Giuseppe Attardi and Antonio Cisternino. Self reflection for adaptive programming. Lecture Notes in Computer Science, 2487:50–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2487/24870050.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2487/24870050.pdf.

Back:2002:DSS

[1732] Godmar Back. DataScript — A specification and scripting language for binary data. Lecture Notes in Computer Science, 2487:66–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2487/24870066.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2487/24870066.pdf.

Balat:2002:MTD

[1733] Vincent Balat and Olivier Danvy.

Memoization in type-directed partial evaluation. Lecture Notes in Computer Science, 2487:78–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2487/24870078.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2487/24870078.pdf.

Barbeau:2002:PSD

[1734] Michel Barbeau and Francis Bordeleau. A protocol stack development tool using generative programming. Lecture Notes in Computer Science, 2487:93–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2487/24870093.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2487/24870093.pdf.

Brichau:2002:BCA

[1735] Johan Brichau, Kim Mens, and Kris De Volder. Building composable aspectspecific languages with logic metaprogramming. Lecture Notes in Computer Science, 2487:110–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2487/24870110.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2487/24870110.pdf.

Butler:2002:ARF

[1736] Greg Butler. Architectural refactoring in framework evolution: a case study. Lecture Notes in Computer Science, 2487:128–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2487/24870128.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2487/24870128.pdf.

Chin:2002:TMP

[1737] Wei-Ngan Chin and Zhenjiang Hu. Towards a modular program derivation via fusion and tupling. Lecture Notes in Computer Science, 2487: 140–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2487/24870140.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2487/24870140.
                                        pdf.

Czarnecki:2002:GPE

[1738] Krzysztof Czarnecki, Thomas Bednasch, Peter Unger, and Ulrich Eisenecker. Generative programming for embedded software: An industrial experience report. Lecture Notes in Computer Science, 2487: 156–??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2487/24870156.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2487/24870156.
pdf.

Douence:2002:FDR

[1739] Rémi Douence, Pascal Fradet, and Mario Südholt. A framework for the detection and resolution of aspect interactions. Lecture Notes in Computer Science, 2487:173–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2487/24870173.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2487/24870173.pdf.

Elrad:2002:AOM

[1740] Tzilla Elrad, Omar Aldawud, and Atef Bader. Aspect-oriented modeling: Bridging the gap between implementation and design. Lecture Notes in Computer Science, 2487: 189–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2487/24870189.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2487/24870189.
pdf.

Kiselyov:2002:MCS

[1741] Oleg Kiselyov. Macros that compose: Systematic macro programming. Lecture Notes in Computer Science, 2487:202–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2487/24870202.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2487/24870202.pdf.

Lee:2002:PTA

[1742] Chin Soon Lee. Program termination analysis in polynomial time. Lecture Notes in Computer Science, 2487:218–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2487/24870218.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2487/24870218.pdf.

Neema:2002:GSQ

[1743] Sandeep Neema, Ted Bapty, Jeff Gray, and Aniruddha Gokhale. Generators for synthesis of QoS adaptation in distributed real-time embedded systems. Lecture Notes in Computer Science, 2487:236–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2487/24870236.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2487/24870236.pdf.

Noga:2002:OCM

[1744] Markus Noga and Florian Krüper. Optimizing content management system pipelines. Lecture Notes in Computer Science, 2487:252–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2487/24870252.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2487/24870252.pdf.

Saraiva:2002:CBP

[1745] João Saraiva. Component-based programming for higher-order attribute grammars. Lecture Notes in Computer Science, 2487:268–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2487/24870268.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2487/24870268.pdf.

Tanter:2002:AJS

[1746] Éric Tanter, Marc Ségura-Devillechaise, Jacques Noyé, and José Piquer. Altering Java semantics via bytecode manipulation. Lecture Notes in Computer Science, 2487:283–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2487/24870283.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2487/24870283.pdf.

Visser:2002:MPC

[1747] Eelco Visser. Meta-programming with concrete object syntax. Lecture Notes in Computer Science, 2487: 299–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2487/24870299.htm;   http:
//link.springer.de/link/service/

                                        series/0558/papers/2487/24870299.
                                        pdf.

Wang:2002:MDC

[1748] Zhenghao Wang and Richard R. Muntz. Managing dynamic changes in multi-stage program generation systems. Lecture Notes in Computer Science, 2487:316–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2487/24870316.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2487/24870316.pdf.

Anonymous:2002:AIbo

[1749] Anonymous. Author index. Lecture Notes in Computer Science, 2487: 335–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2487/2487auth.pdf.

Nishikawa:2002:UES

[1750] Atsushi Nishikawa, Daiji Negoro, Haruhiko Kakutani, Fumio Miyazaki, Mitsugu Sekimoto, Masayoshi Yasui, Shuji Takiguchi, and Morito Monden. Using an endoscopic solo surgery simulator for quantitative evaluation of human-machine interface in robotic camera positioning systems. Lecture Notes in Computer Science, 2488:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880001.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2488/24880001.pdf.

Krupa:2002:ADP

[1751] A. Krupa, M. de Mathelin, C. Doignon, J. Gangloff, G. Morel, L. Soler, J. Leroy, and J. Marescaux. Automatic 3-D positioning of surgical instruments during robotized laparoscopic surgery using automatic visual feedback. Lecture Notes in Computer Science, 2488:9–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880009.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880009.pdf.

Berkelman:2002:DCC

[1752] Peter J. Berkelman, Philippe Cinquin, Jocelyne Troccaz, Jean-Marc Ayoubi, and Christian Létoublon. Development of a compact cabledriven laparoscopic endoscope manipulator. Lecture Notes in Computer Science, 2488:17–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880017.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880017.pdf.

Mourgues:2002:FCA

[1753] Fabien Mourgues and Ève Coste-Manière. Flexible calibration of actuated stereoscopic endoscope for overlay in robot assisted surgery. Lecture Notes in Computer Science, 2488:25–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880025.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880025.pdf.

Cotin:2002:MLS

[1754] Stephane Cotin, Nicholas Stylopoulos, Mark Ottensmeyer, Paul Neumann, David Rattner, and Steven Dawson. Metrics for laparoscopic skills trainers: The weakest link! Lecture Notes in Computer Science, 2488:35–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880035.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880035.pdf.

Yamauchi:2002:SSE

[1755] Yasushi Yamauchi, Juli Yamashita, Osamu Morikawa, Ryoichi Hashimoto, Masaaki Mochimaru, Yukio Fukui, Hiroshi Uno, and Kazunori Yokoyama. Surgical skill evaluation by force data for endoscopic sinus surgery training system. Lecture Notes in Computer Science, 2488:44–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880044.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880044.pdf.

Jinno:2002:DMS

[1756] Makoto Jinno, Nobuto Matsuhira, Takamitsu Sunaoshi, Takehiro Hato, Toyomi Miyagawa, Yasuhide Morikawa, Toshiharu Furukawa, Soji Ozawa, Masaki Kitajima, and Kazuo Nakazawa. Development of a master slave combined manipulator for laparoscopic surgery — functional model and its evaluation. Lecture Notes in Computer Science, 2488:52–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880052.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880052.pdf.

Koizumi:2002:DTD

[1757] Naoshi Koizumi, Kazuki Sumiyama, Naoki Suzuki, Asaki Hattori, Hisao Tajiri, and Akihiko Uchiyama. Development of three-dimensional endoscopic ultrasound system with optical tracking. Lecture Notes in Computer Science, 2488:60–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880060.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880060.pdf.

Hu:2002:RTH

[1758] Tie Hu, Andres E. Castellanos, Gregory Tholey, and Jaydev P. Desai. Real-time haptic feedback in laparoscopic tools for use in gastro-intestinal surgery. Lecture Notes in Computer Science, 2488:66–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2488/24880066.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880066.pdf.

Kobayashi:2002:SOR

[1759] Yuki Kobayashi, Shingo Chiyoda, Kouichi Watabe, Masafumi Okada, and Yoshihiko Nakamura. Small occupancy robotic mechanisms for endoscopic surgery. Lecture Notes in Computer Science, 2488:75–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880075.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880075.pdf.

Tajima:2002:DMC

[1760] Fujio Tajima, Kousuke Kishi, Kouji Nishizawa, Kazutoshi Kan, Yasuhiro Nemoto, Haruo Takeda, Shin ichiro Umemura, Hiroshi Takeuchi, Masakatsu G. Fujie, Takeyoshi Dohi, Ken ichi Sudo, and Shin ichi Takamoto. Development of MR compatible surgical manipulator toward a unified support system for diagnosis and treatment of heart disease. Lecture Notes in Computer Science, 2488:83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880083.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880083.pdf.

Fichtinger:2002:TPB

[1761] Gabor Fichtinger, Axel Krieger, Robert C. Susil, Attila Tanacs, Louis L. Whitcomb, and Ergin Atalar. Transrectal prostate biopsy inside closed MRI scanner with remote actuation, under real-time image guidance. Lecture Notes in Computer Science, 2488:91–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880091.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880091.pdf.

Kim:2002:NCM

[1762] Daeyoung Kim, Etsuko Kobayashi, Takeyoshi Dohi, and Ichiro Sakuma. A new, compact MR-compatible surgical manipulator for minimally invasive liver surgery. Lecture Notes in Computer Science, 2488: 99–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2488/24880099.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2488/24880099.
pdf.

Miyata:2002:MGF

[1763] Nobuhiko Miyata, Etsuko Kobayashi, Daeyoung Kim, Ken Masamune, Ichiro Sakuma, Naoki Yahagi, Takayuki Tsuji, Hiroshi Inada, Takeyoshi Dohi, Hiroshi Iseki, and Kintomo Takakura. Micro-grasping forceps manipulator for MR-guided neurosurgery. Lecture Notes in Computer Science, 2488:107–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880107.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880107.pdf.

Koseki:2002:EMT

[1764] Yoshihiko Koseki, Toshikatsu Washio, Kiyoyuki Chinzei, and Hiroshi Iseki. Endoscope manipulator for transnasal neurosurgery, optimized for and compatible to vertical field open MRI. Lecture Notes in Computer Science, 2488:114–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880114.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880114.pdf.

Hong:2002:MAN

[1765] Jae-Sung Hong, Takeyoshi Dohi, Makoto Hasizume, Kozo Konishi, and Nobuhiko Hata. A motion adaptable needle placement instrument based on tumor specific ultrasonic image segmentation. Lecture Notes in Computer Science, 2488:122–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880122.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880122.pdf.

Masuda:2002:EWT

[1766] Kohji Masuda, Norihiko Tateishi, Yasuyuki Suzuki, Eizen Kimura, Ying Wie, and Ken Ishihara. Experiment of wireless tele-echography system by controlling echographic diagnosis robot. Lecture Notes in Computer Science, 2488:130–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880130.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880130.pdf.

Vilchis:2002:ETT

[1767] Adriana Vilchis, Jocelyne Troccaz, Philippe Cinquin, Agnes Guerraz, Franck Pellisier, Pierre Thorel, Bertrand Tondu, Fabien Courrèges, Gérard Poisson, Marc Althuser, and Jean-Marc Ayoubi. Experiments with the TER tele-echography robot. Lecture Notes in Computer Science, 2488:138–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880138.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880138.pdf.

Gerovichev:2002:EVH

[1768] Oleg Gerovichev, Panadda Marayong, and Allison M. Okamura. The effect of visual and haptic feedback on manual and teleoperated needle insertion. Lecture Notes in Computer Science, 2488:147–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880147.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880147.pdf.

Kitagawa:2002:ASM

[1769] Masaya Kitagawa, Allison M. Okamura, Brian T. Bethea, Vincent L. Gott, and William A. Baumgartner. Analysis of suture manipulation forces for teleoperation with force feedback. Lecture Notes in Computer Science, 2488:155–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880155.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880155.pdf.

Ikuta:2002:RMS

[1770] Koji Ikuta, Keiji Sasaki, Keiichi Yamamoto, and Takayuki Shimada. Remote microsurgery system for deep and narrow space — development of new surgical procedure and micro-robotic tool. Lecture Notes in Computer Science, 2488:163–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880163.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880163.pdf.

Ikuta:2002:HFR

[1771] Koji Ikuta, Shinichi Daifu, Takahiko Hasegawa, and Humihiro Higashikawa. Hyper-finger for remote minimally invasive surgery in deep area. Lecture Notes in Computer Science, 2488: 173–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2488/24880173.htm;   http:

//link.springer.de/link/service/
series/0558/papers/2488/24880173.
pdf;   http://link.springer.de/
link/service/series/0558/papers/
2488/2488173a.pdf;   http://link.
springer.de/link/service/series/
0558/papers/2488/2488173b.pdf.

Ikuta:2002:SAC

[1772] Koji Ikuta, Hironobu Ichikawa, and Katsuya Suzuki. Safety-active catheter with multiple-segments driven by micro-hydraulic actuators. Lecture Notes in Computer Science, 2488: 182–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2488/24880182.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2488/24880182.
pdf.

Ohashi:2002:SCH

[1773] Kota Ohashi, Nobuhiko Hata, Tomoko Matsumura, Naoki Yahagi, Ichiro Sakuma, and Takeyoshi Dohi. A stem cell harvesting manipulator with flexible drilling unit for bone marrow transplantation. Lecture Notes in Computer Science, 2488:192–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880192.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880192.pdf.

Banovac:2002:LTB

[1774] Filip Banovac, Neil Glossop, David Lindisch, Daigo Tanaka, Elliot Levy, and Kevin Cleary. Liver tumor biopsy in a respiring phantom with the assistance of a novel electromagnetic navigation device. Lecture Notes in Computer Science, 2488:200–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880200.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880200.pdf.

Han:2002:NIM

[1775] Lianghao Han, Michael Burcher, and J. Alison Noble. Non-invasive measurement of biomechanical properties of in vivo soft tissues. Lecture Notes in Computer Science, 2488: 208–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2488/24880208.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2488/24880208.
pdf.

Kataoka:2002:MTF

[1776] Hiroyuki Kataoka, Toshikatsu Washio, Kiyoyuki Chinzei, Kazuyuki Mizuhara, Christina Simone, and Allison M. Okamura. Measurement of the tip and friction force acting on a needle during penetration. Lecture Notes in Computer Science, 2488: 216–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2488/24880216.htm;   http:
//link.springer.de/link/service/

series/0558/papers/2488/24880216.
pdf.

Hanafusa:2002:CFE

[1777] Akihiko Hanafusa, Tsuneshi Isomura, Yukio Sekiguchi, Hajime Takahashi, and Takeyoshi Dohi. Contact force evaluation of orthoses for the treatment of malformed ears. Lecture Notes in Computer Science, 2488:224–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880224.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880224.pdf.

Iyun:2002:CAC

[1778] O. Iyun, D. P. Borschneck, and R. E. Ellis. Computer-assisted correction of bone deformities using a 6-DOF parallel spatial mechanism. Lecture Notes in Computer Science, 2488:232–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880232.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880232.pdf.

Otake:2002:DDH

[1779] Yoshito Otake, Keisuke Hagio, Naoki Suzuki, Asaki Hattori, Nobuhiko Sugano, Kazuo Yonenobu, and Takahiro Ochi. Development of 4-dimensional human model system for the patient after total hip arthroplasty. Lecture Notes in Computer Science, 2488:241–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880241.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880241.pdf.

Tokuyasu:2002:DTS

[1780] Tatsushi Tokuyasu, Shin’ichiro Oota, Ken’ichi Asami, Tadashi Kitamura, Gen’ichi Sakaguchi, Tadaaki Koyama, and Masashi Komeda. Development of a training system for cardiac muscle palpation. Lecture Notes in Computer Science, 2488:248–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880248.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880248.pdf.

Jakopec:2002:PRE

[1781] Matjaž Jakopec, Simon J. Harris, Ferdinando Rodriguez y Baena, Paula Gomes, Justin Cobb, and Brian L. Davies. Preliminary results of an early clinical experience with the Acrobot^{T M} system for total knee replacement surgery. Lecture Notes in Computer Science, 2488:256–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880256.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880256.pdf.

Kimura:2002:PBT

[1782] Asako Kimura (Nishitai), Jon Camp, Richard Robb, and Brian Davis. A prostate brachytherapy training rehearsal system — simulation of deformable needle insertion. Lecture Notes in Computer Science, 2488:264–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880264.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880264.pdf.

Adhami:2002:VSC

[1783] Louai Adhami and Ève Coste-Manière.

A versatile system for computer integrated mini-invasive robotic surgery. Lecture Notes in Computer Science, 2488:272–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880272.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880272.pdf;   http://link.
springer.de/link/service/series/
0558/papers/2488/2488272a.pdf;
http://link.springer.de/link/service/
series/0558/papers/2488/2488272b.
pdf;   http://link.springer.de/
link/service/series/0558/papers/
2488/2488272c.pdf.

Bruyns:2002:MST

[1784] Cynthia Bruyns and Mark Ottensmeyer. Measurements of soft-tissue mechanical properties to support development of a physically based virtual animal model. Lecture Notes in Computer Science, 2488:282–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880282.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880282.pdf.

Cuadra:2002:VTM

[1785] M. Bach Cuadra, B. Platel, E. Solanas, T. Butz, and J.-Ph. Thiran. Validation of tissue modelization and classification techniques in T1-weighted MR brain images. Lecture Notes in Computer Science, 2488:290–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880290.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880290.pdf.

Warfield:2002:VIS

[1786] Simon K. Warfield, Kelly H. Zou, and William M. Wells. Validation of image segmentation and expert quality with an expectation-maximization algorithm. Lecture Notes in Computer Science, 2488:298–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880298.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880298.pdf.

Tanner:2002:VVP

[1787] C. Tanner, J. A. Schnabel, A. Degenhard, A. D. Castellano-Smith, C. Hayes, M. O. Leach, D. R. Hose, D. L. G. Hill, and D. J. Hawkes. Validation of volume-preserving nonrigid registration: Application to contrast-enhanced MR-mammography. Lecture Notes in Computer Science, 2488:307–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880307.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880307.pdf.

Zou:2002:SVA

[1788] Kelly H. Zou, William M. Wells III, Michael R. Kaus, Ron Kikinis, Ferenc A. Jolesz, and Simon K. Warfield. Statistical validation of automated probabilistic segmentation against composite latent expert ground truth in MR imaging of brain tumors. Lecture Notes in Computer Science, 2488:315–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880315.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880315.pdf.

Bardinet:2002:PVP

[1789] É. Bardinet, P. Cathier, A. Roche, N. Ayache, and D. Dormont. A posteriori validation of pre-operative planning in functional neurosurgery by quantification of brain pneumocephalus. Lecture Notes in Computer Science, 2488:323–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880323.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2488/24880323.pdf.

Ma:2002:ATA

[1790] B. Ma, J. F. Rudan, and R. E. Ellis. Affine transformations and atlases: Assessing a new navigation tool for knee arthroplasty. Lecture Notes in Computer Science, 2488:331–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880331.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880331.pdf.

Hagio:2002:ERS

[1791] Keisuke Hagio, Nobuhiko Sugano, Masaki Takashina, Takashi Nishii, Hideki Yoshikawa, and Takahiro Ochi. Effectiveness of the ROBODOC system during total hip arthroplasty in preventing intraoperative pulmonary embolism. Lecture Notes in Computer Science, 2488:339–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880339.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880339.pdf.

Chen:2002:MIS

[1792] James Z. Chen, Stephen M. Pizer, Edward L. Chaney, and Sarang Joshi. Medical image synthesis via Monte Carlo simulation. Lecture Notes in Computer Science, 2488:347–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880347.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880347.pdf.

Timoner:2002:PIS

[1793] Samson J. Timoner, Pollina Golland, Ron Kikinis, Martha E. Shenton, W. Eric L. Grimson, and William M. Wells III. Performance issues in shape classification. Lecture Notes in Computer Science, 2488:355–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880355.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880355.pdf.

Prima:2002:SAL

[1794] Sylvain Prima, Nicholas Ayache, Andrew Janke, Simon J. Francis, Douglas L. Arnold, and D. Louis Collins. Statistical analysis of longitudinal MRI data: Applications for detection of disease activity in MS. Lecture Notes in Computer Science, 2488:363–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880363.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880363.pdf.

Moon:2002:ABT

[1795] Nathan Moon, Elizabeth Bullitt, Koen van Leemput, and Guido Gerig. Automatic brain and tumor segmentation. Lecture Notes in Computer Science, 2488:372–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880372.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880372.pdf.

Cuadra:2002:ABS

[1796] M. Bach Cuadra, J. Gomez, P. Hagmann, C. Pollo, J.-G. Villemure, B. M. Dawant, and J.-Ph. Thiran. Atlas-based segmentation of pathological brains using a model of tumor growth. Lecture Notes in Computer Science, 2488:380–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880380.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880380.pdf.

Gering:2002:RDN

[1797] David T. Gering, W. Eric L. Grimson, and R. Kikinis. Recognizing deviations from normalcy for brain tumor segmentation. Lecture Notes in Computer Science, 2488:388–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880388.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880388.pdf.

Hastreiter:2002:VRN

[1798] P. Hastreiter, R. Naraghi, B. Tomandl, M. Bauer, and R. Fahlbusch. 3Dvisualization and registration for neurovascular compression syndrome analysis. Lecture Notes in Computer Science, 2488:396–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880396.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880396.pdf.

Baert:2002:GWR

[1799] S. A. M. Baert, E. B. van der Kraats, and W. J. Niessen. 3D guide wire reconstruction from biplane image sequences for 3D navigation in endovascular interventions. Lecture Notes in Computer Science, 2488:404–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880404.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880404.pdf.

Iserhardt-Bauer:2002:SAI

[1800] S. Iserhardt-Bauer, P. Hastreiter, B. Tomandl, N. Köstner, M. Schempershofe, U. Nissen, and T. Ertl. Standardized analysis of intracranial aneurysms using digital video sequences. Lecture Notes in Computer Science, 2488:411–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880411.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2488/24880411.pdf.

McLaughlin:2002:DAU

[1801] Robert A. McLaughlin and J. Alison Noble. Demarcation of aneurysms using the seed and cull algorithm. Lecture Notes in Computer Science, 2488:419–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880419.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880419.pdf.

Cachia:2002:GPC

[1802] A. Cachia, J.-F. Mangin, D. Rivière, D. Papadopoulos-Orfanos, I. Bloch, and J. Régis. Gyral parcellation of the cortical surface using geodesic voronoï diagrams. Lecture Notes in Computer Science, 2488:427–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880427.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880427.pdf.

Bjornemo:2002:RSW

[1803] Mats Björnemo, Anders Brun, Ron Kikinis, and Carl-Fredrik Westin. Regularized stochastic white matter tractography using diffusion tensor MRI. Lecture Notes in Computer Science, 2488:435–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880435.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880435.pdf.

Hutton:2002:SSC

[1804] Chloe Hutton, Enrico De Vita, and Robert Turner. Sulcal segmentation for cortical thickness measurements. Lecture Notes in Computer Science, 2488:443–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880443.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880443.pdf.

Jaume:2002:LBS

[1805] Sylvain Jaume, Benoît Macq, and Simon K. Warfield. Labeling the brain surface using a deformable multiresolution mesh. Lecture Notes in Computer Science, 2488:451–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880451.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880451.pdf.

ODonnell:2002:NAE

[1806] Lauren O’Donnell, Steven Haker, and Carl-Fredrik Westin. New approaches to estimation of white matter connectivity in diffusion tensor MRI: Elliptic PDEs and geodesics in a tensor-warped space. Lecture Notes in Computer Science, 2488:459–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2488/24880459.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880459.pdf.

Flandin:2002:IDS

[1807] Guillaume Flandin, Ferath Kherif, Xavier Pennec, Grégoire Malandain, Nicholas Ayache, and Jean-Baptiste Poline. Improved detection sensitivity in functional MRI data using a brain parcelling technique. Lecture Notes in Computer Science, 2488:467–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880467.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880467.pdf.

Cointepas:2002:SGB

[1808] Y. Cointepas, C. Poupon, D. Le Bihan, and J.-F. Mangin. A spin glass based framework to untangle fiber crossing in MR diffusion based tracking. Lecture Notes in Computer Science, 2488:475–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880475.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880475.pdf.

Sturm:2002:AAL

[1809] Bernhard Sturm, Dominik Meier, and Elizabeth Fisher. Automated approximation of lateral ventricular shape in magnetic resonance images of multiple sclerosis patients.

Lecture Notes in Computer Science, 2488:483–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880483.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880483.pdf.

Studholme:2002:ICA

[1810] C. Studholme, V. Cardenas, A. Maudsley, and M. Weiner. An intensity consistent approach to the cross sectional analysis of deformation tensor derived maps of brain shape. Lecture Notes in Computer Science, 2488:492–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880492.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880492.pdf.

Aubert-Broche:2002:DIH

[1811] Bérengère Aubert-Broche, Christophe Grova, Pierre Jannin, Irène Buvat, Habib Benali, and Bernard Gibaud. Detection of inter-hemispheric asymmetries of brain perfusion in SPECT. Lecture Notes in Computer Science, 2488:500–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880500.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880500.pdf.

Golland:2002:DAI

[1812] Polina Golland, Bruce Fischl, Mona Spiridon, Nancy Kanwisher, Randy L. Buckner, Martha E. Shenton, Ron Kikinis, Anders Dale, and W. Eric L. Grimson. Discriminative analysis for image-based studies. Lecture Notes in Computer Science, 2488: 508–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2488/24880508.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2488/24880508.
                                        pdf.

Cocosco:2002:AGT

[1813] Chris A. Cocosco, Alex P. Zijdenbos, and Alan C. Evans. Automatic generation of training data for brain tissue classification from MRI. Lecture Notes in Computer Science, 2488:516–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880516.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880516.pdf.

Hojjat:2002:PIG

[1814] A. Hojjat, D. Collie, and A. C. F. Colchester. The putamen intensity gradient in CJD diagnosis. Lecture Notes in Computer Science, 2488:524–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880524.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2488/24880524.pdf.

Hill:2002:DBA

[1815] D. L. G. Hill, J. V. Hajnal, D. Rueckert, S. M. Smith, T. Hartkens, and K. McLeish. A dynamic brain atlas. Lecture Notes in Computer Science, 2488:532–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880532.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880532.pdf.

Koikkalainen:2002:MLD

[1816] Juha Koikkalainen and Jyrki Lötjönen.

Model library for deformable modelbased segmentation of 3-D brain MRimages. Lecture Notes in Computer Science, 2488:540–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880540.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880540.pdf.

Bardinet:2002:CRH

[1817] É. Bardinet, S. Ourselin, D. Dormont, G. Malandain, D. Tandé, K. Parain, N. Ayache, and J. Yelnik. Co-registration of histological, optical and MR data of the human brain. Lecture Notes in Computer Science, 2488:548–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880548.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880548.pdf.

Tsagaan:2002:ASM

[1818] Baigalmaa Tsagaan, Akinobu Shimizu, Hidefumi Kobatake, and Kunihisa Miyakawa. An automated segmentation method of kidney using statistical information. Lecture Notes in Computer Science, 2488: 556–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2488/24880556.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2488/24880556.
                                        pdf.

Pohl:2002:INR

[1819] Kilian M. Pohl, William M. Wells, Alexandre Guimond, Kiyoto Kasai, Martha E. Shenton, Ron Kikinis, W. Eric L. Grimson, and Simon K. Warfield. Incorporating non-rigid registration into expectation maximization algorithm to segment MR images. Lecture Notes in Computer Science, 2488:564–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880564.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880564.pdf.

Tek:2002:SMS

[1820] Hüseyin Tek, Martin Bergtholdt, Dorin Comaniciu, and James Williams. Segmentation of 3D medical structures using robust ray propagation. Lecture Notes in Computer Science, 2488:572–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880572.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880572.pdf.

Wyatt:2002:MMJ

[1821] Paul P. Wyatt and J. Alison Noble.

MAP MRF joint segmentation and registration. Lecture Notes in Computer Science, 2488:580–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880580.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880580.pdf.

Yang:2002:SND

[1822] Jing Yang, Lawrence H. Staib, and James S. Duncan. Statistical neighbor distance influence in active contours. Lecture Notes in Computer Science, 2488:588–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880588.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880588.pdf.

Lapeer:2002:AWC

[1823] R. J. Lapeer, A. C. Tan, and R. Aldridge. Active watersheds: Combining 3D watershed segmentation and active contours to extract abdominal organs from MR images. Lecture Notes in Computer Science, 2488:596–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880596.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880596.pdf.

Wink:2002:CIP

[1824] O. Wink, R. Kemkers, S. J. Chen, and J. D. Carroll. Coronary intervention planning using hybrid 3D reconstruction. Lecture Notes in Computer Science, 2488:604–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880604.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880604.pdf.

Gao:2002:DMB

[1825] Jianxin Gao, Nick Ablitt, Andrew Elkington, and Guang-Zhong Yang. Deformation modelling based on PLSR for cardiac magnetic resonance perfusion imaging. Lecture Notes in Computer Science, 2488:612–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880612.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880612.pdf.

Montillo:2002:ASL

[1826] Albert Montillo, Dimitris Metaxas, and Leon Axel. Automated segmentation of the left and right ventricles in 4D cardiac SPAMM images. Lecture Notes in Computer Science, 2488:620–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880620.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880620.pdf.

Shi:2002:SFE

[1827] Pengcheng Shi and Huafeng Liu.

Stochastic finite element framework for cardiac kinematics function and material property analysis. Lecture Notes in Computer Science, 2488:634–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880634.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880634.pdf.

Lorenzo-Valdes:2002:ABS

[1828] M. Lorenzo-Valdés, G. I. Sanchez-Ortiz, R. Mohiaddin, and D. Rueckert. Atlas-based segmentation and tracking of 3D cardiac MR images using non-rigid registration. Lecture Notes in Computer Science, 2488:642–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880642.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880642.pdf.

Noble:2002:MDR

[1829] Nicholas M. I. Noble, Derek L. G. Hill, Marcel Breeuwer, Julia A. Schnabel, David J. Hawkes, Frans A. Gerritsen, and Reza Razavi. Myocardial delineation via registration in a polar coordinate system. Lecture Notes in Computer Science, 2488:651–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880651.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880651.pdf.

Bansal:2002:IIR

[1830] Ravi Bansal and Gareth Funka-Lea. Integrated image registration for cardiac MR perfusion data. Lecture Notes in Computer Science, 2488:659–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880659.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880659.pdf.

Yezzi:2002:ASC

[1831] Anthony Yezzi and Allen Tannenbaum. 4D active surfaces for cardiac analysis. Lecture Notes in Computer Science, 2488:667–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880667.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880667.pdf.

Fukumoto:2002:CDS

[1832] Ichiro Fukumoto. A computer diagnosing system of dementia using smooth pursuit oculogyration. Lecture Notes in Computer Science, 2488:674–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880674.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880674.pdf.

Lin:2002:CMS

[1833] Ning Lin, Weichuan Yu, and James S.

Duncan. Combinative multi-scale level set framework for echocardiographic image segmentation. Lecture Notes in Computer Science, 2488: 682–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2488/24880682.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2488/24880682.
pdf.

Pednekar:2002:AHS

[1834] A. Pednekar, I. A. Kakadiaris, V. Zavaletta, R. Muthupillai, and S. Flamm. Automatic hybrid segmentation of dual contrast cardiac MR data. Lecture Notes in Computer Science, 2488:690–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880690.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880690.pdf.

Noe:2002:EPV

[1835] Aljaž Noe and James C. Gee. Efficient partial volume tissue classification in MRI scans. Lecture Notes in Computer Science, 2488:698–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880698.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880698.pdf.

Hu:2002:VSS

[1836] Zhenhua Hu, Dimitris Metaxas, and Leon Axel. In-vivo strain and stress estimation of the left ventricle from MRI images. Lecture Notes in Computer Science, 2488:706–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880706.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880706.pdf.

Sermesant:2002:BMC

[1837] M. Sermesant, C. Forest, X. Pennec, H. Delingette, and N. Ayache. Biomechanical model construction from different modalities: Application to cardiac images. Lecture Notes in Computer Science, 2488:714–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880714.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880714.pdf.

Rao:2002:CCM

[1838] A. Rao, G. I. Sanchez-Ortiz, R. Chandrashekara, M. Lorenzo-Valdés, R. Mohiaddin, and D. Rueckert. Comparison of cardiac motion across subjects using non-rigid registration. Lecture Notes in Computer Science, 2488:722–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880722.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880722.pdf.

Claridge:2002:CTH

[1839] Ela Claridge, Symon Cotton, Per Hall, and Marc Moncrieff. From colour to tissue histology: Physics based interpretation of images of pigmented skin lesions. Lecture Notes in Computer Science, 2488:730–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880730.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880730.pdf.

Ntziachristos:2002:VMI

[1840] Vasilis Ntziachristos, Jorge Ripoll, Edward Graves, and Ralph Weissleder. In-vivo molecular investigations of live tissues using diffracting sources. Lecture Notes in Computer Science, 2488:739–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880739.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2488/24880739.pdf.

Wei:2002:ADN

[1841] Guo-Qing Wei, Li Fan, and JianZhong Qian. Automatic detection of nodules attached to vessels in lung CT by volume projection analysis. Lecture Notes in Computer Science, 2488:746–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880746.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880746.pdf.

Park:2002:LRS

[1842] Kyoungju Park, Dimitris N. Metaxas, and Leon Axel. LV-RV shape modeling based on a blended parameterized model. Lecture Notes in Computer Science, 2488:753–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880753.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880753.pdf.

Gee:2002:CRP

[1843] James Gee, Tessa Sundaram, Ichiro Hasegawa, Hidemasa Uematsu, and Hiroto Hatabu. Characterization of regional pulmonary mechanics from serial MRI data. Lecture Notes in Computer Science, 2488:762–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880762.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880762.pdf.

Lin:2002:UVB

[1844] Michael P. Lin, Christian Devita, James C. Gee, and Murray Grossman. Using voxel-based morphometry to examine atrophy-behavior correlates in Alzheimer’s disease and frontotemporal dementia. Lecture Notes in Computer Science, 2488:770–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880770.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880770.pdf.

Vermeer:2002:DWS

[1845] Koen Vermeer, Frans Vos, Hans Lemij, and Albert Vossepoel. Detecting wedge shaped defects in polarimetric images of the retinal nerve fiber layer. Lecture Notes in Computer Science, 2488:777–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2488/24880777.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2488/24880777.pdf.

Guimond:2002:ASI

[1846] Alexandre Guimond, Svetlana Egorova, Ronald J. Killiany, Marilyn S. Albert, and Charles R. G. Guttmann. Automatic statistical identification of neuroanatomical abnormalities between different populations. Lecture Notes in Computer Science, 2488:

785–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2488/24880785.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2488/24880785.
                                        pdf.

Kawata:2002:EBA

[1847] Yoshiki Kawata, Noboru Niki, Hironobu Ohmatsu, and Noriyuki Moriyama. Example-based assisting approach for pulmonary nodule classification in 3-D thoracic CT images. Lecture Notes in Computer Science, 2488:793–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2488/24880793.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2488/24880793.pdf.

Anonymous:2002:AIbp

[1848] Anonymous. Author index. Lecture Notes in Computer Science, 2488: 801–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2488/2488auth.pdf.

Kitaoka:2002:ANL

[1849] Hiroko Kitaoka, Yongsup Park, Juerg Tschirren, Joseph Reinhardt, Milan Sonka, Goeffrey McLennan, and Eric A. Hoffman. Automated nomenclature labeling of the bronchial tree in 3D-CT lung images. Lecture Notes in Computer Science, 2489:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890001.pdf.

Tschirren:2002:SSB

[1850] J. Tschirren, K. Palágyi, J. M.

Reinhardt, E. A. Hoffman, and M. Sonka. Segmentation, skeletonization, and branchpoint matching — A fully automated quantitative evaluation of human intrathoracic airway trees. Lecture Notes in Computer Science, 2489:12–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890012.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890012.pdf.

Harders:2002:IVE

[1851] Matthias Harders, Simon Wildermuth, Dominik Weishaupt, and Gábor Székely. Improving virtual endoscopy for the intestinal tract. Lecture Notes in Computer Science, 2489:20–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890020.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890020.pdf.

Yasuba:2002:FNC

[1852] Chihiro Yasuba, Syoji Kobashi, Katsuya Kondo, Yutaka Hata, Seturo Imawaki, and Makoto Ishikawa. Finding a non-continuous tube by fuzzy inference for segmenting the MR cholangiography inage. Lecture Notes in Computer Science, 2489: 28–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2489/24890028.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2489/24890028.
pdf.

vanBemmel:2002:LSB

[1853] C. M. van Bemmel, L. J. Spreeuwers, M. A. Viergever, and W. J. Niessen. Level-set based carotid artery segmentation for stenosis grading. Lecture Notes in Computer Science, 2489:36–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890036.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890036.pdf.

Figl:2002:PBC

[1854] Michael Figl, Wolfgang Birkfellner, Franz Watzinger, Felix Wanschitz, Johann Hummel, Rudolf Hanel, Rolf Ewers, and Helmar Bergmann. PCbased control unit for a head mounted operating microscope for augmented reality visualization in surgical navigation. Lecture Notes in Computer Science, 2489:44–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890044.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890044.pdf.

Morikawa:2002:TDM

[1855] Shigehiro Morikawa, Toshiro Inubushi, Yoshimasa Kurumi, Shigeyuki Naka, Koichiro Sato, Tohru Tani, Nobuhiko Hata, Viswanathan Seshan, and Hasnine A. Haque. Technical developments for MR-guided microwave thermocoagulation therapy of liver tumors. Lecture Notes in Computer Science, 2489:52–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890052.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890052.pdf.

Livyatan:2002:RAC

[1856] H. Livyatan, Z. Yaniv, and L. Joskowicz. Robust automatic C-arm calibration for fluoroscopy-based navigation: a practical approach. Lecture Notes in Computer Science, 2489: 60–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2489/24890060.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2489/24890060.
pdf.

Finnis:2002:APB

[1857] Kirk W. Finnis, Yves P. Starreveld, Andrew G. Parrent, Abbas F. Sadikot, and Terry M. Peters. Application of a population based electrophysiological database to the planning and guidance of deep brain stereotactic neurosurgery. Lecture Notes in Computer Science, 2489:69–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890069.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890069.pdf.

Masamune:2002:IOS

[1858] Ken Masamune, Gabor Fichtinger, Anton Deguet, Daisuke Matsuka, and Russell Taylor. An image overlay system with enhanced reality for percutaneous therapy performed inside CT scanner. Lecture Notes in Computer Science, 2489:77–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890077.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890077.pdf.

Liao:2002:HRS

[1859] Hongen Liao, Nobuhiko Hata, Makoto Iwahara, Susumu Nakajima, Ichiro Sakuma, and Takeyoshi Dohi. Highresolution stereoscopic surgical display using parallel integral videography and multi-projector. Lecture Notes in Computer Science, 2489: 85–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2489/24890085.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2489/24890085.
pdf.

Take:2002:TDD

[1860] Noriyuki Take, Yukio Kosugi, and Toshimitsu Musha. Three-dimensional display for multi-sourced activities and their relations in the human brain by information flow between estimated dipoles. Lecture Notes in Computer Science, 2489:93–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890093.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890093.pdf.

Baert:2002:GWT

[1861] S. A. M. Baert and W. J. Niessen.

2D guide wire tracking during endovascular interventions. Lecture Notes in Computer Science, 2489: 101–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2489/24890101.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2489/24890101.
pdf.

Furushiro:2002:SMS

[1862] Naomichi Furushiro, Tomoharu Saito, Yoshitaka Masutani, and Ichiro Sakuma. Specification method of surface measurement for surgical navigation: Ridgeline based organ registration. Lecture Notes in Computer Science, 2489:109–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890109.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2489/24890109.pdf.

Sauer:2002:ARN

[1863] F. Sauer, A. Khamene, and S. Vogt. An augmented reality navigation system with a single-camera tracker: System design and needle biopsy phantom trial. Lecture Notes in Computer Science, 2489:116–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2489/24890116.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2489/24890116.pdf.

Sasama:2002:NLG

[1864] Toshihiko Sasama, Nobuhiko Sugano, Yoshinobu Sato, Yasuyuki Momoi, Tsuyoshi Koyama, Yoshikazu Nakajima, Ichiro Sakuma, Masakatsu Fujie, Kazuo Yonenobu, Takahiro Ochi, and Shinichi Tamura. A novel laser guidance system for alignment of linear surgical tools: Its principles and performance evaluation as a man-machine system. Lecture Notes in Computer Science, 2489:125–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2489/24890125.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2489/24890125.pdf.

Sakuma:2002:NHI

[1865] Ichiro Sakuma, Yuichi Takai, Etsuko Kobayashi, Hiroshi Inada, Katsuhiko Fujimoto, and Tekehide Asano. Navigation of high intensity focused ultrasound applicator with an integrated three-dimensional ultrasound imaging system. Lecture Notes in Computer Science, 2489:133–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890133.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890133.pdf.

Ourselin:2002:RRM

[1866] Sébastien Ourselin, Radu Stefanescu, and Xavier Pennec. Robust registration of multi-modal images: Towards real-time clinical applications. Lecture Notes in Computer Science, 2489:140–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890140.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890140.pdf.

Nakamoto:2002:USU

[1867] Masahiko Nakamoto, Yoshinobu Sato, Masaki Miyamoto, Yoshikazu Nakamjima, Kozo Konishi, Mitsuo Shimada, Makoto Hashizume, and Shinichi Tamura. 3D ultrasound system using a magneto-optic hybrid tracker for augmented reality visualization in laparoscopic liver surgery. Lecture Notes in Computer Science, 2489:148–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890148.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2489/24890148.pdf.

Gobbi:2002:IIO

[1868] David G. Gobbi and Terry M. Peters. Interactive intra-operative 3D ultrasound reconstruction and visualization. Lecture Notes in Computer Science, 2489:156–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2489/24890156.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2489/24890156.pdf.

Hata:2002:PPM

[1869] Nobuhiko Hata, Junichi Tokuda, Shigeo Morikawa, and Takeyoshi Dohi. Projection profile matching for intraoperative MRI registration embedded in MR imaging sequence. Lecture Notes in Computer Science, 2489:164–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2489/24890164.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2489/24890164.pdf.

Megali:2002:NTS

[1870] Giuseppe Megali, Oliver Tonet, Marcello Mazzoni, Paolo Dario, Alberto Vascellari, and Maurilio Marcacci. A new tool for surgical training in knee arthroscopy. Lecture Notes in Computer Science, 2489:170–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890170.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890170.pdf.

Nakao:2002:CVS

[1871] Megumi Nakao, Tomohiro Kuroda, Hiroshi Oyama, Masaru Komori, Tetsuya Matsuda, and Takashi Takahashi. Combining volumetric soft tissue cuts for interventional surgery simulation. Lecture Notes in Computer Science, 2489:178–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890178.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890178.pdf.

Tiede:2002:VEU

[1872] Ulf Tiede, Norman von Sternberg-Gospos, Paul Steiner, and Karl Heinz Höhne. Virtual endoscopy using cubic QuickTime-VR panorama views. Lecture Notes in Computer Science, 2489:186–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890186.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890186.pdf.

Chihoub:2002:HLS

[1873] A. Chihoub. High level simulation & modeling for medical applications — ultrasound case. Lecture Notes in Computer Science, 2489:193–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2489/24890193.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2489/24890193.pdf.

Sierra:2002:GPS

[1874] Raimundo Sierra, Gabor Székely, and Michael Bajka. Generation of pathologies for surgical training simulators. Lecture Notes in Computer Science, 2489:202–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2489/24890202.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2489/24890202.pdf.

Aharon:2002:CDA

[1875] Shmuel Aharon and Christophe Lenglet. Collision detection algorithm for deformable objects using OpenGL. Lecture Notes in Computer Science, 2489:211–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2489/24890211.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2489/24890211.pdf.

Paloc:2002:OMV

[1876] Celine Paloc, Fernando Bello, Richard I. Kitney, and Ara Darzi. Online multiresolution volumetric mass spring model for real time soft tissue deformation. Lecture Notes in Computer Science, 2489:219–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890219.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890219.pdf.

Hanafusa:2002:ODS

[1877] Akihiko Hanafusa, Tsuneshi Isomura, Yukio Sekiguchi, Hajime Takahashi, and Takeyoshi Dohi. Orthosis design system for malformed ears based on spline approximation. Lecture Notes in Computer Science, 2489:227–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890227.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890227.pdf.

Forest:2002:CSM

[1878] C. Forest, H. Delingette, and N. Ayache. Cutting simulation of manifold volumetric meshes. Lecture Notes in Computer Science, 2489: 235–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2489/24890235.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2489/24890235.
pdf.

Alderliesten:2002:SGW

[1879] Tanja Alderliesten, Maurits K. Konings, and Wiro J. Niessen. Simulation of guide wire propagation for minimally invasive vascular interventions. Lecture Notes in Computer Science, 2489:245–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890245.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890245.pdf.

DiMaio:2002:NIM

[1880] S. P. DiMaio and S. E. Salcudean. Needle insertion modelling for the interactive simulation of percutaneous procedures. Lecture Notes in Computer Science, 2489:253–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890253.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890253.pdf.

Kawakami:2002:AAL

[1881] Hideo Kawakami, Nobuhiko Sugano, Takashi Nagaoka, Keisuke Hagio, Kazuo Yonenobu, Hideki Yoshikawa, Takahiro Ochi, Asaki Hattori, and Naoki Suzuki. 3D analysis of the alignment of the lower extremity in high tibial osteotomy. Lecture Notes in Computer Science, 2489:261–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890261.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890261.pdf.

Lehmann:2002:SIO

[1882] G. Lehmann, D. Habets, D. W.

Holdsworth, T. Peters, and M. Drangova. Simulation of intra-operative 3D coronary angiography for enhanced minimally invasive robotic cardiac intervention. Lecture Notes in Computer Science, 2489:268–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890268.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890268.pdf.

Martelli:2002:CIA

[1883] S. Martelli and A. Visani. Computer investigation into the anatomical location of the axes of rotation in the normal knee. Lecture Notes in Computer Science, 2489:276–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890276.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890276.pdf.

Szczerba:2002:MMV

[1884] Dominik Szczerba and Gábor Székely.

Macroscopic modeling of vascular systems. Lecture Notes in Computer Science, 2489:284–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890284.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890284.pdf.

Manduca:2002:STD

[1885] Armando Manduca, David S. Lake, and Richard L. Ehman. Spatiotemporal directional filtering for improved inversion of MR elastography images. Lecture Notes in Computer Science, 2489:293–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890293.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890293.pdf.

Masutani:2002:RBR

[1886] Yoshitaka Masutani. RBF-based representation of volumetric data: Application in visualization and segmentation. Lecture Notes in Computer Science, 2489:300–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890300.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890300.pdf.

Martelli:2002:AMK

[1887] S. Martelli, F. Acquaroli, V. Pinskerova, A. Spettol, and A. Visani. An anatomical model of the knee joint obtained by computer dissection. Lecture Notes in Computer Science, 2489:308–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890308.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890308.pdf.

Chabanas:2002:MPS

[1888] Matthieu Chabanas, Christophe Marecaux, Yohan Payan, and Franck Boutault. Models for planning and simulation in computer assisted orthognatic surgery. Lecture Notes in Computer Science, 2489: 315–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2489/24890315.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2489/24890315.
pdf.

Luboz:2002:SER

[1889] Vincent Luboz, Annaig Pedrono, Pascal Swider, Frank Boutault, and Yohan Payan. Simulation of the exophthalmia reduction using a finite element model of the orbital soft tissues. Lecture Notes in Computer Science, 2489:323–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890323.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890323.pdf.

Kukuk:2002:RTD

[1890] Markus Kukuk and Bernhard Geiger.

A real-time deformable model for flexible instruments inserted into tubular structures. Lecture Notes in Computer Science, 2489:331–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890331.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2489/24890331.pdf.

Li:2002:MHO

[1891] Zirui Li, Chee-Kong Chui, Yiyu Cai, Shantha Amrith, Poh-Sun Goh, James H. Anderson, Jeremy Teo, Cherine Liu, Irma Kusuma, Yee-Shin Siow, and Wieslaw L. Nowinski. Modeling of the human orbit from MR images. Lecture Notes in Computer Science, 2489:339–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2489/24890339.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2489/24890339.pdf.

deBruin:2002:AHQ

[1892] P. W. de Bruin, P. M. van Meeteren, F. M. Vos, A. M. Vossepoel, and F. H. Post. Accurate and high quality triangle models from 3D grey scale images. Lecture Notes in Computer Science, 2489:348–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2489/24890348.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2489/24890348.pdf.

Hayashibe:2002:IFS

[1893] Mitsuhiro Hayashibe, Naoki Suzuki, Asaki Hattori, and Yoshihiko Nakamura. Intraoperative fast 3D shape recovery of abdominal organs in laparoscopy. Lecture Notes in Computer Science, 2489:356–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890356.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890356.pdf.

Fleute:2002:IAM

[1894] M. Fleute, S. Lavallée, and L. Desbat. Integrated approach for matching statistical shape models with intra-operative 2D and 3D data. Lecture Notes in Computer Science, 2489:364–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890364.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890364.pdf.

Paulsen:2002:BTS

[1895] Rasmus Paulsen, Rasmus Larsen, Claus Nielsen, Søren Laugesen, and Bjarne Ersbøll. Building and testing a statistical shape model of the human ear canal. Lecture Notes in Computer Science, 2489:373–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890373.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890373.pdf.

Dubb:2002:SCC

[1896] Abraham Dubb, Brian Avants, Ruben Gur, and James Gee. Shape characterization of the corpus callosum in schizophrenia using template deformation. Lecture Notes in Computer Science, 2489:381–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2489/24890381.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2489/24890381.pdf.

Fan:2002:PSD

[1897] Shao Fan, Ling Keck Voon, and Ng Wan Sing. 3D prostate surface detection from ultrasound images based on level set method. Lecture Notes in Computer Science, 2489:389–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2489/24890389.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2489/24890389.pdf.

Martin:2002:BAV

[1898] Marcos Martín and Carlos Alberola. A Bayesian approach to in vivo kidney ultrasound contour detection using Markov random fields. Lecture Notes in Computer Science, 2489:397–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2489/24890397.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2489/24890397.pdf.

Watanabe:2002:LSB

[1899] Masao Watanabe, Ron Kikinis, and Carl-Fredrik Westin. Level set based integration of segmentation and computational fluid dynamics for flow correction in phase contrast angiography. Lecture Notes in Computer Science, 2489:405–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890405.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890405.pdf.

Osareh:2002:CEC

[1900] Alireza Osareh, Majid Mirmehdi, Barry Thomas, and Richard Markham. Comparative exudate classification using support vector machines and neural networks. Lecture Notes in Computer Science, 2489:413–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890413.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890413.pdf.

Lamecker:2002:SSM

[1901] Hans Lamecker, Thomas Lange, and Martin Seebass. A statistical shape model for the liver. Lecture Notes in Computer Science, 2489:421–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890421.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890421.pdf.

Larsen:2002:SSA

[1902] Rasmus Larsen, Klaus Baggesen Hilger, and Mark C. Wrobel. Statistical 2D and 3D shape analysis using non-euclidean metrics. Lecture Notes in Computer Science, 2489:428–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890428.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890428.pdf.

Vohra:2002:KFS

[1903] N. Vohra, B. C. Vemuri, A. Rangarajan, R. L. Gilmore, S. N. Roper, and C. M. Leonard. Kernel Fisher for shape based classification in epilepsy. Lecture Notes in Computer Science, 2489:436–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890436.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890436.pdf.

Hilger:2002:NRS

[1904] Klaus B. Hilger, Mikkel B. Stegmann, and Rasmus Larsen. A noise robust statistical texture model. Lecture Notes in Computer Science, 2489:444–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890444.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890444.pdf.

Mohamed:2002:CSB

[1905] Ashraf Mohamed, Christos Davatzikos, and Russell Taylor. A combined statistical and biomechanical model for estimation of intra-operative prostate deformation. Lecture Notes in Computer Science, 2489:452–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890452.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890452.pdf.

Tomazevic:2002:GSR

[1906] Dejan Tomaževič, Boštjan Likar, and Franjo Pernuš. “Gold Standard” 2D/ 3D registration of X-ray to CT and MR images. Lecture Notes in Computer Science, 2489:461–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890461.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890461.pdf.

Rohlfing:2002:NIS

[1907] Torsten Rohlfing and Calvin R. Maurer, Jr. A novel image similarity measure for registration of 3-D MR images and X-ray projection images. Lecture Notes in Computer Science, 2489:469–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890469.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890469.pdf.

Imamura:2002:RPC

[1908] Hiroshi Imamura, Noriaki Ida, Naozo Sugimoto, Shigeru Eiho, Shin ichi Urayama, Katsuya Ueno, and Kanji Inoue. Registration of preoperative CTA and intraoperative fluoroscopic images for assisting aortic stent grafting. Lecture Notes in Computer Science, 2489:477–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2489/24890477.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2489/24890477.pdf.

Nakajima:2002:PAO

[1909] Yoshikazu Nakajima, Yuichi Tamura, Yoshinobu Sato, Takahito Tashiro, Nobuhiko Sugano, Kazuo Yonenobu, Hideki Yoshikawa, Takahiro Ochi, and Shinichi Tamura. Preoperative analysis of optimal imaging orientation in fluoroscopy for voxel-based 2-D/3-D registration. Lecture Notes in Computer Science, 2489:485–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2489/24890485.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2489/24890485.pdf.

Masumoto:2002:NSM

[1910] Jun Masumoto, Yoshinobu Sato, Masatoshi Hori, Takamichi Murakami, Takeshi Johkoh, Hironobu Nakamura, and Shinichi Tamura. A new similarity measure for nonrigid volume registration using known joint distribution of target tissue: Application to dynamic CT data of the liver. Lecture Notes in Computer Science, 2489:493–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890493.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890493.pdf.

Hipwell:2002:IBR

[1911] John H. Hipwell, Graeme P. Penney, Tim C. Cox, James V. Byrne, and David J. Hawkes. 2D-3D intensity based registration of DSA and MRA — A comparison of similarity measures. Lecture Notes in Computer Science, 2489:501–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890501.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890501.pdf.

Kherif:2002:MBS

[1912] Ferath Kherif, Guillaume Flandin, Philippe Ciuciu, Habib Benali, Olivier Simon, and Jean-Baptiste Poline. Model based spatial and temporal similarity measures between series of functional magnetic resonance images. Lecture Notes in Computer Science, 2489:509–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890509.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890509.pdf.

McLaughlin:2002:CIB

[1913] Robert A. McLaughlin, John Hipwell, David J. Hawkes, J. Alison Noble, James V. Byrne, and Tim Cox. A comparison of 2D-3D intensity-based registration and feature-based registration for neurointerventions. Lecture Notes in Computer Science, 2489:517–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2489/24890517.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2489/24890517.pdf.

Chung:2002:MMI

[1914] Albert C. S. Chung, William M. Wells III, Alexander Norbash, and W. Eric L. Grimson. Multi-modal image registration by minimising Kullback–Leibler distance. Lecture Notes in Computer Science, 2489:525–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2489/24890525.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2489/24890525.pdf.

Sinha:2002:CSR

[1915] Tuhin K. Sinha, David M. Cash, Robert J. Weil, Robert L. Galloway, and Michael I. Miga. Cortical surface registration using texture mapped point clouds and mutual information. Lecture Notes in Computer Science, 2489:533–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2489/24890533.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890533.pdf.

DAgostino:2002:VFM

[1916] E. D’Agostino, F. Maes, D. Vandermeulen, and P. Suetens. A viscous fluid model for multimodal non-rigid image registration using mutual information. Lecture Notes in Computer Science, 2489:541–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890541.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890541.pdf.

Soza:2002:NRR

[1917] Grzegorz Soza, Michael Bauer, Peter Hastreiter, Christopher Nimsky, and Günther Greiner. Non-rigid registration with use of hardwarebased 3D Bézier functions. Lecture Notes in Computer Science, 2489: 549–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2489/24890549.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2489/24890549.
pdf.

Nielsen:2002:BWL

[1918] M. Nielsen, P. Johansen, A. D. Jackson, and B. Lautrup. Brownian warps: a least committed prior for non-rigid registration. Lecture Notes in Computer Science, 2489:557–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2489/24890557.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2489/24890557.pdf.

Hartkens:2002:UPS

[1919] T. Hartkens, D. L. G. Hill, A. D. Castellano-Smith, D. J. Hawkes, C. R. Maurer, Jr., A. J. Martin, W. A. Hall, H. Liu, and C. L. Truwit. Using points and surfaces to improve voxel-based non-rigid registration. Lecture Notes in Computer Science, 2489:565–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2489/24890565.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2489/24890565.pdf.

Nain:2002:IPP

[1920] Delphine Nain, Steven Haker, W. Eric L. Grimson, Eric Cosman, Jr., William W. Wells, Hoon Ji, Ron Kikinis, and Carl-Fredrik Westin. Intra-patient prone to supine colon registration for synchronized virtual colonoscopy. Lecture Notes in Computer Science, 2489:573–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2489/24890573.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2489/24890573.pdf.

Suarez:2002:NRU

[1921] Eduardo Suárez, Carl-Fredrik Westin, Eduardo Rovaris, and Juan Ruiz-Alzola. Nonrigid registration using regularized matching weighted by local structure. Lecture Notes in Computer Science, 2489:581–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890581.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890581.pdf.

Hellier:2002:ISR

[1922] P. Hellier, J. Ashburner, I. Corouge, C. Barillot, and K. J. Friston. Inter subject registration of functional and anatomical data using SPM. Lecture Notes in Computer Science, 2489:590–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890590.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890590.pdf.

Pommert:2002:EIQ

[1923] Andreas Pommert and Karl Heinz Höhne. Evaluation of image quality in medical volume visualization: The state of the art. Lecture Notes in Computer Science, 2489: 598–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2489/24890598.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2489/24890598.
pdf.

Wang:2002:SWV

[1924] Zhenlan Wang, Chee-Kong Chui, Chuan-Heng Ang, and Wieslaw L. Nowinski. Shear-warp volume rendering algorithms using linear level octree for PC-based medical simulation. Lecture Notes in Computer Science, 2489:606–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2489/24890606.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2489/24890606.pdf.

McGraw:2002:LIC

[1925] T. McGraw, B. C. Vemuri, Z. Wang, Y. Chen, M. Rao, and T. Mareci. Line integral convolution for visualization of fiber tract maps from DTI. Lecture Notes in Computer Science, 2489:615–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2489/24890615.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2489/24890615.pdf.

Pommert:2002:AIT

[1926] Andreas Pommert, Ulf Tiede, and Karl Heinz Höhne. On the accuracy of isosurfaces in tomographic volume visualization. Lecture Notes in Computer Science, 2489:623–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2489/24890623.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2489/24890623.pdf.

Hayashi:2002:MDU

[1927] Yuichiro Hayashi, Kensaku Mori, Jun ichi Hasegawa, Yasuhito Suenaga, and Jun ichiro Toriwaki. A method for detecting undisplayed regions in virtual colonoscopy and its application to quantitative evaluation of fly-through methods. Lecture Notes in Computer Science, 2489:631–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890631.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890631.pdf.

Rosch:2002:RMC

[1928] Peter Rösch, Thomas Netsch, Marcel Quist, and Jürgen Weese. 3D respiratory motion compensation by template propagation. Lecture Notes in Computer Science, 2489: 639–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2489/24890639.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2489/24890639.
pdf.

Schmanske:2002:EOM

[1929] Brian M. Schmanske and Murray H.

Loew. An efficient observer model for assessing signal detection performance of lossy-compressed images. Lecture Notes in Computer Science, 2489:647–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2489/24890647.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2489/24890647.pdf.

Corouge:2002:SMP

[1930] Isabelle Corouge and Christian Barillot. Statistical modeling of pairs of sulci in the context of neuroimaging probabilistic atlas. Lecture Notes in Computer Science, 2489: 655–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2489/24890655.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2489/24890655.
                                        pdf.

Freire:2002:TSA

[1931] L. Freire and J.-F. Mangin. Twostage alignment of fMRI time series using the experiment profile to discard activation-related bias. Lecture Notes in Computer Science, 2489:663–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2489/24890663.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2489/24890663.pdf.

Wang:2002:RTD

[1932] Fei Wang, Thomas E. Davis, and Baba C. Vemuri. Real-time DRR generation using cylindrical harmonics. Lecture Notes in Computer Science, 2489:671–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890671.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890671.pdf.

Yamagishi:2002:SPM

[1933] Tetsuya Yamagishi, Karl Heinz Höhne, Taku Saito, Kimihiko Abe, Jiro Ishida, Ryuko Nishimura, and Tadashi Kudo. Strengthening the potential of magnetic resonance cholangiopancreatography (MRCP) by a combination of high-resolution data acquisition and omni-directional stereoscopic viewing. Lecture Notes in Computer Science, 2489:679–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2489/24890679.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2489/24890679.pdf.

Anonymous:2002:AIbq

[1934] Anonymous. Author index. Lecture Notes in Computer Science, 2489: 687–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2489/2489auth.pdf.

Auth:2002:SAX

[1935] Gunnar Auth and Eitel von Maur.

A software architecture for XMLbased metadata interchange in data warehouse systems. Lecture Notes in Computer Science, 2490:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2490/24900001.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2490/24900001.pdf.

Klettke:2002:MXD

[1936] Meike Klettke, Lars Schneider, and Andreas Heuer. Metrics for XML document collections. Lecture Notes in Computer Science, 2490:15–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2490/24900015.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2490/24900015.pdf.

Lujan-Mora:2002:WOA

[1937] Sergio Luján-Mora, Enrique Medina, and Juan Trujillo. A Web-oriented approach to manage multidimensional models through XML schemas and XSLT. Lecture Notes in Computer Science, 2490:29–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2490/24900029.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2490/24900029.pdf.

Bertino:2002:ESD

[1938] Elisa Bertino, Giovanna Guerrini, Marco Mesiti, and Luigi Tosetto. Evolving a set of DTDs according to a dynamic set of XML documents. Lecture Notes in Computer Science, 2490:45–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2490/24900045.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2490/24900045.pdf.

Kempa:2002:XBA

[1939] Martin Kempa and Volker Linnemann.

XML-based applications using XML schema. Lecture Notes in Computer Science, 2490:67–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2490/24900067.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2490/24900067.pdf.

Kha:2002:SNS

[1940] Dao Dinh Kha, Masatoshi Yoshikawa, and Shunsuke Uemura. A structural numbering scheme for XML data. Lecture Notes in Computer Science, 2490:91–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2490/24900091.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2490/24900091.pdf.

Olteanu:2002:XLF

[1941] Dan Olteanu, Holger Meuss, Tim Furche, and François Bry. XPath: Looking forward. Lecture Notes in Computer Science, 2490:109– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/

                                        0558/bibs/2490/24900109.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2490/24900109.
                                        pdf.

Paparizos:2002:GX

[1942] Stelios Paparizos, Shurug Al-Khalifa, H. V. Jagadish, Laks Lakshmanan, Andrew Nierman, Divesh Srivastava, and Yuqing Wu. Grouping in XML. Lecture Notes in Computer Science, 2490:128–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2490/24900128.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2490/24900128.pdf.

Ludascher:2002:TLT

[1943] Bertram Ludäscher, Ilkay Altintas, and Amarnath Gupta. Time to leave the trees: From syntactic to conceptual querying of XML. Lecture Notes in Computer Science, 2490:148–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2490/24900148.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2490/24900148.pdf.

Norvaag:2002:ATQ

[1944] Kjetil Nørvåg. Algorithms for temporal query operators in XML databases. Lecture Notes in Computer Science, 2490:169–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2490/24900169.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2490/24900169.pdf.

Pankowski:2002:XSX

[1945] Tadeusz Pankowski. XML-SQL:

An XML query language based on SQL and path tables. Lecture Notes in Computer Science, 2490: 184–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2490/24900184.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2490/24900184.
pdf.

Kudrass:2002:MXD

[1946] Thomas Kudrass and Matthias Conrad. Management of XML documents in object-relational databases. Lecture Notes in Computer Science, 2490:210–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2490/24900210.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2490/24900210.pdf.

Lewis:2002:EXR

[1947] Bryn Lewis. Extraction of XML from relational databases. Lecture Notes in Computer Science, 2490: 228–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2490/24900228.htm;   http:
//link.springer.de/link/service/

                                        series/0558/papers/2490/24900228.
                                        pdf.

Psaila:2002:EEI

[1948] Giuseppe Psaila. ERX: An experience in integrating entity-relationship models, relational databases, and XML technologies. Lecture Notes in Computer Science, 2490:242– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2490/24900242.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2490/24900242.
                                        pdf.

Runapongsa:2002:SQX

[1949] Kanda Runapongsa and Jignesh M. Patel. Storing and querying XML data in object-relational DBMSs. Lecture Notes in Computer Science, 2490:266–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2490/24900266.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2490/24900266.pdf.

Ardizzone:2002:PQI

[1950] Edoardo Ardizzone, Marco La Cascia, Umberto Maniscalco, Daniele Peri, and Roberto Pirrone. Population and query interface for a content-based video database. Lecture Notes in Computer Science, 2490:286–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2490/24900286.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2490/24900286.pdf.

Zhou:2002:SSC

[1951] Junyu Zhou and Wallapak Tavanapong. ShotWeave: a shot clustering technique for story browsing for large video databases. Lecture Notes in Computer Science, 2490: 299–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2490/24900299.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2490/24900299.
pdf.

Petkovic:2002:EDS

[1952] Milan Petković, Vojkan Mihajlović, and Willem Jonker. Extending a DBMS to support content-based video retrieval: a formula 1 case study. Lecture Notes in Computer Science, 2490:318–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2490/24900318.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2490/24900318.pdf.

Vachon:2002:PVS

[1953] Eric Vachon. A pre-viewing step in video retrieval. Lecture Notes in Computer Science, 2490:342– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2490/24900342.htm;   http:

                                        //link.springer.de/link/service/
                                        series/0558/papers/2490/24900342.
                                        pdf.

Schmitt:2002:MLW

[1954] Ingo Schmitt, Nadine Schulz, and Gunter Saake. Multi-level weighting in multimedia retrieval systems. Lecture Notes in Computer Science, 2490:353–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2490/24900353.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2490/24900353.pdf.

Balko:2002:TEC

[1955] Sören Balko, Ingo Schmitt, and Gunter Saake. Towards enhanced compression techniques for efficient high-dimensional similarity search in multimedia databases. Lecture Notes in Computer Science, 2490: 365–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2490/24900365.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2490/24900365.
                                        pdf.

Henrich:2002:RAS

[1956] Andreas Henrich. A relaxed algorithm for similarity queries performed with high-dimensional access structures. Lecture Notes in Computer Science, 2490:376–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2490/24900376.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2490/24900376.pdf.

Won:2002:ACE

[1957] Youjip Won and Kyungsun Cho.

Adaptive cycle extension in multimedia document retrieval. Lecture Notes in Computer Science, 2490: 391–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2490/24900391.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2490/24900391.
pdf.

Fernandez:2002:IIC

[1958] Gregory Fernandez, Abdelouaheb Meckaouche, Philippe Peter, and Chabane Djeraba. Intelligent image clustering. Lecture Notes in Computer Science, 2490:406–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2490/24900406.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2490/24900406.pdf.

Malerba:2002:MHP

[1959] Donato Malerba, Floriana Esposito, and Michelangelo Ceci. Mining HTML pages to support document sharing in a cooperative system. Lecture Notes in Computer Science, 2490:420–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2490/24900420.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2490/24900420.pdf.

Zimmermann:2002:MCI

[1960] Birgit Zimmermann, Melanie Gnasa, and Karin Harbusch. Modeling a corporate information system to improve knowledge management. Lecture Notes in Computer Science, 2490:435–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2490/24900435.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2490/24900435.pdf.

Ghandeharizadeh:2002:DWS

[1961] Shahram Ghandeharizadeh, Frank Sommers, Kuntal Joisher, and Esam Alwagait. A document as a Web service: Two complementary frameworks. Lecture Notes in Computer Science, 2490:450–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2490/24900450.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2490/24900450.pdf.

Andreadis:2002:MAD

[1962] Alessandro Andreadis, Giuliano Benelli, Pasquale Fedele, and Giovanni Giambene. Mobile access to distributed data sources. Lecture Notes in Computer Science, 2490: 462–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2490/24900462.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2490/24900462.
pdf.

Wack:2002:CNM

[1963] Maxime Wack, Nathanael Cottin, and Rachid Bouyekhf. Cooperative navigation in multimedia systems. Lecture Notes in Computer Science, 2490:472–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2490/24900472.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2490/24900472.pdf.

Cho:2002:ALR

[1964] K. W. Cho, Y. S. Ryu, Youjip Won, and Kern Koh. ABM: Looping reference-aware cache management scheme for media-on-demand server. Lecture Notes in Computer Science, 2490:484–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2490/24900484.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2490/24900484.pdf.

Soldar:2002:ARR

[1965] Goran Soldar and Dan Smith. An architecture for retrieval of RDFdescribed scientific data semantics. Lecture Notes in Computer Science, 2490:501–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2490/24900501.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2490/24900501.pdf.

Schimkat:2002:LDM

[1966] Ralf-Dieter Schimkat and Wolfgang Küchlin. Living documents — micro servers for documents. Lecture Notes in Computer Science, 2490:512–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2490/24900512.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2490/24900512.pdf.

Gudes:2002:RCM

[1967] Ehud Gudes, Carmel Domshlak, and Natalia Orlov. Remote conferencing with multimedia objects. Lecture Notes in Computer Science, 2490:526–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2490/24900526.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2490/24900526.pdf.

Dekeyser:2002:MQO

[1968] Stijn Dekeyser. Multi query optimization using query pack trees. Lecture Notes in Computer Science, 2490:544–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2490/24900544.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2490/24900544.pdf.

Schmitt:2002:IPU

[1969] Bethina Schmitt. Impact and potential of user profiles used for distributed query processing based on literature services. Lecture Notes in Computer Science, 2490:555–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2490/24900555.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2490/24900555.pdf.

Tatbul:2002:QDL

[1970] Nesime Tatbul. QoS-driven load shedding on data streams. Lecture Notes in Computer Science, 2490: 566–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2490/24900566.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2490/24900566.
pdf.

Bright:2002:PBD

[1971] Laura Bright. Profile-based data delivery for Web applications. Lecture Notes in Computer Science, 2490: 577–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2490/24900577.htm;   http:
//link.springer.de/link/service/

                                        series/0558/papers/2490/24900577.
                                        pdf.

Carminati:2002:SPX

[1972] Barbara Carminati. Secure publishing of XML documents. Lecture Notes in Computer Science, 2490: 587–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2490/24900587.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2490/24900587.
                                        pdf.

Klein:2002:SEO

[1973] Michel Klein. Supporting evolving ontologies on the Internet. Lecture Notes in Computer Science, 2490:597–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2490/24900597.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2490/24900597.pdf.

Garcia-Banuelos:2002:AIC

[1974] Luciano García-Bañuelos. An adaptable infrastructure for customized persistent object management. Lecture Notes in Computer Science, 2490: 607–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2490/24900607.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2490/24900607.
                                        pdf.

Serrano-Alvarado:2002:DAM

[1975] Patricia Serrano-Alvarado. Defining an adaptable mobile transaction service. Lecture Notes in Computer Science, 2490:616–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2490/24900616.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2490/24900616.pdf.

Katchaounov:2002:QPS

[1976] Timour Katchaounov. Query processing in self-profiling composable peerto-peer mediator databases. Lecture Notes in Computer Science, 2490: 627–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2490/24900627.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2490/24900627.
pdf.

Balko:2002:CTB

[1977] Sören Balko. Compression techniques based on concave cluster geometries for efficient high-dimensional nearest neighbor retrieval. Lecture Notes in Computer Science, 2490: 638–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2490/24900638.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2490/24900638.
pdf.

Anonymous:2002:AIbr

[1978] Anonymous. Author index. Lecture Notes in Computer Science, 2490: 651–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2490/2490auth.pdf.

Pnueli:2002:ESC

[1979] Amir Pnueli. Embedded systems: Challenges in specification and verification. Lecture Notes in Computer Science, 2491:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2491/24910001.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2491/24910001.pdf.

Leveson:2002:ADS

[1980] Nancy G. Leveson. An approach to designing safe embedded software. Lecture Notes in Computer Science, 2491:15–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2491/24910015.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2491/24910015.pdf.

Harel:2002:CBRa

[1981] David Harel. Can behavioral requirements be executed? (and why would we want to do so?). Lecture Notes in Computer Science, 2491:30–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2491/24910030.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2491/24910030.pdf.

Koo:2002:PBE

[1982] T. John Koo, Judith Liebman, Cedric Ma, Benjamin Horowitz, Alberto Sangiovanni-Vincentelli, and Shankar Sastry. Platform-based embedded software design for multivehicle multi-modal systems. Lecture Notes in Computer Science, 2491: 32–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2491/24910032.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2491/24910032.
pdf.

Kirsch:2002:GBH

[1983] Christoph M. Kirsch, Marco A. A.

Sanvido, Thomas A. Henzinger, and Wolfgang Pree. A Giottobased helicopter control system. Lecture Notes in Computer Science, 2491:46–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2491/24910046.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2491/24910046.pdf.

Kirsch:2002:PRT

[1984] Christoph M. Kirsch. Principles of real-time programming. Lecture Notes in Computer Science, 2491:

61–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2491/24910061.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2491/24910061.
                                        pdf.

Henzinger:2002:TSC

[1985] Thomas A. Henzinger, Christoph M. Kirsch, Rupak Majumdar, and Slobodan Matic. Time-safety checking for embedded programs. Lecture Notes in Computer Science, 2491: 76–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2491/24910076.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2491/24910076.
                                        pdf.

Gossler:2002:CMM

[1986] Gregor Gössler and Alberto Sangiovanni-Vincentelli. Compositional modeling in metropolis. Lecture Notes in Computer Science, 2491:93–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2491/24910093.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2491/24910093.pdf.

deAlfaro:2002:TI

[1987] Luca de Alfaro, Thomas A. Henzinger, and Mariëlle Stoelinga. Timed interfaces. Lecture Notes in Computer Science, 2491:108–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2491/24910108.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2491/24910108.pdf.

Tripakis:2002:DSA

[1988] Stavros Tripakis. Description and schedulability analysis of the software architecture of an automated vehicle control system. Lecture Notes in Computer Science, 2491: 123–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2491/24910123.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2491/24910123.
pdf.

Cofer:2002:FMA

[1989] Darren Cofer and Murali Rangarajan.

Formal modeling and analysis of advanced scheduling features in an avionics RTOS. Lecture Notes in Computer Science, 2491:138–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2491/24910138.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2491/24910138.pdf.

Buttazzo:2002:SAE

[1990] Giorgio C. Buttazzo. Scalable applications for energy-aware processors. Lecture Notes in Computer Science, 2491:153–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2491/24910153.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2491/24910153.pdf.

Sha:2002:UES

[1991] Lui Sha. Upgrading embedded software in the field: Dependability and survivability. Lecture Notes in Computer Science, 2491:166–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2491/24910166.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2491/24910166.pdf.

Mok:2002:RTV

[1992] Aloysius K. Mok and Alex Xiang Feng. Real-time virtual resource: a timely abstraction for embedded systems. Lecture Notes in Computer Science, 2491:182–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2491/24910182.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2491/24910182.pdf.

Delsart:2002:JLM

[1993] Bertrand Delsart, Vania Joloboff, and Eric Paire. JCOD: a lightweight modular compilation technology for embedded Java. Lecture Notes in Computer Science, 2491:197–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2491/24910197.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2491/24910197.pdf.

Ritzau:2002:DMO

[1994] Tobias Ritzau and Peter Fritzson. Decreasing memory overhead in hard real-time garbage collection. Lecture Notes in Computer Science, 2491:213–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2491/24910213.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2491/24910213.pdf.

ColindeVerdiere:2002:SPS

[1995] Vincent Colin de Verdière, Sébastien Cros, Christian Fabre, Romain Guider, and Sergio Yovine. Speedup prediction for selective compilation of embedded Java programs. Lecture Notes in Computer Science, 2491:227–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2491/24910227.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2491/24910227.pdf.

Halbwachs:2002:SMA

[1996] Nicolas Halbwachs and Siwar Baghdadi. Synchronous modelling of asynchronous systems. Lecture Notes in Computer Science, 2491: 240–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/

                                        0558/bibs/2491/24910240.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2491/24910240.
                                        pdf.

Benveniste:2002:PLT

[1997] Albert Benveniste, Paul Caspi, Paul Le Guernic, Hervé Marchand, Jean-Pierre Talpin, and Stavros Tripakis. A protocol for loosely time-triggered architectures. Lecture Notes in Computer Science, 2491:252–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2491/24910252.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2491/24910252.pdf.

Girault:2002:APG

[1998] Alain Girault and Clément Ménier. Automatic production of globally asynchronous locally synchronous systems. Lecture Notes in Computer Science, 2491:266–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2491/24910266.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2491/24910266.pdf.

Schmidt:2002:ARM

[1999] Douglas C. Schmidt. Adaptive and reflective middleware for distributed real-time and embedded systems. Lecture Notes in Computer Science, 2491:282–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2491/24910282.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2491/24910282.pdf.

Caspi:2002:TAT

[2000] Paul Caspi and Albert Benveniste.

Toward an approximation theory for computerised control. Lecture Notes in Computer Science, 2491: 294–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2491/24910294.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2491/24910294.
pdf.

Desoli:2002:NFD

[2001] Giuseppe Desoli, Nikolay Mateev, Evelyn Duesterwald, Paolo Faraboschi, and Josh Fisher. A new facility for dynamic control of program execution: DELI. Lecture Notes in Computer Science, 2491:305–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2491/24910305.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2491/24910305.pdf.

Qin:2002:DTA

[2002] Wei Qin, Subramanian Rajagopalan, Manish Vachharajani, Hangsheng Wang, Xinping Zhu, David August, Kurt Keutzer, Sharad Malik, and Li-Shiuan Peh. Design tools for application specific embedded processors. Lecture Notes in Computer Science, 2491:319–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2491/24910319.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2491/24910319.pdf.

Engblom:2002:PPT

[2003] Jakob Engblom and Bengt Jonsson. Processor pipelines and their properties for static WCET analysis. Lecture Notes in Computer Science, 2491:334–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2491/24910334.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2491/24910334.pdf.

Theiling:2002:IBI

[2004] Henrik Theiling. ILP-based interprocedural path analysis. Lecture Notes in Computer Science, 2491: 349–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2491/24910349.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2491/24910349.
                                        pdf.

Zambreno:2002:ECT

[2005] Joseph Zambreno, Mahmut Taylan Kandemir, and Alok Choudhary. Enhancing compiler techniques for memory energy optimizations. Lecture Notes in Computer Science, 2491:364–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2491/24910364.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2491/24910364.pdf.

Bertin:2002:FOR

[2006] Valérie Bertin, Jean-Marc Daveau, Philippe Guillaume, Thierry Lepley, Denis Pilat, Claire Richard, Miguel Santana, and Thomas Thery. FlexCC2: An optimizing retargetable C compiler for DSP processors. Lecture Notes in Computer Science, 2491:382–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2491/24910382.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2491/24910382.pdf.

Selic:2002:PPB

[2007] Bran Selic. Physical programming: Beyond mere logic. Lecture Notes in Computer Science, 2491: 399–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2491/24910399.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2491/24910399.
pdf.

Balarin:2002:PIP

[2008] F. Balarin, L. Lavagno, C. Passerone, and Y. Watanabe. Processes, interfaces and platforms. embedded software modeling in metropolis. Lecture Notes in Computer Science, 2491:407–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2491/24910407.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2491/24910407.pdf.

Anonymous:2002:AIbs

[2009] Anonymous. Author index. Lecture Notes in Computer Science, 2491: 417–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2491/2491auth.pdf.

Galam:2002:SCF

[2010] Serge Galam. Spontaneous coalition forming. why some are stable? Lecture Notes in Computer Science, 2493:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2493/24930001.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2493/24930001.pdf.

Bandman:2002:SSD

[2011] Olga Bandman. Simulating spatial dynamics by probabilistic cellular automata. Lecture Notes in Computer Science, 2493:10–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2493/24930010.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2493/24930010.pdf.

Nagel:2002:CAM

[2012] Kai Nagel. Cellular automata models for transportation applications. Lecture Notes in Computer Science, 2493:20–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2493/24930020.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2493/24930020.pdf.

Jimenez-Morales:2002:EAS

[2013] Francisco Jiménez-Morales. An evolutionary approach to the study of nontrivial collective behavior in cellular automata. Lecture Notes in Computer Science, 2493:32–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2493/24930032.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2493/24930032.pdf.

Tomassini:2002:AEA

[2014] Marco Tomassini and Mattias Venzi.

Artificially evolved asynchronous cellular automata for the density task. Lecture Notes in Computer Science, 2493:44–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2493/24930044.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2493/24930044.pdf.

Ganguly:2002:ECA

[2015] Niloy Ganguly, Pradipta Maji, Sandip Dhar, Biplab K. Sikdar, and P. Pal Chaudhuri. Evolving cellular automata as pattern classifier. Lecture Notes in Computer Science, 2493: 56–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2493/24930056.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2493/24930056.
pdf.

Umeo:2002:EMS

[2016] Hiroshi Umeo, Masashi Maeda, and Norio Fujiwara. An efficient mapping scheme for embedding any onedimensional firing squad synchronization algorithm onto two-dimensional arrays. Lecture Notes in Computer Science, 2493:69–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2493/24930069.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2493/24930069.pdf.

Cattaneo:2002:CSG

[2017] Gianpiero Cattaneo and Alberto Dennunzio. Chaotic subshifts generated by one dimensional elementary CA. the role of transitivity. Lecture Notes in Computer Science, 2493:82–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2493/24930082.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2493/24930082.pdf.

Muhlenbein:2002:SAC

[2018] Heinz Mühlenbein and Robin Höns.

Stochastic analysis of cellular automata and the voter model. Lecture Notes in Computer Science, 2493: 92–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2493/24930092.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2493/24930092.
pdf.

Makowiec:2002:UCP

[2019] Danuta Makowiec and Piotr Gnaciński.

Universality class of probabilistic cellular automata. Lecture Notes in Computer Science, 2493:104– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2493/24930104.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2493/24930104.
pdf.

Succi:2002:KAL

[2020] Sauro Succi. Kinetic approach to lattice quantum mechanics. Lecture Notes in Computer Science, 2493: 114–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2493/24930114.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2493/24930114.
pdf.

Stauffer:2002:ESR

[2021] André Stauffer and Moshe Sipper. Emergence of self-replicating loops in an interactive, hardwareimplemented game-of-life environment. Lecture Notes in Computer Science, 2493:123–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2493/24930123.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2493/24930123.pdf.

Azpeitia:2002:SER

[2022] Iker Azpeitia and Jesús Ibáñez. Spontaneous emergence of robust cellular replicators. Lecture Notes in Computer Science, 2493:132–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2493/24930132.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2493/24930132.pdf.

Roli:2002:EMS

[2023] Andrea Roli and Franco Zambonelli. Emergence of macro spatial structures in dissipative cellular automata. Lecture Notes in Computer Science, 2493:144–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2493/24930144.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2493/24930144.pdf.

Bandini:2002:ECS

[2024] Stefania Bandini, Sara Manzoni, and Carla Simone. Enhancing cellular spaces by multilayered multi agent situated systems. Lecture Notes in Computer Science, 2493: 156–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2493/24930156.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2493/24930156.
pdf.

Serra:2002:PRT

[2025] Roberto Serra and Marco Villani. Perturbing the regular topology of cellular automata: Implications for the dynamics. Lecture Notes in Computer Science, 2493:168–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2493/24930168.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2493/24930168.pdf.

Marchese:2002:PPM

[2026] Fabio M. Marchese. A path-planner for mobile robots of generic shape with multilayered cellular automata. Lecture Notes in Computer Science, 2493:178–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2493/24930178.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2493/24930178.pdf.

Droz:2002:DPE

[2027] Michel Droz and Andrzej Pekalski. Dynamics of populations in extended systems. Lecture Notes in Computer Science, 2493:190–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2493/24930190.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2493/24930190.pdf.

Bandini:2002:SVP

[2028] Stefania Bandini and Giulio Pavesi. Simulation of vegetable populations dynamics based on cellular automata. Lecture Notes in Computer Science, 2493:202–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2493/24930202.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2493/24930202.pdf.

Schonfisch:2002:FMM

[2029] Birgitt Schönfisch and Michael Kinder. A fish migration model. Lecture Notes in Computer Science, 2493: 210–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2493/24930210.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2493/24930210.
                                        pdf.

Albuquerque:2002:PCA

[2030] Paul Albuquerque and Alexandre Dupuis. A parallel cellular ant colony algorithm for clustering and sorting. Lecture Notes in Computer Science, 2493:220–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2493/24930220.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2493/24930220.pdf.

Marconi:2002:MLG

[2031] Stefan Marconi and Bastien Chopard.

A multiparticle lattice gas automata model for a crowd. Lecture Notes in Computer Science, 2493: 231–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2493/24930231.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2493/24930231.
pdf.

Schadschneider:2002:CAC

[2032] Andreas Schadschneider, Ansgar Kirchner, and Katsuhiro Nishinari. CA approach to collective phenomena in pedestrian dynamics. Lecture Notes in Computer Science, 2493: 239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2493/24930239.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2493/24930239.
pdf.

Bagnoli:2002:OFP

[2033] Franco Bagnoli, Fabio Franci, and Raúl Rechtman. Opinion formation and phase transitions in a probabilistic cellular automaton with two absorbing states. Lecture Notes in Computer Science, 2493:249–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2493/24930249.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2493/24930249.pdf.

Mukherjee:2002:CAB

[2034] Monalisa Mukherjee, Niloy Ganguly, and P. Pal Chaudhuri. Cellular automata based authentication (CAA). Lecture Notes in Computer Science, 2493:259–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2493/24930259.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2493/24930259.pdf.

Maji:2002:CAM

[2035] Pradipta Maji, Niloy Ganguly, Sourav Saha, Anup K. Roy, and P. Pal Chaudhuri. Cellular automata machine for pattern recognition. Lecture Notes in Computer Science, 2493:270–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2493/24930270.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2493/24930270.pdf.

Sloot:2002:CAM

[2036] Peter Sloot, Fan Chen, and Charles Boucher. Cellular automata model of drug therapy for HIV infection. Lecture Notes in Computer Science, 2493:282–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2493/24930282.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2493/24930282.pdf.

Weimar:2002:CAA

[2037] Jörg R. Weimar. Cellular automata approaches to enzymatic reaction networks. Lecture Notes in Computer Science, 2493:294–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2493/24930294.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2493/24930294.pdf.

DAmbrosio:2002:MSF

[2038] Donato D’Ambrosio, Salvatore Di Gregorio, Giulio Iovine, Valeria Lupiano, Rocco Rongo, and William Spataro. Modelling surface flows for macroscopic phenomena by cellular automata: An application to debris flows. Lecture Notes in Computer Science, 2493:304–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2493/24930304.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2493/24930304.pdf.

Marinosson:2002:SFA

[2039] Sigurur F. Marinósson, Roland Chrobok, Andreas Pottmeier, Joachim Wahle, and Michael Schreckenberg. Simulation framework for the Autobahn traffic in North Rhine–Westphalia. Lecture Notes in Computer Science, 2493:315–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2493/24930315.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2493/24930315.pdf.

Cheng:2002:CAB

[2040] Jianquan Cheng and Ian Masser. Cellular automata based temporal process understanding of urban growth. Lecture Notes in Computer Science, 2493:325–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2493/24930325.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2493/24930325.pdf.

Blecic:2002:PAI

[2041] Ivan Blecic, Arnaldo Cecchini, Paola Rizzi, and Giuseppe A. Trunfio. Playing with automata. an innovative perspective for gaming simulation. Lecture Notes in Computer Science, 2493:337–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2493/24930337.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2493/24930337.pdf.

Rabino:2002:UCA

[2042] Giovanni A. Rabino and Alessandra Laghi. Urban cellular automata: The inverse problem. Lecture Notes in Computer Science, 2493: 349–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2493/24930349.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2493/24930349.
pdf.

Yacoubi:2002:RCC

[2043] Samira El Yacoubi, Abdelhaq El Jai, and Nezha Ammor. Regional controllability with cellular automata models. Lecture Notes in Computer Science, 2493:357–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2493/24930357.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2493/24930357.pdf.

Anonymous:2002:AIbt

[2044] Anonymous. Author index. Lecture Notes in Computer Science, 2493: 369–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2493/2493auth.pdf.

Das:2002:QNM

[2045] S. Das, K. Yamada, H. Yu, S. S.

Lee, and M. Gerla. A QoS network management system for robust and reliable multimedia services. Lecture Notes in Computer Science, 2496:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2496/24960001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2496/24960001.pdf.

Bhushan:2002:FAM

[2046] B. Bhushan, T. Gringel, C. Ryan, E. Leray, E. de Leastar, and J. Cloney. Federated accounting management system architecture for multimedia service usage management. Lecture Notes in Computer Science, 2496:12–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2496/24960012.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2496/24960012.pdf.

Gay:2002:PBQ

[2047] Valérie Gay, Sandrine Duflos, Brigitte Kervella, Gladys Diaz, and Eric Horlait. Policy-based quality of service and security management for multimedia services on IP networks in the RTIPA project. Lecture Notes in Computer Science, 2496:25–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2496/24960025.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2496/24960025.pdf.

vanLe:2002:FMS

[2048] M. van Le, B. J. F. van Beijnum, and B. L. de Goede. Formal modeling of service session management. Lecture Notes in Computer Science, 2496:36–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2496/24960036.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2496/24960036.pdf.

Zheng:2002:NRM

[2049] Bing Zheng and Mohammed Atiquzzaman. Network requirement for management of multimedia over wireless channel. Lecture Notes in Computer Science, 2496:49–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2496/24960049.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2496/24960049.pdf.

Adwankar:2002:ASM

[2050] Sandeep Adwankar and Venu Vasudevan. Agile systems manager for enterprise wireless networks. Lecture Notes in Computer Science, 2496:62–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2496/24960062.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2496/24960062.pdf.

Haider:2002:SPH

[2051] K. Haider and H. S. Al-Raweshidy. System performance of HiperLAN/2. Lecture Notes in Computer Science, 2496:77–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2496/24960077.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2496/24960077.pdf.

Aboobaker:2002:SMC

[2052] Nadeem Aboobaker, David Chanady, Mario Gerla, and M. Y. Sanadidi. Streaming media congestion control using bandwidth estimation. Lecture Notes in Computer Science, 2496:89–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2496/24960089.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2496/24960089.pdf.

Mendes:2002:SPS

[2053] Paulo Mendes, Henning Schulzrinne, and Edmundo Monteiro. Signaling protocol for session-aware popularity-based resource allocation. Lecture Notes in Computer Science, 2496:101–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2496/24960101.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2496/24960101.pdf.

Tan:2002:SVS

[2054] Guang Tan, Hai Jin, and Liping Pang. A scalable video server using intelligent network attached storage. Lecture Notes in Computer Science, 2496:114–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2496/24960114.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2496/24960114.pdf.

Wakamiya:2002:PCM

[2055] Naoki Wakamiya, Masayuki Murata, and Hideo Miyahara. On proxycaching mechanisms for cooperative video streaming in heterogeneous environments. Lecture Notes in Computer Science, 2496:127–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2496/24960127.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2496/24960127.pdf.

Gu:2002:UCA

[2056] Qiang Gu and Alan Marshall. Using CORBA’s advanced services to enhance the integrity of QoS management programmable networks. Lecture Notes in Computer Science, 2496:140–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2496/24960140.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2496/24960140.pdf.

Chesterfield:2002:RMM

[2057] Julian Chesterfield, Bill Fenner, and Lee Breslau. Remote multicast monitoring using the RTP MIB. Lecture Notes in Computer Science, 2496:154–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2496/24960154.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2496/24960154.pdf.

Achir:2002:ATE

[2058] Nadjib Achir, Nazim Agoulmine, Mauro Fonseca, Yacine Ghamri-Doudane, and Guy Pujolle. Active technology as an efficient approach to control DiffServ networks: The DACA architecture. Lecture Notes in Computer Science, 2496:170–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2496/24960170.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2496/24960170.pdf.

Shi:2002:EIP

[2059] Weidong Shi, Xiaotong Zhuang, Indrani Paul, and Karsten Schwan. Efficient implementation of packet scheduling algorithm on high-speed programmable network processors. Lecture Notes in Computer Science, 2496:184–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2496/24960184.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2496/24960184.pdf.

Yuksel:2002:DDC

[2060] Murat Yuksel and Shivkumar Kalyanaraman. Distributed dynamic capacity contracting: a congestion pricing framework for Diff-Serv. Lecture Notes in Computer Science, 2496:198–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2496/24960198.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2496/24960198.pdf.

Das:2002:FSA

[2061] A. Das, D. Dutta, and A. Helmy. Fair stateless aggregate traffic marking using active queue management techniques. Lecture Notes in Computer Science, 2496:211–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2496/24960211.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2496/24960211.pdf.

Seok:2002:DMS

[2062] Seung-Joon Seok, Seok-Min Hong, Sung-Hyuck Lee, and Chul-Hee Kang. A dynamic marking scheme of assured service for alleviating unfairness among service level agreements. Lecture Notes in Computer Science, 2496:224–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2496/24960224.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2496/24960224.pdf.

Tham:2002:MTC

[2063] Chen-Khong Tham and Yong Liu. Minimizing transmission costs through adaptive marking in differentiated services networks. Lecture Notes in Computer Science, 2496:237– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2496/24960237.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2496/24960237.
                                        pdf.

Ahmed:2002:DQA

[2064] Toufik Ahmed, Ahmed Mehaoua, and Raouf Boutaba. Dynamic QoS adaptation using COPS and network monitoring feedback. Lecture Notes in Computer Science, 2496: 250–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2496/24960250.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2496/24960250.
                                        pdf.

Kulju:2002:DIA

[2065] William Kulju and Hanan Lutfiyya. Design and implementation of an application layer protocol for reducing UDP traffic based on user hints and policies. Lecture Notes in Computer Science, 2496:263–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2496/24960263.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2496/24960263.pdf.

Ravindran:2002:MFS

[2066] Govindan Ravindran, Muhammad Jaseemudin, and Abdallah Rayhan. A management framework for service personalization. Lecture Notes in Computer Science, 2496: 276–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2496/24960276.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2496/24960276.
pdf.

Kim:2002:EAU

[2067] Dae gun Kim, Seung-Jin Lee, Lynn Choi, and Chul-Hee Kang. Efficient access using hierarchical WML decks for multimedia services under wireless and mobile networks. Lecture Notes in Computer Science, 2496:289–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2496/24960289.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2496/24960289.pdf.

Floris:2002:RSB

[2068] Andrea Floris and Luca Veltri.

Roaming scenarios based on SIP. Lecture Notes in Computer Science, 2496:302–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2496/24960302.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2496/24960302.pdf.

Scholl:2002:LWC

[2069] Jeremiah Scholl and Peter Parnes. Low-weight congestion control for multi-sender applications. Lecture Notes in Computer Science, 2496: 315–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2496/24960315.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2496/24960315.
                                        pdf.

Peng:2002:RBV

[2070] Jun Peng and Biplab Sikdar. Routingbased video multicast congestion control. Lecture Notes in Computer Science, 2496:328–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2496/24960328.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2496/24960328.pdf.

Gan:2002:RED

[2071] Yung-Sze Gan and Chen-Khong Tham. Random early detection assisted layered multicast. Lecture Notes in Computer Science, 2496: 341–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2496/24960341.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2496/24960341.
pdf.

Anonymous:2002:AIbu

[2072] Anonymous. Author index. Lecture Notes in Computer Science, 2496: 355–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2496/2496auth.pdf.

Crowcroft:2002:PPP

[2073] Jon Crowcroft and Ian Pratt. Peer to peer: Peering into the future. Lecture Notes in Computer Science, 2497:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2497/24970001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2497/24970001.pdf.

Mascolo:2002:MCM

[2074] Cecilia Mascolo, Licia Capra, and Wolfgang Emmerich. Mobile computing middleware. Lecture Notes in Computer Science, 2497:20–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2497/24970020.htm;   http://link.
springer.de/link/service/series/
0558/papers/2497/24970020.pdf.

Molva:2002:NSM

[2075] Refik Molva and Alain Pannetrat. Network security in the multicast framework. Lecture Notes in Computer Science, 2497:59–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2497/24970059.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2497/24970059.pdf.

Gantenbein:2002:CCA

[2076] Dieter Gantenbein and Luca Deri. Categorizing computing assets according to communication patterns. Lecture Notes in Computer Science, 2497:83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2497/24970083.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2497/24970083.pdf.

Basagni:2002:RAH

[2077] Stefano Basagni. Remarks on ad hoc networking. Lecture Notes in Computer Science, 2497:101–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2497/24970101.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2497/24970101.pdf.

Davide:2002:CTV

[2078] Fabrizio Davide, Pierpaolo Loreti, Massimiliano Lunghi, Giuseppe Riva, and Francesco Vatalaro. Communications through virtual technologies. Lecture Notes in Computer Science, 2497:124–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2497/24970124.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2497/24970124.pdf.

Rouskas:2002:TON

[2079] George N. Rouskas and Harry G.

Perros. A tutorial on optical networks. Lecture Notes in Computer Science, 2497:155–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2497/24970155.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2497/24970155.pdf.

Anonymous:2002:AIbv

[2080] Anonymous. Author index. Lecture Notes in Computer Science, 2497: 195–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2497/2497auth.pdf.

Carbonell:2002:ARL

[2081] Jaime Carbonell, Katharina Probst, Erik Peterson, Christian Monson, Alon Lavie, Ralf Brown, and Lori Levin. Automatic rule learning for resource-limited MT. Lecture Notes in Computer Science, 2499: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2499/24990001.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2499/24990001.
                                        pdf.

Carl:2002:THI

[2082] Michael Carl, Andy Way, and Reinhard Schäler. Toward a hybrid integrated translation environment. Lecture Notes in Computer Science, 2499:11–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2499/24990011.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2499/24990011.pdf.

Chuang:2002:ABS

[2083] Thomas C. Chuang, GN You, and Jason S. Chang. Adaptive bilingual sentence alignment. Lecture Notes in Computer Science, 2499:21–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2499/24990021.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2499/24990021.pdf.

Dorr:2002:DMU

[2084] Bonnie J. Dorr, Lisa Pearl, Rebecca Hwa, and Nizar Habash. DUSTer: a method for unraveling cross-language divergences for statistical word-level alignment. Lecture Notes in Computer Science, 2499:31–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2499/24990031.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2499/24990031.pdf.

Foster:2002:TPF

[2085] George Foster, Philippe Langlais, and Guy Lapalme. Text prediction with fuzzy alignments. Lecture Notes in Computer Science, 2499: 44–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2499/24990044.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2499/24990044.
pdf.

Varea:2002:EIM

[2086] Ismael García Varea, Franz J. Och, Hermann Ney, and Francisco Casacuberta. Efficient integration of maximum entropy lexicon models within the training of statistical alignment models. Lecture Notes in Computer Science, 2499:54–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2499/24990054.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2499/24990054.pdf.

Gdaniec:2002:UWF

[2087] Claudia Gdaniec and Esmé Manandise. Using word formation rules to extend MT lexicons. Lecture Notes in Computer Science, 2499: 64–??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2499/24990064.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2499/24990064.
                                        pdf.

Gough:2002:EBM

[2088] Nano Gough, Andy Way, and Mary Hearne. Example-based machine translation via the Web. Lecture Notes in Computer Science, 2499:74–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2499/24990074.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2499/24990074.pdf.

Habash:2002:HTD

[2089] Nizar Habash and Bonnie Dorr. Handling translation divergences: Combining statistical and symbolic techniques in generation-heavy machine translation. Lecture Notes in Computer Science, 2499:84–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2499/24990084.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2499/24990084.pdf.

Kim:2002:KCM

[2090] Changhyun Kim, Munpyo Hong, Yinxia Huang, Young Kil Kim, Sung Il Yang, Young Ae Seo, and Sung-Kwon Choi. Korean-Chinese machine translation based on verb patterns. Lecture Notes in Computer Science, 2499:94–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2499/24990094.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2499/24990094.pdf.

Langlais:2002:MEB

[2091] Philippe Langlais and Michel Simard.

Merging example-based and statistical machine translation: An experiment. Lecture Notes in Computer Science, 2499:104–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2499/24990104.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2499/24990104.pdf.

Lee:2002:CAW

[2092] Hyo-Kyung Lee. Classification approach to word selection in machine translation. Lecture Notes in Computer Science, 2499:114–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2499/24990114.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2499/24990114.pdf.

Menezes:2002:BCT

[2093] Arul Menezes. Better contextual translation using machine learning. Lecture Notes in Computer Science, 2499:124–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2499/24990124.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2499/24990124.pdf.

Moore:2002:FAS

[2094] Robert C. Moore. Fast and accurate sentence alignment of bilingual corpora. Lecture Notes in Computer Science, 2499:135–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2499/24990135.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2499/24990135.pdf.

Nyberg:2002:DSK

[2095] Eric Nyberg, Teruko Mitamura, and Kathryn Baker. Deriving semantic knowledge from descriptive texts using an MT system. Lecture Notes in Computer Science, 2499:145–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2499/24990145.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2499/24990145.pdf.

Soricut:2002:ULM

[2096] Radu Soricut, Kevin Knight, and Daniel Marcu. Using a large monolingual corpus to improve translation accuracy. Lecture Notes in Computer Science, 2499:155–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2499/24990155.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2499/24990155.pdf.

Utsuro:2002:SAC

[2097] Takehito Utsuro, Takashi Horiuchi, Yasunobu Chiba, and Takeshi Hamamoto. Semi-automatic compilation of bilingual lexicon entries from cross-lingually relevant news articles on WWW news sites. Lecture Notes in Computer Science, 2499:165–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2499/24990165.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2499/24990165.pdf.

Weerasinghe:2002:BLB

[2098] Ruvan Weerasinghe. Bootstrapping the lexicon building process for machine translation between ‘new’ languages. Lecture Notes in Computer Science, 2499:177–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2499/24990177.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2499/24990177.pdf.

Clarke:2002:REI

[2099] Anthony Clarke, Elisabeth Maier, and Hans-Udo Stadler. A report on the experiences of implementing an MT system for use in a commercial environment. Lecture Notes in Computer Science, 2499:187–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2499/24990187.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2499/24990187.pdf.

Morland:2002:GMG

[2100] Verne Morland. Getting the message in: a global company’s experience with the new generation of low-cost, high performance machine translation systems. Lecture Notes in Computer Science, 2499:195–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2499/24990195.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2499/24990195.pdf.

Rychtyckyj:2002:AMT

[2101] Nestor Rychtyckyj. An assessment of machine translation for vehicle assembly process planning at Ford Motor Company. Lecture Notes in Computer Science, 2499:207–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2499/24990207.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2499/24990207.pdf.

Abir:2002:FME

[2102] Eli Abir, Steve Klein, David Miller, and Michael Steinbaum. Fluent machines’ EliMT system. Lecture Notes in Computer Science, 2499: 216–??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2499/24990216.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2499/24990216.
pdf.

Akers:2002:LTV

[2103] Glenn A. Akers. LogoMedia TRANSLATE^{T M}, version 2.0. Lecture Notes in Computer Science, 2499:220–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2499/24990220.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2499/24990220.pdf.

Bender:2002:NIM

[2104] Howard J. Bender. Natural intelligence in a machine translation system. Lecture Notes in Computer Science, 2499:224–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2499/24990224.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2499/24990224.pdf.

Benjamin:2002:TNL

[2105] Bryce Benjamin, Kevin Knight, and Daniel Marcu. Translation by the numbers: Language weaver. Lecture Notes in Computer Science, 2499:229–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2499/24990229.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2499/24990229.pdf.

Blekhman:2002:NFP

[2106] Michael Blekhman, Andrei Kursin, and Alla Rakova. A new family of the PARS translation systems. Lecture Notes in Computer Science, 2499:232–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2499/24990232.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2499/24990232.pdf.

Dolan:2002:MMM

[2107] William B. Dolan, Jessie Pinkham, and Stephen D. Richardson. MSR-MT: The Microsoft Research Machine Translation System. Lecture Notes in Computer Science, 2499: 237–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2499/24990237.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2499/24990237.
                                        pdf.

Lavie:2002:NSS

[2108] Alon Lavie, Lori Levin, Robert Frederking, and Fabio Pianesi. The NE-SPOLE! speech-to-speech translation system. Lecture Notes in Computer Science, 2499:240–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2499/24990240.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2499/24990240.pdf.

Mitamura:2002:KMS

[2109] Teruko Mitamura, Eric Nyberg, Kathy Baker, Peter Cramer, Jeongwoo Ko, David Svoboda, and Michael Duggan. The KANTOO MT system: Controlled language checker and lexical maintenance tool. Lecture Notes in Computer Science, 2499:244–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2499/24990244.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2499/24990244.pdf.

Montgomery:2002:AST

[2110] Christine A. Montgomery and Naicong Li. Approaches to spoken translation. Lecture Notes in Computer Science, 2499:248–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2499/24990248.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2499/24990248.pdf.

Anonymous:2002:AIbw

[2111] Anonymous. Author index. Lecture Notes in Computer Science, 2499: 253–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2499/2499auth.pdf.

Farwer:2002:A

[2112] Berndt Farwer. 1 ω-automata. Lecture Notes in Computer Science, 2500:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2500/25000003.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2500/25000003.pdf.

Mazala:2002:IG

[2113] René Mazala. 2 infinite games. Lecture Notes in Computer Science, 2500:23–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2500/25000023.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2500/25000023.pdf.

Roggenbach:2002:DBA

[2114] Markus Roggenbach. 3 determinization of Büchi-automata. Lecture Notes in Computer Science, 2500: 43–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2500/25000043.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2500/25000043.
                                        pdf.

Klaedtke:2002:CBA

[2115] Felix Klaedtke. 4 complementation of Büchi automata using alternation. Lecture Notes in Computer Science, 2500:61–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2500/25000061.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2500/25000061.pdf.

Schwoon:2002:DCS

[2116] Stefan Schwoon. 5 determinization and complementation of Streett automata. Lecture Notes in Computer Science, 2500:79–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2500/25000079.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2500/25000079.pdf.

Kusters:2002:MDP

[2117] Ralf Küsters. 6 memoryless determinacy of parity games. Lecture Notes in Computer Science, 2500: 95–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2500/25000095.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2500/25000095.
pdf.

Klauck:2002:APG

[2118] Hartmut Klauck. 7 algorithms for parity games. Lecture Notes in Computer Science, 2500:107–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2500/25000107.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2500/25000107.pdf.

Niessner:2002:NTA

[2119] Frank Nießner. 8 nondeterministic tree automata. Lecture Notes in Computer Science, 2500:135–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2500/25000135.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2500/25000135.pdf.

Kirsten:2002:ATA

[2120] Daniel Kirsten. 9 alternating tree automata and parity games. Lecture Notes in Computer Science, 2500:153–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2500/25000153.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2500/25000153.pdf.

Zappe:2002:MCA

[2121] Júlia Zappe. 10 modal μ-calculus and alternating tree automata. Lecture Notes in Computer Science, 2500: 171–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2500/25000171.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2500/25000171.
                                        pdf.

Alberucci:2002:SMC

[2122] Luca Alberucci. 11 strictness of the modal μ-calculus hierarchy. Lecture Notes in Computer Science, 2500:185–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2500/25000185.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2500/25000185.pdf.

Weyer:2002:DSS

[2123] Mark Weyer. 12 decidability of S1S and S2S. Lecture Notes in Computer Science, 2500:207–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2500/25000207.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2500/25000207.pdf.

Reinhardt:2002:CTL

[2124] Klaus Reinhardt. 13 the complexity of translating logic to finite automata. Lecture Notes in Computer Science, 2500:231–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2500/25000231.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2500/25000231.pdf.

Rohde:2002:EPM

[2125] Philipp Rohde. 14 expressive power of monadic second-order logic and modal μ-calculus. Lecture Notes in Computer Science, 2500:239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2500/25000239.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2500/25000239.pdf.

Leucker:2002:PRG

[2126] Martin Leucker. 15 prefix-recognizable graphs and monadic logic. Lecture Notes in Computer Science, 2500: 263–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2500/25000263.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2500/25000263.
pdf.

Berwanger:2002:MTT

[2127] Dietmar Berwanger and Achim Blumensath. 16 the monadic theory of tree-like structures. Lecture Notes in Computer Science, 2500: 285–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2500/25000285.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2500/25000285.
pdf.

Cachat:2002:TWT

[2128] Thierry Cachat. 17 two-way tree automata solving pushdown games. Lecture Notes in Computer Science, 2500:303–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2500/25000303.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2500/25000303.pdf.

Rasch:2002:IGL

[2129] Thoralf Räsch. 18 introduction to guarded logics. Lecture Notes in Computer Science, 2500:321– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2500/25000321.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2500/25000321.
pdf.

Berwanger:2002:AGF

[2130] Dietmar Berwanger and Achim Blumensath. 19 automata for guarded fixed point logics. Lecture Notes in Computer Science, 2500:343– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2500/25000343.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2500/25000343.
pdf.

Fritz:2002:SFP

[2131] Carsten Fritz. 20 some fixed point basics. Lecture Notes in Computer Science, 2500:359–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2500/25000359.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2500/25000359.pdf.

Anonymous:2002:L

[2132] Anonymous. Literature. Lecture Notes in Computer Science, 2500: 365–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2500/2500lite.pdf.

Anonymous:2002:SI

[2133] Anonymous. Symbol index. Lecture Notes in Computer Science, 2500: 377–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2500/2500symb.pdf.

Anonymous:2002:Ia

[2134] Anonymous.  Index.  Lecture

Notes in Computer Science, 2500: 381–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2500/2500indx.pdf.

Anonymous:2002:IP

[2135] Anonymous. Introduction to Part I. Lecture Notes in Computer Science, 2500:??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2500/2500inp1.pdf.

Anonymous:2002:IPIa

[2136] Anonymous. Introduction to Part II. Lecture Notes in Computer Science, 2500:??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2500/2500inp2.pdf.

Anonymous:2002:IPIb

[2137] Anonymous. Introduction to Part III. Lecture Notes in Computer Science, 2500:??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2500/2500inp3.pdf.

Anonymous:2002:IPIc

[2138] Anonymous. Introduction to Part IV. Lecture Notes in Computer Science, 2500:??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2500/2500inp4.pdf.

Anonymous:2002:IPVa

[2139] Anonymous. Introduction to Part V.

Lecture Notes in Computer Science, 2500:??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2500/2500inp5.pdf.

Anonymous:2002:IPVb

[2140] Anonymous. Introduction to Part VI. Lecture Notes in Computer Science, 2500:??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2500/2500inp6.pdf.

Anonymous:2002:IPVc

[2141] Anonymous. Introduction to Part VII. Lecture Notes in Computer Science, 2500:??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2500/2500inp7.pdf.

Anonymous:2002:IPVd

[2142] Anonymous. Introduction to Part VIII. Lecture Notes in Computer Science, 2500:??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2500/2500inp8.pdf.

Anonymous:2002:NP

[2143] Anonymous. Notes on Part I. Lecture Notes in Computer Science, 2500:??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2500/2500nop1.pdf.

Anonymous:2002:NPIa

[2144] Anonymous. Notes on Part II. Lecture Notes in Computer Science, 2500:??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2500/2500nop2.pdf.

Anonymous:2002:NPIb

[2145] Anonymous. Notes on Part III. Lecture Notes in Computer Science, 2500:??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2500/2500nop3.pdf.

Anonymous:2002:NPIc

[2146] Anonymous. Notes on Part IV.

Lecture Notes in Computer Science, 2500:??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2500/2500nop4.pdf.

Anonymous:2002:NPVa

[2147] Anonymous. Notes on Part V.

Lecture Notes in Computer Science, 2500:??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2500/2500nop5.pdf.

Anonymous:2002:NPVb

[2148] Anonymous. Notes on Part VI.

Lecture Notes in Computer Science, 2500:??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2500/2500nop6.pdf.

Anonymous:2002:NPVc

[2149] Anonymous. Notes on Part VII.

Lecture Notes in Computer Science, 2500:??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2500/2500nop7.pdf.

Anonymous:2002:NPVd

[2150] Anonymous. Notes on Part VIII.

Lecture Notes in Computer Science, 2500:??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2500/2500nop8.pdf.

Lenstra:2002:ABF

[2151] Arjen K. Lenstra, Adi Shamir, Jim Tomlinson, and Eran Tromer. Analysis of Bernstein’s factorization circuit. Lecture Notes in Computer Science, 2501:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2501/25010001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2501/25010001.pdf.

Lucks:2002:VCS

[2152] Stefan Lucks. A variant of the Cramer– Shoup cryptosystem for groups of unknown order. Lecture Notes in Computer Science, 2501:27–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2501/25010027.htm;   http://link.
springer.de/link/service/series/
0558/papers/2501/25010027.pdf.

Bosma:2002:LBX

[2153] Wieb Bosma, James Hutton, and Eric R. Verheul. Looking beyond XTR. Lecture Notes in Computer Science, 2501:46–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2501/25010046.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2501/25010046.pdf.

Ogata:2002:BRM

[2154] Wakaha Ogata and Kaoru Kurosawa. Bounds for robust metering schemes and their relationship with A-code. Lecture Notes in Computer Science, 2501:64–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2501/25010064.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2501/25010064.pdf.

Hanaoka:2002:USA

[2155] Goichiro Hanaoka, Junji Shikata, Yumiko Hanaoka, and Hideki Imai. Unconditionally secure anonymous encryption and group authentication. Lecture Notes in Computer Science, 2501:81–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2501/25010081.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2501/25010081.pdf.

Dent:2002:AWR

[2156] Alexander W. Dent. Adapting the weaknesses of the random oracle model to the generic group model. Lecture Notes in Computer Science, 2501:100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2501/25010100.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2501/25010100.pdf.

Chang:2002:IBO

[2157] Yan-Cheng Chang, Chun-Yun Hsiao, and Chi-Jen Lu. On the impossibilities of basing one-way permutations on central cryptographic primitives. Lecture Notes in Computer Science, 2501:110–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2501/25010110.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2501/25010110.pdf.

Damgaard:2002:SHI

[2158] Ivan Damgård and Eiichiro Fujisaki. A statistically-hiding integer commitment scheme based on groups with hidden order. Lecture Notes in Computer Science, 2501:125–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2501/25010125.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2501/25010125.pdf.

Hong:2002:EOT

[2159] Dowon Hong, Ku-Young Chang, and Heuisu Ryu. Efficient oblivious transfer in the bounded-storage model. Lecture Notes in Computer Science, 2501:143–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2501/25010143.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2501/25010143.pdf.

Barkan:2002:HMW

[2160] Elad Barkan and Eli Biham. In how many ways can you write rijndael? Lecture Notes in Computer Science, 2501:160–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2501/25010160.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2501/25010160.pdf.

Park:2002:SRL

[2161] Sangwoo Park, Soo Hak Sung, Seongtaek Chee, E-Joong Yoon, and Jongin Lim. On the security of Rijndael-like structures against differential and linear cryptanalysis. Lecture Notes in Computer Science, 2501:176–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2501/25010176.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2501/25010176.pdf.

Katz:2002:TCB

[2162] Jonathan Katz and Moti Yung.

Threshold cryptosystems based on factoring. Lecture Notes in Computer Science, 2501:192–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2501/25010192.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2501/25010192.pdf.

Abe:2002:NID

[2163] Masayuki Abe, Ronald Cramer, and Serge Fehr. Non-interactive distributed-verifier proofs and proving relations among commitments. Lecture Notes in Computer Science, 2501:206–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2501/25010206.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2501/25010206.pdf.

Srinathan:2002:ASC

[2164] K. Srinathan, M. V. N. Ashwin Kumar, and C. Pandu Rangan. Asynchronous secure communication tolerating mixed adversaries. Lecture Notes in Computer Science, 2501:224–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2501/25010224.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2501/25010224.pdf.

Kim:2002:ABA

[2165] Jongsung Kim, Dukjae Moon, Wonil Lee, Seokhie Hong, Sangjin Lee, and Seokwon Jung. Amplified boomerang attack against reduced-round SHA-CAL. Lecture Notes in Computer Science, 2501:243–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2501/25010243.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2501/25010243.pdf.

Biham:2002:EDL

[2166] Eli Biham, Orr Dunkelman, and Nathan Keller. Enhancing differentiallinear cryptanalysis. Lecture Notes in Computer Science, 2501:254– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2501/25010254.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2501/25010254.
pdf.

Courtois:2002:CBC

[2167] Nicolas T. Courtois and Josef Pieprzyk.

Cryptanalysis of block ciphers with overdefined systems of equations. Lecture Notes in Computer Science, 2501:267–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://eprint.iacr.org/2002/044/
;   http://link.springer.de/link/
service/series/0558/bibs/2501/25010267.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2501/25010267.pdf.

Klimov:2002:ANC

[2168] Alexander Klimov, Anton Mityagin, and Adi Shamir. Analysis of neural cryptography. Lecture Notes in Computer Science, 2501: 288–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2501/25010288.htm;   http:

                                        //link.springer.de/link/service/
                                        series/0558/papers/2501/25010288.
                                        pdf.

Catalano:2002:HHL

[2169] Dario Catalano, Phong Q. Nguyen, and Jacques Stern. The hardness of Hensel lifting: The case of RSA and discrete logarithm. Lecture Notes in Computer Science, 2501:299–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2501/25010299.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2501/25010299.pdf.

Gaudry:2002:CCS

[2170] Pierrick Gaudry. A comparison and a combination of SST and AGM algorithms for counting points of elliptic curves in characteristic 2. Lecture Notes in Computer Science, 2501:311–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2501/25010311.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2501/25010311.pdf.

Koga:2002:GFT

[2171] Hiroki Koga. A general formula of the (t, n)-threshold visual secret sharing scheme. Lecture Notes in Computer Science, 2501:328–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2501/25010328.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2501/25010328.pdf.

DArco:2002:USR

[2172] Paolo D’Arco and Douglas R. Stinson. On unconditionally secure robust distributed key distribution centers. Lecture Notes in Computer Science, 2501:346–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2501/25010346.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2501/25010346.pdf.

Granboulan:2002:SSR

[2173] Louis Granboulan. Short signatures in the random oracle model. Lecture Notes in Computer Science, 2501:364–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2501/25010364.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2501/25010364.pdf.

Hevia:2002:PSG

[2174] Alejandro Hevia and Daniele Micciancio. The provable security of graphbased one-time signatures and extensions to algebraic signature schemes. Lecture Notes in Computer Science, 2501:379–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2501/25010379.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2501/25010379.pdf.

Bellare:2002:TSB

[2175] Mihir Bellare and Gregory Neven. Transitive signatures based on factoring and RSA. Lecture Notes in Computer Science, 2501:397– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2501/25010397.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2501/25010397.
pdf.

Abe:2002:SVK

[2176] Masayuki Abe, Miyako Ohkubo, and Koutarou Suzuki. 1-out-of-n signatures from a variety of keys. Lecture Notes in Computer Science, 2501:415–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2501/25010415.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2501/25010415.pdf.

Asano:2002:RSM

[2177] Tomoyuki Asano. A revocation scheme with minimal storage at receivers. Lecture Notes in Computer Science, 2501:433–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2501/25010433.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2501/25010433.pdf.

Golle:2002:OME

[2178] Philippe Golle, Sheng Zhong, Dan Boneh, Markus Jakobsson, and Ari Juels. Optimistic mixing for exitpolls. Lecture Notes in Computer Science, 2501:451–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2501/25010451.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2501/25010451.pdf.

Gupta:2002:ICN

[2179] Kishan Chand Gupta and Palash Sarkar. Improved construction of nonlinear resilient S-boxes. Lecture Notes in Computer Science, 2501: 466–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2501/25010466.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2501/25010466.
pdf.

Carlet:2002:UBN

[2180] Claude Carlet and Aline Gouget.

An upper bound on the number of m-resilient Boolean functions. Lecture Notes in Computer Science, 2501:484–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2501/25010484.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2501/25010484.pdf.

Bresson:2002:GDH

[2181] Emmanuel Bresson, Olivier Chevassut, and David Pointcheval. Group Diffie–Hellman key exchange secure against dictionary attacks. Lecture Notes in Computer Science, 2501: 497–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2501/25010497.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2501/25010497.
pdf.

Namprempre:2002:SCB

[2182] Chanathip Namprempre. Secure channels based on authenticated encryption schemes: a simple characterization. Lecture Notes in Computer Science, 2501:515–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2501/25010515.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2501/25010515.pdf.

Zhang:2002:IBB

[2183] Fangguo Zhang and Kwangjo Kim. ID-based blind signature and ring signature from pairings. Lecture Notes in Computer Science, 2501: 533–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2501/25010533.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2501/25010533.
pdf.

Gentry:2002:HIB

[2184] Craig Gentry and Alice Silverberg.

Hierarchical ID-based cryptography. Lecture Notes in Computer Science, 2501:548–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2501/25010548.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2501/25010548.pdf.

Yung:2002:CI

[2185] Moti Yung. Crypto-integrity. Lecture Notes in Computer Science, 2501: 567–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2501/25010567.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2501/25010567.
pdf.

Matsumoto:2002:GCS

[2186] Tsutomu Matsumoto. Gummy and conductive silicone rubber fingers. Lecture Notes in Computer Science, 2501:574–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2501/25010574.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2501/25010574.pdf.

Anonymous:2002:AIbx

[2187] Anonymous. Author index. Lecture Notes in Computer Science, 2501: 577–??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2501/2501auth.pdf.

Backes:2002:CPN

[2188] Michael Backes and Birgit Pfitzmann. Computational probabilistic non-interference. Lecture Notes in Computer Science, 2502:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2502/25020001.htm;   http://link.
springer.de/link/service/series/
0558/papers/2502/25020001.pdf.

Kurosawa:2002:BSA

[2189] Kaoru Kurosawa and Wakaha Ogata. Bit-slice auction circuit. Lecture Notes in Computer Science, 2502: 24–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2502/25020024.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2502/25020024.
pdf.

Biskup:2002:CPT

[2190] Joachim Biskup and Piero Bonatti. Confidentiality policies and their enforcement for controlled query evaluation. Lecture Notes in Computer Science, 2502:39–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2502/25020039.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2502/25020039.pdf.

Wang:2002:CBI

[2191] Lingyu Wang, Duminda Wijesekera, and Sushil Jajodia. Cardinality-based inference control in sum-only data cubes. Lecture Notes in Computer Science, 2502:55–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2502/25020055.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2502/25020055.pdf.

Smith:2002:OAP

[2192] Sean W. Smith. Outbound authentication for programmable secure coprocessors. Lecture Notes in Computer Science, 2502:72–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2502/25020072.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2502/25020072.pdf.

Gomulkiewicz:2002:HWA

[2193] Marcin Gomułkiewicz and Mirosław Kutyłowski. Hamming weight attacks on cryptographic hardware — breaking masking defense. Lecture Notes in Computer Science, 2502: 90–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2502/25020090.htm;   http:
//link.springer.de/link/service/

series/0558/papers/2502/25020090.
pdf.

Hogben:2002:FCR

[2194] Giles Hogben, Tom Jackson, and Marc Wilikens. A fully compliant research implementation of the P3P standard for privacy protection: Experiences and recommendations. Lecture Notes in Computer Science, 2502:104–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2502/25020104.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2502/25020104.pdf.

Giles:2002:ADW

[2195] James Giles, Reiner Sailer, Dinesh Verma, and Suresh Chari. Authentication for distributed Web caches. Lecture Notes in Computer Science, 2502:126–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2502/25020126.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2502/25020126.pdf.

Broadfoot:2002:ASA

[2196] Philippa Broadfoot and Gavin Lowe. Analysing a stream authentication protocol using model checking. Lecture Notes in Computer Science, 2502: 146–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2502/25020146.htm;   http:

//link.springer.de/link/service/
series/0558/papers/2502/25020146.
pdf.

Heather:2002:ET

[2197] James Heather and Steve Schneider. Equal to the task? Lecture Notes in Computer Science, 2502: 162–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2502/25020162.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2502/25020162.
pdf.

Mok:2002:TRB

[2198] Aloysius K. Mok and Weijiang Yu.

TINMAN: a resource bound security checking system for mobile code. Lecture Notes in Computer Science, 2502:178–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2502/25020178.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2502/25020178.pdf.

Santen:2002:CPR

[2199] Thomas Santen, Maritta Heisel, and Andreas Pfitzmann. Confidentialitypreserving refinement is compositional — sometimes. Lecture Notes in Computer Science, 2502: 194–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2502/25020194.htm;   http:
//link.springer.de/link/service/

                                        series/0558/papers/2502/25020194.
                                        pdf.

vonOheimb:2002:FSA

[2200] David von Oheimb and Volkmar Lotz. Formal security analysis with interacting state machines. Lecture Notes in Computer Science, 2502:212–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2502/25020212.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2502/25020212.pdf.

Koch:2002:DSG

[2201] Manuel Koch, Luigi V. Mancini, and Francesco Parisi-Presicce. Decidability of safety in graph-based models for access control. Lecture Notes in Computer Science, 2502: 229–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2502/25020229.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2502/25020229.
                                        pdf.

Wang:2002:IPD

[2202] Xinyuan Wang, Douglas S. Reeves, and S. Felix Wu. Inter-packet delay based correlation for tracing encrypted connections through stepping stones. Lecture Notes in Computer Science, 2502:244–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2502/25020244.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2502/25020244.pdf.

Lee:2002:LFD

[2203] Sin Yeung Lee, Wai Lup Low, and Pei Yuen Wong. Learning fingerprints for a database intrusion detection system. Lecture Notes in Computer Science, 2502:264–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2502/25020264.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2502/25020264.pdf.

Anonymous:2002:AIby

[2204] Anonymous. Author index. Lecture Notes in Computer Science, 2502: 281–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2502/2502ind.pdf.

Weber:2002:CMO

[2205] Ron Weber. Conceptual modelling and ontology: Possibilities and pitfalls. Lecture Notes in Computer Science, 2503:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2503/25030001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2503/25030001.pdf.

Pigneur:2002:OMB

[2206] Yves Pigneur. An ontology for m-business models. Lecture Notes in Computer Science, 2503:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2503/25030003.htm;   http://link.
springer.de/link/service/series/
0558/papers/2503/25030003.pdf.

Bider:2002:MDB

[2207] Ilia Bider and Paul Johannesson. Modeling dynamics of business processes: Key for building next generation of business information systems. Lecture Notes in Computer Science, 2503:7–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2503/25030007.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2503/25030007.pdf.

Guarino:2002:ODCa

[2208] Nicola Guarino and Luc Schneider. Ontology-driven conceptual modelling. Lecture Notes in Computer Science, 2503:10–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2503/25030010.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2503/25030010.pdf.

Henderson-Sellers:2002:AOM

[2209] Brian Henderson-Sellers. Advanced OO modelling: Metamodels and notations for the new millennium. Lecture Notes in Computer Science, 2503:11–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2503/25030011.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2503/25030011.pdf.

Guarino:2002:ODCb

[2210] Nicola Guarino and Luc Schneider. Ontology-driven conceptual modelling: Advanced concepts. Lecture Notes in Computer Science, 2503: 12–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2503/25030012.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2503/25030012.
pdf.

Grefen:2002:WME

[2211] Paul Grefen. Workflow management in electronic commerce. Lecture Notes in Computer Science, 2503:13–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2503/25030013.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2503/25030013.pdf.

Mayr:2002:DWN

[2212] Heinrich C. Mayr. Do we need an ontology of ontologies? Lecture Notes in Computer Science, 2503: 15–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/

                                        0558/bibs/2503/25030015.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2503/25030015.
                                        pdf.

VandeWeghe:2002:DCD

[2213] Nico Van de Weghe. Development of a conceptual data model for digital spatio-temporal geographical information, with application to several themes and GIS. Lecture Notes in Computer Science, 2503:16–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2503/25030016.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2503/25030016.pdf.

Silva:2002:USR

[2214] Ozifrankly Silva, Tatiana Tavares, Márcia Lucena, and Guido Souza Filho. Using semantic rules database to dynamically set up the ICSpace virtual building. Lecture Notes in Computer Science, 2503:17–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2503/25030017.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2503/25030017.pdf.

Bekaert:2002:TOOa

[2215] Pieter Bekaert and Bert Van Nuffelen. On the transformation of object oriented conceptual models to logical theories: From EROOS to IDlogic. Lecture Notes in Computer Science, 2503:18–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2503/25030018.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2503/25030018.pdf.

Thalheim:2002:CCD

[2216] Bernhard Thalheim. Component construction of database schemes. Lecture Notes in Computer Science, 2503:20–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2503/25030020.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2503/25030020.pdf.

Vigna:2002:MSE

[2217] Sebastiano Vigna. Multirelational semantics for extended entity-relationship schemata with applications. Lecture Notes in Computer Science, 2503: 35–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2503/25030035.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2503/25030035.
pdf.

Cheung:2002:MMC

[2218] S. C. Cheung, Patrick C. K. Hung, and Dickson K. W. Chiu. A metamodel for e-contract template variable dependencies facilitating e-negotiation. Lecture Notes in Computer Science, 2503:50–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2503/25030050.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2503/25030050.pdf.

Guizzardi:2002:GOF

[2219] Giancarlo Guizzardi, Heinrich Herre, and Gerd Wagner. On the general ontological foundations of conceptual modeling. Lecture Notes in Computer Science, 2503:65–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2503/25030065.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2503/25030065.pdf.

Klein:2002:FCC

[2220] Michel Klein, Atanas Kiryakov, Damyan Ognyanov, and Dieter Fensel. Finding and characterizing changes in ontologies. Lecture Notes in Computer Science, 2503:79–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2503/25030079.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2503/25030079.pdf.

Bowers:2002:SSI

[2221] Shawn Bowers, Lois Delcambre, and David Maier. Superimposed schematics: Introducing E-R structure for in-situ information selections. Lecture Notes in Computer Science, 2503:90–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2503/25030090.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2503/25030090.pdf.

Lowe:2002:WEU

[2222] David Lowe, Brian Henderson-Sellers, and Alice Gu. Web extensions to UML: Using the MVC triad. Lecture Notes in Computer Science, 2503:105–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2503/25030105.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2503/25030105.pdf.

Liu:2002:WDM

[2223] Zehua Liu, Feifei Li, and Wee Keong Ng. Wiccap data model: Mapping physical Websites to logical views. Lecture Notes in Computer Science, 2503:120–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2503/25030120.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2503/25030120.pdf.

daSilva:2002:RQS

[2224] Altigran S. da Silva, Irna M. R. Evangelista Filha, Alberto H. F. Laender, and David W. Embley. Representing and querying semistructured Web data using nested tables with structural variants. Lecture Notes in Computer Science, 2503:135–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2503/25030135.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2503/25030135.pdf.

Bekaert:2002:TOOb

[2225] Pieter Bekaert, Bert Van Nuffelen, Maurice Bruynooghe, David Gilis, and Marc Denecker. On the transformation of object-oriented conceptual models to logical theories. Lecture Notes in Computer Science, 2503:152–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2503/25030152.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2503/25030152.pdf.

Giorgini:2002:RGM

[2226] Paolo Giorgini, John Mylopoulos, Eleonora Nicchiarelli, and Roberto Sebastiani. Reasoning with goal models. Lecture Notes in Computer Science, 2503:167–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2503/25030167.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2503/25030167.pdf.

Gupta:2002:RSI

[2227] Amarnath Gupta, Bertram Ludäscher, and Maryann E. Martone. Registering scientific information sources for semantic mediation. Lecture Notes in Computer Science, 2503:

182–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2503/25030182.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2503/25030182.
pdf.

Lujan-Mora:2002:MMU

[2228] Sergio Luján-Mora, Juan Trujillo, and Il-Yeol Song. Multidimensional modeling with UML package diagrams. Lecture Notes in Computer Science, 2503:199–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2503/25030199.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2503/25030199.pdf.

Moody:2002:CEL

[2229] Daniel L. Moody. Comparative evaluation of large data model representation methods: The analyst’s perspective. Lecture Notes in Computer Science, 2503:214–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2503/25030214.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2503/25030214.pdf.

Hui:2002:ECR

[2230] Bowen Hui and Eric Yu. Extracting conceptual relationships from specialized documents. Lecture Notes in Computer Science, 2503: 232–??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2503/25030232.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2503/25030232.
                                        pdf.

Chen:2002:TMD

[2231] Jun Chen, Songting Chen, and Elke A. Rundensteiner. A transactional model for data warehouse maintenance. Lecture Notes in Computer Science, 2503:247–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2503/25030247.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2503/25030247.pdf.

Pletch:2002:ADM

[2232] Andrew Pletch, Chih yang Tsai, and Charles Matula. The account data model. Lecture Notes in Computer Science, 2503:263–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2503/25030263.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2503/25030263.pdf.

Cheng:2002:SMH

[2233] Kai Cheng and Yahiko Kambayashi. A semantic model for hypertext data caching. Lecture Notes in Computer Science, 2503:276–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2503/25030276.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2503/25030276.pdf.

Ciarlini:2002:USN

[2234] Angelo E. M. Ciarlini and Antonio L. Furtado. Understanding and simulating narratives in the context of information systems. Lecture Notes in Computer Science, 2503:291–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2503/25030291.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2503/25030291.pdf.

Hakimpour:2002:GSG

[2235] Farshad Hakimpour and Andreas Geppert. Global schema generation using formal ontologies. Lecture Notes in Computer Science, 2503: 307–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2503/25030307.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2503/25030307.
pdf.

Embley:2002:AEO

[2236] David W. Embley, Cui Tao, and Stephen W. Liddle. Automatically extracting ontologically specified data from HTML tables of unknown structure. Lecture Notes in Computer Science, 2503:322–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2503/25030322.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2503/25030322.pdf.

Cali:2002:EPD

[2237] Andrea Calì, Diego Calvanese, Giuseppe De Giacomo, and Maurizio Lenzerini. On the expressive power of data integration systems. Lecture Notes in Computer Science, 2503:338– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2503/25030338.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2503/25030338.
                                        pdf.

Parsons:2002:PBS

[2238] Jeffrey Parsons and Yair Wand. Property-based semantic reconciliation of heterogeneous information sources. Lecture Notes in Computer Science, 2503:351–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2503/25030351.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2503/25030351.pdf.

Santini:2002:CIM

[2239] Simone Santini and Amarnath Gupta. Conceptual integration of multiple partial geometric models. Lecture Notes in Computer Science, 2503: 365–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2503/25030365.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2503/25030365.
pdf.

Moody:2002:EQP

[2240] Daniel L. Moody, Guttorm Sindre, Terje Brasethvik, and Arne Sølvberg. Evaluating the quality of process models: Empirical testing of a quality framework. Lecture Notes in Computer Science, 2503:380–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2503/25030380.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2503/25030380.pdf.

Pernici:2002:DQW

[2241] Barbara Pernici and Monica Scannapieco. Data quality in Web information systems. Lecture Notes in Computer Science, 2503:397–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2503/25030397.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2503/25030397.pdf.

Cherfi:2002:CMQ

[2242] Samira Si-Said Cherfi, Jacky Akoka, and Isabelle Comyn-Wattiau. Conceptual modeling quality — from EER to UML schemas evaluation. Lecture Notes in Computer Science, 2503:414–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2503/25030414.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2503/25030414.pdf.

Elmasri:2002:CMC

[2243] Ramez Elmasri, Yu-Chi Wu, Babak Hojabri, Charley Li, and Jack Fu. Conceptual modeling for customized XML schemas. Lecture Notes in Computer Science, 2503:429–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2503/25030429.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2503/25030429.pdf.

Biskup:2002:FCM

[2244] Joachim Biskup and Ralf Menzel. A flexible cost model for abstract object-oriented database schemas. Lecture Notes in Computer Science, 2503:444–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2503/25030444.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2503/25030444.pdf.

Chen:2002:DVX

[2245] Ya Bing Chen, Tok Wang Ling, and Mong Li Lee. Designing valid XML views. Lecture Notes in Computer Science, 2503:463–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2503/25030463.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2503/25030463.pdf.

Anonymous:2002:AIbz

[2246] Anonymous. Author index. Lecture Notes in Computer Science, 2503: 479–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2503/2503ind.pdf.

Abasolo:2002:CCB

[2247] Chema Abásolo, Enric Plaza, and Josep-Lluís Arcos. Components for case-based reasoning systems. Lecture Notes in Computer Science, 2504:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2504/25040001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2504/25040001.pdf.

Aguilar-Martin:2002:ATP

[2248] J. Aguilar-Martin, N. Agell, M. Sánchez, and F. Prats. Analysis of tensions in a population based on the adequacy concept. Lecture Notes in Computer Science, 2504:17–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2504/25040017.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2504/25040017.pdf.

Monferrer:2002:QV

[2249] M. Teresa Escrig Monferrer and Francisco Toledo Lobo. Qualitative velocity. Lecture Notes in Computer Science, 2504:29–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2504/25040029.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2504/25040029.pdf.

Fiol-Roig:2002:KRM

[2250] Gabriel Fiol-Roig. Knowledge representation model for dynamic processes. Lecture Notes in Computer Science, 2504:40–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2504/25040040.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2504/25040040.pdf.

Martinez:2002:DCB

[2251] Iliana Gutiérrez Martínez and Rafael E. Bello Pérez. A decision case-based system, that reasons in uncertainty conditions. Lecture Notes in Computer Science, 2504:54–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2504/25040054.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2504/25040054.pdf.

Museros:2002:MMQ

[2252] Lledó Museros and M. Teresa Escrig Monferrer. Modeling motion qualitatively: Integrating space and time. Lecture Notes in Computer Science, 2504:64–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2504/25040064.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2504/25040064.pdf.

Ortega:2002:QCT

[2253] J. A. Ortega, F. J. Cuberos, R. M.

Gasca, M. Toro, and J. Torres. Qualitative comparison of temporal series. QSI. Lecture Notes in Computer Science, 2504:75–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2504/25040075.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2504/25040075.pdf.

Pacheco:2002:IOM

[2254] Julio Pacheco, M. Teresa Escrig Monferrer, and Francisco Toledo Lobo. Integrating 3D orientation models. Lecture Notes in Computer Science, 2504:88–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2504/25040088.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2504/25040088.pdf.

Vallespi:2002:IRC

[2255] Carles Vallespí, Elisabet Golobardes, and Joan Martí. Improving reliability in classification of microcalcifications in digital mammograms using case-based reasoning. Lecture Notes in Computer Science, 2504:101–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2504/25040101.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2504/25040101.pdf.

Valls:2002:FQP

[2256] Aïda Valls and Vicenç Torra. Fusion of qualitative preferences with different vocabularies. Lecture Notes in Computer Science, 2504: 113–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2504/25040113.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2504/25040113.
                                        pdf.

Vilcahuaman:2002:FHS

[2257] Raúl Vilcahuamán, Joaquim Meléndez, and Josep Lluis de la Rosa. FU-TURA: Hybrid system for electric load forecasting by using casebased reasoning and expert system. Lecture Notes in Computer Science, 2504:125–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2504/25040125.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2504/25040125.pdf.

Alsinet:2002:MRS

[2258] T. Alsinet, R. Béjar, A. Cabiscol, C. Fernàndez, and F. Manyà. Minimal and redundant SAT encodings for the all-interval-series problem. Lecture Notes in Computer Science, 2504:139–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2504/25040139.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2504/25040139.pdf.

Riera:2002:PCM

[2259] Daniel Riera, Miquel A. Piera, and Antoni Guasch. PN to CSP methodology: Improved bounds. Lecture Notes in Computer Science, 2504:145–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2504/25040145.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2504/25040145.pdf.

Salido:2002:DNB

[2260] Miguel A. Salido and Federico Barber. Disjunction of non-binary and numeric constraint satisfaction problems. Lecture Notes in Computer Science, 2504:159–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2504/25040159.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2504/25040159.pdf.

Alonso:2002:EAD

[2261] Laura Alonso, Irene Castellón, Karina Gibert, and Lluís Padró. An empirical approach to discourse markers by clustering. Lecture Notes in Computer Science, 2504:173–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2504/25040173.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2504/25040173.pdf.

Bacardit:2002:RII

[2262] Jaume Bacardit and Josep Maria Garrell. The role of interval initialization in a GBML system with rule representation and adaptive discrete intervals. Lecture Notes in Computer Science, 2504:184–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2504/25040184.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2504/25040184.pdf.

Bressan:2002:FSS

[2263] Marco Bressan and Jordi Vitrià. Feature subset selection in an ICA space. Lecture Notes in Computer Science, 2504:196–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2504/25040196.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2504/25040196.pdf.

Domingo-Ferrer:2002:VDB

[2264] Josep Domingo-Ferrer and Vicenç Torra. Validating distance-based record linkage with probabilistic record linkage. Lecture Notes in Computer Science, 2504:207–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2504/25040207.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2504/25040207.pdf.

Molina:2002:EFS

[2265] Luis Carlos Molina, Lluís Belanche, and Àngela Nebot. Evaluating feature selection algorithms. Lecture Notes in Computer Science, 2504: 216–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2504/25040216.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2504/25040216.
pdf.

Rodas:2002:KMK

[2266] Jorge Rodas, Karina Gibert, and J. Emilio Rojo. KDSM methodology for knowledge discovery from illstructured domains presenting very short and repeated serial measures with blocking factor. Lecture Notes in Computer Science, 2504: 228–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2504/25040228.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2504/25040228.
pdf.

Sanchez:2002:FTP

[2267] J. S. Sánchez, R. Barandela, and F. J. Ferri. On filtering the training prototypes in nearest neighbour classification. Lecture Notes in Computer Science, 2504:239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2504/25040239.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2504/25040239.pdf.

Traore:2002:CAS

[2268] Adama Traoré, Monique Polit, Cecilio Angulo, and Andreu Català. Classification of abnormal situations in a waste water treatment plant. Lecture Notes in Computer Science, 2504:249–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2504/25040249.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2504/25040249.pdf.

Bautista:2002:ACR

[2269] Joaquín Bautista and Jordi Pereira. Ant colonies for the RCPS problem. Lecture Notes in Computer Science, 2504:257–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2504/25040257.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2504/25040257.pdf.

Busquets:2002:MAA

[2270] Dídac Busquets, Ramon López de Màntaras, C. Sierra, and T. G. Dietterich. A multi-agent architecture integrating learning and fuzzy techniques for landmark-based robot navigation. Lecture Notes in Computer Science, 2504:269–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2504/25040269.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2504/25040269.pdf.

Julian:2002:SAR

[2271] V. Julian, C. Carrascosa, M. Rebollo, J. Soler, and V. Botti. SIMBA: An approach for real-time multi-agent systems. Lecture Notes in Computer Science, 2504:282–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2504/25040282.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2504/25040282.pdf.

Ontanon:2002:CCB

[2272] Santiago Ontañón and Enric Plaza.

Cooperative case bartering for casebased reasoning agents. Lecture Notes in Computer Science, 2504: 294–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2504/25040294.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2504/25040294.
pdf.

Vazquez-Salceda:2002:IOT

[2273] Javier Vázquez-Salceda, Ulises Cortés, and Julian Padget. Integrating the organ and tissue allocation processes through an agent-mediated electronic institution. Lecture Notes in Computer Science, 2504:309–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2504/25040309.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2504/25040309.pdf.

Canero:2002:MAG

[2274] Cristina Cañero, Eduardo Nofrerías, Josefina Mauri, and Petia Radeva. Modelling the acquisition geometry of a C-arm angiography system for 3D reconstruction. Lecture Notes in Computer Science, 2504: 322–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2504/25040322.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2504/25040322.
                                        pdf.

Guillamet:2002:NNM

[2275] David Guillamet and Jordi Vitrià. Nonnegative matrix factorization for face recognition. Lecture Notes in Computer Science, 2504:336–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2504/25040336.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2504/25040336.pdf.

Iniguez:2002:EPP

[2276] Pedro Iñiguez and Jan Rosell. Efficient path planning using harmonic functions computed on a non-regular grid. Lecture Notes in Computer Science, 2504:345–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2504/25040345.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2504/25040345.pdf.

Llado:2002:ITP

[2277] Xavier Lladó, Joan Martí, and Maria Petrou. Image texture prediction using colour photometric stereo. Lecture Notes in Computer Science, 2504:355–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2504/25040355.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2504/25040355.pdf.

Munoz:2002:RBC

[2278] Xavier Muñoz, Jordi Freixenet, Xavier Cufí, and Joan Martí. Regionboundary cooperative image segmentation based on active regions. Lecture Notes in Computer Science, 2504:364–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2504/25040364.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2504/25040364.pdf.

Pla:2002:SPP

[2279] Filiberto Pla and V. Javier Traver. Spherical panoramas for pan-tilt camera motion compensation in spacevariant images. Lecture Notes in Computer Science, 2504:375– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2504/25040375.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2504/25040375.
                                        pdf.

Ponsa:2002:UPG

[2280] Daniel Ponsa and Xavier Roca. Unsupervised parameterisation of Gaussian mixture models. Lecture Notes in Computer Science, 2504: 388–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2504/25040388.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2504/25040388.
                                        pdf.

Radeva:2002:BCI

[2281] Petia Radeva, Marco Bressan, A. Tovar, and Jordi Vitrià. Bayesian classification for inspection of industrial products. Lecture Notes in Computer Science, 2504:399–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2504/25040399.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2504/25040399.pdf.

Rotger:2002:IEC

[2282] D. Rotger, Petia Radeva, Josefina Mauri, and E. Fernandez-Nofrerias. Internal and external coronary vessel images registration. Lecture Notes in Computer Science, 2504:408–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2504/25040408.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2504/25040408.pdf.

Tous:2002:TDB

[2283] Francesc Tous, Agnés Borràs, Robert Benavente, Ramon Baldrich, Maria Vanrell, and Josep Lladós. Textual descriptors for browsing people by visual appearance. Lecture Notes in Computer Science, 2504: 419–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2504/25040419.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2504/25040419.
pdf.

Anonymous:2002:AIca

[2284] Anonymous. Author index. Lecture Notes in Computer Science, 2504: 431–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2504/2504auth.pdf.

Ghezzi:2002:UDE

[2285] Carlo Ghezzi. Ubiquitous, decentralized, and evolving software:

Challenges for software engineering. Lecture Notes in Computer Science, 2505:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2505/25050001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2505/25050001.pdf.

Harel:2002:CBRb

[2286] David Harel. Can behavioral requirements be executed? (and why would we want to do so?). Lecture Notes in Computer Science, 2505:6–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2505/25050006.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2505/25050006.pdf.

Milner:2002:BMM

[2287] Robin Milner. Bigraphs as a model for mobile interaction. Lecture Notes in Computer Science, 2505: 8–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2505/25050008.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2505/25050008.
pdf.

Baldan:2002:ABG

[2288] Paolo Baldan and Barbara König. Approximating the behaviour of graph transformation systems. Lecture Notes in Computer Science, 2505:

14–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2505/25050014.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2505/25050014.
pdf.

Bardohl:2002:TSA

[2289] Roswitha Bardohl, Claudia Ermel, and Julia Padberg. Transforming specification architectures by GenGED. Lecture Notes in Computer Science, 2505:30–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2505/25050030.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2505/25050030.pdf.

Bauderon:2002:DGS

[2290] Michel Bauderon and Frédérique Carrère. Decomposing graphs with symmetries. Lecture Notes in Computer Science, 2505:45–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/link/service/
series/0558/bibs/2505/25050045.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2505/25050045.pdf.

Beck:2002:GTV

[2291] J. Christopher Beck, Patrick Prosser, and Evgeny Selensky. Graph transformations for the vehicle routing and job shop scheduling problems. Lecture Notes in Computer Science, 2505:60–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2505/25050060.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2505/25050060.pdf.

Fernandez:2002:CVG

[2292] Maribel Fernández and Ian Mackie. Call-by-value λ-graph rewriting without rewriting. Lecture Notes in Computer Science, 2505:75–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.de/link/service/
                                        series/0558/bibs/2505/25050075.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2505/25050075.pdf.

Gerber:2002:TML

[2293] Anna Gerber, Michael Lawley, Kerry Raymond, Jim Steel, and Andrew Wood. Transformation: The missing link of MDA. Lecture Notes in Computer Science, 2505: 90–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2505/25050090.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2505/25050090.
                                        pdf.

Godard:2002:TDD

[2294] Emmanuel Godard, Yves Métivier, Mohamed Mosbah, and Afif Sellami. Termination detection of distributed algorithms by graph relabelling systems. Lecture Notes in Computer Science, 2505:106–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2505/25050106.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2505/25050106.pdf.

Gyapay:2002:GTT

[2295] Szilvia Gyapay, Reiko Heckel, and Dániel Varró. Graph transformation with time: Causality and logical clocks. Lecture Notes in Computer Science, 2505:120–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2505/25050120.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2505/25050120.pdf.

Habel:2002:RGT

[2296] Annegret Habel and Detlef Plump.

Relabelling in graph transformation. Lecture Notes in Computer Science, 2505:135–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2505/25050135.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2505/25050135.pdf.

Hage:2002:EGT

[2297] Jurriaan Hage, Tero Harju, and Emo Welzl. Euler graphs, triangle-free graphs and bipartite graphs in switching classes. Lecture Notes in Computer Science, 2505:148–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2505/25050148.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2505/25050148.pdf.

Heckel:2002:CTA

[2298] Reiko Heckel, Jochen Malte Küster, and Gabriele Taentzer. Confluence of typed attributed graph transformation systems. Lecture Notes in Computer Science, 2505:161–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2505/25050161.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2505/25050161.pdf.

Hoffmann:2002:ACS

[2299] Berthold Hoffmann. Abstraction and control for shapely nested graph transformation. Lecture Notes in Computer Science, 2505:177–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2505/25050177.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2505/25050177.pdf.

Klempien-Hinrichs:2002:HSB

[2300] Renate Klempien-Hinrichs. Hyperedge substitution in basic atom-replacement languages. Lecture Notes in Computer Science, 2505:192–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2505/25050192.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2505/25050192.pdf.

Knirsch:2002:DGT

[2301] Peter Knirsch and Sabine Kuske. Distributed graph transformation units. Lecture Notes in Computer Science, 2505:207–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2505/25050207.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2505/25050207.pdf.

Koch:2002:DPG

[2302] Manuel Koch and Francesco Parisi-Presicce. Describing policies with graph constraints and rules. Lecture Notes in Computer Science, 2505: 223–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2505/25050223.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2505/25050223.
pdf.

deLara:2002:CAM

[2303] Juan de Lara and Hans Vangheluwe.

Computer aided multi-paradigm modelling to process Petri-nets and statecharts. Lecture Notes in Computer Science, 2505:239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2505/25050239.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2505/25050239.pdf.

Reis:2002:UGT

[2304] Carla A. Lima Reis, Rodrigo Quites Reis, Marcelo M. de Abreu, Heribert Schlebbe, and Daltro J. Nunes. Using graph transformation as the semantical model for software process execution in the apsee environment. Lecture Notes in Computer Science, 2505:254–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2505/25050254.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2505/25050254.pdf.

Marburger:2002:GBR

[2305] André Marburger and Bernhard Westfechtel. Graph-based reengineering of telecommunication systems. Lecture Notes in Computer Science, 2505:270–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2505/25050270.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2505/25050270.pdf.

Mens:2002:FBP

[2306] Tom Mens, Serge Demeyer, and Dirk Janssens. Formalising behaviour preserving program transformations. Lecture Notes in Computer Science, 2505:286–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2505/25050286.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2505/25050286.pdf.

Minas:2002:UDD

[2307] Mark Minas and Frank Strüber.

Unparsing of diagrams with Dia-Gen. Lecture Notes in Computer Science, 2505:302–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2505/25050302.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2505/25050302.pdf.

Montanari:2002:LOG

[2308] Ugo Montanari and Leila Ribeiro.

Linear ordered graph grammars and their algebraic foundations. Lecture Notes in Computer Science, 2505:317–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2505/25050317.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2505/25050317.pdf.

Padberg:2002:RIG

[2309] Julia Padberg and Bettina E. Enders. Rule invariants in graph transformation systems for analyzing safety-critical systems. Lecture Notes in Computer Science, 2505: 334–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2505/25050334.htm;   http:
//link.springer.de/link/service/

series/0558/papers/2505/25050334.
pdf.

Pfaltz:2002:ITL

[2310] John L. Pfaltz. Incremental transformation of lattices: a key to effective knowledge discovery. Lecture Notes in Computer Science, 2505: 351–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2505/25050351.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2505/25050351.
pdf.

Szuba:2002:GGB

[2311] Janusz Szuba, Andy Schürr, and Adam Borkowski. GraCAD — graph-based tool for conceptual design. Lecture Notes in Computer Science, 2505:363–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2505/25050363.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2505/25050363.pdf.

Varro:2002:FSU

[2312] Dániel Varró. A formal semantics of UML statecharts by model transition systems. Lecture Notes in Computer Science, 2505:378–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2505/25050378.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2505/25050378.pdf.

Woo:2002:HVO

[2313] Sung-Ho Woo and Sung-Bong Yang.

Hierarchical vertex ordering. Lecture Notes in Computer Science, 2505: 393–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2505/25050393.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2505/25050393.
pdf.

Baresi:2002:TIG

[2314] Luciano Baresi and Reiko Heckel. Tutorial introduction to graph transformation: a software engineering perspective. Lecture Notes in Computer Science, 2505:402–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2505/25050402.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2505/25050402.pdf.

Harju:2002:TDC

[2315] Tero Harju, Ion Petre, and Grzegorz Rozenberg. Tutorial on DNA computing and graph transformation — computational nature of gene assembly in ciliates. Lecture Notes in Computer Science, 2505:430–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2505/25050430.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2505/25050430.pdf.

Plump:2002:TWS

[2316] Detlef Plump. TERMGRAPH 2002 — workshop survey. Lecture Notes in Computer Science, 2505: 435–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2505/25050435.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2505/25050435.
                                        pdf.

Mens:2002:WGB

[2317] Tom Mens, Andy Schürr, and Gabriele Taentzer. Workshop on graph-based tools. Lecture Notes in Computer Science, 2505:440–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2505/25050440.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2505/25050440.pdf.

Bottoni:2002:WGT

[2318] Paolo Bottoni and Mark Minas. Workshop on graph transformation and visual modeling techniques. Lecture Notes in Computer Science, 2505:445–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2505/25050445.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2505/25050445.pdf.

Heckel:2002:WSE

[2319] Reiko Heckel, Tom Mens, and Michel Wermelinger. Workshop on software evolution through transformations: Towards uniform support throughout the software life-cycle. Lecture Notes in Computer Science, 2505:450–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2505/25050450.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2505/25050450.pdf.

Courcelle:2002:WLG

[2320] Bruno Courcelle and Pascal Weil.

Workshop on logic, graph transformations and discrete structures. Lecture Notes in Computer Science, 2505:455–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2505/25050455.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2505/25050455.pdf.

Anonymous:2002:AIcb

[2321] Anonymous. Author index. Lecture Notes in Computer Science, 2505: 459–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2505/2505auth.pdf.

Nantel:2002:MRI

[2322] Jacques Nantel. More research is indeed needed in E-commerce; where were business academicians when we needed them? Lecture Notes in Computer Science, 2506: 1–??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2506/25060001.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2506/25060001.
                                        pdf.

Stokes:2002:CCM

[2323] Ellen Stokes. Cool to critical: Managing Web services now. Lecture Notes in Computer Science, 2506: 2–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2506/25060002.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2506/25060002.
                                        pdf.

Rodosek:2002:EQM

[2324] Gabi Dreo Rodosek and Metin Feridun. Enforcing QoS: Myth or reality? Lecture Notes in Computer Science, 2506:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2506/25060003.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2506/25060003.pdf.

Daly:2002:MSL

[2325] David Daly, Gautam Kar, and William H. Sanders. Modeling of service-level agreements for composed services. Lecture Notes in Computer Science, 2506:4–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2506/25060004.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2506/25060004.pdf.

Han:2002:ANM

[2326] Se-Hee Han, Myung-Sup Kim, Hong-Taek Ju, and James Won-Ki Hong. The architecture of NG-MON: a passive network monitoring system for high-speed IP networks. Lecture Notes in Computer Science, 2506:16–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2506/25060016.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2506/25060016.pdf.

Sahai:2002:ASM

[2327] Akhil Sahai, Vijay Machiraju, Mehmet Sayal, Aad van Moorsel, and Fabio Casati. Automated SLA monitoring for Web services. Lecture Notes in Computer Science, 2506: 28–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2506/25060028.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2506/25060028.
pdf.

Diao:2002:OQS

[2328] Yixin Diao, Joseph L. Hellerstein, and Sujay Parekh. Optimizing quality of service using fuzzy control. Lecture Notes in Computer Science, 2506:42–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2506/25060042.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2506/25060042.pdf.

Oh:2002:ITM

[2329] Yoon-Jung Oh, Hong-Taek Ju, Mi-Jung Choi, and James Won-Ki Hong. Interaction translation methods for XML/SNMP gateway. Lecture Notes in Computer Science, 2506: 54–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2506/25060054.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2506/25060054.
                                        pdf.

Keller:2002:MAR

[2330] Alexander Keller, Andreas Köppel, and Karl Schopmeyer. Measuring application response times with the CIM metrics model. Lecture Notes in Computer Science, 2506:66–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2506/25060066.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2506/25060066.pdf.

Rodosek:2002:QAI

[2331] Gabi Dreo Rodosek. Quality aspects in IT service management. Lecture Notes in Computer Science, 2506:82–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2506/25060082.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2506/25060082.pdf.

Anerousis:2002:RNM

[2332] Nikos Anerousis. Replication and notification management in a knowledge delivery network. Lecture Notes in Computer Science, 2506: 94–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2506/25060094.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2506/25060094.
pdf.

Liotta:2002:DSA

[2333] Antonio Liotta, Alvin Yew, Chris Bohoris, and George Pavlou. Delivering service adaptation with 3G technology. Lecture Notes in Computer Science, 2506:108–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2506/25060108.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2506/25060108.pdf.

Giblin:2002:RCB

[2334] Chris Giblin, Sean Rooney, and Anthony Bussani. Remote code browsing, a network based computation utility. Lecture Notes in Computer Science, 2506:121–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2506/25060121.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2506/25060121.pdf.

Zeng:2002:PSC

[2335] Yijun Zeng and Omar Cherkaoui. Performance study of COPS over TLS and IPsec secure session. Lecture Notes in Computer Science, 2506: 133–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2506/25060133.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2506/25060133.
                                        pdf.

Brenner:2002:CCB

[2336] Michael Brenner, Igor Radisic, and Martina Schollmeyer. A criteria catalog based methodology for analyzing service management processes. Lecture Notes in Computer Science, 2506:145–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2506/25060145.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2506/25060145.pdf.

Duflos:2002:CSP

[2337] Sandrine Duflos, Gladys Diaz, Valérie Gay, and Eric Horlait. A comparative study of policy specification languages for secure distributed applications. Lecture Notes in Computer Science, 2506:157–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2506/25060157.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2506/25060157.pdf.

Burgess:2002:TDT

[2338] Mark Burgess. Two dimensional time-series for anomaly detection and regulation in adaptive systems. Lecture Notes in Computer Science, 2506:169–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2506/25060169.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2506/25060169.pdf.

Roelle:2002:HFS

[2339] Harald Roelle. A hot-failover state machine for gateway services and its application to a Linux firewall. Lecture Notes in Computer Science, 2506:181–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2506/25060181.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2506/25060181.pdf.

Steinder:2002:DFL

[2340] Małgorzata Steinder and Adarsh Sethi. Distributed fault localization in hierarchically routed networks. Lecture Notes in Computer Science, 2506:195–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2506/25060195.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2506/25060195.pdf.

Anonymous:2002:AIcc

[2341] Anonymous. Author index. Lecture Notes in Computer Science, 2506: 209–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2506/2506auth.pdf.

Bar-Joseph:2002:EDD

[2342] Ziv Bar-Joseph, Idit Keidar, and Nancy Lynch. Early-delivery dynamic atomic broadcast. Lecture Notes in Computer Science, 2508: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2508/25080001.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2508/25080001.
                                        pdf.

Goldwasser:2002:SCA

[2343] Shafi Goldwasser and Yehuda Lindell. Secure computation without agreement. Lecture Notes in Computer Science, 2508:17–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2508/25080017.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2508/25080017.pdf.

Agrawal:2002:LRE

[2344] D. Agrawal, A. El Abbadi, A. Mostéfaoui, M. Raynal, and M. Roy. The lord of the rings: Efficient maintenance of views at data warehouses. Lecture Notes in Computer Science, 2508:33–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2508/25080033.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2508/25080033.pdf.

Mostefaoui:2002:CBP

[2345] A. Mostéfaoui, S. Rajsbaum, M. Raynal, and M. Roy. Condition-based protocols for set agreement problems. Lecture Notes in Computer Science, 2508:48–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2508/25080048.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2508/25080048.pdf.

Friedman:2002:DAR

[2346] R. Friedman, A. Mostéfaoui, S. Rajsbaum, and M. Raynal. Distributed agreement and its relation with error-correcting codes. Lecture Notes in Computer Science, 2508:63–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2508/25080063.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2508/25080063.pdf.

Koukopoulos:2002:SCU

[2347] D. Koukopoulos, M. Mavronicolas, S. Nikoletseas, and P. Spirakis. On the stability of compositions of universally stable, greedy contention-resolution protocols. Lecture Notes in Computer Science, 2508:88–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2508/25080088.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2508/25080088.pdf.

Ioannidou:2002:TSS

[2348] Kleoni Ioannidou. Transformations of self-stabilizing algorithms. Lecture Notes in Computer Science, 2508:103–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2508/25080103.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2508/25080103.pdf.

Vitanyi:2002:SWF

[2349] Paul Vitányi. Simple wait-free multireader registers. Lecture Notes in Computer Science, 2508:118– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2508/25080118.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2508/25080118.
                                        pdf.

Dutta:2002:EUC

[2350] Partha Dutta, Svend Frølund, Rachid Guerraoui, and Bastian Pochon. An efficient universal construction for message-passing systems. Lecture Notes in Computer Science, 2508: 133–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2508/25080133.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2508/25080133.
pdf.

Frolund:2002:RDB

[2351] Svend Frølund and Fernando Pedone.

Ruminations on domain-based reliable broadcast. Lecture Notes in Computer Science, 2508:148–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2508/25080148.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2508/25080148.pdf.

Stupp:2002:STD

[2352] Gideon Stupp. Stateless termination detection. Lecture Notes in Computer Science, 2508:163–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2508/25080163.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2508/25080163.pdf.

Lynch:2002:RAR

[2353] Nancy Lynch and Alex A. Shvartsman. R AMBO : a reconfigurable atomic memory service for dynamic networks. Lecture Notes in Computer Science, 2508:173–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2508/25080173.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2508/25080173.pdf.

Anker:2002:AHM

[2354] Tal Anker, Danny Dolev, and Ilya Shnayderman. Ad hoc membership for scalable applications. Lecture Notes in Computer Science, 2508:191–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2508/25080191.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2508/25080191.pdf.

Huang:2002:ABP

[2355] Yongqiang Huang and Hector Garcia-Molina. Assignment-based partitioning in a condition monitoring system. Lecture Notes in Computer Science, 2508:206–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2508/25080206.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2508/25080206.pdf.

Merritt:2002:TBS

[2356] Michael Merritt, Omer Reingold, Gadi Taubenfeld, and Rebecca N. Wright. Tight bounds for shared memory systems accessed by Byzantine processes. Lecture Notes in Computer Science, 2508:222–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2508/25080222.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2508/25080222.pdf.

Delporte-Gallet:2002:FDL

[2357] Carole Delporte-Gallet, Hugues Fauconnier, and Rachid Guerraoui. Failure detection lower bounds on registers and consensus. Lecture Notes in Computer Science, 2508:237–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2508/25080237.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2508/25080237.pdf.

Dourisboure:2002:ICR

[2358] Yon Dourisboure and Cyril Gavoille.

Improved compact routing scheme for chordal graphs. Lecture Notes in Computer Science, 2508:252– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2508/25080252.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2508/25080252.
pdf.

Harris:2002:PMW

[2359] Timothy L. Harris, Keir Fraser, and Ian A. Pratt. A practical multi-word compare-and-swap operation. Lecture Notes in Computer Science, 2508:265–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2508/25080265.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2508/25080265.pdf.

Gartner:2002:FDS

[2360] Felix C. Gärtner and Stefan Pleisch. Failure detection sequencers: Necessary and sufficient information about failures to solve predicate detection. Lecture Notes in Computer Science, 2508:280–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2508/25080280.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2508/25080280.pdf.

Chlebus:2002:BWC

[2361] Bogdan S. Chlebus, Leszek Gasieniec, Dariusz R. Kowalski, and Alex A. Shvartsman. Bounding work and communication in robust cooperative computation. Lecture Notes in Computer Science, 2508:295–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2508/25080295.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2508/25080295.pdf.

Martin:2002:MBS

[2362] Jean-Philippe Martin, Lorenzo Alvisi, and Michael Dahlin. Minimal Byzantine storage. Lecture Notes in Computer Science, 2508:311– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2508/25080311.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2508/25080311.
pdf.

Attiya:2002:WFS

[2363] Hagit Attiya and Zvi Avidor. Waitfree n-set consensus when inputs are restricted. Lecture Notes in Computer Science, 2508:326–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2508/25080326.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2508/25080326.pdf.

Herlihy:2002:ROP

[2364] Maurice Herlihy, Victor Luchangco, and Mark Moir. The repeat offender problem: a mechanism for supporting dynamic-sized, lock-free data structures. Lecture Notes in Computer Science, 2508:339–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2508/25080339.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2508/25080339.pdf.

Aguilera:2002:IFF

[2365] Marcos K. Aguilera, Gérard Le Lann, and Sam Toueg. On the impact of fast failure detectors on real-time fault-tolerant systems. Lecture Notes in Computer Science, 2508:354–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2508/25080354.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2508/25080354.pdf.

Anonymous:2002:AIcd

[2366] Anonymous. Author index. Lecture Notes in Computer Science, 2508: 371–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2508/2508auth.pdf.

Campagnolo:2002:CRR

[2367] Manuel Lameiras Campagnolo. The complexity of real recursive functions. Lecture Notes in Computer Science, 2509:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2509/25090001.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2509/25090001.pdf.

Copeland:2002:HCR

[2368] B. Jack Copeland. Hypercomputation in the Chinese Room. Lecture Notes in Computer Science, 2509:15–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2509/25090015.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2509/25090015.pdf.

DeHon:2002:VLS

[2369] André DeHon. Very large scale spatial computing. Lecture Notes in Computer Science, 2509:27–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2509/25090027.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2509/25090027.pdf.

Ogihara:2002:MMD

[2370] Mitsunori Ogihara and Animesh Ray.

The minimum-model DNA computation on a sequence of probe arrays. Lecture Notes in Computer Science, 2509:38–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2509/25090038.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2509/25090038.pdf.

Ohya:2002:ITA

[2371] Masanori Ohya. An information theoretic approach to the study of genome sequences: An application to the evolution of HIV. Lecture Notes in Computer Science, 2509:50–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2509/25090050.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2509/25090050.pdf.

Ozawa:2002:HQT

[2372] Masanao Ozawa. Halting of quantum Turing machines. Lecture Notes in Computer Science, 2509: 58–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2509/25090058.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2509/25090058.
                                        pdf.

Siwak:2002:FA

[2373] Paweł Siwak. Filtrons of automata. Lecture Notes in Computer Science, 2509:66–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2509/25090066.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2509/25090066.pdf.

Toffoli:2002:MHC

[2374] Tommaso Toffoli. A man and his computer: An issue of adaptive fitness and personal satisfaction. Lecture Notes in Computer Science, 2509:86–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2509/25090086.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2509/25090086.pdf.

Amano:2002:EDP

[2375] Masami Amano, Kazuo Iwama, and Rudy Raymond H. P. Exploiting the difference in probability calculation between quantum and probabilistic computations. Lecture Notes in Computer Science, 2509:100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2509/25090100.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2509/25090100.pdf.

Arulanandham:2002:IBS

[2376] Joshua J. Arulanandham. Implementing bead-sort with P systems. Lecture Notes in Computer Science, 2509:115–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2509/25090115.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2509/25090115.pdf.

Diaz:2002:SAR

[2377] C. Graciani Díaz, F. J. Martín Mateos, and Mario J. Pérez Jiménez. Specification of Adleman’s restricted model using an automated reasoning system: Verification of Lipton’s experiment. Lecture Notes in Computer Science, 2509:126–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2509/25090126.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2509/25090126.pdf.

Giavitto:2002:DST

[2378] Jean-Louis Giavitto and Olivier Michel. Data structure as topological spaces. Lecture Notes in Computer Science, 2509:137–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2509/25090137.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2509/25090137.pdf.

Gruau:2002:BBT

[2379] Frédéric Gruau and Philippe Malbos. The blob: a basic topological concept for “Hardware-Free” distributed computation. Lecture Notes in Computer Science, 2509:151–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2509/25090151.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2509/25090151.pdf.

Imai:2002:ELU

[2380] Katsunobu Imai, Kenji Fujita, Chuzo Iwamoto, and Kenichi Morita. Embedding a logically universal model and a self-reproducing model into number-conserving cellular automata. Lecture Notes in Computer Science, 2509:164–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2509/25090164.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2509/25090164.pdf.

Jimenez:2002:GDS

[2381] Álvaro Romero Jiménez and Mario J. Pérez Jiménez. Generation of Diophantine sets by computing P systems with external output. Lecture Notes in Computer Science, 2509:176–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2509/25090176.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2509/25090176.pdf.

Kameda:2002:ACE

[2382] Atsushi Kameda, Nobuo Matsuura, Masahito Yamamoto, and Azuma Ohuchi. An analysis of computational efficiency of DNA computing. Lecture Notes in Computer Science, 2509:191–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2509/25090191.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2509/25090191.pdf.

Kawakami:2002:CCQ

[2383] Takeshi Kawakami. Communication and computation by quantum games. Lecture Notes in Computer Science, 2509:199–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2509/25090199.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2509/25090199.pdf.

Krishna:2002:PTP

[2384] Shankara Narayanan Krishna and Raghavan Rama. On the power of tissue P systems working in the minimal mode. Lecture Notes in Computer Science, 2509:208–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2509/25090208.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2509/25090208.pdf.

Lee:2002:RCA

[2385] Jia Lee, Ferdinand Peper, Susumu Adachi, Kenichi Morita, and Shinro Mashiko. Reversible computation in asynchronous cellular automata. Lecture Notes in Computer Science, 2509:220–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2509/25090220.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2509/25090220.pdf.

Niwa:2002:GPP

[2386] Jumpei Niwa, Keiji Matsumoto, and Hiroshi Imai. General-purpose parallel simulator for quantum computing. Lecture Notes in Computer Science, 2509:230–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2509/25090230.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2509/25090230.pdf.

Shimono:2002:TAE

[2387] Toshiyuki Shimono. Towards additivity of entanglement of formation. Lecture Notes in Computer Science, 2509:252–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2509/25090252.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2509/25090252.pdf.

Sosik:2002:MCW

[2388] Petr Sosík and Jiří Matýsek. Membrane computing: When communication is enough. Lecture Notes in Computer Science, 2509:264– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2509/25090264.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2509/25090264.
pdf.

Umeo:2002:SNG

[2389] Hiroshi Umeo, Masaya Hisaoka, Koshi Michisaka, Koji Nishioka, and Masashi Maeda. Some new generalized synchronization algorithms and their implementations for large scale cellular automata. Lecture Notes in Computer Science, 2509:276–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2509/25090276.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2509/25090276.pdf.

Wiedermann:2002:RCN

[2390] Jiří Wiedermann and Jan van Leeuwen.

Relativistic computers and nonuniform complexity theory. Lecture Notes in Computer Science, 2509: 287–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2509/25090287.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2509/25090287.
                                        pdf.

Yamakami:2002:QOP

[2391] Tomoyuki Yamakami. Quantum optimization problems. Lecture Notes in Computer Science, 2509: 300–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2509/25090300.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2509/25090300.
                                        pdf.

Yamasaki:2002:AAT

[2392] Tomohiro Yamasaki, Hirotada Kobayashi, and Hiroshi Imai. An analysis of absorbing times of quantum walks. Lecture Notes in Computer Science, 2509:315–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2509/25090315.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2509/25090315.pdf.

Anonymous:2002:AIce

[2393] Anonymous. Author index. Lecture Notes in Computer Science, 2509: 331–??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2509/2509auth.pdf.

Homayounpour:2002:SMD

[2394] M. Mehdi Homayounpour. Speaker model and decision threshold updating in speaker verification. Lecture Notes in Computer Science, 2510: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2510/25100001.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2510/25100001.
pdf.

Yakhno:2002:ACH

[2395] Tatyana Yakhno and Evren Tekin.

Application of constraint hierarchy to timetabling problems. Lecture Notes in Computer Science, 2510: 11–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2510/25100011.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2510/25100011.
pdf.

Canadas:2002:IST

[2396] José Joaquín Cañadas, Isabel María del Águila, Alfonso Bosch, and Samuel Túnez. An intelligent system for therapy control in a distributed organization. Lecture Notes in Computer Science, 2510:19–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2510/25100019.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100019.pdf.

Jin:2002:DLP

[2397] Xiaoming Jin, Yuchang Lu, and Chunyi Shi. Discovering local patterns from multiple temporal sequences. Lecture Notes in Computer Science, 2510:27–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100027.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100027.pdf.

Kratky:2002:GFE

[2398] Michal Krátký, Jaroslav Pokorný, Tomáš Skopal, and Václav Snášel. The geometric framework for exact and similarity querying XML data. Lecture Notes in Computer Science, 2510:35–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100035.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100035.pdf.

Han:2002:MSE

[2399] Kyungsook Han and Hyoungguen Kim. A mobile system for extracting and visualizing protein-protein interactions. Lecture Notes in Computer Science, 2510:47–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2510/25100047.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100047.pdf.

Lee:2002:DTR

[2400] Jun Wook Lee, Yong Joon Lee, Hey Kyu Kim, Bu Hun Hwang, and Keun Ho Ryu. Discovering temporal relation rules mining from interval data. Lecture Notes in Computer Science, 2510:57–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100057.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100057.pdf.

Chalechale:2002:AIR

[2401] Abdolah Chalechale and Alfred Mertins.

An abstract image representation based on edge pixel neighborhood information (EPNI). Lecture Notes in Computer Science, 2510:67–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2510/25100067.htm;   http://link.
springer.de/link/service/series/
0558/papers/2510/25100067.pdf.

Yoon:2002:MEB

[2402] H. S. Yoon, G. S. Lee, S. H. Kim, and J. Y. Chang. Motion estimation based on temporal correlations. Lecture Notes in Computer Science, 2510:75–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2510/25100075.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100075.pdf.

Son:2002:NBM

[2403] NamRye Son, YoJin Yang, GueeSang Lee, and SungJu Park. A new boundary matching algorithm based on edge detection. Lecture Notes in Computer Science, 2510:84–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100084.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100084.pdf.

Breiteneder:2002:LWI

[2404] Christian Breiteneder, Horst Eidenberger, Geert Fiedler, and Markus Raab. Lookmark: a 2.5D Web information visualization system. Lecture Notes in Computer Science, 2510:93–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100093.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100093.pdf.

Shouraki:2002:DLS

[2405] Saeed Bagheri Shouraki and Gholamreza Haffari. Different local search algorithms in STAGE for solving bin packing problem. Lecture Notes in Computer Science, 2510:102–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100102.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100102.pdf.

Neelanarayanan:2002:PFB

[2406] V. Neelanarayanan, N. Satyanarayana, N. Subramanian, and E. Usha Rani. A prototype for functionality based network management system. Lecture Notes in Computer Science, 2510:110–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100110.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100110.pdf.

Habibi:2002:CSS

[2407] J. Habibi, E. Chiniforooshan, A. HeydarNoori, M. Mirzazadeh, M. A. Safari, and H. R. Younesy. Coaching a soccer simulation team in RoboCup environment. Lecture Notes in Computer Science, 2510:117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100117.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100117.pdf.

Long:2002:IIR

[2408] Dongyang Long. Improving information retrieval system security via an optimal maximal coding scheme. Lecture Notes in Computer Science, 2510:127–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100127.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100127.pdf.

Iglesias:2002:NSB

[2409] A. Iglesias. A new scheme based on semiconductor lasers with phaseconjugate feedback for cryptographic communications. Lecture Notes in Computer Science, 2510:135– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2510/25100135.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2510/25100135.
                                        pdf.

Kim:2002:PAA

[2410] Hyun-Sung Kim and Kee-Young Yoo. Parallel algorithm and architecture for public-key cryptosystem. Lecture Notes in Computer Science, 2510:145–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100145.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100145.pdf.

Jalili:2002:SVS

[2411] Rasool Jalili and Mohsen Rezvani. Specification and verification of security policies in firewalls. Lecture Notes in Computer Science, 2510:154–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100154.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100154.pdf.

Kim:2002:ISB

[2412] Daehee Kim and Yo-Sung Ho. Image segmentation based on shape space modeling. Lecture Notes in Computer Science, 2510:164–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100164.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100164.pdf.

Won:2002:HFS

[2413] Youjip Won, Jinyoun Park, and Sangback Ma. HERMES: File system support for multimedia streaming in information home appliance. Lecture Notes in Computer Science, 2510:172–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100172.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100172.pdf.

Suh:2002:MVR

[2414] Jae-Won Suh and Yo-Sung Ho. Motion vector recovery for error concealment based on macroblock distortion modeling. Lecture Notes in Computer Science, 2510:180–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100180.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100180.pdf.

Kim:2002:MCR

[2415] JeongWon Kim, YoungUhg Lho, YoungJu Kim, KwangBaek Kim, and SeungWon Lee. A memory copy reduction scheme for networked multimedia service in Linux kernel. Lecture Notes in Computer Science, 2510:188–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100188.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100188.pdf.

Yook:2002:HMM

[2416] Dongsuk Yook. Hidden Markov model and neural network hybrid. Lecture Notes in Computer Science, 2510:196–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100196.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100196.pdf.

Mahramian:2002:NNB

[2417] Mehran Mahramian, Nasser Yazdani, Karim Faez, and Hassan Taheri. Neural network based algorithms for IP lookup and packet classification. Lecture Notes in Computer Science, 2510:204–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100204.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100204.pdf.

Ashouri:2002:NLP

[2418] K. Ashouri, M. Amini, and M. H.

Savoji. Non-linear prediction of speech signal using artificial neural nets. Lecture Notes in Computer Science, 2510:212–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100212.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100212.pdf.

Shim:2002:WDA

[2419] Won Bo Shim and Seog Park. Web document access control using twolayered storage structures with RBAC server. Lecture Notes in Computer Science, 2510:220–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100220.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100220.pdf.

Gogolla:2002:DUD

[2420] Martin Gogolla and Mark Richters. Development of UML descriptions with USE. Lecture Notes in Computer Science, 2510:228–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2510/25100228.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100228.pdf.

Utami:2002:FID

[2421] Dewi Utami, Hadi Suwastio, and Bambang Sumadjudin. FPGA implementation of digital chaotic cryptography. Lecture Notes in Computer Science, 2510:239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100239.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100239.pdf.

Jeon:2002:SRS

[2422] Jeonghee Jeon, Choongwon Kim, and Yo-Sung Ho. Stereo for recovering sharp object boundaries. Lecture Notes in Computer Science, 2510:248–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100248.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100248.pdf.

Celik:2002:PVP

[2423] Cengiz Celik. Priority vantage points structures for similarity queries in metric spaces. Lecture Notes in Computer Science, 2510:256–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100256.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2510/25100256.pdf.

Saryazdi:2002:HPI

[2424] Saeid Saryazdi and Mostafa Jafari. A high performance image coding using uniform morphological sampling, residues classifying, and vector quantization. Lecture Notes in Computer Science, 2510:264–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100264.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100264.pdf.

Haghighat:2002:GAS

[2425] A. T. Haghighat, K. Faez, M. Dehghan, A. Mowlaei, and Y. Ghahremani. A genetic algorithm for Steiner tree optimization with multiple constraints using Prüfer number. Lecture Notes in Computer Science, 2510:272–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100272.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100272.pdf.

Sharifi:2002:NTP

[2426] Mohsen Sharifi, S. F. Noorani, and F. Orooji. A new technique for participation of non-CORBA independent persistent objects in OTS transactions. Lecture Notes in Computer Science, 2510:281–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100281.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100281.pdf.

Salimifard:2002:CMW

[2427] Khodakaram Salimifard and Mike B. Wright. Compositional modelling of workflow processes. Lecture Notes in Computer Science, 2510: 289–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2510/25100289.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2510/25100289.
                                        pdf.

Lee:2002:EMI

[2428] In-Gi Lee, Minsoo Lee, and Hwan-Seung Yong. EDMIS: Metadata interchange system for OLAP. Lecture Notes in Computer Science, 2510: 297–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2510/25100297.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2510/25100297.
                                        pdf.

Jun:2002:EMC

[2429] Woochun Jun and Le Gruenwald. An efficient method for controlling access in object-oriented databases. Lecture Notes in Computer Science, 2510:306–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2510/25100306.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100306.pdf.

Jung:2002:EIX

[2430] Jong-Seok Jung, Dong-Ik Oh, Yong-Hae Kong, and Jong-Keun Ahn. Extracting information from XML documents by reverse generating a DTD. Lecture Notes in Computer Science, 2510:314–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100314.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100314.pdf.

Hongwei:2002:MXS

[2431] Sun Hongwei, Zhang Shusheng, Zhou Jingtao, and Wang Jing. Mapping XML-schema to relational schema. Lecture Notes in Computer Science, 2510:322–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100322.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100322.pdf.

Kim:2002:FMR

[2432] Jaehoon Kim and Seog Park. Flexible modification of relational schema by X2RMap in storing XML into relations. Lecture Notes in Computer Science, 2510:330–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2510/25100330.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100330.pdf.

Hofreiter:2002:BIA

[2433] Birgit Hofreiter and Christian Huemer. B2B integration — aligning ebXML and ontology approaches. Lecture Notes in Computer Science, 2510:339–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100339.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100339.pdf.

Wang:2002:ABS

[2434] Zhou Wang and Jochen Seitz. An agent based service discovery architecture for mobile environments. Lecture Notes in Computer Science, 2510:350–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100350.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100350.pdf.

Lee:2002:LMU

[2435] Dong Chun Lee. Location management using multicasting HLR in mobile networks. Lecture Notes in Computer Science, 2510:358–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100358.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2510/25100358.pdf.

Roh:2002:PEP

[2436] Jae-Sung Roh, Chang-Heon Oh, Heau-Jo Kang, and Sung-Joon Cho. Packet error probability of multi-carrier CDMA system in fast/slow correlated fading plus interference channel. Lecture Notes in Computer Science, 2510:366–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100366.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100366.pdf.

Shin:2002:DLC

[2437] Min-Woo Shin, Nak-Keun Joo, and Hyeong-Seok Lim. A distributed low-cost dynamic multicast routing algorithm with delay constraints. Lecture Notes in Computer Science, 2510:377–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100377.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100377.pdf.

Lee:2002:NBR

[2438] Won-Jong Lee, Hyung-Rae Kim, Woo-Chan Park, Jung-Woo Kim, Tack-Don Han, and Sung-Bong Yang. A new bandwidth reduction method for distributed rendering systems. Lecture Notes in Computer Science, 2510:387–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100387.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100387.pdf.

Ahmet:2002:NNB

[2439] Çinar Ahmet and Arslan Ahmet. Neural networks based mesh generation method in 2-D. Lecture Notes in Computer Science, 2510: 395–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2510/25100395.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2510/25100395.
                                        pdf.

Ghabeli:2002:IDU

[2440] Leila Ghabeli and Hamidreza Amindavar. Image denoising using hidden Markov models. Lecture Notes in Computer Science, 2510: 402–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2510/25100402.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2510/25100402.
                                        pdf.

Marx:2002:ADD

[2441] Maarten Marx and Mehdi Dastani. Anaphoric definitions in description logic. Lecture Notes in Computer Science, 2510:410–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2510/25100410.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100410.pdf.

Hoque:2002:SQH

[2442] Abu Sayed M. Latiful Hoque. Storage and querying of high dimensional sparsely populated data in compressed representation. Lecture Notes in Computer Science, 2510: 418–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2510/25100418.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2510/25100418.
pdf.

Rodrigues:2002:GFT

[2443] Luís Rodrigues, Hugo Miranda, Ricardo Almeida, João Martins, and Pedro Vicente. The GlobData faulttolerant replicated distributed object database. Lecture Notes in Computer Science, 2510:426–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100426.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100426.pdf.

Kim:2002:LSE

[2444] Sungrim Kim and Yong ik Yoon.

A levelized schema extraction for XML document using user-defined graphs. Lecture Notes in Computer Science, 2510:434–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100434.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100434.pdf.

Nachouki:2002:EIA

[2445] Gilles Nachouki and Mohamed Quafafou. Extracting, interconnecting, and accessing heterogeneous data sources: An XML query based approach. Lecture Notes in Computer Science, 2510:442–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100442.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100442.pdf.

Beigy:2002:CAC

[2446] Hamid Beigy and Mohammadreza Meybodi. Call admission control in cellular mobile networks: a learning automata approach. Lecture Notes in Computer Science, 2510: 450–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2510/25100450.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2510/25100450.
                                        pdf.

Seok:2002:AFC

[2447] Seung-Joon Seok, Sung-Min Hong, and Chul-Hee Kang. An adaptive flow control scheme for improving TCP performance in wireless Internet. Lecture Notes in Computer Science, 2510:458–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100458.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100458.pdf.

Hong:2002:ATP

[2448] Choong Seon Hong, YingXia Niu, and Jae-Jo Lee. An adaptive TCP protocol for lossy mobile environment. Lecture Notes in Computer Science, 2510:466–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100466.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100466.pdf.

Park:2002:DIA

[2449] Sung-Yong Park, Jihoon Yang, and Yoonhee Kim. Design and implementation of application-level multicasting services over ATM networks. Lecture Notes in Computer Science, 2510:479–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100479.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100479.pdf.

Tashakori:2002:BPS

[2450] Masoud Tashakori, Mohammadreza Meybodi, and Farhad Oroumchian. Bon: The Persian stemmer. Lecture Notes in Computer Science, 2510: 487–??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2510/25100487.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2510/25100487.
                                        pdf.

Krottmaier:2002:CFF

[2451] Harald Krottmaier. Current and future features of digital journals. Lecture Notes in Computer Science, 2510:495–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100495.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100495.pdf.

Brisaboa:2002:SLP

[2452] Nieves R. Brisaboa, José R. Paramá, Miguel R. Penabad, Ángeles S. Places, and Francisco J. Rodríguez. Solving language problems in a multilingual digital library federation. Lecture Notes in Computer Science, 2510:503–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100503.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100503.pdf.

Yazdani:2002:PIL

[2453] Nasser Yazdani and Nazila Salimi. Performing IP lookup on very high line speed. Lecture Notes in Computer Science, 2510:511–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100511.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100511.pdf.

Seok:2002:SMA

[2454] Seung-Joon Seok, Sung-Min Hong, and Chul-Hee Kang. A study of marking aggregated TCP and UDP flows using generalized marking scheme. Lecture Notes in Computer Science, 2510:519–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100519.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100519.pdf.

Anastasiades:2002:TGI

[2455] Panagiotes S. Anastasiades. Towards the global information society: The enactment of a regulatory framework as a factor of transparency and social cohesion. Lecture Notes in Computer Science, 2510:527–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100527.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100527.pdf.

Santos:2002:ASW

[2456] Juan M. Santos, Luis Anido, Martín Llamas, and Judith S. Rodríguez. On the application of the Semantic Web concepts to adaptive Elearning. Lecture Notes in Computer Science, 2510:536–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100536.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100536.pdf.

Xinogalos:2002:IPE

[2457] Stelios Xinogalos and Maya Satratzemi. An integrated programming environment for teaching the object-oriented programming paradigm. Lecture Notes in Computer Science, 2510: 544–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2510/25100544.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2510/25100544.
                                        pdf.

Jahankhani:2002:CLC

[2458] Hamid Jahankhani, John A. Lynch, and Jonathan Stephenson. The current legislation covering E-learning provisions for the visually impaired in the EU. Lecture Notes in Computer Science, 2510:552–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100552.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100552.pdf.

Rodionov:2002:MCS

[2459] Alexey S. Rodionov and Hyunseung Choo. Monte Carlo soft handoff modeling. Lecture Notes in Computer Science, 2510:560–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100560.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100560.pdf.

Qu:2002:QPAb

[2460] ZhaoWei Qu, Choong Seon Hong, and Sungyoung Lee. A QoS provision architecture for mobile IPv6 over MPLS using HMAT. Lecture Notes in Computer Science, 2510:569–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100569.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100569.pdf.

Arablouei:2002:NPM

[2461] Reza Arablouei and Ayaz Ghorbani.

A new propagation model for cellular mobile radio communications in urban environments including tree effects. Lecture Notes in Computer Science, 2510:580–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100580.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100580.pdf.

Kim:2002:SMA

[2462] Seongyeol Kim and Ilyong Chung.

A secure mobile agent system applying identity-based digital signature scheme. Lecture Notes in Computer Science, 2510:588–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100588.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100588.pdf.

Moon:2002:TTA

[2463] Il-Young Moon, Jae-Sung Roh, and Sung-Joon Cho. Transmission time analysis of WAP over CDMA system using turbo code scheme. Lecture Notes in Computer Science, 2510: 597–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2510/25100597.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2510/25100597.
                                        pdf.

Anido:2002:UNT

[2464] Luis Anido, Fernando Aguado, Olga Folgueiras, Judith S. Rodríguez, Juan M. Santos, and Manuel Caeiro. On the use of new technologies in health care. Lecture Notes in Computer Science, 2510:607–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100607.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100607.pdf.

Lee:2002:HQS

[2465] Dong Chun Lee, Il-Sun Hwang, and Robert Young Chul Kim. Hybrid queuing strategy to reduce call blocking in multimedia wireless networks. Lecture Notes in Computer Science, 2510:615–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100615.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100615.pdf.

Jang:2002:DBS

[2466] Kil-Woong Jang, Sung-Ho Hwang, and Ki-Jun Han. A dynamic backoff scheme to guarantee QoS over IEEE 802.11 wireless local area networks. Lecture Notes in Computer Science, 2510:624–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100624.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100624.pdf.

Roh:2002:PES

[2467] Jae-Sung Roh, Choon-Gil Kim, and Sung-Joon Cho. Performance evaluation of serial/parallel block coded CDMA system with complex spreading in near/far multiple-access interference and multi-path Nakagami fading channel. Lecture Notes in Computer Science, 2510:632–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100632.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100632.pdf.

Beigy:2002:LAB

[2468] Hamid Beigy and Mohammadreza Meybodi. A learning automata based dynamic guard channel scheme. Lecture Notes in Computer Science, 2510:643–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100643.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100643.pdf.

Mahbub:2002:DSS

[2469] Khaled Mahbub and M. S. J. Hashmi. Dynamic system simulation on the Web. Lecture Notes in Computer Science, 2510:651–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100651.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100651.pdf.

Sanghi:2002:UPI

[2470] Dheeraj Sanghi, Pankaj Jalote, and Puneet Agarwal. Using proximity information for load balancing in geographically distributed Web server systems. Lecture Notes in Computer Science, 2510:659–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100659.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100659.pdf.

Akhgar:2002:STA

[2471] Babak Akhgar, Jawed Siddiqi, and Mehrdad Naderi. Strategie tool for assessment of the supply and demand relationship between ASPs and SMEs for competitive advantage. Lecture Notes in Computer Science, 2510:667–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100667.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100667.pdf.

Broersen:2002:TCD

[2472] Jan Broersen, Mehdi Dastani, Zhisheng Huang, and Leendert van der Torre. Trust and commitment in dynamic logic. Lecture Notes in Computer Science, 2510:677–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100677.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100677.pdf.

Bandini:2002:MHM

[2473] Stefania Bandini, Sara Manzoni, and Carla Simone. Modelling heterogeneity in multi agent systems. Lecture Notes in Computer Science, 2510:685–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100685.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100685.pdf.

Lee:2002:PAG

[2474] Yong Kyu Lee, Shin Woo Kim, Min Jung Ko, and Sung Eun Park. Pricing agents for a group buying system. Lecture Notes in Computer Science, 2510:693–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100693.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100693.pdf.

Mayoh:2002:ECM

[2475] Brian Mayoh. Evolution of cooperation in multiagent systems. Lecture Notes in Computer Science, 2510:701–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100701.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100701.pdf.

Kim:2002:DWB

[2476] Jaehun Kim and Sungwon Jung. A dynamic window-based approximate shortest path re-computation method for digital road map databases in mobile environments. Lecture Notes in Computer Science, 2510: 711–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2510/25100711.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2510/25100711.
                                        pdf.

Tkachuk:2002:WBP

[2477] Mykola V. Tkachuk, Heinrich C.

Mayr, Dmytro V. Kuklenko, and Michail D. Godlevsky. Web-based process control systems: Architectural patterns, data models, and services. Lecture Notes in Computer Science, 2510:721–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100721.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100721.pdf.

Ghandeharizadeh:2002:CTE

[2478] Shahram Ghandeharizadeh, Shan Gao, and Chris Gahagan. A comparison of techniques to estimate response time for data placement. Lecture Notes in Computer Science, 2510:730–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100730.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100730.pdf.

Hsiao:2002:URT

[2479] Sung-Jung Hsiao, Kuo-Chin Fan, Wen-Tsai Sung, and Shih-Ching Ou. Using a real-time Web-based pattern recognition system to search for component patterns database. Lecture Notes in Computer Science, 2510:739–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100739.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100739.pdf.

Seo:2002:ACA

[2480] Joo-Hwan Seo and Ki-Jun Han. An adaptive call admission control to support flow handover in wireless ad hoc networks. Lecture Notes in Computer Science, 2510:747–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100747.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100747.pdf.

Maitra:2002:DOP

[2481] Madhubanti Maitra, Ranjan Kumar Pradhan, Debasish Saha, and Amitava Mukherjee. Design of optimal LA in personal communication services network using simulated annealing technique. Lecture Notes in Computer Science, 2510:755–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100755.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100755.pdf.

Seo:2002:SBP

[2482] Dae-Hee Seo, Im-Yeong Lee, Dong ik Oh, and Doo soon Park. Secure Bluetooth Piconet using non-anonymous group key. Lecture Notes in Computer Science, 2510:766–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2510/25100766.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100766.pdf.

Lee:2002:DBA

[2483] Tae-Jin Lee and Yongsuk Kim. Differentiated bandwidth allocation and power saving for wireless personal area networks. Lecture Notes in Computer Science, 2510:778–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100778.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100778.pdf.

Nawrocki:2002:CEP

[2484] Jerzy R. Nawrocki, Michał Jasiñski, Bartosz Walter, and Adam Wojciechowski. Combining extreme programming with ISO 9000. Lecture Notes in Computer Science, 2510: 786–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2510/25100786.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2510/25100786.
pdf.

Choi:2002:CCU

[2485] Wan-Kyoo Choi, Il-Yong Chung Sung-Joo Lee, and Hong-Sang Yoon. The class cohesion using the reference graph G1 and G2. Lecture Notes in Computer Science, 2510:795–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2510/25100795.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100795.pdf.

Haberlein:2002:POI

[2486] Tobias Häberlein and Thomas Gantner. Process-oriented interactive simulation of software acquisition projects. Lecture Notes in Computer Science, 2510:806–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100806.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100806.pdf.

Espinoza:2002:ADP

[2487] Félix Agustín Castro Espinoza, Gustavo Núñez Esquer, and Joel Suárez Cansino. Automatic design patterns identification of C++ programs. Lecture Notes in Computer Science, 2510:816–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100816.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100816.pdf.

Dastani:2002:SMD

[2488] Mehdi Dastani and Leendert van der Torre. Specifying the merging of desires into goals in the context of beliefs. Lecture Notes in Computer Science, 2510:824–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2510/25100824.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100824.pdf.

Lee:2002:ICP

[2489] Deok-Gyu Lee, Im-Yeong Lee, Jong-Keun Ahn, and Yong-Hae Kong. The illegal copy protection using hidden agent. Lecture Notes in Computer Science, 2510:832–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100832.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100832.pdf.

Kim:2002:MAB

[2490] Tae-Kyung Kim, Dong-Young Lee, and T. M. Chung. Mobile agentbased misuse intrusion detection rule propagation model for distributed system. Lecture Notes in Computer Science, 2510:842–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100842.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100842.pdf.

Diaz:2002:CLD

[2491] Josep Díaz, Jaroslav Nešetřil, Maria Serna, and Dimitrios M. Thilikos. H-colorings of large degree graphs. Lecture Notes in Computer Science, 2510:850–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2510/25100850.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100850.pdf.

Lee:2002:HSG

[2492] Hyeong-Ok Lee, Jong-Seok Kim, Eunseuk Oh, and Hyeong-Seok Lim. Hyper-star graph: a new interconnection network improving the network cost of the hypercube. Lecture Notes in Computer Science, 2510:858–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100858.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100858.pdf.

Raynal:2002:SCL

[2493] Michel Raynal. Sequential consistency as lazy linearizability. Lecture Notes in Computer Science, 2510: 866–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2510/25100866.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2510/25100866.
                                        pdf.

Kim:2002:EFT

[2494] Cheol Kim, Jung Choi, and Hyeong-Seok Lim. Embedding full ternary trees into recursive circulants. Lecture Notes in Computer Science, 2510: 874–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/

0558/bibs/2510/25100874.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2510/25100874.
pdf.

Jang:2002:HPS

[2495] Kil-Woong Jang, Sun-Woo Lee, and Ki-Jun Han. A handoff priority scheme for TDMA/FDMA-based cellular networks. Lecture Notes in Computer Science, 2510:883–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100883.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100883.pdf.

Misic:2002:DTB

[2496] Jelena Mišić and Vojislav B. Mišiç.

On delay times in a Bluetooth Piconet: The impact of different scheduling policies. Lecture Notes in Computer Science, 2510:891–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100891.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100891.pdf.

Misra:2002:IPS

[2497] I. Saha Misra, S. Karmakar, M. S. Mahapatra, P. S. Bhattacharjee, D. Saha, and A. Mukhertjee. Intelligent paging strategy in 3G personal communication systems. Lecture Notes in Computer Science, 2510:899–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2510/25100899.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100899.pdf.

Lim:2002:EMA

[2498] Shinyoung Lim and Youjin Song. Experience from mobile application service framework in WIP. Lecture Notes in Computer Science, 2510:907–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100907.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100907.pdf.

Utsumi:2002:EAI

[2499] Satoshi Utsumi, Salahuddin M. S. Zabir, and Norio Shiratori. An efficient approach to improve TCP performance over wireless networks. Lecture Notes in Computer Science, 2510:916–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100916.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100916.pdf.

Pezeshk:2002:EHC

[2500] Ali Pezeshk and Babak H. Khalaj. Extended hexagonal constellations as a means of multicarrier PAPR reduction. Lecture Notes in Computer Science, 2510:926–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2510/25100926.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100926.pdf.

Kim:2002:AAC

[2501] Yoonhee Kim and Sung-Yong Park.

Adaptive application-centric management in meta-computing environments. Lecture Notes in Computer Science, 2510:937–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100937.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100937.pdf.

Park:2002:WFD

[2502] Sung-Hoon Park. The weakest failure detector for solving election problems in asynchronous distributed systems. Lecture Notes in Computer Science, 2510:945–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100945.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100945.pdf.

Fauthoux:2002:LFS

[2503] David Fauthoux and Jean-Paul Bahsoun. From lens to flow structure. Lecture Notes in Computer Science, 2510:953–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100953.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2510/25100953.pdf.

Perez:2002:ALA

[2504] Jennifer Pérez, José A. Carsí, and Isidro Ramos. ADML: a language for automatic generation of migration plans. Lecture Notes in Computer Science, 2510:965–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100965.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100965.pdf.

Korhonen:2002:CUD

[2505] Kalle Korhonen. Considerations for using domain-specific modeling in the analysis phase of software development process. Lecture Notes in Computer Science, 2510:975–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2510/25100975.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2510/25100975.pdf.

Dastani:2002:ONA

[2506] Mehdi Dastani, Virginia Dignum, and Frank Dignum. Organizations and normative agents. Lecture Notes in Computer Science, 2510: 982–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2510/25100982.htm;   http:
                                        //link.springer.de/link/service/

series/0558/papers/2510/25100982.
pdf.

Far:2002:FAB

[2507] Behrouz Homayoun Far. A framework for agent-based software development. Lecture Notes in Computer Science, 2510:990–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25100990.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25100990.pdf.

Marik:2002:AAT

[2508] V. Marík, M. Pechoucek, J. Vokrínek, and A. Ríha. Application of agent technologies in extended enterprise production planning. Lecture Notes in Computer Science, 2510: 998–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2510/25100998.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2510/25100998.
pdf.

Halavati:2002:ZAE

[2509] Ramin Halavati and Saeed Bagheri Shouraki. Zamin: An artificial ecosystem. Lecture Notes in Computer Science, 2510:1008–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2510/25101008.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2510/25101008.pdf.

Anonymous:2002:AIcf

[2510] Anonymous. Author index. Lecture Notes in Computer Science, 2510: 1017–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2510/2510auth.pdf.

Janic:2002:IQP

[2511] Milena Janic, Fernando Kuipers, Xiaoming Zhou, and Piet Van Mieghem. Implications for QoS provisioning based on traceroute measurements. Lecture Notes in Computer Science, 2511:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2511/25110003.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2511/25110003.pdf.

Mendes:2002:RDA

[2512] Paulo Mendes, Henning Schulzrinne, and Edmundo Monteiro. A receiverdriven adaptive mechanism based on the popularity of scalable sessions. Lecture Notes in Computer Science, 2511:15–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2511/25110015.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2511/25110015.pdf.

Ozkasap:2002:LSB

[2513] Öznur Özkasap. Large-scale behavior of end-to-end epidemic message loss recovery. Lecture Notes in Computer Science, 2511:25–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2511/25110025.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2511/25110025.pdf.

Sander:2002:EDS

[2514] Volker Sander and Markus Fidler.

Evaluation of a differentiated services based implementation of a premium and an Olympic service. Lecture Notes in Computer Science, 2511:36–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2511/25110036.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2511/25110036.pdf.

Seok:2002:UAS

[2515] Seung-Joon Seok, Sung-Hyuck Lee, Seok-Min Hong, and Chul-Hee Kang. Unfairness of assured service and a rate adaptive marking strategy. Lecture Notes in Computer Science, 2511:47–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2511/25110047.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2511/25110047.pdf.

Cano:2002:CBM

[2516] Maria-Dolores Cano, Fernando Cerdan, Joan Garcia-Haro, and Josemaria Malgosa-Sanahuja. Countersbased modified traffic conditioner. Lecture Notes in Computer Science, 2511:57–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2511/25110057.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2511/25110057.pdf.

Zander:2002:HQI

[2517] Sebastian Zander and Georg Carle. High quality IP video streaming with adaptive packet marking. Lecture Notes in Computer Science, 2511: 68–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2511/25110068.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2511/25110068.
                                        pdf.

Filali:2002:SSS

[2518] Fethi Filali and Walid Dabbous. SBQ: a simple scheduler for fair bandwidth sharing between unicast and multicast flows. Lecture Notes in Computer Science, 2511:78–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2511/25110078.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2511/25110078.pdf.

Ngin:2002:CTA

[2519] Hoon-Tong Ngin and Chen-Khong Tham. A control-theoretical approach for fair share computation in core-stateless networks. Lecture Notes in Computer Science, 2511: 90–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2511/25110090.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2511/25110090.
pdf.

Ossipov:2002:SSO

[2520] Evgueni Ossipov and Gunnar Karlsson. SOS: Sender oriented signaling for a simplified guaranteed service. Lecture Notes in Computer Science, 2511:100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2511/25110100.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2511/25110100.pdf.

Bauer:2002:PMB

[2521] Daniel Bauer, Sean Rooney, Paolo Scotton, Sonja Buchegger, and Ilias Iliadis. The performance of measurement-based overlay networks. Lecture Notes in Computer Science, 2511:115–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2511/25110115.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2511/25110115.pdf.

Quoitin:2002:URC

[2522] Bruno Quoitin, Steve Uhlig, and Olivier Bonaventure. Using redistribution communities for interdomain traffic engineering. Lecture Notes in Computer Science, 2511: 125–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2511/25110125.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2511/25110125.
pdf.

Abrahamsson:2002:MPR

[2523] Henrik Abrahamsson, Bengt Ahlgren, Juan Alonso, Anders Andersson, and Per Kreuger. A multi-path routing algorithm for IP networks based on flow optimisation. Lecture Notes in Computer Science, 2511: 135–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2511/25110135.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2511/25110135.
pdf.

Shen:2002:PMP

[2524] Jing Shen, Jiaoying Shi, and Jon Crowcroft. Proactive multi-path routing. Lecture Notes in Computer Science, 2511:145–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2511/25110145.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2511/25110145.pdf.

deSousa:2002:PIR

[2525] Paulo de Sousa. Premium IP:

On the road to ambient networking. Lecture Notes in Computer Science, 2511:157–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2511/25110157.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2511/25110157.pdf.

Hegde:2002:SDG

[2526] Nidhi Hegde and Konstantin E.

Avrachenkov. Service differentiation and guarantees for TCP-based elastic traffic. Lecture Notes in Computer Science, 2511:159–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2511/25110159.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2511/25110159.pdf.

Siris:2002:SDT

[2527] Vasilios A. Siris, Bob Briscoe, and Dave Songhurst. Service differentiation in third generation mobile networks. Lecture Notes in Computer Science, 2511:169–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2511/25110169.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2511/25110169.pdf.

Trimintzios:2002:PDT

[2528] Panos Trimintzios, Paris Flegkas, and George Pavlou. Policy-driven traffic engineering for intra-domain quality of service provisioning. Lecture Notes in Computer Science, 2511: 179–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2511/25110179.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2511/25110179.
pdf.

Balaouras:2002:CCS

[2529] Pantelis Balaouras and Ioannis Stavrakakis. A congestion control scheme for continuous media streaming applications. Lecture Notes in Computer Science, 2511:194–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2511/25110194.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2511/25110194.pdf.

Anjali:2002:NPS

[2530] Tricha Anjali, Caterina Scoglio, Jaudelice C. de Oliveira, Leonardo C. Chen, Ian F. Akyildiz, Jeff A. Smith, George Uhl, and Agatino Sciuto. A new path selection algorithm for MPLS networks based on available bandwidth estimation. Lecture Notes in Computer Science, 2511:205–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2511/25110205.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2511/25110205.pdf.

Sanchez-Lopez:2002:PQM

[2531] Sergio Sánchez-López, Xavier Masip-Bruin, Josep Solé-Pareta, and Jordi Domingo-Pascual. Providing QoS in MPLS-ATM integrated environment. Lecture Notes in Computer Science, 2511:215–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2511/25110215.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2511/25110215.pdf.

Pau:2002:BMF

[2532] Louis-François Pau and Jeroen Dits.

Business modeling framework for personalisation in mobile business services. Lecture Notes in Computer Science, 2511:227–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2511/25110227.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2511/25110227.pdf.

Hwang:2002:CCT

[2533] Junseok Hwang, Praveen Aravamudham, Elizabeth Liddy, Jeffrey Stanton, and Ian MacInnes. Charging control and transaction accounting mechanisms using IRTL (information resource transaction layer) middleware for P2P services. Lecture Notes in Computer Science, 2511:239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2511/25110239.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2511/25110239.pdf.

Papaioannou:2002:DIC

[2534] Thanasis G. Papaioannou and George D. Stamoulis. Design and implementation of a charging and accounting architecture for QoS-differentiated VPN services to mobile users. Lecture Notes in Computer Science, 2511: 250–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2511/25110250.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2511/25110250.
                                        pdf.

Romeral:2002:MDS

[2535] Ricardo Romeral, Alberto García-Martínez, Ana B. García, Arturo Azcorra, and Manuel Álvarez-Campana. MIRA: a distributed and scalable WAN/LAN real-time measurement platform. Lecture Notes in Computer Science, 2511:263–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2511/25110263.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2511/25110263.pdf.

Welzl:2002:TCC

[2536] Michael Welzl. Traceable congestion control. Lecture Notes in Computer Science, 2511:273–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2511/25110273.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2511/25110273.pdf.

Sureka:2002:AGV

[2537] Ashish Sureka and Peter R. Wurman. Applying the generalized Vickrey auction to pricing reliable multicasts. Lecture Notes in Computer Science, 2511:283–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2511/25110283.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2511/25110283.pdf.

Nassif:2002:ISE

[2538] Lilian N. Nassif, Luiz Henrique A. Correia, Carlos Frederico M. C. Cavalcanti, José Marcos Nogueira, Antonio A. F. Loureiro, and Geraldo Robson Mateus. InterQoS — strategy enterprise game for price and QoS negotiation on the Internet. Lecture Notes in Computer Science, 2511:293–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2511/25110293.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2511/25110293.pdf.

Lepler:2002:RPU

[2539] Jörg H. Lepler and Karsten Neuhoff.

Resource pricing under a marketbased reservation protocol. Lecture Notes in Computer Science, 2511: 303–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2511/25110303.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2511/25110303.
                                        pdf.

Fulp:2002:EIN

[2540] Errin W. Fulp and Douglas S. Reeves. The economic impact of network pricing intervals. Lecture Notes in Computer Science, 2511: 315–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2511/25110315.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2511/25110315.
                                        pdf.

Jagannathan:2002:PRP

[2541] Srinivasan Jagannathan and Kevin C. Almeroth. Pricing and resource provisioning for delivering E-content on-demand with multiple levels-of-service. Lecture Notes in Computer Science, 2511:325–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2511/25110325.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2511/25110325.pdf.

Bagnulo:2002:PAA

[2542] Marcelo Bagnulo, Bernardo Alarcos, María Calderón, and Marifeli Sedano. Providing authentication & authorization mechanisms for active service charging. Lecture Notes in Computer Science, 2511:337–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2511/25110337.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2511/25110337.pdf.

Anonymous:2002:AIcg

[2543] Anonymous. Author index. Lecture Notes in Computer Science, 2511: 347–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2511/2511indx.pdf.

Alpuente:2002:IDS

[2544] M. Alpuente, S. Escobar, B. Gramlich, and S. Lucas. Improving on-demand strategy annotations. Lecture Notes in Computer Science, 2514:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2514/25140001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2514/25140001.pdf.

Apt:2002:FOL

[2545] K. R. Apt and C. F. M. Vermeulen. First-order logic as a constraint programming language. Lecture Notes in Computer Science, 2514:19–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2514/25140019.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2514/25140019.pdf.

Autexier:2002:MFS

[2546] Serge Autexier and Dieter Hutter. Maintenance of formal software developments by stratified verification. Lecture Notes in Computer Science, 2514:36–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2514/25140036.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2514/25140036.pdf.

Beckmann:2002:NUM

[2547] Arnold Beckmann. A note on universal measures for weak implicit computational complexity. Lecture Notes in Computer Science, 2514: 53–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2514/25140053.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2514/25140053.
                                        pdf.

Bollig:2002:ECM

[2548] Benedikt Bollig, Martin Leucker, and Philipp Lucas. Extending compositional message sequence graphs. Lecture Notes in Computer Science, 2514:68–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2514/25140068.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2514/25140068.pdf.

Brotherston:2002:SIU

[2549] James Brotherston, Anatoli Degtyarev, Michael Fisher, and Alexei Lisitsa. Searching for invariants using temporal resolution. Lecture Notes in Computer Science, 2514:86–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2514/25140086.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2514/25140086.pdf.

Castellini:2002:PPF

[2550] Claudio Castellini and Alan Smaill.

Proof planning for feature interactions: a preliminary report. Lecture Notes in Computer Science, 2514: 102–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2514/25140102.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2514/25140102.
pdf.

Dastani:2002:EBF

[2551] Mehdi Dastani and Leendert van der Torre. An extension of BDI_{CTL} with functional dependencies and components. Lecture Notes in Computer Science, 2514:115–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2514/25140115.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2514/25140115.pdf.

Dowek:2002:BLP

[2552] Gilles Dowek, Thérèse Hardin, and Claude Kirchner. Binding logic: Proofs and models. Lecture Notes in Computer Science, 2514:130– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2514/25140130.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2514/25140130.
                                        pdf.

Edelkamp:2002:DAT

[2553] Stefan Edelkamp and Peter Leven. Directed automated theorem proving. Lecture Notes in Computer Science, 2514:145–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2514/25140145.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2514/25140145.pdf.

Fan:2002:FSB

[2554] Xiaocong Fan and John Yen. A framework for splitting BDI agents. Lecture Notes in Computer Science, 2514:160–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2514/25140160.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2514/25140160.pdf.

Ferrari:2002:CDE

[2555] Mauro Ferrari, Camillo Fiorentini, and Guido Fiorino. On the complexity of disjunction and explicit definability properties in some intermediate logics. Lecture Notes in Computer Science, 2514:175–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2514/25140175.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2514/25140175.pdf.

Fontaine:2002:UBC

[2556] Pascal Fontaine and E. Pascal Gribomont. Using BDDs with combinations of theories. Lecture Notes in Computer Science, 2514: 190–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2514/25140190.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2514/25140190.
pdf.

Grandi:2002:EDL

[2557] Fabio Grandi. On expressive description logics with composition of roles in number restrictions. Lecture Notes in Computer Science, 2514:202–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2514/25140202.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2514/25140202.pdf.

Greco:2002:QOD

[2558] Gianluigi Greco, Sergio Greco, Irina Trubtsyna, and Ester Zumpano. Query optimization of disjunctive databases with constraints through binding propagation. Lecture Notes in Computer Science, 2514:216–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2514/25140216.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2514/25140216.pdf.

Guglielmi:2002:NCE

[2559] Alessio Guglielmi and Lutz Straßburger. A non-commutative extension of MELL. Lecture Notes in Computer Science, 2514:231–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2514/25140231.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2514/25140231.pdf.

Guller:2002:PSF

[2560] Dušan Guller. Procedural semantics for fuzzy disjunctive programs. Lecture Notes in Computer Science, 2514:247–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2514/25140247.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2514/25140247.pdf.

Kupferman:2002:PS

[2561] Orna Kupferman, Nir Piterman, and Moshe Y. Vardi. Pushdown specifications. Lecture Notes in Computer Science, 2514:262–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2514/25140262.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2514/25140262.pdf.

Kutsia:2002:TPS

[2562] Temur Kutsia. Theorem proving with sequence variables and flexible arity symbols. Lecture Notes in Computer Science, 2514:278–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2514/25140278.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2514/25140278.pdf.

McIver:2002:GPQ

[2563] A. K. McIver and C. C. Morgan. Games, probability, and the quantitative μ-calculus qMμ. Lecture Notes in Computer Science, 2514:292–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2514/25140292.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2514/25140292.pdf.

Niehren:2002:PTR

[2564] Joachim Niehren and Mateu Villaret. Parallelism and tree regular constraints. Lecture Notes in Computer Science, 2514:311–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2514/25140311.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2514/25140311.pdf.

Preining:2002:GLC

[2565] Norbert Preining. Gödel logics and Cantor-Bendixon analysis. Lecture Notes in Computer Science, 2514:327–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2514/25140327.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2514/25140327.pdf.

Richardson:2002:SPP

[2566] Julian Richardson. A semantics for proof plans with applications to interactive proof planning. Lecture Notes in Computer Science, 2514:337–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2514/25140337.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2514/25140337.pdf.

Santo:2002:IBF

[2567] José Espírito Santo. An isomorphism between a fragment of sequent calculus and an extension of natural deduction. Lecture Notes in Computer Science, 2514:352–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2514/25140352.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2514/25140352.pdf.

Siekmann:2002:PDI

[2568] J. Siekmann, C. Benzmüller, A. Fiedler, A. Meier, and M. Pollet. Proof development with ΩMEGA: \sqrt{2} is irrational. Lecture Notes in Computer Science, 2514:367–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2514/25140367.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2514/25140367.pdf.

Strassburger:2002:LSL

[2569] Lutz Straßburger. A local system for linear logic. Lecture Notes in Computer Science, 2514:388–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2514/25140388.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2514/25140388.pdf.

Strecker:2002:ITC

[2570] Martin Strecker. Investigating typecertifying compilation with isabelle. Lecture Notes in Computer Science, 2514:403–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2514/25140403.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2514/25140403.pdf.

Syme:2002:ATS

[2571] Don Syme and Andrew D. Gordon. Automating type soundness proofs via decision procedures and guided reductions. Lecture Notes in Computer Science, 2514:418–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2514/25140418.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2514/25140418.pdf.

Tessaris:2002:ASR

[2572] Sergio Tessaris and Ian Horrocks. Abox satisfiability reduced to terminological reasoning in expressive description logics. Lecture Notes in Computer Science, 2514:435–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2514/25140435.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2514/25140435.pdf.

Vaucheret:2002:FPb

[2573] Claudio Vaucheret, Sergio Guadarrama, and Susana Muñoz. Fuzzy prolog. Lecture Notes in Computer Science, 2514:450–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2514/25140450.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2514/25140450.pdf.

Anonymous:2002:AIch

[2574] Anonymous. Author index. Lecture Notes in Computer Science, 2514: 465–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2514/2514auth.pdf.

Hernandez:2002:OTS

[2575] Enrique Hernández, Joan Vila, Sergio Saez, and Silvia Terrasa. Optimal traffic shaping with deterministic guarantees using GPS scheduling. Lecture Notes in Computer Science, 2515:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2515/25150001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2515/25150001.pdf.

Arias:2002:EVS

[2576] J. R. Arias, F. J. Suárez, D. F.

García, X. X. García, and V. G. García. Evaluation of video server capacity with regard to quality of the service in interactive news-on-demand systems. Lecture Notes in Computer Science, 2515:15–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2515/25150015.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2515/25150015.pdf.

Lacan:2002:WFS

[2577] J. Lacan, L. Lancérica, and L. Dairaine. When FEC speed up data access in P2P networks. Lecture Notes in Computer Science, 2515: 26–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2515/25150026.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2515/25150026.
pdf.

Eide:2002:SIM

[2578] Viktor S. Wold Eide, Frank Eliassen, Ole-Christoffer Granmo, and OlavLysne. Scalable independent multi-level distribution in multimedia content analysis. Lecture Notes in Computer Science, 2515:37–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2515/25150037.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2515/25150037.pdf.

Garcia:2002:STD

[2579] A. B. García, E. García, M. Álvarez-Campana, J. Berrocal, and E. Vázquez. A simulation tool for dimensioning and performance evaluation of the UMTS terrestrial radio access network. Lecture Notes in Computer Science, 2515:49–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2515/25150049.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2515/25150049.pdf.

Indulska:2002:VHB

[2580] Jadwiga Indulska and Sasitharan Balasubramaniam. Vertical handover based adaptation for multimedia applications in pervasive systems. Lecture Notes in Computer Science, 2515:61–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2515/25150061.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2515/25150061.pdf.

Bagnulo:2002:ADI

[2581] Marcelo Bagnulo, Ignacio Soto, Alberto García-Martínez, and Arturo Azcorra. Avoiding DAD for improving real-time communication in MIPv6 environments. Lecture Notes in Computer Science, 2515:73–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2515/25150073.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2515/25150073.pdf.

Eikerling:2002:DOC

[2582] Heinz-Josef Eikerling and Frank Berger. Design of OSGi compatible middleware components for mobile multimedia applications. Lecture Notes in Computer Science, 2515: 80–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/

                                        0558/bibs/2515/25150080.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2515/25150080.
                                        pdf.

Ngoc:2002:FQA

[2583] Nam Pham Ngoc, Gauthier Lafruit, Geert Deconinck, and Rudy Lauwereins. A fast QoS adaptation algorithm for MPEG-4 multimedia applications. Lecture Notes in Computer Science, 2515:92–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2515/25150092.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2515/25150092.pdf.

Kim:2002:GAB

[2584] Eun Yi Kim and Se Hyun Park. A genetic algorithm-based segmentation for automatic VOP generation. Lecture Notes in Computer Science, 2515:106–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2515/25150106.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2515/25150106.pdf.

Lim:2002:AMS

[2585] Dong-Keun Lim and Yo-Sung Ho. Adaptive motion search based on block difference and motion magnitude. Lecture Notes in Computer Science, 2515:118–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2515/25150118.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2515/25150118.pdf.

Owezarski:2002:DII

[2586] Philippe Owezarski. Does IPv6 improve the scalability of the Internet? Lecture Notes in Computer Science, 2515:130–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2515/25150130.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2515/25150130.pdf.

Exposito:2002:DNQ

[2587] E. Exposito, P. Sénac, D. Garduno, M. Diaz, and M. Urueña. Deploying new QoS aware transport services. Lecture Notes in Computer Science, 2515:141–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2515/25150141.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2515/25150141.pdf.

Dorfinger:2002:RCL

[2588] Peter Dorfinger, Christof Brandauer, and Ulrich Hofmann. A rate controller for long-lived TCP flows. Lecture Notes in Computer Science, 2515:154–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2515/25150154.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2515/25150154.pdf.

Cagnazzo:2002:LCS

[2589] Marco Cagnazzo, Giovanni Poggi, and Luisa Verdoliva. Low-complexity scalable video coding through table lookup VQ and index coding. Lecture Notes in Computer Science, 2515:166–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2515/25150166.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2515/25150166.pdf.

Leszczuk:2002:AVS

[2590] Mikolaj Leszczuk and Zdzislaw Papir. Accuracy vs. speed trade-off in detecting of shots in video content for abstracting digital video libraries. Lecture Notes in Computer Science, 2515:176–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2515/25150176.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2515/25150176.pdf.

Mostefaoui:2002:DVD

[2591] Ahmed Mostefaoui and Loic Favory. Distributed video documents indexing and content-based retrieving. Lecture Notes in Computer Science, 2515:190–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2515/25150190.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2515/25150190.pdf.

Zuehlke:2002:VBB

[2592] Mario Zuehlke and Hartmut Koenig.

Voting based bandwidth management in multiparty video conferences. Lecture Notes in Computer Science, 2515:202–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2515/25150202.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2515/25150202.pdf.

Ryu:2002:CNR

[2593] Yeonseung Ryu. Considering non-real-time traffic in real-time packet scheduling. Lecture Notes in Computer Science, 2515:216–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2515/25150216.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2515/25150216.pdf.

Lijding:2002:FRT

[2594] Maria Eva Lijding, Pierre Jansen, and Sape Mullender. A flexible real-time hierarchical multimedia archive. Lecture Notes in Computer Science, 2515:229–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2515/25150229.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2515/25150229.pdf.

Sargento:2002:TBS

[2595] Susana Sargento and Rui Valadas. Tradeoffs between signaling and resource utilization in DiffServ networks with hierarchical aggregation. Lecture Notes in Computer Science, 2515:241–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2515/25150241.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2515/25150241.pdf.

Park:2002:RDE

[2596] Kyung-Joon Park, Won Jong Jeon, Klara Nahrstedt, Tamer Basar, and Chong-Ho Choi. Robust delay estimation for Internet multimedia applications. Lecture Notes in Computer Science, 2515:255–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2515/25150255.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2515/25150255.pdf.

Park:2002:SMM

[2597] Jin Park, Young-Joo Suh, and Shin-Gak Kang. Supporting mobile multicast in mobile networks by considering host mobility. Lecture Notes in Computer Science, 2515: 263–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/

0558/bibs/2515/25150263.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2515/25150263.
pdf.

Lin:2002:FAS

[2598] Jinsong Lin, Glenn Glazer, Richard Guy, and Rajive Bagrodia. Fast asynchronous streaming handoff. Lecture Notes in Computer Science, 2515: 274–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2515/25150274.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2515/25150274.
pdf.

Choi:2002:PCA

[2599] Jungwan Choi and Youjip Won. Power constraints: Another dimension of complexity in continuous media playback. Lecture Notes in Computer Science, 2515:288–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2515/25150288.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2515/25150288.pdf.

Song:2002:RSM

[2600] Minseok Song and Heonshik Shin.

Replica striping for multi-resolution video servers. Lecture Notes in Computer Science, 2515:300–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/link/service/
series/0558/bibs/2515/25150300.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2515/25150300.pdf.

On:2002:AQR

[2601] Giwon On, Jens Schmitt, and Ralf Steinmetz. On availability QoS for replicated multimedia service and content. Lecture Notes in Computer Science, 2515:313–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2515/25150313.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2515/25150313.pdf.

Lohse:2002:OMA

[2602] Marco Lohse, Michael Repplinger, and Philipp Slusallek. An open middleware architecture for networkintegrated multimedia. Lecture Notes in Computer Science, 2515: 327–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2515/25150327.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2515/25150327.
                                        pdf.

Chong:2002:SAB

[2603] Cheun Ngen Chong, René van Buuren, Pieter H. Hartel, and Geert Kleinhuis. Security attributes based digital rights management. Lecture Notes in Computer Science, 2515:339–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2515/25150339.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2515/25150339.pdf.

Uruena:2002:ANA

[2604] Manuel Urueña, David Larrabeiti, María Calderón, Arturo Azcorra, Jens E. Kristensen, Lars Kroll Kristensen, Ernesto Exposito, David Garduno, and Michel Diaz. An active network approach to support multimedia relays. Lecture Notes in Computer Science, 2515:353–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2515/25150353.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2515/25150353.pdf.

Schulzrinne:2002:IMS

[2605] Henning Schulzrinne. Internet multimedia — streaming media (tutorial). Lecture Notes in Computer Science, 2515:365–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2515/25150365.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2515/25150365.pdf.

Smirnow:2002:IMT

[2606] Mikhail Smirnow. Internet middleware (tutorial). Lecture Notes in Computer Science, 2515:366–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2515/25150366.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2515/25150366.pdf.

Owezarski:2002:INM

[2607] Philippe Owezarski. IP network monitoring and measurements: Techniques and experiences (tutorial). Lecture Notes in Computer Science, 2515:367–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2515/25150367.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2515/25150367.pdf.

Pereira:2002:MRS

[2608] Fernando Pereira. Media representation standards for the new millennium (tutorial). Lecture Notes in Computer Science, 2515:368–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2515/25150368.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2515/25150368.pdf.

Anonymous:2002:AIci

[2609] Anonymous. Author index. Lecture Notes in Computer Science, 2515: 371–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2515/2515auth.pdf.

Yung:2002:DLC

[2610] Kwong H. Yung. Detecting long connection chains of interactive terminal sessions. Lecture Notes in Computer Science, 2516:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2516/25160001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2516/25160001.pdf.

Donoho:2002:MSS

[2611] David L. Donoho, Ana Georgina Flesia, Umesh Shankar, Vern Paxson, Jason Coit, and Stuart Staniford. Multiscale stepping-stone detection: Detecting pairs of jittered interactive streams by exploiting maximum tolerable delay. Lecture Notes in Computer Science, 2516:17–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2516/25160017.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2516/25160017.pdf.

Apap:2002:DMS

[2612] Frank Apap, Andrew Honig, Shlomo Hershkop, Eleazar Eskin, and Sal Stolfo. Detecting malicious software by monitoring anomalous Windows registry accesses. Lecture Notes in Computer Science, 2516: 36–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2516/25160036.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2516/25160036.
pdf.

Tan:2002:UAB

[2613] Kymie M. C. Tan, Kevin S. Killourhy, and Roy A. Maxion. Undermining an anomaly-based intrusion detection system using common exploits. Lecture Notes in Computer Science, 2516:54–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2516/25160054.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2516/25160054.pdf.

Ning:2002:AII

[2614] Peng Ning, Yun Cui, and Douglas S. Reeves. Analyzing intensive intrusion alerts via correlation. Lecture Notes in Computer Science, 2516:74–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2516/25160074.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2516/25160074.pdf.

Porras:2002:MIB

[2615] Phillip A. Porras, Martin W. Fong, and Alfonso Valdes. A mission-impact-based approach to INFOSEC alarm correlation. Lecture Notes in Computer Science, 2516:95–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2516/25160095.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2516/25160095.pdf.

Morin:2002:MFD

[2616] Benjamin Morin, Ludovic Mé, Hervé Debar, and Mireille Ducassé. M2D2: a formal data model for IDS alert correlation. Lecture Notes in Computer Science, 2516:115–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2516/25160115.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2516/25160115.pdf.

Johnston:2002:DLF

[2617] Steven R. Johnston. Development of a legal framework for intrusion detection. Lecture Notes in Computer Science, 2516:138–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2516/25160138.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2516/25160138.pdf.

Just:2002:LUA

[2618] James E. Just, James C. Reynolds, Larry A. Clough, Melissa Danforth, Karl N. Levitt, Ryan Maglich, and Jeff Rowe. Learning unknown attacks — A start. Lecture Notes in Computer Science, 2516:158–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2516/25160158.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2516/25160158.pdf.

Debar:2002:EDC

[2619] Hervé Debar and Benjamin Morin. Evaluation of the diagnostic capabilities of commercial intrusion detection systems. Lecture Notes in Computer Science, 2516:177–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2516/25160177.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2516/25160177.pdf.

Goldman:2002:SMI

[2620] Robert P. Goldman. A stochastic model for intrusions. Lecture Notes in Computer Science, 2516: 199–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2516/25160199.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2516/25160199.
                                        pdf.

Gorodetski:2002:AAC

[2621] Vladimir Gorodetski and Igor Kotenko. Attacks against computer network: Formal grammar-based framework and simulation tool. Lecture Notes in Computer Science, 2516:219– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2516/25160219.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2516/25160219.
                                        pdf.

Hall:2002:CVH

[2622] Mike Hall and Kevin Wiley. Capacity verification for high speed network intrusion detection systems. Lecture Notes in Computer Science, 2516:239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2516/25160239.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2516/25160239.pdf.

Lee:2002:PAR

[2623] Wenke Lee, João B. D. Cabrera, Ashley Thomas, Niranjan Balwalli, Sunmeet Saluja, and Yi Zhang. Performance adaptation in realtime intrusion detection systems. Lecture Notes in Computer Science, 2516:252–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2516/25160252.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2516/25160252.pdf.

Toth:2002:ABO

[2624] Thomas Toth and Christopher Kruegel.

Accurate buffer overflow detection via abstract payload execution. Lecture Notes in Computer Science, 2516: 274–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2516/25160274.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2516/25160274.
pdf.

Zimmermann:2002:IRF

[2625] Jacob Zimmermann, Ludovic Mé, and Christophe Bidan. Introducing reference flow control for detecting intrusion symptoms at the OS level. Lecture Notes in Computer Science, 2516:292–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2516/25160292.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2516/25160292.pdf.

Lippmann:2002:EIV

[2626] Richard Lippmann, Seth Webster, and Douglas Stetson. The effect of identifying vulnerabilities and patching software on the utility of network intrusion detection. Lecture Notes in Computer Science, 2516:307–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2516/25160307.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2516/25160307.pdf.

Anonymous:2002:AIcj

[2627] Anonymous. Author index. Lecture Notes in Computer Science, 2516: 327–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2516/2516auth.pdf.

Melham:2002:ASI

[2628] Thomas F. Melham and Robert B. Jones. Abstraction by symbolic indexing transformations. Lecture Notes in Computer Science, 2517: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2517/25170001.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2517/25170001.
pdf.

Das:2002:CEB

[2629] Satyaki Das and David L. Dill.

Counter-example based predicate discovery in predicate abstraction. Lecture Notes in Computer Science, 2517:19–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2517/25170019.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2517/25170019.pdf.

Chauhan:2002:AAR

[2630] Pankaj Chauhan, Edmund Clarke, James Kukula, Samir Sapra, Helmut Veith, and Dong Wang. Automated abstraction refinement for model checking large state spaces using SAT based conflict analysis. Lecture Notes in Computer Science, 2517:33–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2517/25170033.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2517/25170033.pdf.

Moon:2002:SCF

[2631] In-Ho Moon, Hee Hwan Kwak, James Kukula, Thomas Shiple, and Carl Pixley. Simplifying circuits for formal verification using parametric representation. Lecture Notes in Computer Science, 2517:52–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2517/25170052.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2517/25170052.pdf.

Yang:2002:GST

[2632] Jin Yang and Carl-Johan H. Seger. Generalized symbolic trajectory evaluation – abstraction in action. Lecture Notes in Computer Science, 2517:70–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2517/25170070.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2517/25170070.pdf.

Somenzi:2002:ASS

[2633] Fabio Somenzi, Kavita Ravi, and Roderick Bloem. Analysis of symbolic SCC hull algorithms. Lecture Notes in Computer Science, 2517: 88–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2517/25170088.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2517/25170088.
                                        pdf.

Wang:2002:SDD

[2634] Chao Wang and Gary D. Hachtel.

Sharp disjunctive decomposition for language emptiness checking. Lecture Notes in Computer Science, 2517: 106–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2517/25170106.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2517/25170106.
pdf.

Aagaard:2002:RMS

[2635] Mark D. Aagaard, Nancy A. Day, and Meng Lou. Relating multi-step and single-step microprocessor correctness statements. Lecture Notes in Computer Science, 2517:123–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2517/25170123.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2517/25170123.pdf.

Lahiri:2002:MVO

[2636] Shuvendu K. Lahiri, Sanjit A. Seshia, and Randal E. Bryant. Modeling and verification of out-of-order microprocessors in UCLID. Lecture Notes in Computer Science, 2517:142–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2517/25170142.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2517/25170142.pdf.

Strichman:2002:SPL

[2637] Ofer Strichman. On solving Presburger and linear arithmetic with SAT. Lecture Notes in Computer Science, 2517:160–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2517/25170160.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2517/25170160.pdf.

Ganesh:2002:DPA

[2638] Vijay Ganesh, Sergey Berezin, and David L. Dill. Deciding Presburger arithmetic by model checking and comparisons with other methods. Lecture Notes in Computer Science, 2517:171–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2517/25170171.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2517/25170171.pdf.

Ayari:2002:QDQ

[2639] Abdelwaheb Ayari and David Basin. QUBOS: Deciding quantified Boolean logic using propositional satisfiability solvers. Lecture Notes in Computer Science, 2517:187–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2517/25170187.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2517/25170187.pdf.

Penna:2002:ETL

[2640] Giuseppe Della Penna, Benedetto Intrigila, Enrico Tronci, and Marisa Venturini Zilli. Exploiting transition locality in the disk based mur_{φ} verifier. Lecture Notes in Computer Science, 2517:202–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2517/25170202.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2517/25170202.pdf.

Sole:2002:TTC

[2641] Marc Solé and Enric Pastor. Traversal techniques for concurrent systems. Lecture Notes in Computer Science, 2517:220–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2517/25170220.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2517/25170220.pdf.

Frisch:2002:FBE

[2642] Alan Frisch, Daniel Sheridan, and Toby Walsh. A fixpoint based encoding for bounded model checking. Lecture Notes in Computer Science, 2517:238–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2517/25170238.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2517/25170238.pdf.

Ciardo:2002:UEV

[2643] Gianfranco Ciardo and Radu Siminiceanu. Using edge-valued decision diagrams for symbolic generation of shortest paths. Lecture Notes in Computer Science, 2517: 256–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2517/25170256.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2517/25170256.
pdf.

Sawada:2002:MVS

[2644] Jun Sawada and Ruben Gamboa. Mechanical verification of a square root algorithm using Taylor’s theorem. Lecture Notes in Computer Science, 2517:274–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2517/25170274.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2517/25170274.pdf.

Chatterjee:2002:SVF

[2645] Prosenjit Chatterjee and Ganesh Gopalakrishnan. A specification and verification framework for developing weak shared memory consistency protocols. Lecture Notes in Computer Science, 2517:292–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2517/25170292.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2517/25170292.pdf.

vanderMeulen:2002:MCD

[2646] Meine van der Meulen. Model checking the design of an unrestricted, stuck-at fault tolerant, asynchronous sequential circuit using SMV. Lecture Notes in Computer Science, 2517:310–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2517/25170310.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2517/25170310.pdf.

Sharp:2002:FDU

[2647] Richard Sharp. Functional design using behavioural and structural components. Lecture Notes in Computer Science, 2517:324–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2517/25170324.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2517/25170324.pdf.

McKeever:2002:CHD

[2648] Steve McKeever, Wayne Luk, and Arran Derbyshire. Compiling hardware descriptions with relative placement information for parametrised libraries. Lecture Notes in Computer Science, 2517:342–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2517/25170342.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2517/25170342.pdf.

Carmona:2002:IOC

[2649] Josep Carmona and Jordi Cortadella. Input/output compatibility of reactive systems. Lecture Notes in Computer Science, 2517:360–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2517/25170360.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2517/25170360.pdf.

Harel:2002:SPB

[2650] David Harel, Hillel Kugler, Rami Marelly, and Amir Pnueli. Smart play-out of behavioral requirements. Lecture Notes in Computer Science, 2517:378–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2517/25170378.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2517/25170378.pdf.

Anonymous:2002:AIck

[2651] Anonymous. Author index. Lecture Notes in Computer Science, 2517: 399–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2517/2517auth.pdf.

Bagchi:2002:BSL

[2652] Amitabha Bagchi, Adam L. Buchsbaum, and Michael T. Goodrich. Biased skip lists. Lecture Notes in Computer Science, 2518:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2518/25180001.htm;   http://link.
springer.de/link/service/series/
0558/papers/2518/25180001.pdf.

Sadakane:2002:SED

[2653] Kunihiko Sadakane. Space-efficient data structures for flexible text retrieval systems. Lecture Notes in Computer Science, 2518:14–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2518/25180014.htm;   http://link.
springer.de/link/service/series/
0558/papers/2518/25180014.pdf.

Iacono:2002:KIO

[2654] John Iacono. Key independent optimality. Lecture Notes in Computer Science, 2518:25–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2518/25180025.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2518/25180025.pdf.

Pettie:2002:CAC

[2655] Seth Pettie. On the comparisonaddition complexity of all-pairs shortest paths. Lecture Notes in Computer Science, 2518:32–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2518/25180032.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2518/25180032.pdf.

Boliac:2002:CWG

[2656] Rodica Boliac and Vadim Lozin. On the clique-width of graphs in hereditary classes. Lecture Notes in Computer Science, 2518:44–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2518/25180044.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2518/25180044.pdf.

Dietzfelbinger:2002:PRM

[2657] Martin Dietzfelbinger. The probability of a rendezvous is minimal in complete graphs. Lecture Notes in Computer Science, 2518:55–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2518/25180055.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2518/25180055.pdf.

Cai:2002:MVP

[2658] Jin-Yi Cai. On the minimum volume of a perturbed unit cube. Lecture Notes in Computer Science, 2518:67–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2518/25180067.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2518/25180067.pdf.

Guha:2002:NDB

[2659] Sumanta Guha, Paula Josiah, Anoop Mittal, and Son Dinh Tran. Nondelaunay-based curve reconstruction. Lecture Notes in Computer Science, 2518:79–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2518/25180079.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2518/25180079.pdf.

vanKreveld:2002:CCS

[2660] Marc van Kreveld and BettinaSpeckmann. Cutting a country for smallest square fit. Lecture Notes in Computer Science, 2518:91–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2518/25180091.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2518/25180091.pdf.

Dang:2002:EPT

[2661] Zhe Dang, Oscar H. Ibarra, and Zhi-Wei Sun. On the emptiness problem for two-way NFA with one reversal-bounded counter. Lecture Notes in Computer Science, 2518:103–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2518/25180103.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2518/25180103.pdf.

Kobayashi:2002:QMP

[2662] Hirotada Kobayashi and Keiji Matsumoto. Quantum multi-prover interactive proof systems with limited prior entanglement. Lecture Notes in Computer Science, 2518: 115–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2518/25180115.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2518/25180115.
                                        pdf.

Cheng:2002:SRC

[2663] Qi Cheng. Some remarks on the Lconjecture. Lecture Notes in Computer Science, 2518:128–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2518/25180128.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2518/25180128.pdf.

Wolle:2002:FNR

[2664] Thomas Wolle. A framework for network reliability problems on graphs of bounded treewidth. Lecture Notes in Computer Science, 2518: 137–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2518/25180137.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2518/25180137.
                                        pdf.

Galluccio:2002:FAA

[2665] Anna Galluccio and Guido Proietti. A faster approximation algorithm for 2-edge-connectivity augmentation. Lecture Notes in Computer Science, 2518:150–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2518/25180150.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2518/25180150.pdf.

Brandstadt:2002:TSC

[2666] A. Brandstädt, F. F. Dragan, H.-O. Le, and V. B. Le. Tree spanners on chordal graphs: Complexity, algorithms, open problems. Lecture Notes in Computer Science, 2518:163–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2518/25180163.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2518/25180163.pdf.

Jansen:2002:AFP

[2667] Klaus Jansen and Roberto Solis-Oba.

An asymptotic fully polynomial time approximation scheme for bin covering. Lecture Notes in Computer Science, 2518:175–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2518/25180175.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2518/25180175.pdf.

Blaser:2002:IAA

[2668] Markus Bläser and Bodo Manthey. Improved approximation algorithms for max-2SAT with cardinality constraint. Lecture Notes in Computer Science, 2518:187–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2518/25180187.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2518/25180187.pdf.

Karuno:2002:BAT

[2669] Yoshiyuki Karuno and Hiroshi Nagamochi. A better approximation for the two-stage assembly scheduling problem with two machines at the first stage. Lecture Notes in Computer Science, 2518:199–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2518/25180199.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2518/25180199.pdf.

Iacono:2002:Q

[2670] John Iacono and Stefan Langerman. Queaps. Lecture Notes in Computer Science, 2518:211–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2518/25180211.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2518/25180211.pdf.

Brodal:2002:FHC

[2671] Gerth Stølting Brodal and Rolf Fagerberg. Funnel heap — A cache oblivious priority queue. Lecture Notes in Computer Science, 2518: 219–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2518/25180219.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2518/25180219.
pdf.

Hartline:2002:CHI

[2672] Jason D. Hartline, Edwin S. Hong, Alexander E. Mohr, William R. Pentney, and Emily C. Rocke. Characterizing history independent data structures. Lecture Notes in Computer Science, 2518:229–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2518/25180229.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2518/25180229.pdf.

Raman:2002:FFP

[2673] Venkatesh Raman, Saket Saurabh, and C. R. Subramanian. Faster fixed parameter tractable algorithms for undirected feedback vertex set. Lecture Notes in Computer Science, 2518:241–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2518/25180241.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2518/25180241.pdf.

Stege:2002:APC

[2674] Ulrike Stege, Iris van Rooij, Alex Hertel, and Philipp Hertel. An O(pn + 1.151^{p})-algorithm for p-profit cover and its practical implications for vertex cover. Lecture Notes in Computer Science, 2518:249–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2518/25180249.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2518/25180249.pdf.

Demaine:2002:ESF

[2675] Erik D. Demaine, Mohammad Taghi Hajiaghayi, and Dimitrios M. Thilikos. Exponential speedup of fixed-parameter algorithms on K_{3,3}- minor-free or K_{5}-minor-free graphs. Lecture Notes in Computer Science, 2518:262–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2518/25180262.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2518/25180262.pdf.

Ahn:2002:CPD

[2676] Hee-Kap Ahn, Otfried Cheong, and René van Oostrum. Casting a polyhedron with directional uncertainty. Lecture Notes in Computer Science, 2518:274–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2518/25180274.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2518/25180274.pdf.

Cheng:2002:HSM

[2677] Siu-Wing Cheng, Tamal K. Dey, and Sheung-Hung Poon. Hierarchy of surface models and irreducible triangulation. Lecture Notes in Computer Science, 2518:286–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2518/25180286.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2518/25180286.pdf.

Yang:2002:ACT

[2678] Boting Yang, Cao An Wang, and Francis Chin. Algorithms and complexity for tetrahedralization detections. Lecture Notes in Computer Science, 2518:296–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2518/25180296.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2518/25180296.pdf.

Huang:2002:ACC

[2679] Chun-Hsi Huang and Xin He. Averagecase communication-optimal parallel parenthesis matching. Lecture Notes in Computer Science, 2518: 308–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2518/25180308.htm;   http:
//link.springer.de/link/service/

series/0558/papers/2518/25180308.
pdf.

Clementi:2002:ORP

[2680] Andrea E. F. Clementi, Angelo Monti, and Riccardo Silvestri. Optimal F -reliable protocols for the do-all problem on single-hop wireless networks (extended abstract). Lecture Notes in Computer Science, 2518:320–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2518/25180320.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2518/25180320.pdf.

Caragiannis:2002:NRE

[2681] Ioannis Caragiannis, Christos Kaklamanis, and Panagiotis Kanellopoulos. New results for energy-efficient broadcasting in wireless networks. Lecture Notes in Computer Science, 2518:332–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2518/25180332.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2518/25180332.pdf.

Kato:2002:IAM

[2682] Ryo Kato, Keiko Imai, and Takao Asano. An improved algorithm for the minimum Manhattan network problem. Lecture Notes in Computer Science, 2518:344–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

series/0558/bibs/2518/25180344.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2518/25180344.pdf.

Gudmundsson:2002:ADO

[2683] Joachim Gudmundsson, Christos Levcopoulos, Giri Narasimhan, and Michiel Smid. Approximate distance oracles revisited. Lecture Notes in Computer Science, 2518: 357–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2518/25180357.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2518/25180357.
pdf.

Aloupis:2002:FSC

[2684] Greg Aloupis, Erik D. Demaine, Vida Dujmović, Jeff Erickson, Stefan Langerman, Henk Meijer, Joseph O’Rourke, Mark Overmars, Michael Soss, Ileana Streinu, and Godfried T. Toussaint. Flat-state connectivity of linkages under dihedral motions. Lecture Notes in Computer Science, 2518:369–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2518/25180369.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2518/25180369.pdf.

Grigoriev:2002:PSI

[2685] Alexander Grigoriev and Gerhard J.

Woeginger. Project scheduling with irregular costs: Complexity, approximability, and algorithms.

Lecture Notes in Computer Science, 2518:381–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2518/25180381.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2518/25180381.pdf.

Bampis:2002:SID

[2686] Evripidis Bampis, Massimiliano Caramia, Jiří Fiala, Aleksei V. Fishkin, and Antonio Iovanella. Scheduling of independent dedicated multiprocessor tasks. Lecture Notes in Computer Science, 2518:391–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2518/25180391.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2518/25180391.pdf.

Miranda:2002:AMT

[2687] Antonio Miranda, Luz Torres, and Jianer Chen. On the approximability of multiprocessor task scheduling problems. Lecture Notes in Computer Science, 2518:403–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2518/25180403.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2518/25180403.pdf.

Biedl:2002:BDI

[2688] Therese Biedl and Dana F. Wilkinson. Bounded-degree independent sets in planar graphs (extended abstract). Lecture Notes in Computer Science, 2518:416–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2518/25180416.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2518/25180416.pdf.

Makino:2002:MER

[2689] Kazuhisa Makino, Yushi Uno, and Toshihide Ibaraki. Minimum edge ranking spanning trees of threshold graphs. Lecture Notes in Computer Science, 2518:428–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2518/25180428.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2518/25180428.pdf.

Ito:2002:FTT

[2690] Hiro Ito, Hiroshi Nagamochi, Yosuke Sugiyama, and Masato Fujita. File transfer tree problems. Lecture Notes in Computer Science, 2518: 441–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2518/25180441.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2518/25180441.
pdf.

Arvind:2002:AAS

[2691] V. Arvind and Venkatesh Raman.

Approximation algorithms for some parameterized counting problems. Lecture Notes in Computer Science, 2518:453–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2518/25180453.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2518/25180453.pdf.

Avidor:2002:AMS

[2692] Adi Avidor and Uri Zwick. Approximating MINk-SAT. Lecture Notes in Computer Science, 2518: 465–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2518/25180465.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2518/25180465.
                                        pdf.

Fujiwara:2002:ACC

[2693] Hiroshi Fujiwara and Kazuo Iwama. Average-case competitive analyses for ski-rental problems. Lecture Notes in Computer Science, 2518: 476–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2518/25180476.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2518/25180476.
                                        pdf.

Ambuhl:2002:CPI

[2694] Christoph Ambühl and Uli Wagner. On the clique problem in intersection graphs of ellipses. Lecture Notes in Computer Science, 2518:489–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2518/25180489.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2518/25180489.pdf.

Lingas:2002:GAB

[2695] Andrzej Lingas. A geometric approach to Boolean matrix multiplication. Lecture Notes in Computer Science, 2518:501–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2518/25180501.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2518/25180501.pdf.

Papadopoulou:2002:MMV

[2696] Evanthia Papadopoulou and D. T. Lee.

The min-max Voronoi diagram of polygons and applications in VLSI manufacturing. Lecture Notes in Computer Science, 2518:511–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2518/25180511.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2518/25180511.pdf.

Chowdhury:2002:IDO

[2697] Rezaul Alam Chowdhury and Vijaya Ramachandran. Improved distance oracles for avoiding link-failure. Lecture Notes in Computer Science, 2518:523–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2518/25180523.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2518/25180523.pdf.

Jurdzinski:2002:PAW

[2698] Tomasz Jurdziński and Grzegorz Stachowiak. Probabilistic algorithms for the wakeup problem in single-hop radio networks. Lecture Notes in Computer Science, 2518:535–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2518/25180535.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2518/25180535.pdf.

Shikaripura:2002:SME

[2699] Vivek Shikaripura and Ajay D. Kshemkalyani. A simple, memoryefficient bounded concurrent timestamping algorithm. Lecture Notes in Computer Science, 2518:550– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2518/25180550.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2518/25180550.
                                        pdf.

Buchheim:2002:CMS

[2700] Christoph Buchheim and Seok-Hee Hong. Crossing minimization for symmetries. Lecture Notes in Computer Science, 2518:563–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2518/25180563.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2518/25180563.pdf.

Erten:2002:SEP

[2701] Cesim Erten and Stephen G. Kobourov.

Simultaneous embedding of a planar graph and its dual on the Grid. Lecture Notes in Computer Science, 2518:575–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2518/25180575.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2518/25180575.pdf.

Vitanyi:2002:MIE

[2702] Paul Vitányi. Meaningful information (extended abstract). Lecture Notes in Computer Science, 2518: 588–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2518/25180588.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2518/25180588.
pdf.

Sandholm:2002:MCS

[2703] Tuomas Sandholm and Subhash Suri. Market clearing with supply and demand curves. Lecture Notes in Computer Science, 2518: 600–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2518/25180600.htm;   http:
//link.springer.de/link/service/

                                        series/0558/papers/2518/25180600.
                                        pdf.

Ito:2002:PTS

[2704] Takehiro Ito, Xiao Zhou, and Takao Nishizeki. Partitioning trees of supply and demand. Lecture Notes in Computer Science, 2518:612– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2518/25180612.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2518/25180612.
                                        pdf.

Dehne:2002:MVR

[2705] Frank Dehne, Rolf Klein, and Raimund Seidel. Maximizing a Voronoi region: The convex case. Lecture Notes in Computer Science, 2518:624–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2518/25180624.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2518/25180624.pdf.

Devroye:2002:RT

[2706] Luc Devroye. Random tries. Lecture Notes in Computer Science, 2518: 635–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2518/25180635.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2518/25180635.
                                        pdf.

Pippenger:2002:EAC

[2707] Nicholas Pippenger. Expected acceptance counts for finite automata with almost uniform input. Lecture Notes in Computer Science, 2518:636–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2518/25180636.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2518/25180636.pdf.

Pach:2002:MDP

[2708] János Pach and Géza Tóth. Monotone drawings of planar graphs. Lecture Notes in Computer Science, 2518:647–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2518/25180647.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2518/25180647.pdf.

Anonymous:2002:AIcl

[2709] Anonymous. Author index. Lecture Notes in Computer Science, 2518: 655–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2518/2518auth.pdf.

Rahal:2002:DEC

[2710] Amira Rahal, Qiang Zhu, and Per-Åke Larson. Developing evolutionary cost models for query optimization in a dynamic multidatabase environment. Lecture Notes in Computer Science, 2519:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190001.pdf.

Demey:2002:CML

[2711] Jan Demey, Mustafa Jarrar, and Robert Meersman. A conceptual markup language that supports interoperability between business rule modeling systems. Lecture Notes in Computer Science, 2519: 19–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2519/25190019.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2519/25190019.
pdf.

Borgida:2002:DDL

[2712] Alex Borgida and Luciano Serafini. Distributed description logics: Directed domain correspondences in federated information sources. Lecture Notes in Computer Science, 2519:36–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190036.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190036.pdf.

Heinlein:2002:SCW

[2713] Christian Heinlein. Synchronization of concurrent workflows using interaction expressions and coordination protocols. Lecture Notes in Computer Science, 2519:54–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190054.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190054.pdf.

Chen:2002:CRA

[2714] Qiming Chen and Meichun Hsu. CPM revisited — an architecture comparison. Lecture Notes in Computer Science, 2519:72–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190072.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190072.pdf.

Colombo:2002:MCC

[2715] Enzo Colombo, Chiara Francalanci, and Barbara Pernici. Modeling coordination and control in crossorganizational workflows. Lecture Notes in Computer Science, 2519: 91–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2519/25190091.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2519/25190091.
pdf.

Lauzac:2002:VPI

[2716] Susan Weissman Lauzac and Panos K.

Chrysanthis. View propagation and inconsistency detection for cooperative mobile agents. Lecture Notes in Computer Science, 2519: 107–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2519/25190107.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2519/25190107.
                                        pdf.

Balke:2002:RTT

[2717] Wolf-Tilo Balke, Ulrich Güntzer, and Werner Kießling. On realtime top k querying for mobile services. Lecture Notes in Computer Science, 2519:125–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190125.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190125.pdf.

Jiao:2002:MAM

[2718] Yu Jiao and Ali R. Hurson. Mobile agents in mobile data access systems. Lecture Notes in Computer Science, 2519:144–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190144.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190144.pdf.

Blake:2002:UAC

[2719] M. Brian Blake. Using agent control and communication in a distributed workflow information system. Lecture Notes in Computer Science, 2519:163–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190163.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190163.pdf.

Winoto:2002:EAO

[2720] Pinata Winoto, Gordon McCalla, and Julita Vassileva. An extended alternating-offers bargaining protocol for automated negotiation in multi-agent systems. Lecture Notes in Computer Science, 2519:179–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190179.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190179.pdf.

Herrero:2002:HBP

[2721] Pilar Herrero and Angélica de Antonio. A human based perception model for cooperative intelligent virtual agents. Lecture Notes in Computer Science, 2519:195–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190195.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190195.pdf.

Suryanarayana:2002:DAC

[2722] Girish Suryanarayana and Richard Taylor. A decentralized algorithm for coordinating independent peers: An initial examination. Lecture Notes in Computer Science, 2519:213–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190213.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190213.pdf.

Vassileva:2002:SPP

[2723] Julita Vassileva. Supporting peerto-peer user communities. Lecture Notes in Computer Science, 2519: 230–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2519/25190230.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2519/25190230.
                                        pdf.

Heutelbeck:2002:CSS

[2724] Dominic Heutelbeck. Context spaces — self-structuring distributed networks for contextual messaging and resource discovery. Lecture Notes in Computer Science, 2519:248–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190248.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190248.pdf.

Wang:2002:SCL

[2725] Weigang Wang, Jörg M. Haake, and Martin Wessner. Supporting cooperative learning in distributed project teams. Lecture Notes in Computer Science, 2519:266–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190266.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190266.pdf.

Simoff:2002:DEV

[2726] Simeon J. Simoff and Robert P. Biuk-Aghai. Discovering emergent virtual work processes in collaborative systems. Lecture Notes in Computer Science, 2519:286–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190286.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190286.pdf.

Shen:2002:FMA

[2727] Haifeng Shen and Chengzheng Sun.

Flexible merging for asynchronous collaborative systems. Lecture Notes in Computer Science, 2519:304– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2519/25190304.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2519/25190304.
pdf.

Madria:2002:MVT

[2728] Sanjay Kumar Madria, Mohammed Baseer, and Sourav S. Bhowmick. A multi-version transaction model to improve data availability in mobile computing. Lecture Notes in Computer Science, 2519:322–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190322.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190322.pdf.

Wombacher:2002:FTP

[2729] Andreas Wombacher and Bendick Mahleko. Finding trading partners to establish ad-hoc business processes. Lecture Notes in Computer Science, 2519:339–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190339.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190339.pdf.

Murata:2002:RWD

[2730] Takahiro Murata and Naftaly H. Minsky. Regulating work in digital enterprises: a flexible managerial framework. Lecture Notes in Computer Science, 2519:356–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190356.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190356.pdf.

Fauvet:2002:CQD

[2731] Marie-Christine Fauvet, Marlon Dumas, and Boualem Benatallah. Collecting and querying distributed traces of composite service executions. Lecture Notes in Computer Science, 2519:373–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190373.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190373.pdf.

Lai:2002:AWB

[2732] Wei-Cheng Lai, Gerard Sychay, and Edward Chang. An architecture of a Web-based collaborative image search engine. Lecture Notes in Computer Science, 2519:391–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190391.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190391.pdf.

Gancarski:2002:PPA

[2733] Stéphane Gançarski, Hubert Naacke, Esther Pacitti, and Patrick Valduriez. Parallel processing with autonomous databases in a cluster system. Lecture Notes in Computer Science, 2519:410–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190410.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190410.pdf.

Amann:2002:QXS

[2734] Bernd Amann, Catriel Beeri, Irini Fundulaki, and Michel Scholl. Querying XML sources using an ontology-based mediator. Lecture Notes in Computer Science, 2519:429–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190429.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190429.pdf.

Badia:2002:HPM

[2735] Antonio Badia and Sanjay Kumar Madria. Handling partial matches in semistructured data with cooperative query answering techniques. Lecture Notes in Computer Science, 2519:449–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190449.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190449.pdf.

Manolescu:2002:EQD

[2736] Ioana Manolescu, Luc Bouganim, Françoise Fabret, and Eric Simon. Efficient querying of distributed resources in mediator systems. Lecture Notes in Computer Science, 2519:468–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190468.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190468.pdf.

Mecella:2002:MDQ

[2737] Massimo Mecella, Monica Scannapieco, Antonino Virgillito, Roberto Baldoni, Tiziana Catarci, and Carlo Batini. Managing data quality in cooperative information systems. Lecture Notes in Computer Science, 2519:486–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190486.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190486.pdf.

Deshpande:2002:AFT

[2738] Umesh Deshpande, Arobinda Gupta, and Anupam Basu. Adaptive fault tolerant hospital resource scheduling. Lecture Notes in Computer Science, 2519:503–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190503.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190503.pdf.

Simoff:2002:MLD

[2739] Simeon J. Simoff. Mining ’living’ data — providing context information to a negotiation process. Lecture Notes in Computer Science, 2519:521–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190521.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190521.pdf.

Barthelmess:2002:NPE

[2740] P. Barthelmess and C. A. Ellis. The neem platform: An evolvable framework for perceptual collaborative applications. Lecture Notes in Computer Science, 2519:522–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190522.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190522.pdf.

Baina:2002:MPS

[2741] Karim Baïna, Samir Tata, and Khalid Benali. A model for process service interaction. Lecture Notes in Computer Science, 2519:524–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190524.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190524.pdf.

Jamil:2002:LDI

[2742] Hasan M. Jamil and Giovanni A. Modica. Leveraging dynamic inheritance in complex ontology representation. Lecture Notes in Computer Science, 2519:525–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190525.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190525.pdf.

Rivera:2002:DAG

[2743] G. Rivera and M. C. Norrie. A database approach to global document spaces: Replacing files with shared, connected objects. Lecture Notes in Computer Science, 2519:526–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190526.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190526.pdf.

Maier:2002:PRI

[2744] F. Maier, D. Nute, W. D. Potter, J. Wang, M. Twery, and H. M. Rauscher. PROLOG/RDBMS integration in the NED intelligent information system. Lecture Notes in Computer Science, 2519:528–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190528.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190528.pdf.

Lee:2002:NLM

[2745] Dong Chun Lee, Hyun Cheul Shin, and Jeom Goo Kim. New location management for reducing HLR overhead traffic in mobile networks. Lecture Notes in Computer Science, 2519:529–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190529.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190529.pdf.

Nabar:2002:FOF

[2746] Pranav S. Nabar, Amit L. Padalkar, and Rushikesh K. Joshi. A filter object framework for MICO.

Lecture Notes in Computer Science, 2519:531–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190531.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190531.pdf.

Klefstad:2002:DPM

[2747] Raymond Klefstad, Arvind S. Krishna, and Douglas C. Schmidt. Design and performance of a modular portable object adapter for distributed, realtime, and embedded CORBA applications. Lecture Notes in Computer Science, 2519:549–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190549.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190549.pdf.

Deshpande:2002:DPA

[2748] Mayur Deshpande, Douglas C. Schmidt, Carlos O’Ryan, and Darrell Brunsch. Design and performance of asynchronous method handling for CORBA. Lecture Notes in Computer Science, 2519:568–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190568.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190568.pdf.

Siddhartha:2002:WSI

[2749] Pradyumna Siddhartha and Shubhashis Sengupta. Web services interoperability: a practitioner’s experience. Lecture Notes in Computer Science, 2519:587–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190587.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190587.pdf.

Gillmann:2002:CWE

[2750] Michael Gillmann, Joachim Hertel, Christoph G. Jung, Günther Kaufmann, and Michael Wolber. Cooking the Web-ERP. Lecture Notes in Computer Science, 2519: 602–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2519/25190602.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2519/25190602.
pdf.

Baker:2002:WSC

[2751] Seán Baker. Web services and CORBA. Lecture Notes in Computer Science, 2519:618–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190618.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190618.pdf.

Gokhale:2002:CDG

[2752] Aniruddha Gokhale and Balachandran Natarajan. Composing and deploying Grid middleware Web services using model driven architecture. Lecture Notes in Computer Science, 2519:633–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190633.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190633.pdf.

Long:2002:DPE

[2753] Brad Long. A design pattern for efficient retrieval of large data sets from remote data sources. Lecture Notes in Computer Science, 2519:650–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190650.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190650.pdf.

Atzmon:2002:RPD

[2754] Hilla Atzmon, Roy Friedman, and Roman Vitenberg. Replacement policies for a distributed object caching service. Lecture Notes in Computer Science, 2519:661–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190661.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190661.pdf.

Rafaelsen:2002:DPM

[2755] Hans Ole Rafaelsen and Frank Eliassen. Design and performance of a media gateway trader. Lecture Notes in Computer Science, 2519: 675–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2519/25190675.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2519/25190675.
pdf.

Beznosov:2002:OSA

[2756] Konstantin Beznosov. Object security attributes: Enabling applicationspecific access control in middleware. Lecture Notes in Computer Science, 2519:693–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190693.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190693.pdf.

deOliveiraDias:2002:IOV

[2757] Gláucia de Oliveira Dias and Alba Cristina Magalhaes Alves de Melo. Integrating optimistic virtual synchrony to a CORBA object group service. Lecture Notes in Computer Science, 2519:711–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190711.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190711.pdf.

Cotroneo:2002:ICB

[2758] D. Cotroneo, A. Mazzeo, L. Romano, and S. Russo. Implementing a CORBA-based architecture for leveraging the security level of existing applications. Lecture Notes in Computer Science, 2519:723–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190723.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190723.pdf.

Felber:2002:RRT

[2759] Pascal Felber and Priya Narasimhan. Reconciling replication and transactions for the end-to-end reliability of CORBA applications. Lecture Notes in Computer Science, 2519: 737–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2519/25190737.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2519/25190737.
                                        pdf.

Zinky:2002:RPM

[2760] John Zinky, Joseph Loyall, and Richard Shapiro. Runtime performance modeling and measurement of adaptive distributed object applications. Lecture Notes in Computer Science, 2519:755–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190755.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190755.pdf.

David:2002:IAM

[2761] Pierre-Charles David and Thomas Ledoux. An infrastructure for adaptable middleware. Lecture Notes in Computer Science, 2519: 773–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2519/25190773.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2519/25190773.
pdf.

Gutierrez-Nolasco:2002:RMF

[2762] Sebastian Gutierrez-Nolasco and Nalini Venkatasubramanian. A reflective middleware framework for communication in dynamic environments. Lecture Notes in Computer Science, 2519:791–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190791.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190791.pdf.

Biegel:2002:DPB

[2763] Gregory Biegel, Vinny Cahill, and Mads Haahr. A dynamic proxy based architecture to support distributed Java objects in a mobile environment. Lecture Notes in Computer Science, 2519:809–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190809.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190809.pdf.

Reinehr:2002:CBW

[2764] Leonardo Hartleben Reinehr and Maria Beatriz Felgar de Toledo. A CORBAbased workflow management system for wireless communication environments. Lecture Notes in Computer Science, 2519:827–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190827.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190827.pdf.

Paal:2002:CDC

[2765] Stefan Paal, Reiner Kammüller, and Bernd Freisleben. Customizable deployment, composition, and hosting of distributed Java applications. Lecture Notes in Computer Science, 2519:845–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190845.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190845.pdf.

Rakotonirainy:2002:RDP

[2766] Andry Rakotonirainy and Greg Groves. Resource discovery for pervasive environments. Lecture Notes in Computer Science, 2519:866–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190866.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190866.pdf.

Cervieri:2002:ASS

[2767] Alexandre Cervieri, Rômulo Silva de Oliveira, and Cláudio F. Resin Geyer. An adaptive scheduling service for real-time CORBA. Lecture Notes in Computer Science, 2519: 884–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2519/25190884.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2519/25190884.
pdf.

Corsaro:2002:DPJ

[2768] Angelo Corsaro and Douglas C.

Schmidt. The design and performance of the jRate real-time Java implementation. Lecture Notes in Computer Science, 2519:900–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190900.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190900.pdf.

Gill:2002:EDB

[2769] Christopher D. Gill, Fred Kuhns, Douglas C. Schmidt, and Ron K. Cytron. Empirical differences between COTS middleware scheduling strategies. Lecture Notes in Computer Science, 2519:922–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190922.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190922.pdf.

Coronato:2002:ABR

[2770] Antonio Coronato, Marco Cinquegrani, and Giuseppe De Pietro. Adding business rules and constraints in component based applications. Lecture Notes in Computer Science, 2519:948–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190948.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190948.pdf.

Teixeira:2002:CCM

[2771] Sandra Teixeira, Pedro Vicente, Alexandre Pinto, Hugo Miranda, Luis Rodrigues, Jorge Martins, and António Rito-Silva. Configuring the communication middleware to support multi-user object-oriented environments. Lecture Notes in Computer Science, 2519:965–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190965.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25190965.pdf.

Kalibera:2002:DCS

[2772] Tomás̆ Kalibera and Petr Tuma. Distributed component system based on architecture description: The SOFA experience. Lecture Notes in Computer Science, 2519:981–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25190981.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2519/25190981.pdf.

Farias:2002:CEP

[2773] Andrés Farías and Mario Südholt. On components with explicit protocols satisfying a notion of correctness by construction. Lecture Notes in Computer Science, 2519:995–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25190995.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25190995.pdf.

Yang:2002:WFO

[2774] Guizhen Yang and Michael Kifer.

Well-founded optimism: Inheritance in frame-based knowledge bases. Lecture Notes in Computer Science, 2519:1013–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25191013.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25191013.pdf.

Heymans:2002:DOL

[2775] S. Heymans and D. Vermeir. A defeasible ontology language. Lecture Notes in Computer Science, 2519: 1033–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2519/25191033.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2519/25191033.
pdf.

Yang:2002:SAI

[2776] Guizhen Yang and Michael Kifer. On the semantics of anonymous identity and reification. Lecture Notes in Computer Science, 2519: 1047–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2519/25191047.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2519/25191047.
                                        pdf.

Pan:2002:EDSa

[2777] Jeff Z. Pan and Ian Horrocks. Extending datatype support in Web ontology reasoning. Lecture Notes in Computer Science, 2519:1067–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25191067.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25191067.pdf.

Motik:2002:CMA

[2778] Boris Motik, Alexander Maedche, and Raphael Volz. A conceptual modeling approach for semanticsdriven enterprise applications. Lecture Notes in Computer Science, 2519: 1082–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2519/25191082.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2519/25191082.
                                        pdf.

Guizzardi:2002:TOF

[2779] Giancarlo Guizzardi, Heinrich Herre, and Gerd Wagner. Towards ontological foundations for UML conceptual models. Lecture Notes in Computer Science, 2519:1100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25191100.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25191100.pdf.

Bai:2002:TSO

[2780] Yun Bai. Towards secure object oriented database systems. Lecture Notes in Computer Science, 2519: 1118–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2519/25191118.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2519/25191118.
pdf.

Kalfoglou:2002:IFB

[2781] Yannis Kalfoglou and Marco Schorlemmer. Information-flow-based ontology mapping. Lecture Notes in Computer Science, 2519:1132–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25191132.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25191132.pdf.

Bechhofer:2002:SSA

[2782] Sean Bechhofer, Leslie Carr, Carole Goble, Simon Kampa, and Timothy Miles-Board. The semantics of semantic annotation. Lecture Notes in Computer Science, 2519:1152–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25191152.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25191152.pdf.

Buccafurri:2002:UBB

[2783] Francesco Buccafurri, Gianluca Lax, Domenico Rosaci, and Domenico Ursino. A user behavior-based agent for improving Web usage. Lecture Notes in Computer Science, 2519:1168–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25191168.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25191168.pdf.

Stojanovic:2002:UOE

[2784] Nenad Stojanovic and Ljiljana Stojanovic. Usage-oriented evolution of ontology-based knowledge management systems. Lecture Notes in Computer Science, 2519:1186– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2519/25191186.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2519/25191186.
pdf.

Sure:2002:OGO

[2785] York Sure, Juergen Angele, and Steffen Staab. OntoEdit: Guiding ontology development by methodology and inferencing. Lecture Notes in Computer Science, 2519:1205–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25191205.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25191205.pdf.

Singh:2002:OMC

[2786] Push Singh, Thomas Lin, Erik T.

Mueller, Grace Lim, Travell Perkins, and Wan Li Zhu. Open mind common sense: Knowledge acquisition from the general public. Lecture Notes in Computer Science, 2519:1223–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25191223.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25191223.pdf.

Jarrar:2002:FOE

[2787] Mustafa Jarrar and Robert Meersman. Formal ontology engineering in the DOGMA approach. Lecture Notes in Computer Science, 2519: 1238–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2519/25191238.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2519/25191238.
pdf.

Garelli:2002:DMM

[2788] Francesco Garelli and Carlo Ferrari. A dynamic model for mapping XML elements in a object-oriented fashion. Lecture Notes in Computer Science, 2519:1255–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25191255.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25191255.pdf.

Xu:2002:KDH

[2789] Haiyuan Xu, Quanyuan Wu, Huaimin Wang, Guogui Yang, and Yan Jia. KF-Diff+: Highly efficient change detection algorithm for XML documents. Lecture Notes in Computer Science, 2519:1273–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25191273.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25191273.pdf.

Lawrence:2002:NXD

[2790] Ramon Lawrence. Naming in XML documents. Lecture Notes in Computer Science, 2519:1287–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25191287.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25191287.pdf.

Corbett:2002:ROU

[2791] Dan Corbett. Reasoning with ontologies by using knowledge conjunction in conceptual graphs. Lecture Notes in Computer Science, 2519: 1304–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2519/25191304.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2519/25191304.
pdf.

Katz:2002:NLA

[2792] Boris Katz, Jimmy Lin, and Dennis Quan. Natural language annotations for the Semantic Web. Lecture Notes in Computer Science, 2519:1317–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25191317.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25191317.pdf.

Nam:2002:MIA

[2793] Young-Kwang Nam, Joseph Goguen, and Guilian Wang. A metadata integration assistant generator for heterogeneous distributed databases. Lecture Notes in Computer Science, 2519:1332–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25191332.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25191332.pdf.

Kerschberg:2002:IWS

[2794] Larry Kerschberg, Wooju Kim, and Anthony Scime. Intelligent Web search via personalizable meta-search agents. Lecture Notes in Computer Science, 2519:1345–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25191345.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25191345.pdf.

Dinh:2002:DIS

[2795] Thang Le Dinh and M. Leonard. Defining information system components. Lecture Notes in Computer Science, 2519:1359–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25191359.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25191359.pdf.

Anido:2002:DRS

[2796] Luis Anido, Judith Rodríguez, Manuel Caeiro, and Juan Santos. Discovering resources in the Semantic Web. Lecture Notes in Computer Science, 2519:1360–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2519/25191360.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2519/25191360.pdf.

Sotnykova:2002:GIS

[2797] Anastasiya Sotnykova. Geodata interoperation via semantic correspondences. Lecture Notes in Computer Science, 2519:1361–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25191361.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25191361.pdf.

Pan:2002:EDSb

[2798] Jeff Z. Pan and Ian Horrocks. Extending datatype support in Web ontology reasoning. Lecture Notes in Computer Science, 2519:1362–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2519/25191362.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2519/25191362.pdf.

Anonymous:2002:AIcm

[2799] Anonymous. Author index. Lecture Notes in Computer Science, 2519: 1365–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2519/2519auth.pdf.

Toivonen:2002:ORA

[2800] Santtu Toivonen and Heikki Helin.

Options for reusing agent conversations. Lecture Notes in Computer Science, 2521:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2521/25210001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2521/25210001.pdf.

Lopes:2002:MMS

[2801] Rui Pedro Lopes and José Luis Oliveira. Multi-management schemes for MAF platforms. Lecture Notes in Computer Science, 2521:11–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2521/25210011.htm;   http://link.
springer.de/link/service/series/
0558/papers/2521/25210011.pdf.

Delgado:2002:ANM

[2802] Jaime Delgado, Isabel Gallego, Roberto García, and Rosa Gil. An architecture for negotiation with mobile agents. Lecture Notes in Computer Science, 2521:21–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2521/25210021.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2521/25210021.pdf.

Renganarayanan:2002:IAE

[2803] Vidya Renganarayanan, Abdelsalam (Sumi) Helal, and Amar Nalla. Internet agents for effective collaboration. Lecture Notes in Computer Science, 2521:33–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2521/25210033.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2521/25210033.pdf.

Dillenseger:2002:PET

[2804] Bruno Dillenseger, Anne-Marie Tagant, and Laurent Hazard. Programming and executing telecommunication service logic with moorea reactive mobile agents. Lecture Notes in Computer Science, 2521:48–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2521/25210048.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2521/25210048.pdf.

Lee:2002:GEA

[2805] Seongkee Lee and Taiyun Kim.

Generic engineering approach for agent-based system development. Lecture Notes in Computer Science, 2521:58–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2521/25210058.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2521/25210058.pdf.

Rossier:2002:EIM

[2806] Daniel Rossier and Rudolf Scheurer.

An ecosystem-inspired mobile agent middleware for active network management. Lecture Notes in Computer Science, 2521:73–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2521/25210073.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2521/25210073.pdf.

Kim:2002:EMM

[2807] Insu Kim, Keecheon Kim, Sunyoung Han, and Hyukjoon Lee. An enhanced mobility management mechanism for active access networks. Lecture Notes in Computer Science, 2521:83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2521/25210083.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2521/25210083.pdf.

Barba:2002:AAN

[2808] Antonio Barba and Ernesto Sánchez. An architecture for active network performance management based on intelligent agents. Lecture Notes in Computer Science, 2521:94–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2521/25210094.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2521/25210094.pdf.

Khedr:2002:ABC

[2809] Mohamed Khedr, Ahmed Karmouch, Ramiro Liscano, and Tom Gray. Agentbased context-aware ad hoc communication. Lecture Notes in Computer Science, 2521:105–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2521/25210105.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2521/25210105.pdf.

Ahmed:2002:DMS

[2810] Mohamed Ahmed, Hamid Harroud, Roger Impey, and Ahmed Karmouch. Designing multimedia service agents for mobile users. Lecture Notes in Computer Science, 2521: 119–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2521/25210119.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2521/25210119.
pdf.

Umezawa:2002:MAB

[2811] Takeshi Umezawa, Ichiro Satoh, and Yuichiro Anzai. A mobile agent-based framework for configurable sensor networks. Lecture Notes in Computer Science, 2521:128–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2521/25210128.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2521/25210128.pdf.

Gonzalez-Valenzuela:2002:EMS

[2812] Sergio Gonzalez-Valenzuela and Son T.

Vuong. Evaluation of migration strategies for mobile agents in network routing. Lecture Notes in Computer Science, 2521:141–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2521/25210141.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2521/25210141.pdf.

Bohoris:2002:HAN

[2813] Christos Bohoris, George Pavlou, and Antonio Liotta. A hybrid approach to network performance monitoring based on mobile agents and CORBA. Lecture Notes in Computer Science, 2521:151–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2521/25210151.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2521/25210151.pdf.

Sugar:2002:MAD

[2814] Robert Sugar and Sandor Imre. Mobile agent distribution in a game-theoretic approach. Lecture Notes in Computer Science, 2521:163–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2521/25210163.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2521/25210163.pdf.

Korba:2002:ACM

[2815] Larry Korba, Ronggong Song, and George Yee. Anonymous communications for mobile agents. Lecture Notes in Computer Science, 2521: 171–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2521/25210171.htm;   http:
                                        //link.springer.de/link/service/

series/0558/papers/2521/25210171.
pdf.

Robles:2002:ISA

[2816] Sergi Robles, Joan Mir, Joan Ametller, and Joan Borrell. Implementation of secure architectures for mobile agents in MARISM-A. Lecture Notes in Computer Science, 2521: 182–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2521/25210182.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2521/25210182.
pdf.

Mota:2002:QSM

[2817] Telma Mota, Stylianos Gouveris, George Pavlou, Angelos Michalas, and John Psoroulas. Quality of service management in IP networks using mobile agent technology. Lecture Notes in Computer Science, 2521:193–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2521/25210193.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2521/25210193.pdf.

Dharmalingam:2002:TQS

[2818] Kalaiarul Dharmalingam and Martin Collier. Transparent QoS support of network applications using netlets. Lecture Notes in Computer Science, 2521:206–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2521/25210206.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2521/25210206.pdf.

Bravo:2002:FBQ

[2819] Alberto Montilla Bravo and Marisol García Valls. FIPA-based QoS negotiator for nomadic agents. Lecture Notes in Computer Science, 2521: 216–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2521/25210216.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2521/25210216.
                                        pdf.

Reuter:2002:SNM

[2820] Emmanuel Reuter and Françoise Baude. System and network management itineraries for mobile agents. Lecture Notes in Computer Science, 2521:227–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2521/25210227.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2521/25210227.pdf.

Johansson:2002:FMA

[2821] Stefan Johansson, Paul Davidsson, and Martin Kristell. Four multi-agent architectures for intelligent network load management. Lecture Notes in Computer Science, 2521:239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2521/25210239.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2521/25210239.pdf.

Yang:2002:AMI

[2822] Kun Yang, Alex Galis, Telma Mota, and Stylianos Gouveris. Automated management of IP networks through policy and mobile agents. Lecture Notes in Computer Science, 2521:249–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2521/25210249.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2521/25210249.pdf.

Tarkoma:2002:FAM

[2823] Sasu Tarkoma and Mikko Laukkanen. Facilitating agent messaging on PDAs. Lecture Notes in Computer Science, 2521:259–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2521/25210259.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2521/25210259.pdf.

Wang:2002:MAD

[2824] Zhou Wang and Jochen Seitz. Mobile agents for discovering and accessing services in nomadic environments. Lecture Notes in Computer Science, 2521:269–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2521/25210269.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2521/25210269.pdf.

Olougouna:2002:EMA

[2825] Edgar A. Olougouna, Samuel Pierre, and Roch H. Glitho. An extensible mobile-agent-based framework for coordinating distributed information retrieval applications. Lecture Notes in Computer Science, 2521:281–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2521/25210281.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2521/25210281.pdf.

Sygkouna:2002:SIA

[2826] Irene Sygkouna, Maria Strimpakou, Francisco Valera, Anastasia Kaltabani, Luis Bellido, Enrique Vazquez, and Miltiades Anagnostou. Seamless incorporation of agents in an E-commerce intermediation platform. Lecture Notes in Computer Science, 2521:292–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2521/25210292.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2521/25210292.pdf.

Zapf:2002:UAO

[2827] Michael Zapf, Rolf Reinema, Ruben Wolf, and Sven Türpe. UNITE — an agent-oriented teamwork environment. Lecture Notes in Computer Science, 2521:302–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2521/25210302.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2521/25210302.pdf.

Anonymous:2002:AIcn

[2828] Anonymous. Author index. Lecture Notes in Computer Science, 2521: 317–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2521/2521auth.pdf.

Akrivas:2002:CSQ

[2829] Giorgos Akrivas, Manolis Wallace, Giorgos Stamou, and Stefanos Kollias. Context — sensitive query expansion based on fuzzy clustering of index terms. Lecture Notes in Computer Science, 2522:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2522/25220001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2522/25220001.pdf.

Badea:2002:III

[2830] Liviu Badea and Doina Tilivea.

Intelligent information integration as a constraint handling problem. Lecture Notes in Computer Science, 2522:12–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2522/25220012.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2522/25220012.pdf.

Balaban:2002:EPM

[2831] Mira Balaban and Steffen Jurk. Effect preservation as a means for achieving update consistency. Lecture Notes in Computer Science, 2522:28–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2522/25220028.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2522/25220028.pdf.

terBekke:2002:RQP

[2832] J. H. ter Bekke and J. A. Bakker. Recursive queries in product databases. Lecture Notes in Computer Science, 2522:44–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2522/25220044.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2522/25220044.pdf.

Benslimane:2002:CBL

[2833] Djamal Benslimane, Mohand-Saïd Hacid, Evimaria Terzi, and Farouk Toumani. A class-based logic language for ontologies. Lecture Notes in Computer Science, 2522: 56–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2522/25220056.htm;   http:
                                        //link.springer.de/link/service/

series/0558/papers/2522/25220056.
pdf.

Bertossi:2002:CAI

[2834] Leopoldo Bertossi, Jan Chomicki, Alvaro Cortés, and Claudio Gutiérrez. Consistent answers from integrated data sources. Lecture Notes in Computer Science, 2522:71–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/link/service/
series/0558/bibs/2522/25220071.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2522/25220071.pdf.

Bresciani:2002:KBQ

[2835] Paolo Bresciani and Paolo Fontana.

A knowledge-based query system for biological databases. Lecture Notes in Computer Science, 2522: 86–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2522/25220086.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2522/25220086.
pdf.

Bulskov:2002:MSC

[2836] Henrik Bulskov, Rasmus Knappe, and Troels Andreasen. On measuring similarity for conceptual querying. Lecture Notes in Computer Science, 2522:100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2522/25220100.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2522/25220100.pdf.

Chountas:2002:RQT

[2837] Panagiotis Chountas, Ilias Petrounias, Krassimir Atanassov, Vassilis Kodogiannis, and Elia El-Darzi. Representation & querying of temporal conflict. Lecture Notes in Computer Science, 2522:112–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2522/25220112.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2522/25220112.pdf.

Ciaccia:2002:AFS

[2838] Paolo Ciaccia and Wilma Penzo. Adding flexibility to structure similarity queries on XML data. Lecture Notes in Computer Science, 2522: 124–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2522/25220124.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2522/25220124.
pdf.

Dang:2002:EIL

[2839] Tran Khanh Dang, Josef Küng, and Roland Wagner. An efficient incremental lower bound approach for solving approximate nearest-neighbor problem of complex vague queries. Lecture Notes in Computer Science, 2522:140–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

series/0558/bibs/2522/25220140.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2522/25220140.pdf.

Delgado:2002:ARE

[2840] M. Delgado, M. J. Martín-Bautista, D. Sánchez, J. M. Serrano, and M. A. Vila. Association rule extraction for text mining. Lecture Notes in Computer Science, 2522: 154–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2522/25220154.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2522/25220154.
pdf.

DellAcqua:2002:UPI

[2841] Pierangelo Dell’Acqua, Luís Moniz Pereira, and Aida Vitória. User preference information in query answering. Lecture Notes in Computer Science, 2522:163–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2522/25220163.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2522/25220163.pdf.

Dubois:2002:BFQ

[2842] Didier Dubois and Henri Prade. Bipolarity in flexible querying. Lecture Notes in Computer Science, 2522: 174–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2522/25220174.htm;   http:

                                        //link.springer.de/link/service/
                                        series/0558/papers/2522/25220174.
                                        pdf.

Fomichov:2002:TKC

[2843] Vladimir A. Fomichov. Theory of K-calculuses as a powerful and flexible mathematical framework for building ontologies and designing natural language processing systems. Lecture Notes in Computer Science, 2522:183–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2522/25220183.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2522/25220183.pdf.

Garruzzo:2002:XCX

[2844] Salvatore Garruzzo, Stefano Modafferi, Domenico Rosaci, and Domenico Ursino. X-compass: An XML agent for supporting user navigation on the Web. Lecture Notes in Computer Science, 2522:197–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2522/25220197.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2522/25220197.pdf.

Kang:2002:PSJ

[2845] Myoung-Soo Kang, Seung-Kyu Ko, Kyun Koh, and Yoon-Chul Choy. A parallel spatial join processing for distributed spatial databases. Lecture Notes in Computer Science, 2522:212–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2522/25220212.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2522/25220212.pdf.

Kapetanios:2002:QCT

[2846] E. Kapetanios and P. Groenewoud.

Query construction through meaningful suggestions of terms. Lecture Notes in Computer Science, 2522: 226–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2522/25220226.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2522/25220226.
pdf.

Keogh:2002:AVQ

[2847] Eamonn Keogh, Harry Hochheiser, and Ben Shneiderman. An augmented visual query mechanism for finding patterns in time series data. Lecture Notes in Computer Science, 2522:240–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2522/25220240.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2522/25220240.pdf.

Koprulu:2002:STQ

[2848] Mesru Köprülü, Nihan Kesim Çiçekli, and Adnan Yazıcı. Spatio-temporal querying in video databases. Lecture Notes in Computer Science, 2522: 251–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2522/25220251.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2522/25220251.
                                        pdf.

Krajci:2002:SBU

[2849] S. Krajči, R. Lencses, J. Medina, M. Ojeda-Aciego, and P. Vojtáš. A similarity-based unification model for flexible querying. Lecture Notes in Computer Science, 2522: 263–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2522/25220263.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2522/25220263.
                                        pdf.

Lalmas:2002:FMD

[2850] Mounia Lalmas. A formal model for data fusion. Lecture Notes in Computer Science, 2522:274–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2522/25220274.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2522/25220274.pdf.

Munoz-Escoi:2002:FMC

[2851] Francesc D. Muñoz-Escoí, Luis Irún-Briz, Pablo Galdámez, José M. Bernabéu-Aubán, Jordi Bataller, M. Carmen Bañuls, and Hendrik Decker. Flexible management of consistency and availability of networked data replications. Lecture Notes in Computer Science, 2522:

289–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2522/25220289.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2522/25220289.
pdf.

Parimala:2002:QMS

[2852] N. Parimala and T. V. Vijay Kumar.

Querying multidatabase systems using SIQL. Lecture Notes in Computer Science, 2522:301–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2522/25220301.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2522/25220301.pdf.

Sicilia:2002:ERD

[2853] Miguel-Angel Sicilia, Elena García, Paloma Díaz, and Ignacio Aedo. Extending relational data access programming libraries for fuzziness: The fJDBC framework. Lecture Notes in Computer Science, 2522: 314–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2522/25220314.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2522/25220314.
pdf.

Stuckenschmidt:2002:ATQ

[2854] Heiner Stuckenschmidt and Frank van Harmelen. Approximating terminological queries. Lecture Notes in Computer Science, 2522:329–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2522/25220329.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2522/25220329.pdf.

deTre:2002:CAE

[2855] Guy de Tré, Rita de Caluwe, Jörg Verstraete, and Axel Hallez. Conjunctive aggregation of extended possibilistic truth values and flexible database querying. Lecture Notes in Computer Science, 2522:344–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2522/25220344.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2522/25220344.pdf.

Trigoni:2002:IQF

[2856] Agathoniki Trigoni. Interactive query formulation in semistructured databases. Lecture Notes in Computer Science, 2522:356–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2522/25220356.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2522/25220356.pdf.

Villadsen:2002:PQA

[2857] Jørgen Villadsen. Paraconsistent query answering systems. Lecture Notes in Computer Science, 2522: 370–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2522/25220370.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2522/25220370.
pdf.

Anonymous:2002:AIco

[2858] Anonymous. Author index. Lecture Notes in Computer Science, 2522: 385–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2522/2522auth.pdf.

Thorpe:2002:URS

[2859] Simon Thorpe. Ultra-rapid scene categorization with a wave of spikes. Lecture Notes in Computer Science, 2525:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250001.pdf.

Hansen:2002:BMS

[2860] Thorsten Hansen and Heiko Neumann. A biologically motivated scheme for robust junction detection. Lecture Notes in Computer Science, 2525:16–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250016.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250016.pdf.

Kolesnik:2002:ITS

[2861] Marina Kolesnik, Alexander Barlit, and Evgeny Zubkov. Iterative tuning of simple cells for contrast invariant edge enhancement. Lecture Notes in Computer Science, 2525:27–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250027.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250027.pdf.

Pellegrino:2002:HSF

[2862] Felice Andrea Pellegrino, Walter Vanzella, and Vincent Torre. How the spatial filters of area V1 can be used for a nearly ideal edge detection. Lecture Notes in Computer Science, 2525:38–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250038.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250038.pdf.

Grigorescu:2002:ICD

[2863] Cosmin Grigorescu, Nicolai Petkov, and Michel A. Westenberg. Improved contour detection by nonclassical receptive field inhibition. Lecture Notes in Computer Science, 2525:50–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250050.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2525/25250050.pdf.

Hugues:2002:CDS

[2864] Etienne Hugues, Florent Guilleux, and Olivier Rochel. Contour detection by synchronization of integrate-and-fire neurons. Lecture Notes in Computer Science, 2525:60–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250060.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250060.pdf.

Ishihara:2002:RSS

[2865] Yukio Ishihara and Satoru Morita.

Reading speed and superiority of right visual field on foveated vision. Lecture Notes in Computer Science, 2525:70–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250070.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250070.pdf.

Mundhenk:2002:MCI

[2866] T. Nathan Mundhenk and Laurent Itti. A model of contour integration in early visual cortex. Lecture Notes in Computer Science, 2525:80–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250080.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2525/25250080.pdf.

Santos:2002:CCC

[2867] Luís M. Santos and J. M. Hans du Buf. Computational cortical cell models for continuity and texture. Lecture Notes in Computer Science, 2525:90–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250090.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250090.pdf.

Thielscher:2002:NMH

[2868] Axel Thielscher, Anna Schuboe, and Heiko Neumann. A neural model of human texture processing: Texture segmentation vs. visual search. Lecture Notes in Computer Science, 2525:99–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250099.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250099.pdf.

Ouadfel:2002:UIS

[2869] Salima Ouadfel and Mohamed Batouche. Unsupervised image segmentation using a colony of cooperating ants. Lecture Notes in Computer Science, 2525:109–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250109.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2525/25250109.pdf.

Wundrich:2002:IRG

[2870] Ingo J. Wundrich, Christoph von der Malsburg, and Rolf P. Würtz. Image reconstruction from Gabor magnitudes. Lecture Notes in Computer Science, 2525:117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250117.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250117.pdf.

Bernardino:2002:BSA

[2871] Alexandre Bernardino and José Santos-Victor. A binocular stereo algorithm for log-polar foveated systems. Lecture Notes in Computer Science, 2525:127–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250127.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250127.pdf.

Dahlem:2002:RIO

[2872] Markus A. Dahlem and Florentin Wörgötter. Rotation-invariant optical flow by gaze-depended retino-cortical mapping. Lecture Notes in Computer Science, 2525:137–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250137.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2525/25250137.pdf.

Zanker:2002:AMS

[2873] Johannes M. Zanker and Jochen Zeil. An analysis of the motion signal distributions emerging from locomotion through a natural environment. Lecture Notes in Computer Science, 2525:146–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250146.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250146.pdf.

Giese:2002:PBM

[2874] Martin A. Giese. Prototypes of biological movements in brains and machines. Lecture Notes in Computer Science, 2525:157–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250157.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250157.pdf.

Franz:2002:IIE

[2875] Matthias O. Franz and Javaan S. Chahl. Insect-inspired estimation of self-motion. Lecture Notes in Computer Science, 2525:171–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250171.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250171.pdf.

Langer:2002:TTO

[2876] Michael S. Langer and Richard Mann. Tracking through optical snow. Lecture Notes in Computer Science, 2525:181–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250181.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250181.pdf.

Ben-Shahar:2002:CVF

[2877] Ohad Ben-Shahar, Patrick S. Huggins, and Steven W. Zucker. On computing visual flows with boundaries: The case of shading and edges. Lecture Notes in Computer Science, 2525:189–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250189.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250189.pdf.

Kalberer:2002:BMS

[2878] Gregor A. Kalberer, Pascal Müller, and Luc Van Gool. Biological motion of speech. Lecture Notes in Computer Science, 2525:199–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250199.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250199.pdf.

Kersten:2002:OPG

[2879] Daniel Kersten. Object perception: Generative image models and Bayesian inference. Lecture Notes in Computer Science, 2525:207–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2525/25250207.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2525/25250207.pdf.

Kimia:2002:RPM

[2880] Benjamin Kimia and Amir Tamrakar. The role of propagation and medial geometry in human vision. Lecture Notes in Computer Science, 2525:219–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2525/25250219.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2525/25250219.pdf.

Elder:2002:ESC

[2881] James H. Elder. Ecological statistics of contour grouping. Lecture Notes in Computer Science, 2525: 230–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2525/25250230.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2525/25250230.
                                        pdf.

Kruger:2002:SSO

[2882] Norbert Krüger and Florentin Wörgötter. Statistics of second order multi-modal feature events and their exploitation in biological and artificial visual systems. Lecture Notes in Computer Science, 2525:239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250239.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250239.pdf.

Sinha:2002:QRR

[2883] Pawan Sinha. Qualitative representations for recognition. Lecture Notes in Computer Science, 2525: 249–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2525/25250249.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2525/25250249.
pdf.

Oliva:2002:SCD

[2884] Aude Oliva and Antonio Torralba.

Scene-centered description from spatial envelope properties. Lecture Notes in Computer Science, 2525: 263–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2525/25250263.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2525/25250263.
pdf.

Knoblich:2002:VCH

[2885] Ulf Knoblich, Maximilian Riesenhuber, David J. Freedman, Earl K. Miller, and Tomaso Poggio. Visual categorization: How the monkey brain does it. Lecture Notes in Computer Science, 2525:273–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2525/25250273.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2525/25250273.pdf.

Maass:2002:NAT

[2886] Wolfgang Maass, Robert Legenstein, and Henry Markram. A new approach towards vision suggested by biologically realistic neural microcircuit models. Lecture Notes in Computer Science, 2525:282–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2525/25250282.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2525/25250282.pdf.

Bolme:2002:ILC

[2887] David S. Bolme and Bruce A. Draper. Interpreting LOC cell responses. Lecture Notes in Computer Science, 2525:294–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2525/25250294.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2525/25250294.pdf.

Bayerl:2002:NMV

[2888] Pierre Bayerl and Heiko Neumann. Neural mechanisms of visual flow integration and segregation — insights from the pinna-brelstaff illusion and variations of it. Lecture Notes in Computer Science, 2525: 301–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2525/25250301.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2525/25250301.
pdf.

Kogo:2002:RSS

[2889] Naoki Kogo, Christoph Strecha, Rik Fransen, Geert Caenen, Johan Wagemans, and Luc Van Gool. Reconstruction of subjective surfaces from occlusion cues. Lecture Notes in Computer Science, 2525:311–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250311.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250311.pdf.

Kruger:2002:EOR

[2890] Norbert Krüger, Thomas Jäger, and Christian Perwass. Extraction of object representations from stereo image sequences utilizing statistical and deterministic regularities in visual data. Lecture Notes in Computer Science, 2525:322–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250322.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250322.pdf.

Cheoi:2002:MEO

[2891] Kyungjoo Cheoi and Yillbyung Lee. A method of extracting objects of interest with possible broad application in computer vision. Lecture Notes in Computer Science, 2525: 331–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2525/25250331.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2525/25250331.
                                        pdf.

deCunha:2002:MUI

[2892] Darryl de Cunha, Leila Eadie, Benjamin Adams, and David Hawkes. Medical ultrasound image similarity measurement by human visual system (HVS) modelling. Lecture Notes in Computer Science, 2525: 340–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2525/25250340.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2525/25250340.
                                        pdf.

Friedrich:2002:SPD

[2893] Gil Friedrich and Yehezkel Yeshurun. Seeing people in the dark: Face recognition in infrared images. Lecture Notes in Computer Science, 2525:348–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2525/25250348.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2525/25250348.pdf.

Neumann:2002:MIC

[2894] Titus R. Neumann. Modeling insect compound eyes: Space-variant spherical vision. Lecture Notes in Computer Science, 2525:360–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250360.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250360.pdf.

Park:2002:FEG

[2895] Kang Ryoung Park, Jeong Jun Lee, and Jaihie Kim. Facial and eye gaze detection. Lecture Notes in Computer Science, 2525:368–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250368.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250368.pdf.

Loos:2002:CLO

[2896] Hartmut S. Loos and Christoph von der Malsburg. 1-click learning of object models for recognition. Lecture Notes in Computer Science, 2525:377–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250377.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250377.pdf.

Serre:2002:ROS

[2897] Thomas Serre, Maximilian Riesenhuber, Jennifer Louie, and Tomaso Poggio. On the role of objectspecific features for real world object recognition in biological vision. Lecture Notes in Computer Science, 2525:387–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250387.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250387.pdf.

Hamker:2002:ODN

[2898] Fred H. Hamker and James Worcester. Object detection in natural scenes by feedback. Lecture Notes in Computer Science, 2525:398–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250398.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250398.pdf.

Koike:2002:SGS

[2899] Takahiko Koike and Jun Saiki. Stochastic guided search model for search asymmetries in visual search tasks. Lecture Notes in Computer Science, 2525:408–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250408.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250408.pdf.

Park:2002:BIS

[2900] Sang-Jae Park, Jang-Kyoo Shin, and Minho Lee. Biologically inspired saliency map model for bottomup visual attention. Lecture Notes in Computer Science, 2525:418– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2525/25250418.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2525/25250418.
pdf.

Sun:2002:HSO

[2901] Yaoru Sun and Robert Fisher. Hierarchical selectivity for object-based visual attention. Lecture Notes in Computer Science, 2525:427–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250427.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250427.pdf.

Tsotsos:2002:AML

[2902] John K. Tsotsos, Marc Pomplun, Yueju Liu, Julio C. Martinez-Trujillo, and Evgueni Simine. Attending to motion: Localizing and classifying motion patterns in image sequences. Lecture Notes in Computer Science, 2525:439–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250439.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250439.pdf.

Navalpakkam:2002:GOA

[2903] Vidhya Navalpakkam and Laurent Itti. A goal oriented attention guidance model. Lecture Notes in Computer Science, 2525:453–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2525/25250453.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2525/25250453.pdf.

Ramstrom:2002:VAU

[2904] Ola Ramström and Henrik I. Christensen. Visual attention using game theory. Lecture Notes in Computer Science, 2525:462–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2525/25250462.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2525/25250462.pdf.

Walther:2002:ASO

[2905] Dirk Walther, Laurent Itti, Maximilian Riesenhuber, Tomaso Poggio, and Christof Koch. Attentional selection for object recognition — A gentle way. Lecture Notes in Computer Science, 2525:472–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2525/25250472.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2525/25250472.pdf.

vanderWilligen:2002:AOT

[2906] Robert Frans van der Willigen and Mark von Campenhausen. Audiooculomotor transformation. Lecture Notes in Computer Science, 2525: 480–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2525/25250480.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2525/25250480.
pdf.

Graf:2002:GCH

[2907] Arnulf B. A. Graf and Felix A. Wichmann. Gender classification of human faces. Lecture Notes in Computer Science, 2525:491–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250491.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250491.pdf.

Hwang:2002:FRP

[2908] Bon-Woo Hwang and Seong-Whan Lee.

Face reconstruction from partial information based on a morphable face model. Lecture Notes in Computer Science, 2525:501–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250501.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250501.pdf.

Kim:2002:DFC

[2909] Jeounghoon Kim. Dynamics of face categorization. Lecture Notes in Computer Science, 2525:511– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2525/25250511.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2525/25250511.
pdf.

Kumar:2002:RED

[2910] Vinay P. Kumar and Tomaso Poggio. Recognizing expressions by direct estimation of the parameters of a pixel morphable model. Lecture Notes in Computer Science, 2525:519–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250519.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250519.pdf.

Ilg:2002:MMS

[2911] Winfried Ilg and Martin Giese. Modeling of movement sequences based on hierarchical spatial-temporal correspondence of movement primitives. Lecture Notes in Computer Science, 2525:528–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250528.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250528.pdf.

Giese:2002:ASS

[2912] Martin A. Giese, Barbara Knappmeyer, and Heinrich H. Bülthoff. Automatic synthesis of sequences of human movements by linear combination of learned example patterns. Lecture Notes in Computer Science, 2525:538–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250538.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250538.pdf.

Bray:2002:AHM

[2913] Alistair Bray. An adaptive hierarchical model of the ventral visual pathway implemented on a mobile robot. Lecture Notes in Computer Science, 2525:548–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250548.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250548.pdf.

Chung:2002:NRP

[2914] Daesu Chung, Reid Hirata, T. Nathan Mundhenk, Jen Ng, Rob J. Peters, Eric Pichon, April Tsui, Tong Ventrice, Dirk Walther, Philip Williams, and Laurent Itti. A new robotics platform for neuromorphic vision: Beobots. Lecture Notes in Computer Science, 2525:558–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250558.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2525/25250558.pdf.

Natale:2002:LAO

[2915] Lorenzo Natale, Sajit Rao, and Giulio Sandini. Learning to act on objects. Lecture Notes in Computer Science, 2525:567–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2525/25250567.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2525/25250567.pdf.

Rushton:2002:EDV

[2916] Simon K. Rushton, Jia Wen, and Robert S. Allison. Egocentric direction and the visual guidance of robot locomotion background, theory and implementation. Lecture Notes in Computer Science, 2525:576–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2525/25250576.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2525/25250576.pdf.

Zufferey:2002:EVB

[2917] Jean-Christophe Zufferey, Dario Floreano, Matthijs van Leeuwen, and Tancredi Merenda. Evolving visionbased flying robots. Lecture Notes in Computer Science, 2525:592– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2525/25250592.htm;   http:

//link.springer.de/link/service/
series/0558/papers/2525/25250592.
pdf.

Kang:2002:ODC

[2918] Seonghoon Kang and Seong-Whan Lee. Object detection and classification for outdoor walking guidance system. Lecture Notes in Computer Science, 2525:601–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2525/25250601.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250601.pdf.

Yu:2002:UHB

[2919] Chen Yu and Dana H. Ballard. Understanding human behaviors based on eye-head-hand coordination. Lecture Notes in Computer Science, 2525: 611–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2525/25250611.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2525/25250611.
pdf.

Sturzl:2002:VBH

[2920] Wolfgang Stürzl and Hanspeter A.

Mallot. Vision-based homing with a panoramic stereo sensor. Lecture Notes in Computer Science, 2525: 620–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2525/25250620.htm;   http:
//link.springer.de/link/service/

                                        series/0558/papers/2525/25250620.
                                        pdf.

Edelman:2002:ULV

[2921] Shimon Edelman, Nathan Intrator, and Judah S. Jacobson. Unsupervised learning of visual structure. Lecture Notes in Computer Science, 2525:629–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2525/25250629.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2525/25250629.pdf.

Schwaninger:2002:RFC

[2922] Adrian Schwaninger, Janek S. Lobmaier, and Stephan M. Collishaw. Role of featural and configural information in familiar and unfamiliar face recognition. Lecture Notes in Computer Science, 2525:643–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2525/25250643.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2525/25250643.pdf.

Wallraven:2002:VBR

[2923] Christian Wallraven, Adrian Schwaninger, Sandra Schuhmacher, and Heinrich H. Bülthoff. View-based recognition of faces in man and machine: Re-visiting inter-extra-ortho. Lecture Notes in Computer Science, 2525:651–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2525/25250651.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2525/25250651.pdf.

Anonymous:2002:AIcp

[2924] Anonymous. Author index. Lecture Notes in Computer Science, 2525: 661–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2525/2525auth.pdf.

Alonso:2002:KDU

[2925] Fernando Alonso, África López-Illescas, Loïc Martínez, Cesar Montes, and Juan P. Valente. Knowledge discovery using medical data mining. Lecture Notes in Computer Science, 2526:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2526/25260001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2526/25260001.pdf.

Cammarota:2002:ASP

[2926] Camillo Cammarota, Giuseppe Guarini, and Maria Ambrosini. Analysis of stationary periods of heart rate via symbolic dynamics. Lecture Notes in Computer Science, 2526: 13–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2526/25260013.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2526/25260013.
pdf.

Caputo:2002:MDU

[2927] B. Caputo, E. La Torre, and G. E. Gigante. Microcalcification detection using a kernel Bayes classifier. Lecture Notes in Computer Science, 2526:20–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2526/25260020.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2526/25260020.pdf.

Kaczmarek:2002:UCT

[2928] Elzbieta Kaczmarek, Edyta Nieruchalska, Michal Michalak, Aldona Wozniak, and Wieslawa Salwa-Zurawska. The use of classification trees for analysis of kidney images in computer assisted microscopy. Lecture Notes in Computer Science, 2526:32–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2526/25260032.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2526/25260032.pdf.

Paetz:2002:NFB

[2929] Jürgen Paetz and Björn Arlt. A neuro-fuzzy based alarm system for septic shock patients with a comparison to medical scores. Lecture Notes in Computer Science, 2526: 42–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2526/25260042.htm;   http:
                                        //link.springer.de/link/service/

series/0558/papers/2526/25260042.
pdf.

Sauerbrei:2002:DFR

[2930] Willi Sauerbrei and Patrick Royston.

Determination of functional relationships for continuous variables by using a multivariable fractional polynomial approach. Lecture Notes in Computer Science, 2526:53–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2526/25260053.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2526/25260053.pdf.

Aksenova:2002:NOM

[2931] Tetyana I. Aksenova, Olga K. Chibirova, Alim-Louis Benabid, and Alessandro E. P. Villa. Nonlinear oscillation models for spike separation. Lecture Notes in Computer Science, 2526:61–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2526/25260061.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2526/25260061.pdf.

Baratta:2002:MGI

[2932] G. Baratta, F. Barcellona, G. Lucidi, and A. M. Bersani. Modeling glucoseinsulin behavior in ill patients (DM Type2). Lecture Notes in Computer Science, 2526:71–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2526/25260071.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2526/25260071.pdf.

Cho:2002:EAM

[2933] Yongho Cho, Hosung Kim, Myoung ju Jeon, Dongjoo Kang, Hyungsik Choi, Jong min Lee, In young Kim, and Sun I. Kim. ECG acquisition and management system for knowledge discovery in database: Data modeling, design, and implementation. Lecture Notes in Computer Science, 2526:79–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2526/25260079.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2526/25260079.pdf.

Fazi:2002:CBC

[2934] Paola Fazi, Daniela Luzi, Fabrizio L. Ricci, and Marco Vignetti. The conceptual basis of WITH, a collaborative writer system of clinical trials. Lecture Notes in Computer Science, 2526:86–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2526/25260086.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2526/25260086.pdf.

Jager:2002:PMC

[2935] Bernd Jäger and Karl-Ernst Biebler. Problems of the model choice in classes of mendelian inheritance models. Lecture Notes in Computer Science, 2526:98–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2526/25260098.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2526/25260098.pdf.

Shim:2002:CMC

[2936] Eun Bo Shim, Chan Hyun Youn, Thomas Heldt, Roger D. Kamm, and Roger G. Mark. Computational modeling of the cardiovascular system after fontan procedure. Lecture Notes in Computer Science, 2526:105–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2526/25260105.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2526/25260105.pdf.

Allegrini:2002:SLT

[2937] Paolo Allegrini, Rita Balocchi, Santi Chillemi, Paolo Grigolini, Luigi Palatella, and Giacomo Raffaelli. Short- and long-term statistical properties of heartbeat time-series in healthy and pathological subjects. Lecture Notes in Computer Science, 2526:115–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2526/25260115.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2526/25260115.pdf.

Barbaro:2002:IOR

[2938] Vincenzo Barbaro, Pietro Bartolini, Giovanni Calcagnini, Federica Censi, Antonio Michelucci, and Samantha Poli. An index of organization of the right atrium during atrial fibrillation: Effects of internal cardioversion. Lecture Notes in Computer Science, 2526:127–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2526/25260127.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2526/25260127.pdf.

Jankowski:2002:MAE

[2939] S. Jankowski, J. Tijink, G. Vumbaca, M. Balsi, and G. Karpinski. Morphological analysis of ECG holter recordings by support vector machines. Lecture Notes in Computer Science, 2526:134–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2526/25260134.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2526/25260134.pdf.

Pagani:2002:MSA

[2940] Jacopo Pagani, Maria Pia Villa, Giovanni Calcagnini, Enrica Lombardozzi, Federica Censi, Samantha Poli, Pietro Bartolini, Vincenzo Barbaro, and Roberto Ronchetti. Monitoring of sleep apnea in children using pulse transit time. Lecture Notes in Computer Science, 2526:144–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2526/25260144.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2526/25260144.pdf.

Zbilut:2002:LNE

[2941] Joseph P. Zbilut, Peter A. Santucci, Shiow-Ying Yang, and Janice L. Podolski. Linear and nonlinear evaluation of ventricular arrhythmias. Lecture Notes in Computer Science, 2526:151–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2526/25260151.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2526/25260151.pdf.

Zimatore:2002:ISD

[2942] Giovanna Zimatore, Alessandro Giuliani, Stavros Hatzopoulos, Alessandro Martini, and Alfredo Colosimo. Invariant and subject-dependent features of otoacoustic emissions. Lecture Notes in Computer Science, 2526: 158–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2526/25260158.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2526/25260158.
pdf.

Balsi:2002:STE

[2943] M. Balsi, V. Cimagalli, G. Cruccu, G. D. Iannetti, A. Londei, and P. L. Romanelli. Short term evaluation of brain activities in fMRI data by spatiotemporal independent component analysis. Lecture Notes in Computer Science, 2526:167–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2526/25260167.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2526/25260167.pdf.

Behnke:2002:IIS

[2944] Michael Behnke, Tim Eckmanns, and Henning Rüden. Implementation of an IT system for the support of a hospital infection surveillance system. Lecture Notes in Computer Science, 2526:177–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2526/25260177.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2526/25260177.pdf.

Capozza:2002:ANN

[2945] M. Capozza, G. D. Iannetti, J. J. Marx, G. Cruccu, and N. Accornero. An artificial neural network for 3D localization of brainstem functional lesions. Lecture Notes in Computer Science, 2526:186–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2526/25260186.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2526/25260186.pdf.

Lobbert:2002:DRV

[2946] Sebastian Löbbert and Steffen Märkle. Dependent rendering: Visualizing multiple properties of an object. Lecture Notes in Computer Science, 2526:198–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2526/25260198.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2526/25260198.pdf.

Walter:2002:ADM

[2947] Thomas Walter and Jean-Claude Klein. Automatic detection of microaneurysms in color fundus images of the human retina by means of the bounding box closing. Lecture Notes in Computer Science, 2526:210–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2526/25260210.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2526/25260210.pdf.

Anonymous:2002:AIcq

[2948] Anonymous. Author index. Lecture Notes in Computer Science, 2526: 221–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2526/2526indx.pdf.

Bressan:2002:INB

[2949] Marco Bressan and Jordi Vitrià.

Improving naive Bayes using classconditional ICA. Lecture Notes in Computer Science, 2527:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2527/25270001.htm;   http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2527/25270001.pdf.

Pons-Porrata:2002:DET

[2950] Aurora Pons-Porrata, Rafael Berlanga-Llavori, and José Ruíz-Shulcloper. Detecting events and topics by using temporal references. Lecture Notes in Computer Science, 2527: 11–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2527/25270011.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2527/25270011.
                                        pdf.

Guo:2002:ANS

[2951] Gongde Guo, Hui Wang, and David Bell. Asymmetric neighbourhood selection and support aggregation for effective classification. Lecture Notes in Computer Science, 2527: 21–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2527/25270021.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2527/25270021.
                                        pdf.

Alcobe:2002:ILT

[2952] Josep Roure Alcobé. Incremental learning of tree augmented naive Bayes classifiers. Lecture Notes in Computer Science, 2527:32–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270032.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2527/25270032.pdf.

Macias-Macias:2002:CPG

[2953] Miguel Macías-Macías, Carlos J.

García-Orellana, Horacio M. González-Velasco, Ramón Gallardo-Caballero, and Antonio Serrano-Pérez. A comparison of PCA and GA selected features for cloud field classification. Lecture Notes in Computer Science, 2527:42–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270042.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270042.pdf.

Perez:2002:FNC

[2954] José Ramón Quevedo Pérez, María Dolores García, and Elena Montañés. Filtering noisy continuous labeled examples. Lecture Notes in Computer Science, 2527:50–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270050.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270050.pdf.

Abidi:2002:DAH

[2955] Syed Sibte Raza Abidi. Designing adaptive hypermedia for Internet portals: a personalization strategy featuring case base reasoning with compositional adaptation. Lecture Notes in Computer Science, 2527:60–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270060.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270060.pdf.

Solorio:2002:ICA

[2956] Thamar Solorio and Olac Fuentes. Improving classification accuracy of large test sets using the ordered classification algorithm. Lecture Notes in Computer Science, 2527:70–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270070.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270070.pdf.

Soares:2002:CSS

[2957] Carlos Soares and Pavel Brazdil. A comparative study of some issues concerning algorithm recommendation using ranking methods. Lecture Notes in Computer Science, 2527:80–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270080.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270080.pdf.

Brito:2002:PCF

[2958] Luís Brito and José Neves. Properties and complexity in feasible logicbased argumentation for electronic commerce. Lecture Notes in Computer Science, 2527:90–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270090.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270090.pdf.

Fdez-Riverola:2002:HCM

[2959] Florentino Fdez-Riverola and Juan M.

Corchado. A hybrid CBR model for forecasting in complex domains. Lecture Notes in Computer Science, 2527:101–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270101.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270101.pdf.

Truck:2002:GMI

[2960] Isis Truck, Amel Borgi, and Herman Akdag. Generalized modifiers as an interval scale: Towards adaptive colorimetric alterations. Lecture Notes in Computer Science, 2527:111–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270111.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270111.pdf.

Ramirez:2002:CGR

[2961] R. Carolina Medina Ramírez, Olivier Corby, and Rose Dieng-Kuntz. A conceptual graph and RDF(S) approach for representing and querying document content. Lecture Notes in Computer Science, 2527:

121–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2527/25270121.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2527/25270121.
                                        pdf.

Moreno:2002:AOM

[2962] Ginés Moreno. Automatic optimization of multi-paradigm declarative programs. Lecture Notes in Computer Science, 2527:131–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270131.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270131.pdf.

Cordero:2002:SLE

[2963] Pablo Cordero, Manolo Enciso, Angel Mora, and Inmaculada P. de Guzmán. SL_{F D} logic: Elimination of data redundancy in knowledge representation. Lecture Notes in Computer Science, 2527:141–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270141.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270141.pdf.

Olmedo-Aguirre:2002:IID

[2964] Oscar Olmedo-Aguirre and Guillermo Morales-Luna. Indeed: Interactive deduction on Horn clause theories. Lecture Notes in Computer Science, 2527:151–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270151.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270151.pdf.

deGuzman:2002:RTR

[2965] Inma P. de Guzmán, Manuel Ojeda-Aciego, and Agustín Valverde. Restricted Δ-trees and reduction theorems in multiple-valued logics. Lecture Notes in Computer Science, 2527: 161–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2527/25270161.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2527/25270161.
pdf.

Ceballos:2002:MCA

[2966] R. Ceballos, R. M. Gasca, Carmelo Del Valle, and Miguel Toro. Max-CSP approach for software diagnosis. Lecture Notes in Computer Science, 2527:172–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270172.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270172.pdf.

deCampos:2002:LSM

[2967] Luis Miguel de Campos, Juan Manuel Fernández-Luna, and Jose Miguel Puerta. Local search methods for learning Bayesian networks using a modified neighborhood in the space of DAGs. Lecture Notes in Computer Science, 2527:182–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270182.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270182.pdf.

Gutierrez-Naranjo:2002:QMM

[2968] Miguel A. Gutiérrez-Naranjo, José A. Alonso-Jiménez, and Joaquín Borrego-Díaz. A quasi-metric for machine learning. Lecture Notes in Computer Science, 2527:193–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270193.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270193.pdf.

Estruch:2002:SEL

[2969] Victor Estruch, Cesar Ferri, Jose Hernández-Orallo, and Maria Jose Ramírez-Quintana. Shared ensemble learning using multi-trees. Lecture Notes in Computer Science, 2527:204–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270204.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270204.pdf.

Cano:2002:GAC

[2970] J. R. Cano, O. Cordón, F. Herrera, and L. Sánchez. A GRASP algorithm for clustering. Lecture Notes in Computer Science, 2527:214–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270214.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270214.pdf.

Monekosso:2002:APQ

[2971] Ndedi Monekosso and Paolo Remagnino. An analysis of the pheromone Q-learning algorithm. Lecture Notes in Computer Science, 2527:224–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270224.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270224.pdf.

Ruiz:2002:SEF

[2972] Roberto Ruiz, Jesús S. Aguilar-Ruiz, and José C. Riquelme. SOAP: Efficient feature selection of numeric attributes. Lecture Notes in Computer Science, 2527:233–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270233.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270233.pdf.

Sicilia:2002:DFR

[2973] Miguel Ángel Sicilia, José Antonio Gutiérrez, and Elena García. Designing fuzzy relations in orthogonal persistence object-oriented database engines. Lecture Notes in Computer Science, 2527:243–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270243.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270243.pdf.

FernandezdelPozo:2002:IFO

[2974] Juan A. Fernández del Pozo and Concha Bielza. An interactive framework for open queries in decision support systems. Lecture Notes in Computer Science, 2527:254–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270254.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270254.pdf.

Garza:2002:IFD

[2975] Luis E. Garza, Francisco Cantú, and Salvador Acevedo. Integration of fault detection and diagnosis in a probabilistic logic framework. Lecture Notes in Computer Science, 2527:265–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270265.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270265.pdf.

Bosch:2002:SPT

[2976] Alfonso Bosch, Francisco Guil, Carmen Martinez, and Roque Marin. Series-parallel and tree-decomposition approaches for fuzzy constraint networks. Lecture Notes in Computer Science, 2527:275–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270275.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270275.pdf.

Velloso:2002:CDU

[2977] Maria Luiza F. Velloso and Flávio J.

de Souza. Change-detection using contextual information and fuzzy entropy principle. Lecture Notes in Computer Science, 2527:285–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270285.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270285.pdf.

Alcala:2002:ISL

[2978] Rafael Alcalá, Jorge Casillas, Oscar Cordón, and Francisco Herrera. Improving simple linguistic fuzzy models by means of the weighted COR methodology. Lecture Notes in Computer Science, 2527:294–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270294.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270294.pdf.

Ortega:2002:SAS

[2979] Juan Antonio Ortega, Rafael M.

Gasca, Miguel Toro, and Jesús Torres. A semiquantitative approach to study semiqualitative systems. Lecture Notes in Computer Science, 2527:303–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270303.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270303.pdf.

deToro:2002:MOO

[2980] Francisco de Toro, Eduardo Ros, Sonia Mota, and Julio Ortega. Multiobjective optimization evolutionary algorithms applied to paroxysmal atrial fibrillation diagnosis based on the k-nearest neighbours classifier. Lecture Notes in Computer Science, 2527:313–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270313.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270313.pdf.

Mendes:2002:PSG

[2981] Alexandre Mendes, Paulo França, Pablo Moscato, and Vinícius Garcia. Population studies for the gate matrix layout problem. Lecture Notes in Computer Science, 2527: 319–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2527/25270319.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2527/25270319.
pdf.

Affenzeller:2002:NGH

[2982] Michael Affenzeller. New generic hybrids based upon genetic algorithms. Lecture Notes in Computer Science, 2527:329–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270329.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270329.pdf.

Ribeiro:2002:GAB

[2983] Simone J. M. Ribeiro and João C. P.

da Silva. Genetic algorithms and biological images restoration: Preliminary report. Lecture Notes in Computer Science, 2527:340–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270340.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270340.pdf.

Bacardit:2002:EMA

[2984] Jaume Bacardit and Josep Maria Garrell. Evolution of multi-adaptive discretization intervals for a rule-based genetic learning system. Lecture Notes in Computer Science, 2527: 350–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2527/25270350.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2527/25270350.
pdf.

Cutello:2002:IAC

[2985] Vincenzo Cutello and Giuseppe Nicosia. An immunological approach to combinatorial optimization problems. Lecture Notes in Computer Science, 2527:361–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270361.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270361.pdf.

Garcia:2002:GAS

[2986] Jose Manuel Garcia, Sebastian Lozano, Fernando Guerrero, and Ignacio Eguia. A genetic algorithm for solving a production and delivery scheduling problem with time windows. Lecture Notes in Computer Science, 2527:371–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270371.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270371.pdf.

Cordon:2002:PSC

[2987] O. Cordón, F. Herrera, J. de la Montaña, A. M. Sánchez, and P. Villar. A prediction system for cardiovascularity diseases using genetic fuzzy rule-based systems. Lecture Notes in Computer Science, 2527:381–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270381.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2527/25270381.pdf.

Herrera:2002:MCP

[2988] F. Herrera, M. Lozano, E. Pérez, A. M. Sánchez, and P. Villar. Multiple crossover per couple with selection of the two best offspring: An experimental study with the BLX-α crossover operator for real-coded genetic algorithms. Lecture Notes in Computer Science, 2527:392–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270392.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270392.pdf.

Aguilar:2002:ARF

[2989] Jose Aguilar. Adaptive random fuzzy cognitive maps. Lecture Notes in Computer Science, 2527: 402–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2527/25270402.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2527/25270402.
pdf.

Osuna:2002:CHF

[2990] Edgar Osuna and Osberth De Castro. Convex hull in feature space for support vector machines. Lecture Notes in Computer Science, 2527: 411–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2527/25270411.htm;   http:

                                        //link.springer.de/link/service/
                                        series/0558/papers/2527/25270411.
                                        pdf.

Rojas:2002:ANN

[2991] Fernando Rojas, M. R. Alvarez, C. G. Puntonet, and Ruben Martin-Clemente. Applying neural networks and genetic algorithms to the separation of sources. Lecture Notes in Computer Science, 2527: 420–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2527/25270420.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2527/25270420.
                                        pdf.

Aligue:2002:NAP

[2992] Francisco J. López Aligué, Isabel Acevedo Sotoca, and Ignacio Alvarez Troncoso. A neural associative pattern classifier. Lecture Notes in Computer Science, 2527:430–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270430.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270430.pdf.

Nunez:2002:RER

[2993] Haydemar Núñez, Cecilio Angulo, and Andreu Català. Rule extraction from radial basis function networks by using support vectors. Lecture Notes in Computer Science, 2527:440–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270440.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270440.pdf.

Vazquez:2002:EPA

[2994] Elisa Guerrero Vázquez, Joaquín Pizarro Junquera, Andrés Yáñez Escolano, and Pedro L. Riaño Galindo. Empirical performance assessment of nonlinear model selection techniques. Lecture Notes in Computer Science, 2527:450–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270450.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270450.pdf.

Merino:2002:ENN

[2995] Enrique Dominguez Merino and José Muñoz Perez. An efficient neural network algorithm for the p-median problem. Lecture Notes in Computer Science, 2527:460–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270460.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270460.pdf.

Preciado:2002:ICN

[2996] Víctor M. Preciado. Improving cellular nonlinear network computational capabilities. Lecture Notes in Computer Science, 2527:470–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270470.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270470.pdf.

Diez:2002:LAP

[2997] J. Díez, J. J. del Coz, O. Luaces, F. Goyache, J. Alonso, A. M. Peña, and A. Bahamonde. Learning to assess from pair-wise comparisons. Lecture Notes in Computer Science, 2527:481–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270481.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270481.pdf.

Montanes:2002:FTS

[2998] Elena Montañés, José R. Quevedo, Maria M. Prieto, and César O. Menéndez. Forecasting time series combining machine learning and Box-Jenkins time series. Lecture Notes in Computer Science, 2527: 491–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2527/25270491.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2527/25270491.
                                        pdf.

Crespo:2002:GSN

[2999] J. L. Crespo, R. J. Duro, and F. López-Peña. Gaussian synapse networks for hyperspectral image segmentation. Lecture Notes in Computer Science, 2527:500–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270500.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270500.pdf.

Merida-Casermeiro:2002:AMR

[3000] Enrique Mérida-Casermeiro, José Muñoz-Pérez, and M. A. García-Bernal. An associative multivalued recurrent network. Lecture Notes in Computer Science, 2527: 509–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2527/25270509.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2527/25270509.
pdf.

Rocco:2002:MLM

[3001] Claudio M. Rocco and José A.

Moreno. Machine learning models for online dynamic security assessment of electric power systems. Lecture Notes in Computer Science, 2527:519–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270519.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270519.pdf.

Grana:2002:EDH

[3002] Manuel Graña, Bogdan Raducanu, Peter Sussner, and Gerhard Ritter. On endmember detection in hyperspectral images with morphological associative memories. Lecture Notes in Computer Science, 2527:526–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270526.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270526.pdf.

Garcia:2002:ALM

[3003] Cristina García and José Alí Moreno. Application of learning machine methods to 3D object modeling. Lecture Notes in Computer Science, 2527:536–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270536.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270536.pdf.

Gimenez-Lugo:2002:EIA

[3004] Gustavo A. Giménez-Lugo, Analia Amandi, Jaime Simão Sichman, and Daniela Godoy. Enriching information agents’ knowledge by ontology comparison: a case study. Lecture Notes in Computer Science, 2527:546–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270546.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270546.pdf.

Lopes:2002:NAA

[3005] Fernando Lopes, Nuno Mamede, A. Q. Novais, and Helder Coelho. Negotiation among autonomous computational agents. Lecture Notes in Computer Science, 2527:556–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270556.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270556.pdf.

Lahlouhi:2002:IAD

[3006] Ammar Lahlouhi, Zaidi Sahnoun, Med Lamine Benbrahim, and Abdelouahab Boussaha. Interface agents development in MASA for human integration in multiagent systems. Lecture Notes in Computer Science, 2527:566–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270566.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270566.pdf.

Bianchi:2002:CDR

[3007] Reinaldo A. C. Bianchi and Anna H. R. Costa. Comparing distributed reinforcement learning approaches to learn agent coordination. Lecture Notes in Computer Science, 2527:575–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270575.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270575.pdf.

LopezyLopez:2002:ESA

[3008] Fabiola López y López and Michael Luck. Empowered situations of autonomous agents. Lecture Notes in Computer Science, 2527:585– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2527/25270585.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2527/25270585.
                                        pdf.

daSilva:2002:DAM

[3009] João L. T. da Silva and Yves Demazeau. Distributed agenda management through decentralised vowels co-ordination approach. Lecture Notes in Computer Science, 2527: 596–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2527/25270596.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2527/25270596.
                                        pdf.

Gomez-Sanz:2002:MMA

[3010] Jorge J. Gómez-Sanz and Juan Pavón. Meta-modelling in agent oriented software engineering. Lecture Notes in Computer Science, 2527: 606–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2527/25270606.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2527/25270606.
                                        pdf.

dosSantos:2002:MAS

[3011] Nelson dos Santos, Flávio Miguel Varejão, and Orivaldo de Lira Tavares. Multi-agent systems and network management — A positive experience on Unix environments. Lecture Notes in Computer Science, 2527:616–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270616.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270616.pdf.

Vauvert:2002:FSO

[3012] Guillaume Vauvert and Amal El Fallah-Seghrouchni. Formal specification of opinions applied to the consensus problem. Lecture Notes in Computer Science, 2527: 625–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2527/25270625.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2527/25270625.
pdf.

Vilares:2002:PNB

[3013] Jésus Vilares, F. Mario Barcala, Miguel A. Alonso, Jorge Graña, and Manuel Vilares. Practical NLP-based text indexing. Lecture Notes in Computer Science, 2527: 635–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2527/25270635.htm;   http:
//link.springer.de/link/service/

                                        series/0558/papers/2527/25270635.
                                        pdf.

Munoz:2002:DDR

[3014] Rafael Muñoz and Andrés Montoyo. Definite description resolution enrichment with WordNet domain labels. Lecture Notes in Computer Science, 2527:645–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270645.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270645.pdf.

Molina:2002:HMM

[3015] Antonio Molina, Ferran Pla, and Encarna Segarra. A hidden Markov model approach to word sense disambiguation. Lecture Notes in Computer Science, 2527:655–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270655.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270655.pdf.

Castro:2002:SCA

[3016] María José Castro and Emilio Sanchis. A simple connectionist approach to language understanding in a dialogue system. Lecture Notes in Computer Science, 2527:664–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270664.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2527/25270664.pdf.

Carreras:2002:WCS

[3017] Xavier Carreras, Lluís Màrquez, and Lluís Padró. Wide-coverage Spanish named entity extraction. Lecture Notes in Computer Science, 2527:674–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270674.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270674.pdf.

Penas:2002:TRT

[3018] Anselmo Peñas, Felisa Verdejo, and Julio Gonzalo. Terminology retrieval: Towards a synergy between thesaurus and free text searching. Lecture Notes in Computer Science, 2527:684–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270684.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270684.pdf.

Alonso:2002:MPT

[3019] Miguel A. Alonso, Vicente Carrillo, and Víctor J. Díaz. Mixed parsing of tree insertion and tree adjoining grammars. Lecture Notes in Computer Science, 2527:694–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270694.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2527/25270694.pdf.

Roman:2002:TOD

[3020] Norton Trevisan Roman and Ariadne Maria Brito Rizzoni Carvalho. Task oriented dialogue processing using multiagents theory. Lecture Notes in Computer Science, 2527:704–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270704.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270704.pdf.

Neco:2002:AAN

[3021] Ramón P. Ñeco, Óscar Reinoso, José M. Azorín, José M. Sabater, M. Asunción Vicente, and Nicolás García. Automatic adaptation of a natural language interface to a robotic system. Lecture Notes in Computer Science, 2527:714–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270714.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270714.pdf.

Gounon:2002:SCT

[3022] Patricia Gounon and Benoît Lemaire. Semantic comparison of texts for learning environments. Lecture Notes in Computer Science, 2527: 724–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/

0558/bibs/2527/25270724.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2527/25270724.
pdf.

Read:2002:PMP

[3023] Timothy Read, Elena Bárcena, Beatriz Barros, and Felisa Verdejo. I-PETER: Modelling personalised diagnosis and material selection for an online English course. Lecture Notes in Computer Science, 2527:734–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270734.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270734.pdf.

Rosatelli:2002:ABS

[3024] Marta C. Rosatelli, John A. Self, and Tulio V. D. Christofoletti. An agent-based system for supporting learning from case studies. Lecture Notes in Computer Science, 2527:745–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270745.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270745.pdf.

Webber:2002:EDC

[3025] Carine Webber and Sylvie Pesty. Emergent diagnosis via coalition formation. Lecture Notes in Computer Science, 2527:755–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2527/25270755.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270755.pdf.

Gama:2002:AB

[3026] João Gama and Gladys Castillo. Adaptive Bayes. Lecture Notes in Computer Science, 2527:765–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270765.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270765.pdf.

Campos:2002:DSA

[3027] Antonio M. Campos and Daniel F. García. A dynamic scheduling algorithm for real-time expert systems. Lecture Notes in Computer Science, 2527:775–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270775.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270775.pdf.

Jayme:2002:PKB

[3028] Salvador Carlos De Lara Jayme, Raul Garduño Ramírez, Marino Sánchez Parra, Luis Castelo Cuevas, and Marco Antonio Carretero R. A process knowledge-based controller for maneuverability improvement of a nonlinear industrial process. Lecture Notes in Computer Science, 2527:785–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270785.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270785.pdf.

deJesusGonzalez-Noriega:2002:AOD

[3029] Luis de Jesús González-Noriega and Pablo H. Ibargüengoytia. An architecture for online diagnosis of gas turbines. Lecture Notes in Computer Science, 2527:795–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270795.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270795.pdf.

Sebastia:2002:SVP

[3030] Laura Sebastia, Eva Onaindia, and Eliseo Marzal. STeLLa v2.0: Planning with intermediate goals. Lecture Notes in Computer Science, 2527:805–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270805.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270805.pdf.

Varela:2002:SHS

[3031] Ramiro Varela and Elena Soto.

Scheduling as heuristic search with state space reduction. Lecture Notes in Computer Science, 2527: 815–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/

                                        0558/bibs/2527/25270815.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2527/25270815.
                                        pdf.

Sapena:2002:DIO

[3032] Oscar Sapena and Eva Onaindía. Domain-independent online planning for STRIPS domains. Lecture Notes in Computer Science, 2527: 825–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2527/25270825.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2527/25270825.
                                        pdf.

Valle:2002:PBM

[3033] Carmelo Del Valle, Miguel Toro, Rafael Ceballos, and Jesús S. Aguilar-Ruiz. A pomset-based model for estimating workcells’ setups in assembly sequence planning. Lecture Notes in Computer Science, 2527:835–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270835.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270835.pdf.

Pomares:2002:NMS

[3034] H. Pomares, I. Rojas, and J. González. New methodology for structure identification of fuzzy controllers in real time. Lecture Notes in Computer Science, 2527:845–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2527/25270845.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270845.pdf.

Avila-Garcia:2002:CVB

[3035] Orlando Avila-García and Lola Cañamero.

Comparing a voting-based policy with winner-takes-all to perform action selection in motivational agents. Lecture Notes in Computer Science, 2527:855–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270855.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270855.pdf.

Valverde:2002:BAB

[3036] Francisco L. Valverde, Nicolás Guil, and Jose Muñoz. Bayesian approach based on geometrical features for validation and tuning of solution in deformable models. Lecture Notes in Computer Science, 2527:865–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270865.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270865.pdf.

DiazdeLeon:2002:RCA

[3037] Rocío Díaz de León and L. Enrique Sucar. Recognition of continuous activities. Lecture Notes in Computer Science, 2527:875–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270875.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270875.pdf.

Gracia:2002:KCS

[3038] Luis Gracia and Josep Tornero. Kinematic control system for car-like vehicles. Lecture Notes in Computer Science, 2527:882–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270882.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270882.pdf.

Lorenzo:2002:HBS

[3039] Javier Lorenzo and Mario Hernández. Habituation based on spectrogram analysis. Lecture Notes in Computer Science, 2527:893–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2527/25270893.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2527/25270893.pdf.

Canas:2002:DSH

[3040] José M. Cañas and Vicente Matellán. Dynamic schema hierarchies for an autonomous robot. Lecture Notes in Computer Science, 2527: 903–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2527/25270903.htm;   http:
                                        //link.springer.de/link/service/

series/0558/papers/2527/25270903.
pdf.

Saez:2002:MCL

[3041] Juan Manuel Sáez and Francisco Escolano. Monte Carlo localization in 3D maps using stereo vision. Lecture Notes in Computer Science, 2527:913–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270913.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270913.pdf.

Merchan:2002:CSS

[3042] P. Merchán, A. Adán, S. Salamanca, and C. Cerrada. 3D complex scenes segmentation from a single range image using virtual exploration. Lecture Notes in Computer Science, 2527:923–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270923.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2527/25270923.pdf.

Lozano:2002:RII

[3043] Miguel Angel Lozano and Francisco Escolano. Recognizing indoor images with unsupervised segmentation and graph matching. Lecture Notes in Computer Science, 2527:933–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2527/25270933.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2527/25270933.pdf.

Lopez-Martinez:2002:VBS

[3044] M. López-Martínez, M. Vargas, and F. R. Rubio. Vision-based system for the safe operation of a solar power tower plant. Lecture Notes in Computer Science, 2527: 943–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2527/25270943.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2527/25270943.
                                        pdf.

Anonymous:2002:AIcr

[3045] Anonymous. Author index. Lecture Notes in Computer Science, 2527: 953–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2527/2527auth.pdf.

Brandes:2002:SDO

[3046] Ulrik Brandes, Markus Eiglsperger, Michael Kaufmann, and Dorothea Wagner. Sketch-driven orthogonal graph drawing. Lecture Notes in Computer Science, 2528:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2528/25280001.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2528/25280001.pdf.

Kaufmann:2002:MMM

[3047] Michael Kaufmann and Roland Wiese.

Maintaining the mental map for circular drawings. Lecture Notes in Computer Science, 2528:12–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2528/25280012.htm;   http://link.
springer.de/link/service/series/
0558/papers/2528/25280012.pdf.

Diehl:2002:GTC

[3048] Stephan Diehl and Carsten Görg.

Graphs, they are changing. Lecture Notes in Computer Science, 2528: 23–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2528/25280023.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2528/25280023.
pdf.

Cornelsen:2002:DGT

[3049] Sabine Cornelsen, Thomas Schank, and Dorothea Wagner. Drawing graphs on two and three lines. Lecture Notes in Computer Science, 2528:31–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2528/25280031.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2528/25280031.pdf.

Dujmovic:2002:PWT

[3050] Vida Dujmović, Pat Morin, and David R. Wood. Path-width and threedimensional straight-line grid drawings of graphs. Lecture Notes in Computer Science, 2528:42–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2528/25280042.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2528/25280042.pdf.

Biedl:2002:DOP

[3051] Therese Biedl. Drawing outerplanar graphs in O(n log n) area. Lecture Notes in Computer Science, 2528:54–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2528/25280054.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2528/25280054.pdf.

Binucci:2002:CLO

[3052] Carla Binucci, Walter Didimo, Giuseppe Liotta, and Maddalena Nonato. Computing labeled orthogonal drawings. Lecture Notes in Computer Science, 2528:66–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2528/25280066.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2528/25280066.pdf.

Bachl:2002:CDI

[3053] Sabine Bachl and Franz-Josef Brandenburg. Computing and drawing isomorphic subgraphs. Lecture Notes in Computer Science, 2528:

74–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2528/25280074.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2528/25280074.
pdf.

Abelson:2002:GTM

[3054] David Abelson, Seok-Hee Hong, and Donald E. Taylor. A group-theoretic method for drawing graphs symmetrically. Lecture Notes in Computer Science, 2528:86–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2528/25280086.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2528/25280086.pdf.

Healy:2002:BCA

[3055] Patrick Healy and Nikola S. Nikolov.

A branch-and-cut approach to the directed acyclic graph layering problem. Lecture Notes in Computer Science, 2528:98–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2528/25280098.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2528/25280098.pdf.

Fiala:2002:GSD

[3056] Jivrí Fiala, Jan Kratochvíl, and Andrzej Proskurowski. Geometric systems of disjoint representatives. Lecture Notes in Computer Science, 2528:110–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2528/25280110.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2528/25280110.pdf.

Dujmovic:2002:EFP

[3057] Vida Dujmović and Sue Whitesides. An efficient fixed parameter tractable algorithm for 1-sided crossing minimization. Lecture Notes in Computer Science, 2528:118–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2528/25280118.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2528/25280118.pdf.

Barth:2002:SEB

[3058] Wilhelm Barth, Michael Jünger, and Petra Mutzel. Simple and efficient bilayer cross counting. Lecture Notes in Computer Science, 2528: 130–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2528/25280130.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2528/25280130.
                                        pdf.

Giacomo:2002:OST

[3059] Emilio Di Giacomo, Giuseppe Liotta, and Maurizio Patrignani. Orthogonal 3D shapes of theta graphs. Lecture Notes in Computer Science, 2528:142–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2528/25280142.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2528/25280142.pdf.

Eppstein:2002:STG

[3060] David Eppstein. Separating thickness from geometric thickness. Lecture Notes in Computer Science, 2528: 150–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2528/25280150.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2528/25280150.
pdf.

Giacomo:2002:BEP

[3061] Emilio Di Giacomo, Walter Didimo, Giuseppe Liotta, and Stephen K. Wismath. Book embeddings and pointset embeddings of series-parallel digraphs. Lecture Notes in Computer Science, 2528:162–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2528/25280162.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2528/25280162.pdf.

Chanda:2002:CEP

[3062] Amrita Chanda and Ashim Garg.

Compact encodings of planar orthogonal drawings. Lecture Notes in Computer Science, 2528:174– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2528/25280174.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2528/25280174.
                                        pdf.

Sykora:2002:FLC

[3063] Ondrej Sýkora, László A. Székely, and Imrich Vrt’o. Fractional lengths and crossing numbers. Lecture Notes in Computer Science, 2528: 186–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2528/25280186.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2528/25280186.
                                        pdf.

Carmel:2002:DDG

[3064] Liran Carmel, David Harel, and Yehuda Koren. Drawing directed graphs using one-dimensional optimization. Lecture Notes in Computer Science, 2528:193–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2528/25280193.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2528/25280193.pdf.

Harel:2002:GDH

[3065] David Harel and Yehuda Koren. Graph drawing by high-dimensional embedding. Lecture Notes in Computer Science, 2528:207–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2528/25280207.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2528/25280207.pdf.

Gutwenger:2002:APT

[3066] Carsten Gutwenger, Michael Jünger, Sebastian Leipert, Petra Mutzel, Merijam Percan, and René Weiskircher. Advances in C-planarity testing of clustered graphs. Lecture Notes in Computer Science, 2528:220– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2528/25280220.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2528/25280220.
pdf.

Raitner:2002:HLH

[3067] Marcus Raitner. HGV: a library for hierarchies, graphs, and views. Lecture Notes in Computer Science, 2528:236–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2528/25280236.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2528/25280236.pdf.

Rahman:2002:RDP

[3068] Md. Saidur Rahman, Takao Nishizeki, and Shubhashis Ghosh. Rectangular drawings of planar graphs. Lecture Notes in Computer Science, 2528:244–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2528/25280244.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2528/25280244.pdf.

Miura:2002:ERD

[3069] Kazuyuki Miura, Ayako Miyazawa, and Takao Nishizeki. Extended rectangular drawings of plane graphs with designated corners. Lecture Notes in Computer Science, 2528: 256–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2528/25280256.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2528/25280256.
                                        pdf.

Teoh:2002:RTV

[3070] Soon Tee Teoh and Kwan-Liu Ma. RINGS: a technique for visualizing large hierarchies. Lecture Notes in Computer Science, 2528: 268–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2528/25280268.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2528/25280268.
                                        pdf.

Forster:2002:ACR

[3071] Michael Forster. Applying crossing reduction strategies to layered compound graphs. Lecture Notes in Computer Science, 2528:276–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2528/25280276.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2528/25280276.pdf.

Eschbach:2002:CRW

[3072] Thomas Eschbach, Wolfgang Günther, Rolf Drechsler, and Bernd Becker. Crossing reduction by windows optimization. Lecture Notes in Computer Science, 2528:285–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2528/25280285.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2528/25280285.pdf.

Pach:2002:GGN

[3073] János Pach, Rom Pinchasi, Gábor Tardos, and Géza Tóth. Geometric graphs with no self-intersecting path of length three. Lecture Notes in Computer Science, 2528:295–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2528/25280295.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2528/25280295.pdf.

Newton:2002:TNH

[3074] Matthew Newton, Ondrej Sýkora, and Imrich Vrt’o. Two new heuristics for two-sided bipartite graph drawing. Lecture Notes in Computer Science, 2528:312–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2528/25280312.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2528/25280312.pdf.

Garg:2002:SLD

[3075] Ashim Garg and Adrian Rusu. Straight-line drawings of binary trees with linear area and arbitrary aspect ratio. Lecture Notes in Computer Science, 2528:320–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2528/25280320.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2528/25280320.pdf.

Chen:2002:SAO

[3076] Ho-Lin Chen, Chien-Chih Liao, Hsueh-I Lu, and Hsu-Chun Yen. Some applications of orderly spanning trees in graph drawing. Lecture Notes in Computer Science, 2528:332–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2528/25280332.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2528/25280332.pdf.

Buchheim:2002:IWA

[3077] Christoph Buchheim, Michael Jünger, and Sebastian Leipert. Improving Walker’s algorithm to run in linear time. Lecture Notes in Computer Science, 2528:344–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2528/25280344.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2528/25280344.pdf.

Bachl:2002:SDO

[3078] Walter Bachl. Semi-dynamic orthogonal drawings of planar graphs. Lecture Notes in Computer Science, 2528:354–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2528/25280354.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2528/25280354.pdf.

Diguglielmo:2002:GLW

[3079] Gilles Diguglielmo, Eric Durocher, Philippe Kaplan, Georg Sander, and Adrian Vasiliu. Graph layout for workflow applications with ILOG JViews. Lecture Notes in Computer Science, 2528:362–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2528/25280362.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2528/25280362.pdf.

Han:2002:IDV

[3080] Kyungsook Han, Byong-Hyon Ju, and Jong H. Park. InterViewer: Dynamic visualization of protein-protein interactions. Lecture Notes in Computer Science, 2528:364–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2528/25280364.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2528/25280364.pdf.

Baburin:2002:SMS

[3081] Danil E. Baburin. Some modifications of sugiyama approach. Lecture Notes in Computer Science, 2528:366–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2528/25280366.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2528/25280366.pdf.

Dogrusoz:2002:FCM

[3082] Ugur Dogrusoz and Burkay Genc. A framework for complexity management in graph visualization. Lecture Notes in Computer Science, 2528: 368–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2528/25280368.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2528/25280368.
pdf.

Byun:2002:PAP

[3083] Yanga Byun, Euna Jeong, and Kyungsook Han. A partitioned approach to protein interaction mapping. Lecture Notes in Computer Science, 2528:370–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2528/25280370.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2528/25280370.pdf.

Streinu:2002:CPR

[3084] Ileana Streinu and Elif Tosun.

Camera position reconstruction and tight direction networks. Lecture Notes in Computer Science, 2528: 372–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2528/25280372.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2528/25280372.
pdf.

Mutton:2002:DPS

[3085] Paul Mutton and Peter Rodgers.

Demonstration of a preprocessor for the spring embedder. Lecture Notes in Computer Science, 2528: 374–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2528/25280374.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2528/25280374.
pdf.

Brandenburg:2002:GDC

[3086] Franz J. Brandenburg. Graphdrawing contest report. Lecture Notes in Computer Science, 2528: 376–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2528/25280376.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2528/25280376.
pdf.

Arvo:2002:TIG

[3087] James Arvo. Techniques for interactive graph drawing. Lecture Notes in Computer Science, 2528: 380–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2528/25280380.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2528/25280380.
                                        pdf.

Ruskey:2002:DVD

[3088] Frank Ruskey. Drawing Venn diagrams. Lecture Notes in Computer Science, 2528:381–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2528/25280381.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2528/25280381.pdf.

Anonymous:2002:AIcs

[3089] Anonymous. Author index. Lecture Notes in Computer Science, 2528: 383–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2528/2528auth.pdf.

Nunez:2002:EPT

[3090] Manuel Núñez and Ismael Rodríguez. Encoding PAMR into (timed) EF-SMs. Lecture Notes in Computer Science, 2529:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2529/25290001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2529/25290001.pdf.

vonBochmann:2002:SCS

[3091] Gregor von Bochmann. Submodule construction for specifications with input assumptions and output guarantees. Lecture Notes in Computer Science, 2529:17–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2529/25290017.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2529/25290017.pdf.

Brower:2002:CWC

[3092] Ronald W. Brower and Kenneth S.

Stevens. Congruent weak conformance, a partial order among processes. Lecture Notes in Computer Science, 2529:34–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2529/25290034.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2529/25290034.pdf.

Wang:2002:SSS

[3093] Farn Wang and Karsten Schmidt. Symmetric symbolic safety-analysis of concurrent software with pointer data structures. Lecture Notes in Computer Science, 2529:50–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2529/25290050.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2529/25290050.pdf.

Bosnacki:2002:NDF

[3094] Dragan Bošnački. A nested depth first search algorithm for model checking with symmetry reduction. Lecture Notes in Computer Science, 2529:65–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2529/25290065.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2529/25290065.pdf.

Turner:2002:PTT

[3095] Kenneth J. Turner and Qian Bing. Protocol techniques for testing radiotherapy accelerators. Lecture Notes in Computer Science, 2529: 81–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2529/25290081.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2529/25290081.
                                        pdf.

Pickin:2002:STS

[3096] Simon Pickin, Claude Jard, Yves Le Traon, Thierry Jéron, Jean-Marc Jézéquel, and Alain Le Guennec. System test synthesis from UML models of distributed software. Lecture Notes in Computer Science, 2529: 97–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/

0558/bibs/2529/25290097.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2529/25290097.
pdf.

Deussen:2002:FTP

[3097] Peter H. Deussen and Stephan Tobies.

Formal test purposes and the validity of test cases. Lecture Notes in Computer Science, 2529:114–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2529/25290114.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2529/25290114.pdf.

Reiff-Marganiec:2002:ULD

[3098] Stephan Reiff-Marganiec and Kenneth J. Turner. Use of logic to describe enhanced communications services. Lecture Notes in Computer Science, 2529:130–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2529/25290130.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2529/25290130.pdf.

Livadas:2002:FVR

[3099] Carolos Livadas and Nancy A. Lynch.

A formal venture into reliable multicast territory. Lecture Notes in Computer Science, 2529:146–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2529/25290146.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2529/25290146.pdf.

Turner:2002:MSS

[3100] Kenneth J. Turner. Modelling SIP services using CRESS. Lecture Notes in Computer Science, 2529: 162–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2529/25290162.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2529/25290162.
pdf.

Tynjala:2002:VRD

[3101] Teemu Tynjälä, Sari Leppänen, and Vesa Luukkala. Verifying reliable data transmission over UMTS radio interface with high level petri nets. Lecture Notes in Computer Science, 2529:178–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2529/25290178.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2529/25290178.pdf.

Kwiatkowska:2002:VRB

[3102] Marta Kwiatkowska and Gethin Norman. Verifying randomized Byzantine agreement. Lecture Notes in Computer Science, 2529:194–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2529/25290194.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2529/25290194.pdf.

Armando:2002:ASC

[3103] Alessandro Armando and Luca Compagna. Automatic SAT-compilation of protocol insecurity problems via reduction to planning. Lecture Notes in Computer Science, 2529: 210–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2529/25290210.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2529/25290210.
pdf.

Amla:2002:VSM

[3104] Nina Amla, E. Allen Emerson, Kedar S. Namjoshi, and Richard J. Trefler. Visual specifications for modular reasoning about asynchronous systems. Lecture Notes in Computer Science, 2529:226–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2529/25290226.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2529/25290226.pdf.

Audemard:2002:BMC

[3105] G. Audemard, A. Cimatti, A. Kornilowicz, and R. Sebastiani. Bounded model checking for timed systems. Lecture Notes in Computer Science, 2529:243–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2529/25290243.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2529/25290243.pdf.

DuVarney:2002:CWT

[3106] Daniel C. DuVarney and S. Purushothaman Iyer. C Wolf — A toolset for extracting models from C programs. Lecture Notes in Computer Science, 2529:260–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2529/25290260.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2529/25290260.pdf.

Garavel:2002:NGS

[3107] Hubert Garavel and Frédéric Lang. NTIF: a general symbolic model for communicating sequential processes with data. Lecture Notes in Computer Science, 2529:276–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2529/25290276.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2529/25290276.pdf.

Verdejo:2002:BTL

[3108] Alberto Verdejo. Building tools for LOTOS symbolic semantics in maude. Lecture Notes in Computer Science, 2529:292–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2529/25290292.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2529/25290292.pdf.

Giannakopoulou:2002:STI

[3109] Dimitra Giannakopoulou and Flavio Lerda. From states to transitions: Improving translation of LTL formulae to Büchi automata. Lecture Notes in Computer Science, 2529:308–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2529/25290308.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2529/25290308.pdf.

Kristensen:2002:CSL

[3110] Lars Michael Kristensen and Thomas Mailund. A compositional sweepline state space exploration method. Lecture Notes in Computer Science, 2529:327–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2529/25290327.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2529/25290327.pdf.

Ribet:2002:CPS

[3111] Pierre-Olivier Ribet, François Vernadat, and Bernard Berthomieu. On combining the persistent sets method with the covering steps graph method. Lecture Notes in Computer Science, 2529:344–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2529/25290344.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2529/25290344.pdf.

Melamed:2002:IVT

[3112] Victor Melamed, Harry Stuimer, David Wilkins, Lawrence Chang, Kevin Normoyle, and Sutikshan Bhutani. Innovative verification techniques used in the implementation of a thirdgeneration 1.1GHz 64b microprocessor. Lecture Notes in Computer Science, 2529:360–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2529/25290360.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2529/25290360.pdf.

Bogdanov:2002:MTA

[3113] Andrej Bogdanov, Stephen J. Garland, and Nancy A. Lynch. Mechanical translation of I/O automaton specifications into first-order logic. Lecture Notes in Computer Science, 2529:364–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2529/25290364.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2529/25290364.pdf.

Sakunkonchak:2002:VEB

[3114] Thanyapat Sakunkonchak and Masahiro Fujita. Verification of event-based synchronization of SpecC description using difference decision diagrams. Lecture Notes in Computer Science, 2529:369–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2529/25290369.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2529/25290369.pdf.

Palmer:2002:DPO

[3115] Robert Palmer and Ganesh Gopalakrishnan. A distributed partial order reduction algorithm. Lecture Notes in Computer Science, 2529: 370–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2529/25290370.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2529/25290370.
pdf.

Anonymous:2002:AIct

[3116] Anonymous. Author index. Lecture Notes in Computer Science, 2529: 371–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2529/2529auth.pdf.

Cesa-Bianchi:2002:EI

[3117] Nicolo Cesa-Bianchi, Masayuki Numao, and Rudiger Reischuk. Editors’ introduction. Lecture Notes in Computer Science, 2533:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2533/25330001.htm;   http://link.
springer.de/link/service/series/
0558/papers/2533/25330001.pdf.

Hayashi:2002:MBLa

[3118] Susumu Hayashi. Mathematics based on learning. Lecture Notes in Computer Science, 2533:7–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2533/25330007.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2533/25330007.pdf.

Kruse:2002:DMGa

[3119] Rudolf Kruse and Christian Borgelt. Data mining with graphical models. Lecture Notes in Computer Science, 2533:22–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2533/25330022.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2533/25330022.pdf.

Shawe-Taylor:2002:EGMa

[3120] John Shawe-Taylor, Chris Williams, Nello Cristianini, and Jaz Kandola. On the eigenspectrum of the gram matrix and its relationship to the operator eigenspectrum. Lecture Notes in Computer Science, 2533:23–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2533/25330023.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2533/25330023.pdf.

Widmer:2002:SHFa

[3121] Gerhard Widmer. In search of the horowitz factor: Interim report on a musical discovery project. Lecture Notes in Computer Science, 2533:41–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2533/25330041.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2533/25330041.pdf.

Witten:2002:LSSa

[3122] Ian H. Witten. Learning structure from sequences, with applications in a digital library. Lecture Notes in Computer Science, 2533:42–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2533/25330042.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2533/25330042.pdf.

Amano:2002:LMB

[3123] Kazuyuki Amano and Akira Maruoka.

On learning monotone Boolean functions under the uniform distribution. Lecture Notes in Computer Science, 2533:57–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2533/25330057.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2533/25330057.pdf.

Servedio:2002:LEM

[3124] Rocco A. Servedio. On learning embedded midbit functions. Lecture Notes in Computer Science, 2533:69–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2533/25330069.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2533/25330069.pdf.

Bshouty:2002:MAC

[3125] Nader H. Bshouty and Lynn Burroughs. Maximizing agreements and CoAgnostic learning. Lecture Notes in Computer Science, 2533: 83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2533/25330083.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2533/25330083.
                                        pdf.

Gavinsky:2002:OSA

[3126] Dmitry Gavinsky. Optimally-smooth adaptive boosting and application to agnostic learning. Lecture Notes in Computer Science, 2533:98–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2533/25330098.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2533/25330098.pdf.

Kivinen:2002:LMC

[3127] Jyrki Kivinen, Alex J. Smola, and Robert C. Williamson. Large margin classification for moving targets. Lecture Notes in Computer Science, 2533:113–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2533/25330113.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2533/25330113.pdf.

Forster:2002:SPD

[3128] Jürgen Forster and Hans Ulrich Simon. On the smallest possible dimension and the largest possible margin of linear arrangements representing given concept classes uniform distribution. Lecture Notes in Computer Science, 2533:128–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2533/25330128.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2533/25330128.pdf.

Kobler:2002:GDA

[3129] Johannes Köbler and Wolfgang Lindner. A general dimension for approximately learning Boolean functions. Lecture Notes in Computer Science, 2533:139–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2533/25330139.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2533/25330139.pdf.

Kobler:2002:CLC

[3130] Johannes Köbler and Wolfgang Lindner. The complexity of learning concept classes with polynomial general dimension. Lecture Notes in Computer Science, 2533:149–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2533/25330149.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2533/25330149.pdf.

Kalnishkan:2002:APC

[3131] Yuri Kalnishkan and Michael V. Vyugin. On the absence of predictive complexity for some games. Lecture Notes in Computer Science, 2533:164–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2533/25330164.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2533/25330164.pdf.

Grieser:2002:CQI

[3132] Gunter Grieser, Klaus P. Jantke, and Steffen Lange. Consistency queries in information extraction. Lecture Notes in Computer Science, 2533:173–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2533/25330173.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2533/25330173.pdf.

Suzuki:2002:OTT

[3133] Yusuke Suzuki, Takayoshi Shoudai, Tomoyuki Uchida, and Tetsuhiro Miyahara. Ordered term tree languages which are polynomial time inductively inferable from positive data. Lecture Notes in Computer Science, 2533:188–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2533/25330188.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2533/25330188.pdf.

Grieser:2002:RII

[3134] Gunter Grieser. Reflective inductive inference of recursive functions. Lecture Notes in Computer Science, 2533:203–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2533/25330203.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2533/25330203.pdf.

Jain:2002:CEL

[3135] Sanjay Jain, Wolfram Menzel, and Frank Stephan. Classes with easily learnable subclasses. Lecture Notes in Computer Science, 2533: 218–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2533/25330218.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2533/25330218.
pdf.

Harizanov:2002:LVS

[3136] Valentina S. Harizanov and Frank Stephan. On the learnability of vector spaces. Lecture Notes in Computer Science, 2533:233–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2533/25330233.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2533/25330233.pdf.

Martin:2002:LLT

[3137] Eric Martin, Arun Sharma, and Frank Stephan. Learning, logic, and topology in a common framework. Lecture Notes in Computer Science, 2533:248–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2533/25330248.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2533/25330248.pdf.

Furnkranz:2002:PBH

[3138] Johannes Fürnkranz. A pathology of bottom-up hill-climbing in inductive rule learning. Lecture Notes in Computer Science, 2533: 263–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2533/25330263.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2533/25330263.
                                        pdf.

Fronhofer:2002:MRH

[3139] Bertram Fronhöfer and Akihiro Yamamoto. Minimised residue hypotheses in relevant logic. Lecture Notes in Computer Science, 2533: 278–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2533/25330278.htm;   http:
                                        //link.springer.de/link/service/

series/0558/papers/2533/25330278.
pdf.

Uemura:2002:CLC

[3140] Jin Uemura and Masako Sato. Compactness and learning of classes of unions of erasing regular pattern languages. Lecture Notes in Computer Science, 2533:293–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2533/25330293.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2533/25330293.pdf.

Reidenbach:2002:NRI

[3141] Daniel Reidenbach. A negative result on inductive inference of extended pattern languages. Lecture Notes in Computer Science, 2533: 308–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2533/25330308.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2533/25330308.
pdf.

Schmitt:2002:RNN

[3142] Michael Schmitt. RBF neural networks and Descartes’ rule of signs. Lecture Notes in Computer Science, 2533:321–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2533/25330321.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2533/25330321.pdf.

Vovk:2002:AOT

[3143] Vladimir Vovk. Asymptotic optimality of transductive confidence machine. Lecture Notes in Computer Science, 2533:336–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2533/25330336.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2533/25330336.pdf.

Dasgupta:2002:EPA

[3144] Sanjoy Dasgupta, Elan Pavlov, and Yoram Singer. An efficient PAC algorithm for reconstructing a mixture of lines. Lecture Notes in Computer Science, 2533:351–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2533/25330351.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2533/25330351.pdf.

Har-Peled:2002:CCN

[3145] Sariel Har-Peled, Dan Roth, and Dav Zimak. Constraint classification: a new approach to multiclass classification. Lecture Notes in Computer Science, 2533:365–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2533/25330365.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2533/25330365.pdf.

Braess:2002:HAM

[3146] Dietrich Braess, Jürgen Forster, Tomas Sauer, and Hans U. Simon. How to achieve minimax expected Kullback–Leibler distance from an unknown finite distribution. Lecture Notes in Computer Science, 2533:380–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2533/25330380.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2533/25330380.pdf.

Lindgren:2002:CIR

[3147] Tony Lindgren and Henrik Boström.

Classification with intersecting rules. Lecture Notes in Computer Science, 2533:395–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2533/25330395.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2533/25330395.pdf.

Coulom:2002:FNN

[3148] Rémi Coulom. Feedforward neural networks in reinforcement learning applied to high-dimensional motor control. Lecture Notes in Computer Science, 2533:403–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2533/25330403.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2533/25330403.pdf.

Anonymous:2002:AIcu

[3149] Anonymous. Author index. Lecture Notes in Computer Science, 2533: 415–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2533/2533auth.pdf.

Hayashi:2002:MBLb

[3150] Susumu Hayashi. Mathematics based on learning. Lecture Notes in Computer Science, 2534:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2534/25340001.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2534/25340001.pdf.

Kruse:2002:DMGb

[3151] Rudolf Kruse and Christian Borgelt. Data mining with graphical models. Lecture Notes in Computer Science, 2534:2–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2534/25340002.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2534/25340002.pdf.

Shawe-Taylor:2002:EGMb

[3152] John Shawe-Taylor, Chris Williams, Nello Cristianini, and Jaz Kandola. On the eigenspectrum of the gram matrix and its relationship to the operator eigenspectrum. Lecture Notes in Computer Science, 2534:12–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2534/25340012.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2534/25340012.pdf.

Widmer:2002:SHFb

[3153] Gerhard Widmer. In search of the horowitz factor: Interim report on a musical discovery project. Lecture Notes in Computer Science, 2534:13–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2534/25340013.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2534/25340013.pdf.

Witten:2002:LSSb

[3154] Ian H. Witten. Learning structure from sequences, with applications in a digital library. Lecture Notes in Computer Science, 2534:33–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2534/25340033.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2534/25340033.pdf.

Palopoli:2002:DFS

[3155] Luigi Palopoli and Giorgio Terracina.

Discovering frequent structured patterns from string databases: An application to biological sequences. Lecture Notes in Computer Science, 2534:34–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2534/25340034.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2534/25340034.pdf.

Devaney:2002:DHP

[3156] Judith E. Devaney and John G. Hagedorn. Discovery in hydrating plaster using machine learning methods. Lecture Notes in Computer Science, 2534:47–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2534/25340047.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2534/25340047.pdf.

Saito:2002:RQM

[3157] Kazumi Saito, Stephen Bay, and Pat Langley. Revising qualitative models of gene regulation. Lecture Notes in Computer Science, 2534:59–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2534/25340059.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2534/25340059.pdf.

Ghazizadeh:2002:SSE

[3158] Shayan Ghazizadeh and Sudarshan S. Chawathe. SEuS: Structure Extraction Using Summaries. Lecture Notes in Computer Science, 2534: 71–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2534/25340071.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2534/25340071.
pdf.

Inenaga:2002:DBV

[3159] Shunsuke Inenaga, Hideo Bannai, Ayumi Shinohara, Masayuki Takeda, and Setsuo Arikawa. Discovering best variable-length-don’t-care patterns. Lecture Notes in Computer Science, 2534:86–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2534/25340086.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2534/25340086.pdf.

Ting:2002:SEC

[3160] Kai Ming Ting. A study on the effect of class distribution using cost-sensitive learning. Lecture Notes in Computer Science, 2534:98–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2534/25340098.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2534/25340098.pdf.

Hilario:2002:MCA

[3161] Melanie Hilario. Model complexity and algorithm selection in classification. Lecture Notes in Computer Science, 2534:113–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2534/25340113.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2534/25340113.pdf.

Elomaa:2002:EPL

[3162] Tapio Elomaa and J. T. Lindgren. Experiments with projection learning. Lecture Notes in Computer Science, 2534:127–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2534/25340127.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2534/25340127.pdf.

Peng:2002:IDC

[3163] Yonghong Peng, Peter A. Flach, Carlos Soares, and Pavel Brazdil. Improved dataset characterisation for metalearning. Lecture Notes in Computer Science, 2534:141–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2534/25340141.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2534/25340141.pdf.

Frank:2002:RCL

[3164] Eibe Frank, Geoffrey Holmes, Richard Kirkby, and Mark Hall. Racing committees for large datasets. Lecture Notes in Computer Science, 2534:153–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2534/25340153.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2534/25340153.pdf.

Ferri:2002:EMC

[3165] C. Ferri, J. Hernández-Orallo, and M. J. Ramírez-Quintana. From ensemble methods to comprehensible models. Lecture Notes in Computer Science, 2534:165–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2534/25340165.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2534/25340165.pdf.

Danks:2002:LCS

[3166] David Danks. Learning the causal structure of overlapping variable sets. Lecture Notes in Computer Science, 2534:178–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2534/25340178.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2534/25340178.pdf.

Holena:2002:ELR

[3167] Martin Holeňa. Extraction of logical rules from data by means of piecewise-linear neural networks. Lecture Notes in Computer Science, 2534:192–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2534/25340192.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2534/25340192.pdf.

Saito:2002:SNN

[3168] Kazumi Saito and Ryohei Nakano. Structuring neural networks through bidirectional clustering of weights. Lecture Notes in Computer Science, 2534:206–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2534/25340206.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2534/25340206.pdf.

Maruyama:2002:TDA

[3169] Osamu Maruyama, Takayoshi Shoudai, and Satoru Miyano. Toward drawing an atlas of hypothesis classes: Approximating a hypothesis via another hypothesis model. Lecture Notes in Computer Science, 2534: 220–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2534/25340220.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2534/25340220.
                                        pdf.

Okada:2002:DSU

[3170] Takashi Okada. Datascape survey using the cascade model. Lecture Notes in Computer Science, 2534: 233–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2534/25340233.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2534/25340233.
                                        pdf.

Ichise:2002:LHS

[3171] Ryutaro Ichise, Daniel Shapiro, and Pat Langley. Learning hierarchical skills from observation. Lecture Notes in Computer Science, 2534:247–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2534/25340247.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2534/25340247.pdf.

Ruosaari:2002:IAD

[3172] Salla Ruosaari and Jaakko Hollmén.

Image analysis for detecting faulty spots from microarray images. Lecture Notes in Computer Science, 2534: 259–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2534/25340259.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2534/25340259.
pdf.

deHoon:2002:IGR

[3173] Michiel de Hoon, Seiya Imoto, and Satoru Miyano. Inferring gene regulatory networks from time-ordered gene expression data using differential equations. Lecture Notes in Computer Science, 2534:267–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2534/25340267.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2534/25340267.pdf.

Masotti:2002:DTC

[3174] Daniele Masotti. DNA-tract curvature profile reconstruction: a fragment flipping algorithm. Lecture Notes in Computer Science, 2534: 275–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2534/25340275.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2534/25340275.
                                        pdf.

Kitamoto:2002:EMM

[3175] Asanobu Kitamoto. Evolution map: Modeling state transition of typhoon image sequences by spatio-temporal clustering. Lecture Notes in Computer Science, 2534:283–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2534/25340283.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2534/25340283.pdf.

Halova:2002:SSR

[3176] Jaroslava Halova, Premysl Zak, Pavel Stopka, Tomoaki Yuzuri, Yukino Abe, Kazuhisa Sakakibara, Hiroko Suezawa, and Minoru Hirota. Structuresweetness relationships of aspartame derivatives by GUHA. Lecture Notes in Computer Science, 2534: 291–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2534/25340291.htm;   http:
                                        //link.springer.de/link/service/

series/0558/papers/2534/25340291.
pdf.

Fang:2002:HAC

[3177] Xiaoshan Fang and Huanye Sheng.

A hybrid approach for Chinese named entity recognition. Lecture Notes in Computer Science, 2534: 297–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2534/25340297.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2534/25340297.
pdf.

Moon:2002:EWS

[3178] Yoo-Jin Moon, Minkoo Kim, Youngho Hwang, Pankoo Kim, and Kijoon Choi. Extraction of word senses from human factors in knowledge discovery. Lecture Notes in Computer Science, 2534:302–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2534/25340302.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2534/25340302.pdf.

Li:2002:EPD

[3179] Fang Li, Huanye Sheng, and Dongmo Zhang. Event pattern discovery from the stock market bulletin. Lecture Notes in Computer Science, 2534:310–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2534/25340310.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2534/25340310.pdf.

Yang:2002:ECU

[3180] Jihoon Yang and Sung-Yong Park. Email categorization using fast machine learning algorithms. Lecture Notes in Computer Science, 2534: 316–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2534/25340316.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2534/25340316.
                                        pdf.

Haraguchi:2002:DMA

[3181] Makoto Haraguchi, Shigetora Nakano, and Masaharu Yoshioka. Discovery of maximal analogies between stories. Lecture Notes in Computer Science, 2534:324–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2534/25340324.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2534/25340324.pdf.

Yamada:2002:AWG

[3182] Yasuhiro Yamada, Daisuke Ikeda, and Sachio Hirokawa. Automatic wrapper generation for multilingual Web resources. Lecture Notes in Computer Science, 2534:332–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2534/25340332.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2534/25340332.pdf.

Bao:2002:CMK

[3183] Yongguang Bao and Naohiro Ishii.

Combining multiple K-nearest neighbor classifiers for text classification by reducts. Lecture Notes in Computer Science, 2534:340–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2534/25340340.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2534/25340340.pdf.

Kontos:2002:ACK

[3184] John Kontos, Areti Elmaoglou, and Ioanna Malagardi. ARISTA causal knowledge discovery from texts. Lecture Notes in Computer Science, 2534:348–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2534/25340348.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2534/25340348.pdf.

Rousseaux:2002:KDA

[3185] Francis Rousseaux and Alain Bonardi.

Knowledge discovery as applied to music: Will music Web retrieval revolutionize musicology? Lecture Notes in Computer Science, 2534: 356–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2534/25340356.htm;   http:
//link.springer.de/link/service/

                                        series/0558/papers/2534/25340356.
                                        pdf.

Maruster:2002:PMD

[3186] Laura Maruster, A. J. M. M. (Ton) Weijters, W. M. P. (Wil) van der Aalst, and Antal van den Bosch. Process mining: Discovering direct successors in process logs. Lecture Notes in Computer Science, 2534:364–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2534/25340364.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2534/25340364.pdf.

Nakamura:2002:EAC

[3187] Makoto Nakamura and Satoshi Tojo. The emergence of artificial Creole by the EM algorithm. Lecture Notes in Computer Science, 2534: 374–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2534/25340374.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2534/25340374.
                                        pdf.

Lartillot:2002:GMP

[3188] Olivier Lartillot. Generalized musical pattern discovery by analogy from local viewpoints. Lecture Notes in Computer Science, 2534:382–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2534/25340382.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2534/25340382.pdf.

Fu:2002:UGA

[3189] Zhiwei Fu. Using genetic algorithmsbased approach for better decision trees: a computational study. Lecture Notes in Computer Science, 2534:390–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2534/25340390.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2534/25340390.pdf.

Hoppner:2002:HFA

[3190] Frank Höppner. Handling feature ambiguity in knowledge discovery from time series. Lecture Notes in Computer Science, 2534:398–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2534/25340398.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2534/25340398.pdf.

Zhao:2002:CFM

[3191] Haiyan Zhao, Zhenjiang Hu, and Masato Takeichi. A compositional framework for mining longest ranges. Lecture Notes in Computer Science, 2534:406–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2534/25340406.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2534/25340406.pdf.

Jorge:2002:PPO

[3192] Alipio Jorge, João Poças, and Paulo Azevedo. Post-processing operators for browsing large sets of association rules. Lecture Notes in Computer Science, 2534:414–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2534/25340414.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2534/25340414.pdf.

Matsuda:2002:MPS

[3193] Takashi Matsuda, Hiroshi Motoda, Tetsuya Yoshida, and Takashi Washio. Mining patterns from structured data by beam-wise graph-based induction. Lecture Notes in Computer Science, 2534:422–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2534/25340422.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2534/25340422.pdf.

Krogel:2002:FSP

[3194] Mark-A. Krogel and Stefan Wrobel. Feature selection for propositionalization. Lecture Notes in Computer Science, 2534:430–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2534/25340430.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2534/25340430.pdf.

Narahashi:2002:SCB

[3195] Masaki Narahashi and Einoshin Suzuki. Subspace clustering based on compressibility. Lecture Notes in Computer Science, 2534:435– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2534/25340435.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2534/25340435.
pdf.

Magnani:2002:ETD

[3196] Lorenzo Magnani, Matteo Piazza, and Riccardo Dossena. The extratheoretical dimension of discovery. extracting knowledge by abduction. Lecture Notes in Computer Science, 2534:441–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2534/25340441.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2534/25340441.pdf.

Saito:2002:DPW

[3197] Hitomi Saito and Kazuhisa Miwa. Discovery process on the WWW: Analysis based on a theory of scientific discovery. Lecture Notes in Computer Science, 2534:449–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2534/25340449.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2534/25340449.pdf.

Piscopo:2002:IVD

[3198] Carlotta Piscopo and Mauro Birattari. Invention vs. discovery. Lecture Notes in Computer Science, 2534:457–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2534/25340457.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2534/25340457.pdf.

Anonymous:2002:AIcv

[3199] Anonymous. Author index. Lecture Notes in Computer Science, 2534: 463–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2534/2534auth.pdf.

Arumugam:2002:AWM

[3200] Subramanian Arumugam, Abdelsalam (Sumi) Helal, and Amar Nalla. aZIMAS: Web mobile agent system. Lecture Notes in Computer Science, 2535:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2535/25350001.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2535/25350001.pdf.

Delamaro:2002:MCN

[3201] Márcio Delamaro and Gian Pietro Picco. Mobile code in .NET: a porting experience. Lecture Notes in Computer Science, 2535:16–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2535/25350016.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2535/25350016.pdf.

Hayashi:2002:MAL

[3202] Hisashi Hayashi, Kenta Cho, and Akihiko Ohsuga. Mobile agents and logic programming. Lecture Notes in Computer Science, 2535: 32–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2535/25350032.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2535/25350032.
pdf.

Roth:2002:EMS

[3203] Volker Roth. Empowering mobile software agents. Lecture Notes in Computer Science, 2535:47–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2535/25350047.htm;   http://link.
springer.de/link/service/series/
0558/papers/2535/25350047.pdf.

Vigna:2002:IDS

[3204] Giovanni Vigna, Bryan Cassell, and Dave Fayram. An intrusion detection system for aglets. Lecture Notes in Computer Science, 2535:64–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2535/25350064.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2535/25350064.pdf.

Stoops:2002:FGI

[3205] Luk Stoops, Tom Mens, and Theo D’Hondt. Fine-grained interlaced code loading for mobile systems. Lecture Notes in Computer Science, 2535:78–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2535/25350078.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2535/25350078.pdf.

Ahn:2002:ISR

[3206] JinHo Ahn, Sung-Gi Min, and Chong-Sun Hwang. Improving scalability of replicated services in mobile agent systems. Lecture Notes in Computer Science, 2535:93–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2535/25350093.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2535/25350093.pdf.

Grimstrup:2002:TIM

[3207] Arne Grimstrup, Robert Gray, David Kotz, Maggie Breedy, Marco Carvalho, Thomas Cowin, Daria Chacón, Joyce Barton, Chris Garrett, and Martin Hofmann. Toward interoperability of mobile-agent systems. Lecture Notes in Computer Science, 2535:106–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2535/25350106.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2535/25350106.pdf.

Zaini:2002:MIS

[3208] Norliza Zaini and Luc Moreau. Mobile intermediaries supporting information sharing between mobile users. Lecture Notes in Computer Science, 2535:121–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2535/25350121.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2535/25350121.pdf.

Cao:2002:MAE

[3209] Jiannong Cao, Xianbing Wang, and Jie Wu. A mobile agent enabled fully distributed mutual exclusion algorithm. Lecture Notes in Computer Science, 2535:138–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2535/25350138.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2535/25350138.pdf.

Binder:2002:USM

[3210] Walter Binder and Balázs Lichtl. Using a secure mobile object kernel as operating system on embedded devices to support the dynamic upload of applications. Lecture Notes in Computer Science, 2535:154–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2535/25350154.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2535/25350154.pdf.

Carvalho:2002:SFD

[3211] Marco Carvalho and Maggie Breedy. Supporting flexible data feeds in dynamic sensor grids through mobile agents. Lecture Notes in Computer Science, 2535:171–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2535/25350171.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2535/25350171.pdf.

Satoh:2002:PML

[3212] Ichiro Satoh. Physical mobility and logical mobility in ubiquitous computing environments. Lecture Notes in Computer Science, 2535: 186–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2535/25350186.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2535/25350186.
                                        pdf.

Anonymous:2002:AIcw

[3213] Anonymous. Author index. Lecture Notes in Computer Science, 2535: 203–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2535/2535auth.pdf.

Verbeke:2002:FPP

[3214] Jerome Verbeke, Neelakanth Nadgir, Greg Ruetsch, and Ilya Sharapov. Framework for peer-to-peer distributed computing in a heterogeneous, decentralized environment. Lecture Notes in Computer Science, 2536:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2536/25360001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2536/25360001.pdf.

Li:2002:GBM

[3215] Yaohang Li and Michael Mascagni.

Grid-based Monte Carlo application. Lecture Notes in Computer Science, 2536:13–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2536/25360013.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2536/25360013.pdf.

vonLaszewski:2002:GSB

[3216] Gregor von Laszewski, Branko Ruscic, Patrick Wagstrom, Sriram Krishnan, Kaizar Amin, Sandeep Nijsure, Sandra Bittner, Reinhardt Pinzon, John C. Hewson, Melita L. Morton, Mike Minkoff, and Al Wagner. A Grid service-based active thermochemical table framework. Lecture Notes in Computer Science, 2536:25–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2536/25360025.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2536/25360025.pdf.

Allen:2002:GEA

[3217] Gabrielle Allen, Dave Angulo, Tom Goodale, Thilo Kielmann, André Merzky, Jarek Nabrzysky, Juliusz Pukacki, Michael Russell, Thomas Radke, Ed Seidel, John Shalf, and Ian Taylor. GridLab: Enabling applications on the Grid. Lecture Notes in Computer Science, 2536:39–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2536/25360039.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2536/25360039.pdf.

Bell:2002:SDG

[3218] William H. Bell, David G. Cameron, Luigi Capozza, A. Paul Millar, Kurt Stockinger, and Floriano Zini. Simulation of dynamic Grid replication strategies in OptorSim. Lecture Notes in Computer Science, 2536: 46–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2536/25360046.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2536/25360046.
                                        pdf.

Dogan:2002:SIT

[3219] Atakan Dogan and Füsun Özgüner. Scheduling independent tasks with QoS requirements in Grid computing with time-varying resource prices. Lecture Notes in Computer Science, 2536:58–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2536/25360058.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2536/25360058.pdf.

Lee:2002:TAL

[3220] Soo-Young Lee and Jun Huang. A theoretical approach to load balancing of a target task in a temporally and spatially heterogeneous Grid computing environment. Lecture Notes in Computer Science, 2536:70–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2536/25360070.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2536/25360070.pdf.

David:2002:SCT

[3221] Romaric David, Stephane Genaud, Arnaud Giersch, Benjamin Schwarz, and Eric Violard. Source code transformations strategies to loadbalance Grid applications. Lecture Notes in Computer Science, 2536: 82–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2536/25360082.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2536/25360082.
pdf.

Perez:2002:PCC

[3222] Christian Pérez, Thierry Priol, and André Ribes. A parallel CORBA component model for numerical code coupling. Lecture Notes in Computer Science, 2536:88–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2536/25360088.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2536/25360088.pdf.

Mayer:2002:MBG

[3223] Anthony Mayer, Stephen McGough, Murtaza Gulamali, Laurie Young, Jim Stanton, Steven Newhouse, and John Darlington. Meaning and behaviour in Grid oriented components. Lecture Notes in Computer Science, 2536:100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2536/25360100.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2536/25360100.pdf.

Chang:2002:TGC

[3224] Bor-Yuh Evan Chang, Karl Crary, Margaret DeLap, Robert Harper, Jason Liszka, Tom Murphy VII, and Frank Pfenning. Trustless Grid computing in ConCert. Lecture Notes in Computer Science, 2536: 112–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2536/25360112.htm;   http:
                                        //link.springer.de/link/service/

series/0558/papers/2536/25360112.
pdf.

Hoschek:2002:UPP

[3225] Wolfgang Hoschek. A unified peerto-peer database framework for scalable service and resource discovery. Lecture Notes in Computer Science, 2536:126–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2536/25360126.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2536/25360126.pdf.

Li:2002:GRD

[3226] Wei Li, Zhiwei Xu, Fangpeng Dong, and Jun Zhang. Grid resource discovery based on a routing-transferring model. Lecture Notes in Computer Science, 2536:145–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2536/25360145.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2536/25360145.pdf.

Faerman:2002:RAS

[3227] Marcio Faerman, Adam Birnbaum, Henri Casanova, and Fran Berman. Resource allocation for steerable parallel parameter searches. Lecture Notes in Computer Science, 2536: 157–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2536/25360157.htm;   http:
//link.springer.de/link/service/

                                        series/0558/papers/2536/25360157.
                                        pdf.

Ramakrishnan:2002:AFG

[3228] Lavanya Ramakrishnan, Helen Rehn, Jay Alameda, Rachana Ananthakrishnan, Madhusudhan Govindaraju, Aleksander Slominski, Kay Connelly, Von Welch, Dennis Gannon, Randall Bramley, and Shawn Hampton. An authorization framework for a Grid based component architecture. Lecture Notes in Computer Science, 2536:169–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2536/25360169.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2536/25360169.pdf.

Lorch:2002:SSA

[3229] Markus Lorch and Dennis Kafura. Supporting secure ad-hoc user collaboration in Grid environments. Lecture Notes in Computer Science, 2536:181–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2536/25360181.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2536/25360181.pdf.

Sundaram:2002:XBP

[3230] Babu Sundaram and Barbara M. Chapman. XML-based policy engine framework for usage policy management in Grids. Lecture Notes in Computer Science, 2536:194–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2536/25360194.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2536/25360194.pdf.

Keahey:2002:FGA

[3231] K. Keahey and V. Welch. Finegrain authorization for resource management in the Grid environment. Lecture Notes in Computer Science, 2536:199–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2536/25360199.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2536/25360199.pdf.

Lee:2002:ARS

[3232] Byoung-Dai Lee and Jon B. Weissman. Adaptive resource scheduling for network services. Lecture Notes in Computer Science, 2536: 207–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2536/25360207.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2536/25360207.
pdf.

Ernemann:2002:EAM

[3233] Carsten Ernemann, Volker Hamscher, Achim Streit, and Ramin Yahyapour. Enhanced algorithms for multi-site scheduling. Lecture Notes in Computer Science, 2536:219–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2536/25360219.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2536/25360219.pdf.

YarKhan:2002:ESU

[3234] Asim YarKhan and Jack J. Dongarra. Experiments with scheduling using simulated annealing in a Grid environment. Lecture Notes in Computer Science, 2536:232–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2536/25360232.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2536/25360232.pdf.

Verma:2002:PSG

[3235] Dinesh Verma, Sambit Sahu, Seraphin Calo, Manid Beigi, and Isabella Chang. A policy service for GRID computing. Lecture Notes in Computer Science, 2536:243–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2536/25360243.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2536/25360243.pdf.

Silberstein:2002:DDN

[3236] Mark Silberstein, Michael Factor, and Dean Lorenz. DYNAMO — DirectorY, Net Archiver and MOver. Lecture Notes in Computer Science, 2536:256–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2536/25360256.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2536/25360256.pdf.

Baker:2002:GRM

[3237] Mark Baker and Garry Smith.

GridRM: a resource monitoring architecture for the Grid. Lecture Notes in Computer Science, 2536: 268–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2536/25360268.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2536/25360268.
pdf.

Seymour:2002:OGR

[3238] Keith Seymour, Hidemoto Nakada, Satoshi Matsuoka, Jack Dongarra, Craig Lee, and Henri Casanova. Overview of GridRPC: a remote procedure call API for Grid computing. Lecture Notes in Computer Science, 2536:274–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2536/25360274.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2536/25360274.pdf.

Smith:2002:DQP

[3239] Jim Smith, Anastasios Gounaris, Paul Watson, Norman W. Paton, Alvaro A. A. Fernandes, and Rizos Sakellariou. Distributed query processing on the Grid. Lecture Notes in Computer Science, 2536:279–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2536/25360279.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2536/25360279.pdf.

Vazhkudai:2002:UDT

[3240] Sudharshan Vazhkudai and Jennifer M. Schopf. Using disk throughput data in predictions of end-to-end Grid data transfers. Lecture Notes in Computer Science, 2536:291–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2536/25360291.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2536/25360291.pdf.

Nallipogu:2002:ITR

[3241] Elsie Nallipogu, Füsun Özgüner, and Mario Lauria. Improving the throughput of remote storage access through pipelining. Lecture Notes in Computer Science, 2536:305–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2536/25360305.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2536/25360305.pdf.

Anonymous:2002:AIcx

[3242] Anonymous. Author index. Lecture Notes in Computer Science, 2536: 317–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2536/2536auth.pdf.

Streit:2002:STJ

[3243] Achim Streit. A self-tuning job scheduler family with dynamic policy switching. Lecture Notes in Computer Science, 2537:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2537/25370001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2537/25370001.pdf.

Snell:2002:PBB

[3244] Quinn O. Snell, Mark J. Clement, and David B. Jackson. Preemption based backfill. Lecture Notes in Computer Science, 2537:24–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2537/25370024.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2537/25370024.pdf.

Krevat:2002:JSB

[3245] Elie Krevat, José G. Castaños, and José E. Moreira. Job scheduling for the BlueGene/L system. Lecture Notes in Computer Science, 2537:38–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2537/25370038.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2537/25370038.pdf.

Srinivasan:2002:SRS

[3246] Srividya Srinivasan, Rajkumar Kettimuthu, Vijay Subramani, and Ponnuswamy Sadayappan. Selective reservation strategies for backfill job scheduling. Lecture Notes in Computer Science, 2537:55–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2537/25370055.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2537/25370055.pdf.

Lawson:2002:MQB

[3247] Barry G. Lawson and Evgenia Smirni. Multiple-queue backfilling scheduling with priorities and reservations for parallel systems. Lecture Notes in Computer Science, 2537:72–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2537/25370072.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2537/25370072.pdf.

Ward:2002:SJP

[3248] William A. Ward, Jr., Carrie L. Mahood, and John E. West. Scheduling jobs on parallel systems using a relaxed backfill strategy. Lecture Notes in Computer Science, 2537: 88–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2537/25370088.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2537/25370088.
pdf.

Chiang:2002:IMA

[3249] Su-Hui Chiang, Andrea Arpaci-Dusseau, and Mary K. Vernon. The impact of more accurate requested runtimes on production job scheduling performance. Lecture Notes in Computer Science, 2537:103–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2537/25370103.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2537/25370103.pdf.

Ernemann:2002:ESG

[3250] Carsten Ernemann, Volker Hamscher, and Ramin Yahyapour. Economic scheduling in Grid computing. Lecture Notes in Computer Science, 2537:128–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2537/25370128.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2537/25370128.pdf.

Czajkowski:2002:SPN

[3251] Karl Czajkowski, Ian Foster, Carl Kesselman, Volker Sander, and Steven Tuecke. SNAP: a protocol for negotiating service level agreements and coordinating resource management in distributed systems. Lecture Notes in Computer Science, 2537:153–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2537/25370153.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2537/25370153.pdf.

Bucur:2002:LVG

[3252] Anca I. D. Bucur and Dick H. J. Epema. Local versus global schedulers with processor co-allocation in multicluster systems. Lecture Notes in Computer Science, 2537: 184–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2537/25370184.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2537/25370184.
                                        pdf.

Pinchak:2002:PHP

[3253] Christopher Pinchak, Paul Lu, and Mark Goldenberg. Practical heterogeneous placeholder scheduling in overlay metacomputers: Early experiences. Lecture Notes in Computer Science, 2537:205–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2537/25370205.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2537/25370205.pdf.

Nitzberg:2002:CAS

[3254] Bill Nitzberg and Jennifer M. Schopf. Current activities in the scheduling and resource management area of the Global Grid Forum. Lecture Notes in Computer Science, 2537: 229–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/

0558/bibs/2537/25370229.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2537/25370229.
pdf.

Anonymous:2002:AIcy

[3255] Anonymous. Author index. Lecture Notes in Computer Science, 2537: 237–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2537/2537ind.pdf.

Konig-Ries:2002:RDD

[3256] Birgitta König-Ries, Kia Makki, S. A. M. Makki, Charles E. Perkins, Niki Pissinou, Peter Reiher, Peter Scheuermann, Jari Veijalainen, Alexander Wolf, and Ouri Wolfson. Research direction for developing an infrastructure for mobile & wireless systems: Consensus report of the NSF workshop held on October 15, 2001 in Scottsdale, Arizona. Lecture Notes in Computer Science, 2538:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2538/25380001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2538/25380001.pdf.

Perkins:2002:RDM

[3257] Charles E. Perkins. Recent developments with mobile IPv6. Lecture Notes in Computer Science, 2538: 38–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/

                                        0558/bibs/2538/25380038.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2538/25380038.
                                        pdf.

Kumar:2002:AHW

[3258] P. R. Kumar. Ad hoc wireless networks: Analysis, protocols, architecture, and convergence. Lecture Notes in Computer Science, 2538:39–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2538/25380039.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2538/25380039.pdf.

Vingralek:2002:SCW

[3259] Radek Vingralek. Supporting Ecommerce in wireless networks. Lecture Notes in Computer Science, 2538: 40–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2538/25380040.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2538/25380040.
                                        pdf.

Wolfson:2002:MMO

[3260] Ouri Wolfson, Sam Chamberlain, Kostas Kalpakis, and Yelena Yesha. Modeling moving objects for location based services. Lecture Notes in Computer Science, 2538: 46–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2538/25380046.htm;   http:

//link.springer.de/link/service/
series/0558/papers/2538/25380046.
pdf.

Carzaniga:2002:CBN

[3261] Antonio Carzaniga and Alexander L.

Wolf. Content-based networking: a new communication infrastructure. Lecture Notes in Computer Science, 2538:59–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2538/25380059.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2538/25380059.pdf.

Kumar:2002:FIC

[3262] Mohan Kumar and Sajal Das. Flexible integrated cache for efficient information access in mobile computing environments. Lecture Notes in Computer Science, 2538:69–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2538/25380069.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2538/25380069.pdf.

Yee:2002:DCM

[3263] Wai Gen Yee and Shamkant B. Navathe. Design considerations for mobile client-server database applications. Lecture Notes in Computer Science, 2538:78–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2538/25380078.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2538/25380078.pdf.

Weissman:2002:SEM

[3264] Jon Weissman and Zhi-Li Zhang. Smart environments: Middleware building blocks for pervasive network computing (A position paper). Lecture Notes in Computer Science, 2538:90–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2538/25380090.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2538/25380090.pdf.

Joshi:2002:AMM

[3265] Anupam Joshi, Timothy Finin, and Yelena Yesha. Agents, mobility, and M-services: Creating the next generation applications and infrastructure on mobile ad-hoc networks. Lecture Notes in Computer Science, 2538:106–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2538/25380106.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2538/25380106.pdf.

Goni:2002:AMS

[3266] A. Goñi, A. Illarramendi, E. Mena, Y. Villate, and J. Rodríguez. ANTARC-TICA: a multiagent system for Internet data services in a wireless computing framework. Lecture Notes in Computer Science, 2538:119–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2538/25380119.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2538/25380119.pdf.

Wegner:2002:CIM

[3267] Lutz Wegner, Morad Ahmad, Stefan Fröhlich, Christian Schmidt, and Wilfried Evers. A collaborative infrastructure for mobile and wireless systems. Lecture Notes in Computer Science, 2538:136–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2538/25380136.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2538/25380136.pdf.

Davies:2002:RIS

[3268] Nigel Davies, Keith Cheverst, Christos Efstratiou, and Adrian Friday. The rationale for infrastructure support for adaptive and contextaware applications: a position paper. Lecture Notes in Computer Science, 2538:146–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2538/25380146.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2538/25380146.pdf.

Veijalainen:2002:MWI

[3269] Jari Veijalainen and Tom Gross. Mobile wireless interfaces: In search for the limits. Lecture Notes in Computer Science, 2538:153–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2538/25380153.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2538/25380153.pdf.

Hatefi:2002:PSR

[3270] Farid Hatefi, Forouzan Golshani, and Gamze Seckin. Prospects of secure realtime video transmission over CDPD networks. Lecture Notes in Computer Science, 2538:164–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2538/25380164.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2538/25380164.pdf.

Lee:2002:MLB

[3271] Yugyung Lee, Nitin Prabhu, and E. K. Park. Middleware for location based information services in mobile environment. Lecture Notes in Computer Science, 2538:173–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2538/25380173.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2538/25380173.pdf.

Anonymous:2002:AIcz

[3272] Anonymous. Author index. Lecture Notes in Computer Science, 2538: 183–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2538/2538indx.pdf.

Borner:2002:VID

[3273] Katy Börner and Chaomei Chen. Visual interfaces to digital libraries: Motivation, utilization, and sociotechnical challenges. Lecture Notes in Computer Science, 2539:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2539/25390001.htm;   http://link.
springer.de/link/service/series/
0558/papers/2539/25390001.pdf.

Buchanan:2002:SHR

[3274] George Buchanan, Ann Blandford, Matt Jones, and Harold Thimbleby. Spatial hypertext as a reader tool in digital libraries. Lecture Notes in Computer Science, 2539: 13–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2539/25390013.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2539/25390013.
pdf.

Christoffel:2002:ALE

[3275] Michael Christoffel and Bethina Schmitt. Accessing libraries as easy as a game. Lecture Notes in Computer Science, 2539:25–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2539/25390025.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2539/25390025-1.pdf;   http://link.
springer.de/link/service/series/
0558/papers/2539/25390025-2.pdf;

                                         http://link.springer.de/link/service/
                                        series/0558/papers/2539/25390025.
                                        pdf.

Monroy:2002:ITV

[3276] Carlos Monroy, Rajiv Kochumman, Richard Furuta, and Eduardo Urbina. Interactive Timeline Viewer (ItLv): a tool to visualize variants among documents. Lecture Notes in Computer Science, 2539:39–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2539/25390039.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2539/25390039.pdf.

Weiss-Lijn:2002:EEI

[3277] Mischa Weiss-Lijn, Janet T. McDonnell, and Leslie James. An empirical evaluation of the interactive visualization of metadata to support document use. Lecture Notes in Computer Science, 2539:50–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2539/25390050.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2539/25390050.pdf.

Eick:2002:VAW

[3278] Stephen G. Eick. Visual analysis of Website browsing patterns. Lecture Notes in Computer Science, 2539:65–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2539/25390065.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2539/25390065.pdf.

Graham:2002:ETP

[3279] Adrian Graham, Hector Garcia-Molina, Andreas Paepcke, and Terry Winograd. Extreme temporal photo browsing. Lecture Notes in Computer Science, 2539:81–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2539/25390081.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2539/25390081.pdf.

Christel:2002:ANV

[3280] Michael G. Christel. Accessing news video libraries through dynamic information extraction, summarization, and visualization. Lecture Notes in Computer Science, 2539: 98–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2539/25390098.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2539/25390098.
pdf.

Spoerri:2002:HNV

[3281] Anselm Spoerri. Handwritten notes as a visual interface to index, edit and publish audio/video highlights. Lecture Notes in Computer Science, 2539:116–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2539/25390116.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2539/25390116.pdf.

Buzydlowski:2002:TCO

[3282] Jan W. Buzydlowski, Howard D. White, and Xia Lin. Term cooccurrence analysis as an interface for digital libraries. Lecture Notes in Computer Science, 2539: 133–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2539/25390133.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2539/25390133.
                                        pdf.

Boyack:2002:IVH

[3283] Kevin W. Boyack, Brian N. Wylie, and George S. Davidson. Information visualization, human-computer interaction, and cognitive psychology: Domain visualizations. Lecture Notes in Computer Science, 2539: 145–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2539/25390145.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2539/25390145.
                                        pdf.

Skupin:2002:GTM

[3284] André Skupin. On geometry and transformation in map-like information visualization. Lecture Notes in Computer Science, 2539:161–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2539/25390161.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2539/25390161.pdf.

Cai:2002:GVI

[3285] Guoray Cai. GeoVIBE: a visual interface for geographic digital libraries. Lecture Notes in Computer Science, 2539:171–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2539/25390171.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2539/25390171.pdf.

Ong:2002:IIV

[3286] Teong Joo Ong, John J. Leggett, Hugh D. Wilson, Stephan L. Hatch, and Monique D. Reed. Interactive information visualization in the digital flora of texas. Lecture Notes in Computer Science, 2539: 188–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2539/25390188.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2539/25390188.
pdf.

Ancona:2002:VEA

[3287] Dan Ancona, Mike Freeston, Terry Smith, and Sara Fabrikant. Visual explorations for the Alexandria Digital Earth Prototype. Lecture Notes in Computer Science, 2539: 199–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2539/25390199.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2539/25390199-
                                        1.pdf;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2539/25390199-2.pdf;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2539/25390199.pdf.

Shen:2002:LPB

[3288] Rao Shen, Jun Wang, and Edward A. Fox. A lightweight protocol between digital libraries and visualization systems. Lecture Notes in Computer Science, 2539:217–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2539/25390217.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2539/25390217.pdf.

Chen:2002:TTP

[3289] Chaomei Chen and Katy Börner. Top ten problems in visual interfaces to digital libraries. Lecture Notes in Computer Science, 2539: 226–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2539/25390226.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2539/25390226.
                                        pdf.

Anonymous:2002:AIda

[3290] Anonymous. Author index. Lecture Notes in Computer Science, 2539: 233–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2539/2539auth.pdf.

Fouchal:2002:CTT

[3291] Hacène Fouchal. Conformance testing techniques for timed systems. Lecture Notes in Computer Science, 2540:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2540/25400001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2540/25400001.pdf.

Reussner:2002:CCF

[3292] Ralf Reussner. Counter-constrained finite state machines: a new model for component protocols with resourcedependencies. Lecture Notes in Computer Science, 2540:20–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/link/service/
series/0558/bibs/2540/25400020.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2540/25400020.pdf.

Kucera:2002:ECI

[3293] Antonín Kucera and Petr Jancar.

Equivalence-checking with infinitestate systems: Techniques and results. Lecture Notes in Computer Science, 2540:41–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2540/25400041.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2540/25400041.pdf.

Kamareddine:2002:FTT

[3294] Fairouz Kamareddine. On functions and types: a tutorial. Lecture Notes in Computer Science, 2540: 74–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2540/25400074.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2540/25400074.
                                        pdf.

Rhodes:2002:DSM

[3295] Philip J. Rhodes, R. Daniel Bergeron, and Ted M. Sparr. Database support for multisource multiresolution scientific data. Lecture Notes in Computer Science, 2540:94–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2540/25400094.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2540/25400094.pdf.

Assfalg:2002:SAI

[3296] Jürgen Assfalg, Marco Bertini, Carlo Colombo, Alberto Del Bimbo, and Walter Nunziati. Semantic annotation and indexing of news and sports videos. Lecture Notes in Computer Science, 2540:115–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2540/25400115.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2540/25400115.pdf.

Sapino:2002:MPD

[3297] Maria Luisa Sapino. Multimedia presentations databases (extended abstract). Lecture Notes in Computer Science, 2540:132–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2540/25400132.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2540/25400132.pdf.

Worring:2002:IIR

[3298] Marcel Worring, Andrew Bagdanov, Jan v. Gemert, Jan-Mark Geusebroek, Hoang Minh, Guus Schreiber, Cees Snoek, Jeroen Vendrig, Jan Wielemaker, and Arnold Smeulders. Interactive indexing and retrieval of multimedia content. Lecture Notes in Computer Science, 2540:135–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2540/25400135.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2540/25400135.pdf.

Golshani:2002:MBA

[3299] Forouzan Golshani, Youngchoon Park, and Sethuraman Panchanathan. A model-based approach to semanticbased retrieval of visual information. Lecture Notes in Computer Science, 2540:149–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2540/25400149.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2540/25400149.pdf.

Dubois:2002:BPL

[3300] Didier Dubois and Henri Prade. Bipolarity in possibilistic logic and fuzzy rules (extended abstract). Lecture Notes in Computer Science, 2540:168–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2540/25400168.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2540/25400168.pdf.

Jeffery:2002:DMC

[3301] Keith G. Jeffery. Data management challenges for GRID computing. Lecture Notes in Computer Science, 2540:174–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2540/25400174.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2540/25400174.pdf.

Kovacs:2002:HSS

[3302] Kornél Kovács, András Kocsor, and László Tóth. Hungarian speech synthesis using a phase exact HNM approach. Lecture Notes in Computer Science, 2540:181–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2540/25400181.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2540/25400181.pdf.

Bartak:2002:MRT

[3303] Roman Barták. Modelling resource transitions in constraint-based scheduling. Lecture Notes in Computer Science, 2540:186–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2540/25400186.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2540/25400186.pdf.

Subramani:2002:SFR

[3304] K. Subramani. A specification framework for real-time scheduling. Lecture Notes in Computer Science, 2540:195–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2540/25400195.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2540/25400195.pdf.

Kutrib:2002:STD

[3305] Martin Kutrib and Jan-Thomas Löwe. String transformation for n-dimensional image compression. Lecture Notes in Computer Science, 2540:208–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2540/25400208.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2540/25400208.pdf.

Iliopoulos:2002:VDP

[3306] Costas S. Iliopoulos and Manal Mohamed. Validation and decomposition of partially occluded images. Lecture Notes in Computer Science, 2540:218–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2540/25400218.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2540/25400218.pdf.

Katarzyniak:2002:SCA

[3307] Radoslaw Piotr Katarzyniak and Ngoc Thanh Nguyen. Solving conflicts of agent knowledge states in multiagent systems. Lecture Notes in Computer Science, 2540:231–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2540/25400231.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2540/25400231.pdf.

Alagar:2002:SVS

[3308] Vangalur S. Alagar and Kasi Periyasamy. Specification and verification of secure business transaction systems. Lecture Notes in Computer Science, 2540:240–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2540/25400240.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2540/25400240.pdf.

Kisiel-Dorohinicki:2002:AOM

[3309] Marek Kisiel-Dorohinicki. Agentoriented model of simulated evolution. Lecture Notes in Computer Science, 2540:253–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2540/25400253.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2540/25400253.pdf.

Gebala:2002:RSC

[3310] Maciej Gebala. The reconstruction of some 3D convex polyominoes from orthogonal projections. Lecture Notes in Computer Science, 2540:262–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2540/25400262.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2540/25400262.pdf.

Midrijanis:2002:CPV

[3311] Gatis Midrijanis. The complexity of probabilistic versus quantum finite automata. Lecture Notes in Computer Science, 2540:273–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2540/25400273.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2540/25400273.pdf.

Ultes-Nitsche:2002:DWN

[3312] Ulrich Ultes-Nitsche. Do we need liveness? — approximation of liveness properties by safety properties. Lecture Notes in Computer Science, 2540:279–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2540/25400279.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2540/25400279.pdf.

Anonymous:2002:AIdb

[3313] Anonymous. Author index. Lecture Notes in Computer Science, 2540: 289–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2540/2540ind.pdf.

Barkowsky:2002:CI

[3314] Thomas Barkowsky. Chapter 1: Introduction. Lecture Notes in Computer Science, 2541:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2541/25410001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2541/25410001.pdf.

Barkowsky:2002:CSA

[3315] Thomas Barkowsky. Chapter 2: State of the art. Lecture Notes in Computer Science, 2541:21–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2541/25410021.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2541/25410021.pdf.

Barkowsky:2002:CMD

[3316] Thomas Barkowsky. Chapter 3:

MIRAGE — developing the model. Lecture Notes in Computer Science, 2541:65–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2541/25410065.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2541/25410065.pdf.

Barkowsky:2002:CVM

[3317] Thomas Barkowsky. Chapter 4: Visual mental image construction in detail. Lecture Notes in Computer Science, 2541:95–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2541/25410095.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2541/25410095.pdf.

Barkowsky:2002:CMI

[3318] Thomas Barkowsky. Chapter 5:

MIRAGE implementation. Lecture Notes in Computer Science, 2541: 117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2541/25410117.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2541/25410117.
pdf.

Barkowsky:2002:CCO

[3319] Thomas Barkowsky. Chapter 6:

Conclusion and outlook. Lecture Notes in Computer Science, 2541: 139–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2541/25410139.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2541/25410139.
pdf.

Anonymous:2002:B

[3320] Anonymous. Bibliography. Lecture Notes in Computer Science, 2541: 157–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2541/25410001.pdf.

Anonymous:2002:Ib

[3321] Anonymous.  Index.  Lecture

Notes in Computer Science, 2541: 169–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2541/25410001.pdf.

Shalaby:2002:SSN

[3322] Nadia Shalaby, Yitzchak Gottlieb, Mike Wawrzoniak, and Larry Peterson. Snow on silk: a NodeOS in the Linux kernel. Lecture Notes in Computer Science, 2546:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2546/25460001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2546/25460001.pdf.

Keller:2002:PDE

[3323] Ralph Keller, Lukas Ruf, Amir Guindehi, and Bernhard Plattner. PromethOS: a dynamically extensible router architecture supporting explicit routing. Lecture Notes in Computer Science, 2546:20–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2546/25460020.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2546/25460020.pdf.

Bos:2002:OCC

[3324] Herbert Bos and Bart Samwel. The OKE corral: Code organisation and reconfiguration at runtime using active linking. Lecture Notes in Computer Science, 2546:32–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2546/25460032.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2546/25460032.pdf.

Donnelly:2002:LTT

[3325] Austin Donnelly. Lightweight thread tunnelling in network applications. Lecture Notes in Computer Science, 2546:48–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2546/25460048.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2546/25460048.pdf.

Martin:2002:RRB

[3326] Sylvain Martin and Guy Leduc. RADAR: Ring-based adaptive discovery of active neighbour routers. Lecture Notes in Computer Science, 2546:62–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2546/25460062.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2546/25460062.pdf.

Bossardt:2002:ISD

[3327] Matthias Bossardt, Takashi Egawa, Hideki Otsuki, and Bernhard Plattner. Integrated service deployment for active networks. Lecture Notes in Computer Science, 2546: 74–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2546/25460074.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2546/25460074.
pdf.

Solarski:2002:CBD

[3328] Marcin Solarski, Matthias Bossardt, and Thomas Becker. Componentbased deployment and management of services in active networks. Lecture Notes in Computer Science, 2546: 87–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2546/25460087.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2546/25460087.
pdf.

Rogers:2002:AAN

[3329] Craig Milo Rogers. ANQL — an active networks query language. Lecture Notes in Computer Science, 2546:99–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2546/25460099.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2546/25460099.pdf.

Moore:2002:PLM

[3330] Jonathan T. Moore, Jessica Kornblum Moore, and Scott Nettles. Predictable, lightweight management agents. Lecture Notes in Computer Science, 2546:111–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2546/25460111.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2546/25460111.pdf.

Anagnostakis:2002:OPM

[3331] Kostas G. Anagnostakis, Michael Greenwald, Sotiris Ioannidis, and Stefan Miltchev. Open packet monitoring on FLAME: Safety, performance, and applications. Lecture Notes in Computer Science, 2546: 120–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2546/25460120.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2546/25460120.
pdf.

Prehofer:2002:ANM

[3332] Christian Prehofer and Qing Wei. Active networks for 4G mobile communication: Motivation, architecture, and application scenarios. Lecture Notes in Computer Science, 2546:132–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2546/25460132.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2546/25460132.pdf.

Song:2002:EAU

[3333] Seong-Kyu Song, Stephen Shannon, Michael Hicks, and Scott Nettles. Evolution in action: Using active networking to evolve network support for mobility. Lecture Notes in Computer Science, 2546:146–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2546/25460146.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2546/25460146.pdf.

Fuhrmann:2002:AIA

[3334] Thomas Fuhrmann, Till Harbaum, Marcus Schöller, and Martina Zitterbart. AMnet 2.0: An improved architecture for programmable networks. Lecture Notes in Computer Science, 2546:162–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2546/25460162.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2546/25460162.pdf.

Baumgartner:2002:DIP

[3335] Florian Baumgartner, Torsten Braun, and Bharat Bhargava. Design and implementation of a pythonbased active network platform for network management and control. Lecture Notes in Computer Science, 2546:177–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2546/25460177.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2546/25460177.pdf.

Bond:2002:DSS

[3336] Mary Bond, James Griffioen, Chetan Singh Dhillon, and Kenneth L. Calvert. Designing service-specific execution environments. Lecture Notes in Computer Science, 2546:191–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2546/25460191.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2546/25460191.pdf.

Bagnulo:2002:RRO

[3337] Marcelo Bagnulo, Bernardo Alarcos, María Calderón, and Marifeli Sedano. ROSA: Realistic open security architecture for active networks. Lecture Notes in Computer Science, 2546:204–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/

                                        series/0558/bibs/2546/25460204.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2546/25460204.pdf.

Sehgal:2002:FCB

[3338] Amit Sehgal, Kenneth L. Calvert, and James Griffioen. A flexible concast-based grouping service. Lecture Notes in Computer Science, 2546:216–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2546/25460216.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2546/25460216.pdf.

Smith:2002:PRD

[3339] Paul Smith, Steven Simpson, and David Hutchison. Programmable resource discovery using peer-to-peer networks. Lecture Notes in Computer Science, 2546:229–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2546/25460229.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2546/25460229.pdf.

Ogose:2002:FID

[3340] Go Ogose, Jyunya Yoshida, Tae Yoneda, and Tadashi Ohta. Feature interaction detection in active networks. Lecture Notes in Computer Science, 2546:241–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2546/25460241.

htm;   http://link.springer.de/
link/service/series/0558/papers/
2546/25460241.pdf.

Schmid:2002:FDS

[3341] Stefan Schmid, Tim Chart, Manolis Sifalakis, and Andrew C. Scott. Flexible, dynamic, and scalable service composition for active routers. Lecture Notes in Computer Science, 2546:253–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2546/25460253.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2546/25460253.pdf.

Anonymous:2002:AIdc

[3342] Anonymous. Author index. Lecture Notes in Computer Science, 2546: 267–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2546/2546indx.pdf.

Czajkowski:2002:RMS

[3343] Grzegorz Czajkowski and Jan Vitek.

Resource management for safe languages. Lecture Notes in Computer Science, 2548:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2548/25480001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2548/25480001.pdf.

Czarnecki:2002:GP

[3344] Krzysztof Czarnecki, Kasper Østerbye, and Markus Völter. Generative programming. Lecture Notes in Computer Science, 2548:15–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2548/25480015.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2548/25480015.pdf.

Michiels:2002:TEL

[3345] Isabel Michiels, Jürgen Börstler, and Kim B. Bruce. Tools and environments for learning object-oriented concepts. Lecture Notes in Computer Science, 2548:30–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2548/25480030.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2548/25480030.pdf.

Perez:2002:WPS

[3346] Miguel A. Pérez and Pedro J. Clemente. 12th Workshop for PhD Students in Object Oriented Systems. Lecture Notes in Computer Science, 2548:44–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2548/25480044.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2548/25480044.pdf.

Pastor:2002:WOS

[3347] Oscar Pastor, Daniel Schwabe, Gustavo Rossi, and Luis Olsina. Weboriented software technology. Lecture Notes in Computer Science, 2548: 55–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2548/25480055.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2548/25480055.
pdf.

Bosch:2002:COP

[3348] Jan Bosch, Clemens Szyperski, and Wolfgang Weck. Component-oriented programming. Lecture Notes in Computer Science, 2548:70–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/link/service/
series/0558/bibs/2548/25480070.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2548/25480070.pdf.

Beugnard:2002:CCA

[3349] Antoine Beugnard, Salah Sadou, Laurence Duchien, and Eric Jul. Concrete communication abstractions of the next 701 distributed object systems. Lecture Notes in Computer Science, 2548:79–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2548/25480079.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2548/25480079.pdf.

Kniesel:2002:USE

[3350] Günter Kniesel, Joost Noppen, Tom Mens, and Jim Buckley. Unanticipated software evolution. Lecture Notes in Computer Science, 2548: 92–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2548/25480092.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2548/25480092.
pdf.

Lumpe:2002:CL

[3351] Markus Lumpe, Jean-Guy Schneider, Bastiaan Schönhage, and Thomas Genssler. Composition languages. Lecture Notes in Computer Science, 2548:107–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2548/25480107.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2548/25480107.pdf.

Arevalo:2002:IW

[3352] Gabriela Arévalo, Andrew Black, Yania Crespo, Michel Dao, Erik Ernst, Peter Grogono, Marianne Huchard, and Markku Sakkinen. The inheritance workshop. Lecture Notes in Computer Science, 2548:117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2548/25480117.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2548/25480117.pdf.

Speck:2002:MBS

[3353] Andreas Speck, Elke Pulvermüller, Ragnhild Van Der Straeten, Ralf H. Reussner, and Matthias Clauß. Modelbased software reuse. Lecture Notes in Computer Science, 2548: 135–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2548/25480135.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2548/25480135.
pdf.

Piattini:2002:QAO

[3354] Mario Piattini, Fernando Brito e Abreu, Geert Poels, and Houari A. Sahraoui. Quantitative approaches in object-oriented software engineering. Lecture Notes in Computer Science, 2548:147–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2548/25480147.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2548/25480147.pdf.

Davis:2002:MPO

[3355] Kei Davis, Yannis Smaragdakis, and Jörg Striegnitz. Multiparadigm programming with object-oriented languages. Lecture Notes in Computer Science, 2548:154–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2548/25480154.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2548/25480154.pdf.

DHondt:2002:KBO

[3356] Maja D’Hondt, Kim Mens, and Ellen Van Paesschen. Knowledgebased object-oriented software engineering. Lecture Notes in Computer Science, 2548:160–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2548/25480160.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2548/25480160.pdf.

Gal:2002:OOO

[3357] Andreas Gal, Olaf Spinczyk, and Dario Alvarez. Object-orientation and operating systems. Lecture Notes in Computer Science, 2548: 174–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2548/25480174.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2548/25480174.
                                        pdf.

Araujo:2002:ITU

[3358] João Araújo, Jonathan Whittle, Ambrosio Toval, and Robert France. Integration and transformation of UML models. Lecture Notes in Computer Science, 2548:184–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2548/25480184.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2548/25480184.pdf.

Bryce:2002:MOS

[3359] Ciarán Bryce. Mobile object systems. Lecture Notes in Computer Science, 2548:192–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2548/25480192.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2548/25480192.pdf.

Meuter:2002:FRC

[3360] Wolfgang De Meuter, Pascal Costanza, Martine Devos, and Dave Thomas. Feyerabend: Redefining computing. Lecture Notes in Computer Science, 2548:197–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2548/25480197.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2548/25480197.pdf.

Drossopoulou:2002:FTJ

[3361] Sophia Drossopoulou, Susan Eisenbach, Gary T. Leavens, Arnd Poetzsch-Heffter, and Erik Poll. Formal techniques for Java-like programs. Lecture Notes in Computer Science, 2548:203–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2548/25480203.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2548/25480203.pdf.

Murillo:2002:PS

[3362] Juan Manuel Murillo and Fernando Sánchez. Poster session. Lecture Notes in Computer Science, 2548:211–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2548/25480211.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2548/25480211.pdf.

Anonymous:2002:AIdd

[3363] Anonymous. Author index. Lecture Notes in Computer Science, 2548: 223–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2548/2548indx.pdf.

Benko:2002:CS

[3364] Igor Benko and Jo Ebergen. Composing snippets. Lecture Notes in Computer Science, 2549:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer.
                                        de/link/service/series/0558/bibs/
                                        2549/25490001.htm;   http://link.
                                        springer.de/link/service/series/
                                        0558/papers/2549/25490001.pdf.

Josephs:2002:PAD

[3365] Mark B. Josephs and Dennis P. Furey. A programming approach to the design of asynchronous logic blocks. Lecture Notes in Computer Science, 2549:34–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2549/25490034.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2549/25490034.pdf.

Varshavsky:2002:GGA

[3366] Victor Varshavsky and Vyacheslav Marakhovsky. GALA (globally asynchronous — locally arbitrary) design. Lecture Notes in Computer Science, 2549:61–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2549/25490061.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2549/25490061.pdf.

Carmona:2002:SRS

[3367] Josep Carmona, Jordi Cortadella, and Enric Pastor. Synthesis of reactive systems: Application to asynchronous circuit design. Lecture Notes in Computer Science, 2549:108–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2549/25490108.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2549/25490108.pdf.

Vogler:2002:DAC

[3368] Walter Vogler and Ralf Wollowski. Decomposition in asynchronous circuit design. Lecture Notes in Computer Science, 2549:152–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2549/25490152.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2549/25490152.pdf.

LaRue:2002:FPM

[3369] William LaRue, Sherry Solden, and Bishnupriya Bhattacharya. Functional and performance modeling of concurrency in VCC. Lecture Notes in Computer Science, 2549: 191–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2549/25490191.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2549/25490191.
                                        pdf.

Balarin:2002:MDH

[3370] Felice Balarin, Luciano Lavagno, Claudio Passerone, Alberto Sangiovanni-Vincentelli, Marco Sgroi, and Yosinori Watanabe. Modeling and designing heterogeneous systems. Lecture Notes in Computer Science, 2549:228–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2549/25490228.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2549/25490228.pdf.

Moller:2002:TVA

[3371] Jesper Møller, Henrik Hulgaard, and Henrik Reif Andersen. Timed verification of asynchronous circuits. Lecture Notes in Computer Science, 2549:274–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2549/25490274.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2549/25490274.pdf.

Beerel:2002:PAA

[3372] Peter A. Beerel and Aiguo Xie. Performance analysis of asynchronous circuits using Markov chains. Lecture Notes in Computer Science, 2549: 313–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2549/25490313.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2549/25490313.
pdf.

Anonymous:2002:AIde

[3373] Anonymous. Author index. Lecture Notes in Computer Science, 2549: 345–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2549/2549ind.pdf.

Bassi:2002:LBS

[3374] Alessandro Bassi, Micah Beck, Terry Moore, and James S. Plank. The logistical backbone: Scalable infrastructure for global data Grids. Lecture Notes in Computer Science, 2550:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2550/25500001.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2550/25500001.pdf.

Plateau:2002:GCR

[3375] Brigitte Plateau. The Grid: Challenges and research issues. Lecture Notes in Computer Science, 2550: 13–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2550/25500013.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2550/25500013.
                                        pdf.

Shimojo:2002:PVC

[3376] Shinji Shimojo, Junzo Kamahara, Takahiro Nishizawa, and Hideo Miyahara. Personalized video for contents delivery network. Lecture Notes in Computer Science, 2550: 15–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2550/25500015.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2550/25500015.
                                        pdf.

Jin:2002:DCC

[3377] Hai Jin and Kai Hwang. Distributed checkpointing on clusters with dynamic striping and staggering. Lecture Notes in Computer Science, 2550:19–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2550/25500019.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2550/25500019.pdf.

Thoai:2002:SRR

[3378] Nam Thoai, Dieter Kranzlmüller, and Jens Volkert. Shortcut replay: a replay technique for debugging long-running parallel programs. Lecture Notes in Computer Science, 2550:34–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2550/25500034.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2550/25500034.pdf.

Cogis:2002:LCP

[3379] Olivier Cogis, Jean-Claude König, and Jérôme Palaysi. On the list colouring problem. Lecture Notes in Computer Science, 2550:47–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2550/25500047.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2550/25500047.pdf.

Subramani:2002:DML

[3380] K. Subramani. On determining the minimum length, tree-like resolution refutation of 2SAT, and extended 2SAT formulas. Lecture Notes in Computer Science, 2550:57–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2550/25500057.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2550/25500057.pdf.

Sumari:2002:DSS

[3381] Putra Sumari and Hailiza Kamarulhaili. Data storage and stream caching for video on demand servers. Lecture Notes in Computer Science, 2550:66–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2550/25500066.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2550/25500066.pdf.

Liu:2002:ACM

[3382] Zhen Liu, Mark Squillante, Cathy Xia, S. Yu, Li Zhang, Naceur Malouch, and Paul Dantzig. Analysis of caching mechanisms from sporting event Web sites. Lecture Notes in Computer Science, 2550:76–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2550/25500076.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2550/25500076.pdf.

Ho:2002:SBP

[3383] Dac Phuong Ho, Yong Meng Teo, and Johan Prawira Gozali. Solving the Nbody problem with the ALiCE Grid System. Lecture Notes in Computer Science, 2550:87–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2550/25500087.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2550/25500087.pdf.

Kaishima:2002:EPD

[3384] Takeshi Kaishima, Yuko Mizuno-Matsumoto, Susumu Date, and Shinji Shimojo. The experience and practice of developing a brain functional analysis system using ICA on a Grid environment. Lecture Notes in Computer Science, 2550:98–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2550/25500098.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2550/25500098.pdf.

Combes:2002:SAN

[3385] Philippe Combes, Frédéric Lombard, Martin Quinson, and Frédéric Suter. A scalable approach to network enabled servers. Lecture Notes in Computer Science, 2550:110–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2550/25500110.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2550/25500110.pdf.

Fujikawa:2002:SSM

[3386] Kazutoshi Fujikawa, Masato Hori, Shinji Shimojo, and Hideo Miyahara. A server selection method based on communication delay and communication frequency among users for networked virtual environments. Lecture Notes in Computer Science, 2550:125–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2550/25500125.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2550/25500125.pdf.

Maimour:2002:TAA

[3387] Moufida Maimour and CongDuc Pham. Towards an application-aware multicast communication framework for computational Grids. Lecture Notes in Computer Science, 2550: 140–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2550/25500140.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2550/25500140.
                                        pdf.

Serban:2002:DRA

[3388] Rares Serban, Chadi Barakat, and Walid Dabbous. Dynamic resource allocation in core routers of a Diffserv network. Lecture Notes in Computer Science, 2550:153–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2550/25500153.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2550/25500153.pdf.

Tamura:2002:PAF

[3389] Hitomi Tamura, Kenji Kawahara, and Yuji Oie. Performance analysis of flow loss probability and link utilization in MPLS networks for traffic engineering. Lecture Notes in Computer Science, 2550:168–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2550/25500168.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2550/25500168.pdf.

Tientanopajai:2002:MQR

[3390] Kitt Tientanopajai and Kanchana Kanchanasut. Multiconstraint QoS routing using a path-vector protocol. Lecture Notes in Computer Science, 2550:181–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2550/25500181.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2550/25500181.pdf.

Tonoike:2002:DIW

[3391] Masatsugu Tonoike, Takaaki Komura, Kenji Fujikawa, Yasuo Okabe, and Katsuo Ikeda. Design and implementation of a Web-based Internet broadcasting system on QoS multicast routers. Lecture Notes in Computer Science, 2550:195–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2550/25500195.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2550/25500195.pdf.

Ueno:2002:WBM

[3392] Daiki Ueno, Tatsuo Nakajima, Ichiro Satoh, and Kouta Soejima. Webbased middleware for home entertainment. Lecture Notes in Computer Science, 2550:206–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2550/25500206.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2550/25500206.pdf.

Shyamasundar:2002:MES

[3393] R. K. Shyamasundar and Siddharth B. Deshmukh. MicroBill: An efficient secure system for subscription based services. Lecture Notes in Computer Science, 2550:220–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2550/25500220.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2550/25500220.pdf.

Anonymous:2002:AIdf

[3394] Anonymous. Author index. Lecture Notes in Computer Science, 2550: 233–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2550/2550auth.pdf.

Richards:2002:CLA

[3395] Debbie Richards, Kathrin Boettger, and Oscar Aguilera. A controlled language to assist conversion of use case descriptions into concept lattices. Lecture Notes in Computer Science, 2557:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2557/25570001.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2557/25570001.pdf.

Min:2002:PDC

[3396] Kyongho Min, William H. Wilson, and Yoo-Jin Moon. Preferred document classification for a highly inflectional/derivational language. Lecture Notes in Computer Science, 2557:12–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570012.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570012.pdf.

Zukerman:2002:EQP

[3397] Ingrid Zukerman, Bhavani Raskutti, and Yingying Wen. Experiments in query paraphrasing for information retrieval. Lecture Notes in Computer Science, 2557:24–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570024.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570024.pdf.

Vos:2002:DDM

[3398] Marina De Vos and Dirk Vermeir. Dynamic decision-making in logic programming and game theory. Lecture Notes in Computer Science, 2557:36–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570036.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570036.pdf.

Fusaoka:2002:LRT

[3399] Akira Fusaoka. On a linear representation theory for quantitative belief change. Lecture Notes in Computer Science, 2557:48–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2557/25570048.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2557/25570048.pdf.

Liu:2002:TSC

[3400] Chuchang Liu and Maris A. Ozols. Trust in secure communication systems — the concept, representations, and reasoning techniques. Lecture Notes in Computer Science, 2557:60–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2557/25570060.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2557/25570060.pdf.

Brennan:2002:FFN

[3401] Jane Brennan and Eric Martin. Foundations for a formalism of nearness. Lecture Notes in Computer Science, 2557:71–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2557/25570071.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2557/25570071.pdf.

Choi:2002:SSR

[3402] Seungyeob Choi and Manfred Kerber. Semantic selection for resolution in clause graphs. Lecture Notes in Computer Science, 2557: 83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2557/25570083.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2557/25570083.
pdf.

Dawson:2002:MCT

[3403] Jeremy E. Dawson and Rajeev Goré.

Machine-checking the timed interval calculus. Lecture Notes in Computer Science, 2557:95–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570095.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570095.pdf.

Mayer:2002:MPU

[3404] Wolfgang Mayer and Markus Stumptner. Modeling programs with unstructured control flow for debugging. Lecture Notes in Computer Science, 2557:107–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570107.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570107.pdf.

Kornienko:2002:MLF

[3405] Lara Kornienko, David L. Dowe, and David W. Albrecht. Message length formulation of support vector machines for binary classification — A preliminary scheme. Lecture Notes in Computer Science, 2557:119–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570119.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570119.pdf.

Tan:2002:MID

[3406] Peter J. Tan and David L. Dowe. MML inference of decision graphs with multiway joins. Lecture Notes in Computer Science, 2557:131–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570131.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570131.pdf.

Agusta:2002:MCC

[3407] Yudi Agusta and David L. Dowe. MML clustering of continuous-valued data using Gaussian and t distributions. Lecture Notes in Computer Science, 2557:143–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570143.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570143.pdf.

Kim:2002:OKB

[3408] Sang-Woon Kim and B. John Oommen. Optimizing kernel-based nonlinear subspace methods using prototype reduction schemes. Lecture Notes in Computer Science, 2557:155–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570155.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570155.pdf.

Wobcke:2002:IRP

[3409] Wayne Wobcke. Intention and rationality for PRS-like agents. Lecture Notes in Computer Science, 2557: 167–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2557/25570167.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2557/25570167.
pdf.

Seo:2002:MSD

[3410] Hee Suk Seo and Tae Ho Cho. Modeling and simulation for detecting a distributed denial of service attack. Lecture Notes in Computer Science, 2557:179–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570179.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570179.pdf.

Debenham:2002:KDP

[3411] John Debenham. Knowledge-driven processes can be managed. Lecture Notes in Computer Science, 2557: 191–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2557/25570191.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2557/25570191.
pdf.

Chen:2002:AMA

[3412] Juei-Nan Chen, Yueh-Min Huang, and William C. Chu. Adaptive multi-agent decision making using analytical hierarchy process. Lecture Notes in Computer Science, 2557:203–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570203.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570203.pdf.

Sherwood:2002:APS

[3413] Rob Sherwood, Steve Chien, Rebecca Castano, and Gregg Rabideau. Autonomous planning and scheduling on the TechSat 21 mission. Lecture Notes in Computer Science, 2557:213–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570213.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570213.pdf.

Ashmore:2002:ODR

[3414] Mark Ashmore and Nick Barnes.

Omni-drive robot motion on curved paths: The fastest path between two points is not a straight-line. Lecture Notes in Computer Science, 2557:225–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570225.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570225.pdf.

Gopalan:2002:IID

[3415] Raj P. Gopalan and Grant Lee.

Indexing of image databases using untrained 4D holographic memory model. Lecture Notes in Computer Science, 2557:237–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570237.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570237.pdf.

Chen:2002:FOC

[3416] Zaiping Chen, Qiang Gao, Chao Dong, Hongjin Liu, Peng Zhang, and Zhenlin Xu. The flux-oriented control of an induction machine utilizing an online controller parameter adaptation scheme. Lecture Notes in Computer Science, 2557:249–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570249.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2557/25570249.pdf.

Dowman:2002:MAC

[3417] Mike Dowman. Modelling the acquisition of colour words. Lecture Notes in Computer Science, 2557: 259–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2557/25570259.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2557/25570259.
                                        pdf.

Hope:2002:BIR

[3418] Lucas R. Hope and Kevin B. Korb. Bayesian information reward. Lecture Notes in Computer Science, 2557:272–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2557/25570272.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2557/25570272.pdf.

Jung:2002:PUP

[3419] Kyung-Yong Jung and Jung-Hyun Lee. Prediction of user preference in recommendation system using associative user clustering and Bayesian estimated value. Lecture Notes in Computer Science, 2557:284–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2557/25570284.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2557/25570284.pdf.

George:2002:AIU

[3420] Sarah George and Ingrid Zukerman. Argument interpretation using minimum message length. Lecture Notes in Computer Science, 2557:297–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570297.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570297.pdf.

Loveard:2002:GPC

[3421] Thomas Loveard and Vic Ciesielski.

Genetic programming for classification: An analysis of convergence behaviour. Lecture Notes in Computer Science, 2557:309–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570309.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570309.pdf.

vanRemortel:2002:LID

[3422] Piet van Remortel, Tom Lenaerts, and Bernard Manderick. Lineage and induction in the development of evolved genotypes for non-uniform 2D CAs. Lecture Notes in Computer Science, 2557:321–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570321.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570321.pdf.

Cornforth:2002:EOB

[3423] David Cornforth. Evolution in the orange box — A new approach to the sphere-packing problem in CMAC-based neural networks. Lecture Notes in Computer Science, 2557:333–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570333.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570333.pdf.

Kosoresow:2002:FWC

[3424] Andrew P. Kosoresow and Matthew P. Johnson. Finding worst-case instances of, and lower bounds for, online algorithms using genetic algorithms. Lecture Notes in Computer Science, 2557:344–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570344.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570344.pdf.

Xu:2002:AAF

[3425] Shuxiang Xu and Ming Zhang. An adaptive activation function for higher order neural networks. Lecture Notes in Computer Science, 2557:356–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570356.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570356.pdf.

Sun:2002:ALA

[3426] Jian Sun, Rui-Min Shen, and Fan Yang. An adaptive learning algorithm aimed at improving RBF network generalization ability. Lecture Notes in Computer Science, 2557: 363–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2557/25570363.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2557/25570363.
pdf.

Chen:2002:NNO

[3427] Zaiping Chen, Jun Li, Youjun Yue, Qiang Gao, Hui Zhao, and Zhenlin Xu. A neural network online training algorithm based on compound gradient vector. Lecture Notes in Computer Science, 2557:374–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570374.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570374.pdf.

Kim:2002:AWT

[3428] Cheol-Ki Kim and Eui-Young Cha. Applications of wavelet transform and artificial neural networks to pattern recognition for environmental monitoring. Lecture Notes in Computer Science, 2557:385–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570385.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2557/25570385.pdf.

Gao:2002:AKV

[3429] Junbin Gao, Steve Gunn, and Jaz Kandola. Adapting kernels by variational approach in SVM. Lecture Notes in Computer Science, 2557:395–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2557/25570395.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2557/25570395.pdf.

Verbeeck:2002:LRP

[3430] Katja Verbeeck, Ann Nowé, Tom Lenaerts, and Johan Parent. Learning to reach the Pareto optimal Nash equilibrium as a team. Lecture Notes in Computer Science, 2557:407–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2557/25570407.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2557/25570407.pdf.

Zhou:2002:CMA

[3431] Weidong Zhou and Richard Coggins. Computational models of the amygdala and the orbitofrontal cortex: a hierarchical reinforcement learning system for robotic control. Lecture Notes in Computer Science, 2557:419–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2557/25570419.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570419.pdf.

Hofstedt:2002:GAB

[3432] Petra Hofstedt. A general approach for building constraint languages. Lecture Notes in Computer Science, 2557:431–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570431.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570431.pdf.

Ghose:2002:MSP

[3433] Aditya Ghose and Peter Harvey. Metric SCSPs: Partial constraint satisfaction via semiring CSPs augmented with metrics. Lecture Notes in Computer Science, 2557:443–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570443.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570443.pdf.

Wilke:2002:HGA

[3434] Peter Wilke, Matthias Gröbner, and Norbert Oster. A hybrid genetic algorithm for school timetabling. Lecture Notes in Computer Science, 2557:455–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570455.
htm;   http://link.springer.de/

                                        link/service/series/0558/papers/
                                        2557/25570455.pdf.

Yao:2002:GSM

[3435] Wensheng Yao, Baiyan Li, and Jinyuan You. Genetic scheduling on minimal processing elements in the Grid. Lecture Notes in Computer Science, 2557:465–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2557/25570465.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2557/25570465.pdf.

Wang:2002:PSC

[3436] Dianhui Wang, N. K. Lee, T. S. Dillon, and N. J. Hoogenraad. Protein sequences classification using modular RBF neural networks. Lecture Notes in Computer Science, 2557:477–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2557/25570477.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2557/25570477.pdf.

Jie:2002:FES

[3437] Sun Jie, G. S. Hong, M. Rahman, and Y. S. Wong. Feature extraction and selection in tool condition monitoring system. Lecture Notes in Computer Science, 2557:487–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2557/25570487.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2557/25570487.pdf.

Karras:2002:RME

[3438] D. A. Karras and B. G. Mertzios.

A robust meaning extraction methodology using supervised neural networks. Lecture Notes in Computer Science, 2557:498–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570498.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570498.pdf.

Frayman:2002:SRP

[3439] Yakov Frayman, Bernard F. Rolfe, and Geoffrey I. Webb. Solving regression problems using competitive ensemble models. Lecture Notes in Computer Science, 2557:511–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570511.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570511.pdf.

Matsumoto:2002:LFU

[3440] Satoshi Matsumoto, Takayoshi Shoudai, Tetsuhiro Miyahara, and Tomoyuki Uchida. Learning of finite unions of tree patterns with internal structured variables from queries. Lecture Notes in Computer Science, 2557: 523–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2557/25570523.htm;   http:

                                        //link.springer.de/link/service/
                                        series/0558/papers/2557/25570523.
                                        pdf.

Gopalan:2002:TMM

[3441] Raj P. Gopalan and Yudho Giri Sucahyo. TreeITL-Mine: Mining frequent itemsets using pattern growth, tid intersection, and prefix tree. Lecture Notes in Computer Science, 2557:535–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2557/25570535.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2557/25570535.pdf.

Zhang:2002:CLP

[3442] Dongmo Zhang and Norman Foo. Convergency of learning process. Lecture Notes in Computer Science, 2557:547–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2557/25570547.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2557/25570547.pdf.

Bain:2002:SFC

[3443] Michael Bain. Structured features from concept lattices for unsupervised learning and classification. Lecture Notes in Computer Science, 2557:557–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2557/25570557.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2557/25570557.pdf.

Thornton:2002:TFP

[3444] John Thornton, Wayne Pullan, and Justin Terry. Towards fewer parameters for SAT clause weighting algorithms. Lecture Notes in Computer Science, 2557:569–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570569.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570569.pdf.

Kravchuk:2002:IVR

[3445] Olena Kravchuk, Wayne Pullan, John Thornton, and Abdul Sattar. An investigation of variable relationships in 3-SAT problems. Lecture Notes in Computer Science, 2557:579–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570579.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570579.pdf.

Slater:2002:MMR

[3446] Andrew Slater. Modelling more realistic SAT problems. Lecture Notes in Computer Science, 2557: 591–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2557/25570591.htm;   http:
//link.springer.de/link/service/

series/0558/papers/2557/25570591.
pdf.

Thornton:2002:TLL

[3447] John Thornton, Stuart Bain, Abdul Sattar, and Duc Nghia Pham. A two level local search for MAX-SAT problems with hard and soft constraints. Lecture Notes in Computer Science, 2557:603–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570603.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570603.pdf.

Karras:2002:SPB

[3448] D. A. Karras and V. Zorkadis. Strong pseudorandom bit sequence generators using neural network techniques and their evaluation for secure communications. Lecture Notes in Computer Science, 2557:615–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570615.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570615.pdf.

Laghari:2002:SFR

[3449] Mohammad Shakeel Laghari. Surface feature recognition of wear debris. Lecture Notes in Computer Science, 2557:627–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570627.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2557/25570627.pdf.

Karras:2002:IDD

[3450] D. A. Karras and B. G. Mertzios.

Improved defect detection using novel wavelet feature extraction involving principal component analysis and neural network techniques. Lecture Notes in Computer Science, 2557:638–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570638.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570638.pdf.

Matsuhara:2002:EMT

[3451] Masafumi Matsuhara, Kenji Araki, and Koji Tochinai. Effectiveness for machine translation method using inductive learning on number representation. Lecture Notes in Computer Science, 2557:648–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570648.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570648.pdf.

Williams:2002:EEC

[3452] Graham Williams, Rohan Baxter, Chris Kelman, Chris Rainsford, Hongxing He, Lifang Gu, Deanne Vickers, and Simon Hawkins. Estimating episodes of care using linked medical claims data. Lecture Notes in Computer Science, 2557: 660–??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2557/25570660.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2557/25570660.
                                        pdf.

Tran:2002:AMF

[3453] Cong Tran, Ajith Abraham, and Lakhmi Jain. Adaptation of a Mamdani fuzzy inference system using neuro-genetic approach for tactical air combat decision support system. Lecture Notes in Computer Science, 2557:672–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2557/25570672.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2557/25570672.pdf.

Arotaritei:2002:ORN

[3454] Dragos Arotaritei and Mircea G. Negoita. Optimization of recurrent NN by GA with variable length genotype. Lecture Notes in Computer Science, 2557:681–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2557/25570681.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2557/25570681.pdf.

Wang:2002:TFN

[3455] Dianhui Wang and Tharam S. Dillon. Theoretical foundation for nonlinear edge-preserving regularized learning image restoration.

Lecture Notes in Computer Science, 2557:693–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570693.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570693.pdf.

Han:2002:ACB

[3456] Peng Han, Ruimin Shen, Fan Yang, and Qiang Yang. The application of case based reasoning on Q&A system. Lecture Notes in Computer Science, 2557:704–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570704.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570704.pdf.

Chau:2002:DIT

[3457] K. W. Chau, F. Albermani, and S. L. Chan. Development of an intelligent tutoring system on design of liquid retaining structures. Lecture Notes in Computer Science, 2557:714–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570714.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570714.pdf.

Chau:2002:RTP

[3458] K. W. Chau and C. T. Cheng.

Real-time prediction of water stage with artificial neural network approach. Lecture Notes in Computer Science, 2557:715–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2557/25570715.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2557/25570715.pdf.

Kerber:2002:DMC

[3459] Manfred Kerber and Martin Pollet. On the design of mathematical concepts. Lecture Notes in Computer Science, 2557:716–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2557/25570716.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2557/25570716.pdf.

Liu:2002:CTT

[3460] Chuchang Liu and Maris A. Ozols. Consistency of trust theories. Lecture Notes in Computer Science, 2557: 717–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2557/25570717.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2557/25570717.
                                        pdf.

McCabe:2002:AIS

[3461] Alan McCabe. An artificially intelligent sports tipper. Lecture Notes in Computer Science, 2557: 718–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2557/25570718.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2557/25570718.
pdf.

Brown:2002:SSU

[3462] Marianne Brown. Selecting semantics for use with semantic pruning of linear deductions. Lecture Notes in Computer Science, 2557:719–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570719.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570719.pdf.

Chau:2002:CFW

[3463] Kwokwing Chau. Calibration of flow and water quality modeling using genetic algorithm. Lecture Notes in Computer Science, 2557: 720–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2557/25570720.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2557/25570720.
pdf.

Jelinek:2002:CML

[3464] Herbert Jelinek, David Cornforth, Patricia Waley, Eduardo Fernandez, and Wayne Robinson. A comparison of machine learning approaches for the automated classification of dementia. Lecture Notes in Computer Science, 2557:721–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2557/25570721.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2557/25570721.pdf.

Governatori:2002:DLP

[3465] Guido Governatori, Vineet Padmanabhan, and Abdul Sattar. A defeasible logic of policy-based intention (extended abstract). Lecture Notes in Computer Science, 2557: 723–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2557/25570723.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2557/25570723.
                                        pdf.

Richer:2002:SOT

[3466] Toby J. Richer and Dan R. Corbett. A self-organizing territorial approach to multi-robot search and surveillance. Lecture Notes in Computer Science, 2557:724–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2557/25570724.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2557/25570724.pdf.

Mohamed:2002:KBT

[3467] K. A. Mohamed, A. Datta, and R. Kozera. Knowledge-based techniques for constraints satisfaction in resource allocation problems. Lecture Notes in Computer Science, 2557:725–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570725.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570725.pdf.

Nabeshima:2002:ESP

[3468] Hidetomo Nabeshima, Koji Iwanuma, and Katsumi Inoue. Effective SAT planning by speculative computation. Lecture Notes in Computer Science, 2557:726–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2557/25570726.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2557/25570726.pdf.

Anonymous:2002:AIdg

[3469] Anonymous. Author index. Lecture Notes in Computer Science, 2557: 729–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2557/2557auth.pdf.